A Complete Bibliography of Publications in Theory of Probability and its Applications: 1956–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 May 2024 Version 3.01

Title word cross-reference

(1 ≤ p < ∞) [Vak66]. (B, ρ) [KZ96]. (B, S) [KS95]. (c_{n}) [VV88]. (F, F_{1}) [Chu94]. (r, s) [MP93b]. 1 < p ≤ 2 [Küh81]. 1 ≤ P < ∞ [Nor93b]. 2 ≤ q < ∞ [Küh81]. 3 [Lut78]. 5 [Ken95]. A [Cho93]. α [Chi89b, KS85b]. B [BGS61]. β [Nai82]. C(0, ∞) [BS74]. C(α) [VZ80]. C_{0} [KS91]. C_{p} [PT91]. χ^{2} [Bor78c, Cad99, Gik56, Ing88, Ric64, Sel85, SP57, SP61, Tum56, Bor77c]. D [BG73, Dau76, Nov88, QT97, VS91, VN94, VN99]. D[0, 1] [Rao63b]. Δ [Asa88, Lut78]. Δ + b · ∇ [KS91]. ℓ [IN78]. F [BYK88]. g [Rya79]. Γ [Pag65, SP57, Küh81]. GI/G/1 [DM68]. GI/G/1 → G′/1/∞ [KK82]. GI|G|1 [Rie80]. GI|M|1 [Ivn77]. G|G|1 [Zol77b]. G|G|∞ [Zol78d]. H [Ul’87, Sac71b]. H^{∞} [Pav80]. I_{0} [Fel85, Fel86, Ula84]. k [BS73b, Gik57, Kel73, Yak96]. kt + a [Mir69]. L [Bra80, Ibr57]. L^{0} [Leb96b]. L^{2}(Ω, F, P) [Ale92a]. l^{p} [Fat97]. L^{p}(R^{d}) [KS91]. L_{2} [Nad73a, Nad74a, Sed77]. L_{p} [CDM93, Ing87, Ing95, Mus69, Vak64, Vak66, Nor93b]. l_{p}(1 ≤ p ≤ +∞) [TT80]. l_{p}(x), 1 ≤ p < +∞ [KT77]. l_{p}, 1 ≤ p < +∞ [Tie77]. L_{Φ} [Meg86]. L_{r} [ZKC91]. M [Abr87b, Abr88, Gur89b, Lis94, MA83, PR99, She80, Sun96, Sun99]. M/G/1 [Gri92b]. M/G/1/P S [Gri95]. mA [Sac71a]. R^{∞} [Med98]. G [Jur83]. G_{⇕}(⊔) [Ste70a]. G_{⇕}(§|⟨) [Ste73]. AU + V u = 0 [Kha59]. L [Chi95]. N [Zai81, Zai92, Ara81, Kor66, Kor68, SS62, Sol67, Vil68, Yag57]. OL [Yam85]. ω^{2} [Bol90, Bor85d, Bor86b, Csö75, Ing88, Khm80, KMT78, LS96, Mar76b, Mog77c]. P [Chu92a, Top67, Khr96b, Khr96a, Khr98, Küh81, MR93, Tsa79, Ula84, Vla81]. p(0 ≤ p ≤ 2) [Mus80b]. p ≥ 2 [Fat97]. q [Küh81]. R [Ul’77]. R^{d} [Jur83, Nag80, Cho87]. R^{K} [SM79, Nys87, Sen84]. R^{m} [MP78b, Nik88a, Pis82]. R^{m∗} [Pis98]. R^{N} [Pus92]. R^{p} [Sem85]. ρ [Gap92]. s [Bel65, Bel69a, Bel69b, KC75a, ZM80]. S^{2} [Bak89a]. σ [Gri79, Kry80, Pit87a, Pit87b, Sko92, Tor88, TF62]. σ^{h+1} = σ^{h} [Sac71b]. Σf(x2^{n}) [Mos70]. ∑ f(2^{k}t) [Ibr67a]. ∑ f(t2^{k}) [LM71]. T [Tik93, Lev75a]. T^{2} [OCG93]. τ [Gel74]. U [Ale91b, BG96, BPS97, EL99, Gru86, Gru88, IL94, Kha88, Kha93, Kha94, KA99, KB86, MA83, PS92b, Ron82, Ron86b, Saz92, Van83]. UB [Bor92c]. UH [KB92]. V [Vol96b]. ε [Ero58, Fai79, Fai80, KY84]. φ [Deo77, Lin92, RS88, Ute91, Yan76, Zub75]. φ^{4} [Ble86]. w [PP96]. w_{n+1} = f(w_{n}, ξ_{n}) [Bor89a]. Ĉ(R^{d}) [KS91]. Z [BGS61, CH93]. Z^{2} [Ste91]. Z^{n} [Haf73].

-Adic [Khr96b, Khr96a, Khr98]. -Algebra [Gri79, Pit87a, Pit87b, TF62]. -Algebras [Kry80, Sko92, Tor88]. -Amarts [KZ96]. -Branching [Yan76, Zub75]. -Chains [Bel69b]. -Connectedness [Kel73]. -Consistency [VV88]. -Continuity [Top67]. -Cycles [Bel69a]. -Cyclic [Rya79]. -Dependent [MA83, She80, Sun96, Sun99, PR99]. -Dimensional [Dau76, Ken95, Kor66, Kor68, Lut78, Sol67, Yag57, QT97]. -Distribution [BGS61, Pag65]. -Divisibility [Lis94]. -Entropy [Ero58, MP93b]. -Estimates [Gur89b]. -Estimators [Tik93]. -Factorizations [Chi89b]. -Fields [Bra80]. -Fold [Ara81, Zai81, Zai92]. -Function [SP57]. -Functions [PP96, Ula84]. -Guarantee [VN99]. -Guaranty [Nov88]. -Hierarchical [Ble86]. -Independence [KY84].

-Infinitely [Vol96b]. -Isotonic [CDM93]. -Market [KS95]. -Max [MR93]. -Meeting [Lut78]. -Mixing [Deo77, Gap92, Lin92, RS88, Ute91]. -Norm [Nad74a, ZKC91, Nad73a]. -Optimal [Fai79, Fai80, BG73, Cho93]. -Person [Vil68]. -Processes [EL99]. -Random [Asa88]. -Representation [CH93]. -Risk [VN94]. -Sample [Sac71a]. -Scheme [Abr87b, Abr88]. -Semigroups [KS91]. -Sequential [OCG93]. -Space [SS62]. -Stability [Tsa79]. -Stable [Jur83]. -Statistic [LS96]. -Statistics [BS73b, Bor85d, Bor86b, Bor92c, KB92, Ale91b, BG96, BPS97, Gru86, Gru88, IL94, Kha88, Kha93, Kha94, KA99, KB86, MA83, PS92b, Ron82, Ron86b, Saz92, Van83]. -Summing [Küh81, Vla81]. -Symmetric [KS85b]. -Test [Csö75, PT91]. -Tests [VZ80]. -Tuple [IN78]. -Tuples [Bel65, KC75a, ZM80]. -Uniformity [Top67]. -Valued [Leb96b, Rao63b].

175th [Ano97d]. 1960 [Ano61g]. 1961 [Ano62e]. 1962 [Ano63c]. 1963 [Ano64d]. 1964 [Ano65d, Bor65a]. 1965 [Ano67f]. 1965-1966 [Ano67f]. 1966 [Ano67f]. 1970 [Ano71c]. 1972 [Ano73k]. 1975 [Ano77h, Ano78g, Ano78j, Ano79g]. 1976 [Ano78i, Ano78h]. 1977 [Ano78k, Ano79f, Ano79i, Ano79h]. 1978 [Ano79j, Ano80e, Ano80f, Ano80i, Ano80h]. 1979 [Ano80g, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano81k, Ano82l, Ano82n]. 1980 [Ano82q, Ano82g, Ano82p, Ano82o, Ano82m, Ano82h, Ano82i, Ano82j]. 1981 [Ano82k, Ano83e, Ano83d, Ano83c, Ano83f, Ano84h]. 1982 [Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p]. 1982-1983 [Ano85q]. 1983 [Ano85q, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o]. 1984 [Ano85r, Ano87k, Ano87n]. 1985 [Ano87j, Ano87m]. 1999 [Ano99d].

20th [Ano99d]. 23 [Bak89b]. 2nd [Ano93a].

32 [Bak89b]. 33 [Fry90, Ul′91]. 34 [Shu91c]. 39 [SZ98].

62118a [Bak89b]. 652 [Leo82].

80h [Bak89b]. 80th [Shi93]. 89k [Ul′91]. 89m [JK90].

90a [Fry90]. 90b [Ul′91]. 90m [Bak89b]. 91d [Shu91c]. 97c [SZ98].

A. [Ano73g, Ara76, AS74, Bin90, Bor65a, GS63, JK90, Lin64, Lis96, Pro60, SS90, Vak90]. Aat [Ano65d]. Abel [Gap65]. Abelian [Fel79, FF82, Fel88, Fel90, Fel92, Fel93, Fel98, Fel99, Klo61, PS64, She99, ZK76]. Above [Nos80, BN69]. Absence [But92b]. Absolute [BK89, BMS96, Dav89b, EN75, Ers72a, IC98, Ivn93, Lif83, Lod71, Mog74, Pra99, Sal98, Sat73a, Sko65, SY73, VIL69, Ven68, Ven61, Yam95, Zol66a, Bol97]. Absolutely [Ber84, Gir60a, Shu91a]. Absorbing [Bev58, Mai75b, Man64a]. Absorption [Lev80, Nag68a, Nag71b, Pec81]. Abstract [Nas57, Per59, Roz59b]. Academician [Ano88h]. Academy [Ano61g, Ano62e, Ano63c, Ano64d, Ano65d, Ano67f, Ano68f, Ano69h, Ano71c, Ano71g, Ano73h, Ano73k, Ano74j, Ano75a, Ano75k, Ano76f, Ano76g, Ano77h, Ano78i, Ano78h, Ano78k, Ano79f, Ano79i, Ano79j, Ano80f, Ano80i, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82m, Ano82h, Ano82j, Ano82k, Ano83c, Ano83f, Ano84h, Ano84j, Ano84l, Ano84i, Ano85q, Ano85p, Ano85k, Ano85n, Ano85o, Ano85r, Ano86k, Ano87k, Ano87n, Ano87j, Ano87m, Ano87l, Ano88l, Leo82]. Accelerated [DK93, Zul76]. Acceptance [Mik56]. Access [Kha90b, Mik96a].

Accompanying [Ara81, BGZ98, Haz96, IP74, Stu68, Zai84]. According [Mir84c, Sav64]. Account [Lin61a, Lin61b, Lin62, Nag69a, Nag69b, Nag69c, Sho79, Sho83]. Accumulation [Bag79]. Accuracy [Bar86, Ibr66, Ibr94, Kir91, Kor89b, Kor89a, Kor57, Kor91, Lif76a, Lot89, MN89, Mil79a, Roz87, Saz90, Shi80, Shi82, Tik92a, Tik94, Vin86a, Vin86b, Yur83, Zai84, Zal83, Zal89, ZSU89c, Zal90a, ZM79b]. Achieved [Sht66]. Across [Bel68, GK68, Rag86]. Action [Bas58, Fre73, VT73b, KL93]. Actuarial [Ano95c]. Acyclic [Lis76]. Adaptation [Lep93]. Adaptive [Aga80, DF93, GN93, Gor85, Lep91, Lep93, Lep92]. Addendum [Ano85a, Bag89a, Ber70a, Bez80, Bor79c, Bor80b, CvZ89, Chi65b, Dob66, Dyn68a, Eng85, Erm66, Fed70, Gam91, Gam92a, Gih91, Gor80, Gri78, Gri94b, Ios70, Kog66, Kog70, Kor68, Kra80, Kru71a, KA87, Kry73a, Lab79, Lan90, Leo82, Liv80, Lum71, Mal77, Mal92a, Mal95, Mal74b, MP78a, Mat85, Mat86a, Mat65b, Nag69a, Nag69d, NR76, Nag77, NK82, Nag84a, Nag85, Nev72, Nov96, Osi74, Ost67, Pau79, Pet66a, Rak90, Rez80, Rez71, Ryz72, Sad67a, Sad67b, San66, Saz90, Saz92, SC64a, Sev66a, Shi82, Shi67, Shu66a, Shu91b, Shu87, Sli64, Sol67, Tru66, Tur91, Ul’79a, Val82, VT82, Yag68, Yur72, Yur78a, YF80, Yus91a, Zai85, Zai88a, Zol66b, Zol78a, Zol78b, Zol84a]. Addition [MFY74, She97]. Additive [Afa91a, Bag79, Bor81a, Bui82, BZ77, Dir81, DK84a, DK84b, Dyn60, Fal78, MR96, Por67, Roz60a, Shi64b, Shu80b, Sko61a, Vla79, Vol93, Vol59, Yan70, Zak90a, JK90, KY97, Ver94]. Additively [Mna86]. Adic [Khr96b, Khr96a, Khr98]. Adjoined [Kry87b]. Adjoint [Shu70b]. Adjusted [BN94]. Adjustment [Kah93]. Admissibility [Dev69, Fin71, Kak78, Kle74a, Ruk67]. Admissible [BNS66, Ber84, Chi67, Sko70b, Syt69].

Admitting [Chi73a, Chi81a, Chi81b, KR75, Sap88]. Advances [Kal93a]. Affine [Oza86]. After [Khi56b, KK96, Mil82, Lev75a]. After-Effect [KK96, Mil82]. After-Effects [Khi56b]. Against [Ing83, Ing89, Pen93]. Age [Chi72, Jag68, Sev64, Sev66a, Sev68, Vin64, Yag68]. Age-Dependent [Chi72, Jag68, Sev64, Sev66a, Sev68, Vin64, Yag68]. Aids [WS93, Góm93]. Aleksandrovich [Ano94a]. Alexander [Gne60]. Algebra [Gri79, Ism73, Pit87a, Pit87b, Sap88, Ste69a, TF62, Tut67]. Algebraic [Gir92, Puk82, Vis86]. Algebras [Kry80, LS97, LSS99, Sko92, Tor88, QT97]. Algorithmic [She93b, Vov92]. Algorithms [Kha90b, KU88a, Lot96]. All-Union [Ano63b, Ano64c, Ano80k, Ano81b, Ano81a, Ano82a, Ano89a, NK79, Pro83, Ano59a]. Allocated [Mor75]. Allocating [ZM79b]. Allocation [IT79, IM66, IM89, Kol71, Kol76, Mak96a, Mik78, Mik80, Mik98, Sel99, Sev66b, VM83]. Allowance [LNF89]. Allowing [Jag68, Yag68]. Almost [Bas77, CN87, Gap80, Gap88, Gap98, Gla63, HR97, Jak98, Kho71a, Kle96, KZ96, Mus70, Pre89, Pre98, Sko79, Sta84, ZKC91, IK74a, Wol76]. Almost-Periodic [Kho71a]. Almost-Periodically [Gla63]. Along [Kuz80, Ven65]. Alphabet [Per59, Qua99]. Alternative [Chi64b, Ing83]. Alternatives [Ber86a, Ing89, Kud78, Mar69, Tim84, TC86]. Amarts [Dzh89, Dzh95, KZ96, Zie92]. American [SKKM95a, SKKM95b]. Among [Fra93, Shl89]. Amount [Tin62a, TP58]. Analogies [Zub75]. Analogs [Bul85b, KMM85]. Analogue [FM92, Ibr89, KL77, Lis96, Ost65, Ost67, Pit71b, ST74, Shi86, Zol66c]. Analogues [AB93, Gri94a, Pro68b, Sad66b, Sad67b, Tep95]. Analysis [Ale74a, Ano61b, BK65, Bor66, Bor95, BD91a, BD91b, Fed69, Fed70, Gir88, GS65, Hid70, Hid83, IK74a, Kal77a, Kel77, KS57, Kor59, Kor62, Kor90b, Lot80a, Lot80b, LN91, Mir79a, NS70, Par70, Roz59b, Rug93, Sem88a, Shi64b, Szt93, Tsi76d, Vol61b, Vol74, Vor65, Gil93a, Gil93b]. Analytic [Bel59, BM63, Jes80, LE68, Roz60b, Seĭ91, Vor61]. Analytical [Kag90, Zin58]. Analyticity [Ven82]. Ancestor [Zub76]. Andrei [Ano88h, Shi90]. Angular [Mir79b, Mir84c]. Anniversary [Ano82r, SS90, Ano97d]. Announcement [Ano91a, EH93]. ANOVA [Kul87, PP93]. Answer [Soz64]. Antanovich [KGPS91]. Anticipation [Dor96a]. Any [Kuz81]. Appearance [Bel65]. Applicability [Lët93, Roz61, Ver66]. Applicable [Ver64a]. Application [Amb65, Ano59b, Ano60a, Bor79d, Bor85d, Chi67, Chi74b, Cow93, Dor87, Gam89, Góm93, Kha65, Kir94, Kog65, Kog66, Le93, Lif82, Lif85b, Lin57b, Mil97, Nov91, Pin91a, Ric64, Rog60, RR64, Sev57a, Sol66, Tut73, Vin68, Vor65, Wal93, Yag56, Yur71a, Yus88]. Applications [AHE85, AI99, Bol65a, Bor79c, Bor79e, Bor89a, Bor86b, CPP98, Chi64d, Fil62, Har91, Ken87, Kon85, Mal74a, Mal77, Mar85, Pav91b, Rac85a, Rak88, Rak90, RS94, Ryc84, Sad66b, Sad67b, Sal80, San66, Sie80, Ste89, Ste91, Tou94, Vak94a, Vin77, VZ80, Ano90b]. Applied [Nik91, Vla56]. Approach [Che56, Chi90, Dyn98, IP74, KD87, Khm82, Khr96b, Kry87a, Nak88, Nos91, Tsi83, Tsi86, Tsi87, VS91, Vov92, Svi90]. Approaching [Kud78]. Approximate [GR67, Lin66, Mil75, Zam70]. Approximated [Cle98]. Approximately [Bak98, LPS67]. Approximating [Zol78c]. Approximation [AL91, Ara81, BD93, Bar86, BGZ98, BB87, Bor79b, Bor89c, BPS97, CDC82, Dub91, Erm66, Fal92, GP71, GK75, Ger96, Gla65, Hag74, Ibr66, Ivk68, Kir91, Kor89b, Kor89a, Kor90a, KB86, Kor78a, Kor84, Kra62, Kra69, Kra71, KLP89, LS97, Lif76a, LS65, Lot96, Lot89, Mes60, Mik98, Mil86, Mir88, Mol96, Muk98, Nev73a, Nev86, Nov95, Nov96, Pre84, Pre86, PW87, PS89, Roo99, Roz99a, Roz87, RS80, Sal80, Saz90, Shi80, Shi82, Shu91d, Sob73a, Sor78, SA94, Stu60, Tik92a, Tik94, Tsa58b, Tut68a, VTGM85, Vin86a, Vin86b, Yag60, Yur78b, Yur78c, Yur83, Zai84, Zai88b, Zal83, ZS84, Zal89, ZSU89a, ZSU89b, ZSU89c, Zal90a, Zol77a, Zol78a, Zol78b, ZM79b, Zul76, Ivk94, Kle97, Ul′91]. Approximations [Dev69, EZ60, Mes61, Mil88]. April [Ano62d, Ano64g, Ano67g, Ano73i, Ano77g, Ano82k, Ano87o]. Arbitrage [KK95a, KK95b]. Arbitrary [Chi76a, Fai83, GD71, Gus84, IM65a, KS80, Kha68, Kha69a, KLP89, Leo78, Lev86b, Lya87a, Mak67b, Mak70b, PP96, Son92, Yan96]. Arc [Bre65, Nos87]. Arc-Sin [Bre65]. Area [KT69]. Argument [Il’76a, Pro68a, Zam70]. Arising [Das79, Dob58d, Ign82, Kag88, Kle79, Pro92, SS62, VS77, VS78]. Arithmetic [Kur88, LN91, MPY70, Ost87, Rog88b]. Armed [Yus91b, Pre91]. Array [Bet89, Sīl92]. Arrays [Ute91, QT97]. Arrival [Bra91, KLP89]. Arriving [Ivn93]. Ars [Yus87]. Aspects [EK94]. Associated [BS68a, Bul94, Bul96, Chu92b, CZ79, Gru86, Gus76a, Gus76b, LSS99, Mik74, Mol88, PP96, PR69a, PS61, VT73a, Vla56, Vro99, ZM74]. Association [PS75, XP89]. Assumptions [BG96]. Asuming [Ros80]. Asymmetric [Mik70b, Nag84b, Roz98b, Vas80, VT92]. Asymmetrical [KN75]. Asymptotic [AB85, Ale67, Ale96, Amb76, Amb79, Are77, Arn81, AN93, Bab85, BD97a, BT87, Ber78, Ber87, Ber74, Ber76, Bet86, Bik69b, Bik69a, BB81, Bor82a, Bor85a, BK65, Bor66, BD91a, Bun70, Bur97, CdHY97, Chi64b, Chi65a, Chi67, Chi73a, Chi74a, Chi74b, Chi80a, Chi80b, Chi81a, Chi81b, Chi89a, CV65, Chi71b, Csö75, DI71, Den97, DP82, Dob62a, Dob63, Dob64, DN88, DL86, Dzh75, Efr80, Erm90, Erm96, Erm98, Fil62, Gal71, Gär75, Gik56, God82, Gol75, Gus76a, Gus76b, Gus96, Ibr64, Ibr67c, IS68, IK73b, IK73a, IK74a, IH76a, IK79, Ign82, Ing83, Iva77, IM65a, IM66, Ivc73, IL76b, IL78, IM89, IT89, JR91, KK95a, Kal67, Kar81, Kel73, Kha88, Kha90a, Kha67, Kho71b, Kho71c]. Asymptotic [Kol88, Kon74a, Kon79, KNS94, Kor59, Kor61b, Kor62, Kor90b, Koz77a, KA83, KA85, KA87, Kut76, KY74, Lep88, Lep90, Lev74, Lev75a, Lev78, Lev81, Lev86a, Lot80a, Lot80b, Lot86, Lot88, Mal64, Mal69b, Mal69a, Mal87a, Mal93, Mal94, Mal95, Mat65a, Mat65b, Mik75b, Mik80, Mir84a, Mir69, Mit67, MTV91, Mor75, Mos70, NS70, Nag70a, Nag82, Nev79, Nik85, Nik88b, Nos85, Nos88, Nov88, Osi69, Osi71, OT95, PR87, Pag98, PvZ97, PK87, Pav78, PR99, Pen77, Per92, Pet59a, Pet68, PP82, Pit69a, PT91, PS69, Rag93, Riv96, Ron82, Roz76, Roz78, San66, Sar90, Sen74, SC64a, SC64b, Sev67a, Sev99, Sgi92, SS65a, Shu85a, Shu86, Shu96, Smo97, Sta65, Szt93, Tim99, Tsa77, Tum56, Ul’87, Vat74, Vat77a]. Asymptotic [VM97, VT76, VS78, Vin95, Vis64, VN99, Vov92, V’y77, Wol65, Yak86, Yak87, Yak88, Zäh84, Zäh86, Zin88, Zol61b, Zol65a, Zol66b, Zol92a, Ego97, Mak88, Bur75b]. Asymptotical [BM92]. Asymptotically [Ali95a, Ber86a, Ber87, Bol63, BS68b, Bor76, Bor99, Chi86a, Dzh71, Erm89, Gin89a, Ing84, Ing86, Kho79, Kli62, Kol88, KS84, KS85a, Kor94a, Kus68, KP71, Lep92, Lep93, Mal86, Mal90b, Nev81, Pav91a, Pav92, PT93, Ras77, Rog73, Sat95a, Sat96, Son92, Ter93, VT92, VN94, Zol64b, Sil79]. Asymptotics [Bak96d, IJL91, Kif77, Rog77a, Sgi98, Tik93, Tik98, Top83, Yak96, Yur92a, Yur92b, Yur93a]. Asympyotic [VS77]. Atmospheric [Mon58]. Atomic [MW99]. Attainment [Sem90]. Attracted [God79, Pet74a, Vas80]. Attracting [Ste71]. Attraction [Amo91a, Gri91a, Gri94c, GK96b, Jur83, Kru93, MR93, Roz90c, Roz98b, Sak56, Shl89, Zin86, dH95]. Attrition [Yus67]. Australian [WS93]. Automata [Ose90]. Automaton [Fed69, Fed70]. Automorphism [Abr59, Rok61]. Autoregression [Bol83, Kap74, KP97, Per92, Bol97]. Autoregressions [Kos99]. Autoregressive [Bra94, CFZ95, Mal93, Mal95, Nov91, Sal93a]. Autoregressive-Moving-Average [Bra94]. Available [Bas60, Vol65]. Average [Bol90, Bra94, Fai80, GE91, Iva60, Kis62, Kit82, Kit86, MF71, Nos79, Oni90, Ruk67, Sev62, Vol79, Yus82]. Averages [Gap98]. Averaging [Gih91, Gik90, Kha63, Ver99b, Ver99a, Yur89, Kle97]. Award [Ano98a]. Axiomatic [Khr96b, Vil63, Vil68]. Axioms [van94]. Axis [Afa87, Her91, Ros80, Roz94b, Shu81, Shu82b, Tit82].

B [Bak89b, KKSY95, Ul′91]. B. [Amb76, Chi87a, Kol62, ZK61]. Back [WS93]. Back-Projection [WS93]. Background [Bur79]. Backward [Leb74, Roz84b]. Bagirov [Bak89b]. Bahadur [BDEM92, Nik85, Ral97]. Bahadur-Kiefer [BDEM92]. Ball [Kol66, Kol67, Lut78, SC64a, SC64b, Sev67b, Tik92a, Tik94, Yur83]. Balls [Bor91c, CV65, Lin90, Pau91, Sen84, Sen87]. Banach [AV82, BO96b, BR85, Ben87, BS81, Bul74b, Bul74a, BP78, Bul85a, Cap98, Chu92b, Chu94, Chu95, DEO84, EO86, Gia96, Kol92, KA83, KA85, KA87, Küh81, LP75, LTC80, Lin90, MP92, Met74, Met75, Mus85, Pau77, Pau79, Pau91, Pin79, PRR87, Raj73, Sak85, Sie90, Sko70a, TT82, TT83, Thu85, TTC78, Tie79, Tie80, Vla80, Woy74, Zal89, Zal90a]. Banach-Space [AV82]. Banach-Space-Valued [Sak85]. Banach-Valued [Chu95, DEO84, EO86]. Band [Bor79c]. Bandit [Yus91b, Pre91]. Bandwidth [Tsy88]. Barrier [Man64a].

Bartlett [Kah93]. Based [Art89, Chi61a, DK93, GP79, KR90, Mal86, PC96, Pet56a, Plu85, Ron86a, Ron89, Sap79, Tik88, ZKC91]. Basic [Nas57, Sav99, Shi95]. Basis [Kha66a, Pav98, Sac71a]. Batch [BN96]. Baxter [Ara72, Gor89]. Bayes [Bul77, IK73a, IH76b, Ing84, Lev75b, Mik56, Ol′89, Shi64b, Str78]. Bayesian [IK74b, Lan88, LS65, MP93b, Par70, PP93, Rug93]. bayesiennes [Lan90]. Be [Bor83a, Eag77, Mal80, Son92, Vla80, Kag90, Kop96, Ver64a]. Before [Der93]. Behavior [Amb79, Are77, AN93, Bab85, BD97a, BT87, Bet86, Bet88, BB81, Bor82a, Bor85a, Büh68, Chi80a, Chi71b, DI71, Fil62, Gär75, Gol75, Gor71, GPU98, Haz95, Ibr64, IS68, IK73b, IK73a, IM66, Ivc73, Kha93, Khr98, Kif73, Kol69a, Kol87, Kol88, KNS94, Kor61b, Kov86, Koz77a, Kru75, KA83, KA85, KA87, KU88b, Kul67, KH81, Kuz80, Leb96a, Lev75b, Lot86, Mal64, Mos70, Nag82, Nev64, Pet68, PP82, Rog68a, Rog72, Roz78, Sar90, Saz65, Sev67a, Sev99, Sgi92, Sha80a, SS65a, Shu69a, Shu69b, Sin83, Tar85, Tut62, Vat77a, VT76, Vve85b, WR69, Yak88, Zak91, Zin88, Zol61b, Zol64a, Zol65a, Zol66b, Zub83, Den97, Ego97, Mak88]. Behrens [CS71]. Being [Asa91]. Bellman [Rom59, Sag86, Vat77d, Vat77b, Vat77c, Vat79, Vat80, Vat81, Vat82, Vat87a, Vat87b, VS89, VS90, Vat94b, V’y78a, Yus88]. Bellman-Harris [Sag86, Vat77d, Vat77b, Vat77c, Vat79, Vat80, Vat81, Vat82, Vat87a, Vat87b, Vat94b, V’y78a]. Belong [Kru93]. Belonging [IK83]. Below [Mir69, Sev96]. Berlin [JK90]. Bernoulli [Ano88a, Ano90b, Ano90h, Ano93a, Ano90a, Cho84, Khr98, Kor59, Liv75, Liv80, Mas74a, Pos61, Pre86, Yus87]. Bernshtein [Ano69f, Gin66, KY86, KS60, Sem85]. Bernstein [BK71, Fel87, Pro68c, Pro68d, Pro68b, Yur70, Zol68b]. Berry [BG96, Gri94a, Zol66c]. Berry-Esséen [BG96]. Bessel [DSS94, Pro68a]. Best [Nik77, PS73, PS76, Sap79, Zer70, Zol64b]. Beta [Vol70]. Beta-Distribution [Vol70]. Bethe [BG91]. Between [Gam91, Pet66b, Sid57, Top67, Val74b, Vin99, Vol67, Amb65, BR85, Gam68, Gam92a, Gap89a, Gel74, GP79, IR70, Kar65, Mak65, Mik58, Muk96, NV85, PP96, Pos61, Val73, Val82, Vol65, Zub75]. Beyond [GS88]. Bhattacharyya [Sha80b]. Bias [Cle98, Kos88, Sir56]. Bibliographies [Gen59]. Bibliography [Ano58a, Ano58b, Ano58c, Ano59d, Ano61c, Ano62b, Ano64b, Ano65b, Ano66e, Ano66f, Ano67d, Ano67e, Ano68d, Ano68e, Ano69d, Ano69e, Ano70e, Ano71d, Ano71e, Ano72b, Ano73e, Ano73f, Ano74f, Ano75g, Ano75h, Ano76c, Ano77e, Ano77f, Ano78e, Ano80d, Ano81f, Ano81g, Ano82f, Ano84f, Ano86i, Ano87h, Ano88i, Ano88j, Ano89h, Ano89i, Ano89j, Ano90i, Ano91g, Ano91h, Ano91i, Ano95d, Ano96c, Ano96d, Ano97e, Ano97f, Ano98b, Ano98c, Ano99c]. Bicompact [Klo59, Klo61, ZK76, Mak67a]. Bilinear [SB67]. Billiards [Bun74]. Bimatrix [Vor58]. Bin [Sev62]. Binary [Dob62b, Kis62, Mar69]. Binomial [BM79, Ivc75b, Kov59, KS95, Lut90, Lut91, Nik80a, Pre84, Shi86, Sir56, Sor78, Tim99, Tsa58b, Tur91, Vol69, WW92, YTS89, Shi97]. Biology [Ano59b, Ano60a]. Birth [But92a, Eag69, Kuz74, Ano97d]. Birthday [Pro91b, Shi93, Kol62, KKSY95, KGPS91]. Bistochastic [Mak70a]. Bivariate [MW59]. Blackwell [Bor64a]. Blends [UU87]. Blocks [Gir96]. blockwise [Gap95]. blockwise-independent [Gap95]. blockwise-orthogonal [Gap95]. BMO [Pav80]. Bochner [PS61]. Bol’shev [Ano79a, KC79]. Book [Ano66a]. Books [Ano66b, Ano66c, Ano67a, Ano67b, Ano67c, Ano68a, Ano68b, Ano68c, Ano69a, Ano69b, Ano69c, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano72a, Ano73a, Ano73b, Ano73c, Ano73d, Ano74b, Ano74c, Ano74d, Ano74e, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79b, Ano79c, Ano79d, Ano79e, Ano80a, Ano80b, Ano80c, Ano81c, Ano81d, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano84a, Ano84b, Ano85d, Ano85e, Ano85f, Ano85g, Ano86a, Ano86b, Ano86c, Ano86d, Ano87f, Ano87b, Ano87c, Ano87d, Ano87e, Ano88b, Ano88c, Ano88d, Ano88e, Ano89d, Ano89e, Ano89f, Ano90d, Ano90e, Ano90f, Ano91b, Ano91c, Ano91d, Ano92a, Ano93b, Ano95a, Ano95b, Ano97a, Ano97b]. Boolean [Bal74, Bal96, Kov67, Koz71b, Mas99, Mol92]. Bootstrap [LS96, Tho93a]. Borel [BS74, Kuz81, Pus92, Sob70, Yus76]. Boris [Ano94a, Shi93, V^{+}97]. Borokov [FM92]. Borovkov [Ara76, AS74]. Both [BR94, SKKM95a, SKKM95b]. Bound [Ali87, Ara83, Bag88, Bag89a, GO92, Kha79a, Mat85, Mat84, Mat86a, Mat86b, Mik91, Sun99, AP97, Bak89b, Kha79b]. Boundaries [Bev58, BK65, Leb71, Lot80a, Lot80b, Mol67c, Mol67b, Rag91]. Boundary [AR94, Ali92, Anu80, Bei65, Bel68, Bon92, Bor82b, BR64, Bor67a, Dok92, Dyn67, Fre63, Fro89, GR84, GS88, GP85, GN87, Kha58, Kif73, Kog65, Kog66, Kor74, KSS77, Kor90b, KC83, Kry66b, Mai75b, Maz69, Mil96, Mog74, Nag70b, Nag70c, Nag87, Nag85, Nov83, Pit69b, Pre67, Rag86, Rag93, Rog66a, Ron86c, Rot67, Roz60b, Roz88, Roz91, Roz96, Roz67, Sah75, Sap92, Shu68, Shu69a, Shu69b, Shu70a, Shu73, Shu79a, Shu82b, Sko64a, Ven59, Ver82, Mil97, PS98, Shu81]. Boundary-Value [Bor67a, Fro89]. Bounded [Afa65, AM69, Ali96, Bev58, Bla83, Bor81a, Bor60b, Bor60a, Bor61, Bul75, Fre66, Gru73, Gru76, KT69, Kuz83, Lar94, Lin59c, Mai75b, MR96, Nov82, Rog88a, Rog88b, Sac73b, Saf77, Sev96, Sie80, Sko61b, Sko62, Son92, Zol66c, Zub73]. boundedness [Bau97]. Bounds [Ark87, Ark88, Ark90, Bag89b, Bor96, Bul75, Dau88, Fer98, Gap92, Gol92a, IK82, Jur72, Jur74a, Jur74b, Mal74a, Mal77, Mir90a, Ost91, Pav81a, PU90, RR99, Rug93, Ryc84, Str78, Ver88a, Vol78a, Vol79, Vol80, Vol81, VZ80, V’y88, Whi60, Yur93b, Zue88, Lep92]. Boxes [Chi64c, CV65, Sev67b, VC66]. Branch [Ano86k, Ano87l]. Branching [BR79, BR84, BZ84, BM90, BM91, BD97a, BD93, Bev58, Big93, Bor86a, Bor89b, Bor95, Büh68, Chi57, Chi59, Chi61b, Chi64d, Chi65b, Chi70b, Chi70c, Chi70d, Chi71b, Chi72, Chi73b, Dit91, ES75, GC87, Gri87, Gri91b, Gri92b, Gri92c, HSVJ67, Jag68, Jan76, Kal82, Kal99, KN94, Kav91, KW71, Kha68, Kha69b, Kha69a, Koz77a, Lab72, LN68, Lam72, Mai75b, Mai79, Mak82b, Nag67, Nag75, NK81, NK82, Pol78, Rak87, Rak88, Rak90, Rak93, RS70, Sag83, Sag86, Sag91, Sag96, SG62, Sel69, Sev57b, Sev61, Sev64, Sev66a, Sev67a, Sev68, SC71, SZ74, Sev90, Sev96, Sev99, She87, Shu77d, Shu80b, Sko64b, Top83, Top89, Vat74, Vat77a, Vat77d, Vat77b, Vat77c, Vat78, Vat79, Vat80]. Branching [Vat81, Vat82, Vat87a, Vat87b, VS90, VS92, Vat94b, Vat95, VW99, Vin64, Vin88, Vol86, V’y77, V’y78a, V’y78b, Yag68, Yak81, Yak86, Yan76, YM84, Zol57b, Zub72, Zub73, Zub75, Zub83, GP97, Sev59, VS89]. Bridge [KS87]. Brobability [Ano81a]. Brownian [Tou92, Amb65, Ara99, Bor90b, Bor90a, Che57a, DV96, Doo57, Dzh94, Fer98, GV59, Hid83, IK64, Kif76, Kir94, KS87, McK63, Mol67a, Nas88, Ouk91, PS98, Ron86c, Roz73, Ski56, Sko61a, Ven68, ZI70]. browniens [Tou92]. Bruijn [Mór93]. BTL [Glu73]. BTL-Systems [Glu73]. Bucy [Lip76]. Burgers’ [BM92, LO96]. Busy [Bas58]. Busy-Signals [Bas58].

C. [Ost69]. Calculating [Koz81, Nag86a, TC86]. Calculation [BGS61, Chi86a, Chi64c, Kah93, KS87, LS59, LNF89, Val73, Val74b, Val82, VTGM85, Vla56]. Calculus [DD99, Smo89, JK90].

Calibrated [BRR98]. Calls [VP64]. Canonical [Eag69, IR71, Sir73, Sir78, Vak94a, Gir95, JK90]. Cantelli [Saz63, Shi86]. Capacity [Kab78, Pre89, Sev62, Tsa58a]. Carleman [Iva70]. Carlo [Bor92b, EZ60, Erm66, Fet74, Koz73, Koz81, Mik70a, Sob60, Tur71, VTGM85, Vla56]. carries [She97]. Cartesian [Ost83]. Case [AS98, Ark88, Ark71, Bab85, Ber81, Ber87, Bor79b, Che70, DJ99, Dyc73, Ebr71b, Epa68, Fed71b, Gal78a, Glo79b, Gur92, Gus76a, Gus76b, Had78, HP94, IK73b, IK74a, IM79, Ivk68, Kal78, Kol67, Kol88, Kra67, KMT78, Kut78, Lin59b, LS84, Lip97, Mal90b, Mna86, Mos73, Mos91, Nag68a, NR73, Oni90, Per59, Pro68c, Rod88, Rot76a, Saz65, Tik92b, VT92, Vin99, WW68, Wol76, Yus76, Mak88]. Cases [Par70]. catastrophe [KY97]. Category [Nat73]. Cauchy [Bla64, Das94, Fel92, Kha60, ST74, Sar90]. Celebration [Shi93]. Cells [Bol68, Kol71, Kol76, Mik80, Mik98, Sev66b, VM83, ZM79b]. Cellular [Ose90]. Censored [Góm93, Tik85, Tik98, Tim84]. Censoring [Tik92b]. Censorship [Csö84]. Center [Ano95c, Sen87, Ste71]. Centering [FTV92, PR85]. Central [FLR97, Fra93, Kud84, RS98, Rog60, Rot70a, Ryc84, Sei87a, Ali87, Ali90, AR86, Bha70, BR65, BZ77, Bul96, Bun74, CPU93, CPP98, DEO84, Dob56a, Dob56b, DM80a, Eli89, Eng84, Eng85, EO86, Fom83, Gal71, Gam81, Gap70, Gil82, Gir80, Gir82, Gir98, Gol83, Hat75, Her91, Her92, Ibr63a, Ibr67a, Ibr75a, JM94, KS64, Kel90, Kha94, KA99, Khr91, KR92, Kol92, Kru82, Kub96, KT77, LM71, Lët93, Lif76a, Lif76b, Lif79, Lin59a, LS81b, LS81a, LS82, Maĭ90, MA83, Mal87a, MO83, Mat85, Mat84, Mat86a, Mik76, Mik94, Nai82, Nor93b, OP67, OR85, Par64, Pau77, Pau79, Pet66b, PL83, Pit78, Pro92, Rez81, Rie77, RS76a, Rot70b, Rot78a, Rot78b, Roz60a, Roz61, Roz76]. Central [Roz78, Roz79, RS76b, RS80, Sad66b, Sad67b, Sad68, Sal79, Saz67, SM79, Sem85, Sen81, Sen82, Sen84, Sen86b, Sen86a, Sen87, Sen90, Sen92, Sen93, She80, Shi80, Shi82, Sho79, Sho83, SF84, SI67, Sun96, Sun99, Tik81, Tom71, Ul’77, Ul’79a, Ul’79b, Ute91, Ven90, Vir70, Vol96a, Zai87a, Zai88a, Zal86, Zol66a, Zup92]. Central-Server [Sei87a]. Century [Gal95]. Certain [AB85, Ale91a, Ben76b, Bor85d, Chi71a, Cra65, DR78, Gab75, Gär76, Gir60a, Gor85, JR89, Jur72, Kal76, Kif75, Leh65, LTC80, Lum71, MN77, Mog74, Mus73, Nag76b, Nev73a, Nhu68, Oli93, Pau77, Pau79, Saf92, Sko64a, Sla78, Ste69b, Vin95, Vir74, Yur71a, Yur71b, Yur72, Zhu70, Zol61a]. Cesaro [Gap65]. Chain [Amb79, Arn81, Bel67a, Bra91, Chi76a, Chi76b, Cow93, Fai79, Fai80, Fil84, Fri69, Gik58, Glo70, GW98, Gri79, Kry65, LK74, Mak80, Mih56, Pre67, Sam82, Shp68, Son92, Vas70, Vol58b, Vol60b, Zas73]. Chains [Ash68, AS82, AN93, Bel65, Bel69b, Bel78, Bez75, Bez78, Bez71, Bez80, Bez82, Bor81b, Bor91a, Bor92a, BK97, Chi83, Coh62, Dav74b, Dob56a, Dob56b, GR67, Goo59, Gur68, Ibr75b, Iva65, Kar86a, Kar86b, KS61, Kon70, KS84, Kor94b, Lif76b, Lif79, Mak65, Mak70a, Mal87b, Mal90a, Mal92a, Mes58, Mit88, Mol67c, Nag57, Nag61, Ose65, Pit69b, Pop80, Sar61, Shp66, Shu58, Shu82a, Shu85b, SF84, Sko63a, Sko87, Tsi76d, VMS95, VJK59, Vol57, DRS89]. Chance [CKR68, Khi56a, Khi56b]. Chance-Constrained [CKR68]. Change [Ali92, Bor99, CL94, Dar85, DB88, Dar96, Kal93a, Kha84c, Leb96a, Maĭ91, Yak87]. Change-Point [Bor99, CL94, Dar96, Maĭ91]. Channel [Dob58b, Dob62a, Dob58d, Glo79b, Kab78, Kat75, Pre89, Sei82, SS65a, Tsa58a]. Channels [Ano59c, Bor65c, Dob63, RR64, Tin62b]. Chaos [Bor98, Tan70, JK90]. Chaotic [Asa91, Vov88]. Characteristic [Ara81, BS73b, Bik69a, Bor85d, Bor86b, Bra86, Csö84, Gab75, Gam92b, Ibr94, Il’76a, Jur72, KS87, Koz71b, MSV99, Mkr81, Ost89, Pav82, Sad66a, Sad67a, Saz58, SE96, Sgi99, SK92, Tar90b, Tri90, Ush82, Ver64b, Vin90, Vir84, Yur71a, ZL64a]. Characteristics [BN69, BN96, Bur75a, Dar85, Dau88, GV93, IR70, Ivn77, Kru74, Lot86, Nos88, PR69b, ST73, Sen68, Ste69b, Ven86]. Characterization [BD97b, BL75, Bet86, Bol65a, BU84, BMS96, CPU93, Das94, DR78, Ebe86, Fel88, Fel92, Fel99, FM92, Fry89, Fry90, KS67, Kha65, Kle74c, Kle81, KN90, LTC80, Mis91, Mkr81, Mus77, Nik85, PS85, Plu85, Pra99, PF57, Pro65a, Pus93, Rao76, Rie80, RS82, Sha80b, Ski56, Ste94, Tro93, Vak94b, Vis86, Vla84, Vor61, Yan85, Zin80]. Characterizations [Bon74, CS71, Nhu68, Nik91, Pla96, Ruk76, Yan89]. Characterizing [Bra88, IC98, Kle79, PS92a]. Charter [Ano90b]. Chauvenet [BU75, IK79, Pag86]. Chebyshev [Ano97d, BO96a, Ibr67c, Whi58]. Chebyshev-Cramér [Ibr67c]. Chebyshev’S [Ark71]. Checking [BP73]. Chentsov [BBD^{+}94]. Cherkasov [Shi64a]. Chernoff [FM92, Pra93, Ron86a, Sal93b]. Chernoff-Type [Pra93]. Chernov [Sal99]. Chevchenko [Ano80h]. Chi [Bak96e, BM79, Chi71a, Erm98, GR85, Mir90b, Nik74, Ron85]. Chi-Square [Bak96e, Chi71a, Erm98, GR85, Mir90b, Nik74, Ron85]. Choice [GZ66, Nik77, PS73, PS76, Sed77, Tsy88]. Choosing [Ger73, PT93]. Chung [Mar87b, Mog80]. Church [Pos61]. Circle [Bra94, Nev64, Tep95, ZI70]. Circular [Gir85]. Class [BN83, Ben76a, Ben76b, Ber84, Bor65b, Bor78c, Bor79e, Bul84, Bun74, Chi74a, Chi95, Das79, DT75, ES75, Fel85, Fel86, Fok77, Fre66, Fre73, Gär76, Gir60a, Hau93, HJ78, Ibr57, Ibr63a, Ibr65a, IK83, Ios68, Ios70, IL76b, JM94, Kag88, Kak78, Klo62, KS86, Kov59, Kov67, Kov86, Kru73b, Kud84, Kud86, Lev76, Mak95, Mal91, MP58, Mik65, Mik70a, PR99, Pet59b, Pit84, Pro65a, PS89, Qua99, Rac85b, Rao89, Saf92, Sha70, Shk74, Shu67, Shu68, Shu77b, Son67, Sre70, Teu75, Tim84, Ula84, Vat81, Vat82, Yak88, Yam96, Yus74, Zal86, Zan99, Zez75, Zhu70, Zin65, Zol61b, Zol64a, Bor77c, Mak88]. Classes [Alb93, Ber85, Ber86a, Ber86b, Ber87, Dob57, Eid73, Fai87, Faj88, Gap78, Gor85, Iva70, Kag89, LTC80, Lu’93, Rez68, Rez71, Rog65, Sza73, Vla84, Yag57]. Classical [BU75, DM67, IM65b, IM66, Kol66, Kol67, Mac71b, Rot76a, Rot78a, Rot78b, Sel99, SC64a, SC64b, Sev67b, Sho79, Tho91, Wol76, Zol92b]. Classification [AL81, Bol57, Kha84a, MS79, Nas57, Ser79, Vol78a, AP97]. Classifications [Vol78b]. Classifiers [Kha84b]. Close [Chi64b, Khm75, Mak82b, NR73, NS89, Pit83, RU88, V’y77]. Closed [Ivn98, Mol88, Mol91, Mol92, Sei87a]. Closedness [RT98]. Closely [Vol67]. Closeness [Bas93, Gam84, Kab86, Mir86a, ZS84, Zol65b]. Closest [Zub76]. CLT [RR99, Vak94b]. CME [Yam96]. Coalition [Vor60, Vor67]. Coalition-Free [Vor60]. Coalitional [Vil68]. Code [Kis62, Koz81]. Coded [Kis62]. Codes [Dob62b, Dob63]. Coefficient [Gel74, Kut89, Por75a, Por75b, Sar70, Sha80a, Tik82, ZI74]. Coefficients [Aly88a, Aly88b, Anu91, Bra94, EK72, Fre64, Fre67, Gir92, IK99, KF91, Kho69, Kho71c, Kle85, Kry91, Kul83, LR63, Mal69b, Mat65a, Mat65b, Mel79, PS93, Por79, Pur89, Ras77, Roz99b, TP59, IM71a, IM71b]. Coexcessive [Shu69a, Shu69b]. Coexistence [Dob73]. Coherent [BRR98]. Coincidence [MS88, Tor88]. Collection [Ber74]. Collectives [Pos61]. Collinearity [Le93]. Colloquium [Ano63b, Ano80k, Ano81a, Ano82a, Ano88g, Ano89k, Ano90j]. Combinations [EN75, Gri89, Gri94a]. Combinatorial [KC74, Sol66, Ste69a]. Combinatoric [Cow93]. Combinatorics [Ano90b, BBM59]. Commemorating [Bol81]. Commemoration [Ano96a]. Comments [Pit69b]. Commercial [Mor75]. Committee [Ano90b, Ano90h]. Common [Kar86a, Kar86b, PG80, Roz90c, Roz98b, Tom90, Zub76]. Communication [Bor89a, CPP98, Cle98, Dob62a, Dyn98, Fer98, Ivn98, Jak98, Mik98, Pis98, PP95, Pre89, RR64, SZ98]. Communications [Ano97g, Ano97h, Ano97i, Ano97j, IS98, Ost98, Rad98]. Commutative [EM68, Gol83, Rok61]. Commutativity [VJK59]. Compact [EM68, Fai76, Fel79, FF82, Kai86, Kap74, Klo62, Mak71a, Mak73, ML65, MR89, Mus85, PS64, Pes94, Rok61, Sav99, Shl81, Shu73, Sko74b, Min91]. Compactness [Ers79, Kru99, Leb81, Leb82, Nik80b, Shu92, TTC78]. Compacts [Mol96]. Company [Vin99]. Comparability [Fal89, KFS77]. Comparative [BD91b]. Comparing [Fin93, Lin64]. Comparison [Aiv59, Bak96b, Bar89, Ber70a, Ber70b, Bol62, BP78, DP82, Gal83, Ing88, Ivc75b, Kle86, Kra94, Lip66, Pag68, Pit82, Pol63]. Comparisons [Glu73]. Complete [Ber86a, Ber86b, Ber87, Kon79, McK93, She88b, Tim84, Yus76, Zez75, Ver99b]. Completely [Vas80]. Completeness [Ale92a, DS85]. Completion [Pla96, Shp68]. Complex [Ara68, Fal88, Fel65, Mak91, Mak96a, Mit88, Ruk67, SK93, Szt93, VK97]. Complexes [Haf73, IM66]. Complexity [She93b, V’y88]. Component [Che70, Chi86b, Dub91, ES69a, ES69b, Leb74, Mai75a, Sem90]. Components [Bas77, Bor61, Bor81b, Glo70, Kor66, Kor68, Lip67, LS68b, Mik75b, Pav89a, Sem90, Sol67, Son92, Vla87, Yan79, Zol68a]. Composite [Ber81, Ber85, Ber86a, Ber76, Dzh77, Ing83, Ing84, Ing86, Pav88, Pav91b, Sil73, VN99, Mak88]. Composition [Dor90b, Ibr56a, Kho83, Mak71a]. Compositions [KT59, Tut62, Tut64, Vir74]. Compound [Bel65, Bel67a, Bel69b, Bor76, Kor99, Kor74, Kor76, Pir68, Tuc63, Zub87]. Compounding [Tuc63]. Computable [Muc88]. Computation [Mar76a, Mar76b, Pis65, Tur71]. Computational [Pro83]. Computing [KC83]. Concentrated [Ost89]. Concentration [AK95, Kru93, Lar94, MR80, MR83a, Mir90a, Pet70, PY81, PY78, PY79, Rog61a, Rog88b, Rog94, Sal97, Saz66, SU90b, Zig82, NK97]. Concentrations [Maĭ95]. Concept [ML66, RR64, Shu62]. Concepts [Pos61, Shi95]. Concerning [Ano64a, Bul64a, Her91, Kif75, Mol67a, Ros75, Sol66, Tak57, VTF73, Zol61a]. Condition [Dob58a, Gor83, Gri83, Gur89a, Gus96, Ibr65b, Ibr70, Ivn82, Ivn90, Kag90, Kak78, Leb77b, Leo86, Mir74, Mir75, Nag69a, Nag69b, Nag69c, Plu85, RS93, Sac71b, Sta84, Zub73]. Conditional [Afa79, Afa91a, Ave72, Bez75, Bez78, Bez71, Bez80, Bez82, Bor78a, CS60, Cow93, Dob68, Dob70, Döh81, Dok95, DE77, Evs87, FV61, Gur92, Kif73, Kni66, Kri64, Kry73b, Kud85, KST61, Pus93, Rén56, Rom87, Ros67, Roz73, Roz80, Shi67, Sid57, Str60, Zie92]. Conditionally [Cho84, Kha61a]. Conditioned [LN68]. Conditions [AL77, AR94, Ale63, Ber76, Bla83, Bor81a, Bor74a, But92b, Che89, Dav74b, Dob68, Dob64, Dyn67, Eag77, Efr80, Fak69, GK84a, Gur68, HP94, Ibr65b, Ibr70, Ibr84, IK93, KK82, Kha67, Kle74b, KR60, Kor80, Kra74, Kra90, Kul83, LO96, Leo78, Lin59b, Lin59c, Lin59a, LS81a, Mak68, Mal80, Mal85a, MV89, MV94, MP77, MP78a, Mar80, Mir78a, Mir78b, Nag71a, Nag73, Nem91, Nov80, Pit69b, Sak56, SL93, Ser61, Sev61, Shi63a, Sho96, Sho99, Shu79a, Sil74, Sko64a, VC82, Vat87b, Ven59, Ver66, Ver64a, Ver82, Vla80, Yan76, Mil97]. Cone [Pav81b]. Cones [Nos83a, Nos83b]. Conference [Ano59a, Ano59b, Ano60a, Ano60b, Ano61a, Ano64c, Ano65a, Ano74a, Ano78f, Ano90c].

Confidence [Ara68, BO98, BP91, Bog80, Bol65b, BL63, GR67, Kah93, Lum69, PG80, Pag83, Pav81a, Tho93a, VS91]. Configuration [Dob64, Smo89]. Configurations [Smo91]. Conflicting [Fed77, Yak91]. Confluence [KS57]. Congestion [Bas60]. Congratulations [Iva92, Pro91b]. Congress [Ano88a, Ano93a]. Conjectandi [Yus87]. Conjecture [Mar87a, Pav80]. Conjugate [Nag99, Sch79]. Connected [Ark65, CPU93, ES77, Fak82, Gol65, Gri74, GZ69, Ivc76, Kan88, Kle81, Kuz75, Mih56, Nev73b, Pit71a, Shu62, Tak76, Tsi76b]. Connectedness [Kel73, Ste70a]. Connecting [Bol57, Kot64]. Connection [Amb65, Gam68, Gam91, Gam92a, IR70, Kag88, Pos61, Pro92, Top67]. Connections [Ano59c]. Connectivity [Ivc74a]. Consecutive [VIL69]. Consequences [Bag92]. Considered [Amb76]. Consistency [BCK^{+}92, CDM93, Gap86, Hau93, Lev86b, Pes94, VV88, Zer70]. Consistent [Kop96, Kul85b, Mol92, Sko74a, Vor62, Vor63b]. Constancy [Fel99]. Constant [IS98, Ign82, Kle80, Kor94a, Lep94, VTGM85, MP97]. Constants [Kla85, Zol64b]. Constrained [CKR68]. Constraints [AM67, Fri72, Kar69, KNS94, Piu94, Piu98]. Constructed [Kud84, Tik98]. Constructing [Nik88a]. Construction [Ber80, Bol65b, BG73, Elb76, GR67, Gur65, Kon70, Kry65, KH81, Kuz74, Kuz78, Ter93, Vol61a, Zer87]. Containing [BP91, Chi74a, Vat94a]. Context [AT93, MP93b]. Contiguity [KK95a]. Contiguous [Ber86a]. Continual [Pol61, Tep95]. Continuation [FV61, Her92, RS76a]. Continuity [Akh80, Akh81, Ano84g, Ben76a, Bor73a, BS89, But92b, Dav89b, Dob58a, DM67, Ers72a, Gri74, JR89, Lif83, Lod71, ML65, Sei82, Ser61, Shu80a, Sko65, SY73, Sko74b, Sol94, Top67, VTT81, Ven68, Ven61, Yam95, Yur74, Zol75, Zol76, Zol77b, Tsi97, Zol78d]. Continuous [Aga80, AT86, CN87, Che60, Chi71a, Dam65, Dyn59, Dzh95, Eng75, Ers79, Fak70, Gap89a, Gir60a, Gla63, IP89, KS69, Kel77, KS61, Kha77, Kif71b, Kli62, Kog68, Kor78a, Kra71, Lot96, Mai79, Mak67a, Nas96b, Nev73a, Nik74, Nik80b, Nov80, Oza86, Per59, Pit69b, PR85, RK79, Sat96, Sko63b, Sko66, Smi69, Sok63a, Sok63b, Su99, TV98, TP57, TP59, Ven65, Vol59, Whi58, YF79, Yus78, YF80, Yus91b, Zub73, SKKM95b]. Continuous-Time [Gap89a, Pit69b, Yus91b, Zub73]. Continuously [RS70, Shu91a]. Contour [KT69]. Contoured [Kur87]. Contours [Mal75, Sac73a]. Contraction [BS87]. Contrast [Chi74b]. Contributions [Luk68]. Control [Aga80, Ano90c, Anu80, Bul64b, Che64a, Che64b, Che67b, Che70, Fai76, Fai80, Fri72, GE91, Kog68, Kog69, Kog70, Kry72, MP70, Man64b, Man67, Mik56, NP99, Pit71a, Piu94, Rya76, Saf77, VM91, Wit71, Yus76, Yus88]. Controllable [Kit82, Yus80]. Controlled [Ali96, Bar77, Chi76a, Chi76b, Dyn65, Eng71, Evs89, Fai76, Fai83, Fed69, Fed70, GK76, Glo79a, Gor85, GE91, Kit86, KS86, Kry65, Kry87a, Kry87b, Mih81, Piu98, SZ74, Yus76, Yus78, YF80, Yus81, Yus82, Yus90, Yus91a, OK97]. Controls [Yus90, Yus91a]. Convergence [AL77, Akh80, Akh81, AK95, Ale74b, Ale89, AS83, Ali87, Ali90, Ara82a, Ara82b, Ara83, AR86, Bar88, Bau88, Bau96, BPRS94, BJL98, Bez82, Bha70, Big93, Bor77a, Bor78a, Bor84a, Bor82b, Bor67c, Bor67b, Bor74a, Bor74b, BS74, Bor78d, Bor80b, BS81, Bor81b, Bor83b, Bor85b, Bor85c, Bor86a, Bor91c, Bul81a, BS88, BS92, BV88, Bul96, Cad99, Che56, Chi70b, Cho84, Dau88, Dav68, Dav81, Dav82b, Die77, DF93, Ego73, EN75, Ego80, Fäh83, Fal88, FTV92, Fom83, Gam66, GR78, Gap68, Gap80, Gap82, Gap98, Gap73, Gil82, GH93, GK84a, Gor84a, Gor76, Gor78a, Gor80, Gor84b, Gri89, Gri63, GM84, Gri96b, Gru86, Haf73, HP94, Haz96, Her91, Her92, IMS93b, IMS93a, IM87, KLS84, Kal78, Kap74]. Convergence [Kar75, Kar80, Kel77, Kha77, Kha85, KA99, Khm81, Kho83, KR92, Kle96, Klo64, Kni66, Kol66, Kol56a, KP83, KP84, Kon82, Kor86b, Kor95, Kor96, Kor98, Kor99, KB89, KB92, Kor78a, Kor80, Kri64, KR90, Kru71b, Kru74, Kru86, Kru89, Kru92, KZ96, KS89, Kub96, Kud98, Kul83, Lev73, Lif76b, LS82, Log91, Lya87a, Mak66, Mak67b, Mak68, Mak70a, Mal94, MU83, Mal85b, Mar85, Mar65, Mat85, Mat84, Mat86a, Mat86b, Mik75c, Mik78, Mil88, Mir77, Mir81, Mna88a, Mna88b, Mol96, Mos91, Mus80a, Mus70, Nad74a, Nag68a, Nag70b, Nag70c, Nag70d, Nag71b, Nag86b, Nag85, Nak81, Nas57, Nev71, Nev72, Nos83b, OR89, Orl75, PN92, Pan94, Pas99, Pau77, Pau79, Pec74b, Pis80, Pop80, Pro56, PS89]. Convergence [Rac86, RR93, Rad98, RR99, Rog66c, RS76a, Ros80, Ros95, Rot70a, Rot70b, Rot78a, Rot78b, RS93, Roz81b, Roz90b, RS76b, Sah75, Sar90, Sat73a, Sat73b, Saz67, Saz69, Sem88b, Sen81, Sen82, Sen84, Sen86b, Sen87, Sen89, Sen90, Sen92, Sen93, Sev67b, She80, She71, Sho96, Sho99, Sho79, Sho83, Sie87, Sil74, SM62, SF84, Sko80b, Sla85, Sla87, Su99, Sun99, Ten91, TT82, Tie77, Tie80, Tik81, Top67, Tou94, Ul’77, Ul’79a, Ul’79b, VC71, VC82, VJ64, Vro99, V′y98a, Whi64, ZA84, Zai85, Zal86, Zas73, ZKC91, Zup92, Min91, Nad73a, Tou92]. Convergence-Rate [Saz69]. Convergent [Cho84, Kal67, Mak71a]. Converse [Mar81, Smi90b]. Convex [Bul67, DV96, Kar75, Khr86, Mol96, Pav81a, VT78, Vla84]. Convexity [Nos91]. Convolution [Den92, FF82, Haz96, Lin57a, Mak66, Mak68, Ost65, Ost67, Par67, Rog88a, RU88, Tit82, Yan89]. Convolutions [AK95, Ara81, MR89, Rog88b, RS93, Sat95b, Sie87, Ula88, Yur78b, Yur78c, Zai81, Zai92, Zol78b, Min91]. Coordinate [Che60]. Coordinate-Homogeneity [Che60]. Coordinates [She86, Tut73]. Córdoba [Ibr89]. Corners [Nos83a]. Corput [Yur71a]. correct [AP97]. Corrected [KR87]. Correctedness [Kor86a]. Correction [Bor65a, Cle98, Rya66, Vin94]. Corrections [Ano59c, Ano85b, Rya76]. Correctness [Che81]. Correlated [Gad90, Gap86, Gla63, Hur75, HR97]. Correlation [Gap82, Gel74, Hur75, Mik75a, Ryz71, Ryz72, Sar70, SO85, Vas70]. Correspond [Sko57b]. Correspondences [DE77]. Corresponding [Ale63, Apo67, Dau76, Ers72a, Gor84a, Kra74, Lod71, Nik80b, Roz63, Roz64a, Roz73, Sko60, SY73, Sok84]. Corrosion [Nos77]. Cost [Bul67, Elb76, Kit82, Kit86, Raf90]. Cotype [Küh81, Mus80b]. Countable [Bez78, Bez80, Dob56c, GS85, Gri79, Kel90, Lin59c, Pop80, Vol57, YF79, Yus78, YF80]. Countably [Vla79]. Countably-Additive [Vla79]. Counter [Tak56]. Counterexample [Pav80]. Counterexamples [Anu91]. Counting [Gru86, KLS84, NV85, PN92]. Counts [Cow93]. Coupled [Rez77a, Rez78, Rez80]. Coupled-Mode [Rez77a, Rez78, Rez80]. Coupling [Nor93a, Tho91, Tho93b]. Covariance [Bal94, Bar85, CPP98, OT95, Sen86b, Sen87, Ser96, VT78, Vak94a, YK86]. Covariances [Bul75]. Covering [Mik97, Rév94]. Cox [Kor99, Ter93]. Craméer [IR71]. Cramer [Kho74, Lin56b, Nag69a, Nag99, Bol61, Chi70a, Ibr67c, IL94, Kag88, Mal68, Nag69b, Nag69c, Roz99b, Sha61, She88a]. Cramér-Type [IL94]. Creation [SS90]. Criteria [Aiv59, AI99, BS68b, Che60, Fai83, Fed71a, Gap78, Gap97, GS65, Hau93, Kru93, Mar85, Mus80a, Shu80a, Ver83, Yus91b].

Criterion [Ben76a, Bor83a, Fai80, Gik57, Glu73, Kit82, Kit86, Kor78a, Lët93, Mal91, Mir85, Nak81, Tri90, Tum56, Yus82, dH95]. Critical [BR79, BM90, BY96, BPRS94, Ble86, Chi71b, Dit91, Kav91, Kol87, Koz77a, Mak81, Mak82b, MS88, Men88, NV75, NK81, NK82, Rak87, Rak89, Sag83, Sag86, Sag91, Sev67a, Shu77d, Ste88b, Top83, Top89, TV98, Vat77a, Vat77d, Vat77b, Vat77c, Vat78, Vat79, Vat81, Vat82, Vat87a, VS90, VS92, Vin88, Vol86, V’y77, YM84, Zub83, VS89]. Criticality [GC87]. Criticism [Ano59d]. Cross [Gur92, Mye93, SO85]. Cross-Correlation [SO85]. Cross-Validation [Mye93]. Crossing [GN87, KC83, Nag94, Nov83, Rag93]. Crossings [Bel67b, BJL98, Bul61, Iva60, Mal74a, Mal75, Mal69a, Mal77, Pit78]. Crump [Gri92b]. Crump-Mode-Jagers [Gri92b]. Cubic [Bag89b, Mak97]. Cumulants [Bak98, Bul75, Fry89, Shl89, Ute90, Fry90]. Cumulative [AN74, BZKR70, Mik75c, NP69, Nev71, Nev72, Sīl92, Zin71]. Curious [KK99]. Curvature [Kif76]. Curvatures [Nos85]. Curve [Ale92b, Nos79]. Curves [Cra64, Nad65, Nad70]. Curvilinear [GR84, GS88, Har91, KC83, Nov82, Rot67]. Customer [Ivn93]. Customers [Ivn93, Mal91]. Cutpoints [JP97]. Cuts [BNK96]. Cycles [Bel69a, Kol71]. Cyclic [Dub69, Gru76, Rya79]. Cylindrical [LP75, TT83, Vla79, Vla80, Vla81, Vla87, Vla96].

D [Shi64a]. D. [Fri70, Luk68]. Damage [Bag79]. Damping [Kor80]. Dantzig [Luk68]. Darmois [Fel93, Fel98, Kag88, Kag89]. Darmois-Skitovich [Kag88]. Data [AK70, BM92, Cai93, CL94, Fäh78, Góm93, IM81, LS68a, Man68, PP95, Ryz71, Ryz72, Tim84, Yus76, ZKC91]. Data-Based [ZKC91]. Davidon [BD83]. Death [But92a, Eag69, Kuz74]. Decay [Ara68, Gap82, SU90a, SU90b]. December [Ano57b, Ano59e, Ano60c, Ano61e, Ano62c, Ano64f, Ano65c, Ano67g, Ano68g, Ano69g, Ano74h, Ano75j, Ano76d, Ano83d, Ano85j, Ano85n, Ano85m, Ano85o, Ano87k, Ano87n, Ano87i, Ano87o, Ano87l, Ano88l, GIM^{+}97]. Decision [Bog82, Bun70, Fak69, Gor81, KK76, Kov59, Pit69a, Pit71a, Rom65, Ryk66, VGR84, Yus74, Yus90, Yus91a]. Decision-Making [Pit69a, VGR84]. Decisions [BP73, Chi76a, Die77]. Decisive [OCG93]. Decomposable [IM81, Khm81, VS90, Zub83, VS89]. Decomposed [Son92]. Decomposing [UU87]. Decomposition [Dam65, Fai87, Fel77, FF82, Fel83, Gal81, Lin56b, Lin57a, Pan91, Pan95, Rob63, Rog94, SZ98, Zol68a]. Decompositional [Chi87b, Chi95]. Decompositions [Abr87b, Kud76, Mac69, Mac71a, MW99, Ost65, Ost67, Sev72, Ush84, Yan79]. Decrease [Bar88, Sch79]. Decreasing [BR79, BR84, Rak87, Rak89]. Defect [Mal90b, V’y88]. Defective [Kor58, Kor61a]. Deficiency [Chi86a]. Defined [Ara90, Bor84b, Bor81b, But77, Has66, Kap74, Mak66, Mak80, Pre67, Sie80, Vin86a, Vin86b, Vol58b, Vol60b, Zas73, Pus97]. Definite [Fre68, Ros91]. Definition [Kle77, Nos80, TP58, Vil63, Vil68, Vin74, Yus60, Rad97]. Degeneracy [Lev74, Yan76, Zub73]. Degenerate [AK95, Bla64, Bor74a, Gir59, Glo70, Jur72, Kle85, KB89, MO69, Pin69, Ron82, Ron86b, Roz80, Roz84b, Son67, Vak66]. Degenerating [Dok99, Fre69, Kha58, Sar73]. Degeneration [Lev75a, Ras77, Vat74]. Degree [Bra87, Mal74b, Rot74, Yan96]. Degrees [Ivc73]. Delayed [Mir78c, Mir80b, VD93]. Delocalization [AB99]. Demand [Vit70]. Dense [Pav80]. Densities [Ale91a, Ale94, Chi80a, Gol94, HJ94, IH75b, Kha79a, Kha61c, KP83, Kuz83, Lax63, Man68, Pet56b, Rog88a, Rog88b, Ros91, Roz66a, Roz66b, Rud93, She71, SM62, Sta65, Tik88, Zol95, Kha79b]. Density [Ale74b, Ale81, Ale89, Ale77, Ara86, BO98, Bei65, BR83, Bor81a, Dha78, DT74, DT75, Dok95, Dza70, DM75, Dzh71, Efr86, Efr90, Epa69, Erm89, Erm90, Erm95, Gap80, Gin89a, Gin89b, Ibr67b, IS68, IK73a, IK83, Ibr94, Ing87, Ing95, Kha84a, Kha85, Kha61b, Kif77, Kon73, Kon79, KP84, Kos89, LS69, MO91, Mos91, Nad65, Nad70, Nad74a, Nad74b, Nad77, Nad93, Pin91a, Rag86, Ras77, Roz62, SE96, Tik85, Tsi76c, ZKC91, Zhu81, Zhu84, Bau97, Bor93, Nad73a]. Denumberable [Dyn67]. Denumerable [Dob57, KKR89]. Department [Ano60c, BGS90, MS90, MT90, Sev90, SS90, Zhu90, Gne90]. Departures [McK93]. Dependence [Dok93, Gap82, IR70, Kul83, Llo62, Pen93, Rus96, Sat96, Sen87, Ver99b, Vor61]. Dependences [PL83]. Dependencies [Leo86]. Dependency [Kab96, Sen86b]. Dependent [Amb76, Ber70a, Ber70b, Bon98, Bor79b, Bor67b, But92b, Chi72, CP91, DEO84, Eri76, EO86, Gap70, Ger73, Gri91a, Gri92a, Gri94b, Gri96a, Gri74, Her92, Ibr63a, Ibr75a, Ios68, Ios70, Ivn98, Jag68, Kag89, Kha88, Kha93, Kha94, Kru88, Leb71, Leo61, Lif85a, MA83, Mik75c, Nak88, Pet74b, Rak88, Rak90, Rak93, Ron80, Sal90, Sal71, Sar67, Sev64, Sev66a, Sev68, Sev73, She80, Sun96, Sun99, Tik81, Tik87, Tik92a, Tik94, Vin64, Yag68, Zup92, de 93, PR99, Sho97]. Depending [Bla62, Evs87, Fre64, KLP89, Lab72, Leb96b, Rya76, Sko81]. Depth [Mit88]. Derivation [Whi82]. Derivative [Fre63, Kor73, Nas88, Roz67]. Derivatives [Chi80a, DT74, Glo79a, IK81, VT76]. Derived [Coh62, Ouk91, Vol59]. Descendants [KN94, Nag67]. Describe [Bas58]. Describes [Ven86]. Describing [Dob73]. Description [Ave72, Dob68, Koz77b, LL94, Seĭ91, Tar90b, Ula84]. Design [DF93, Fed71a, Fed71b, Fre75, Koz74b, Páz68, Per92, Raf90, Sed77, SS88b, Ter93, Tsi85]. Designing [Vol73]. Designs [BG73, Cho93, MF71, Mal74b, Tsu81]. désintegration [Lan90]. Detect [McK93]. Detecting [AVR95, Gut64, PR87, Pag98]. Detection [BD91b, Bur79, BB91, Dar76, DB81, DB88, Dar96, Der93, Dob58d, Erm91, GD71, LS71, Lin66, LPS67, Mai75a, PC96, Par87, Rod88, Sal90, Shi63b, Sir63a, Sir63b, Vor65, Vos82]. Determinants [Gir80, Gir82, Gir98, Kha65, Sla78, Zhu68]. Determination [Ale64, Bog67, IS74, Jes81, Rie81, Ula88]. Determined [Dyn60, KU88b, Rao76]. Determining [Ibr78, Kar65, RU88, Tit82]. Deterministic [Bla89, KM65, Mai75a, MR96, Yus90, Yus91a]. Detween [Kha66a]. Development [Ano75a, Gal95]. Developments [BGW75, Ken87, She93b]. Deviation [Ale84, Amo80, BK97, DM80b, GO92, Gol65, Kha85, Kon78, Kon79, KP83, Kor59, MA79, Mos70, Nad77, Nad93, Nag82, Nag94, Nyb86, Osi79, PS86, PS89, Sha96, Sla79, Van83, Yur92a, Yur92b, Bol97]. Deviations [Ale91b, Ale79, Ale80, Amo88a, Amo91b, Ark87, Ark88, Ark90, BO96b, Bak96a, Bak98, Ben87, Big93, Ble86, Bor67a, BM93a, BM99, Ebr71b, Ebr71a, EL99, Fre74, GS88, Gär76, Gär77, God79, God82, IL94, Jur74a, KN75, KP84, Kon85, Kor62, Lin61a, Lin61b, Lin62, Lip97, Mak81, Mak82b, Mak95, Mak96b, Mal92b, Mal97, Mog75, Mog77a, Mog77c, Mos73, Nag65, Nag69a, Nag69b, Nag69c, Nag71a, NV75, Nag84b, Nag99, Osi73, Osi74, Osi75, Pet65, Pet68, Pin81, Pit83, Puk94, Ric57, Ric58, Ric64, Roz82, Roz86, Roz87, Roz90c, Roz94b, Roz98b, Ryc84, San66, Sau92, Sho79, Sho83, SMI90a, Sla82, SA94, Tim99, Top89, VT82, Ven77a, Ven77b, Ven80, Ven83, Ver91, Ver92, Ver99b, Ver99a]. Deviations [VG90, Vol61b, Yur71b, Yur72, Yur93a, Yur93b, Zal90b, Fat97, Ver94]. Diagonality [Sha80b]. Diameter [Nik64, Zol66c]. Difference [Nik80a, PS93, Sch93a]. Differences [IM65a, VIL69]. Different [CV65, IN78, Ivn93]. Differentiability [Bog89, BVT90, Sko57b, Sko60]. Differentiable [Mir74, Mir75, Mir78b, Rez77b, Sob73a, Sob73b]. Differential [Anu91, Ara90, BN83, Bor78b, Dor96a, Dub91, Dub74, Dyn98, EW99, EK72, Fed85, Fro89, Gin66, Glo79a, Gyö95, Has66, IK99, KS69, KK99, Kha58, Kha63, Kha66b, Khr86, Kle86, Kog65, Kog66, Kol69b, KS91, Kul74, KK96, Leb73, Leb77b, Leb77a, Leb78, Leb86, LR63, Mat65a, Mat65b, Med98, Mel79, Mil79a, Mil86, Mil88, Mil96, Mil75, Nor88, PR69b, Por75b, Por80, Roz84a, Roz88, ST74, Saf92, Sar73, Sar90, Sko81, Smo89, Sok84, Tsa77, Tsa79, Tsi76a, VT76, Ver80, Ver99b, Ver99a, Vet81, Zam70, Den97]. Differentials [LSS99]. Differentiation [Dav76a]. Differently [Mos73]. Differing [Kis62]. Diffusing [Bev58]. Diffusion [Ali96, Bab64, Bel95, Ben76b, BS68a, Bla62, Bor67c, Bor86a, Che57b, Che96, Chi70b, Chi86b, Cle98, Dok87, Dok99, Dub91, Ers70, Ers72a, Ers75a, Fal92, Fre63, Fre67, Gir59, Glo70, Isa85, Kha58, Kha60, Kha63, Kif73, Kif77, Kog69, Kog70, Kon82, KLP89, Kry66a, Kry80, Kry87a, Kry87b, Kul67, Kul68, Kul76, KH81, Kul83, Kul85a, Kut78, Kut79, Kut85, Leb71, Lev73, LS65, Lip67, LS68b, Mai75b, Mai79, Mak95, Man64b, Man64a, Mih81, Mis91, Mol64, Mol68, MO69, Nev64, Pin69, Pis80, Pis82, Pis95, Pis98, Por75a, Por79, Pur89, Rao89, Rez77b, Roz80, Roz84b, Sar73, Sev61, Sha80a, Shi64a, Sko61b, Sko62, Sko64b, Sob73a, Sob73b, Son67, Tik82, Tor88, Tsi78a, Ven59, Ven61, Ver99a, Bel94]. diffusion [Ers74, Kle97, KL93]. Diffusion-Type [Chi86b, Pis82, Pis98].

Diffusional [AL91]. Diffusions [Koc90, LN68, Tol99, Yur89]. Digits [Koz81]. Digraph [Lis76]. Dimension [Lus99, Pro91a, Sem88a]. Dimensional [Ark71, AR86, Bal94, Bas58, Bas77, BR64, BR65, BK97, But92a, But95, Che96, Dau76, Dob65, Dob66, Dob73, Dob74, Dyn59, Fom83, Gam78, GK76, Gla58, Jam78, Kag88, Kal62, KL83, Ken95, Kha61c, Kho83, Khr91, KNS94, Kor92, Kor66, Kor68, Koz74a, KMT78, Kul83, Kul85a, Lët93, Lev79, Lut78, Mal85b, McK63, Mog77a, Muk78, NV85, OR85, PRR82, PY81, Ric58, Ric64, Rot70a, Rot78a, Rot78b, Sad66b, Sad67b, Saz65, Sen81, Sen82, Sen86a, Shi80, Shi82, Shu76c, Sin83, Sol67, SU90a, Tsi83, Tsi86, Tsi87, Vak67, Ven61, Ver82, Vol57, Vol59, Yag57, YK87, Yur70, Zas67, Zin80, Zol77a, Zol78a, JK90, Mat61, QT97, Tol99]. Dimensionally [Rén56]. Direct [BR94, LS97, LSS99, Lin57b, LS68b, Mol67c, Nag67, Whi82]. Directional [Fre63, Kor73]. Dirichlet [Amb65, EW99, Fre64, Fre65, Fre66, Ign82]. Discipline [Gri91b]. Disciplines [Pav91a, Pav92]. Discontinuities [Che56, Lya87a, Sko67]. Discontinuous [AS98, Dyn58, Hur75, IH75b, Kor78b, Kor88b, Mis89, Mos91, Nov75, Tan70, Tik85, KL93]. Discounted [Piu98]. discounting [Pre91]. Discrepancies [Bar91]. Discrepancy [GP79]. Discrete [AB99, Ano87o, Bor63, Das94, Dob62a, Döh80, Dzh89, Ero58, Fai87, Faj88, Fel98, GS76, Ger79, Gur92, Kal78, Kel77, Lip97, Lod71, Mir90b, MW99, Nov79, Per59, Ryz71, Ryz72, Sav99, Sel69, Sve87, Vat77b, Vat80, Yus80, SKKM95a]. Discrete-Time [MW99, Sve87]. Discretized [Cle98]. Discriminant [Sem88a]. Discriminating [Khm75]. Discrimination [Fra93, Vol75, VN99, Zez75]. Disintegration [Lan88]. Disorder [BD91a, BD91b, Dar76, DB81, DP82, GR71, Kap91, Kha87, Mel89, Nik99a, Nov91, Par87, RS83, Ron85, Sal90, Shi65, Sir63a, Sir63b, Vor65, Vos82]. Dispersion [Leo61, Rog61b]. Displacements [Syt69]. Disposal [Mik77]. Disruption [Ron88]. Dissipative [Bel63, ZI74]. Distance [Abr77, BR85, Fer98, Jur75a, PP96, Sal99, Val74b, Val82, Vat94a, Zai87a, Zai88a, Zai92, Zub76, Val73]. Distances [NV85, Zai87b]. Distant [VS92]. Distinct [Sac71a]. Distinction [Vol65]. Distinctive [Nos79]. Distinguish [Vol67]. Distinguishable [Kol88]. Distinguished [Kha66a]. Distinguishing [Aiv65]. Distorted [BM93b]. Distributed [AV82, Bag88, Bag89a, Bag89b, BP91, Bei64, Bei65, Bet88, Bik69b, IH76a, IH76b, Kag90, Kle87, Kry68, Mar87b, Mar76a, Mos73, Pon79b, Pon86, PY81, Pro93, Ros80, She99, Sla82, Sla85, Sza73, VD93, Zin76, Bak89b]. Distributins [PS64]. Distribution [Afa65, Anu81, Ara75, Ark64a, Arn81, Bab64, BYK88, Bak89a, Bak96b, Bak96c, Bak96e, Bal68, Bal74, BNS66, Bar86, Bar88, Bas60, BL75, Bel67a, Bel69a, Bel69b, Bel78, BR83, Ber70a, Ber70b, Bla83, Bog80, Bol83, Bol94, Bol68, Bol57, BGS61, Bol64a, Bol65a, Bor60a, Bor70, BU84, Cad99, Che67a, Chi73a, Chi74b, Chi81a, Chi81b, Chi87b, Chi89b, CPU96, Dev69, Dob64, Dob58d, DK93, Dok69, Dub69, Ebe86, EH93, Erm95, Fed77, Fel85, Fel88, Fel92, Fil61, Fil62, Fok74, Fok77, Fry89, Gab75, GO92, Gam78, Ger67, Gri74, Gru76, Gur92, Gus69, GK69, Gus84, Ibr66, Ibr78, IK83, Ibr94, Ing87, Ing95, Iva77, IN78, Ivc71b, Ivn93, JL56, Jes80, Kai86, Kat78, Kha85, Kha88, Kha89b, Kho83, Khr91]. Distribution [Kle81, KN90, KT69, Kop96, Kor58, Kor61a, Kor59, Kos85, Kot64, Kov60, Kov67, Kov73, Koz71a, Koz73, KS97, KMT78, Kud74, Kud78, Kul67, Kuz83, Lei82, Lev80, Lif87, Lin56a, Log91, Mac69, Maĭ95, Mak82a, MU83, Man68, Mar76a, Mar78, Mar65, Mas75, Mas99, Mat78, Mes61, MSV99, Mik65, Mik74, Mik78, Mik98, Mir86a, Mir86b, Mir88, Mkr81, Mol91, MP79, Nad64b, Nad77, Nad93, Nag70d, Nag71a, NR73, NM76, NP78, NS89, Nag99, NP69, Nev71, Nev72, Nev88a, Nos87, Nov92, Nov97a, OCG93, Oli93, Orl75, Osi69, Osi71, Ost91, Ost98, Pag65, PRR82, PS85, Par67, Pav78, PU85a, Pec81, Pen77, Pet56a, Pin79, Pit82, Pop68, Pre86, Pro65b, Pro94, Pro74, Rag93, Ric64, Rod88, Rog64, Rog66c]. Distribution [Rog71, Rog73, Ron85, Ron86a, Roo99, Ros75, Ros80, RS82, Rot76c, Roz66b, Roz90c, Roz98b, Rud93, Ruk67, Sac71a, Sac71b, Sal98, Sal71, San66, Sch89, Sen74, Sev67b, Sev94, Sgi99, Sht65b, Sht66, Shu79b, Shu91d, Sid91, Sie80, Sin57, Sir56, Sko74a, Sla78, SL90, Smo97, Sor78, Sta97, Ste69c, Stu60, Stu68, Tik87, Tim88, Tit82, Tou94, Tru66, Tsa58b, Tsi76c, Tuc63, Tum56, Tur91, Tyu70b, Ula88, Vat94a, Vin68, Vin74, Vin76, Vin96, Vis86, Vit70, Vol60a, Vol58b, Vol69, Vol70, Whi82, XP89, YK86, YK87, YTS89, Yus73, Zai92, Zas67, Zhi63, ZL64a, Zin71, ZK77, Zol58, ZI89, Bau97, Fry90, KY97, Mak88, MP97, OK97, She97, Shi97, TC86]. Distributions [Aiv65, AT93, Ara81, AN93, BM91, Bag88, Bag89a, Bag89b, Bag89c, Bak90, BD78, BdH79, Bas93, BD97b, Bik69b, Bol62, Bol64a, BZKR70, BM79, Bon74, Bor77a, Bor90b, Bor90a, Bor60b, Bor91c, BU94, Bor96, BK97, Bra87, Bra86, Bra88, BB91, Bur75a, CdHY97, Che66, Chi64a, Chi71a, Chi95, Cho87, Cra56, DM68, Dav68, Dav81, Dav82b, Dav89b, Dav91, DJD86, Dob70, Dok95, Fel76, Fel77, Fel79, FF82, Fel83, Fel86, Fel87, Fel99, Fry76, Gam66, Gam68, GP71, Gam84, Gam86, Gam91, Gam92a, Gol65, Gol85b, Góm93, GPU98, Gri92c, Gru88, Gur92, Gus76b, Ibr56a, Ibr57, IP74, IH75b, IK81, Ivc75a, Ivc71b, Ivc75b, Ivc76, Jam78, Jes81, JP75, Jur72, Jur75b, KLS84, Kag88, Kal67, Kan88, Kap74]. Distributions [KT59, Ken95, Kha65, Kha66a, Kha89a, Kha61a, Kha61b, Khr96a, KR87, Kle74c, KR75, Kle79, KMM85, Klo59, Klo61, Klo62, Klo64, Kol66, KC75a, KC75b, Kon74a, KP83, KP84, Kor89b, Kor89a, Kor90a, Kor93, Kor98, Kor99, Kor62, KB92, Kov59, Kra79, Kra80, Kra90, Kru70, Kru71a, Kru71b, Kru73b, KA83, KA85, KA87, KU88b, Kru92, Kud76, Kud85, Kud98, Kul68, KS85b, Kur87, Lar94, Leb81, Leb82, Lif83, Lis98, Lot80a, Lot80b, Mac71a, Mak82a, Mak66, Mak67a, Mak68, Mak71b, Mak75, Mak80, MA78, Mar76b, Mas74a, Mes60, Mir88, Mir90b, Mit62, Mkr81, Muk85, NS89, Nas57, Nev73b, Nik74, Nik80a, Nos85, Nos88, Orl73, Ost83, Ost87, Ost89, OT95, PG80, Pan91, PP96, Pav84, Pav89b, PR69a, Pec77]. Distributions [Pet59a, Pet62, Pet74a, Pin91a, Plu85, Pol61, Pon79a, Pon79b, Pon86, Pre82, Pre84, PF57, Pro58, Pro65a, Pro68d, Pus93, Rao89, Rog66a, Rog88b, RU88, Ros67, RS82, RS86, Roz57, Roz66a, Roz80, Ruk76, Ryz69, Sat96, Sat95b, ST66, Sei87b, Seĭ89, Sel85, Shu82b, Sie90, Sin76, Sko65, Sre70, Sta65, Ste69b, Sve87, Sza73, Teu75, Tik85, Tik98, Tim99, Tim84, Tut71, Ugl89, Ul’87, Urb56, Ush82, UU85a, UU87, Vak64, Vak66, Vak90, VIL69, Val74b, Val82, Vat77b, Vat80, Vin99, Vir74, Vol96b, Vol65, Vol75, VP83, Yak88, Yam85, Yam96, Yan96, Yan79, Yur71a, Yur78a, Zai81, Zai84, Zai88b, Zin58, Zin65, ZK92, Zol65a, Zol65b, Zol66b, ZK76, Zol77a, Zol78a, Zol78c, ZM74, Zub74, Zub76]. Distributions [dH95, Bak89b, GP97, Val73]. Distributionsby [IC98]. Diversity [MTP93]. Divisibility [Lis94, SK92, Zem93, Zol67b]. Divisible [Bag92, BP93, Den89, Fel86, FST95, Fry76, Ibr56b, Ibr78, IM87, Jes80, Jes81, JR89, Jur75b, Khm84, Kho87, KMM85, Kru70, Kru71a, KA83, KA85, KA87, Lin58, Lin59b, Lin59c, Mak71b, Mar70, Mes60, Mes61, Mir88, Mna88a, Ost83, PS64, Par67, Pre84, Pro58, RS82, Sat95b, Sie87, SMI90a, Sko65, Stu60, Tsa58b, UU85b, Vol96b, Yak88, Yam96, Yur78a, Zai92, Zol58, Zol61b, ZK76]. Divisors [Ost83]. DNA [Cow93, Saw93]. Dobrushin [GIM^{+}97, Fal78, Soz64, Val74a]. Does [Kle80, Nag69a, Nag69b, Nag69c]. Doléans’ [Mis89]. Dolnik [Gas63]. Domain [Ali96, Anu91, BBM59, Bev58, Bun63, Das94, Kha58, Mai75b, Nys87, Roz90c, Roz98b, Tim99, Tsi78a, Tsi78b, Zin86, dH95]. Domains [Gri91a, Gri94c, GK96b, Jur83, KS84, Kru93, MR93, Zak90b]. Donsker [Bak96d, Bor84a, BB87]. Double [Ivc75a]. Doubly [Góm93, GV93, Mak65, RR93]. Drift [Afa79, Fok74, Kir94, Kul76, Kut89, Lev72, Lev73, Por75a, Por75b, Por79, Por80, Sha80a, Tik82, Tsi78a, Yus90, KL93]. Drifting [Kor86c]. Driving [Leb86]. Dstribution [Are77]. Dual [Shu78b, VMS95, Vla80]. Dudley [Jam78]. Due [Chi70a, God82, Koz73, Rog69]. Duplications [Mik74, ZM74]. Duration [LNF89, NM76, Tik92b, Vol60a]. Dynamic [Bla89, Fai82, Fai87, Faj88, Fre74, Fri70, Ose65, Pit69a, Pit71a, Sin60, Tut74]. Dynamical [BPRS94, Dau76, Gri75, Gri78, Kif75, Kor80, SS88a]. Dynamics [EK82]. Dynamo [MT85]. Dynkin [CS61, KKSY95, Pit69b]. Dyson [Nai82].

E. [Haz96, KKSY95, KC79]. Each [Kis62]. Eaton [Lis98]. Economic [EK82]. Edges [Kar95, Sev94]. Edgeworth [CvZ85, CvZ89, Kal93b]. Editor [Bor77b]. Editorial [Ano62a]. editors [Bak89b, Ul′91, Ano63a, Ano75b, Ano77a, Ano85b, Ano99a, Bur96]. Effect [KK96, Mil82]. Effective [LS68b, V′y98a, Yag60]. Effectiveness [Kha66a]. Effects [Bor95, Khi56b]. Efficiency [Erm90, GW93, Ing83, Kho71b, Lev76, Mal69b, Mal86, Nik88b, Ron86a, Vol79, VN99, Vov92, Wol65, Zhu81, Zhu84, Zue98]. Efficient [BR94, Chi86a, Dzh71, Gin89a, GK96a, Mal86, Mal90b, Nem91, Nev81, Ras77]. Effort [Ark64a]. Egorov [Tsi76b]. Eigen [Vla56]. Eigen-Function [Vla56]. Eigenfunctions [VT73a]. Eigenmeasures [VT73a]. Eigenvalue [PS92a, VT76, Vla56]. Eigenvalues [BYK88, Dub69, Gir91, YK86, YK87]. Elementary [KK95b, KD87, Vak94b]. Elements [BR85, Bul81a, De 82, Dor87, Gir83, GE93, KS64, KM89, LS97, Mak96a, Sac71a, Sac71b, Shu76b, Tie77]. Ellipsoids [Ing95]. Elliptic [Fre69, Gir86, Gor89, Kha58, Kha63, Lev83, Mil79b, PS88]. Elliptically [Kur87]. Embedded [Fai87, Naw82]. Embedding [Ibr74, MP91]. Emergence [Vol64]. Emigration [Kav91, Vat78, Vin88]. Empiric [Ral97]. Empiric-Quantile [Ral97]. Empirical [Ale74a, BYK88, Fil62, GO92, Kon79, KP83, KP84, Kon85, KC83, Kov86, KR90, Maĭ90, Mol88, Mol91, Nor91, Nor93b, Ron86a, SS92, Shi97, Tsi97]. Empty [Chi64c, Mik80, Mik98, Sev67b, VM83, VC66]. Energy [BP89b, Glo82]. Enlargement [Gal94]. Entire [IK83, Ost89]. Entitled [Ano87o]. Entrance [Maz69]. Entropic [Kol92]. Entropy [Abr59, Bas59, Dob58c, Dob60a, Ero58, Mat78, MP93b, Qua99, Rok61, RR64, Sav64, Zub74]. Entry [KS84, KS85a]. Envelope [BN69]. Envelopes [Sha93]. Environment [Afa91b, But92a, But95, DB93, Dit91, Koz77a, Lët89]. Environments [Gol85a, Lët93]. Epidemic [BD93, Cai93, Kal99, NS70, NM76, WS93]. Epidemicin [Sta97]. Epochs [Sgi92]. Equation [Anu91, Bog67, BM92, Bul87, Das94, Dub74, Dyn98, Fed85, For63, Fre64, Gir96, Gir62, Kab93, Kha59, Khr86, Kli57, Kor78b, Kry72, Kry91, Kul67, Kul68, Kul74, Kul76, Kva96, Leb84, Leb86, LO96, Lev86a, Mis89, PS93, Sal90, Sav99, Sgi99, Shu76c, Sob60, Tsi76a, Vak67, Veĭ93, Ver83, Vla56, Gir95, OK97].

Equations [AR94, AK89, Aly88a, Aly88b, Ara90, Bal74, Bal96, BN83, Bla64, Bor78b, Bor79e, Dav82a, Dor96b, Dor96a, Dub91, DK97, EW99, Ers75b, EK72, Fed82, Fre69, Fro89, Gal79, Gal83, Gin66, Gir92, Glo79a, Gor89, Gyö95, Has66, IK99, IK64, KS69, Kha58, Kha60, Kha63, Kha66b, KV85, Kle85, Kle86, Kog65, Kog66, Kol69b, Kop96, Kov67, Kry69, Kry73a, Kry85, Kul98, Kul83, KK96, LSS99, Leb73, Leb74, Leb77b, Leb77a, Leb78, Lev83, Lev86b, LS68b, LR63, Mak96b, Mas99, Mat65a, Mat65b, Med98, Mel79, Mik97, Mik99, Mil79a, Mil86, Mil88, Mil96, Mil75, Mil79b, Pis95, Pis98, Plu81, PR69b, Por67, Por75b, Por80, Pur89, Rez77a, Rez78, Rez80, Roz66a, Roz66b, Roz84a, Roz88, Roz84b]. Equations [ST74, Saf92, Sar73, Sar90, Sch93a, SC71, Shi67, Shu77b, Sko61b, Sko62, Sko81, Sok84, Tsa77, Tsa79, Tut73, Vas70, Ver80, Ver82, Ver88a, Ver88b, Ver92, Ver99b, Ver99a, Vet81, Yan89, Zam70, Den97, Kle97, Mil97]. Equicontinuous [Ros64]. Equidistribution [Shu96]. Equilbrium [Whi82]. Equilibrium [Bei64, EK82, PS76, Vil68, Vor58]. Equimeasurable [Vla96]. Equiprobable [Bel69a, IT79, VM83]. Equivalence [Asa91, BS89, Chi89a, Kra74, Kra90, Kry80, Roz63, Sok84, Str78]. Equivalent [Ego73, KK95b, Kra79, Kra80]. Equivariance [Lus99]. Equivariant [Erm89, Nik87]. Eratosthenes [Pla92]. Erdos [Bor91b, Nov98]. Ergodic [AM67, AM69, Ali95a, Ali95b, Ano89a, Bez75, Bor79c, Bor79d, CL94, Dir81, DK97, ES77, Fed77, Gap97, Ibr75b, Kha60, Kli77, Kol69b, KS84, Mak65, MS90, Nat73, Naw82, Ose65, Par71, Pus97, Sev57a, Shu58, Shu67, Shu77a, Shu78a, Shu79a, Shu81, Sin63, Sko87, Smi90b, Tem73b, VMS95, Ver91, Vol60a, V′y98a, Zäh83, DRS89]. Ergodicity [Afa88, Akh80, Akh81, Bor79e, Bor89a, Bor91a, Bor92a, Gap89a, Kar86a, Kar86b, Nev69, Ric68, Shi63a, Tem73a]. Erlang [Khi62, Ush78]. Errata [Ano94b]. Erratum [Ano85c, BK93, Fry90, Shu91c]. Error [Bab85, Dob62a, Gol75, Ibr64, IS68, Koz73, Koz81, Man68, Nad74b, SS65a, Yur78c, Yur89, Zol78b]. Errorless [Fre88]. Errors [BU75, Gad90, Gap86, Kol88, LN91, MR96, MS79, Ser79]. Esseen [Gam78, Gri94a, Sad66b, Sad67b, Sal97, Zol66c, BG96, Rog60]. Essentially [Kha61c, Zez75]. Estimate [Ara76, AK70, Bas59, BV88, Bur97, Eng84, Eng85, Fom83, Gam84, Gel64, Gil82, Glo79a, IO66, Kab86, Kha87, Kle74a, Kon78, Kor86c, Kos89, KS97, Kul85b, Kut85, LM71, Mal64, MU83, Mal68, Mik78, Mir69, Mit67, Mol92, Nad77, Nag71b, OP67, Pet66b, Pet70, PT93, Pro92, Rog61a, Rog74, Rot70a, Roz79, Ruk67, Sal97, Saz69, Saz90, Sch65, Sei87a, Sen85, Sen90, Sha70, She71, Sho79, Sir59, Sla87, Tuc63, Ul’77, Yan85, Zal83, ZSU89c, Zas73, ZK92, Zol66a, Zub74, ZM79b, NK97]. Estimates [Abr77, Ale67, Ale74b, Ale89, Ale91a, AN74, BO96b, BR83, BR85, BS73a, Ber74, Bol60, BL66, Bor78d, Bor80b, BS81, Bor81b, BU94, Bor85d, Bor86b, BD91a, Bul77, BK80, Bur75b, But92a, Chi74b, Chi95, DT75, Dob62a, Dok95, Dzh71, Ebe86, Gad90, Gap80, Gap86, GN93, Gur89b, Hat75, Ibr65a, IK81, IK82, Jam78, Jur75a, Kag86b, KC79, Kar80, Kha85, KA99, Kha79a, Kho67, Kho69, Kho71b, Kho71c, Kir91, KR92, Kor89a, Kor88a, Kor66, Kor68, Kos85, Kos88, Kru69, Kry73c, Leb77a, Lev76, Lev81, Log91, LS69, Lum71, Mac71a, Mak82a, Mak75, Mal69b, Mal85b, Men88, Mir86a, Mir86b, Mog74, Nad64b, Nad65, Nad70, Nad74a, Nad74b, Nad93, NR73, NR76, Neu65, Ost98, PU85b, Pit69a, Ras77, Rot78a]. Estimates [Rot78b, RS86, Roz86, Rud93, Sap79, Sat73a, Sav64, Sei82, Sen81, Sen82, Sen86b, Sen86a, Sen87, Sen89, Sen93, Sha70, She88a, Shi80, Shi82, Sho83, SF84, Sko74a, Sol67, SU90a, Tik85, Tik88, Tik98, Ul’79a, Ul’79b, Ush86, VV88, Vol78b, VK86, VK89, Yan79, Zai87a, Zai88a, Zer70, ZKC91, Zhu81, Zhu84, Zig82, Zol75, ZS76, Kha79b, Lep92, Mak88, Nad73a, Zol78d]. Estimating [Ara99, Che67a, Csö84, Dar86, Dza74, Ger96, Gin89b, GC87, Gur89a, IK93, Ibr94, IT83, Kha61b, Kho73a, KC91, KM89, Kud74, Kut89, Lut90, Man68, Nad64a, Nag67, Ol′89, Per92, Pin91a, Pon86, PY81, PY88, Sen84, Sen92, She86, Tsi78a, Roz97]. Estimation [Abd99, Ale70, Ale81, Ale92b, Ara85, Art89, BO98, BCK^{+}92, BD83, BR94, Bol83, Bon92, BL63, Bun63, Cai93, CL94, Che81, Cle98, DP75, Dar85, Dha78, DT74, Dza70, Efr80, Efr86, Efr90, Epa69, Erm89, Erm90, Ers70, FN93, Gam74, GS76, Ger79, Ger67, Gin89a, GK96a, GK80, GK81, Gol92a, Gol94, Góm93, GE91, Háj61, Ibr63b, Ibr67b, IH75a, IH75b, IK83, IK85, INK87, IK88, IK99, IM87, JP75, JR91, KR67, Kah93, Kal78, KS69, Kar95, Khm75, Kho71a, KY86, Kon73, KR76, Kor88b, Kor89c, Kor91, Koz73, Kul76, Kut78, Kut79, Lep88, Lep90, Lep91, Lep93, Lep94, LK74, Lev83, Lut91, Mal93, Mal95, MO91, Mol91, Mye93, Nag68a, NV75, Nem91, Nev79, Nev81]. Estimation [Nik87, Nys87, Oni90, Pag83, PU85a, PS93, Pin91b, PT91, PS92a, Rog88a, Ryz71, Ryz72, Sal93a, Sat73b, Saz75, Sch93a, Sed77, She93a, SK93, SO85, Str78, Tik78, Tik82, Tom71, TP59, Tsi83, Tsi86, Tsi87, Tyu70b, Vil59, Vin90, Vol93, Vol64, VS91, Wal93, WW92, Yur71a, Zhu70, Zhu86, Zue98, Bor93, Lep92]. Estimator [BT87, Fin71, GW98, Iva77, Kak78, Kor94a, Qua99, Wol65, Bol97]. Estimators [AL81, Bul84, Chi74a, Dok69, Dzh75, Gus76a, Gus76b, Gus96, IK73b, IK73a, IK74a, IK74b, IH76b, KP67, Kak78, Kha84a, Kha84b, Kle75, KP97, Lep93, Lev75b, Mal65a, MR96, Mik70a, Mok91, PC96, Sir56, Tik93, VN94, Vov92, WW66, WW68, Wol76, Lus96].

Euclidean [Kho83, SS62, Yam85]. Euler [Aly88a]. European [SKKM95a, SKKM95b]. Evaluated [Bak98]. Evaluating [Dob58c, GV93, Kor57, Rom64]. Evaluation [Ara68, DM75, Kha61c, Pag65, ST73, Yag56]. Event [Sol66]. Events [Bel69a, Döh81, Gel64, Gri79, JP97, Khi56a, Khi56b, Mal74a, Mal77, Mna86, Oso56, Pir68, PS77, Sko92, Sli62, Sli64, Tak56, Tak68, Tak76, VC71, VK86, Zäh80, Zäh84, Zäh86]. Everywhere [Gap88, Gap98]. Evolution [AR94, Bei65, Ose90, Roz94a, Roz99a, Vak90]. Evolutionary [Bul87, Gih91, Gik90]. evolvents [Tsi97]. Evolving [GE93]. Exact [Ale64, Ano85a, Chi95, IS98, Kor94a, Nag61, PU90, PP82, Rog68b, Shi65, Sir59, Tim84, Vve85a, Vve89, Zol57b, OK97, TC86]. Exact-Order [Chi95]. Example [Bel61, Dob56c, GK76, Gir62, KK99, Kut89, Tsi76a, Zai92]. Examples [Kra79, Kra80, Sen86a]. Exceeding [Sac71b]. Excessive [Bui82, Dyn70, Kuz75, Kuz80, Shu62]. Exchange [BV95]. Exchangeability [Kal93a]. Exchangeable [FLR97, JP97, Kur85, Kur88, RS98]. Excursion [Nos87]. Excursions [BN69, GP85, Nos80, Nos91]. Existence [Afa65, Ano85a, Bor81a, BS74, Chi86b, DV96, Dob64, Dob65, Dob66, DM67, Ego73, Ers75a, Fai79, Fed85, Fed77, Gal78b, Gal79, Gap97, Khr86, Kit82, Kle86, Kry91, Kul74, Kuz87, Med98, Pis98, Por75b, Rog68b, Rom65, Sav99, Ver83, Vve85a, Ers74]. Exit [Dok87, Dyn70, GR84, GS88, Gär75, Gus69, MP78b, Nov82, Pec74a, Rag86, Ron86c, Shu66a, Shu79b, Tsi78b, Zhi63]. Exits [Bel68]. Expanding [Nos83b, Shu79b]. Expansion [Bik69a, Chi73a, Chi74a, Chi74b, Chi81a, Chi81b, CvZ85, CvZ89, Chi89a, Csö75, Iva77, Koz84, Kut85, Mir84a, Mit67, Rag93, Smo97, Yak98]. Expansions [Ber87, Bik69b, Bur97, Chi80b, Eag69, Gal71, Gus76a, Gus76b, Ibr67c, Kal67, Kal93b, KF91, Kon79, KY74, Lot88, Mal87a, Mal93, Mal94, Mal95, Nag70a, Osi69, Osi71, Pet59a, Roz76, SB67, Sta65, Tim99, Ul’87, Zol92a, Gus76a, Gus76b]. Expectation [Kor66, Kor68, Kri64, LL94, Mai79, Nag68a, Sol67]. Expectations [BS73a, BVT90, Dok95, DE77, Evs87, Ham66, Sid57, VC82]. Expected [Cow93, Fil84, GP98, IM71a, IM71b, Nag67, Tim89]. Experiment [Bul85b, Koz74b, SL93, Shi97]. Experimental [Raf90, Ter93, Tsu81]. Experimentally [Dob58c]. Experiments [Cho93, CP91, DK93, Fed71a, Fed71b, Fre75, Páz68, Sed77, SS88b, Sok63a, Tsi85, Vol73]. Experiments.II [Sok63b]. Explicit [Par70, Sel69]. Exploratory [McK93]. Explosive [Anu91, Sag96]. Exponential [BO98, BO96b, BCK^{+}92, BNK96, Bor96, Che66, Ger67, Gor84a, Hid83, Jur74a, Jur74b, KP67, Kag86a, Kah93, Kle81, Mis89, MP79, Ost91, PU90, Pus93, Rod88, SU90a, Tsa77, Tsa79, Vin99, Yan85, Yur93b, Zai87a, Zai88a]. Exponentiality [Bol66]. Expositions [Ano90g]. Expression [KS91, Par70]. Expressions [San66]. Extended [Dor96b, Kab76, Nor95, Vla80]. Extending [Fel65, KM80]. Extension [Ers75b, Khr96b, Kru73b, PV64, Pro68d, Tar90a]. Extensions [Dyn68b, Roz77, Shu82a, Vor62, Vor63a]. Exterior [Fre66]. Extinction [Chi71b, Kal82, Koz77a, Lev78, Sev61, Sev67a, Sev96, Sev99, Top83, Vat77a]. Extra [Ano61b]. Extrapolation [For62, For63, Kor57, LS68a, Roz59a, TP57, Vet82]. Extrema [Bak89a, Bak90, Bak96c]. Extremal [Ark65, Fra93, KY65, Kry85, Pro59a, Rog66b, RU57]. Extreme [AB60, BW98, Bur75a, Bur75b, Ego91b, Kol69a, Mik75b, Nag71a, Nag99, Nev79, OR89, Pan95, Pav81b, SZ98, Vol60b, Wal93, dH95]. Extreme-Value [dH95]. Extremes [Gal95]. Extremum [EZ84].

Factor [Ger67, Zai87b]. Factorial [Dyc73, Mal85a, MV94]. Factorization [AS74, Bor70, Fre68, Fry76, Lin58, Lin59b, Lin59c, MW59, SR79, Shm64, Sko92, Vak94a]. Factorizations [Bor81b, Chi89b]. Factors [KR90, LL94]. Faculty [SS90]. Fades [Vol60a]. Falling [CV65, Nik64]. False [Mas99]. Families [BNK96, Che66, DR78, GT77, Gus93, IH75b, KP67, Kag86a, KR75, Kuz83, Leb81, Leb82, Mus85, TTC78, Vor62, Vor63b, Vor64]. Family [Kah93, Kle74c, Kov59, Mak67a, Mir90b, Nik80b, PT93, Shu92, VKR64]. Fast [SS88a, Szt93]. Fast-Switching [SS88a]. Fastest [BD91b, DB88]. Favorable [VS92]. Feature [Sem88a]. Features [Ste88b]. February [Ano57a, Ano69g, Ano71f, Ano74g, Ano76e, Ano84j, Ano84k, Ano85l, Ano85k, Ano85o, Ano85r, Ano87j, Ano87l, Pet59c]. February-December [Ano87l]. February-June [Ano85r]. February-May [Ano74g, Pet59c]. February-October [Ano76e]. Feedback [Dob58b, Fed69, Fed70, Glo79b, Gri95, Kat75, KS86]. Fefferman [Ibr89]. Feldman [KL79]. Feller [Bet89, Gir60b, JR82, MP80, Mol68, Rog69, Sem85, Sho97, Shu82a, Zol67a]. Feynman [JK90, Khr94]. Fiber [Lif82]. Field [Ale89, Bal68, Dob68, DM79, Eli85, For63, GE93, KD87, Khr96a, Mai75a, MV89, MV94, Nag76b, Nos79, Nos80, Nos83a, Nos83b, Nos85, Nos86, Nos88, Ost91, Ost98, Pit87b, Rog58, TP57, MP97]. Fields [Alb93, Ale74a, Amb74, Ara72, Ara90, Ave72, Bra80, Bul96, Che96, Das79, De 82, Deo74, Deo77, DW79, DK84a, DK84b, EO97, Evs78, EO79, Evs89, For62, Gap78, Gap92, Gap98, Gor83, Jur74b, Kap91, Kho73b, Koz77b, Kra74, Kra90, Kuz85, LY79, Lif80, Mal74a, Mal75, Mal77, Mar93, Mis91, MC68, Neu65, Nos91, Nov79, Pit84, Pit87a, PY69, Rao71, Ros67, Roz88, Roz91, Roz94a, Sap92, Shk74, Shl89, SY73, Ste89, Ste91, Sun99, Tem73a, Vro99, Yag57, Yur74, Zäh83, Zäh84, Zäh86, ZT92, Zin79, Dav76b, Ivk94, Tsi97]. Fifteenth [Ano82a]. Fifth [Ano82b]. Fiftieth [SS90, Kol62]. Filter [Lip76, Roz61]. Filtered [GW93, Gus93]. Filtering [Bar91, Bel95, Dav82a, Del96, Glo70, Gol92b, Had78, Kry80, Pur89, Roz84b, Vet81, Vet82, Via91, Won59, Bel94]. Filtration [Ash68, Gal94, Glo79a, Kor57, Leb96a, Lip66, Lip67, Mat62]. Final [Bas58, Chi70d, Kal99, Rom64, Sat96, She87, Vat87a, V’y78b, Vat95]. Finance [Shi95]. Financial [Ano95c, KK95a]. Finding [Gri92c, KC79, Lin56a, Pag86, Tim89]. Fine [Maz69]. Finite [Asa88, Asa91, Ash68, Bak98, Bal68, Bas65, BP70, BP89b, Bor82a, Chi76a, Chi76b, Fai76, Fai79, Fai80, GK81, Hag74, Ing86, IM65a, IT83, Ivn93, Jam78, KS61, KNS66, Kla85, Kry65, Lev86b, Mak65, Mak66, Mak67b, Mak68, Mak70b, Mak71b, Mak75, Mal80, MP92, Mes58, Mes63, Mik97, Mir84a, Par71, Pre67, Ryk66, RS70, Sac73a, Sap79, Sem88a, Sen68, Sev72, Sgi98, She88b, She99, Shi86, Tsa58a, VKR64, Vol58b, Vol60b, V’y88, YF79, YF80, Zai87a, Zai88a, Zin71, GK80, Lin59b]. Finite-Dimensional [Jam78]. Finite-Parameter [Ing86]. Finitely [Gri89, Yan70, JK90]. Finiteness [Mal85a, MV89, MV94, Mir74, Mir75, Mir78a, Mir85]. First [Afa87, Bor65b, Dok87, Fok74, Fok77, Gär75, GK68, Gus69, IN78, Kal69, Kov60, Leb71, MSV99, MN77, Mil96, Mir81, Mog77b, MP78b, Nag94, Nov91, Per92, Rag86, Rog64, Rog71, Rog73, Rot67, Sem90, Sin57, Sob73b, Sol66, Tsi78b, Vat74, VT76, Zol64a, Bar97, PS98, Zhi63, Ano88a, Ano89k]. First- [MSV99]. First-Order [Per92]. first-passage [PS98]. Fisher [BS73b, CS71, LRS65]. Fit [BM79, Chi61a, Chi64b, Chi65a, CS71, GP79, IJL91, Khm82, Kud84, Mir90b, Nik88a, Nik88b, Ron86a, Ron89, Vol80]. Fixed [Dha78, Mak82a, Nos80, Sac71b, Zub87].

Fletcher [BD83]. Floating [LN91]. Floating-Point [LN91]. Flow [Fed69, Fed70, Ivn77, Naw82, Ush78, Veĭ93]. Flows [Bel63, BP70, Fed77, Gur65, Oso56, PS77, Sli62, Sli64, VL76, Yak91]. Fluctuation [Rao89, Rog71]. Fluctuations [Bei65, BPRS94, Nev73b, PR69a]. Fold [Ara81, Zai81, Zai92]. fonctionnelle [Tou92]. Football [KPĬ93]. Forecasting [BRR98, LNF89]. Forest [Pav78, Pav84, Pav89b]. Form [Bor91b, Cad99, Chi74b, Ibr67a, Kab74, Kor88a, LM71, Mog80, Muk77, PY79, Rot76c, Sak56, Sev57b, Tut79]. Former [Gol65]. Forming [Gor71]. Forms [Bag88, Bag89b, Bak90, Bak96b, Bak96c, Bas93, BGZ98, Chi89a, CPU96, GR78, JM94, Kag89, Kag90, Mar76a, Mar78, Mik89, Mik91, Mik96b, Nyb86, Per59, Pon79a, Pon79b, Pon86, Rot76b, RS86, Ten91, Whi60, Whi64, ZI89, Bag89a, Bak89b]. Formula [Bag92, Cow93, DP75, Kho94, Mar78, Nor95, ST74, Sko69, Smo91, Tak77, Vin77]. Formulae [VT73a]. Formulas [AS98, For62, Kel73, Khi62, Kor86a, LS68b, Mis89, Ron82, Sel69, Shi65, Vis64]. Formulation [Nik99a]. Forward [CFZ95, Kva96]. Founded [SS90]. Four [Sen86a]. Fourier [Dor90a, KF91, Sko69]. Fourteenth [Ano81a]. Fourth [Ano81b, OK97]. fourth-order [OK97]. Fractional [MP79]. Framework [VS91]. Fréchet [Mus80a]. Fredholm [Kul98]. Fredholm-Type [Kul98]. Free [Bob64, BB81, Vor60]. Freeway [Veĭ93]. French [HS93, Tou92]. Frequencies [Bel69a, Bel69b, Muc88, VC71, Pus97]. Frequency [Aiv65, Bel78, Cow93, Khr96b, Vol64, Yus73]. Fresh [Nos91]. Fubini [Leb96b]. Fully [Bas60]. Function [Ale77, Ara86, BYK88, Bak96c, BM63, Bor67a, Bul67, BM93b, Cle98, CDC82, DF65, GO92, Gam89, Gam92b, Gap82, Gur89a, Har77, Ibr63b, Ibr65a, Ibr78, Il’76a, IM65a, Kon79, KC91, Kor66, Kor68, Kos85, Kuz81, Kuz87, Mac69, Mak82a, Mal64, Meg86, MSV99, MR80, MR83a, Mir90a, MP79, Nad74b, Nev79, Nys87, Osi69, PU85a, Pet70, PY81, PY78, PY79, PR85, Roz90c, Ryz71, Ryz72, Sal97, Sch65, Shu62, Sil73, Sko74b, SP57, SP61, Sol67, Tri90, Tyu70b, Ven86, Vin86a, Vin86b, Vla56, Yag56, Zhi63, Zig82, NK97, Rad97, Shi97]. Function-Space [Yag56]. Functional [Bor91b, Bul81b, BV88, Cad99, Che67b, CC96, Chi90, Ego91a, Ers75a, Fre73, Gin89b, Ham66, IK85, JM94, Kar65, Kha89b, LS81b, LS81a, Mna86, Mog80, Par87, Rom87, Ron86b, Roz64a, Rus96, SL93, Smo97, Tro93, Tsa77, Tsa79, Vas80, VT73b, Ven90, Ers74, KY97, KL93, Tou92]. Functional-Differential [Tsa77, Tsa79]. Functionally [Sal71]. Functionals [Afa87, Afa91a, Ale81, Ale91b, Ali95a, Ali95b, BM90, BM91, Bak96d, Bor77a, Bor81a, Bor90b, Bor90a, Bui82, BVT90, DI71, Dav74a, Dav81, Dav82b, Dav89a, Dav89b, Dir81, DT74, DT75, DK84a, DK84b, Dok92, Dok95, Dor90a, Dyn60, Fal78, Fil62, Fil84, Gin89a, GK96a, Hid83, IK88, Kab74, Kab75, Kha89a, KS87, Kor61b, KB89, Kor91, Lif82, Lif83, Lif85b, Mil82, Mog74, Mol88, Nem91, Por67, Rac85b, Rad92, Rao71, Rog66a, Rud93, Ryz67, Ryz69, Saz58, Sen74, Shu80b, Shu82b, Sko61a, Sko79, Syt67, Tar90b, Tie79, Ugl89, Vin90, Vol59, Bor93, Szy97, Ver94]. Functionfrom [Roz98b]. Functions [AK95, AT86, Ara81, AI99, Bak96b, Bel58, Bel60, Ben76b, BP89a, BP89b, Ber70a, Ber70b, Bik69a, Bla83, Bog80, BGS61, BZKR70, Bor88a, Bor92a, Bor85d, Bor86b, Bui82, BZ77, Bun70, Chi83, Cho84, Csö84, DR79, Dob58a, Dob60b, DM80a, DM80b, Dyn70, Dzh94, EO97, Fil62, Fre66, Gam78, Gil82, Gir77, GV86, Gol92a, Hid70, Hid83, Hur75, Ibr66, Ibr67a, Ibr74, IK83, Ibr84, Iva60, Jes80, Jes81, Jur72, Kal77a, KL83, Kha90a, Kho73a, Kif76, Kle74a, KS87, Kon73, KP84, Kor84, Kov59, Kov67, Kru93, KST60, KST61, Kuz75, Kuz80, LM71, Leh65, Lep88, Lum69, LS69, Lum71, Maĭ90, Mal90b, Mar76a, Mar78, MW59, Mes61, Mey63, Nad64b, Nad65, Nad70, Nas96a, Nas96b, Nas97, Ol′89, Osi71]. Functions [Ost89, Oza86, PP96, Pav81a, Pes94, Pit82, Pro68a, Rad98, Rog61a, Rog77a, Rog88b, Ron86a, Ros80, Rot76b, Roz59b, Roz60a, Roz81a, Roz90a, Ryz67, Sad66a, Sad67a, Saz66, SE96, Sei87b, Sen74, Sgi98, SK92, Shm64, Shu64, Shu69a, Shu69b, Shu91a, Shu92, Shu77b, Sie80, Sie87, Sil79, Stu68, SU90b, Tik78, Tit82, Tri90, Ula84, Ula88, Urb56, Ush82, VK67, VTT81, VT73b, Ven65, Vin74, Vin90, VR59, VR61, Yur71a, Zol95, BO97, PR97]. Fundamental [Rez78]. Further [BR84, Ken64, Ken87, Lu’93]. Fusions [Cap98, EH93].

G [Ano73j, Ano74i, Ano75i, Ano78g, Ano78j, Ano79g, Ano79h, Ano80e, Ano80h, Ano80g, Ano81k, Ano82o, Ano82i, Ano83d, Ano84k, Ano85l, Ano85m, Ano87i]. Gains [VGR84]. Galton [AS83, BY96, DB93, KN94, KNS66, Mak81, NV75, NK81, NK82, Rak89, TV98, Vat78, Vin88, Zub75]. Gambling [Nov97b]. Game [Elb76, GZ69, Koc98, Lut90, Lut91, Pit75, PS76, Rom61, Sob70, Vil63, Vil68, Yus67, Raz70]. Game-Theoretic [Lut90]. Game-Type [Rom61]. Games [CKR68, Dyu68, Fri74, Kif69, Kif71a, Kif71b, Rom62, She88b, VM89, Vor58, Vor67, Yan64, Yan70]. Gamma [Bon74, Ebe86, Pro94, Vol75]. Gamma-Distribution [Pro94]. Gap [Gap68]. Gaps [Fry89]. Gas [Dau76, Ose90]. Gauss [DJD86, Smo91, VP83]. Gauss-Ostrogradsky [Smo91]. Gauss-Tchebyshev [DJD86]. Gaussian [Fry90, Roz66b, Alb93, Ale63, Ale64, Ale67, Ale70, Ale74b, Ale81, Ale89, Apo67, Ara61, Ara68, Ara86, Ara90, Bak90, Bak96b, Bak96c, Bal94, Bau88, Bau97, Beg69, Bel58, Bel60, Bel66, Bel67b, BN69, Bon92, Bor83a, Bra80, BS88, BS89, Bul61, Bur79, CPU96, CDC82, DP75, Dav89b, Deo74, DW79, Dob60b, DK84a, DK84b, DK93, Dor87, Dza70, DM75, Dzh71, Eli85, Erm91, Fat97, Fel77, Fel79, Fel87, Fel88, Fel99, Fry89, Gap68, Geo80, Gin89b, Gla61, Gol65, GK80, GK81, GK84b, Gol92b, Gol94, GT77, Gor78a, Gor80, Gri78, IV93, Ibr63b, Ibr65b, Ibr66, IR70, Ibr70, IK81, IK85, INK87, IK88, IP89, Kan88, Kat75, Khr91, KT69, KR60, KP83, KP84, Kra74, Kra79, Kra80]. Gaussian [Kra90, Kul98, LO96, Lep91, Lif87, LP75, LTC80, Lin90, Lin57a, Lin66, LPS67, Lip76, Mal74a, Mal75, Mal64, Mal69a, Mal77, Mal80, Mal85a, MV89, MV94, Mar93, MP97, Mik89, Mik91, Mir69, Mir74, Mir75, Mir78a, Mir78b, Mir80a, Mir81, Mir85, MP91, Nev69, Nos85, Nos86, Nos87, Nos88, Nos91, Ost65, Ost67, Ost91, OT95, Pas99, Pau91, Pis65, Pit78, Pit82, PP82, Pit83, Pit87a, Pit87b, Plu85, Pre89, PRR87, Raj73, Rih85, Ros67, Roz62, Roz63, Roz64a, Roz66a, Rya79, Ryz65, Ryz67, Ryz69, Sal93a, Sar67, Sev67b, Shu65, Sko70a, SY73, Sko74a, Sok84, Sol94, Syt69, Tem73a, Tie79, Tsi76c, Tsi76b, Tsi83, Tsi86, TF62, Vak94a, VT73b, Ven82, VT92, Ver64b, Vin90, Vol64, Yur78c, Zäh80]. Gaussian [Zäh84, Zäh86, Zal90a, ZT92, ZK92, Zol78b, ZI89]. General [Abr88, AA68, Alb93, AI99, Bog82, Dyn67, Gir77, Gir88, Kle77, KS86, Leb82, Lin56a, Lin58, Lin59b, Lin59c, Mus73, Ose65, Pag83, Pan87, Pav97, Rak93, Top83, Tro93, Yus74, Zak90a, Gir60b]. Generalisations [Roz67]. Generalization [Chi59, Chi70a, Chu92a, Fal78, Gal78a, Haz96, Iva86, Kag89, Kho74, Kok75, LL94, Lip76, Mar78, Mes63, Nik77, Pet75, Roz98a, Sal99, Sel99, Sko76, Tem73b, Zol67a]. Generalizations [GF91, Stu67, VC66, Whi58]. Generalized [Aiv65, Apo67, Ara90, BP70, Bor83b, Bul77, Che57a, DF65, DM79, DM75, Fel76, Fel83, Fel85, Fok74, Fre65, Gap97, Hid70, Hid83, IK74b, IM89, Jam78, Kag90, Kha88, Kha89a, Kha93, Kha94, KA99, KP97, Lab72, LS97, LSS99, Lev88, Lev75b, Mas74a, Nad93, Per59, Pit87b, Por79, Por80, Rao71, Ron89, Roo99, Roz59a, Roz90a, Roz91, Sgi98, Sor78, Ven68, WW66, WW68, ZT92, RT98]. Generated [Dav68, FS79, KS91, Mal74a, Mal75, Mal77, Mol96, Rén56, Sag91, Shu91a, Shu91b, Smo85, Tut67, Zol76, Shu91c]. Generating [BM63, Koc90, Maĭ90, PR85]. Generation [Kha69b]. Generations [Kha69b, Kha69a]. Generator [Koz81, Mik96a]. Generators [Koz71a, Tan70, TF91]. Geochemical [Pro65b]. Geodesic [Le93]. Geometric [IC59, Lya87a, TT82, Tsi86, Tsi87]. Geometrical [Tsi83]. Geometry [KK99, Le93]. Germs [Din93]. GF [Mik99]. Gibbs [Dob64, Dob73, GS85]. Gibbsian [Ave72, Dob74, Koz77b]. Gibes [KD87]. Gikhman [Ano87a]. Girsanov [DKPF67, Gal78a]. Given [Ano73h, Anu81, Bel64, Bel65, Ers75a, Gur92, GK68, IN78, KC91, Mak68, Nov99, Ros67, Vat94a, V’y88, Ers74]. Glivenko [Saz63, Shi86]. Global [CC96, EO97, EZ84, Kon78, Kru76, Le93, SI67]. Globally [Mar93]. Gnedenko [Chi87a, Kol62, Shi93, V^{+}97, ZK61]. Goncharov [BBM59]. Good [Dzh75, Ger96]. Goodness [BM79, Chi64b, CS71, GP79, IJL91, Khm82, Kud84, Mir90b, Nik88a, Nik88b, Ron86a, Ron89, Vol80]. Goodness-Of-Fit [GP79, BM79, Chi64b, Khm82, Kud84, Mir90b, Nik88a, Ron86a, Ron89, Vol80]. Gradual [Yus90, Yus91a]. Graph [Bur75a, Bur75b, Evs89, Ivc73, Ivc74a, Kol87, Men88, Ste70a, Ste88b]. Graphs [Bal96, Kar95, KKR89, Kel90, Kel73, Kor86a, Lya87b, Mór93, Ste69a, Ste70b, Ste73].

Greater [KR75]. Greatly [Kis62]. Green [Doo57]. Group [Abr59, Ash68, Bas60, Ben76b, Ber74, Dob63, EM68, FF82, Haz95, Kap74, Klo64, Lis96, Mak66, Mak67b, Mak68, Mak70b, ML65, Mik77, Mik78, Mik80, MR89, Mol67b, Par64, Rok61, She99, Shu92, Tut64, Tut71, VM83, ZK76, Min91]. Grouped [Art89, IM81]. Grouping [AK70, Uri93]. Groups [Bol57, But79, Fel76, Fel77, Fel79, Fel83, Fel85, Fel86, Fel87, Fel88, Fel90, Fel92, Fel93, Fel98, Fel99, GS85, Ign82, IT79, Klo59, Klo61, Klo62, Mak65, Mak67a, Mak71a, Mak71b, Mak73, Mak75, Nik87, PS64, ST66, Shl81, Shu91a, Shu96, Tar90b, Tum56, UU85a, VS77, VS78, Vir70, Vir74]. Growing [Leo86, PL83]. Growth [AS92, Ego91b, Pet74b, Zin79]. Guarantee [VN99]. Guaranty [Nov88].

H. [Chi70a, KS89, Val74a]. Haar [Pav98]. Hajek [Bon98, Kon74b, KL79]. Half [Afa87, Anu80, Bla83, Gal95, Her91, Ibr78, KU88b, Ost89, Ros80, Shu81, Shu82b, Tit82, Ula88]. Half-Axis [Afa87, Her91, Shu81, Shu82b, Tit82]. Half-Line [Bla83, Ibr78, KU88b, Ost89, Ula88]. Half-Plane [Anu80]. Halfline [BK86]. Halfspace [Sha80a]. Hammerstein [Elb83]. Hand [Bor78b, Kha66b]. Harmonic [Ben76b, BP89a, BP89b, Dzh94, Hid70, Kif76, Shu91a, Shu92]. Harris [Sag86, Vat77d, Vat77b, Vat77c, Vat79, Vat80, Vat81, Vat82, Vat87a, Vat87b, Vat94b, V’y78a, Mal87b, Mal90a, Mal92a, VS89, VS90]. Harrison [KK95b]. Having [Gin89a, Nas96b, Son92, Tsi76a, Vin76, Ara81]. Hazard [Vin74, Vin76]. Head [Sam82, V′y98b]. Head-Run [Sam82, V′y98b]. Heat [Vak67, Veĭ93]. Heavy [CFZ95, KK82, Nag99, Pav91a, Sam63, VP64]. Heavy-Tailed [CFZ95, Nag99]. Hedging [DKR95, MN99, Nov99, Sch93b]. Height [Pav84, Rog71, Sac71b, Sac73a, Sac73b, Sev94, Vat94a]. Heights [Nag86a]. Helixes [Mas72]. Hermite [Ivk86, SB67, Vis92]. Heteroscedastic [Dok93]. Heuristic [Che56]. Hida [IR71]. Hida-Craméer [IR71]. Hierarchical [Ble86, Nai82, Shn87]. Higgins [Mir77, Mir81]. High [BN69, Nos85, Nos86, Nos87, Nos88, Nos91, Vol64]. High-Frequency [Vol64]. High-Level [Nos87, Nos88, Nos91]. Higher [Ale91a, Ale94, Ale96, Bul75, LS60, Sgi92, Shi60, Shi63a, Sin63, Tei66, VT76, Zez75, Zhu86]. Higher-Order [Ale91a, Ale94, Ale96, Bul75, LS60, Shi60, Zhu86]. Highly [Fre76]. Hilbert [Ul′91, Ale92a, Ali87, Ali90, BPS97, Cra64, Dor90b, Fed82, JK90, KS64, Kir94, Kru71b, Kru73a, Kru74, MO83, Mas72, Mus77, Nag86b, PF57, Pro84, PS92b, RS98, Roz97, Sal80, Sap88, Saz90, Saz92, SS88b, Sen90, Sen92, Sko69, Sko70b, Sko74b, Sko79, Sko80b, Tik92a, Tik94, VK97, Vak99, Vve84, Vve85b, Vve89, Yur92a, Yur92b, Yur93a, Zal83, ZS84, Zal86, ZSU89a, ZSU89b, ZSU89c, Zal90b, Zig82, de 93]. Hilbert-Space-Valued [PS92b, Saz92]. Hill [BT87]. Histogram [ZKC91]. Hit [Rog73]. Hitting [AR86, Shu66b, Tik92a, Tik94, Yur83]. Hodges [Nik88b]. Hold [Nag69a, Nag69b, Nag69c]. Homogeneity [Ber78, Che60, Che80, Dar86, Gik57, Nik88b, Uri93, Vol80]. Homogeneous [Aga80, Ale74a, Ale89, Amb79, AS82, Bra80, BK86, Dob56c, Dob57, Dok87, Eli85, ES69a, ES69b, Gap78, Gap98, Gus69, Gus84, Kai86, Kit82, Kli79a, Kra74, Kra90, Kuz87, LY79, Lif76b, Mak70a, MP70, Nag61, Naw82, Nos85, Nos86, Nos88, PY69, Ron86b, RS70, Sen68, Shu58, Shu89, Shu90, Shu91b, SF84, Sko63b, Sko63a, Sko67, SY73, Sli62, Sli64, Smo97, Tem73a, TP57, Vol57, Vol61a, Vol58b, Vol60b, Yan96, YF79, YF80, Shu91c]. Homographic [Ali92]. Hopf [Roz66b, Roz66a, Tem73b]. Horizon [Nos83a, Nos83b]. Hu [JK90]. Hunt [Glo82]. Hydromagnetic [MT85]. Hyperbolic [DK97, OK97]. Hypergeometric [Bol64a]. Hypoentropy [MTP93]. Hypotheses [Aiv65, Bel67a, Ber81, Ber85, Ber86a, Ber76, Bol90, Bor76, Chi64b, Chi67, Dzh77, Erm95, Gol85b, Ing84, Ing86, Ing87, Ing95, Ivc75c, Khm75, Khm80, Kol88, Kus68, Kut76, Mir78c, Mir80b, Nik88a, PK82, Pav88, Pav91b, Pet56a, VT92, Vol67, Vol69, VN99]. Hypothesis [IV93, IJL91, Ing83, Ing89, IM79, IM81, Kho79, KP71, Nik80a, Tsi85, Tyu70a, Tyu74, Yas84, Zas67, ZK61, Mak88].

I. [Shi64a]. I.I.D. [Vas80]. ibid [Bak89b, Ul′91]. Ideal [Dau76, RR93, Zol78c]. Idempotent [DD99]. Identical [Mak66, Mak68, Pet62, Tit82, Vin94, Vin95, Vin96]. Identically [AV82, Bet88, Bik69b, IH76a, IH76b, Kag90, Kle87, Kry68, PY81, Ros80, Sla82, Sla85, Sza73, Zin76]. Identification [Kos99]. Identities [Bor70, Pec74a]. Identity [AS74, Nov73, Sol66]. if [Ano84l]. Igor [DKPF67]. II [Shu91c, Ul′91, Ale80, Ara82b, BdH79, Bel95, Bel67b, Bor90a, Bor60a, Bor65c, Bor88b, Bor86b, Bur75a, Chi81b, Dav82b, Dob56b, ES69b, Faj88, Gil93b, Gus76b, Ibr70, IM71b, IK73a, IMS93b, Kar86b, Kli79b, Kol88, KP84, Kon70, Kor96, LS60, Lep93, Lin59b, Lin61b, Lot80b, Mal90a, Mal97, Med78, Nag69c, Nag70c, Rot78b, SKKM95b, Shu69b, Shu78a, Shu86, Shu90, Shu82b, Sir63b, Sko60, Sko62, SS66, Ste91, Tik94, Tsi86, VS90, Ven77b, VS78, Vin86b, VR61, Yur92b, Zäh86, ZSU89b, Zhu84]. III [Lin59c, Lin62, Tsi87, VT82, Ven80, Yur93a]. Il’ich [Ano87a]. Image [Kor92]. Images [Zäh83]. Imaginary [Pro68a]. Imbedded [Zäh83]. Imbedding [Par67]. Immigration [BR79, BR84, BZ84, Chi70c, GC87, Jag68, KW71, Nag75, Pol78, Rak87, Rak89, Rak93, Sag83, Vat74, Vat77d, Yag68, Zub72].

Imperfect [Rom62]. Imperfections [Koz73]. Improved [GV93, Kah93, Kos89, Str78]. Improvement [Ara83, Saz69]. Impulse [Yus91a]. Impulses [Rya76]. Impulsive [Gri75, Yus90]. Inaccessible [Roz73]. Inaccurately [Bur79]. Including [BM99]. Incoming [Ush78]. Incomplete [Fäh78, KP67, KY74, LS68a, Sch93b, SP57, Yus76]. Increase [Rog61b, Tum56]. Increases [KSS77, Sem88a]. Increasing [Anu91, Bor78c, Gur65, Nag94, Pur89, Vin74, Vin76, Bor77c]. Increments [Ali92, Bor65b, BK86, BDC68, DI71, Dav74a, DN88, Gal85a, Gal94, Gla61, GN87, GK68, Gus69, GK69, Gus84, Kab75, Kor89c, Lif85b, Lin92, Mak70b, Mog77b, PR69a, Pec77, Rog65, Rog66a, Rog68a, Rog72, Rog92, Sht65b, Sht65a, Sht66, Shu79a, Shu79b, Sko57b, Sko57a, Yam95, Zin79, Zin88, Zol64a, Zol65a, Zol66b]. Incursions [DY68]. Indecomposable [UU85b, UU85a]. Indefinite [LL94]. Independence [Döh81, Kle71, Kle74c, KY84, Mal91, Mik75b, Vor61, Zin58, Zin80]. Independent [Ale77, Ale79, Ale80, Ali92, Amo80, Ant80, AN74, Ara75, Ark90, AB60, Bas59, BR85, BGZ98, Ber70a, Ber70b, Bet88, Bet89, Bik69b, Bor84b, Bor65b, Bor73b, BK86, Bra90, BDC68, BS88, BS92, Cad99, Chi64d, Chu92b, Dar76, DI71, Dav74a, DN88, Ego70, Ego91c, Fra63, FN71, Gal85a, Gal94, Gam82, Gap89b, Gil82, Gir83, Gir96, God79, GN87, Gri74, GK68, Gus69, Gus84, GZ66, Ibr66, Ibr67a, IP74, Ivc74b, JM94, Kab75, Kag90, Kal77b, Kha61a, Kle87, Kle96, Kol56b, Kor86b, Kor89b, Kor89a, Kor90a, Kor98, KSS77, Kor89c, Koz81, Kru71b, Kru73a, Kru74, Kru75, Kru80, Kru99, LM71, Lar94, Lif85b, Lin61a, Lin61b, Lin62, Log91, Mak67b, Mak70b, Mar85, Mar87b, MP93a, Mik74, Mik94, Mik89]. Independent [MR83a, Mog77b, MFY74, Muk85, Muk90, Muk98, Nag70a, Nag70d, Nag77, NE77, Nag78, NP78, Nag98a, NP69, Nev71, Nev72, Nov83, Osi69, Osi71, Osi73, Osi74, Osi79, Pak59, Pan87, PR69a, Pec77, Pet56b, Pet59a, Pet59b, Pet64, Pet65, Pet70, Pin79, PU85b, PU90, Pin91a, PY81, PY88, Pre86, Rag91, Rao71, Rog61b, Rog65, Rog66a, Rog66c, Rog68a, Rog72, Rog92, Ron88, Rot76a, Roz90c, Roz98b, RS76b, Saz75, Sha96, She99, Sho96, Sht65b, Sht65a, Sht66, Shu76a, Shu79a, Shu79b, Sil79, Sin57, Sko57b, Sko57a, Sko80b, Smi69, Smo85, Sre70, Sta65, SU90b, Sza73, Ten91, Ter93, Tie77, Ul’87, Ush86, VIL69, VG90, Vol96a, Vve89, Whi60, Whi64, Yam95, Zai84, Zai88b, Zhu66, Zin79, Zin65, Zin71, Zol64a, Zol65a]. Independent [Zol65b, Zol66b, Zol77a, Zol78c, Zol92a, ZM74, ZM80, Gap95, Kor95, Kor96, NK97]. Index [MTP93]. Indexed [Bul94, IM93, Kar95, Kor93]. Indicators [Lif80, Log91, Mik94, Rak88, Rak90, Vol96a, Zhu88]. Indices [Abd99, Ric65, Vir84, Kor95, Kor96]. Indirect [BR94]. Individual [VW99, V′y98a, V′y98b]. Induced [EN83, Ser79]. Induction [Góm93]. Inequalities [Ant80, Are78, Bak89b, Bor68b, Bor73b, Bul94, CPP98, DM68, DSS94, Ebr71b, Ebr71a, FN71, Har77, Kar86a, Kar86b, Kat78, Kon74b, Lev79, Men88, MR80, MW99, MS86, Nag77, NP78, Nag98a, Nag98b, Nov71a, Pav97, PS86, Pin91a, Pon79a, Pra93, Pre82, Pro68c, Pro68d, Pro68a, Pro68b, PS75, Rog88b, Rog94, Sad66a, Sad67a, Sak85, She88a, She93a, Ush82, Ute90, Yur70, Yur71b, Yur72, Yus88, Zai87b, Zol68b, Zub99]. Inequality [Amo88a, Ark71, BO96a, Ban85, BK71, Bol61, Bon98, BU84, BK89, Bul74b, DJD86, Fed82, Fet74, FM92, Gam78, Gam92b, GF91, Gri94a, IS98, Jur75a, Kho74, Kry71, MR83a, Pet75, Pet76, Pon79b, PY79, Rog60, Sad66b, Sad67b, Sal93b, Sha61, Sha93, Sid91, Tur91, VP83, Whi58, Whi69, YTS89, Zol66c]. Inference [BN94, Gor81, Ivc85, Rao63a, Vol79, PvZ97]. Inferences [MN89]. Inferring [Saw93]. Infinite [AR86, Bal94, Bas77, Bor85a, Háj61, IS74, KL83, Khr91, Kra69, Lev79, Lus99, Oni90, OR85, Qua99, Rak87, Saz65, SK92, Sko80b, SU90a, Tol99, Tsa83, Tsi83, Tsi86, Tsi87, Vat77c, Vat79, Yan70, Yur70, Zem93, Zol77a, Zol78a, JK90]. Infinite- [Tsi86]. Infinite-Dimensional [AR86, Bal94, Bas77, KL83, Khr91, Lev79, OR85, SU90a, Tsi83, Tsi87, Yur70, Zol77a, Zol78a, Tol99, JK90]. Infinitely [Bag92, BP93, Den89, Fel86, FST95, Fry76, Ibr56b, Ibr78, Jes80, Jes81, JR89, Jur75b, Kho87, KMM85, Kru70, Kru71a, KA83, KA85, KA87, Lin58, Lin59b, Lin59c, Mar70, Mes60, Mes61, Ost83, PS64, Par67, Pre84, Pro58, RS82, Sat95b, Sie87, Sko65, Stu60, Tsa58b, UU85b, Vol96b, Yak88, Yam96, Yur78a, Zai92, Zol58, Zol61b, ZK76]. Infinitesimal [Dyn56a, MP58, Pit87a, TF62, Ven86, Zak90b]. Infinity [Pro91a, Zol64a]. Influence [Ego91b, Fra93, Sev62]. Information [Ano59c, Ano60a, Ano61a, Ano64a, Ano65a, Ano89b, Dob58b, Dob60a, Dob62b, Fed82, Glo79b, GW98, Hau93, Kag86a, KL77, Lin59a, Per59, Raf90, Rom62, RR64, She93a, She88b, SS65a, Tin62a, Tin62b, TP58, Yus76]. Information-Theoretic [Lin59a]. Informational [Lev79]. Inhomogeneities [GV59, ZI70]. Inhomogeneous [Sar61]. Initial [Bon92, BM92, LO96, Leo78]. Initial-Boundary [Bon92]. Innovation [Chi86b, IP89, Tsi97]. Input [SS65a, Zak91]. Inputs [Zak90a]. Instant [Gri95]. Institute [Ano61g, Ano62e, Ano63c, Ano64d, Ano65d, Ano67f, Ano68f, Ano69h, Ano71c, Ano71g, Ano73h, Ano73k, Ano74j, Ano75k, Ano76f, Ano76g, Ano77h, Ano78i, Ano78h, Ano78k, Ano79f, Ano79i, Ano79j, Ano80f, Ano80i, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82m, Ano82h, Ano82j, Ano82k, Ano83c, Ano83f, Ano84h, Ano84j, Ano84l, Ano84i, Ano85q, Ano85p, Ano85k, Ano85n, Ano85o, Ano85r, Ano86k, Ano87k, Ano87n, Ano87j, Ano87m, Ano87o, Ano87l, Ano88l, Leo82]. Insurance [EK94, Vin99]. Integer [Bra94, Gam82, Gam86, McD80, She97]. Integer-Valued [Gam82, Gam86, McD80]. Integrability [AI99, JR89, Nov80]. Integrable [Por75b, Sko75]. Integral [Ale84, Bak96d, Beg69, BP89b, Bor77a, Che67b, Dob64, Dok95, Dor96b, FST95, Gam68, Gam91, Gam92a, JL56, JR89, Kab75, Koz81, Kry69, Kry72, Kry73a, KH81, KY74, Met75, Mir76, Nad74b, Nag69a, Nag69b, Nag69c, Nor95, Pav81b, Por67, Roz66a, Roz66b, She88a, Sko76, Smi69, Sob60, VK67, Vla56, JK90, KM95, Met74, Muk96, Rad97, Szy97]. Integrals [BS68a, Bon98, DP75, Dav91, DM96, Dyn60, Gal83, Gal85b, Gap92, Har91, Kab74, Kry71, Kry73c, Kut76, LS97, Leb96b, Nor88, Nov71a, Nov73, Pas99, Pra99, Rad92, Rad98, Rie81, Th93, Tou94, Tur71, VTGM85, Yag56, Zan99]. Integrated [BJL98]. Integration [Dor90a, KL83, Kir94, Leb90, Mil79a, Mil96, Mil75, Mil97]. Integro [BM99, Glo79a]. Integro-Differential [Glo79a]. Integro-Local [BM99]. Intensities [Kru69]. Intensity [BR79, BR84, Vol93]. Interacting [Tsa83]. Interaction [BD97a, Bas77, Kal82, Kal99, MT83, Mil72, ZT92]. Interdependent [KP71]. Interest [BN94]. Interests [Fri69]. Interior [Vor63b]. Intermediate [CdHY97]. Internal [Won59]. International [Ano74a, Ano89c, Ano92d, Ano98d, Ano99d]. Interoutput [Gri95]. Interpolation [Kva96, Leb74, LS68b, Nag76b, Nov79, Pir67, Vet82]. Interpretation [Kal93b, Kla85, Mak97, Mon56]. Interrecord [Ste88a]. Intersection [Fed69, Fed70, Rag91]. Intersections [Bel66]. Interval [BL66, IS74, KC79, PG80, Pec74a, Rog73, Sem90, Shu79b]. Intervals [BO98, BP91, Bor78c, Dar86, GR67, Kab76, Tho93a, Vin99, Bor77c]. Interview [Shi93]. Introduction [Gir88]. Invariability [CS61]. Invariance [Bak96d, Bor78a, Bor84a, BB87, Bor74b, BS81, Bor83b, Bor85b, BM93a, Bul77, Dav70, Gor76, Gor83, Gor84b, Ivn82, Ivn90, Ivn93, Ivn98, KS80, Kru98, LY79, Lif85a, LS84, Pav97, RS88, Sīl92, Smo85, Vol81, Zin86, Szy97]. Invariant [BP91, Gär77, KK76, Koc90, Koz84, Mak75, Mol67b, Nev64, OCG93, Ose90, Ruk76, Shu58, Shu82a, Tol99, Tut61]. Invariants [LS59, Zhu70, Zhu72]. Inventory [Bul64a, Bul64b, Bul67, Rom64, Vil89, Wit71]. Inverse [DK93, Lin57b, Zhd90]. Inversion [AS98, Sko69, Vin77]. Investigating [Gri92b, Tsi91]. Investigation [Ano95c, Bob64, Chi65a, Dob64, Dob74, Gas63, Kha84b, Nos91, Rez77a, Rez80, Tut69a]. Involving [Che64b, Nos91, Ros95, Zas67]. Iosif [Ano87a]. Irregular [Shu73]. Irregularities [Tut69a]. Ising [BG91, Dob65, Dob66, Dob73]. Isolated [Sar73]. Isomorphism [Liv75, Liv80, Zas64]. Isotonic [CDM93]. Isotopic [Ale74a]. Isotropic [LY79, PY69, YK86, YK87]. ItÔ [Dok92, AK89, Aly88a, Aly88b, Dav91, DM96, Dok95, Ers72b, Fed85, KV85, Kry69, Kry73a, Kry91, Nor95, Ver88b, Ver92]. Itô-Volterra [KV85]. Items [Bas65, Kor58, Kor61a]. Iterated [Bul78, Bul81b, BV88, Ego69, Ego70, Ego72, Ego85, Ego91a, Ego91c, Gap65, Gap69, GK75, Ios68, Ios70, Kle87, Kor88a, Kru88, Kul85a, MP77, MP78a, Mar79, Mar81, Mar86a, Mik89, Mik91, Mog80, NP99, Pet66b, Pet71, Rao89, Rez68, Rez71, Roz98a, Vas80, VD93, Vov88, Zhu88]. Iterations [Via91]. Ito [Gir62, LS97]. IV [Ven83].

J [Sil74]. J-Topology [Sil74]. J. [Bol64b]. Jackson [KKR89, Kel90]. Jacobi [GV86, Rez81]. Jagers [Gri92b]. James [Yus87]. Janko [Bol64b]. Janson [Mik92]. Japanese [Ano70f]. Jirina [AS83]. Joint [AT93, Bel69a, Bel69b, Bel78, Gur92, Gus69, GK69, KY97, MSV99, Nev73b, PR69a, Sal71, She97, Vin96]. Jonction [HS93]. Jubilees [GC88]. July [Ano87m, GIM^{+}97]. Jump [AHE85, Bra91, Che89, Fok74, Fok77, Kit86, Kov60, Leb74, Leb87, Mog77b, Piu98, Rog64, Rog72, Yus80, Yus81]. Jumps [Gal78a, God82, KSS77, Pit92, Shu79b, Ver88b, Zin88]. Junction [HS93]. June [Ano85r, Ano87j].

K. [Gir62]. Kadyrov [Bol64b]. Kagan [Lis96]. Kalman [Lip76, Via91]. Kalman-Bucy [Lip76]. Kantorovich [Kai86, Rac85a]. Kantorovich-Rubinshtein [Kai86]. Katz [Pet66a, Stu65]. Kendall [Gel74]. Kernel [Kha89b, Kos89, Mok91, Nad93, Nor95, Ron82, Ron86b, Tsy88]. Kernels [CPP98, FST95, NRV67, Oza86]. Kesten [Lët91, Val74a, ZM79a]. Khasanovich [Ano89g]. Khinchin [SZ98, Gne60, GF91, Kho94, Kok75, Pan95]. Khinchin-Type [SZ98, Pan95]. Khinchine [Kho88]. Kiefer [BDEM92, DN88, Pav91b, Ral97]. Kiefer-Weiss [Pav91b]. Kiev [Ano73j, Ano74i, Ano75i, Ano78g, Ano78j, Ano79g, Ano79h, Ano80e, Ano80h, Ano80g, Ano82o, Ano82i, Ano83e, Ano83d, Ano84k, Ano85j, Ano85l, Ano85m, Ano86l, Ano87i, Ano88k, Ano86j]. Kind [Che56, Mal65b, Pav97, Sko67]. Kinetic [Kli57]. Kingman [Ula84]. Known [Bur79, Kop96, Lum71]. Kolmogorov [Amo88b, Ano88h, Ara82a, Ara82b, Ara83, Asa88, Ban85, Bin90, Bol90, Bon98, Che56, Dar60, Epa68, Epa74a, Epa74b, Erm96, GS63, Ing88, Khr96b, Mak88, Ose90, Pre74, Pro60, Ros75, Roz84a, Roz98a, Sak85, Shi90, Shu77b, SS90, Tie80, TC86, Tur91, Vak90, Vov88, Vov97, Whi69, YTS89, ZA84, Zai85, Zaĭ90, Zol68b]. Kolmogorov-Smirnov [Bol90, Che56, Dar60]. Kolmogorov-Type [Zol68b]. Kolmogorv [Bag92]. Kullback [GW98]. Kurtosis [MRS94].

L [Bak89b, Fal78, Soz64, Val74a]. L. [KC79]. Labeled [Sev94]. Lackwell [Fri70]. lacunae [Fry90]. Lacunary [Tsi76b]. Ladder [Nag86a, Rog71, Rog92]. LAN [Gol92a, IK93]. Language [Sav64]. Laplace [Csö75, Vin74, Vin77, Zub99]. Lapunow [Plu85]. LAQ [Gus96]. Large [AL77, Ale91b, Ale79, Ale80, Ale84, Amb79, Amo88a, Amo91b, Ark87, Ark88, Ark90, AV82, BO96b, Bak96a, Bak98, BGW75, Bel69b, Bel78, Ben87, Big93, Ble86, Bor67a, Bor91b, BM93a, BK97, BM99, BS87, BS88, De 82, DM80b, Ebr71b, Ebr71a, Ego70, Ego72, Ego91c, EL99, Fat97, Fre74, GS88, Gap74, Gap78, Gap82, Gär76, Gär77, God79, God82, IL94, Jur74a, KK95a, KN75, Kru82, Kur85, KT77, Lin61a, Lin61b, Lin62, Lip97, Mak91, Mak81, Mak82b, Mak95, Mak96b, MA79, Mal92b, Mal97, Mar80, Mik96b, Mog77a, Mog77c, Mol88, Mos70, Mos73, Nag65, Nag69a, Nag69b, Nag69c, Nag71a, Nag73, NV75, NV76, Nag82, Nag84b, Nag94, Nag99, Nyb86, Osi73, Osi74, Osi75, Osi79, Pet65, Pet68]. Large [Pet69, Pin81, PS86, PP82, Pit83, Pit92, Pit71b, Pro59b, Puk94, PS89, Rao63b, Ric57, Ric58, Ric64, Roz81b, Roz82, Roz86, Roz87, Roz90b, Roz90c, Roz94b, Roz98b, Sag91, San66, Sau92, Ser80, Sha80a, Sho79, Sho83, SMI90a, Sla82, SA94, Sur63, Tim99, Tut69b, Tut79, Tyu74, Van83, VT82, Ven77a, Ven77b, Ven80, Ven83, Ver66, Ver64a, Ver91, Ver92, Ver99b, Ver99a, VG90, Vol61b, Vol65, VN78, Woy74, YK87, Yur71b, Yur72, Yur92a, Yur92b, Yur93a, Yur93b, Zal90b, ZM79a, Chu94, Gap95, PR97, QT97, Ver94, Vov97]. Large-Deviation [BK97]. Large-Sized [Tyu74]. Lattice [Bik69b, Bor60b, Bor60a, Bor61, CN87, Dob74, Gam64, Gam66, Gam68, GP71, Gam74, Gam91, Gam92a, Kol67, Kor61b, Kor62, MPY70, Mos73, MFY74, Muk92, Ose90, Osi69, Pet62, PY81, Pre67, Roz57, Yak98, ZM95, Zue88]. Latticed [Bor62]. Lattices [BG91]. Laurent [Mus69]. Law [AL77, AS98, Bor91b, BMS96, Bre65, BS87, BS88, Bul78, Bul81b, BV88, CPU93, Dau88, De 82, Ego69, Ego70, Ego72, Ego73, EN75, Ego80, Ego85, Ego91a, Ego91c, Gap68, Gap69, Gap74, Gap78, Gap82, GK75, Gir85, Gir86, God79, Gri89, Ios68, Ios70, JL56, KS67, Kho83, Kle87, Kor88a, Kru82, Kru88, Kul85a, Kur85, KT77, Lin56b, Lut90, Lut91, MP77, MP78a, Mar79, Mar80, Mar81, Mar86a, Mar87b, Mik89, Mik91, Mik96b, Mir86a, Mog80, Nag73, NV76, Nag84b, Nag86b, Nev71, Nev72, Nor98, Ol′89, Pec74b, Pet66b, Pet69, Pet71, Pit71b, Pre86, Pro59b, Ral97, Rao63b, Rao89, Rez68, Rez71, Rih85, Roz81b, Roz90b, Roz90c, Roz98a, Roz98b, Sat73b, Sch79, Sem88b, Ser80, Sev73, SS92, Sur63]. Law [Tut69b, Tut79, Vas80, VD93, Ver66, Ver64a, VN78, Vov88, Yan85, ZM79a, Zhu88, Zin86, Zin80, Zol68a, Gap65, Gap95, PR97, QT97, Tou92, Vov97]. Lawin [NP99]. Laws [AK95, Ara81, AV82, Bag92, BGW75, BGZ98, BW98, Chi87b, Chu95, Dob58d, Gia96, Gri94c, GK96b, Gri96b, Haz96, HJ78, Ibr56b, IC59, Kor90a, Kor98, Kor99, KB92, Kru92, Kru93, Lin57a, Lin58, Lin59b, Lin59c, Mes60, Mes61, Mih56, Mik65, MR93, Mol88, Ost65, Ost67, Pet74a, Sak56, Sat73a, Seĭ91, Stu60, Stu68, Tsa58b, UU85b, Ush84, Woy74, Yan89, Yur78b, Zai84, Zai92, ZL64b, Zin71, Zol61b, Zol67b, Zol95, Chu94]. Layer [Kor90b]. L′vovich [Ano97d, GIM^{+}97]. Leading [BB91, Dob58d]. Least [BL63, Iva77, Kak78, KP97, KC91, Kur87, Lin57b, Mal69b, ZK92, Bol97]. Least-Squares [Kur87, Mal69b]. Lebesgue [Koz84]. Lectures [Ano73h, JK90]. Left [TV98]. Left-Continuous [TV98].

Lehmann [Nik88b, TC86]. Leibler [GW98]. Leight [Gru76]. Lemma [Kha65, KR87, Yur71a]. Length [DI74, Kat78, Kis62, Mal75, Mor75, Nos87, Sam82, Tim89]. Lengths [ZI74]. Leningrad [Ano76f, Ano77h, Ano78h, Ano78k, Ano79i, Ano80i, Ano81h, Ano81i, Ano82q, Ano82l, Ano82h, Ano84h, Ano84l, Ano84i, Ano85o, Ano86k, Ano87n, Ano87l]. Letter [Ano63a, Ano75b, Ano77a, Ano99a, Bak89b, Bor77b, Ul′91]. Letters [Bur96]. Level [Bel66, Bel67b, BN69, BJL98, Bul61, Epa68, Epa74a, Fil84, Fok77, GK68, Gus84, Iva60, Mal74a, Mal75, Mal69a, Mal77, MN77, Nag94, Nos79, Nos80, Nos86, Nos87, Nos88, Nos91, Nov91, Pit78, Sem90, Zol64a]. Levy [Ara99, ZS76, Abr77, Ara72, Bul74b, Che57a, Gor89, Jes81, Jur75a, Kho94, Lif80, Mir90a, Mol67a, MC68, Mus80a, Pet75, Sak85, Sen85, Zai87a, Zai87b, Zai88a, Kho88]. Lévy-Khinchine [Kho88]. Lévy-Prokhorov [Zai88a]. Lie [Mak73, Tut67, Vir70, Vir74]. Liev [Ano81k]. Life [Ano81e, Ano84e, Ano85h, Ano85i, Ano86f, Ano86e, Ano86g, Ano86h, Ano87g, Ano88f, Ano88g, Ano90g, Ano91e, Ano91f, Ano92b, Ano92c, Ano93c, Ano94c, Ano94d, Ano94e, Ano95c, Ano96b, Ano97c, Ano99b, DK93, Zub72]. Life-Periods [Zub72]. Lifetime [Par71]. Lifetimes [VW99, YM84]. Like [Lu’93, Mel83]. Likelihood [AK70, Art89, BCK^{+}92, BN94, Chi74a, CvZ85, CvZ89, Cle98, CDC82, DM75, GW98, Ibr67b, IK73b, IK74b, IH76b, Kha87, KK75, Kor86c, Kut85, Lot86, Lot88, Mar93, Mit67, Mos91, Pis65, Str78, Tar85, Tsi83, Tsi86, Tsi87, Wal93, WW66, WW68, Wol65, PR99]. LIL [Alb93, Kla85]. Limit [Abr87a, Abr88, Afa91a, Afa65, AL77, Ale77, Ale84, AS98, Ali87, Ali90, AK89, Aly88b, Amo91b, Ano60b, Ano63b, Ano64a, Ano65a, Ara82a, Ara82b, Ara83, Ara90, AR86, BZ84, BM90, BM91, BY96, BD78, BdH79, Bar85, Bas60, Bel63, BN96, Bez71, Bha70, Bin90, Bol68, Bor78b, Bor80a, Bor84b, Bor88a, Bor60b, Bor60a, Bor64b, Bor65a, BR65, Bor65c, Bor87, Bor88b, Bor89b, BM99, Bra87, Bre65, BZ77, Bul96, Bun74, BM93b, CPU93, CPP98, Chi64a, Chi90, Chi57, Chi65b, Chi70c, Chi70d, Chi72, Chu95, Cra65, Das79, Dau76, Dau88, Dav74a, Dav89a, DM96, DEO84, Dob56a, Dob56b, Dob60a, DM80a, Eag77, Eli89, Eng84, Eng85, EO86, Fom83, FLR97, Fra63, Gal71, Gam64, Gam66, GR78]. Limit [Gam81, Gam89, Gap70, Gap73, Gel64, Gik58, Gil82, Gir76, Gir77, Gir80, Gir82, Gir83, GV86, Gir91, Gir98, Gla61, Gol65, Gol83, Gol85a, Gol85b, GP85, Gri92a, Gri94b, Gri96a, Gri96b, Gru76, Gur68, Hat75, Haz95, HJ78, Her91, Her92, Ibr62, Ibr63a, Ibr67a, Ibr75a, Isa85, IT79, Ivc75a, Ivc71b, Ivc71a, IM87, JM94, Kal77b, KS64, KT59, KN94, Kar85, Kav91, KW71, Kel90, Kha87, Kha89b, Kha89a, Kha94, KA99, Kha61a, Kha66b, Khm84, Khr91, KR92, Kli79a, Kol66, Kol67, KC74, KC75a, Kol92, Kol56b, KS86, Kon82, Kor93, Kos85, Kov60, Kov67, Kov86, Koz71b, Koz78, Kru73a, Kru76, Kru82, KB97, Kry68, KS89, Kub96, Kud86, Kul67, Kul68, KH81, KK96, KT77, Lab72, LM71, LO96]. Limit [Lët89, Lët93, Lif76a, Lif76b, Lif79, Lin59a, Lin61a, Lin61b, Lin62, LS81b, LS81a, LS82, Mac71b, Maĭ90, Mak65, Mak80, Mal75, MA78, MA83, Mal87a, Mal87b, Mal90a, Mal92a, Mal94, Mal97, MU83, Mal85b, Mal65b, MO83, Mar84, MP93a, Mar76b, Mas74a, Mas60, Mat85, Mat84, Mat86a, Mat86b, McD80, Mes58, MSV99, MN77, Mih56, Mik74, Mik76, Mik77, Mik94, Mik97, Mik99, Mik70b, Mit62, Mna86, MR83b, Mol94, MPY70, Mos73, MFY74, MP79, Muk77, Muk85, Muk92, Nad93, Nag57, Nag61, Nag65, Nag69a, Nag69b, Nag69c, Nag71a, Nag75, NM76, Nag76a, NE77, NK81, NK82, Nag84b, Nag84a, Nai82, Nak88, Nor93b, Nos77, OP67, Osi75, OR85, Oso56, Pan87, Par87, Par64, Pau77, Pau79, Pav82]. Limit [Pav84, Pav89a, Pec77, Pec81, Per59, Pet59b, Pet62, Pet64, Pet66b, PL83, Pis82, Pis95, Pit73, Pit78, Plu92, Pol78, Pop68, Por67, Pre74, Pro56, Pro60, Pro91a, Pro92, Rak88, Rak89, Rak90, RS94, RS98, Rez81, Ric57, Ric58, Ric64, Ric65, Rie77, Rog58, Rog60, Rog66c, Rog74, Rom87, Ron80, RS83, Ron85, Ron86b, Ron88, RS76a, Ros80, Ros95, Rot70b, Rot74, Rot76b, Rot78a, Rot78b, Roz57, Roz60a, Roz61, Roz76, Roz78, Roz79, Roz82, Roz87, Rus96, RS76b, RS80, Ryc84, Ryz65, Ryz67, Ryz69, Sad66b, Sad67b, Sad68, Sal79, Sat73b, Saz67, SM79, Sei87b, Seĭ89, Seĭ91, Sel85, Sem85, Sen81, Sen82, Sen84, Sen86b, Sen86a, Sen87, Sen90, Sen92, Sen93, Sev57b, Sev66b, Sev68, Sev73, She80, She71]. Limit [Shi80, Shi82, Shl81, Sho79, Sho83, Shu76a, Shu90, Shu91b, Shu77d, Shu79b, Sil73, Sil79, Sin62, SF84, Sko56, Sko57a, Sko58, Sko63a, SS65b, SS66, Sla79, Sre70, Sta65, Ste69b, Ste93, SI67, Stu67, Sun96, Sun99, Syt67, Sza71, Sza73, Tik81, Tik85, Tim88, Tom71, Tro91, Tro93, Tut62, Tut65, Tut78, Ul’77, Ul’79a, Ul’79b, Uri93, Ute91, VIL69, Vat77b, Vat77c, Vat80, Vat82, VM83, VT82, Ven77a, Ven77b, Ven80, Ven83, Ven90, VS77, VS78, Vir70, Vir74, Vol57, VR59, VR61, Vol61b, Vol86, Vol96a, Vve85b, Yan96, YK86, YK87, Yus67, Zäh80, ZA84, Zai85, Zai87a, Zai88a, Zaĭ90, Zal86, ZM95, Zin65, Zin71, Zol64b, Zol66a, Zol90, Zol92b, ZM74, Zub74, Zub83, Zup92, GP97, KC75b]. limit [Mak88, Muk96, Shu91c, Svi90, Vat79, Bur75a, IK73a, Shu82b]. Limitations [Sac73a]. Limited [Man67]. Limiting [AT93, BYK88, Bet88, Kha93, Klo61, Kol69a, Kud85, LN68, Mar65, Mas99, Oli93, Saz65, Sin83, Zub76]. Limits [Ara68, Bol65b, Dev69, Lum69, Muc88, Shu89]. Lindberg [Bet89]. Lindeberg [IO66, JR82, Lin59a, Osi66, Zol67a]. Line [Bla83, Ibr78, Ivn90, Kon78, KU88b, Mir69, Ost89, Pit75, Sev62, Ula88, Val74b, Val82, ZI70, ZI71, ZI74, Val73]. Linear [AA68, Bol94, BK80, DT75, DL86, EN75, Fed71a, FN93, Gad90, Gih91, Gin89b, Gir92, GK81, Gol92b, Gor78b, Gri89, Gri94a, Gri94c, Gut64, Háj61, HP94, IK85, IK88, Iva77, Ivk68, Kag90, Kal76, Kha67, Kha61b, Kha61c, Khm81, Kle74b, KS57, Kor57, Kov67, Kov73, Lev86b, Lip66, LS68b, Lum69, Lus96, Mar69, Mat65a, Mat65b, Med98, Mil79b, MT85, Muk78, Mus73, Mus70, Pag83, Pir68, Pit71a, Plu92, PW87, Rad92, Roz61, Roz64b, Roz81a, Roz99a, Sha70, Sko79, Sko80b, Smi69, Tie79, TP57, Tsi76d, VTT81, Vak90, Vla56, Whi60, Yag60, Zhd90, Zin76, Zin80, GK80]. Linearities [Kha61b]. Linearization [Vet81]. Linearly [Anu91, Pre98, Smo85]. Lines [Aly88a, BB87, Gri74, Kru86]. Linnik [Ano65e, Ano73l, Chi70a, Chi95, Ost65, Ost67]. Lobachevsky [GV59, KT59, Tut62, Tut68b]. Local [Ale77, Amo88b, Amo91b, Amo91a, Bas77, Bel60, BJL98, Bor82a, Bor85a, Bor90b, Bor90a, Bor61, BM99, Chi57, Chi65b, Das79, Dav76c, Dav76b, Dav79, Dav89a, DM96, DK84a, DK84b, Dok99, Efr80, EO97, Eli89, EW99, Gam64, Gam66, Gam68, Gam89, Gam91, Gam92a, Geo80, GK84a, Gur68, Her91, IH76a, Kol67, Kon82, Kry68, Lip77, MU83, Mal85b, MT83, Mat86b, McD80, Mir78b, Mit62, MPY70, Mos73, MFY74, MP79, Muk77, Muk85, Muk90, Muk92, Muk98, NE77, Nas88, Nas96a, Nas96b, Nas97, Nik85, Nos85, Nos86, Pav89a, Pet56b, Pet62, Pet64, Rag91, Rak89, Rao71, Ric57, Ric58, Rog68a, Roz57, She71, Sht65a, Shu76a, Sir59, Sko66, Smo85, Tut78, VN99, Zak91, ZT92, ZM95, Zin79, Muk96]. Localization [AB99, Shu62]. Localized [Sav99]. Locally [Faj88, Fel79, FF82, Khr86, Klo62, Kul87, ML65, PS64, Tsa83, VT78, Vla84]. Location [Erm89, Erm90, Fin71, IH75b, Kag86b, Kle74a, KR75, Lin57b, Nev81, Ruk67, Tsi83, Tsi86, Tsi87]. Log [PR99]. Log-likelihood [PR99]. Logarithm [Bul78, Bul81b, BV88, CvZ85, CvZ89, Ego69, Ego70, Ego72, Ego85, Ego91a, Ego91c, Gap65, Gap69, GK75, Ios68, Ios70, Kle87, Kor88a, Kru88, Kul85a, MP77, MP78a, Mar79, Mar81, Mar86a, Mik89, Mik91, Mog80, NP99, Pet66b, Pet71, Rao89, Rez68, Rez71, Roz98a, Vas80, VD93, Vov88, Zhu88]. Logarithmic [Ste88a, Zai87b]. Login [Ano79a]. Loginov [KC79]. Lognormal [Lei82]. loi [Tou92]. Long [Bel69a, Mik74, Mik76, Nov95, Nov96, Pro65b, ST73, VW99, ZM74]. Long-Normal [Pro65b]. Longest [Sam82, V′y98b]. Longuet [Mir77, Mir81]. Longuet-Higgins [Mir77, Mir81]. Look [Ken95, SS95]. Looking [Ken95]. Loop [GE91]. Loss [Lep88, Sap79, Tik78, VP64]. Losses [Tik82, ZK92]. Lossy [Fal88]. Lots [Kor58, Kor61a]. Lower [Alb93, Ali87, Ara83, Ark87, Ark88, Ark90, Bag88, Bag89a, Bag89b, Gol92a, Kha79a, Mat85, Mat84, Mat86a, Mat86b, Mik91, Muc88, Sen86a, Sun99, Vol79, Vol80, Vol81, AP97, Bak89b, Kha79b]. Lowest [Vla56]. LS [Gap86]. Lublin [Ano99d]. Lukacs [Fel90]. Luzin [Tsi76b]. Lyapunov [Bor92a, Mil82, NR73, NR76, Vir84].

M [Hoe64, Lis96]. M. [Bol64b]. machine [Sob58]. Machines [Fre88]. Made [Ano75i, Mir84c]. Main [BP91]. Majorization [Leb87]. Majorized [Shu66a, Shu66b]. Majorizing [Ber80]. Make [Gor81]. Making [Pit69a, VGR84]. Manifold [Dyn98, Fre63, Pus92, MP97]. Manifolds [Kif76, Mol68, Rez77b, Sob73a, Sob73b]. Manipulation [Kal93b]. Mann [Bul85b, Gel74]. Manufactured [Kor58, Kor61a]. Manufacturing [Sir63a, Sir63b, Vor65]. Many [Pav81a, Pav88, Pre65, Rot76c]. Many-Server [Pre65]. Mapping [Gru76, Pav82, Pro74]. Mappings [Den89, Dor90b, GPU98, Gru73, JR89, Kol76, LP75, Pan94, Pav89a, Sac73a, Sac73b, Sev72, Ste69b, Ste71, TT83, VM97]. March [Ano64e, Ano66g]. March-May [Ano64e]. March-November [Ano66g]. Marcinkiewicz [Fel90, Gia96]. Marginal [Mak82a]. Marked [CGM97, KS86]. Market [KS95]. Markets [KK95a, Sch93b]. Markov [Ali95a, Ali95b, Amb79, Ano73g, Ano89a, Anu81, Ara68, Arn81, Ash68, AS82, AN93, Bal94, Bar77, Bas58, Bas77, Bel65, Bel67a, Bel69b, Bel78, Ben76a, Bez75, Bez78, Bez71, Bez80, Bez82, BO97, Bor81b, Bor91a, Bor92a, BU94, BK97, Bra91, Bui82, BS89, Bul87, BP73, Che89, Che60, CS61, CC96, Chi83, Chi76a, Chi76b, CS60, Coh62, Cow93, Dav74b, DJ99, Den89, DKR95, Dob56a, Dob56b, Dob56c, Dob57, Dyn56a, Dyn56b, DY56, Dyn58, Dyn59, Dyn67, Dyn68b, DY68, Dyn70, Dyn71, DRS89, Elb76, Eng71, Eng74, Eng75, EW99, Evs78, ES69a, ES69b, ES77, Fai76, Fai79, Fai80, Fai82, Fai83, Fak71, Fal78, Fil84, Fri69, GR67, Geo80, Gik58, Glo70, Glo79a, God82, Goo59, Gor84a, Gor85, GE93]. Markov [GW98, GS66, Gri79, Gur68, GE91, Ibr75b, Iva65, Kal76, Kal62, Kal69, KS80, Kar85, Kar86a, Kar86b, KS61, Kha77, Kha81, Kha84c, Kha63, KFS77, Kit82, Kit86, Kli79a, Kog68, Kon70, KS84, KS85a, Kor94b, KY64, Kry66b, Kry87b, Kuz74, Kuz75, Kuz80, Kuz81, Leb74, Lev72, Lev73, Lev74, Lev75a, Lev78, Lev80, LK74, Lev75b, Lif76b, Lif79, Lip66, Lip67, LS68a, LS68b, Lip97, Lod71, Mak65, Mak70a, Mak80, Mal91, Mal86, Mal87a, Mal87b, Mal90a, Mal90b, Mal92a, Mal92b, MP58, Mes58, Mey63, Mih56, Mil72, Mir80a, Mir81, Mit88, Mol67c, Mol67b, Nag57, Nag61, Nev69, Nik99b, Nik64, Nik80b, Ose65, Oza86, Par71, Pit73, Pit69b, Pop80, Por67, Pre67, Rom65, RS83, Ros64, Roz77, Roz84a, Roz88]. Markov [Ryk66, Sam82, Sar61, Sat95a, Sat96, Sen68, Sev57a, Shi64b, Shi67, Shp66, Shp68, Shu58, Shu62, Shu64, Shu66a, Shu66b, Shu67, Shu68, Shu70b, Shu76b, Shu77a, Shu78b, Shu78a, Shu80a, Shu82a, Shu85b, Shu77c, Shu85c, Shu87, Sīl92, Sin60, SF84, Sko58, Sko60, Sko63b, Sko63a, Sko64a, Sko66, Sko67, Sko87, Sob73a, Sol94, Son92, Sta84, Str60, Sur63, Svi90, Tak77, Tan70, Tom90, TM79, Tsi76d, Vas70, VMS95, VT73a, VT82, Ven77a, Ven77b, Ven80, Ven83, Ven86, VJK59, Ver94, VGR84, Vol57, Vol58a, Vol59, Vol61a, Vol58b, Vol60b, Vor61, Vor63a, YF79, Yus57, Yus60, Yus74, Yus76, Yus78, YF80, Yus80, Yus81, Yus82, Yus90, Yus91a, Zak90b, Zas73]. Markovian [Ave72, Bor83a, Mar93, Sta97]. Martin [Mol67c, Mol67b, Shu68, Shu69a, Shu69b, Shu70a, Shu73]. Martingale [BP89a, Big93, But95, But92b, Chi90, Dob58a, Ers75a, KK95b, Khm82, Khm84, Kub96, Kul74, Leb84, Leb86, Lu’93, Mel83, Mis91, Pav97, Pav98, Tro93, Uri93, Ers74, Svi90, Vov97]. Martingale-Like [Lu’93, Mel83]. Martingales [Ben62, Bor79a, Cad99, Ego91c, Ers79, Gal77, Gal78b, Gal78a, Gal81, Gal83, GM84, Gus93, Kir91, KR92, Lip76, Lip77, LS84, MW99, MP91, Nov75, Nov80, Nov97b, RR99, Sho99, Sko63b, Sko75, TT82, Tie80]. Maslov [DD99]. Mass [BP91, Bor64b, Bor65a, Bor65c, Khi62, Rac85a, Rog66b, RS70]. Masses [GH93]. Match [Nov95, Nov96]. Math [JK90]. Mathematical [Ano59a, Ano59b, Ano60a, Ano61a, Ano61e, Ano61g, Ano62c, Ano62d, Ano62e, Ano63c, Ano64c, Ano64f, Ano64e, Ano65c, Ano65d, Ano65e, Ano66g, Ano67g, Ano68g, Ano69g, Ano69h, Ano71c, Ano71f, Ano71g, Ano73h, Ano73k, Ano73j, Ano73i, Ano74a, Ano74h, Ano74j, Ano74i, Ano74g, Ano75i, Ano75j, Ano75k, Ano76e, Ano76d, Ano76f, Ano76g, Ano77g, Ano77h, Ano78i, Ano78h, Ano78j, Ano78k, Ano78f, Ano79f, Ano79i, Ano79j, Ano80f, Ano80i, Ano80j, Ano80k, Ano81a, Ano81h, Ano81i, Ano81j, Ano81l, Ano82a, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82o, Ano82m, Ano82h, Ano82j, Ano82k, Ano83e, Ano83d, Ano83c, Ano83f, Ano84h, Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n]. Mathematical [Ano85m, Ano85o, Ano85r, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87o, Ano87l, Ano88a, Ano88g, Ano88k, Ano88l, Ano89k, Ano90b, Ano90j, BGS90, Bol81, Gal95, GKPS60, Kor66, Kor68, Leo82, MS90, MT90, NK79, Pro83, Tut69a, Zhu90, Ano64d, Ano67f, Ano77i]. Mathematics [Ano64g, Ano64d, Ano67f, Ano68f, Ano80j, Ano82n, Ano82m, Ano83c, Ano84c, Ano84d, Ano84j, Ano85k, Ano85n, Ano87k, Ano87j, Ano87m, Ano87o, Ano88l, EK94, SS90]. Matrices [Bal67, Bal68, Bar85, Bez82, Dub69, Fre68, Gir76, Gir77, Gir83, GV86, Gir91, Gir96, Kal77b, Mak65, Mak97, Mik75a, MT85, NRV67, Rez81, Roz60b, Rya79, Ser96, Tut65, Tut68a, Tut73, Tut78, VM97, Vir80, Vir84, YK87]. Matrix [BYK88, Chi83, Dob62a, Gir96, Glo70, KL77, Kov73, Koz71b, MW59, Raf90, Rez78, Sha80b, Shm64, Vil63, YK86]. Matrix-Valued [MW59]. Max [Gri94c, Gri96b, MR93, Pan89, Zem93]. Max-Infinite [Zem93]. Max-Semistable [Gri94c, Gri96b]. Max-Stability [Pan89]. Maxima [AT93, Bab64, Bor60b, Bor61, Gik58, Kar95, Nos85, Pit82, RR93]. Maximal [DSS94, Dzh94, Gir91, HS93, KP84, KB97, Lam72, Lis76, Saz65, Vin68, VG90]. maximale [HS93]. Maximum [Afa91b, Ale77, Ale79, Ale80, Amb79, AN74, Ara75, AK70, Art89, BCK^{+}92, Bor60a, Chi74a, Cra65, Den84, Gam74, Glo82, GW98, GK69, Ibr67b, IK74b, IH76b, IK81, Kha87, Kon79, KP83, Kor59, Kor62, Kor86c, Kut85, Lep94, Lif87, Mit67, Mos91, Nag70a, Nag70d, NE77, NP69, Nev71, Nev72, Nev88a, Nov92, Nov98, Ost91, OT95, Ost98, Pak59, Pan87, Pav78, PY88, Rog66c, Rog88a, Sht65b, Sht66, Shu65, Str78, SU90a, Tik87, TV98, Tsi76c, Tsi83, Tsi86, Tsi87, Ush86, Vat96, Wal93, WW66, WW68, Wol65, Wol76, Yus73, MP97]. Maximum-Likelihood [Kha87]. May [Ano57a, Ano62c, Ano64e, Ano74g, Ano82k, Ano83d, Ano84j, Ano84k, Ano85l, Ano85k, Ano87i, Pet59c, Anu91, Son92, Vla80]. McMillan [Per59]. Mean [Ara99, BD97b, Bul61, Che67a, DB88, DKR95, IS74, KNS66, Kho73a, Kla85, Kle71, Kle74a, Kle80, KY84, KN90, Kru82, MN99, Mil88, Mir78c, Nad74b, Nag71a, Nos83a, Ol′89, PG80, Ryz71, Ryz72, SM62, Tie80, Tsa77, Vil59, Vol80, Vol81, Zol64b, Roz97]. Mean-Square [Mil88, Tsa77]. Mean-Variance [DKR95, MN99]. Means [Bet86, Dob68, IM71a, IM71b, Kov86, Kud84, Kur88, Leo61, Nev88a, Pag68, Pol61, Rie81, VC82, Vol61a, Zai84]. Measurability [Sko80a, Zhd90]. Measurable [Kha96, Sko79, Tie79]. Measure [BS74, Gam86, Gär77, GP79, Kha85, Kon78, Kon85, Leb84, Leb86, Lin90, Llo62, MP92, Nad77, Nev64, Pau91, Pes94, Reb88, Roz62, Tar90a, Vas70, Vla80, Yan79, Ivk94]. Measured [Bar91]. Measurement [Bel95, Páz68, SL93, Bel94]. Measurements [Bog82, Kho87, Mir79b, Mir84c, Rya76]. Measures [AB99, Ale63, Apo67, Bau88, Bau96, Bog89, Bul85a, Cap98, CN87, DF65, DP75, DS85, DR78, Dok93, Ers72a, Ers75b, Fel65, Gal78a, Gal83, Gal85b, Gir60a, GT77, HMV95, Ign82, Jur83, JR89, Kab86, KK95b, Koz84, Kra74, Leb90, Leb96a, Leb96b, LP75, LTC80, Lip97, Lod71, Mak71a, Mak73, MP93b, MR89, Mus73, Mus85, Nik80b, Nor88, Pla96, Pro92, Rac85b, Rad92, Rad98, Raj73, Rén56, Roz63, Roz64a, Sar90, Shu82a, Sin63, Sko57b, Sko60, Sko70b, Sko70a, SY73, Smo89, Sok84, SO85, Syt69, TT80, TT83, Th93, TTC78, Tie79, Tol99, Tro93, Tsi76b, TM79, Tut61, Tut62, Tut64, Tut67, Tut68a, VT78, Ven82, VS77, VS78, VKR64, Vla79, Vla81, Vla87, Vla96, Vor62, Vor63a, Vor63b, V’y88, Fat97]. measures [Min91]. Mechanics [Das79, MS90, SS90]. Media [MTV91]. Median [BD97b, Bur97]. Medium [Koz74a, Sin83]. Meeting [Lut78]. Meetings [Ano61g, Ano63c, Ano64g, Ano64f, Ano64d, Ano64e, Ano65c, Ano65d]. mélange [Tou92]. Mellin [Zol57a]. Memorial [BBD^{+}94]. Memoriam [Shi90]. Memory [BP70, Mit88, SS65a, Tsa58a]. Memoryless [Dob62a]. Meromorphic [SK92]. Messages [Dob62a]. Method [BD83, Ber70a, Ber70b, Bor89b, Bor92b, BL63, Dar76, Dar86, DB88, Dob58c, Dza74, EZ60, Erm66, Fet74, Gam81, Glu73, Gri92b, IMS93b, Kal77a, Kar75, KC79, KP97, KSS77, KC91, Kor66, Kor68, Kos88, Koz73, Koz81, Kur87, LS59, Lif82, Lif85b, Lin57b, Lut90, Lut91, Maĭ91, Mik75a, Mil79a, MPY70, Nag86a, Nik88a, Sal97, San66, Sed77, Sob60, Sol67, Tur71, WS93, Yak98, IMS93a, Roz97, TC86]. Methods [Ale64, Ano59b, Ano60a, Ano87o, BD91b, But95, Che67b, Der93, Dor90a, Dzh71, Gap65, Kar85, Mar86b, Mik56, PP93, PT93, Pro83, Shi63b, Shp68, Tsi91, Vla56, Gil93b]. Metric [Ben62, BS74, Bur75a, Bur75b, Kai86, Kha79a, Roo99, SS62, Sed77, Sen85, Zie93, ZS76, Kha79b]. Metrics [Gor84a, Ing87, Jam78, Nik91, RR91, RR93, Szu83, Zol78c, Zol84a, Zol84b, Zub87]. Meyer [JK90]. Microscopic [Mon56]. Middle [Ivc75a]. Migration [BY96, Fal89, Kru69, NK81, NK82, YM84]. Miller [Fer98]. Miller-Sen [Fer98]. Minimal [BG96, BG73, Gir91, Puk82, Rac85b, RR91, Sha70, Szu83, Tim85, Tim89, Vat94a, Zez75, Zhu66, Zub87]. Minimax [Bog82, Bun70, Bur79, DL86, Erm91, Erm95, Gor81, Ing87, Ing88, Ing89, Ing95, Kor86c, Kor88b, Kor88a, Kor92, Kor94a, Kuz83, LS71, Lep88, Lep90, Lep93, Lev81, Lev83, Lin66, LPS67, Mal74b, Pin91b, Pin91c, Pit71b, Rom62, Sap79, Yan64, Lep92]. Minimaxity [Erm96, Erm98, Lus99]. Minimization [Kar75]. Minimum [BM93b, Chi74b, JP75, KP67, Kha84b, Kis62, MF71, Raf90, Sir56, VN94]. Minimum-Contrast [Chi74b]. Minimum-Variance [KP67]. Mise [Fil62]. Mises [Ale91b, Csö75, Kha89b, Kha89a, Khr96b, KB89, LS96, Orl75, Pos61]. Missing [Ano61b]. Misspecified [Fin93, GW98]. Mixed [BP89a, Bul75, But79, But92b, Pav98, Zhu70]. Mixing [Bak96a, Bel61, Dav74b, Deo77, DM80a, DM80b, DK97, Eag77, Gap92, Gor78b, Gor83, Gor84b, HP94, Ibr65b, Ibr70, Jan93, KR60, KR90, Lin92, Mit88, RS88, Son92, Ute91, VL76, Ver88a, Zhu72]. mixture [Tou92]. Mixtures [Bag89c, Bet86, Kli57, Kor90a, KU88b, Maĭ95]. Mobile [Dok99]. Mode [BD97b, Gri92b, Kon74a, Rez77a, Rez78, Rez80, ST73]. Model [AS92, Bag79, Ble86, BG91, Bol68, Dob73, Gus93, Hau93, KS95, MT85, NS70, Nai82, Nos77, Pag83, Piu98, Pop68, PP95, Rie80, Rya66, ST73, Sta97, Ter93, Tsu81, Vol93, Yus76, Yus80, OK97]. Modelling [Mik65]. Models [Ano81b, Ano82b, Ano84g, Ano86m, Ano89c, Ano90c, Ano92d, Ano93d, Ano98d, Ano99d, BD93, BP93, Dar96, Dob65, Dob66, Dok93, DK93, EK82, Fai87, FN93, Fin93, Fra93, Gad90, GN93, GW93, GW98, Kit82, Kit86, Kra94, MR96, MS88, Mol92, Pes94, PP93, Puk82, RR95, SL93, Shi95, Tut69a, YF79, Yus74, Yus78, YF80, Yus81, Yus82]. Moderate [Amo80, Sla79, Top89]. Modification [DN75, Kag86b]. Modifications [Ale98, Gal78b]. Modified [Tho93a]. moduli [Tsi97]. Moment [BG96, BD91b, Ego73, Gap92, KR76, KST60, Leb77a, Lei82, Maĭ90, Nag98b, Nov71a, Rog60, Rog71, Sei87a, Ute90, Vak94b, Vol64, Zie93]. Moments [BG98, Bor61, BMS96, Bul94, Chi80b, DM79, Dyc73, Ebr71a, Gus76b, IK74b, IC98, Kha61c, Kor86b, Kru89, KS89, Kub96, LS60, Lif76b, Mal80, Mal85a, MV89, MV94, Mil82, Mir74, Mir75, Mir78a, Mir85, Mog74, Nag86a, Nev86, Pet74a, Pet76, Pët93, PU85b, Pra99, Pus93, Ron82, Rot67, Saz75, Sch65, Sem88b, SC71, Shi60, Shi63a, Sla87, Ste88a, Tei66, Vat79, VK86, V’y77, Whi60, Zai87a, Zai88a, ZS84, Zez75, Zhu68]. Monge [Rac85a]. Monotone [Aly88a, Aly88b, KM89, KS97, Kry91]. Monotonicity [KFS77, Kur85, Kur88, Vve89]. Monotonous [Pan94]. Monro [Kra67]. Monte [Bor92b, EZ60, Erm66, Fet74, Koz73, Koz81, Mik70a, Sob60, Tur71, VTGM85, Vla56]. Monte-Carlo [Koz73]. Mosaics [Amb74]. Moscow [Ano57a, Ano59e, Ano60c, Ano61e, Ano62c, Ano62d, Ano64g, Ano66g, Ano67g, Ano68g, Ano69g, Ano71f, Ano73i, Ano74h, Ano74g, Ano75j, Ano76e, Ano76d, Ano77g, Pet59c, Ano61d, Ano64f, Ano64e, Ano65c, Gne90, MS90, MT90, Sev90, SS90, Zhu90]. Most [Bel67a, Kul87]. Motion [Amb65, Ara99, Che57a, Doo57, Dzh94, Fer98, GV59, IK64, Kif76, Kir94, McK63, Mol67a, Ouk91, Ron86c, Ski56, Sko61a, Ven68, ZI70, OK97]. Motions [DV96, Kho83, Roz73]. mouvements [Tou92]. movement [Tou92]. Moves [Kif69]. Moving [Bra94, GP85, Oni90, Vol86]. MR [Bak89b, Fry90, JK90, Shu91c, SZ98, Ul′91]. Multi [BM91, Bas58, BR65, Bul94, Dob58d, Fom83, Gam78, Gla58, Kho83, KMT78, Mal85b, Mat61, MS79, Mog77a, Muk78, Ric58, Ric64, Rot70a, Rot78a, Rot78b, Sch79, Sei82, Sen81, Sen82, Sen86a, Shi80, Shi82, Shu76c, Vol57, ZI70, Zas67, Zin80]. Multi-Channel [Dob58d, Sei82]. Multi-Dimensional [Bas58, BR65, Fom83, Gam78, Gla58, Kho83, KMT78, Mal85b, Mog77a, Muk78, Ric58, Ric64, Rot70a, Rot78a, Rot78b, Sen81, Sen82, Sen86a, Shi80, Shi82, Shu76c, Vol57, Zas67, Zin80, Mat61]. Multi-Indexed [Bul94]. Multi-Type [BM91]. Multi-Variate [MS79, Sch79]. Multi-Wave [ZI70]. Multichannel [Akh80, Bor68a, Bor73a, Ivn93, Pav92]. Multidimensional [Ale94, Ale84, Ark88, Bha70, Bor68b, Bor91a, Bor92a, Dok69, Dyc73, Ebr71b, Fin71, GO92, Gam86, Gam92b, Gia96, Gor76, IV93, IM65b, Jur72, Kon74a, LS68a, Man68, Mat60, Mir88, Mir90a, Mir90b, Osi75, Ost65, Ost67, Pan91, Pre74, Pro68c, Pro68d, Pro68b, Pur89, Rot70b, Sad68, Sag91, Saz67, Saz66, Sem90, Sen93, SC71, She71, Sho79, SS92, Sol94, Tur71, Tyu74, Ven59, Ven68, Vin90, Yam95, Yur89, Zaĭ90, Zol66c, Zol68b, PvZ97]. Multifactor [MO91]. Multilinear [Bas93, Rot76b]. Multimodal [Sat95b]. Multimode [Tut71, ZI71]. Multinomial [Are77, Bel69a, IJL91, Ivc71b, Ivc75c, IL76b, IT89, KC75a, Mat78, Mir86a, Mir88, Ron84, Roo99, SMI90a, Yus73]. Multinornial [San66]. Multiparameter [Dav79]. Multiphase [Vin96]. Multiple [Bor65c, Dav91, DM96, Kha90b, Kle96, KH81, Mik74, Mik96a, Nik99a, Nik99b, VT92]. Multiplication [Lev72, Lev73, Lev74, Lev75a, Lev78, Lev80, Lev88]. Multiplicative [Abr87a, But77, Gap69, Lis98, Mey63, Mol94, Pan91, Vol64]. Multiplicities [Sir73]. Multiplicity [Gru73, KM65, Sir77b]. Multiply [PU85a]. Multiresponse [Fed71b]. Multiserver [Vin95]. Multitype [BM90]. Multivariate [AL81, BYK88, Ber81, Ber87, Bik69a, CdHY97, Che67a, Epa69, Fra93, Gil93a, Gil93b, Kha65, Lis98, Maĭ91, Mik75b, MR93, MRS94, PS75, Roo99, Roz60b, Sol95, Yag60, Zai87a, Zai88a]. Multiwave [ZI74]. Mutual [Glo79b].

N [Bin90, BK71, Gin66, GS63, KC79, KY86, KS60, Pro60, Pro68c, Pro68d, Pro68b, SS90, Vak90, Yur70]. N. [GKPS60, Mal68]. Nadaraya [Ral97]. Naleczow [Ano99d]. Narrow [Amo91a, Roz82, Sla85]. Natanovich [Ano69f]. Natural [BNK96]. Nauka [Ano87f]. Near [Den84, Kol87, Shu69a, Shu69b, Ste88b]. nearly [PvZ97]. Nearness [Glu73]. Necessary [Fak69, Kha67, Khm75, Kul83, LS81a, Mak67a, Mak68, MP77, Mar80, Mir75, Mus81, Nag73, Nem91, Sho99, VC82, Vol78a, Vol78b, MP78a, Ibr65b]. Necessity [Vak94b]. Needed [Vol67]. Negative [Afa79, Afa87, Ara81, BM79, Fre68, Kif76, Kos89, Nov80, Shm64]. Negligible [Shu73]. Neighborhood [Ble86, Nos86]. Neighboring [Aiv65]. Neighbourhood [Tik88]. Network [Afa88, Ivn98, Mal91, Sei87a]. Networks [Bor87, Bor88b, Bor89a, Bor98, Ivn82, Ivn90, KKR89, Kel90, Pav92, Tsa70, Tsi91]. Neumann [Mil97, QT97]. News [Ano73g, Ano81e, Ano84c, Ano84e, Ano84d, Ano85h, Ano85i, Ano86f, Ano86e, Ano86g, Ano86h, Ano87g, Ano88f, Ano88g, Ano90g, Ano91e, Ano91f, Ano92b, Ano92c, Ano93c, Ano94c, Ano94d, Ano94e, Ano95c, Ano96b, Ano97c, Ano99b]. Neyman [Aiv59, Kol88]. Nicholas [Yus87]. Nikolaevich [Ano79a, Ano88h, Shi90]. Nikolaevitch [BBD^{+}94]. Nikolai [Ano66d, BBD^{+}94]. Nilpotent [Tut64, Vir74]. No [Che56, KK95b, Tsi76a, Bak89b, Fry90, Shu91c, SZ98, Ul′91]. No-Arbitrage [KK95b]. Nodes [KC91]. Noise [Ara85, BCK^{+}92, Bol83, Bur79, But77, Che57a, Dob58d, Erm91, GD71, GK81, GK84b, IK85, INK87, IK88, Kra94, Lep91, Lin66, LPS67, Nem91, Pre89, Ras77, Vol64, Won59, GK80]. Noiseless [Kat75]. Noises [ZI71]. noisy [Glo79b]. Nomogram [Bol57, JL56, Kot64, SP57, SP61]. Non [Ara81, Bei64, Bel64, Bel65, Ber76, BS68b, Bor62, BS74, Bun70, Che80, Chi71b, DT75, Epa69, Erm89, Fai82, Fre68, Gil82, GV86, Gir59, Gir62, Gol83, IM71b, IH76a, IH76b, IK82, Iva77, IT79, KM65, Kha79a, Kha61b, Kha61c, Kle74a, KS57, Kog69, Kog70, KC75a, Kon73, KL77, Kos89, Koz77a, Kry68, Leh65, Lei82, LO96, Lev76, Lif76b, LS68b, Mac71b, Mak68, Mak70a, MU83, Mal85b, Man64b, Nad65, Nad70, Nov80, Ost91, Pet62, Rot70a, Rot76a, Rot78a, Rot78b, Ryc84, Sen82, Sen68, Sev67a, Shi64b, Shm64, Sho79, Shu73, Sla82, Sla85, Sob73a, Sta97, Tik78, Tik82, Tyu70a, Ul’77, Vak66, VL76, Vat77a, Vil68, Vol61a, Kha79b]. Non-Additive [Shi64b]. Non-Appearance [Bel65]. Non-Classical [Mac71b, Rot76a, Rot78a, Rot78b, Sho79]. Non-Coalitional [Vil68]. Non-Commutative [Gol83]. Non-Degenerate [Gir59, Vak66]. Non-Deterministic [KM65]. Non-Equilibrium [Bei64]. Non-Equiprobable [IT79]. Non-Existence [BS74]. Non-Extinction [Chi71b, Koz77a, Sev67a, Vat77a]. Non-Gaussian [LO96, Ost91]. Non-Homogeneous [Lif76b, Mak70a, Sen68, Vol61a]. Non-Identical [Mak68, Pet62]. Non-Identically [IH76a, IH76b, Kry68, Sla82, Sla85]. Non-Latticed [Bor62]. Non-Linear [DT75, Iva77, Kha61c, KS57, LS68b]. Non-Linearities [Kha61b]. Non-Markov [Sob73a]. Non-Markovian [Sta97]. Non-Negative [Ara81, Fre68, Kos89, Nov80, Shm64]. Non-Negligible [Shu73]. Non-Occurrence [Bel64]. Non-Occurring [KC75a]. Non-Parametric [BS68b, Bun70, Che80, DT75, Epa69, IK82, Kha79a, Kon73, KL77, Lev76, Nad65, Nad70, Tyu70a, Kha79b]. Non-Quadratic [Kle74a, Tik78, Tik82]. Non-Randomized [Fai82].

Non-Selfadjoint [GV86]. Non-Separable [BS74]. Non-Smooth [Erm89]. Non-Standard [Ber76]. Non-Stopped [Kog70]. Non-Terminating [Kog69, Man64b]. Non-Uniform [Gil82, MU83, Mal85b, Rot70a, Ryc84, Sen82, Ul’77]. Non-Uniqueness [Gir62, Lei82]. Non-Vanishing [VL76]. Non-Verifiability [Leh65]. Non-Zero [IM71b]. Nonanticipating [Nor95]. Nonasymptotic [GO92]. Nonclassical [LS84, Sho83]. Noncommunicating [Son92]. Noncommutative [Den92, Kho94]. Noncompact [Lip97]. Nondegenerate [KB86]. Nondeterminism [Mir78b]. Nondifferentiable [Kut89]. Nonextinction [BD97a]. Nonhomogeneous [But79, Mak73, Mik98, Pag83, Shu76b, Son92, VT92, ZM95]. Nonidentically [Mar87b, VD93]. Nonlinear [BPRS94, Dav82a, Del96, Dyn98, Gur89b, Hau93, Lip66, Mas99, Nov83, Rag86, Rag91, Rag93, Roz99a, San66, Tan70, Vet81, Ivk94, PR97]. Nonlinearity [Bon92]. Nonmetric [Jak98]. Nonnegative [AI99, DRS89]. Nonnested [Fin93]. Nonnuclear [Bal94]. Nonparametric [Abd99, BO98, BD91b, CL94, Dar76, Dar86, DB88, Efr86, Efr90, Erm95, Gik57, Gin89a, Gol92a, GN93, Gol94, IK85, INK87, Ing87, Ing89, Kha84a, Kha85, Kor89c, Kos85, Kos99, Lep94, Maĭ91, MO91, Nad74b, Nem91, Nik88b, Nys87, Pen93, Rud93, Ter93, Tim88, Tsy88]. Nonquadratic [ZK92]. Nonrandom [FTV92]. Nonrecurrent [Nag87]. Nonregular [Kul83, Wal93]. Nonsmooth [Erm90, Kor91, Mel79]. Nonstationary [Ale94, Coh62, Dob56a, Dob56b, HP94, Oli93, OT95, Sun96, Bol97]. Nonsymmetric [Glo82, Lep88]. Nontrivial [Sap88]. Nonuniform [BO96b, Kir91, Kor89a, Sen89]. Nonzero [Mik99, Zai84]. Norm [BP89a, Kor94a, Nad74a, Pav98, Ron86c, SL90, ZKC91, Bau97, Nad73a]. Normal [AL81, AK95, Amo91a, AL91, Bag89c, Bar86, Bar88, BS73a, Bol57, BU84, BPS97, Bra86, CPU93, Che67a, Chi89b, Dev69, Dok69, Ego73, EN75, Gab75, GH93, Gri89, Jur83, KS67, Kha66a, Kir91, Kni66, Kor89b, Kor89a, Kor90a, Kru92, Lin56b, Lin64, LE68, Man68, Mir86a, Mir86b, Mir88, Mkr81, NR73, Nag86b, NS89, Nev71, Nev72, Nhu68, Nyb86, Ol′89, Pag68, PG80, Pec74b, Pon79a, Pos61, PF57, Pro65b, Puk82, PW87, PS89, Pus92, Ros75, RS82, Roz90c, Ruk67, Saz90, Sem88b, Sho99, SS92, Sla85, Tik92a, Tik94, Tyu74, Ul′91, Vak64, Vin86a, Vin86b, Vis86, Vol73, Yur78b, Yur83, Zal83, ZS84, Zal89, ZSU89a, ZSU89b, ZSU89c, Zas67, ZL64a, ZL64b, Zin80, Zol68a]. Normality [Amb76, BM92, CV65, Efr80, Eid73, IH76a, IL76b, IL78, IM89, Kho71c, Kon74a, Kut76, Mal69a, Mar76b, Mik80, Muk78, PR99, PS69, SC64a, SC64b, Ten91, Whi64, ZM79a]. Normalization [Bul78, Mol94]. Normalizations [BS88, Gri94c]. Normalized [GV86, Mar85, Pon86, Sem88b, Sre70, Zin65, Bak89b, Ego97]. Normally [Bag88, Bag89a, Bag89b, Mar76a, Pon79b, Pon86, Pro93, Bak89b]. Normed [Bon98, BS92, Tro91]. Norming [Gri92a, Gri94b, Kla85]. Norms [MC84, Nor98]. Note [Amo91b, AS74, Bor85c, Deo77, Fre65, Goo59, Ibr75a, Kra62, KR90, Kru70, Kru71a, Nev79, NV85, Pau91, Pin69, Shu76c, Sko70a, Ten91, Tsa58a, TP58, Tur91, YTS89]. Notes [Bor73b, JK90, Nad73b, SK92]. November [Ano61e, Ano64f, Ano66g, Ano71f, Ano77g, Ano87m]. November-December [Ano61e, Ano64f]. Nuisance [Ber81, BL66, FN93, PN92, PC96]. Null [BS88]. Number [Amb79, Bak98, Bal74, BNS66, BM63, Bel64, Bel65, Bel66, Bel67b, Bel68, Bel69b, Bel78, BJL98, Bet88, Bet89, BB81, Bor65a, Bor78c, Bor96, BG73, Bul61, Eli85, GP71, Gik58, GK81, Gru76, Háj61, Haz95, IM71a, IM71b, Iva65, Iva60, IN78, IM66, Ivn93, KN94, Khm75, KMM85, KC75a, Kop96, Kor58, Kor61a, Kor89b, Kov67, KS97, Leo86, Lis76, Mal69a, Mal80, Mal85a, MV89, MV94, Man67, Mas74b, Mas75, Mas99, Mes58, Mes63, Mik76, Mik80, Mik96a, Mik97, Mik98, Mik99, Mir74, Mir75, Mir78a, Mir85, Nag67, Nos79, Nos80, Nos83a, Nov92, Nov95, Nov96, OCG93, Pav89b, Pav89a, PL83, Pit78, PS73, Pro74, RS76b, RS70, Sac71a, Sac71b, Saf77, Sag91, Sen68, Sev67b, SK93, Son92, Ste69c]. Number [Sza73, Tom71, Top89, Tsa83, Tum56, TF91, Vat77b, Vat80, VM83, Vat94b, Vin76, Vol57, Vol58b, Vol60b, Vol65, Vol67, Vol69, YF80, Zin71, Bor77c, GK80]. Numbers [AL77, AV82, BGW75, Bol64b, BS87, BS88, De 82, Ego70, Ego72, Ego91c, Gap74, Gap78, Gap82, Kha69b, Kha69a, Khr96a, Koz71a, Koz73, Kru82, Kur85, KT77, Mar80, Mik96b, Mol88, Nag73, NV76, Pet69, Pit71b, Pro59b, Rao63b, Roz81b, Roz90b, Ser80, Sur63, Tut69b, Tut79, Ver66, Ver64a, VN78, Woy74, ZM79a, Chu94, Gap95, PR97, Pus97, QT97, Sob58, Vov97, Bor91b]. Numerical [Dau88, Fai83, Kru74, Mil97].

Obituary [Ano88h, Ano89g, Gne60, V^{+}97]. Object [V’y88]. Objects [Asa91, Ken95, PK82, PS73]. Oblique [Anu80, Roz67, Roz73]. Observable [Fäh83, Tor88]. Observation [Gad90, Gap86, Ivk68, Lut90]. Observations [Ano61b, BR94, Bor79b, BG73, Fra93, Had78, Háj61, IH76a, IH76b, Ivn77, Khm75, KP71, Kva96, Lev75b, Maĭ91, Mal86, MS79, Mir78c, Mir80b, Mir79a, Tsi78a, Tum56, Uba75, VM91, Vol67, Vol69, AP97]. Observed [IS74, Mal90b, Nov79]. Observing [Kor89c]. Obtained [Fal78, Hag74]. Obtaining [Dzh71, Kor66, Kor68, MPY70, San66, Sol67]. occasion [Kol62, KKSY95, KGPS91]. Occupancy [Ivc71a]. Occupation [Lip97]. Occurrence [Bel64, CP91, IN78, Sal90, Sol66, VK86]. Occurring [KS69, KC75a, OT95]. October [Ano60c, Ano62d, Ano64g, Ano68g, Ano74h, Ano75j, Ano76e, Ano76d, Ano83c, Ano85n]. October-December [Ano60c, Ano74h, Ano75j]. Off [LN91]. Omega [Erm96]. Omega-Square [Erm96]. ona [She99]. Once [KA85, KA87, Zak91]. One [Ben76a, Bet86, BZKR70, Bor65b, BK97, But92a, But95, Che96, Dyn59, Epa74a, Epa74b, GK76, GN87, GE91, Kal62, KNS66, Kha61c, Kle74b, Koz74a, Kul83, Kul85a, Lët93, Lis94, Lut90, Mak95, Mal91, Mal85b, Mar84, Mar86a, Mik65, Muk78, Nag82, NV85, Nov83, Par67, PS69, Rao89, Rih85, Rom59, RS82, Roz62, Roz86, Sag91, Sal99, Sel69, SK92, Shu79b, Sin83, Sir73, Ula88, Vak67, Ven61, Ver82, Vol59, Zin88, Zol92a, de 93, Gol65, PS98]. One- [Mal85b]. One-Dependent [de 93]. One-Dimensional [BK97, But92a, But95, Che96, Dyn59, GK76, Kal62, Kha61c, Koz74a, Kul83, Kul85a, Lët93, Muk78, NV85, Sin83, Vak67, Ven61, Ver82, Vol59]. One-Parameter [Par67]. One-Sided [Bet86, BZKR70, Epa74a, Epa74b, GN87, Mar84, Mar86a, Nag82, Nov83, RS82, Roz86, PS98]. One-Signed [Shu79b]. One-Type [Ula88]. Ones [Kel77, Mak82b, Pit83, Pre84, V’y77, Zub87]. Open [Afa88, GE91, Mal91]. Open-Loop [GE91]. Operating [Fre88]. Operation [Che64a]. Operator [BS88, BS92, But92b, Cho87, Dor87, Fed82, HMV95, Kar85, PS92a, Sav99, Sem88b, Sen86b, Sen87, Shu85a, Sie90, VK67, VT76, Zhd90]. Operator-Normalized [Sem88b]. Operator-Normed [BS92]. Operator-Semistable [Cho87]. Operator-Stable [Sie90]. Operator-Valued [VK67]. Operators [But77, DJ99, Dyn56a, Dyn56b, Elb83, Koc90, Kol92, Kri64, Küh81, MP58, Ome97, PS88, Ros64, Shu86, Sko80b, VT78, Vak94a, Vir80, Vla81, Zak90b, Lus96]. Operatorsand [Kul98]. Opposing [Fri69]. Optimal [Aiv59, Ark64b, Bar91, Ber78, Ber76, BS68b, Bor76, Bor99, BG73, Bul84, Cho93, CZ79, Dob62b, DSS94, Dyn68a, Dyn68c, Ebe86, Elb76, Elb83, Eng74, Eng75, Fäh78, Fäh83, Fai79, Fai80, Fak70, Fak71, Fer98, Fri69, Fri72, GD71, Ger79, Glo79a, GS65, GS66, Ing84, Ing86, Kha90b, Kho79, Kif69, Kif71a, Kif71b, Kit82, Koc98, Kog68, Kog69, Kog70, Kov59, Kry65, Kus68, KP71, Läu74, Lep93, Mak96a, MR96, Mih81, Mik56, Mir76, Nik99b, Pav91a, Pav92, Pit71a, PT93, Rom65, Rya66, Rya76, Sal90, Sap88, SS65a, Ter93, Tik78, VT92, Vet81, VM91, Won59, Yak91, Zig66]. Optimality [Bun70, Dob63, Fak69, Fed71a, Gus96, KM80, Lev86a, MP70, Nik85, Nov88, PT91, Pre98, Tsu81]. Optimally [Lep92]. Optimizing [BP73]. Optimum [Ark64a, Bul64a, Bul64b, Bul67, Erm89, Koz74b, Shi63b, Shm64, Vol64]. Option [GP98, KS95, RR95, SS95, KM95]. Optional [Gal78b, Gal85a, Gal85b, Kal93a]. Options [DKR95, Nov99, SKKM95a, SKKM95b]. Order [Ale91a, Ale94, Ale96, Ale98, BD78, BdH79, Bog67, BB87, Bul75, CdHY97, Chi64a, Chi95, Dub74, DL86, EN75, EN83, Ego91b, Gap74, Gap78, Ger73, Gri89, Gri94a, Ivc75a, Ivc71b, KN90, Kol69a, Kol92, Kru80, LS60, Lev81, Lev86a, Mal86, MSV99, Mik75b, Mil79a, Mil88, Mir80a, Mir81, Per92, Pet74b, PT93, PS69, PW87, Shi60, Shi63a, Sin63, Sir78, TTC78, Tik88, VT76, Ver64a, Zhu86, OK97]. Ordered [Cra56, Ego91a, Lic69, Nor93a, PS77, Rén56]. Ordering [Bul67, Kal62]. Orders [Gri91b, Sch79, Zez75]. Ordinal [DF93]. Organization [Ano90h]. Oriented [Evs89]. Orispherical [Tut73]. Orlicz [Bau96, GT77]. Orthogonal [Kan88, Kra79, Kra80, PT91, Sir77b, Tep95, Gap95]. Orthogonality [But77]. Oscillating [RF78]. Ostrogradsky [Smo91]. Other [Kis62, Sem90]. Outcomes [Bel64, Bel78]. Outgoing [Ivn77]. Outlier [Der93]. Outliers [Pag86, PR87, PC96, Pag98, Rod88]. Outlying [Mir79a, Mir79b, Uba75]. Output [Kha61c, Vin96]. Oval [Gya64b]. Overlapping [Kha69b]. Overs [Sīl92]. Overshoots [Nos88]. Overview [HMV95]. Own [SE96].

P [Chi70a, Leo82, San66]. P. [JK90, MC68]. P.-A. [JK90]. Paditz [Her91]. Pafnutiĭ [Ano97d]. Pair [Kni66]. Pairs [SS65a]. Pairwise [BD97a, Glu73]. Palm [Kok75]. Pancheva [SZ98]. Paper [BBM59, Rog60, Roz66b, Shi64a, SZ98, Bor65a, JK90]. Papers [Ano57b, Ano57a, Ano58d, Ano59e, Ano60c, Ano61d, Ano61e, Ano62c, Ano62d, Ano63c, Ano64g, Ano64f, Ano64d, Ano64e, Ano65c, Ano67f, Ano69g, Ano71f, Ano71g, Ano73k, Ano73j, Ano73i, Ano74h, Ano74j, Ano74i, Ano76e, Ano76d, Ano77g, Ano79g, Ano79f, Ano80e, Ano80f, Ano81h, Ano81i, Ano81j, Ano82q, Ano82g, Pet59c, Val74a]. Parabolic [Bla64, Bog67, Kha60, Kha63, Kog65, Kog66, Kor78b, Mil96, Mil79b]. Parallelepipeds [AR86]. Parameter [Ale64, Ara68, Ara85, Bab64, BCK^{+}92, BD83, Bla62, Bon92, Bur97, Che64b, Dzh71, Erm89, Erm90, Fak70, Fin71, Fre64, GS76, Ger79, Gla63, GC87, Gur89a, Has66, IH75b, IK93, Ing86, Iva77, KR67, Kag86b, Kah93, KS69, Kho83, Kle74a, KR75, Kog65, Kog66, KP97, Kor86c, Kru88, Kul76, Kul83, Kul85b, Kut78, Kut79, Kut89, Leb96b, Lep88, LK74, Lic69, Lut90, Lut91, Mal93, Mal95, Mis89, Mit67, Nev81, Nik80a, Nov79, Par67, Per92, PS93, Ruk67, Saf92, Sar90, She88a, Sko81, TP59, VT76, Vol69, WW92, Mak88]. Parameter-Dependent [Kru88].

Parameters [Aga80, Ale70, Art89, BN94, Ber81, Bol57, Bol62, BL66, Bun63, Bur75b, Che57a, Dha78, Dza70, Dza74, Dzh75, Evs87, FN93, GR67, Ger67, Gia96, Gus96, Ibr67b, Jes81, Kho67, Kot64, Kud74, Lum69, Lum71, Mol92, Nik74, Oni90, PN92, PC96, Pav81a, Sal93a, Tik88, Vol70, Vol73]. Parametric [BS68b, Bun70, Che80, DT75, Epa69, IK82, Kah93, Kha79a, Khm80, Kon73, KL77, Leh65, Lev76, Nad65, Nad70, Nik88a, She93a, Tyu70a, Kha79b]. Pareto [XP89]. Paricles [Mik77]. Part [BK97, Kud84, IK99]. Partial [Ara75, Ara90, BN83, Bor88a, BMS96, Chi80a, Chi80b, Chi87a, Chi88, EL99, Fro89, Glo79a, IK99, IC98, KS67, Kal62, Lin92, Mik76, Nov98, Pak59, PR69b, Roz88, Sla87, Vit70, Wit71]. Partially [Fäh83, FLR97, GW93, Lic69, Nov79, RS98, Tor88]. Particle [IM66, Kha68, Kha69a, Nag67, Sag83, VS90]. Particles [BD97a, Bev58, Bol68, Chi70d, Chi73b, Del96, Fil61, Gri87, IT79, Kal82, Kal99, Kha69b, Kol71, Kol76, Mik78, Mik80, Mon56, Pop68, Sag86, Sag91, SG62, Sel69, Sev66b, Top89, Vat77d, Vat77b, Vat79, Vat80, Vat81, VM83, Vat87a, Vat94b, ZM79b, Zub83, VS89, Vat95, Vol86]. Particular [WW68]. Partitions [Gur65, MR83b, Sac74, VM97, VKR64]. Parts [MP79]. Passage [AK89, Aly88b, AI99, Bor65b, Dob60a, GK68, Kal69, Leb71, MN77, Nov91, Rot67, Zol64a, Bar97, PS98]. Passage-Times [AI99]. Passed [Roz61]. Past [Gal95, VS92, Ven86]. Path [Mor75]. Paths [EK82, Tak76]. Pathwise [Dav82a]. Pattern [Bor71, Cow93]. Patterns [CP91, Nov95, Nov96]. Payments [Vin99]. Payoff [Dyu68, GE91]. PDE [BPRS94]. Pearson [Aiv59, Bol63, DN75, Kol88]. Peierl [Sob60]. Percolation [MS88, Roy93, Ste89, Ste91, Zue88, Ste91]. Performed [Páz68]. Period [Sei87a]. Periodic [Fel98, Fre64, Kho71a, WR69]. Periodically [Ale94, Gla63, Hur75, HR97].

Periodograms [Ale96, Ale98]. Periods [Sob64, Zub72]. Permutations [IM65b, Kol71, Pav89a]. Perpendicularity [Ale63, Kra74]. Person [Fri69, Vil68, Yan70]. Perturbation [KR90]. Perturbations [Bon92, Che64a, Che64b, Che67b, DJ99, Dub74, Fre67, Gri75, Gri78, Kif75, Kor80, Rez77a, Rez80, Tut74, VTF73]. Perturbed [Bla89, Fre74, Mar79, Sav99, Sun96]. Petersburg [Raz70]. Petrov [Pet66a, Stu65]. Phase [Dob65, Dob66, KS80, KK82, Kar86a, Kar86b, Shp68, Ste70b, Tom90]. Phases [BG91, Dob73]. Phenomena [AS83, Chi61b, Chi73b, Kor94b, Sev59, She87, V’y78a, V’y78b]. Physical [Koz71a]. Physics [DM67]. Pieces [Smi69]. Piecewise [Kal76, Kha61b, Roz99a, Tsi76d, VTGM85]. Piecewise-Linear [Kal76, Kha61b, Tsi76d]. Pitman [Fin71, Kag86b]. Plane [Amb74, Anu80, Che96, GV59, KT59, Mis91, Nik64, Nos79, Nos83a, Roy93, Ste89, Ste91, Tut62, Tut68b, OK97]. Planning [Sok63a, Sok63b]. Plans [Tik78]. Pliska [KK95b]. Point [BPRS94, Ble86, Bor99, BM93b, CL94, Che81, CS71, Dar96, Das79, GM84, Had78, Haf73, IK81, IM65a, KL83, Kol87, KS86, Koz77b, Kva96, LN91, Maĭ91, Nev79, Nik64, Shu73, Ste88b, Zäh83, Zäh84, Zäh86]. Points [BNS66, DY68, Eli85, Gru76, Kry66b, MS88, Men88, Nos79, Nos85, Nos86, Páz68, PS76, Roz73, Sac71b, Sar73, Shu70a, Vil68, Vor58]. Pointwise [BDC68, Mal91]. Poisson [AK95, BL75, Bol62, Bol65a, BM79, Bor89c, CZ79, Dev69, Fel76, FF82, Fel83, Fel85, Fok74, GR71, Gri63, GV93, Haf73, Ign82, Ivc75b, Kab74, Kab78, Khi56a, Kor74, Kor76, Kru92, Lin59b, Lin59c, Log91, Mac71a, Mar65, MS88, Mik75c, Mik77, Mik78, Mik98, Mik58, Mna88a, Nov95, Nov96, Ost65, Ost67, PS85, Pre86, Pre91, Roo99, Roy93, Sev67b, Sev73, Shu85b, SA94, Tim99, Tuc63, Vol65, Zäh80, ZM79b]. Poissonian [Lin57a]. Poland [Ano99d]. Pole [SK92]. Policies [Bul64a, Bul67, Elb76]. Policy [Fai79, Rom65]. Polya [Vak94b, Vol65, Tri90]. Polyá-Type [Tri90]. Polygonal [Aly88a, BB87, Kru86]. Polyhedral [Mol96]. Polyhedrons [DV96]. Polymerization [Whi82]. Polynomial [EZ60, Erm66, GPU98, Ivc76, Ivk86, Kak78, Kol69a, MF71, Mal74b, Med77, Med78, Mes60, Sal98, Zin58, ZL64b]. Polynomials [BG98, Bra87, DM79, IM71a, IM71b, KR76, Mas74b, Mas75, Pro93, Pro94, Rot74, SB67, Seĭ89, Seĭ91, Tep95, Vis92, Yan96]. Ponomarenko [Bak89b]. Population [BS73a, Hag74, IT83, Lab72, Mir84a, Nhu68, Saw93, Shi86]. Populations [AL81, Ivc85, VW99, Vol73]. Posed [Nik66]. Position [Bab64, Shu79b]. Positive [Chi64d, Gir83, Kha59, Lev83, Lev86a, Ros91, Shu85a, Shu86, Sin57]. Positively [SS93]. Possible [Gor81]. Posterior [Rao89]. Posteriori [BD91a, Dar76, DB81, IK73a, Par87, Sal90, VS91]. Potential [KD87, KSS77, Sav99, Shp66, Shu76b, Sin94]. Potentials [Bui82, DRS89]. Powell [BD83]. Power [Bak96a, Bet86, Bor78c, Chi64b, Chi65a, Chi64c, Epa68, Epa74b, Gri94c, IM65a, JP75, Kar80, Mal90b, MSV99, Pag98, Rao89, Bor77c]. Power-Mixtures [Bet86]. Powerful [Bel67a, Kul87]. Powers [Ale92a, Kut85, Shu85a, Shu86]. Precise [Epa74a, Sid91, Ul’79a, Ul’79b]. Precision [KP97, Roz79]. Predictability [IM93]. Predictable [VV88]. Prediction [AA68, Bab85, CFZ95, Gol75, Ibr64, IS68, Lip67, Pit84, Roz84b, Thu85]. Predictions [WS93]. Preliminary [EH93]. Prescribed [KP97, Lep90, MN89, Raf90, Sem90]. Prescribing [Dob70]. Presence [Ara85, Bei65, BL66, Che67b, Dob58d, Mir80b, Vol64, Won59]. Present [Che64a, VS92]. Presentation [Ano73g]. Presented [Ano57b, Ano57a, Ano58d, Ano59e, Ano60c, Ano61d, Ano61e, Ano61g, Ano62c, Ano62d, Ano62e, Ano63c, Ano64g, Ano64f, Ano64d, Ano64e, Ano65c, Ano65d, Ano66g, Ano67f, Ano67g, Ano68g, Ano68f, Ano69g, Ano69h, Ano71c, Ano71f, Ano71g, Ano73k, Ano73j, Ano73i, Ano74h, Ano74j, Ano74i, Ano74g, Ano75j, Ano75k, Ano76e, Ano76d, Ano76f, Ano76g, Ano77g, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano78j, Ano78k, Ano79g, Ano79f, Ano79i, Ano79h, Ano79j, Ano80e, Ano80f, Ano80i, Ano80h, Ano80g, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano81k, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82o, Ano82m, Ano82h, Ano82i, Ano82j, Ano82k, Ano83e, Ano83d, Ano83c, Ano83f, Ano84h, Ano84j, Ano84k, Ano84l, Ano84i]. Presented [Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o, Ano85r, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87o, Ano87l, Ano88k, Ano88l, Leo82, Pet59c]. Preservation [Jan93, Ol′89, Ste84]. Preserve [Eid73]. Preserving [PS88, Shu77c, Zie93]. Press [Ano87f]. Pressure [DM67]. Price [Mov77]. Prices [RR95]. Pricing [KS95, SS95, SKKM95a, SKKM95b]. Primenen [Bak89b, Fry90, Shu91c, SZ98, Ul′91]. Principle [Bak96d, Bar89, Bor78a, Bor84a, BB87, Bor74b, BS81, Bor83b, Bor85b, BM93a, BS87, Dav70, EL99, Glo82, Gor76, Gor83, Gor84b, KS80, Kha63, Kru98, LY79, Lif85a, LS84, Roz84a, RS88, Sīl92, Smo85, Ver99b, Zin86, Kle97, Szy97]. Principles [Shp66]. Prior [Afa87, Rug93]. Priori [Muc88]. Priority [DI74, Ivn93, Ush78]. Prize [Ano73g, Ano98a]. Probabilistic [ARS^{+}97, Bei64, Dar85, Dyn98, Fre88, Ibr89, Kab93, Kar69, Kor73, Kor78b, Mus77, Nag98a, Nag98b, Nos77, SB67, She88b, Vak67, Vla84, Ome97]. Probabilités [JK90]. Probabilities [Ale91b, Ale79, Ale80, Ale84, Amo80, Amo88a, Ark87, Ark88, Ark90, AR86, Ave72, Bas58, Bez82, Bol68, Bol57, BK97, CV65, Dob68, DM80b, Ebr71b, Ebr71a, FV61, Fre74, Gam82, God79, God82, Ivn82, Ivn90, Ivn98, Kal99, Kal69, Khr98, Kis62, Kol88, Kot64, Kry73b, Mal74a, Mal77, MA79, Mos70, Muk90, Muk98, Mus80a, NV75, Nag82, Nyb86, Osi73, Osi74, Osi79, Pet65, Pet68, PS86, Pir68, Pol63, PS89, Rag91, Roz86, Roz90c, Roz94b, Roz98b, Ryc84, San66, Sat96, Sau92, Sev96, Sev99, She87, Shu66a, Shu66b, SMI90a, Sla79, Tol99, Van83, VC71, Ven68, Ven61, Vol60b, Yak98, Yam95, Yur93b, Zal90b, Bol60]. Probability [Aiv65, Ano57b, Ano57a, Ano58d, Ano59a, Ano59e, Ano60b, Ano60c, Ano61a, Ano61d, Ano63b, Ano64c, Ano64f, Ano64e, Ano65a, Ano65c, Ano65e, Ano69h, Ano70f, Ano71c, Ano71g, Ano73h, Ano73k, Ano74a, Ano74j, Ano74i, Ano74g, Ano75a, Ano75i, Ano76g, Ano78j, Ano78k, Ano78f, Ano79j, Ano80j, Ano80k, Ano81l, Ano82a, Ano82g, Ano82p, Ano82n, Ano82o, Ano82m, Ano83e, Ano83f, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87o, Ano87l, Ano88a, Ano88g, Ano88k, Ano88l, Ano89k, Ano90c, Ano90b, Ano90j, Ant80, BD97a, Bau96, Bel64, Bel65, BGS90, Bra88, BMS96, Bul85a, CN87, Che66, Chi71b, Chi76b, Cra56, DD99, Den89, Den92, DT74, Dob62a, EH93, Epa69, FN71, Gam74, GS63, Gne90]. Probability [Gol83, GPU98, Gut64, HMV95, Ibr57, IC98, Kab86, Kal67, Kal82, KY65, Kel73, Kha61b, Kha61c, Kho87, Kho88, Kho94, Khr96b, Khr96a, Klo59, KT69, KS60, Kol62, Kol94, Kor86a, KC83, Koz77a, KU88b, KST61, KY74, Leo82, Lev74, Lev75a, Lev78, Lev86b, LN91, Mak67a, Mak97, MS90, MT90, Mas60, Mil79b, Mir69, MTV91, Muc88, Mus73, Mus85, Nad74a, Nag68a, Nag71b, Nag77, Nik64, NK79, Nov99, Ost87, PS88, PP82, PY88, Pre98, Pro56, Pro59a, Pro65a, Rac85b, RS94, Rén56, Rih85, Rom69, Ron86c, RR64, Ros91, Roz63, Roz64a, Rud93, Rug93, Ruk76, Sal90, Sav99, Saz65, ST66, SE96, Sch79, Sem90, Sev67a, Sev90, SS65a, Shl81, SS90, Sin76, Sko92, SP57, SP61, Ste70a, Stu67, SU90a]. Probability [Tar90a, Teu75, Tik92a, Tik94, Top83, Tut68a, Tut71, UU85a, Ush86, UU87, Ush84, VT78, Vak66, Vak90, Val74b, Val82, Vat74, Vat77a, Vin99, VP64, VN99, V′y98a, Wol76, Yan79, Yan89, Yur83, Zai84, Zhu90, Zie93, Zol57a, Zol61a, Zol84a, Zol84b, Zol92a, Zub87, AP97, GP97, Nad73a, Val73, Ano61e, Ano62c, Ano62d, Ano64g, Ano66g, Ano67g, Ano68g, Ano68f, Ano69g, Ano71f, Ano73j, Ano73i, Ano74h, Ano75j, Ano75k, Ano76e, Ano76d, Ano76f, Ano77g, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano79g, Ano79f, Ano79i, Ano79h, Ano80e, Ano80f, Ano80i, Ano80h, Ano80g, Ano81h, Ano81i, Ano81j, Ano81k, Ano82q, Ano82l, Ano82h, Ano82i, Ano82j, Ano82k, Ano83d, Ano83c, Ano84h, Ano84j]. Probability [Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o, Ano85r, Ano87i, Pet59c]. Probability-By [Zai84]. Probability-Theoretic [Mil79b]. Problem [AA68, Amb65, Amo88b, AVR95, Ark64a, Bei64, Bel69b, Ber86a, Bla64, Bor82b, Bor71, Bor99, BD91a, Bur82, BB91, BP73, Che81, Che64b, Chi67, Chi87a, CV65, Chi87c, Chi89b, CZ79, Dar99, Dob58d, DN88, Dyn68a, Dyn68c, Elb76, FLR97, Fre64, Fre65, Fre66, Fre75, Fri70, GR71, Ger73, Gin66, GS66, Gru88, GZ66, Her91, IC98, IM65b, IM66, Ivc71a, Kap91, KY65, Kha84a, Kha60, KN75, KMM85, Kog65, Kog66, Kol66, Kol67, Kol76, Kor73, Koz74b, Kry80, Kus68, Kut76, LS71, Lei82, Lep91, Lët91, LS65, Lis98, Liv75, Liv80, Luk68, Lut78, LNF89, Maĭ91, Mal93, Mal95, Mal68, MP70, Mat62, Mel89, MN99, Mil96, Mol64, MP91, Mus69, Nag70b, Nag70c, Nag84b, Nag87, NP99, Nag76b]. Problem [Nag85, Nev69, Nik77, Nik99a, Nik66, Nov91, Ol′89, Ost69, Pan95, Par87, Par66, Pav91b, Pin81, Pir67, Pit69a, Pit71a, PS73, PS76, Pro59a, Rac85a, RK79, RS98, Rie80, RS83, Ron88, Ros75, Roz63, Roz67, RU57, Rya66, Sah75, Sar90, Sel99, SC64a, SC64b, Sev67b, She86, Shi65, Sob73b, Sol66, Sta63, SZ98, Tak58, Tsa70, Ush83, Vak67, VJK59, VJ64, Vin68, Wit71, Yas84, Yus76, Yus91b, Zas64, Zig66, ZK77, Zol61a, Zol68a, Zol78c, ZM79b, Kol71, Pre91, RT98]. Problems [AL81, Ale92b, Ano81b, Ano82b, Ano84g, Ano86m, Ano89c, Ano92d, Ano93d, Ano98d, Ano99d, ARS^{+}97, Ark64b, Ark65, Ber85, Bon92, BR64, BK65, Bor67a, Bul64b, Cow93, CS71, Dar85, Die77, Dok92, Fak69, Fre63, Fri72, Fro89, GS88, GS76, Gol92a, INK87, IK99, Ivc85, Kal77a, Kar75, Kar69, Kel77, Ken64, Kho73b, Kle79, Kol94, Kor74, Kor76, KSS77, Kov61, KST61, LS60, Lin57b, Lot80a, Lot80b, Lum71, Maĭ95, Mal74a, Mal77, MS90, Mol67a, MTV91, Nad73b, Nik85, Nik91, Piu94, Pre67, Pro65b, Rév94, Rog58, Rog66a, Rog66b, Roz64b, Roz81a, Roz88, Roz91, SR79, Shi60, Shi63b, Shi64b, Shu81, Tak57, VTF73, VM91, Vol78a, Vol78b, Yag60, Yan89, Zol75, Mil97]. Procedure [Bor79b, BP73, GV93, Kra67, Pav88, Pav91b, Sed77, ZM79a]. Procedures [GK84a, Gor81, KK76, Kor78a, Kor80, Kor84, Nev73a, Nev86, Pen93, Vol79, Zue98]. Process [AHE85, Ale64, Ale70, Ale74b, Ali95a, Ali95b, Anu80, Anu81, Ara68, Bab64, BD97a, Bal94, BD93, Bar91, Beg69, Bel61, Bel66, Bel67b, Bel68, BN69, BP89b, Bor67b, Bor86a, BK86, Bra94, BVT90, Bul61, But92a, Che57b, Chi71b, Chi86b, CDC82, CZ79, DB93, Den84, Dit91, Dob56c, Dob60b, Dyn60, Dyn68b, Dyn70, Dyn71, Dza70, DM75, Dzh71, Elb76, Eli89, Eph76, Fäh78, Fak71, Fak82, Fer98, Fok74, Fre67, GR71, Gap80, Gap89a, GK76, Gin89b, God82, Gol65, GC87, Gor71, Gri63, GK69, Gus84, GZ69, GV93, Hag74, IV93, Ibr63b, IS74, Iva60, Ivk86, IP89, Kal82, Kal99, KN94, KNS66, Kha69b, Kha69a, Kha77, Kha84c, Kha61c, Khr94, KF67, KS87, Kog68, Kog69, Kog70, Kon82, Kor74]. Process [Kor76, Kor89c, KC83, Kru86, KH81, Kut78, Kut79, Kuz75, Kuz78, Kuz80, Kuz81, Kva96, Lab73, Lab79, Leb74, LS71, Leo61, Lev73, Lev78, Lev80, Lif87, Mai79, Mak81, Mal64, Mal65a, Mal69a, Man67, Mih81, Mik75c, Mik58, Mir69, Mir74, Mir75, Mir77, Mir85, Mir76, Mir78c, Mna88a, Mna88b, Mog77b, MP91, Mos91, Mov77, Nag67, NV75, NK81, NK82, Nev64, Nev69, Nev73b, Nos87, Nov71b, Nov91, OT95, PR69a, Pit78, PP82, Por67, Pur89, Rag86, Rie81, Rom65, Ron85, Roz77, Ryz67, Ryz69, Ryz71, Ryz72, Saf77, Sag86, Sag91, Sag96, Sal93a, Sat96, Sem90, Sev67a, Sha80a, Shi64a, Sht65b, Shu62, Shu65, Shu79b, Sir63a, Sir63b, Sko61a, Smi69, Sun96, Sve87, Tak56, Tak57, Tak76, Tak77]. Process [Tik82, Tsi76c, Tsi78a, TF62, Vat77a, Vat77d, Vat77b, Vat77c, Vat78, Vat79, Vat80, Vat82, Vat87a, VS90, VS92, Vat94b, VT73a, Ven65, Ven61, Ver91, Vil89, Vin64, Vin88, Vol86, Vor65, Vos82, V’y78a, Whi82, Yak87, Yus60, Zhu81, Zhu84, Zub72, BO97, KL93, VS89]. Processes [AA68, Aga80, AL77, Ale63, Ale67, Ale81, Ale94, AS83, Ali92, Ali96, Ano89a, Ano90c, AL91, Apo67, Ara61, Ara86, AI99, BR79, BR84, BZ84, BM90, BM91, BY96, Bar77, Bas58, Bas77, BDEM92, Bel58, Bel59, Bel60, Ben76a, Ben76b, BS68a, BJL98, Bev58, Bla62, Bol68, Bon98, Bor81a, Bor83a, Bor80a, Bor65b, Bor67c, Bor67b, Bor68a, Bor74a, BS74, Bor83b, Bor86a, Bor89b, Bor95, BDC68, BP93, Büh68, Bui82, BS89, Che89, Che56, Che60, Chi90, Chi57, Chi59, Chi61b, Chi64d, Chi65b, Chi70b, Chi70c, Chi70d, Chi72, Chi73b, Chi87c, CS60, Cle98, Cra64, Cra65, Dav68, Dav70, DI71, Dav74a, Dav76c, Dav79, DR79, Dav81, Dav82b, Dav89a, Dob57, DK93, Dok87, Dok92, Dok95]. Processes [Dok99, DN88, DSS94, Dyn56a, Dyn56b, DY56, Dyn58, Dyn59, Dyn67, DY68, Dzh77, Eag69, EL99, Eng71, Eng74, Eng75, EW99, Ers70, Ers72a, Ers72b, ES69a, ES69b, ES75, ES77, Fai76, Fai83, Fak70, Fal88, Fal89, Fok77, Fra63, FS79, Fre63, Fre73, Gal85a, Gal94, Gap74, Gap78, Gap82, GK75, Gär76, Geo80, Ger96, Gir59, Gir60b, Gir60a, Gla58, Gla61, Gla63, Glo79a, Glo82, Gor84a, Gor85, Gor78a, Gor80, GN87, Gri63, GS66, GM84, Gri78, Gri83, Gri87, Gri91b, Gri92b, Gri92c, Gru86, GK68, Gus69, Had78, Haf73, Has66, HSVJ67, Hur75, HR97, Ibr62, Ibr74, Isa85, Iva70, IM93, Ivk68, IR71, Jag68, Jan76, Kab74, Kab75, KLS84, Kal76, KM65, KS69, Kal62, Kal69, KS80, Kar85]. Processes [Kav91, KW71, Kha68, Kha81, Kha93, Kha58, Kha60, Kha63, Kha61b, Kif73, Kif77, KFS77, Kli62, Kli77, Kli79a, KR60, KP83, KP84, KS86, Kor99, KSS77, KS85a, Koz77a, Kra62, Kra69, Kra71, KR90, Kry66a, Kry66b, Kry73b, Kry80, Kry87a, Kry87b, Kul85a, KST60, Kuz74, Lab72, LN68, Lam72, LS97, LSS99, Leb71, Leb87, Lev72, Lev73, Lev74, Lev75a, Lev88, Lif82, Lif83, Lif85b, Lip66, Lip67, LS68a, LS68b, Lip97, Lod71, Lya87a, Lya87b, LNF89, Mak82b, Mak95, Mak70b, Mal74a, Mal77, Mal80, Mal85a, Mal87a, Mal92b, MO83, Man64b, Man64a, Mar70, MW59, MP58, Mat60, Mat61, Mat62, Mel83, Met75, Mik75a, Mil72, Mir78a, Mir78b, Mir80a, Mir81, Mis89, Mol64, Mol67b, Mol68, MO69, Mol96, MS86, Nad73b]. Processes [Nag75, Nas96a, Nas96b, Nas97, Naw82, Nik80b, NV85, Nor91, Nor93b, Ouk91, PN92, PV64, Par71, Pav97, Pec77, Pin69, Pir67, Pis80, Pis82, Pis65, Pit82, Pit83, Pit87a, Plu92, Pol78, Por75a, Por79, PS92a, Pra99, Pre82, Pro56, Qua99, Rak87, Rak88, Rak89, Rak90, Rak93, Ral97, Rao63a, Rao76, Rao89, RK79, Reb88, Rez77b, Rez68, Rez71, Rob63, Rog65, Rog66a, Rog68a, Rog72, Rog92, Rom64, Roz59a, Roz60b, Roz61, Roz63, Roz64a, Roz64b, Roz80, Roz84b, Ryk66, Ryz65, RS70, Saf92, Sag83, Sar73, SG62, Sel69, Sen68, Ser61, Sev57a, Sev57b, Sev61, Sev64, Sev66a, Sev67b, Sev68, SC71, SZ74, Sev90, Sev96, Sev99, She87, Shi63a, Shi67, Sht65a, Sht66, Shu64, Shu66a, Shu66b, Shu67]. Processes [Shu68, Shu70b, Shu76b, Shu77a, Shu78b, Shu78a, Shu80a, Shu77d, Shu77c, Shu79a, Shu80b, Sīl92, Sil74, Sin60, Sin62, Sin63, Sir73, Sir77b, Sir77a, Sir78, Sko56, Sko57b, Sko57a, Sko58, Sko60, Sko61b, Sko62, Sko63b, Sko64a, Sko64b, Sko66, Sko67, Sko80a, Smo85, Sol94, Sol95, Son67, Str60, Su99, Sur63, Sza71, Tan70, Tom90, Top83, Top89, TV98, Tor88, Ugl89, Urb56, Vat74, Vat81, Vat87b, VMS95, VT82, Ven77a, Ven77b, Ven80, Ven59, Ven83, Ven86, Ven90, Ver66, Ver64a, VT92, Ver64b, Vil59, VM89, VGR84, Vol58a, Vol59, Vol61a, VK86, V’y77, V’y78b, Yag68, Yag57, Yag60, Yak81, Yak86, Yam95, Yan76, YM84, Yus57, Yus90, Yus91a, Zäh83, Zäh84, Zäh86, Zak90a, Zak90b, Zak91, Zas64, Zhu70]. Processes [Zhu72, Zin88, Zol57b, Zol58, Zol64a, Zol65a, Zol66b, Zol76, Zub73, Zub75, Zub83, de 93, Bar97, DRS89, GP97, Met74, PvZ97, Sev59, Svi90, Vat95, Ver94, Sko57b]. Processing [BN96]. Processor [Gri92b]. Produced [Koz71a]. Product [Bik69a, But79, But92b, LS97, LSS99, Mol67c, Tut65, Vir84, YK87]. Production [Sev62]. Products [GP79, Gir76, Gir83, Haz95, Kal77b, Mak67b, MT85, Ose65, Ost83, Sko80b, Tut68a, Tut78, Vir80]. profits [KY97]. Programming [Fai82, Fai87, Faj88, Fri70, Pir68, Pit71a]. projecting [Mak88]. Projection [Eph76, Nad77, WS93]. Projections [Bor91c, UU85b]. Projective [PT93, RS94]. Prokhorov [Abr77, Bak96d, Bor84a, BB87, Fer98, Jur75a, Mar87a, Sen85, Zai87a, Zai87b, Zai88a].

Prokhorov-Donsker [Bak96d]. Proof [CPU93, KK95b, Kry91, KL79, Lin59a, Sha61, SS93, Vak94b]. Propagation [Bor98, Tan70]. Properties [AM67, AM69, Aiv59, Ale67, Ale91a, Ale94, Ale96, Arn81, Asa88, Bel60, BM63, BS73b, Ber74, Bez75, Bla89, Bor70, Bor72, Bra86, CC96, Dob60b, Döh80, DK97, Dzh75, Gab75, Gik56, Gir60b, Ibr74, IH76b, Ing88, IT89, Jes80, Kal76, Kha68, Kha90a, Kha60, Kol99, Kol69b, Kor86c, Kos85, KM89, Kry85, Lep90, Lev74, Lev75a, Lev78, LTC80, Lis98, Mal65a, Mkr81, NS89, Nev73a, Nev86, OT95, PR87, Per92, Plu92, Riv96, Rob63, Roz60b, Sap92, SB67, Ser96, Sht65a, Shu58, Shu85a, Shu86, Sko87, SS92, Sli62, Sli64, Ste84, Stu68, TT82, Tsi76b, VM97, Ver64b, Vin95, Vla79, Yak86, Yak87, Zäh83, Zäh84, Zäh86, Zai81, Zie92, Zin58, Pus97]. Property [Bak96e, Bra80, Bul87, CS61, DL86, ES77, Geo80, Gir59, Gor78b, KS67, Kag86a, Kha81, Kle74c, KN90, Lab73, Lab79, Lis94, Mey63, Mol68, Nag78, Ol′89, Pla92, RR91, Sar70, Shu77c, Son92, Tak68, ZL64a, Zol78d]. Proportional [Vol93]. Proposed [ZK61]. Prospective [Plu81]. Proving [Bor89b, Ego85, Gam81, Gam89, Nak88]. Proximity [Mir86b]. Pseudo [Ebr71a, Sob58, Sob64]. Pseudo-Moments [Ebr71a]. Pseudo-Random [Sob64, Sob58]. Pseudobest [Kho71b, Zer70]. Pseudomoments [BK89, Nag98a, Zol79]. Publication [Yus87]. Pure [BG91]. Purely [KM65, Pro68a, Tan70]. Pushkin [Bob64]. Pyke [Pen77].

Quadratic [Bag88, Bag89a, Bak90, Bak96b, Bak96c, BGZ98, Cad99, CPU96, DW79, Fäh83, GR78, Gol92a, JM94, Kha85, Kle74a, Kle80, Man68, Mar76a, Mar78, Mik89, Mik91, Mik96b, Mil82, Nad77, Nad93, Nyb86, Pon79a, Pon79b, Pon86, Pre98, Rot76c, RS86, Sap79, Ten91, Tik78, Tik82, Whi60, Whi64, ZI89, Bak89b]. Quadratures [Gya64a]. Qualitative [Zol75]. Quantile [Nik75, Nor93b, Ral97, Sun96, Tik88]. Quantiles [Mol91, Pag65]. Quantitative [Men88, Sei82, Zol78d]. Quantities [Hat75]. Quantization [IMS93b, IMS93a]. Quantized [Gap89a]. Quantum [Bel94, Bel95, Das79, Den89, Kho73b, Kho79, Kho87, Kho88, Yag56]. Quasi [Bor79a, Fel65, Gih91, Gol92b, Khr91, Kry66a, KST60, Ser80, Zin58]. Quasi-Gaussian [Khr91]. Quasi-Linear [Gih91, Gol92b]. Quasi-Martingales [Bor79a]. Quasi-Measures [Fel65]. Quasi-Moment [KST60]. Quasi-Polynomial [Zin58]. Quasi-Stationary [Ser80]. Quasilinear [Bla64, Gik90]. Quasimoment [KST61]. Quasipolynomial [Rot76b]. Quaternion [Vak99]. Question [Bol66, Lin56a, Soz64]. Questions [Mus73, PY69, Sal71]. Queue [Bas65, DM68, DI74, Gri95, KLP89, Pav91a, Vin68]. Queueing [Afa65, AM69, Afa88, AL77, BN96, Bor87, Bor88b, Bor98, Fal88, Fal92, Gri91b, Gri92c, Ivn82, Ivn90, Ivn93, Ivn98, Kal77a, Koz78, Leo78, Mal91, MT83, Pav92, Pre65, Rie80, Sei82, Sei87a, Tsa83, Tsi91, Ush78, Vin95, Yus88, Zol77b, Zol78d]. Queues [AM67, Ken64, KS86, VJ64, Vis64]. Queuing [Akh80, BP73, DI74, KK82, Kov61, Zol75]. Quickest [Shi63b].

R [Ost69]. R. [Fal78, Rom59, Soz64, Val74a]. Rademacher [Chu92a]. Radial [Tri90]. Radially [Kud76]. Radiation [She86]. Radio [Rez77b, Tut69a]. Radon [DS85, Vla80]. Radonifying [Lus96]. Random [Afa79, Afa87, Afa91a, Afa91b, AB85, AB99, Alb93, Ale92a, Ale70, Ale74a, Ale74b, Ale81, Ale89, Ale79, Ale80, Amb74, Amb76, Amo80, AT86, Ant80, Ara72, AN74, Ara75, Ara61, Ara90, AVR95, Are78, Ark90, Arn81, Asa88, Ave72, AV82, BR84, Bag88, Bag89a, Bag89b, Bak98, Bal67, Bal68, Bal74, BNS66, Bar86, Bas59, BGW75, Bel59, BR85, BGZ98, Ber70a, Ber70b, Ber74, Bet88, Bet89, Big93, Bik69b, Bol59, Bol64b, Bor81a, Bor83a, BB87, Bor78b, Bor80a, Bor82a, Bor84b, Bor85a, Bor60b, Bor60a, Bor62, BR64, Bor67a, Bor68b, Bor72, BS74, Bor96, BM99, Bor85d, Bor86b, Bra87, Bra91, Bra90, BD91b, Bul74b, Bul74a, Bul77, BP78, Bul81a, BS88, BS92, Bul75, BZ77, Bul84, BM92]. Random [Bul96, Bur82, Bur75a, Bur75b, But92a, But95, Che64a, Che64b, Che67b, CC96, Chi57, Chi64d, Chi87c, CPU96, Chu92b, Csö84, Dar76, DB81, Dar85, Dar86, DB88, Das79, Dav76c, Dav79, DR79, Dav81, Dav82b, Dav89a, De 82, Del96, Den84, Deo74, Deo77, DW79, Dit91, Dob68, Dob70, DM79, DK84a, DK84b, Döh81, Dor87, Dor90b, Dub69, Dub74, Dyn65, Dzh77, EO97, EM68, Ego70, Ego91c, Elb83, Eng84, Eng85, Eph76, Eri76, EZ84, Ero58, EK72, Evs78, EO79, Evs87, Evs89, ES77, Fäh83, Fal78, FTV92, FLR97, For62, For63, FN71, GR84, GS88, Gal83, Gal85b, Gam64, GP71, Gam74, Gam86, Gap78, Gap89b, Gap98, Gap73, Gär76, GV59, Gia96, Gik58, Gir76, Gir77, Gir80, Gir82, Gir83]. Random [GV86, Gir91, Gir92, Gir96, Gir98, Gla58, Gla63, God79, Gol85a, Gor71, Gor83, GE93, Gri63, Gri75, Gri83, Gri91b, Gru73, Gru76, Gur92, Gus76a, Har91, Har77, Haz95, Hög93, Ibr63a, IR70, IM71a, IM71b, IP74, Ibr75a, Ibr84, Ios68, Ios70, Ism73, Iva70, Ivc73, Ivc74a, IL78, Ivk68, IR71, JM94, JP97, JR89, Kab96, Kag89, Kag90, Kal93a, KS64, Kap91, KF91, Kar75, Kar95, Kar65, KD87, Kat78, Kel73, Kha65, Kha61a, Kha66b, Kha90b, Kha61b, Kha61c, Kho83, Kif75, KMM85, Kle87, Kle96, KT69, Koc98, Kog68, Kol71, Kol76, Kol87, Kol56b, Kop96, KNS94, Kor86a, Kor86b, Kor89b, Kor89a, Kor90a, Kor93, Kor98, Kor61b, Kor62, Kor90b, KC91, Kor66, Kor68, Kov67, Kov73, Kov86, Koz71a]. Random [Koz71b, Koz73, Koz74a, Koz77a, Koz77b, Koz81, Kra71, KR90, Kru71b, Kru73a, Kru74, Kru75, Kru80, Kru86, Kru89, Kru92, KB97, Kru99, Kry68, KY64, KS89, Kud85, Kul98, Kur85, Kur88, KST60, Kuz74, Kuz85, Lab72, Lar94, LS97, LSS99, Leb84, Leb86, Leb87, Leb90, LL94, Leb96a, Leb96b, LS71, LY79, Lët89, Lët93, Lev86b, Lif83, Lis76, Log91, Lot92, Lum71, Lya87a, Lya87b, LNF89, LR63, Mai75a, Mak82a, Mak67b, Mak70b, Mak80, Mal94, Mal97, Mal65b, Mar93, Mar81, Mar85, Mar86b, MP93a, ML66, Mar76a, Mas74b, Mas75, Mas99, Mat65a, Mat65b, McD80, Mih56, Mik65, Mik74, Mik75c, Mik75a, Mik94, Mik96a, Mik97, Mik99, Mik70b, Mik96b, MR89, Mis91, Mog77a, MP78b, MR83b, MT85, Mol88]. Random [MTV91, Mol91, Mol92, Mol94, Mol96, Mór93, MC68, MPY70, MFY74, Mov77, Muk85, Muk90, Muk92, Muk98, MS86, Mus69, Mus70, Nag70a, Nag70d, Nag77, NE77, Nag78, NP78, Nag87, Nag94, Nag98a, Nag76b, Nak88, Nas57, Nas96b, Naw82, Neu65, NP69, Nev71, Nev72, Nev73b, Nor93a, Nos79, Nos80, Nos83a, Nos83b, Nos85, Nos86, Nos88, Nov92, Nov95, Nov96, Nov97a, Nov79, Nov82, Nov83, Osi69, Osi71, Osi73, Osi74, Osi79, Ost91, Ost98, Pak59, Pan87, Par66, Pav78, Pav82, Pav84, Pav89b, Pav89a, PR99, Pec74a, PR69a, Pec74b, Pec81, PS93, Pet56b, Pet59a, Pet59b, Pet64, Pet65, Pet70, Pet76, PL83, Pin79, PU85b, PU90, Pin91a, Pir67, Pit73, Pit83, Pit84, Pit92, Pit75, Piu94, PR69b, Pon79a, Pon79b, Pon86]. Random [PY69, PY81, Pre67, PS73, Pro73, Pro56, Pro84, Pro91a, Pro93, Pro94, Pro74, PRR87, Rad92, Rad98, Rag91, Rak88, Rak90, Rao63b, Rao71, RS98, Rez77a, Rez80, Rez81, Ric65, Riv96, Rog61b, Rog66c, Rog71, RF78, Rom61, Ron80, Ron88, Rot76c, Roz59a, Roz60a, Roz77, Roz81a, Roz88, Roz90a, Roz91, Roz94a, Roz90c, Roz98b, Rus96, Rya79, RS76b, RS80, Sac71a, Sac73b, Sac74, Saf92, Sak85, Sal90, Sal71, Sar67, Sau92, Sav99, Saz75, Sch89, SS62, Sei87b, Seĭ89, Seĭ91, Sel69, Sem88b, Sev72, Sev73, Sev94, Sgi99, Sha96, She80, She99, Shk74, Shl89, Shn87, Shu76a, Shu89, Shu90, Shu91a, Shu91b, Shu92, Shu96, Shu77c, Shu81, Shu82b, Sid91, Sil73, Sil74, Sil79, Sin57, Sin83, Sin94, Sir77b, Sir78]. Random [Sko60, SS65b, SS66, Sko74b, Sko77, Sko80b, Sla78, Sla82, Sla85, Sli62, Sli64, Smo85, Sob73a, Sob64, Sre70, Sta65, Ste84, Ste69a, Ste69c, Ste69b, Ste70a, Ste70b, Ste71, Ste73, Ste88b, Ste89, Ste91, SO85, Su99, SU90b, Sun96, Sun99, Syt67, Sza71, Sza73, Szu83, Tem73a, Tep95, Th93, Tie77, Tik81, Tik92a, Tik94, Tim85, Tim89, Tom71, TV98, TP57, Tsu81, Tut61, Tut65, Tut68a, Tut68b, Tut69a, Tut73, Tut74, Tut78, TF91, Ugl89, Ul’87, Ush86, Ute91, VK97, Vak99, VTT81, Vas80, Vat96, VTF73, VT73b, Via91, VM91, Vin76, VG90, Vin94, Vin90, Vir80, Vir84, Vol58a, VR59, Vol61a, VR61, Vol58b, Vol60b, Vol96a, Vor64, Vov88, Vve84, Vve85b, Vve89, V′y98a, V′y98b, WR69]. Random [Won59, Woy74, Yag57, Yak98, Yak87, Yan76, YM84, Yan96, YK87, Yur74, Zäh83, Zai84, Zai87a, Zai88a, Zai88b, ZI70, ZI71, ZI74, Zas73, Zhd90, ZM95, Zhu66, Zhu68, Zhu70, Zin79, Zin86, Zin88, Zin65, Zin71, Zol65b, Zol77a, Zol78a, Zol78c, Zol92a, ZM74, Zup92, Bak89b, Bau97, Dav76b, Ego97, Gap95, Ivk94, Kor95, Kor96, Min91, NK97, OK97, PR97, Rad97, Shu91c, Sob58, Vat95]. Random-Number [Mik96a]. Random-Sum [Eng84, Eng85, RS80]. Randomization [Bul85b, Mik75a, SL93, Shi97]. Randomized [Döh80, Fai82, Fai87, GS76, IM89, Mir86a, Mir88, Rom87, SMI90a]. Randomly [Fre74, Kor93, Mor75, Tei66]. Randomness [BB61, KU88a, McK93, She93b, V’y88, van94]. Range [Fil84, PRR82]. Rank [Bal68, Eph76, HP94, KR75, Kov73, Kul87, PS69, PR85, PW87]. Ranks [PK87]. Rao [Bol61, Kho74, Ost69, Sha61, She88a]. Rao-Cramér [She88a]. Rapid [SU90b]. Rare [Mna86, Sol66, VK86, Zäh80, Zäh84, Zäh86]. Rarefying [IP89]. Rate [Akh80, Akh81, Ali87, Ali90, Ara82a, Ara82b, Ara83, AR86, Bar88, Bor77a, Bor78a, Bor84a, Bor82b, Bor74b, Bor78d, Bor80b, BS81, Bor81b, Bor83b, Bor85b, BV88, Bul96, Dau88, Ego73, EN75, Ego80, Ego91b, Fom83, GR78, Gap82, Gil82, Gor76, Gor84b, Gri89, Gri96b, Ibr70, IP74, IM87, Kal78, Kap74, Kar80, Kha85, KR92, KP83, KP84, KB89, LS82, Log91, MU83, Mal85b, Mat85, Mat84, Mat86a, Mat86b, Mik78, Mos91, Nag68a, Nag71b, Nag86b, OR89, Orl75, Pau77, Pau79, Pop80, PS89, RR93, RR99, Rot78a, Rot78b, Roz81b, Roz90b, RS80, Sat73a, Sat73b, Saz67, Saz69, Sch79, Sen81, Sen84, Sen86b, Sen87, Sen89, Sen90, Sen92, Sen93, She80, She71, Sho83, SF84, Sla87, SU90a, Sun99, Tik81]. Rate [Ul’79a, Ul’79b, VL76, VJ64, Ver88a, Vin74, Vin76, Vro99, ZA84, Zai85, Zal86, Zas73, Zup92, Min91]. Ratein [KA99]. Rates [Bha70, Dob62b, Mal94, RS76b, SS65a]. Ratio [Aiv65, Bak96c, CvZ85, CvZ89, DM75, IK73b, KK75, Lot86, Lot88, Mar93, Nov97a, Nyb86, PR99, Pis65, Rom69, Shu89, Shu90, Tar85, VN99]. Rational [Ara86, Dza70, DM75, Dzh71, KS95, Pis65, Pit87b, Yag60]. Ratios [Shu91b, Shu91c]. Rayleigh [Tru66]. Rays [Pav81b]. Reaching [Afa87, Fil84, Tak76]. Real [IM71a, IM71b, Kor94b, Mas74b, Mas75, Rad98]. Real-Valued [Kor94b]. Realization [Mus69]. Realizations [Cow93]. Receding [KS84, KS85a]. Reciprocal [Mir80a]. Recognition [Bor71]. Reconstruction [Ivn77, Kor92, ZK77]. Record [Yak87, Yak96]. Records [Nev88b, Ste93, Ste94]. Recovery [Wit71]. Recurrence [BDC68, Kor86a, Nev69]. Recurrency [RF78]. Recurrent [Bor82a, Bor84b, Bor85a, Gel64, KS80, Kha60, Kon70, Mal87a, Mal92b, Nov79, Sko87, Tak56, Zol76]. Recursive [GK84a, Kor80, Kor89c, Nev81, Sch93a]. Reduced [Sag86, Vat94b, Yak81, Yak86]. Reducing [Yus80]. Reduction [Tik92b, Yus76]. Redundancy [Kis62]. Redundant [Mak91]. Refinement [Bak96d, Ber86b, Bol61, Bor89c, Chu95, Fra63, Gir98, HSVJ67, Kav91, MA83, Mik94, Osi66, SI67, Tho91, Ven90, Vol96a]. Refinements [Nag98b, Whi69, Zol64b]. Refining [Sal79]. Reflecting [DV96]. Reflection [Anu80, AL91, BP89b, Fre63, Kog69, Kog70, Nas96b, Roz73, ZI74]. Reflections [Zol92b]. Reflexive [Met75, Met74]. Refusal [Gor81]. Refusals [Akh81, BB81, Bor73a, Sev57a]. Regenerating [Kli77]. Regeneration [Sei87a, Tsi91]. Regenerative [FS79, VK86]. Region [Bel68, Gär75, Gut64, KM80, Nik64, Sko61b, Sko62]. Regions [Bog80, KS85a]. Regression [Ale92b, Bog67, Bol94, Chi87c, CDM93, Dar96, DP82, Dok93, DL86, Fel99, Gad90, GN93, Gur89a, Gur89b, Hau93, IK82, Iva77, Kho69, Kho71c, Kle74b, Kle80, Kon78, Kor91, Kor94a, Kus68, Mal69b, MF71, Mir79a, Muk78, Nad64a, Nad65, Nad70, Nev79, Nys87, Pag83, Pin91b, Plu92, PT91, PT93, Ras77, RK79, Sed77, SL93, Sok63a, Sok63b, Ter93, TP59, Tsu81, Tsy88, Vol93]. Regular [Bui82, DE77, Dza74, Evs87, Gri78, Gus93, Kry66b, Kul98, Kuz75, Kuz78, Mar86b, Mat61, Pre67, Saz90, Sen74, Shu70a, Tak77, ZSU89c]. Regularity [Che89, Dob68, Gri83, Gri87, Gur68, Kry73b, Lax63, Leb84, Pla96, Sta84, Vat87b]. Regularly [Tik98]. Regulation [Mov77]. Regulator [Pre98]. Rejection [Mir79a, Mir79b]. Related [Bol64a, Bor72, BU84, Bur82, CS71, FM92, IK79, Kag90, KW71, Pec74a, PS76, Rog66a, Saf92, Sal71, Ser80, VM97, Vol67, Yag57, ZL64b].

Relating [Nad73b]. Relation

[Gel74, Kar65, Mak65, Pet66b]. Relations [Shl89, Sid57]. Relationship [Gap89a, Muk96]. Relative [Leb81, Ush78, VC71]. Relatively [Rog77b]. Relativized [Muc88]. Relaxed [Leb78]. Reliability [Abd99, Ano90c, BGS90, Kar69, Kle79, Mak91, MP70, SK93, Szt93]. Reliability-Theoretical [BGS90]. Reliable [Fre76]. Remainder [Eng84, Eng85, Gel64, IO66, LM71, OP67, Pet66b, Rog74, Roz78, Roz79, Sal79, Tom71, Zol66a]. Remains [Mir69]. Remark [BL63, Dav89b, Gap86, Ibr57, Ibr75b, Khm81, Kle71, Lin56b, LRS65, LE68, Mus81, Nor91, Rog60, Rog69, Rom69, Saz58, Sko74a, Sko75, Stu65, Val74a, VN78, Pet66a]. Remarks [BBM59, Bag89c, Bor64a, BS74, CDC82, Kor57, Lam72, Läu74, Lin57b, Mar65, MR83a, Mog77c, Mus85, Nad70, PS86, Plu81, Pro59b, Rot76a, Roz66b, Sar67, Sen89, Sen92, Shu70b, Sil73, Sil79, SZ98, Tie77, Vis92, WR69, Zol68b, JK90]. Removal [Cai93]. Removing [Kos88]. Renewable [Mak91, Mak96a, SK93, Szt93]. Renewal [Bor83b, Dar99, Fra63, Kal78, Kar80, Kar81, KF67, KS88, Mal87a, Mal92b, Nag68b, Nag69d, Nag76a, Nag80, Nag84a, Rog74, Rog77a, SC71, Sgi92, Sgi98, Shu76c, Shu85c, Shu87, Sve87, VK89, Zas73]. Renewals [BM63]. Renyi [Bor91b, VK89, Bon98, Nov98, Ush83, Kon74b]. Repair [SK93, Szt93]. Repeated [CP91, Gri91b]. Repetition [Mik96a]. Repetitions [Iva65, Mik76, ZM80]. Replacement [BB61, IL78, Mir86b]. Report [Ano60b, Ano63b]. Reports [Ano61f, Ano61g, Ano62e, Ano65d, Ano66g, Ano67g, Ano68g, Ano68f, Ano69h, Ano71c, Ano74g, Ano75i, Ano75j, Ano75k, Ano76f, Ano76g, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano78j, Ano78k, Ano79i, Ano79h, Ano79j, Ano80i, Ano80h, Ano80g, Ano80j, Ano81l, Ano81k, Ano82p, Ano82l, Ano82n, Ano82o, Ano82m, Ano82h, Ano82i, Ano82j, Ano82k, Ano83e, Ano83d, Ano83c, Ano83f, Ano84h, Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o, Ano85r, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87o, Ano87l, Ano88k, Ano88l, Leo82, Pro83]. Representability [MP92]. Representable [Ost83]. Representation [Bui82, CH93, Gal77, Glo79a, IR71, Jak98, Kab74, Kab93, KM65, Kha84c, Kho88, Kor73, Kor78b, Lif80, Lip77, PS64, Pav81b, Sko77, Smi69, Sol95, Tut73, Zan99, Zol95]. Representations [Ers72b, IM65a, Kab75, Sir73, Sir78, de 93]. Representing [AT86]. Reproduction [Mik75a, Rak93]. Required [Vol69]. Research [Ano57b, Ano57a, Ano61d, BGS90, MS90, MT90, Pet59c, Sev90, Zhu90]. Residual [Rus96, TP59]. Resolvent [BK86]. Resolvents [Gir96]. Respect [Bar91, But77, Dzh94, Gal79, Gal83, Gal85b, Kit82, Kle74b, Leb84, Leb90, Leb96b, Met75, Nas88, Rad92, Roz62, V’y88, Met74]. Restricted [Her91, Ros95]. Restriction [Ula88]. Restrictions [IK93, Kov61, Sem90]. Results [AB85, BR84, BK65, Büh68, Bul64a, CGM97, DF93, Fak82, Fal78, IK79, Kif75, Kle81, Lut90, MO91, Mir79b, Mir84c, Nor91, Pav98, PS61, Ser80, WW92]. Résumé [Ano71c]. Retarded [Zam70]. Retrial [Fal92]. Retrospective [Dar96]. Return [Vit70]. Reversed [Ano59c, CFZ95]. Reversion [BO96a]. Review [Ano67d, Ano73e, Ano75g, Ano75h, Ano76c, Ano77e, Ano77f, Ano78e, Ano81f, Ano84f]. Reviews [Ano58a, Ano58b, Ano58c, Ano61c, Ano62b, Ano64b, Ano65b, Ano66a, Ano66e, Ano66f, Ano67e, Ano68d, Ano68e, Ano69d, Ano69e, Ano70e, Ano71d, Ano71e, Ano72b, Ano73f, Ano74f, Ano80d, Ano81g, Ano82f, Ano86i, Ano87h, Ano88i, Ano88j, Ano89h, Ano89i, Ano89j, Ano90i, Ano91g, Ano91h, Ano91i, Ano95d, Ano96c, Ano96d, Ano97e, Ano97f, Ano98b, Ano98c, Ano99c, Gen59]. Reward [Fai80, Yus82]. Rewards [Fäh83]. Rice [Mir77, Mir84b, Tru66]. Riemannian [Dyn98]. Riesz [Gap98, TM79]. Right [Bor78b, Kha66b]. Right-Hand [Bor78b, Kha66b]. Rigorous [Men88]. Ring [But77, Lev86b]. Risk [Kha84b, Kle74a, Par70, Pin91c, VN94]. Risks [Gol92a, IK82, Kha79a, Mok91, Kha79b]. Robbins [Kra67, KS89]. Robbins-Monro [Kra67]. Robust [FN93, KK75, Lot96, PC96, PP93, Rug93]. Roland [GIM^{+}97]. Role [AB60, Hid83]. Roof [Dyu68]. Root [Ron86c, Vat94a]. Roots [Kra67, Mas74b, Mas75, Sgi99]. Rosenthal [IS98]. Rotation [Par64]. Rough [Ven77a, Ven77b, Ven80, Ven83]. Round [LN91]. Round-Off [LN91]. Rubinshtein [Kai86]. Ruin [Kor76, Vin99]. Rule [Fri69, GZ66, VM89]. Rules [DSS94, Eng74, Eng75, Fak70, GS66, Zez75]. Run [Bol94, Sam82, V′y98b]. Running [Der93]. Russian [KS95, SS95, BV95, Gas63, GP98].

S [Bak89b]. S. [BK71, Gin66, KY86, KS60, Nik66, Pro68c, Pro68d, Pro68b, Yur70]. Saddlepoint [Shu91d]. Sagdy [Ano89g]. Same [KSS77]. Sample [AB93, AB60, BNS66, Bar85, Bel58, Bel60, BS89, Chi61a, Dob58a, Dob60b, Epa68, Epa74b, Ibr74, Iva60, Kha89b, Kle71, Kle74a, Kle80, KY84, Man68, Mar69, Mir84a, Nag71a, Nev88a, OCG93, PU85a, PS69, Ruk67, Ryz67, Sem88a, Ser96, Sol94, Sun96, Tak76, Tyu74, Urb56, Vol74, Vol78a, Vol78b, Vol79, Vol80, Vol81, YK86, Yur74, Sac71a]. Samples [Art89, Che80, Gik57, IM79, Kha66a, KP71, Lin64, Pag68, Pet56a, Tik85, Tik98, Vol65, Zin76]. Sampling [BB61, BK80, Hag74, IL78, Mir86b, Ser79, Shi86]. sans [Lan90]. Sapogov [Mal68]. Satisfying [For63, Ibr65b, Ibr70]. Sazonov [Ano87p, Ul′91]. Scalar [GP79]. Scale [Art89, KR67, KR90, Kul85b, Nik74]. Scaling [Ste89]. Scanning [Zig66]. Scattering [Mon56]. Scheme [Abr87a, Abr87b, Abr88, Aiv65, Bel69a, Bol83, Bol90, Bol94, Chi87a, Chi88, Die77, Gri96a, IT79, Ivc74b, Ivc75c, IL76b, Ivc76, IL78, IM89, IT89, Kap74, KMM85, Kol69a, KC75a, Kor59, Mas74a, Med77, Med78, Mik75c, Mik77, Mik98, Mir86a, Mir88, Mir84c, Ron84, Sev66b, Sev73, SMI90a, Tor88, VM83, VMS95]. Schemes [BP78, Kel77, Liv75, Liv80, Ron89, Lep93]. School [Ano88g, Ano89a, Ano89k, Ano90j]. School-Colloquium [Ano88g, Ano89k, Ano90j]. School-Seminar [Ano89a]. Schrödinger [Lev86a, RT98, Sav99]. Schwartz [Mus69]. Science [Ano69h, Ano81j]. Sciences [Ano61g, Ano62e, Ano63c, Ano64d, Ano65d, Ano67f, Ano68f, Ano71c, Ano71g, Ano73h, Ano73k, Ano74j, Ano76g, Ano77h, Ano78i, Ano78h, Ano78k, Ano79f, Ano79i, Ano79j, Ano80f, Ano80i, Ano80j, Ano81h, Ano81i, Ano81l, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82m, Ano82h, Ano82j, Ano82k, Ano83c, Ano83f, Ano84h, Ano84j, Ano84l, Ano84i, Ano85q, Ano85p, Ano85k, Ano85n, Ano85o, Ano85r, Ano86k, Ano87k, Ano87n, Ano87j, Ano87m, Ano87l, Ano88l, Leo82, Ano75a, Ano75k, Ano76f]. Scientific [Ano73g, Ano81e, Ano84e, Ano85h, Ano85i, Ano86f, Ano86e, Ano86g, Ano86h, Ano87g, Ano88f, Ano88g, Ano90g, Ano91e, Ano91f, Ano92b, Ano92c, Ano93c, Ano94c, Ano94d, Ano94e, Ano95c, Ano96b, Ano97c, Ano99b, Pro83, Ano57b, Ano57a]. Score [PR85]. Score-Generating [PR85]. Screening [Fre75]. Search [Ark64a, Ark64b, Ark65, EZ84, Gut64, Gya64b, Kar75, LS65, Tsa70]. Searching [Sta63]. Second [Ale98, Ano60a, Ano78f, AS74, BB87, Che56, Dub74, DL86, Ego73, ES69a, ES69b, Gap74, Gap78, Kog65, Kog66, Lev81, Lev86a, Mal86, MSV99, Mil79a, Mir80a, Rot74, Sch65, Sir78, Sko67, TTC78, Vak94b, Vat79, VT76, Ver64a, ZA84, Zai85, Zaĭ90, Mil82]. Second- [Ale98]. Second-Order [BB87, Dub74, DL86, Gap78, Lev86a, Mal86, MSV99, Mil79a, Mir80a, Sir78, TTC78, VT76]. Section [Ano61e, Ano62c, Ano62d, Ano64g, Ano64f, Ano64e, Ano65c, Ano66g, Ano67g, Ano69g, Ano71f, Ano73i, Ano74h, Ano75j, Ano76e, Ano76d, Ano76f, Ano77g, Ano77h, Ano78h, Ano78k, Ano79i, Ano80i, Ano80j, Ano81h, Ano81i, Ano81l, Ano82q, Ano82g, Ano82l, Ano82n, Ano82h, Ano82j, Ano82k, Ano83c, Ano84h, Ano84j, Ano84l, Ano84i, Ano85p, Ano85k, Ano85n, Ano85o, Ano87k, Ano87n, Ano87j, Ano88l, Leo82, ZI74, Ano74g]. Segment [Pec74a]. Segments [Amb74, Kud74]. Selected [Tak56]. Selection [CZ79, Hau93, Mik58]. Self [Sin76, Ego97]. self-normalized [Ego97]. Self-Similar [Sin76]. Selfadjoint [GV86]. Semi [BP73, But79, Ers79, Fai82, Gal78a, Gor78a, Gor80, GM84, KS80, Kha77, Kha81, Kha84c, Kit82, Kit86, KS85a, LS59, Lev72, Lev73, Lev74, Lev80, LS84, Pec74a, PS77, Sīl92, Tom90, VGR84, Yus82, Zhu70, Zhu72, DRS89, Svi90]. Semi-Groups [But79]. Semi-Interval [Pec74a]. Semi-Invariants [LS59, Zhu70, Zhu72]. Semi-Markov [BP73, Fai82, KS80, Kha77, Kha81, Kha84c, Kit82, Kit86, KS85a, Lev72, Lev73, Lev74, Lev80, Sīl92, Tom90, VGR84, Yus82, DRS89, Svi90]. Semi-Martingales [Ers79, Gal78a, GM84, LS84]. Semi-Ordered [PS77]. Semi-Stable [Gor78a, Gor80]. Semicontinuous [Ali92, Kit82]. Semifields [AS82]. Semigroup [MP78b, Pan91, Par67, Sac71b, Ula84]. Semigroups [Den92, Dyn56b, Haz96, KS91, Mak73, Zak90b]. Semimartingale [DV96, Kva96, Leb90, Sha93, Tou92]. Semimartingales [Gal79, Gal85a, Gal85b, Gal94, Had78, Leb81, Leb82, LS81b, LS81a, LS82, Mel89, MN89, Sch93b, Tar85, Tro91, Vet82, Zan99, Tou92]. Séminaire [JK90]. Seminar [Ano57b, Ano57a, Ano58d, Ano61d, Ano64d, Ano67f, Ano68f, Ano73j, Ano75k, Ano76f, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano79f, Ano79i, Ano79h, Ano80e, Ano80f, Ano80i, Ano80h, Ano80g, Ano81b, Ano81h, Ano81i, Ano81j, Ano81k, Ano82b, Ano82q, Ano82l, Ano82h, Ano82i, Ano82j, Ano82k, Ano83d, Ano83c, Ano84h, Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o, Ano85r, Ano86m, Ano87i, Ano87o, Ano89a, Ano89c, Ano92d, Ano93d, Ano98d, Ano99d, Bol81, NK79, Pet59c, Pro83, Ano59e, Ano60c, Ano61g, Ano62e, Ano63c, Ano65d, Ano69h, Ano71c, Ano71g, Ano73h, Ano73k, Ano74j, Ano74i, Ano75i, Ano76g, Ano78j, Ano78k, Ano79j, Ano80j, Ano81l, Ano82g]. Seminar [Ano82p, Ano82n, Ano82o, Ano82m, Ano83e, Ano83f, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87l, Ano88k, Ano88l, Leo82]. Seminars [Ano84g]. Seminary [Ano79g]. Seminorms [Kha96]. Semiordering [Kal69]. Semiparametric [Dok93]. Semisimple [Vir70]. Semistable [Cho87, Gri94c, GK96b, Gri96b]. Sen [Fer98]. Sense [Fre67, Mkr81, Mol67a, Sid57, Ver66]. Sensitive [Yus91b]. Separable [BS74, Bul85a, Iva86, IM79, IM89, Kud86, Kud98, Med77, Med78, Mir86a, Mna86, Mna88b, Ron89, Sie90]. Separated [Zai92]. Separating [VKR64]. September [Ano57b, Ano59e, Ano65c, Ano73i, Ano83c, Ano85j, Ano85m, Ano87k, Ano87n, Ano88l, Ano99d]. September-December [Ano59e, Ano65c, Ano88l]. Sequence [BZ84, Bas59, Bor89a, BD91b, CL94, CP91, Chu92b, Dar76, DB81, Dar85, Dar86, DB88, Dob58c, EM68, Elb83, Fry89, Gik58, GT77, GZ66, Ibr65a, IS68, Ivc75a, Kap74, Kif69, Lin92, ML66, Mik74, Nev88a, PL83, Pis80, Ron88, Sal90, Sin57, Sun96, Tak56, Tak68, Tin62b, V′y98b, Zäh80, Zhu66, ZM74, ZM80, Fry90]. Sequences [Ano85a, Asa88, Bra90, BS88, CFZ95, Chu92b, Cow93, DM80a, DM80b, Dyn65, Ego70, Fäh83, Gap70, Gap74, Gap88, Gap89b, Gap73, Gol92b, Gol94, Gor78b, Gor84b, Gri92a, Gri94b, GE91, Hög93, Ibr65b, Ibr70, Ism73, KL83, Khi56a, Khi56b, Kor93, Leb87, Lu’93, Mak71a, Mal94, Mal97, Mel83, Muc88, Nik99b, Nov95, Nov96, Oli93, Pak59, Piu94, Pos61, Ric65, Ric68, Rog68b, RS88, Saw93, Seĭ91, Ser80, Shn87, Sob64, Ste84, Thu85, Vov88, Vro99, Vve85a, Vve89, V′y98a, Zol76, Kor95, Kor96]. Sequential [AB93, Bor70, BS89, DS85, Efr80, Efr90, Erm89, Erm90, Ers70, Fak69, GR85, Glo70, GK84b, Gol85b, GS65, IH75a, IH75b, II77, Kel77, Kho67, Koz71a, LK74, LS65, Lot86, Lot88, Lot89, Mal93, Mal95, MN89, Mik58, Mik56, Mir78c, Nov88, OCG93, Par70, PK87, Pav88, Pav91b, Páz68, Per92, PS93, Rao89, Rom69, Sen74, Shi64b, Tik78, Tik82, Tsi85, VT92, VN99, Vor65, WW92, Yas84]. Sergei [Ano69f]. Series [Ant80, BP93, Bul74a, BP78, CC96, Cho84, Die77, Dza74, Gap68, Gri96a, Ibr67b, IN78, Ivc74b, JP75, Kle96, Kud74, Kud84, Meg86, Mir77, Mir81, Mir84b, Nag71a, PT91, Roz99b, Tie77, Tie80, TP59, Tsi76b, VMS95, Woy74, Zhu90]. Series-Scheme [Die77]. Server [Bas65, KS86, Pre65, Sei87a]. Servers [BB81, Tsa83]. Service [Bor63, Bor64b, Bor65a, Bor65c, Bor66, Bor68a, Khi62, Kli75, Kli79b, Rog66b, Sam63, Vin94, Vin96, Yak91]. Servicing [Fed77, KLP89]. Serving [Ivn93]. Sessions [Ano57b, Ano57a, Ano59e, Ano60c, Ano61e, Ano62c, Ano62d, Ano66g, Ano67f, Ano67g, Ano68g, Ano68f, Ano69g, Ano69h, Ano71c, Ano71f, Ano71g, Ano73h, Ano73k, Ano73j, Ano73i, Ano74h, Ano74j, Ano74i, Ano74g, Ano75i, Ano75j, Ano75k, Ano76e, Ano76d, Ano76f, Ano76g, Ano77g, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano78j, Ano78k, Ano79g, Ano79f, Ano79i, Ano79h, Ano79j, Ano80e, Ano80f, Ano80i, Ano80h, Ano80g, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano81k, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82o, Ano82m, Ano82h, Ano82i, Ano82j, Ano82k, Ano83e, Ano83d, Ano83c, Ano83f, Ano84h, Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o]. Sessions [Ano85r, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87l, Ano88k, Ano88l, Leo82, Pet59c]. Set [BS73a, Bez78, Bez80, Chi76a, Cho84, Ing89, IM93, Kha68, Kha69a, Kov86, Lic69, Meg86, Mes63, Mik97, MFY74, Nor93a, Rén56, RS82, Rug93, Sac73a, Sko74b, Tak76, Zai92, Vor63b]. Set-Indexed [IM93]. Sets [Aga80, Cra56, Fai76, KY64, Lar94, Lya87b, Mol88, Mol91, Mol92, Mol94, Pes94, Pit73, PRR87, Riv96, Sac74, SS62, Sev72, Sie80, VL76, Vin86a, Vin86b, Vla96, Zal86, Zol66c, Pus97, Tou92]. Setup [VD93]. Sevast’yanov [Ano94a, Smi90b, Amb76]. Seventh [Ano64c]. seventieth [KKSY95]. Several [Bas65, Che57a, Che80, Chi73b, GK84b, Gri87, Kag89, Kra67, McK63, PG80, Pag86, Páz68, Sag83, Sag86, SG62, Tri90, Vat77d, Vat79, Vat80, Vat81, VG90, Zol57b, PR97]. Several-Dimensional [McK63]. Shannon [Sav64]. Shape [Ken87, Ken95, Le93]. Shaped [Ivn98]. Sharing [Gri92b, Kli75, Kli79b]. Sharp [CC96]. Sharper [PY79]. Sharpness [Mal68]. Sheet [Nas88]. Sherman [Kud78]. Shevchenko [Ano73j, Ano74i, Ano75i, Ano78g, Ano78j, Ano79g, Ano79h, Ano80e, Ano80g, Ano81k, Ano82o, Ano82i, Ano83e, Ano83d, Ano84k, Ano85l, Ano85m, Ano86l, Ano86j, Ano87i, Ano88k]. Shift [Art89, Kho83, Mak75, Mar69, Nik74, Rao76, Tho93b, Yus91a]. Shift-Coupling [Tho93b]. Shifts [But77, Mak67a]. Shiryaev [Sal90]. Shiryaev-Optimal [Sal90]. Short [Ano97g, Ano97h, Ano97i, Ano97j, CPP98, Cle98, Dyn98, Fer98, IS98, Ivn98, Jak98, KL79, Mik98, Ost98, Pis98, Rad98, Sha61, SZ98]. Shot [Rya66]. Shrinkage [Sal93a]. Siberian [Ano80j, Ano81l, Ano82g, Ano82n, Ano82j, Ano82k, Ano83c, Ano84j, Ano85p, Ano85k, Ano85n, Ano87k, Ano87j, Ano88l, Leo82]. Side [Bor78b]. Sided [Bet86, BZKR70, DM68, Epa68, Epa74a, Epa74b, GN87, Mar84, Mar86a, Nag82, Nov83, PS92a, RS82, Roz86, Tik85, ZS76, PS98, Zub99]. Sides [Kha66b]. Siebert [Haz96]. Siegel [ZI70]. Sieve [Pla92]. Sign [KSS77, Zin88]. Signal [Bur79, BB91, Erm91, GD71, IK81, Kho71a, Kor88b, Lep94, Lin66, Nem91, Vol64, Won59, Kra94]. Signals [BCK^{+}92, Bas58, Dob58d, Glo79b, GK84b, LPS67]. Signed [Pro92, PR85, PW87, Shu79b]. Significance [Epa68, Epa74a, Sal98]. Signs [Dob60a]. Similar [Bre65, Gik56, Oso56, Sin76]. Similarity [Sko92, Ste89, Ste91]. Simple [Bel67a, Eag77, HP94, IJL91, Ing83, Ing89, IL78, Kab96, Kag86b, Khm75, Kry91, Mir85, Mir78c, Mir80b, PP93, PS75, SS93, Vat96, Zas67]. Simplest [AK70, Tut64]. Simplicity [Vir84]. Simplified [Dob58c, Dzh75]. Simulating [Mik75a]. Simulation [Shk74]. Simultaneously [Tum56]. Sin [Bre65]. Single [AHE85, Bas65, Che64a, Che70, Ivn90, KS86, Mak91, Rya66, Sak56, Ste71]. Single-Line [Ivn90]. Single-Server [KS86]. Single-Shot [Rya66]. Single-Type [Mak91]. Singular [Bab85, CN87, Mat60, Pol61, Roz73, Tik88, Yur71a]. Singularity [Tol99]. Singularly [Shu91a]. Sirazhdinov [Ano89g]. Sixth [Ano61a]. Sixtieth [Pro91b, KGPS91]. Size [Bra91, DM68, Epa74b, Fok77, IT83, Lab72, Mog77b, Pav78, Sta97, Vin68, Vol78a, Vol78b, Vol79, Vol80, Vol81]. Sized [Tyu74]. Sizes [Epa68, Fil61, Rya76]. Skew [Ose65, Ouk91]. Skew-Brownian [Ouk91]. Skewness [MRS94]. Skipping [Kal93a]. Skitovic [Chi70a]. Skitovich [Kag89, Fel93, Fel98, Kag88]. Skorokhod [Bog89, Jak98, Kol99, Kol56a, Reb88]. Slavs [Bob64]. Sliding [Bol90]. SLLN [Vro99]. Slope [Ger96]. Slow [Dub91, Ger96]. Slowly [Leo86, PL83]. Small [Bla62, Che64a, Che64b, Che67b, Chi76b, DJ99, Dob62b, Dub74, Epa68, Epa74b, Fre64, Fre67, GP71, Gri75, Has66, IM79, Kha66a, Kha63, Kif75, Kif77, Kog65, Kog66, Mat65a, Mat65b, Mog75, Nai82, Nev64, Nik64, Pet56a, Rez77a, Rez80, Saf92, Sar90, Sha96, SS65a, Uri93, VTF73, VT76, Ver99a, Vol65, Vol70, Zai84]. Smallest [Pit87b, PS92a]. Smirnov [Bol90, Che56, Dar60, Ano66d, GKPS60, Mak88, Mar78, Orl75, Tim84, TC86]. Smirnov-von [Orl75]. Smooth [Erm89, GK96a, Gol94, GPU98, Gus76a, Gus76b, IK73b, IK74a, Jur74b, Kut78, Mol68, Nys87, Vin86a, Vin86b, Wol76, MP97]. Smoothing [Gam82, Jur75a, Mir84c, Roz84b, Sve87, Zai87b]. Smoothness [Efr86, Gam89, Ibr84, Tol99, Ugl89, Gam86]. Snell [Sha93]. Sobolev [CC96, Roz96]. Society [Ano61e, Ano62c, Ano62d, Ano64g, Ano64f, Ano64e, Ano65c, Ano66g, Ano67g, Ano68g, Ano69g, Ano71f, Ano73i, Ano74h, Ano74g, Ano75j, Ano76e, Ano76d, Ano77g, Ano88a, Ano90b, Ano90a, Ano90h, Ano93a]. Sojourn [Afa65, AM69, DR79, Dok99, Eli89, Gus84, Tak58, Ver91, Vin96]. Solenoidal [Abr59]. Solution [AA68, Anu91, Bei64, BM92, Dav82a, DN88, Dub91, Fed85, Fre65, Gal79, Gir62, Kab93, Kha60, Khr86, Kor73, Kor78b, Kry72, Kry91, Kul67, Kul68, Kul74, Leb77b, Leb77a, Leb78, Leb84, Leb86, Lët91, Mat65a, Mat65b, Mil96, Mil79b, RU57, Sob60, Tsa77, Tsi76a, Ver83, Yan70, Zam70, Den97, Mil97].

Solutions

[Bal74, Bor78b, Bor99, Bul64b, Dor96b, Gyö95, Kha59, Kha66b, KV85, Kle85, Kol69b, Kop96, Kov67, Kry85, Kul83, Lan88, Lan90, LSS99, LO96, Lev83, Lev86a, Mak96b, Mas99, Med98, Mel79, Mik97, Mik99, Mik58, Mik56, Mil86, Mil88, Pis98, Por75b, Roz64b, Sar90, Sok84, Tsa79, Ver80, Ver82, Ver88b, Yag60]. Solvable [Mol67b]. Some [AHE85, Ale64, Ale67, Ale81, Ale92b, ARS^{+}97, AN74, AB93, Ark65, Asa88, AN93, BBM59, BM90, BM91, Bag89c, Bas58, BR83, Bor81a, Bor80a, Bor62, BR64, Bor64b, Bor65a, Bor65c, Bor68b, Bor72, Bor78d, Bor80b, Bor86b, Bre65, BD91a, BD91b, Büh68, BP78, Bul64a, But92a, CGM97, Chi70c, Chi70d, Dar96, Das94, DI71, Dob57, Dob63, Döh80, Dok69, Eag77, Ego72, EN83, Eid73, EK94, Fak82, GS88, Gal77, Gap70, Gap78, Gik56, Gus76a, Gus76b, Har77, Ibr62, IK73a, IK79, INK87, IL94, IT79, IM65b, Ivc76, Ivc85, JK90, Kal93a, Kar69, Kel77, Ken64, Kha84a, Kha90b, Kho73b, Kle79, Kle81, KS87, Kon74b, Kor86a, Kor61b, Kor57, Kor66, Kor68, Kos85, KS88, Kov61, Kra94, Kra79, Kra80]. Some [Kra62, Kry73c, KST61, KY74, LS60, Lin57b, Lip66, Maĭ95, Mal74a, Mal65a, Mal77, MV94, MS90, MT90, Mar65, Mas60, Mkr81, Mol67a, MS86, Nad64b, Nad74b, Nag57, Nag65, Nag68b, Nag69d, Nag76a, NP78, Nag84a, NS89, Nag98b, Nor91, Pav92, Pav97, Pav98, Pec74a, Pec77, Plu92, PY69, Pra93, Pro59b, PS61, Pro65a, Rez68, Rez71, RR99, Rog58, Rog65, Rog66b, Rot74, Rot76a, RS86, Roz66b, Roz81a, Roz91, Roz67, Roz86, Ryz69, Sad66a, Sad67a, Sal71, SG62, Sel69, Sen82, Sen89, Sen92, Shi80, Shi82, Shi60, Shi65, Shi95, Shp68, Shu70b, SS92, Sli62, Sli64, Ste88b, Stu67, Stu68, Syt67, Tak57, Tar90b, Tik88, Tsi76b, Tsi91, Tut69a, Tut69b, Ush82, Vak94a, VTF73, Ver64b, VC66, Vis92, Vla84, VR59]. Some [VR61, Vol61b, Woy74, WW92, Yag57, Zai81, Zie92, Zol68b, Che67b, Ivk94]. Songs [Bob64]. Source [Roz94a, She86]. Sources [Ivn93]. Soviet [Ano70f, Ano90b, Ano90h]. Soviet-Japanese [Ano70f]. Space [Ale92a, Ali87, Ali90, AV82, BO96b, BPS97, Bul74b, Bul74a, CGM97, Chu92b, Cra64, Doo57, Dor90b, Dyn70, Fed82, GS76, Gia96, Hid70, KS64, KS80, KT59, Kar86a, Kar86b, Ken95, Kho83, Khr86, Kir94, Kol99, Kru71b, Kru73a, Kru74, KA83, KA85, KA87, Kuz81, Lin90, MO83, Mas72, MP92, Met75, Mog75, Mus77, Nag86b, Nas88, Oza86, Pav98, Pin79, Pin81, PF57, Pro84, PRR87, PS92b, Rac85b, RS98, Sak85, Sal80, Saz90, Saz92, SS62, SS88b, Sem88a, Sen90, Sen92, Shp68, Sko69, Sko70b, Sko70a, Sko74b, Sko79, Sko80b, Smo89, Smo91, Szu83, TT83, Thu85, TTC78, Tie80, Tik92a, Tik94, Tom90, Tut62, Ven86, Vla80, Vve84, Vve85b, Vve89, Yag56, Yag57, Yur92a, Yur93a, YF79, Yus76, Yus78]. Space [Zal83, ZS84, Zal86, ZSU89a, ZSU89b, ZSU89c, Zig82, de 93, JK90, Met74, Roz97, Ul′91, Zal90b]. Spaces [Bau96, BP89a, Ben62, BR85, Ben87, Bor67a, BS74, BS81, Bra90, BP78, Bul81a, Bul85a, Cap98, CC96, Chu92a, Chu92b, DF65, DS85, GH93, GT77, Jak98, Jam78, Kai86, KL83, Kol92, KS91, Küh81, KT77, Le93, LP75, LTC80, Meg86, Mog80, Mus69, Mus73, Mus80a, Mus80b, Mus85, Mus70, Nor93b, Pau77, Pau79, Pau91, PS77, Raj73, RS94, Rén56, Roz64a, Roz96, Ryk66, SS62, Sen93, Shu58, Shu89, Shu90, Shu91b, Sie90, Sko92, SU90a, Tem73a, TT80, TT82, Tie77, Tie79, VT78, VK97, Vak99, Vak66, Vak90, Vla84, Woy74, Yam85, Yur92b, Zal89, Zal90a, Zol77a, Zol78a, Chu94, Fat97, RT98, Shu91c]. Spacings [BDEM92, Bor88a, Chi61a, Kha90a]. Span [PP82]. Spatial [McK93, VW99]. Spatially [ZT92]. Special [Chi74b, Gur65, IS68, Mal65b, Sev57b, Zol95]. Specifications [Kuz85]. Specified [GW93]. Spectra [Rya79, Zhu86]. Spectral [Ale74a, Ale74b, Ale81, Ale89, Ale91a, Ale94, Ara86, BR83, Dha78, Dza70, Dza74, DM75, Dzh71, Gap80, Gap97, Gin89a, Gin89b, Gir77, GV86, Gol94, Ibr63b, Ibr65a, Ibr67b, IS68, Ivk86, Jes81, Lax63, LS60, Mal64, Neu65, PY69, Ras77, Roz59b, Roz60b, Rud93, Sch65, Ser96, Shi60, Sin63, Sol95, Zhu72, Zhu81, Zhu84, Gir95]. Spectrum [Ibr65b, Ibr70, Lin59b, Lin59c, Mal65a, Pis65, Pit87b, Rez81, Vir84, Yag60]. Speed [Bor85c, Gam66, Gap68, Kol66, Mak66, Nag70b, Nag70c, Nag70d, Nag85, Nev71, Nev72, Rog66c, Rot70a, Rot70b, Sah75, Sen82, Sho79, Ul’77]. Sperner [Mes63]. Sphere [Tut61]. Spherically [Koz84]. Spline [GN93]. Split [Kuz75, Kuz78]. Splitting [Fil61, GE93, Pit87b, EO79]. Springer [JK90]. Sptimal [GZ66]. Square [Bak96e, BM79, Chi71a, Dyu68, Erm96, Erm98, GR85, KP97, Mil88, Mir90b, Nad74b, Nik74, Ron85, Ron86c, Sko75, Tie80, Tsa77, Yan64, Zue88]. Square-Root [Ron86c]. Squares [BL63, Iva77, Kak78, KC91, Kur87, Lin57b, Mal69b, ZK92]. SSSR [Ano61g]. Stability [Abr87b, Ano81b, Ano82b, Ano84g, Ano86m, Ano89c, Ano92d, Ano93d, Ano98d, Ano99d, BL75, Bor78d, Bor79c, Bor79e, Bor79d, Bor80b, Bor81b, Bor85c, Bor89a, Bor91a, Chi87b, Chi95, Fed69, Fed70, Fre74, Fre76, Gab75, Kal77a, Kar86a, Kar86b, Kha67, KY86, Lin59c, Mac69, Mac71a, Mal68, Mkr81, Mov77, Nhu68, Nik91, Nor91, PS85, Pan89, Pan91, Pen93, Pro58, Roz64b, Tin62b, Tsa77, Tsa79, Tsi76d, Ush84, VTF73, Yan85, Yan79, Yan89, Zol68a, Zol90]. Stabilization [Kha60]. Stable [BZKR70, Bra86, BB91, Cap98, Chu92a, Chu95, Dau88, DJ99, Dob58d, Ego80, Gal81, God79, Gor78a, Gor80, HJ94, HMV95, IC59, Jur83, Kho83, KMM85, Klo62, Kor98, Kor99, KB92, Kru73b, Kru93, Lis94, MA78, MU83, Mit62, MR93, MO69, Nag84b, NS89, Nev73b, Pet74a, PS92a, Rag86, Rog73, Rog77b, Roz98b, Sak56, Sat73a, Sat73b, Sie90, Smo89, SL90, Smo97, Sol95, TT80, TT83, Vas80, Zak91, Zan99, Zin86, Zin88, Zol67b, Zol95]. Stably [GM84]. Stage [Bas58]. Standard [Ber76, DN75, Ula84]. Standby [Mak96a, Szt93]. Standstill [Sev62]. Star [Ivn98]. Star-Shaped [Ivn98]. Starting [DY68]. State [Ano73j, Ano74i, Ano75i, Ano78g, Ano78j, Ano79g, Ano79h, Ano80e, Ano80h, Ano80g, Ano81k, Ano82o, Ano82i, Ano83e, Ano83d, Ano84k, Ano85j, Ano85l, Ano85m, Ano86l, Ano86j, Ano87i, Ano88k, Bal94, Bul64b, Dob73, Fai76, Gne90, Gri92c, Ivn93, Ivn98, Ryk66, Sav99, YF79, Yus76, Yus78]. State-Dependent [Ivn98]. Statement [Nag61]. Statements [Zol57b]. States [Aga80, Amb79, Bel69b, Bel78, Bez78, Bez80, Dob56c, Dob74, GS85, Ivn90, Ivn98, Mal91, Mes58, Mik96a, Ose90, Rom64, Sam82, Sen68, Son92, Vol57, Vol58b, Vol60b, Yak87, YF80]. Stationarily [Gol65]. stationarity [Bol97]. Stationary [AA68, Ale74b, Ale94, Apo67, Ara61, Ara68, Bar91, Bel60, Bor72, BK97, Bul61, Chi90, CDC82, Dav68, Dav70, Dob58c, Dob60b, DK84b, Dza70, Dza74, Dzh71, Eli85, Eli89, Fai79, Faj88, Fal88, FS79, Gap74, Gap78, Gap80, Gap82, Gap88, Gap89a, Geo80, Gin89b, Gla58, Gol65, Gol92b, Gol94, Gor71, Gor83, Gri83, GK69, Ibr62, Ibr63b, Ibr65a, Ibr65b, Ibr67b, IS68, Ibr70, Ivn82, Ivn90, Ivn98, Kha93, KF67, Kol69b, KR60, Leo61, Mal64, Mal65a, Mal80, Mal85a, Mal91, MO83, MW59, Mat60, Mat61, Mat62, Mir74, Mir75, Mir78a, Mir78b, Mir80a, Mir81, Mir85, Nad73b, Nag57, Neu65, Nos87, PV64, Pit78, PP82, Pit71a, Rez68, Rez71, Rob63, Rom65, Roz59a, Roz60b, Roz61, Roz63, Roz64a]. Stationary [Roz64b, Ryz65, Ryz67, Ryz69, Ryz71, Ryz72, Ser80, Shi63a, Shu65, Sin60, Sin62, Sin63, Sli62, Sli64, Thu85, Tsa58a, TP59, Vas70, Veĭ93, Ven90, Ver66, Ver64a, Vil59, Vit70, Yag57, Yag60, Yam95, ZT92, Zas64, Zhu72, Zhu81, Zhu84, BO97]. Statistic [Bak89a, Bel67a, Bul85b, Chi73a, Chi81a, Chi81b, CH93, Gel74, HP94, Kle71, Kle74b, KR75, Kle80, KC75b, Kud78, Lin56a, LS96, Mir86a, Orl75, OT95, Pen77, Sel85, Shu91d, Mak88]. Statistical [AHE85, Ale67, Ale91a, Ale92b, AT93, Bag79, Bas59, Beg69, BR83, Bog82, Bol64b, Bol65a, BM93a, Bun63, Bur82, Che81, Das79, Der93, Die77, Dob58d, DM67, Fed69, Fed70, Fil62, Gad90, Gas63, Gir88, Gor81, Gus76a, Gus76b, Gus93, GV93, Gya64a, IK73b, IK73a, IK74a, Ivc85, IT89, KK99, Ken87, Kha84a, Kha84b, Kho73b, Kor88a, Koz73, Koz74b, Kra94, MS90, Mik56, Mon56, PK82, Pes94, Pet56a, Pla92, Sap79, Sap88, SS62, ST73, Ste84, Tik98, Tou94, Vol78a, Vol78b, Vol79, VN94, Yag56, Zas67, Zer87, Zhu86, Zub74]. Statistics [Ale91b, Ano61a, Ano61e, Ano61g, Ano62c, Ano62d, Ano62e, Ano63c, Ano64c, Ano64g, Ano64f, Ano64d, Ano64e, Ano65c, Ano65d, Ano65e, Ano66g, Ano67f, Ano67g, Ano68g, Ano68f, Ano69g, Ano69h, Ano71c, Ano71f, Ano71g, Ano73h, Ano73k, Ano73j, Ano73i, Ano74a, Ano74h, Ano74j, Ano74i, Ano74g, Ano75a, Ano75i, Ano75j, Ano75k, Ano76e, Ano76d, Ano76f, Ano76g, Ano77g, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano78j, Ano78k, Ano78f, Ano79g, Ano79f, Ano79i, Ano79h, Ano79j, Ano80e, Ano80f, Ano80i, Ano80h, Ano80g, Ano80j, Ano80k, Ano81h, Ano81i, Ano81j, Ano81l, Ano81k, Ano82a, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82o, Ano82m, Ano82h, Ano82i, Ano82j, Ano82k, Ano83e, Ano83d, Ano83c, Ano83f]. Statistics [Ano84h, Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j, Ano85l, Ano85k, Ano85n, Ano85m, Ano85o, Ano85r, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87l, Ano88a, Ano88g, Ano88k, Ano88l, Ano90b, Ano90j, Ara61, Ara86, AN93, BD78, BdH79, BG96, BS73b, Bol94, Bol81, Bor88a, Bor85d, Bor86b, BPS97, CdHY97, Chi64a, Dyn68a, Dyn68c, DN75, EN75, EN83, Ego91b, Gik56, GKPS60, GK81, Gri89, Gri94a, Gru86, Gru88, IL94, Iva86, Ivc75a, Ivc71b, IL76b, IM79, IM81, IM87, IM89, Jan76, Kha88, Kha93, Kha94, KA99, Khm81, Khm84, Kho79, Kle74c, Kle75, KY84, KN90, Kol69a, KB86, KMT78, Kru80, Kud86, Kud98, Leo82, Maĭ95, Mak67a, MA78, MA79, MA83]. Statistics [MT90, Mar76b, Med77, Med78, Mik75b, Mir88, Mna86, Mna88a, Mna88b, Mog77c, NK79, PK87, Pro65a, Pro83, Puk82, PS69, PR85, PW87, PS89, PS92b, Ron82, Ron84, Ron86b, Ron89, Ruk76, San66, Saz92, Shi64b, SMI90a, Tik88, Tim84, Tim88, Van83, Yag56, Yur71b, Yur72, ZI74, Zhu66, Zhu90, Zin58, ZL64b, Zin76, Zin80, GK80, TC86, Ano59a, Ano81a, Ano89k, Bor92c, KB92]. Statulyavichus [KGPS91]. Steady [Bal94, Bul64b, Gri92c, Ivn93]. Steady-State [Bul64b, Gri92c]. Stefan [GS66]. Stein [AB93]. Steklov [Ano65d, Ano81j, Ano82p, Ano82m, Ano83f, Ano85q, Ano85o, Ano85r, Ano86k, Ano87n, Ano87m, Ano87o, Ano87l]. Step [Gri63]. Stepanov [Goo59, HR97]. Steps [KS97]. Sticks [Roy93]. Stieltjes [Zol57a]. Stochastic [AR94, Ale63, Ale64, Ale67, Ale94, AK89, Aly88b, Ano81b, Ano82b, Ano84g, Ano86m, Ano89c, Ano92d, Ano93d, Ano98d, Ano99d, Anu91, Ara90, AI99, BN83, Bel66, Bel67b, Bel68, BPRS94, Bev58, Bor79b, Bor79e, Bul87, But79, But92b, Che56, Chi81a, Chi81b, Chi61b, Chi65b, Chi73b, Cra64, Cra65, DP75, Dar99, Dav76a, Dav68, DM96, Dor90a, Dor96b, Dor96a, Dub91, Dyn60, Dzh77, EW99, Ers75b, EK82, Fak70, Fed85, Fok77, Fre73, Fre74, Fri74, Fro89, Gal79, Gal83, Gal85b, GK75, Ger96, Gih91, Gik90, Gin66, Gir60a, Gir62, Gla61, Gla65, Glo79a, GK84a, Gor89, GV93, Gyö95, Has66, Ibr74, IK99, IMS93b, Iva60, IM93, Kab74, Kab76, KM65, KS69, Kha67, Khr86, Khr94, KV85, Kle85]. Stochastic [Kle86, Kli62, Kog65, Kog66, Kol69b, Kor78a, Kor80, Kor84, Kra62, Kra69, Kra71, Kry69, Kry71, Kry72, Kry73a, Kry73b, Kry73c, Kry85, Kul98, Kul67, Kul68, Kul74, Kul76, KH81, Kul83, KK96, Kut76, LS97, Leb73, Leb77b, Leb77a, Leb78, Leb84, Leb86, Leb90, Leb96b, Leo61, Lif82, Llo62, Lot96, Mak96b, Mak65, Mak97, Med98, Mel79, Mel83, Met75, Mik75a, Mil79a, Mil82, Mil86, Mil88, Mil96, Mil75, Mir69, Nad73b, NS70, Nas96a, Nas97, NRV67, Nev73a, Nev86, Nor88, Nor95, Nov71a, Nov73, PV64, PS93, Pis95, Pis98, Por75b, Por80, Pra99, Pre98, Rac85a, Rao63a, Rao76, Rie81, Rob63, Roz84a, Roz88, Roz94a, Roz96, Roz99a, ST74, Sal80, Sat95a, Ser61, Sev57b, Sev61, Shi67, Shi95, Sir73]. Stochastic [Sko56, Sko57b, Sko57a, Sko61b, Sko62, Sko76, Sko80a, Sko81, Smi69, Sok84, Sza71, Tak57, Tak76, Tou94, Tsa77, Tsa79, Tsi76a, TM79, Urb56, VK67, VL76, Veĭ93, VT82, Ven77a, Ven77b, Ven80, Ven83, Ver80, Ver82, Ver83, Ver88a, Ver88b, Ver92, Ver99b, Ver99a, Ver64b, Vet81, Zam70, Zol76, Zol77b, Zul76, Den97, IMS93a, Ivk94, Met74, Mil97, PS98, Sev59]. Stochastically [Bla89]. Stock [BV95]. Stocks [DKR95]. Stopped [BU94, BK86, Kif71a, Kog70, Mal94, Mal97, Sev99, Sza71, Tei66]. Stopping [Anu81, CGM97, DSS94, Dyn68a, Dyn68c, Elb76, Elb83, Eng74, Eng75, Fäh78, Fäh83, Fak70, Fak71, Fri69, Gol85b, GS65, GS66, GZ66, Kif71b, Koc98, Kuz85, Mih81, Mir76, Nik99b, Nov71b, VM89, VM91]. Storage [PP95, Sev62, Zak90a, Zak91]. Straight [Pit75]. Strassen [Nor98]. Strata [Pav89b, Pro74, Ste69c]. Strategies [Ark64b, Fai82, Fai87, Faj88, Fri72, Gor85, Kit82, Kry87b, Nov97b, She88b, Vor60, Yan70, Yus74]. Strategy [Kif69, Kry65]. Stratification [Lif85b]. Stratified [Ivc85]. Stratonovich [LS97]. Stream [Vin96]. Strela [Sob58]. Strength [Ivc74a]. Strengthening [NR73, NR76, Sal93b].

Stress [DK93]. Strict [Sid57]. Strict-Sense [Sid57]. Strictly [Lis94, Smo97, Thu85, Vov97]. Strip [AB85, Bor79d, Lot92, Nos83b, Nov82]. Strong [AL77, BCK^{+}92, BGW75, Bin90, BW98, BS87, BS88, CS61, Dav81, Dav82b, DY56, Dyn59, Ego70, Ego72, Ego91c, EW99, Fed85, Gap74, Gap78, Gap82, Gap86, Gor78b, Gor83, Ibr65b, Ibr70, KLS84, KV85, Kle85, KR60, Kur88, KT77, Lei82, Mar80, Mar84, Mar85, MP93a, Med98, Mel79, Mey63, Mik96b, Mol68, Nag73, NV76, Nik80b, Pet69, Pis80, Pis98, Pro58, Pro59b, Roz81b, Rus96, Ser80, Tsi76a, Tut69b, Ver66, Ver64a, Ver80, Ver82, Ver83, Ver88b, Vol58a, VN78, Yus57, Yus60, ZM79a, Zhu72, Zin86, Gap95, QT97, Vov97]. Strongly [Gap69, Gir60b, Gor84b, Vin76, Yam96]. Strongly-Feller [Gir60b]. Strook [Kul98]. Structurally [Fal88]. Structure [BP91, Ber84, Fed77, KF67, Mon58, MFY74, Nos86, Pit87a, Pit87b, Puk82, Rez78, Sap79, Saw93, Sko66, Ste73, Ste88b, Tie79, TF62, ZK76]. Structures [Sap88, Zer87]. Student [JL56]. Studentized [PRR82]. Study [Kog65, Kog66, Lif82, Lif85b, Mok91, IK73b]. Sub [Pas99]. Sub-Gaussian [Pas99]. Subadditive [Ham66]. Subcritical [Yak86, GP97]. Subexponential [Teu75]. Subgraphs [Men88]. Subject [KR90]. Submartingales [Dam65]. Subordinate [Sir77a]. Subordinators [Rog92]. Subsequences [Jak98, Ste84]. Subsets [Mes63]. Subspaces [Gal81]. Substitution [Gal78a, Gir60a, Kog68, Vol58a, Vol61a]. Subtree [Vat94a]. Successive [Ara75]. Such [Gol65]. Sufficiency [KS67, Kry87b, SS93]. Sufficient [Ara61, Ara86, Che89, Dyn68a, Dyn68c, Fai87, Faj88, Kha96, Kha67, KR75, Kle75, Kul83, LS81a, Mak67a, Mak68, MP77, MP78a, Mar80, Mir74, Mus81, MC84, Nag73, Puk82, Shi64b, SS93, VC82, Vat87b, Zhu66, Ibr70, Lin59b, Lin59c]. Sufficiently [Zhu72]. suites [Tou92]. Sum [Ale77, AN74, AB60, Bar86, Eng84, Eng85, Gri74, JP75, Kru80, Mak82a, Mir84a, NP69, Nev71, Nev72, Osi69, Pet70, PY81, PY88, Pre86, RU88, RS80, She99, Sid91, Sin57, Sir77b, Tik87, Yag56, Yan70, RT98]. Sum-Symmetric [JP75]. Summability [Gap89b]. Summand [Bar88]. Summands [GP71, Her92, Mar87b, Mos73, NR73, Ros80, Sen86b, Sen87, She99, Tom71, VG90]. Summaries [Ano68g, Ano68f, Ano69g, Ano74g, Ano76d, Ano88k]. Summary [Ano57b, Ano57a, Ano58d, Ano59e, Ano60c, Ano61d, Ano61e, Ano61f, Ano61g, Ano62c, Ano62d, Ano62e, Ano63c, Ano64g, Ano64f, Ano64d, Ano64e, Ano65c, Ano65d, Ano66g, Ano67f, Ano67g, Ano69h, Ano71f, Ano71g, Ano73h, Ano73k, Ano73j, Ano73i, Ano74h, Ano74j, Ano74i, Ano75i, Ano75j, Ano75k, Ano76e, Ano76f, Ano76g, Ano77g, Ano77i, Ano77h, Ano78i, Ano78h, Ano78g, Ano78j, Ano78k, Ano79g, Ano79f, Ano79i, Ano79h, Ano79j, Ano80e, Ano80f, Ano80i, Ano80h, Ano80g, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano81k, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82o, Ano82m, Ano82h, Ano82i, Ano82j, Ano82k, Ano83e, Ano83d, Ano83c, Ano83f, Ano84h, Ano84j, Ano84k, Ano84l, Ano84i, Ano85q, Ano85p, Ano85j]. Summary [Ano85l, Ano85k, Ano85n, Ano85m, Ano85o, Ano85r, Ano86k, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87o, Ano87l, Ano88l, Leo82, Pet59c, Pro83]. Summation [BP78, Chi87a, Chi88, Gap65, Gap88, Hag74, ZI71]. Summing [Ivc74b, KMM85, Küh81, Mar86b, Rot76a, Vin76, Vla81]. Sums [Ale79, Ale80, Amb76, Amo80, AT93, Ara75, Ark90, BR85, Ber70a, Ber70b, Bet88, Bet89, Bik69b, Bor84b, Bor88a, Bor60b, Bor60a, Bor61, Bor68b, Bor70, Bor72, Bor73b, BU94, Bor96, BM99, BMS96, BW98, BS87, BS88, BS92, BK80, Bul84, Bul94, Chi74b, Chi64d, Cho84, DM80a, DM80b, Ego91a, Ego91b, Ego91c, EL99, FTV92, FN71, GP71, Gam74, Gam82, Gap68, Gil82, God79, Gor71, Gri63, Gri91a, Ibr66, Ibr67a, IP74, IC98, Kar95, Kat78, Kha61a, Kle87, Kol56b, Kor86b, Kor89b, Kor89a, Kor90a, Kor98, Kru71b, Kru73a, Kru74, Kru75, Kru89, Kru92, KB97, Kru99, Kud85, LM71, Lar94, Lin92, Lin61a, Lin61b, Lin62, Log91, Mar79, Mar85, MP93a, Meg86, Mes61, MSV99, Mik75c]. Sums [Mik94, MR83a, Mos70, Muk85, Muk90, Muk98, Nag70a, Nag70d, Nag77, NE77, Nag78, NP78, Nag98a, Nov97a, Nov98, Nov82, Nov83, Oli93, Osi71, Osi73, Osi74, Osi79, Pak59, Pec74b, Pet56b, Pet59a, Pet64, Pet65, Pet74b, Pin79, PU85b, PU90, Pin91a, Pre67, Pro73, Pro58, Pro84, Pro91a, RR93, Rag91, Rak88, Rak90, Ric68, Rog61b, Rog66c, Ron80, Roz90c, Roz98b, RS76b, Saz75, Sch89, Sem88b, Sev73, Sho96, Shu76a, Sla82, Sla85, Sre70, Sta65, Stu60, SU90b, Sza73, Tei66, Tik92a, Tik94, Tom71, Ul’87, Ush86, VD93, VG90, Vol58b, Vol60b, Vol96a, Vve84, Vve85b, Vve89, WR69, Zai84, Zai88b, Zin86, Zin65, Zin71, Zol65b, Zol77a, Zol78a, Zol78c, Zol92a, BO97, DRS89, Ego97, NK97, Sho97]. Sup [Nor98]. Sup-Norms [Nor98]. Supercritical [DB93, Yak86]. Superharmonic [Shu64]. Superimposed [BCK^{+}92]. Supermartingales [Lic69, Pav81b]. Superposition [Fra63, Sil74, Sil79, Zol58]. Superspace [Khr91, Khr94]. Supplementary [Cho93]. Support [Gyö95, MP92, Raj73, Tut67]. Supports [Bul85a, Tro91]. Supremum [Arn81, Bor70, Bor72, Rud93, Sgi99]. Supremum-Type [Rud93]. Sure [Gap80, Jak98, Kle96, KZ96, Mus70, ZKC91]. Surely [Pre98]. Surface [Dok99]. survival [Gil93a, Gil93b]. Suspended [Syt69]. Suzdal [Ano92d]. Switch [Sīl92]. Switch-Overs [Sīl92]. Switches [Saf77]. Switching [Che64a, Che70, SS88a]. Switchings [Isa85, KS85a, Man67].

Symbols [Kis62, Pos61, SS65a]. Symmetric [Bog80, Bra87, Bra90, Bul85b, Dob62a, Gol85a, Gri79, Ign82, JP75, Kud76, Kul98, KS85b, MO69, NS89, Pre82, Sac71a, Sac71b, Sei87b, Seĭ89, Seĭ91, Shu96, Sid91, Sie80, Stu68, TT80, Th93, VS77, VS78, Yan96]. Symmetrical [KS95, SS65a]. Symmetry [Ale94, Orl73]. Symplectic [Tut71]. Symposium [Ano64a, Ano70f]. Synthesis [Läu74]. System [BN96, BP73, Dau76, Dob70, Dob58d, Dub91, Fed77, Ivn77, Ivn93, KK82, Kha61c, Kov67, Koz78, Leb77b, Leb77a, Leb78, Lev86b, MT83, Mas99, Mat65a, Mat65b, Mik97, Mik99, Ose90, Rez77a, Rez78, Rez80, Sen84, SK93, Szt93, Tsa83, Ush78, Vin94, Vin95, Vin96, Yus88, Den97]. Systematic [Che66, Dob63]. Systems [Afa65, AM69, Aga80, AL77, Akh80, Akh81, Bal74, Bal96, Bas58, Bei64, Bei65, BRR98, Bla89, BB81, Bor63, Bor65c, Bor66, Bor68a, Bor73a, But92b, DI74, Dob74, Fal88, Fal92, Fre74, Fre76, Fro89, Gap69, Gih91, Gik90, Gir92, Gri75, Gri78, Gri91b, Gri92b, Ivn90, Kan88, Kha67, Kha61b, Kif75, Kli75, Kli79b, Kop96, Kor80, KLP89, Läu74, Leb73, Leo78, Mak91, Mak96a, Mil82, Mil86, Mil88, Mil79b, Ose65, Pav98, PP95, Pre65, Saf92, Sam63, SS88a, Sar90, Sei82, Sev57a, Sin60, Tut74, Ver92, Zol75, Zol77b, Zol78d, Glu73].

T [Ano73j, Ano74i, Ano75i, Ano78g, Ano78j, Ano79g, Ano79h, Ano80e, Ano80h, Ano80g, Ano81k, Ano82o, Ano82i, Ano83d, Ano84k, Ano85l, Ano85m, Ano87i]. Tables [BGS61, Bol64b, BZKR70, Tru66]. Tail [Bar88, Ism73]. Tailed [CFZ95, Nag99]. Tails [Bag89b, Tik98]. Taking [Lin61a, Lin61b, Lin62, Mak67b, Nag69a, Nag69b, Nag69c, Sho79, Sho83]. Tandem [Vin94]. Tarieladze [SS93]. Tauberian [Nov97b]. Tchebyshev [DJD86]. Technique [McK93]. Telegraph [Kab93, Tut73]. Telephone [Bas58, Sev57a].

Tends [Pro91a]. Teor [Bak89b, Fry90, Shu91c, SZ98, Ul′91]. Term [Bal96, Eng84, Eng85, Gel64, LM71, Nev88a, OP67, Rog74, Roz78, Roz79, Sal79, Tom71, Zol66a]. Terminating [Kog69, Man64b]. Termination [Chi76b]. Terms [AB60, Ebr71a, Mar86b, Nag71a, Nag98a, Shi63a, Vin74, VK89]. Test [Aiv65, AB93, Bag79, Bel67a, BU75, BM79, Bor78c, Che80, Chi64c, Csö75, Epa68, Epa74a, Epa74b, GR85, IK79, Ing88, II77, KK75, Kha81, Kol88, KM80, Lin64, LRS65, Lot86, Lot88, Lot89, Mal86, Mir78c, Nik74, Nik75, Nov88, OCG93, Pag86, PR87, Pag98, PT91, PS69, Rao89, Rom69, Sal98, Tik92b, Tim88, VC66, Vol69, Vol74, VN99, Bor77c]. Testing [BB61, Ber85, Bol90, Bol66, Chi67, DK93, Erm95, Glu73, GK84b, Gol85b, IV93, Ing87, Ing89, Ing95, Ivc75c, IM79, IM81, Khm80, Kho79, Kus68, KP71, Kut76, Le93, Mal74b, MO91, Mir80b, Muk78, Nik80a, Orl73, PK82, Pav88, Pav91b, Tsi85, Tyu70a, Tyu74, TF91, Uba75, VT92, Vol73, Yas84, Zem93, Bol97]. Tests [Ber78, Ber80, Ber84, Ber86a, Ber86b, Ber87, Ber76, Bol90, Bor76, Che56, Chi61a, Chi64b, Chi65a, Chi67, Chi71a, Chi86a, DP82, Dzh77, Erm96, Erm98, GP79, Goo59, IJL91, Ing83, Ing84, Ing86, IT89, Khm80, Khm82, Koz73, Kud84, Kul87, Kus68, KP71, Kuz83, Mal90b, Mar69, Mar76b, Mir90b, Nik88a, Nik88b, PC96, Ron86a, Ron89, Sen74, Ter93, Tsu81, Tyu70a, Uri93, VV88, Vol80, Vol81, VZ80, Zas67]. th [Yak96]. Their [Ano90b, AB60, AI99, Bet86, Bor79c, Bor79e, Bul81a, Har91, IT89, KU88b, LN68, Mak97, Mar85, Pin91a, Ric64, RU88, Roz96, SE96, SO85, Tan70, Tsi76d, VC71, Vor62, VZ80, de 93, Bor79d, Chi80a, VC82]. Them [Vol59]. Theorem [Afa91a, Amo91b, Ara82a, Ara82b, Ara83, Ara90, Bet89, Ble86, Bor78b, BM93b, Chi67, Chi90, Chi59, Chi64d, Cra56, Cra65, Das79, Dau76, DM96, Deo74, Dir81, Dyc73, Fel98, Fil62, Fra63, Gal78a, Gal81, Gal83, Gam64, Gam66, Gam68, GR78, Gam89, Gam91, Gel64, Gik58, Gir76, Gla61, Gur68, Haz96, Hoe64, Ibr56b, IO66, Ibr75b, Ibr89, Isa85, JR82, KK95b, Kal78, Kar80, Kar81, Kha87, Kha66b, Khr86, KY86, Kli77, Kok75, Kol67, KC74, Kon85, Kon82, Koz71b, Koz78, Kry68, KS89, KL79, Kul98, Kuz85, Lab72, Leb96b, Lët89, Lin56b, Lis96, Mac71a, Mak95, Mak65, Mal75, MU83, Mal68, MP80, ML65, Mas99, Mat86b, McD80, Mes63, Mik77, Mil88, Mna86, Mos73, MP79, Muk77, Nag75, NE77]. Theorem [Nat73, Naw82, Nos77, Nov97b, Osi66, Oso56, Ost65, Ost67, Pan95, Par71, Pet56b, Pet66a, Pet71, Pre74, Pro60, PS61, Rak88, Rak90, RS94, Rog60, Rog69, Rom59, Ron86b, Rot70a, Rot78a, Rot78b, Roz57, Roz98a, Rus96, Sal80, Sat73b, Saz63, Sev57a, Sha96, She71, Shi86, Shu77a, Shu78a, Shu85b, Shu79a, Shu81, Sir59, Sko63a, Sko74b, Smi90b, SS93, Ste89, Ste91, Ste94, Stu65, SZ98, Tem73b, Tie80, TM79, Tut78, Vak94b, Vat79, Vol57, Vol60a, VK89, V′y98a, YK86, Yus67, Zäh80, ZA84, Zai85, Zaĭ90, Zas73, ZM95, Zol67a, Mik92, Sho97, Ali87, Ali90, AR86, Bha70, BR65, BZ77, Bul96, Bun74, CPU93, DEO84, Dob56a, Dob56b, DM80a, Eli89, Eng84, Eng85, EO86, Fel93, Fom83, Gal71]. Theorem [Gam81, Gap70, Gil82, Gir80, Gir82, Gir98, Gol83, Hat75, Her91, Her92, Ibr63a, Ibr67a, Kag89, KS64, Kel90, Kha94, KA99, Khr91, KR92, Kol92, Kru82, Kub96, KT77, LM71, Lët93, Lif76a, Lif76b, Lif79, Lin59a, LS81b, LS81a, LS82, Maĭ90, MA83, Mal87a, MO83, Mat85, Mat84, Mat86a, Mik76, Mik94, Nai82, Nor93b, OP67, OR85, Par64, Pau77, Pau79, Per59, Pet66b, PL83, Pit78, Pro92, Rez81, Rie77, RS76a, Rot70b, Roz60a, Roz61, Roz76, Roz78, Roz79, RS76b, RS80, Sad66b, Sad67b, Sad68, Sal79, Saz67, SM79, Sem85, Sen81, Sen82, Sen84, Sen86b, Sen86a, Sen87, Sen90, Sen92, Sen93, She80, Shi80, Shi82, Sho79, Sho83, SF84, SI67, Sun96, Sun99, Tik81, Tom71]. Theorem [Ul’77, Ul’79a, Ul’79b, Ute91, Ven90, Vir70, Vol96a, Zai87a, Zai88a, Zal86, Zol66a, Zup92]. Theorems [Abr87a, Abr88, Akh80, Akh81, Ale77, Ale84, Ano60b, Ano63b, Ano64a, Ano65a, Ara72, BZ84, BM90, BY96, Bak96b, Bar85, BL75, BDEM92, Bel63, BN96, BG96, Ber86b, Bez71, Bin90, Bor80a, Bor84b, Bor88a, Bor60b, Bor60a, Bor61, Bor62, Bor64a, Bor64b, Bor65a, Bor65c, Bor73a, Bor78d, Bor79c, Bor79e, Bor79d, Bor80b, Bor81b, Bor85c, Bor87, Bor88b, Bor89b, BM99, Bre65, BP78, Chi57, Chi65b, Chi70a, Chi70c, Chi70d, Chi72, Chu95, Dar60, Das94, Dau88, Dav74a, Dav89a, Eag77, Ego72, EN83, Ego85, ES77, Fel90, Gam92a, Gap73, Gär76, Gir77, Gir83, GV86, Gir91, Gol85a, Gor89, GP85, Gri92a, Gri94b, Gri96a, Gri96b, HSVJ67, Ibr62, IK73a, Ibr74, IT79, IM65b, Ivc71a, IM87, Kal77b, KS80, KT59]. Theorems [KN94, Kar85, Kar86a, Kar86b, Kav91, KW71, Kha89a, Khm84, Kle86, Kli79a, Kni66, Kol56b, KS86, KS88, Kru73a, Kru76, KB97, Kud86, KK96, LO96, Lin58, Lin59b, Lin59c, Lin61a, Lin61b, Lin62, Mac71b, Mak65, Mal87b, Mal90a, Mal92a, Mal94, Mal97, Mal85b, Mal65b, Mar84, MP93a, Mas60, Mes58, MN77, Mik97, Mik99, Mik70b, Mit62, MR83b, Mol94, MPY70, MFY74, Muk85, Muk92, Nag57, Nag61, Nag65, Nag68b, Nag69a, Nag69d, Nag69b, Nag69c, Nag76a, Nag80, NK81, NK82, Nag84a, Nag94, Nak88, Ose65, Osi75, Pan87, Par87, Pav82, Pav89a, Pet59b, Pet62, Pet64, Pis82, Pis95, Pit73, Pol78, Pop68, Por67, Pro56, Pro91a, Rak89, Ric57, Ric58, Ric64, Ric65, Rog58, Rog74, Rom87, Rom62, Ron80, RS83]. Theorems [Ron88, Ros80, Ros95, Rot74, Rot76b, Roz82, Roz87, Ryz65, Ryz67, SG62, Sev57b, Sev66b, Sev68, Shl81, Shu76a, Shu67, Shu89, Shu90, Shu91b, Shu77d, Sin62, Sko56, Sko57a, Sko58, SS65b, SS66, Sla79, Sta65, Ste93, Stu67, Syt67, Sza71, Tou94, Tro91, Tro93, Tut65, Tut69b, Uri93, Vat77c, Vat82, VM83, VMS95, VT73b, VT82, Ven77a, Ven77b, Ven80, Ven83, Vir74, VR59, VR61, Vol61b, Vol86, Yan64, YK87, Yus90, Yus91a, Zol57b, Zol64b, Zol90, Zol92b, Muk96, Shu91c, Svi90, CPP98, Ibr75a, JM94, Kag88]. Theoretic [Lin59a, Lut90, Lut91, Mil79b]. Theoretical [BGS90]. Theory [Ano57b, Ano57a, Ano58d, Ano59a, Ano59e, Ano60b, Ano60c, Ano61a, Ano62d, Ano63b, Ano64c, Ano65a, Ano65e, Ano69h, Ano70f, Ano71g, Ano73h, Ano73k, Ano73j, Ano74a, Ano74j, Ano74i, Ano74g, Ano75i, Ano78j, Ano78k, Ano78f, Ano79j, Ano80j, Ano80k, Ano81a, Ano81l, Ano82a, Ano82g, Ano82p, Ano82n, Ano82o, Ano82m, Ano83e, Ano83f, Ano84l, Ano86l, Ano86j, Ano87k, Ano87n, Ano87j, Ano87m, Ano87i, Ano87l, Ano88g, Ano88k, Ano88l, Ano89a, Ano89k, Ano90b, Ano90j, Ark65, Bar77, BGW75, BGS90, Bog82, BU75, Bor64b, Bor65a, Bor65c, Bun63, Che66, Chi57, Chi65b, DP75, Den89, Den92, Dob58d, Eng71, Eng74, Fri70, Gal95, Gin66, GS63, Gol83, Gri92c, Háj61, HSVJ67, Ibr56b, Ign82, Kag90, Kal93a, KM65]. Theory [KY65, Kar69, Ken64, Ken87, Khi62, Khm82, Kho87, Kho88, Kho94, Khr96b, Kir94, Kle79, KS60, Kol62, Kol94, Kry71, Kry73c, KST60, LS60, Leo82, LRS65, LPS67, LR63, MP70, MS90, MT90, Mas60, Mol64, MTV91, MFY74, Mus73, Nag76a, Nag84a, Nat73, Nik87, NK79, OR89, Ose90, PK87, Per59, PY69, Pro56, Pro59a, Puk94, Rak88, Rak90, Rog66b, Rog74, RR64, Roz81a, Roz90a, Sel69, Sev90, Shi60, Shi67, SKKM95a, SKKM95b, Shl81, Shp66, Shu76b, Shu85c, Shu87, SS90, Stu67, Tom90, Vak90, VJK59, VJ64, Ver88a, VS77, VS78, Vil89, Vis64, VK89, Wal93, Won59, Zhu90, Zol57b, Zol57a, Zol92b, Ano71c, Ano76g, Ano88a, Gne90]. Thinning [Riv96]. Third [Ale98, Bra87]. Third-Degree [Bra87]. Third-Order [Ale98]. Thirteenth [Ano80k]. Three [Bol57, Dob65, Dob66, Dob73, Dob74, Tie80]. Three-Dimensional [Dob65, Dob66, Dob73, Dob74]. Threshold [Zue88]. Tight [Leb87, Vla87]. Time [Afa65, AM69, Ali92, Bas60, BJL98, Bor80a, Bor90b, Bor90a, Bor65b, BU94, Bra91, BP93, Cai93, CFZ95, DM68, Dar76, DB81, Dok99, Dza74, Dzh89, Dzh95, Eli89, Eng75, Fre88, Gap89a, Gär75, Gla63, GP98, GS65, GK68, Gus69, Gus84, Ibr67b, IN78, Ivc75c, Ivc76, Kab96, Kal78, Kal93a, Kal69, KS61, Kha87, Kha84c, Kif71b, Kli75, Kli79b, Kog68, Kra71, Leb71, Leo61, Lev75a, Lev80, Lip97, Lod71, Lot96, Mai79, McK63, MN77, MW99, MP78b, Nag94, Nas88, Nov82, Nov83, Nov91, Pit69b, Pre65, Rag86, Rag93, RK79, RS83, Rot67, Sal90, Sel69, Sev62, SKKM95b, Sht66, Shu79b, Sve87, Tak57, Tak58, TP59, Tsi78b, Ver91, Vin95, Vol58a, Vol61a, WR69, YF79, Yus78, YF80, Yus80]. Time [Yus91b, Zhi63, Zhu90, Zol64a, Zub73, Bar97, KY97, PS98, SKKM95a]. Time-Change [Kha84c]. Time-Continuous [Mai79, RK79]. Time-Dependent [Leb71, Leo61]. Time-Sharing [Kli75, Kli79b]. Times [Anu81, AI99, Bor82a, Bor85a, CGM97, Dar85, Dav76c, Dav79, DR79, Döh80, Dok87, Evs78, Fil84, Gol85b, Gri95, KS84, KS85a, Kuz74, Nas96a, Nas96b, Nas97, Nov71b, Per59, Rya76, Ste88a, Tsi91, Vin94, Vin95, Vin96, Yak87, Yak96, Dav76b, EO79]. Together [Sem88b]. Tomilin [Mir84c]. Tomography [ARS^{+}97]. Topological [Bul81a, GH93, Klo59, Kol99, KNS94, ST66, Shl81, Tar90b, Vla79]. Topologically [Sko87]. Topologies [Kha96, Mus81, MC84, SS93]. Topologization [Vor63b]. Topology [Gor84a, Klo64, Maz69, Reb88, Sil74]. Total [KN94, Kol88, Pro92, Top89, Vat94b]. Towns [Mor75]. Trace [Roz96]. Traces [Bar85, MO69]. Traffic [Fed69, Fed70, KK82, Pav91a, Sam63, Veĭ93, VP64].

Trajectories [AVR95, Che56, Chi83, Ibr84, Kuz80, Mog75, Mog77a, Pin81, Ven65]. Trajectory [Kor89c, Rya76]. Transfer [Rom62]. Transference [Rac85a]. Transform [Csö75, Dor90a, Vin74, Vin77]. Transformation [Che57b, Che96, IP89, Leb73, RR91, Shi64a, Sko69]. Transformations [Bol59, Bol63, CS61, Eid73, Hög93, LE68, Shu77c]. Transforming [Gir60a]. Transforms [FST95, Gap97, Zol57a, Zub99]. Transient [Afa91b, Bor95, Chi61b, Dob56c, Kor94b, Nag94, Sev59]. Transition [Bez82, Chi73b, Cow93, Dob62a, Dob65, Dob66, Ivn98, Kif77, Kru69, Kuz81, Kuz87, Mak65, Sat96, She87, Shu77b, Tol99, Ven68, Ven61, V’y78a, V’y78b, Yam95]. Transitional [AS83]. Transitions [Ste70b, Vor64]. Translated [Lin90]. Translations [Sko65, Sko70b]. Transmission [Dob58b, Dob62a, Dob62b, Glo79b, Kat75, RR64, SS65a, ZI70, ZI71, ZI74, Ano59c]. Traveller [Mor75]. Tree [AS92, CGM97, Pav78, Sev94, Ste69c, Tim89]. Trees [Sac71b, Tim85, Vat95]. Trend [KM89, KS97, Kut78, LS71]. Trial [Kal77a]. Trials [Ger73, GZ66, Mik74, ZM74, ZM80]. Triangular [Bet89, Kal77b, Sīl92]. Trimeter [Gas63]. Triple [Ano82r]. Trivial [Tsa77, Tsa79]. Truncated [Epa74a, Gri89, Gri94a, Nag98a, PU85a]. Truncation [Eri76, GS65]. Trunks [Bas60]. Tube [Kle71]. Tubular [KY84]. Tuple [IN78]. Tuples [Bel65, KC75a, ZM80]. Turbulence [Mon58, Ose90]. Turing [Fre88]. Twelfth [Ano90j, NK79]. Twenty [Ano89k]. Twenty-First [Ano89k]. Two [Ale63, Ano90g, AB93, Ark71, Bal96, Bas58, Bel67a, BR64, Che70, Chi89a, DM68, Dar85, Dob65, Dob66, Epa68, FF82, Fri69, GP79, HSVJ67, IR70, Kag88, Kah93, KK82, Kha89b, KR75, Kol56b, KNS94, Kor92, Lin64, Lot80a, Lot80b, Mak82a, Mar69, Mik58, Mir78c, Mir80b, Mis89, Nik91, NV85, Nyb86, Pag68, PRR82, PP96, Par70, Pis95, Pol63, Pop68, PY81, PS92a, Pre82, Rih85, Ron84, Sad66b, Sad67b, Sam82, SR79, ST73, Shu77d, Sid91, Sir73, Tik85, Tou94, VS90, Vis64, Vol67, Vol69, Vol73, Vol74, Yan70, Yus91b, Zol65b, ZS76, Zub83, Zub99, Pre91, VS89]. Two-Armed [Yus91b, Pre91]. Two-Component [Che70]. Two-Dimensional [Ark71, BR64, Dob65, Dob66, Kag88, KNS94, Kor92, PRR82, PY81, Sad66b, Sad67b]. Two-Mode [ST73]. Two-Parameter [Kah93, Mis89]. Two-Person [Fri69, Yan70]. Two-Phase [KK82]. Two-Sample [AB93, Kha89b, Mar69, Vol74]. Two-Sided [DM68, Epa68, PS92a, Tik85, ZS76, Zub99]. Two-Stage [Bas58]. Two-Term [Bal96]. Type [BM91, Bor77a, Chi71a, Chi70d, Chi86b, Chu92a, Dar96, Die77, Dor96a, Ers72a, Fil84, GE91, IS68, Ivk86, Jan93, Kab78, KS80, Kha61b, Kho88, Kol88, Kon74b, KS88, Kra94, KS89, Küh81, Kul98, Lot96, Mak91, Mar76b, Mus80a, NR73, NR76, Ose90, Pan95, Pet56a, Pis80, Pis82, Pis95, Pis98, Pra93, Rom61, Rud93, Sag91, Sel69, She88a, SZ98, Tri90, Tut69b, Ula88, Vat87a, VMS95, Vor61, V’y78b, ZK77, Zol68b, Zol77b, Chu94, Ome97, Szy97, Zol78d, IL94]. Types [Bas65, Chi73b, Erm96, Gri87, Kha68, Kha69a, Oli93, Pan94, Pop68, Rih85, RS70, Sag83, Sag86, SG62, Sch79, SKKM95a, SKKM95b, Vat77d, Vat79, Vat80, Vat81, VS90, Zin71, Zub83, VS89, Vat95]. Typical [Asa91].

U [MA78, MA79]. U-Statistics [MA78, MA79]. U.S.S.R [Ano78k]. U.S.S.R. [Ano73h]. Ul′yanov [Ul′91]. Ultra [Roz94a]. Unbiased [AL81, JP75, KP67, Kle75, KR76, LS69, MR96, PU85a, Sap79, Sed77, SS88b, Sha70, WW92]. Unbiasedness [Gur89a, Kle77].

Unbounded [DJ99, Kif69, Lin59c, Por75a, Rad98, Roz94a, SK93]. Unboundedness [Bel58]. Uncorrelated [Sar67]. Undergo [Fil61]. Underlying [YK86, YK87]. Unequal [Bol68, Vin99]. Unidentifiable [PK82]. Unified [LL94, MP93b, dH95]. Uniform [Ale74b, Ale89, Ara82a, Ara82b, Ara83, BDEM92, Bor91c, Gil82, Gor84a, JR89, Kai86, Kal78, Kar81, Kha66a, Kha79a, Kol67, Kol56b, Kor94a, MU83, Mal85b, Mol88, Nov80, Pas99, Pre74, Pro60, Rot70a, Ryc84, Sen81, Sen82, She71, Tsa58b, Tsi76b, Ul’77, VC71, VC82, ZA84, Zai85, Zai88b, Zaĭ90, Kha79b]. Uniformity [Ber85, Rac86, Top67]. Uniformly [Ark64b, Sap79, She99, Zai92]. Unimodal [BD97b, Bog80, DJD86, Ibr56a, Sat95b, Ush82, Vin76, VP83, Yam96]. Unimodality [IC59]. Unimodular [Shu96]. Unimprovable [Bor96]. Union [Ano63b, Ano64c, Ano80k, Ano81b, Ano81a, Ano82a, Ano89a, NK79, Pro83, Ano59a]. Unions [Mol94]. Unique [Jes81, Ula88]. Uniquely [KU88b]. Uniqueness [Bak96e, Bla83, DV96, Gal79, Gir62, KD87, Kle86, Kul74, Leb77b, Leb78, Leb86, Lei82, SR79, Sie80, Tar90a]. Unit [Dyu68, Mac69, Yan64]. Unitary [HR97]. Units [Mal91, SK93]. Univariate [MR93]. University [Ano59e, Ano60c, Ano78g, Ano78j, Ano79g, Ano79h, Ano80e, Ano80h, Ano80g, Ano81k, Ano82o, Ano82i, Ano83e, Ano83d, Ano84k, Ano85j, Ano85l, Ano85m, Ano86j, Ano87i, Ano88k, Pet59c, Ano61d, Ano73j, Ano74i, Ano75i, Ano86l, Gne90, MS90, MT90, Sev90, SS90, Zhu90]. Unknown [Che67a, Dzh75, Efr86, Lum69, Pav81a, Zez75, Mak88]. Unknowns [Bal96]. Unreliable [BP73]. unstable [PvZ97]. Updating [Pët93]. Upon [Cow93]. Upper [Alb93, Ano85a, EO97, Kor84, Rog68b, Ush86, Vve85a, Vve89, Zig82, Lep92]. Urn [DF93, Ron89]. Use [Che67b, Gri92c, Khm80, Mir84b, WS93].

Used [Shi95]. Using [BR94, Del96, GK81, Kha84b, KP97, Koz81, PS75, Tho93a, GK80]. USSR [Ano62e, Ano63c, Ano64d, Ano65d, Ano67f, Ano71c, Ano73k, Ano75k, Ano76f, Ano77h, Ano78i, Ano78h, Ano79f, Ano79i, Ano79j, Ano80f, Ano80i, Ano80j, Ano81h, Ano81i, Ano81j, Ano81l, Ano82q, Ano82g, Ano82p, Ano82l, Ano82n, Ano82m, Ano82h, Ano82j, Ano82k, Ano83c, Ano83f, Ano84h, Ano84j, Ano84l, Ano84i, Ano85q, Ano85p, Ano85k, Ano85n, Ano85o, Ano85r, Ano87k, Ano87n, Ano87j, Ano87m, Ano87l, Ano88l, Leo82, Ano68f, Ano69h, Ano71g, Ano74j, Ano75a, Ano76g, Ano86k]. Utev [FM92].

V [Ano65d, Ano65e, Ano81j, Ano82p, Ano83f, Ano85q, Ano85o, Ano85r, Ano87n, Ano87m, Ano87o, Ano87l, Chi87a, Chi70a, GKPS60, Kol62, Mar87a, Ost65, Ost67, Ul′91, ZK61]. V. [Chi70a, Hoe64]. Vague [Rac86, Sie87]. Validation [Mye93]. Value [Ara99, Bon92, BR64, Bor67a, Dok92, Epa74a, Fro89, Gin89b, Gus69, IK85, Kog65, Kog66, Mil96, OR89, Rot67, Roz91, Roz67, Sch89, Vil63, Vil68, Wal93, dH95, Mil97]. Valued [AV82, BGW75, Chu95, DEO84, EO86, Gam82, Gam86, Gia96, Gil82, Haz95, Kor94b, Leb96b, MW59, McD80, MW99, MP91, Nad73b, PS92b, Rao63b, Reb88, Sak85, Saz92, Su99, Thu85, Tik92a, Tik94, TM79, VK67]. Values [Bla83, BW98, Cra65, EM68, Fäh83, Ibr78, KS64, KN90, Kru71b, Kru73a, Kru74, Mak67b, Mak70b, Mal90b, Met75, Pan95, Pin79, Pro84, Pro93, Rao71, RS98, RU88, SZ98, Tie77, Tit82, VK97, Vak99, VIL69, Vol60b, Vve84, Vve85b, Zol77a, Zol78a, Met74, Roz97]. Vanishing [VL76]. Varadhan [Kul98]. Variable [Ale92a, Bab64, DK93, Ero58, Fed77, Kha89b, Maĭ95, Mak80, Mit88, Nas88, Pet76, She99]. Variable-Stress [DK93]. Variables [Ale77, Ale79, Ale80, Amb76, Amo80, AT86, Ant80, AN74, Ara75, Ark90, AB60, AV82, Bag88, Bag89a, Bag89b, Bak98, Bak90, Bak96c, Bas59, BGW75, Ber70a, Ber70b, Bet88, Bet89, Bik69b, Bol59, Bor84b, Bor60b, Bor60a, Bor68b, Bor72, Bor73b, Bor96, Bor85d, Bor86b, Bra90, Bul74b, Bul77, Bul75, Bul84, Bul94, Bul96, Chi64d, CPU96, Chu92b, Dar76, DEO84, Dob70, Dzh77, EM68, Ego70, Ego91c, Eri76, EO86, Evs87, FLR97, FN71, Gam64, GP71, Gam74, Gam82, Gap89b, Gap73, Gia96, Gik58, Gil82, God79, Gor71, Gri91a, Gri92a, Gri94b, Gri96a, Gri74, Gur92, Gus76a, Haz95, Ibr63a, Ibr66, Ibr67a, IP74, Ibr75a, Ios68, Ios70, Ivc74b, JM94, Kag89, Kag90, Kar95, Kar65, Kha88, Kha61a]. Variables [KMM85, Kle87, Kle96, Kol56b, Kor86b, Kor89b, Kor89a, Kor90a, Kor98, KR90, Kru71b, Kru73a, Kru74, Kru75, Kru80, Kru99, Kry68, Kud85, Kur85, Kur88, LM71, Lif85a, Lin61a, Lin61b, Lin62, Mak82a, Mak67b, MA83, Mar85, MP93a, Mar76a, McD80, Mik65, Mik74, Mik75c, Mik89, MR83a, MPY70, MFY74, Muk90, Muk92, Muk98, Mus69, Mus70, Nag70a, Nag70d, Nag77, NE77, Nag78, NP78, Nag98a, Nak88, Nas57, NP69, Nev71, Nev72, Nev73b, Nov92, Nov97a, Nov82, Nov83, Nyb86, Osi69, Osi71, Osi73, Osi74, Pak59, Pan87, PR99, PR69a, Pec74b, Pet56b, Pet59a, Pet59b, Pet64, Pet65, Pet70, Pet74b, PL83, Pin79, PU85b, PU90, Pon79a, Pon79b, Pon86, Pre67, Pre86, Pro94, Rag91, Rao63b, RS98, Ric65, Rog61b, Rog66c]. Variables [Ron80, Ron88, Rot76c, Rot76a, Roz90c, Roz98b, Rus96, RS76b, Sak85, Sal90, Sal71, Sar67, Saz75, Sch89, SS62, Sev73, Sha96, She80, Shn87, Sho96, Shu76a, Sin57, Sko77, Sla82, Sla85, Smo85, Sre70, Sta65, SU90b, Sza73, Szu83, Ten91, Tik81, Tik87, Tik92a, Tik94, Tri90, Tsu81, Ul’87, Ute91, VIL69, Vas80, Vin76, VG90, Vin90, Vol58b, Vol60b, Whi60, Whi64, WR69, Yak98, Zai84, Zai87a, Zai88a, Zai88b, Zas73, Zhu66, Zin86, Zin65, Zin71, Zol65b, Zol77a, Zol78a, Zol78c, Zol92a, ZM74, Zup92, Bak89b, DRS89, Ego97, Gap95, NK97, PR97, Sho97]. Variables-Which [Zai84]. Variance [Ano61b, Bor82a, Bor85a, CPP98, DKR95, Eli85, JP75, JR91, KP67, KNS66, Kra69, MF71, Mas74b, MN99, Oni90, Rak87, Sgi98, Sha70, Vat77c, Vol74, BO97, MP97]. Variances [Gor71, Pag83]. Variant [Bor92b, Bul81b, Raz70, Tut78]. Variate [MS79, Sch79]. Variation [Bla83, DW79, Gam84, Kab86, PP96, Pre86, Pro92, Tut68a, Zai92]. Variational [Ivc74b, Kud74, Kud84, Nag71a]. Variogram [Mye93]. Various [Dor90a]. Varying [DB93, Kol67, RS70, Tik98]. Varying-Environment [DB93]. Vasil’evich [Ano66d, Ano87p]. Vavilov [Nik66]. Vector [BNS66, BGW75, Bel68, Gil82, Kat78, Kat75, Kor66, Kor68, Lin66, MV89, Mik75b, MW99, Mit67, MP91, Nad73b, Por80, She88a, SL90, Smo97, Sol67, Su99, Th93, TM79, VGR84, Bau97]. Vector-Valued [BGW75, Gil82, MW99, MP91, Nad73b, Su99, TM79]. Vectors [Are78, Bak96b, Bar86, BGZ98, BM99, BS88, BS92, Gam86, IR70, Kag90, Kha94, Lar94, LE68, Lis94, Mih56, Mik75a, Muk85, Osi79, Pin91a, PY81, PY88, Pro73, Pro84, Pro91a, Pus92, Sau92, Sem88b, Sid91, Ush86, VK97, Vak99, Vve84, Vve85b, Vve89, ZI71, ZI89]. Verifiability [Leh65]. Verification [Ber86a, Pet56a, Yus90, Yus91a]. Verifying [Ber81, Ing86]. Veroyatnost [Bak89b, Fry90, Shu91c, SZ98, Ul′91]. vers [Tou92]. Verse [Bob64]. Version [Amo88b, Cad99, KK75, Pre74, Rie77, SI67, Yur70, Zaĭ90, Pre91, Vov97]. Versions [JR82, Mar84, Tsi76b]. Versus [CFZ95]. Vertices [Ivc73, Lis76, Pav89b, Pro74, Ste69c, Vat94a]. Very [Mna86]. VI [Ano84g]. via [PP96]. View [CS71]. VII [Ano84g]. VIII [Ano86m]. Vilnius [Ano78f]. Vitautas [KGPS91]. Vladimirovich [Ano73l, DKPF67, V^{+}97]. Vladimirovitch [Shi93]. Volatilities [DKR95]. Volterra [KV85, Kle97]. Volume [Bor65a]. Voting [Zue98]. Vs [AB99]. Vyacheslav [Ano87p].

W. [Rog69, Sho97]. Waiting

[DM68, Fak82, GP98, Ivc75c, Ivc76, Pre65, Tak57, Vin95, Bar97, Lab73, Lab79]. Wald [Aiv59, KM80, Lin64, LRS65, Lot89, Rom69]. Walk [Afa79, Afa87, Afa91a, Afa91b, Arn81, Big93, Bor84b, Bor62, Bra91, Den84, Fal78, GR84, GS88, Kab96, KNS94, Kor61b, Kor62, Koz74a, Lët89, Lum71, Mal65b, Mar81, Mik70b, MP78b, MS86, Nag87, Nag94, Par66, Pec74a, Pec81, Pit92, Pit75, Rog71, RF78, Sgi99, Shu92, Sin83, Sin94, Tut68b, Vat96, Vin94, ZM95]. Walks [AB85, Bor82a, Bor85a, BR64, Bor67a, Bor79c, Bor79d, Bor81b, Bur82, But95, Gol85a, Gri74, JP97, Kha77, Koc98, Kor90b, Lët93, Lot92, Mog77a, Mór93, Rog73, Rog77b, Rom87, Rom61, Shu89, Shu90, Shu91a, Shu91b, Shu96, Shu81, Shu82b, SS65b, SS66, Sob73a, Syt67, TV98, VM91, Shu91c]. Walsh [KF91]. Wanderings [Dyn71, Tak77]. Wasserstein [Val73, Val74b, Val82]. Watson [AS83, BY96, DB93, KN94, KNS66, Mak81, NV75, NK81, NK82, Rak89, TV98, Vat78, Vin88, Zub75]. Wave [For63, ZI70]. Waveguide [ST73]. Waveguides [GV59, Rez77b, Tut69a, Tut71]. Way [Bra88, Ego85]. Weak [Bau88, Bau96, BJL98, Bez82, Bha70, Che56, DS85, De 82, Ers79, Fre67, GH93, Gor84a, GM84, HP94, KLS84, Khm81, KB92, KR90, Kru86, Kru99, Kud98, Leb82, Mil86, Muk77, Mus80a, Nak81, Nik80b, Nos83b, PN92, Pit71b, Rac86, Roz90b, Sar90, Sil79, Ste93, Ste94, Su99, Szy97, Top67, Ver82]. Weakening [Reb88]. Weakly [Bon98, Bor79b, Bor67b, DEO84, EO86, Gap70, Gap86, Gri92a, Gri94b, Gri96a, Ing89, Kha96, Kha93, Kha94, Lif85a, Man64a, Nad73b, Tik81, Tik87, Tik92a, Tik94, Zup92]. Weibull [Kot64, Vol75]. Weighing [BG73]. Weighted [BS87, Gap68, Gur89b, Kle87, Kud78, MF71, Mar79, Nor98, Shu76a, SO85]. Weights [LS71]. Weiss [DN88, Pav91b]. Welch [LRS65]. Well [BRR98]. Western [Bob64]. where [Vin99]. Which [Bas58, Kop96, Roz61, SE96, Shu64, Sko57b, Zai84, Dob56c, Ego73, Eid73, Fil61, Gol65, Sed77, SS65a, Shu91b, Ven65, Ven86, Vla80]. White [Bur79, But77, Che57a, Erm91, GK81, GK84b, IK85, INK87, Lep91, Nem91, GK80]. Whitney [Bul85b, Gel74]. Whittle [Cow93]. Whole [Roz94b]. Whose [Che56, Pro91a, Tut67, Urb56, Zai92]. Wide [Mkr81, Sid57, Ver66, Zak90b]. Wide-Sense [Mkr81, Sid57]. Wiener [JK90, Roz66b, Shi64a, Anu80, Anu81, BP89b, BS68a, BJL98, Bor64a, Bor67b, BVT90, Che57b, Dav91, DM96, Den84, Din93, DK93, Dor90a, Dyn60, Fäh78, GZ69, IS74, Ivk86, Kab74, Khr94, KS87, KPĬ93, Kru86, Man67, Mir76, Mir78c, Mna88b, Nov71b, Roz66a, Saf77, Sap92, VTGM85, Ven65, VM89, Vos82]. Wiener-Itô [DM96]. Within [Lep94, VS91, Anu91]. Without [Lep93, BB61, Gur89a, IL78, Khi56b, KK96, Lan88, Mir86b, SS65a, Sko67, VG90]. Wold [Rob63]. Wolf [Ano98a]. Wong [Mir77]. Work [Ano65e, Bin90, GS63, Ken64, KS60, Kol62]. Works [GKPS60]. World [Ano84d, Ano88a, Ano93a].

XVI [Ano93d]. XVIIIth [Ano98d]. XX [Ano88g]. XXII [JK90].

Y. [JK90]. Yakovlevich [Gne60]. Yu. [Ano65e, Chi70a, Mar87a, Ost65, Ost67]. Yule [XP89]. Yurii [Ano73l].

Z [JK90]. Zalesskiĭ [Ul′91]. Zero [CKR68, Gol65, IM71a, IM71b, Kle74b, Mal75, Mal69a, Nag68a, Rih85, Sen87, Yan70]. Zero-One [Rih85]. Zero-Sum [Yan70]. Zero-Zero [CKR68]. Zeroes [Dha78]. Zeros [Gin89a, IM71a, IM71b, Il’76a, Mal80, Mal85a, MV89, MV94, Mir74, Mir75, Mir78a, Mir85, Ost89]. Zolotarev [Hoe64, KMM85]. Zone [Roz82]. Zones [Amo91a, Sla85]. Zygmund [Gia96].

References

Adamyan:1968:GSP

[AA68] V. M. Adamyan and D. Z. Arov. A general solution of a problem in linear prediction of stationary processes. Theory of Probability and its Applications, 13 (3):394–407, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arov:1960:ETS

[AB60] D. Z. Arov and A. A. Bobrov. The extreme terms of a sample and their role in the sum of independent variables. Theory of Probability and its Applications, 5(4):377–396, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:1985:CAR

[AB85] L. G. Afanas’eva and E. V. Bulinskaya. Certain asymptotic results for random walks in a strip. Theory of Probability and its Applications, 29(4):677–693, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arghami:1993:SSA

[AB93] N. R. Arghami and L. Billard.

Some sequential analogues of Stein’s two-sample test. Theory of Probability and its Applications, 37(1):115–117, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 141–144.

Albeverio:1999:LVD

[AB99] S. Albeverio and L. V. Bogachev. Localization vs. delocalization of random discrete measures. Theory of Probability and its Applications, 43(4): 519–538, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97715.

Abdushukurov:1999:NER

[Abd99] A. A. Abdushukurov. On nonparametric estimation of reliability indices. Theory of Probability and its Applications, 43 (1):116–122, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97670.

Abramov:1959:EAS

[Abr59] L. M. Abramov. The entropy of an automorphism of a solenoidal group. Theory of Probability and its Applications, 4(3):231–236, ???? 1959. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Abramov:1977:ELP

[Abr77] V. A. Abramov. Estimates for the Lévy–Prokhorov distance. Theory of Probability and its Applications, 21(2):396–400, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Abramov:1987:LTM

[Abr87a] V. A. Abramov. Limit theorems in the multiplicative scheme. Theory of Probability and its Applications, 31(4):633–644, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Abramov:1987:SDS

[Abr87b] V. A. Abramov. Stability of decompositions in the Mscheme. Theory of Probability and its Applications, 31(1): 123–124, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Abramov:1988:GLT

[Abr88] V. A. Abramov. General limit theorems in the M-scheme. Theory of Probability and its Applications, 32(3):501–503, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:1965:ELD

[Afa65] L. G. Afanas’eva. Existence of a limit distribution in queueing systems with bounded sojourn time. Theory of Probability and its Applications, 10(3):515–522, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasev:1979:CRW

[Afa79] V. I. Afanas’ev. A conditional random walk with a negative drift. Theory of Probability and its Applications, 24(1):192–199, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasev:1987:FRW

[Afa87] V. I. Afanas’ev. On functionals of a random walk prior to first reaching the negative halfaxis. Theory of Probability and its Applications, 31(4):683–687, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:1988:EOQ

[Afa88] L. G. Afanas’eva. On the ergodicity of an open queueing network. Theory of Probability and its Applications, 32(4):710–714, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasev:1991:CLT

[Afa91a] V. I. Afanas′ev. A conditional limit theorem for additive functionals of a random walk. Theory of Probability and its Applications, 35(2):330–336, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 331–337.

Afanasev:1991:MTR

[Afa91b] V. I. Afanas′ev. On a maximum of a transient random walk in random environment. Theory of Probability and its Applications, 35(2):205–215, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 209–219.

Agasandyan:1980:ASH

[Aga80] G. A. Agasandyan. Adaptive systems for homogeneous processes with continuous sets of states and control parameters. Theory of Probability and its Applications, 24(3):521–535, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Al-Hussaini:1985:SJP

[AHE85] A. Al-Hussaini and R. J. Elliott. The single jump process with some statistical applications. Theory of Probability and its Applications, 29 (3):607–613, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aspandiiarov:1999:GCI

[AI99] S. Aspandiiarov and R. Iasnogorodski. General criteria of integrability of functions of passagetimes for nonnegative stochastic processes and their applications. Theory of Probability and its Applications, 43(3):

343–369, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97703.

Aivazian:1959:COP

[Aiv59] S. A. Aivazian. Comparison of the optimal properties of the Wald and Neyman–Pearson criteria. Theory of Probability and its Applications, 4(1):83–89, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aivazyan:1965:DNH

[Aiv65] S. A. Aivazyan. Distinguishing neighboring hypotheses about frequency distributions in the scheme of the generalized probability ratio test. Theory of Probability and its Applications, 10 (4):646–658, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Artamonovskii:1970:EML

[AK70] V. P. Artamonovskii and Kh. B.

Kordonskii. Estimate of maximum likelihood for simplest grouping of data. Theory of Probability and its Applications, 15 (1):128–132, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aliushina:1989:PLI

[AK89] L. A. Aliushina and N. V. Krylov.

Passage to the limit in Itô stochastic equations. Theory of Probability and its Applications, 33(1):1–10, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 3–13.

Alekperov:1995:CCC

[AK95] Sh. O. Alekperov and V. M. Kruglov. Convergence of convolutions of concentration functions to degenerate, normal, and Poisson laws. Theory of Probability and its Applications, 39 (2):197–216, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 248–271.

Akhmarov:1980:RCE

[Akh80] I. Akhmarov. On the rate of convergence in ergodicity and continuity theorems for multichannel queuing systems. Theory of Probability and its Applications, 24(2):423–430, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Akhmarov:1981:CRE

[Akh81] I. Akhmarov. On the convergence rate in ergodicity and continuity theorems for systems with refusals. Theory of Probability and its Applications, 26(1):177–184, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Akhmarov:1977:CCL

[AL77] I. Akhmarov and N. P. Leont’eva.

Conditions for convergence to limit processes and the strong law of large numbers for queueing systems. Theory of Probability and its Applications, 21 (3):545–556, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Abusev:1981:UEC

[AL81] R. A. Abusev and Ya. P.

Lumel’skii. Unbiased estimators and classification problems for multivariate normal populations. Theory of Probability and its Applications, 25(2):377–384, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anulova:1991:DAP

[AL91] S. V. Anulova and B. Sh. Liptser.

Diffusional approximation for processes with the normal reflection. Theory of Probability and its Applications, 35(3): 411–423, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 417–430.

Albin:1993:ULC

[Alb93] J. M. P. Albin. Upper and lower classes for Gaussian random fields and the general LIL. Theory of Probability and its Applications, 37(1):3–6, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 5–10.

Alekseev:1963:CPG

[Ale63] V. G. Alekseev. On conditions for the perpendicularity of Gaussian measures corresponding to two stochastic processes. Theory of Probability and its Applications, 8(3):286–290, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1964:SME

[Ale64] V. G. Alekseev. Some methods for the exact determination of a parameter of a Gaussian stochastic process. Theory of Probability and its Applications, 9(3):466–469, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1967:APS

[Ale67] V. G. Alekseev. On the asymptotic properties of some statistical estimates for Gaussian stochastic processes. Theory of Probability and its Applications, 12(1):1–8, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1970:EPG

[Ale70] V. G. Alekseev. On the estimation of parameters of a Gaussian random process. Theory of Probability and its Applications, 15(1):122–128, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1974:ESA

[Ale74a] V. G. Alekseev. Empirical spectral analysis of homogeneous and isotopic random fields. Theory of Probability and its Applications, 18(2):272–280, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1974:UCE

[Ale74b] V. G. Alekseev. On the uniform convergence of estimates of the spectral density of a stationary Gaussian random process. Theory of Probability and its Applications, 19(1):193–200, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aleshkyavichene:1977:LLT

[Ale77] A. D. Aleshkyavichene. Local limit theorems for the density function of the maximum of a sum of independent variables. Theory of Probability and its Applications, 21(3):449–469, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aleshkyavichene:1979:PLD

[Ale79] A. K. Aleshkyavichene. On the probabilities of large deviations for the maximum of sums of independent random variables. Theory of Probability and its Applications, 24(1):16–33, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aleshkyavichene:1980:PLD

[Ale80] A. K. Aleshkyavichene. On the probabilities of large deviations for the maximum of sums of independent random variables. II. Theory of Probability and its Applications, 24(2):322–337, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1981:ESF

[Ale81] V. G. Alekseev. On estimation of some functionals of spectral density of Gaussian random processes. Theory of Probability and its Applications, 25(2):267–273, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aleshkyavichene:1984:MIL

[Ale84] A. K. Aleshkyavichene. Multidimensional integral limit theorems for large deviation probabilities. Theory of Probability and its Applications, 28(1): 65–88, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1989:UCE

[Ale89] V. G. Alekseev. On the uniform convergence of estimates of the spectral density of a homogeneous Gaussian random field. Theory of Probability and its Applications, 33(4):701–705, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:1991:CPS

[Ale91a]  V. G. Alekseev.  On cer-

tain properties of statistical estimates of higher-order spectral densities. Theory of Probability and its Applications, 35(3): 424–430, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 431–437.

Aleshkevichiene:1991:PLD

[Ale91b] A. K. Aleshkevichiene. Probabilities of large deviations for U-statistics and von Mises functionals. Theory of Probability and its Applications, 35(1):1–14, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 3–14.

Aleksandrov:1992:CPR

[Ale92a] E. L. Aleksandrov. On the completeness of powers of a random variable in Hilbert space L^{2}(Ω, F, P). Theory of Probability and its Applications, 36 (2):344–349, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 337–342.

Alekseev:1992:SPS

[Ale92b] V. G. Alekseev. Some problems of statistical estimation of the regression curve. Theory of Probability and its Applications, 36(2):209–216, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 209–216.

Alekseev:1994:SPH

[Ale94] V. G. Alekseev. Symmetry properties of higher-order spectral densities for multidimensional stationary and periodically nonstationary stochastic processes. Theory of Probability and its Applications, 38(4): 722–726, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 870–874.

Alekseev:1996:APH

[Ale96] V. G. Alekseev. Asymptotic properties of higher-order periodograms. Theory of Probability and its Applications, 40 (3):409–419, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 481–494.

Alekseev:1998:NMS

[Ale98] V. G. Alekseev. New modifications of second- and third-order periodograms. Theory of Probability and its Applications, 42(4): 559–567, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97643.

Aliev:1987:LBC

[Ali87] F. A. Aliev. A lower bound for the convergence rate in the Central Limit Theorem in Hilbert space. Theory of Probability and its Applications, 31(4):730–733, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aliev:1990:CRC

[Ali90] F. A. Aliev. On the convergence rate in the Central Limit Theorem in Hilbert space. Theory of Probability and its Applications, 34(2):361–363, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 407–409.

Alimov:1992:SPI

[Ali92]  D. Alimov.  Semicontinu-

ous processes with independent increments with a boundary and a “Homographic” time change. Theory of Probability and its Applications, 36(3): 557–564, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 542–548.

Alimov:1995:AEM

[Ali95a] D. Alimov. Asymptotically ergodic Markov functionals of an ergodic process. Theory of Probability and its Applications, 39 (4):537–546, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 657–668.

Alimov:1995:MFE

[Ali95b] D. Alimov. Markov functionals of an ergodic Markov process. Theory of Probability and its Applications, 39(3):504–512, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 618–626.

Aliushina:1996:CDP

[Ali96] L. A. Aliushina. On controlled diffusion processes in a bounded domain. Theory of Probability and its Applications, 40(3): 533–538, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 632–638.

Alyushina:1988:EPL

[Aly88a] L. A. Alyushina. Euler polygonal lines for Itô equations with monotone coefficients. Theory of Probability and its Applications, 32(2):340–345, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alyushina:1988:PLS

[Aly88b] L. A. Alyushina. Passage to the limit in stochastic Itô equations with monotone coefficients. Theory of Probability and its Applications, 32(4):741–744, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:1967:EPQ

[AM67] L. G. Afanas’eva and A. V. Martynov. On ergodic properties of queues with constraints. Theory of Probability and its Applications, 12(1):104–109, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:1969:EPQ

[AM69] L. G. Afanas’eva and A. V. Martynov. Ergodic properties of queueing systems with bounded sojourn time. Theory of Probability and its Applications, 14 (1):105–114, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ambartsumian:1965:ACB

[Amb65] R. V. Ambartsumian. On an application of the connection between Brownian motion and the Dirichlet problem. Theory of Probability and its Applications, 10(3):490–493, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ambartsumyan:1974:RFS

[Amb74] R. V. Ambartsumyan. On random fields of segments and random mosaics on the plane. Theory of Probability and its Applications, 18(3):486–498, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ambrosimov:1976:ANS

[Amb76] A. S. Ambrosimov. Asymptotic normality of sums of dependent random variables considered by B. A. Sevastyanov. Theory of Probability and its Applications, 21(1):183–187, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ambrosimov:1979:ABM

[Amb79] A. S. Ambrosimov. On the asymptotic behavior of the maximum in a homogeneous Markov chain with a large number of states. Theory of Probability and its Applications, 23(2):419–426, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Amosova:1980:MDP

[Amo80] N. N. Amosova. On moderate deviation probabilities for sums of independent random variables. Theory of Probability and its Applications, 24(4): 856–863, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Amosova:1988:IPL

[Amo88a] N. N. Amosova. An inequality for probabilities of large deviations. Theory of Probability and its Applications, 32(2):337–340, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Amosova:1988:LVK

[Amo88b] N. N. Amosova. On a local version of Kolmogorov’s problem. Theory of Probability and its Applications, 32(1):175–178, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Amosova:1991:NZL

[Amo91a] N. N. Amosova. Narrow zones of local normal attraction. Theory of Probability and its Applications, 35(1):140–145, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 138–143.

Amosova:1991:NLL

[Amo91b] N. N. Amosova. A note on the local limit theorem for large deviations. Theory of Probability and its Applications, 35(4):758–760, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 754–756.

Arak:1974:SEM

[AN74] T. V. Arak and V. B. Nevzorov.

Some estimates for the maximum cumulative sum of independent random variables. Theory of Probability and its Applications, 18(2):384–387, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Azlarov:1993:ABD

[AN93] T. A. Azlarov and A. A. Narkhuzhaev. On asymptotic behavior of distributions of some statistics for Markov chains. Theory of Probability and its Applications, 37(1):117–119, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 139–141.

Anonymous:1957:SPPb

[Ano57a] Anonymous. Summary of papers presented at the sessions of the Scientific Research Seminar on Probability Theory, (Moscow, February–May, 1957). Theory of Probability and its Applications, 2(4):470–480, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1957:SPPa

[Ano57b] Anonymous. A summary of papers presented at the sessions of the Scientific Research Seminar on the Theory of Probability (September–December 1956). Theory of Probability and its Applications, 2(1):129–137, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1958:RBa

[Ano58a] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 3(1): 106–110, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1958:RBb

[Ano58b] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 3(3): 333–334, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1958:RBc

[Ano58c] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 3(4): 438–440, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1958:SPP

[Ano58d] Anonymous. A summary of papers presented at the seminar on the theory of probability. Theory of Probability and its Applications, 3(2):197–201, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1959:AUC

[Ano59a] Anonymous. The All-Union Conference on the Theory of Probability and Mathematical Statistics. Theory of Probability and its Applications, 4(1):109–112, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1959:CAM

[Ano59b] Anonymous. Conference on the application of mathematical methods in biology. Theory of Probability and its Applications, 4(1):107–109, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1959:CTI

[Ano59c] Anonymous. Corrections “Transmission of information in channels with reversed connections”. Theory of Probability and its Applications, 4(2):230, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1959:CB

[Ano59d] Anonymous. Criticism and bibliography. Theory of Probability and its Applications, 4(2):229, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1959:SPP

[Ano59e] Anonymous. Summary of papers presented at the sessions of the Seminar on Probability Theory, Moscow University (September-December 1958). Theory of Probability and its Applications, 4 (2):218–228, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1960:ISC

[Ano60a] Anonymous. Information on the second conference on the application of mathematical methods in biology. Theory of Probability and its Applications, 5 (1):122–123, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1960:RCL

[Ano60b] Anonymous. Report on the conference on limit theorems of probability theory. Theory of Probability and its Applications, 5 (2):228–238, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1960:SPP

[Ano60c] Anonymous. Summary of papers presented at sessions of the Seminar of the Department of Probability Theory of Moscow University (October-December 1959). Theory of Probability and its Applications, 5(3):332–340, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:ISC

[Ano61a] Anonymous. Information on the sixth conference on the theory of probability and mathematical statistics. Theory of Probability and its Applications, 6 (1):126–130, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:MEO

[Ano61b] Anonymous. Missing and extra observations in analysis of variance. Theory of Probability and its Applications, 6(2):227, ???? 1961. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:RB

[Ano61c] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 6(2): 227–231, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:SPPa

[Ano61d] Anonymous. Summary of papers presented at the probability research seminar, Moscow University. Theory of Probability and its Applications, 6(3):329–334, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:SPPb

[Ano61e] Anonymous. Summary of papers presented at the sessions of the Probability and Statistics Section of the Moscow Mathematical Society (November-December, 1960). Theory of Probability and its Applications, 6(3):334–343, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:SR

[Ano61f] Anonymous. A summary of reports. Theory of Probability and its Applications, 6(2):220–227, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1961:SRP

[Ano61g] Anonymous. Summary of reports presented at meetings of the Seminar on Mathematical Statistics Mathematical Statistics at the Mathematical Institute of the Academy of Sciences, SSSR, 1960. Theory of Probability and its Applications, 6 (4):439–444, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1962:E

[Ano62a] Anonymous. Editorial. Theory of Probability and its Applications, 7(1):1–2, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1962:RB

[Ano62b] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 7(2): 227–235, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1962:SPPa

[Ano62c] Anonymous. Summary of papers presented a the sessions of the Probability and Statistics Section of the Moscow Mathematical Society (December 1960– May 1961). Theory of Probability and its Applications, 7 (2):213–227, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1962:SPPb

[Ano62d] Anonymous. Summary of papers presented at the sessions of the Probability Theory and Mathematical Statistics Section of the Moscow Mathematical Society (October 1961–April 1962). Theory of Probability and its Applications, 7(4):459–466, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1962:SRP

[Ano62e] Anonymous. Summary of reports presented at the Seminar on Mathematical Statistics in the Mathematical Institute of the Academy of Sciences of the USSR, 1961. Theory of Probability and its Applications, 7 (3):341–348, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1963:LE

[Ano63a] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 8(3):334, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1963:RAU

[Ano63b] Anonymous. Report on the allunion colloquium on limit theorems of probability theory. Theory of Probability and its Applications, 8(1):105–119, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1963:SPP

[Ano63c] Anonymous. Summary of papers presented at meetings of the Seminar on Mathematical Statistics at the Mathematical Institute of the Academy of Sciences, USSR, 1962. Theory of Probability and its Applications, 8 (2):212–222, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:ICS

[Ano64a] Anonymous. Information concerning the symposium on limit theorems. Theory of Probability and its Applications, 9(1):167, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:RB

[Ano64b] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 9(3): 515–520, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:SAU

[Ano64c] Anonymous. The seventh allunion conference on probability theory and mathematical statistics. Theory of Probability and its Applications, 9(1):167–168, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:SPPc

[Ano64d] Anonymous. Summary of papers presented at the meetings of the Mathematical Statistics Seminar at the Mathematics Institute of the Academy of Sciences, USSR, 1963. Theory of Probability and its Applications, 9(3):509–513, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:SPPd

[Ano64e] Anonymous. Summary of papers presented at the meetings of the probability and statistics section of the Moscow mathematical society (march-may, 1964). Theory of Probability and its Applications, 9(4):681–687, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:SPPb

[Ano64f] Anonymous. Summary of papers presented at the meetings of the probability and statistics section of the Moscow mathematical society (november-december, 1963). Theory of Probability and its Applications, 9(2):356–359, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1964:SPPa

[Ano64g] Anonymous. Summary of papers presented at the meetings of the Probability and Statistics Section of the Moscow Mathematics Society (October 1962– April 1963). Theory of Probability and its Applications, 9 (1):161–166, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1965:ICL

[Ano65a] Anonymous. Information on the conference on limit theorems in probability theory. Theory of Probability and its Applications, 10(4):697–699, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1965:RB

[Ano65b] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 10(1): 190–192, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1965:SPP

[Ano65c] Anonymous. Summary of papers presented at the meetings of the probability and statistics section of the Moscow mathematical society (september-december, 1964). Theory of Probability and its Applications, 10(2):357–373, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1965:SRP

[Ano65d] Anonymous. Summary of reports presented aat the meetings of the Seminar on Mathematical Statistics of the V. A. Steklov Mathematical Institute of the Academy of Sciences of the USSR, 1964. Theory of Probability and its Applications, 10 (4):691–696, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1965:YVL

[Ano65e] Anonymous. Yu. V. Linnik’s work in probability theory and mathematical statistics. Theory of Probability and its Applications, 10(1):107–119, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:BR

[Ano66a] Anonymous. Book reviews. Theory of Probability and its Applications, 11(1):187–190, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:NBa

[Ano66b] Anonymous. New books. Theory of Probability and its Applications, 11(2):323–324, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:NBb

[Ano66c] Anonymous. New books. Theory of Probability and its Applications, 11(3):494–495, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:NVS

[Ano66d] Anonymous. Nikolai vasil’evich Smirnov. Theory of Probability and its Applications, 11(3): 436–437, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:RBa

[Ano66e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 11(2): 314–319, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:RBb

[Ano66f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 11(3): 488–494, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1966:SRP

[Ano66g] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Section of the Moscow Mathematical Society (March-November 1965). Theory of Probability and its Applications, 11(1):172–186, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:NBa

[Ano67a] Anonymous. New books. Theory of Probability and its Applications, 12(1):157, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:NBb

[Ano67b] Anonymous. New books. Theory of Probability and its Applications, 12(2):344–345, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:NBc

[Ano67c] Anonymous. New books. Theory of Probability and its Applications, 12(3):533, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:RBa

[Ano67d] Anonymous. Review and bibliography. Theory of Probability and its Applications, 12(1):154–156, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:RBb

[Ano67e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 12(3): 531–532, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:SPP

[Ano67f] Anonymous. Summary of papers presented at the sessions of the Mathematical Statistics Seminar of the Mathematics Institute of the Academy of Sciences of the USSR 1965-1966. Theory of Probability and its Applications, 12(2):337–341, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1967:SRP

[Ano67g] Anonymous. Summary of reports presented at the sessions of the Probability and Statistics Section of the Moscow Mathematical Society (December 1966– April 1967). Theory of Probability and its Applications, 12 (4):698–705, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:NBa

[Ano68a] Anonymous. New books. Theory of Probability and its Applications, 13(2):357–358, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:NBb

[Ano68b] Anonymous. New books. Theory of Probability and its Applications, 13(3):545–546, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:NBc

[Ano68c] Anonymous. New books. Theory of Probability and its Applications, 13(4):715, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:RBa

[Ano68d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 13(1): 193–194, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:RBb

[Ano68e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 13(4): 712–714, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:SRPb

[Ano68f] Anonymous. Summaries of reports presented at the sessions of the Probability and Statistics Seminar of the Mathematics Institute of the Academy of Sciences of the USSR. Theory of Probability and its Applications, 13(3):539–544, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1968:SRPa

[Ano68g] Anonymous. Summaries of reports presented at the sessions of the Probability and Statistics Society of the Moscow Mathematical Society (October–December 1967). Theory of Probability and its Applications, 13(3):535–539, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:NBa

[Ano69a] Anonymous. New books. Theory of Probability and its Applications, 14(1):181, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:NBb

[Ano69b] Anonymous. New books. Theory of Probability and its Applications, 14(2):370–371, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:NBc

[Ano69c] Anonymous. New books. Theory of Probability and its Applications, 14(4):727–728, ???? 1969.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:RBa

[Ano69d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 14(1): 176–180, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:RBb

[Ano69e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 14(2): 368–369, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:SNB

[Ano69f] Anonymous. Sergei natanovich bernshtein. Theory of Probability and its Applications, 14 (1):115–123, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:SPP

[Ano69g] Anonymous. Summaries of papers presented at the sessions of the Probability and Statistics Section of the Moscow Mathematical Society (February– December 1968). Theory of Probability and its Applications, 14 (3):530–533, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1969:SRP

[Ano69h] Anonymous. Summary of the reports presented at the sessions of the Seminar on the Theory of Probability and Mathematical Statistics at the Mathematical Institute of the Academy of Science of the USSR. Theory of Probability and its Applications, 14(4):716–725, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1970:NBa

[Ano70a] Anonymous. New books. Theory of Probability and its Applications, 15(1):162, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1970:NBb

[Ano70b] Anonymous. New books. Theory of Probability and its Applications, 15(2):357–358, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1970:NBc

[Ano70c] Anonymous. New books. Theory of Probability and its Applications, 15(3):555–556, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1970:NBd

[Ano70d] Anonymous. New books. Theory of Probability and its Applications, 15(4):728–729, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1970:RB

[Ano70e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 15(3): 552–554, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1970:SJS

[Ano70f] Anonymous. Soviet-japanese symposium on the theory of probability. Theory of Probability and its Applications, 15(1):159, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:NBa

[Ano71a] Anonymous. New books. Theory of Probability and its Applications, 16(2):398–399, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:NBb

[Ano71b] Anonymous. New books. Theory of Probability and its Applications, 16(3):573–574, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:RRP

[Ano71c] Anonymous. Résumé of reports presented at sessions of the Seminar on the theory of Probability and Mathematical Statistics of the Mathematical Institute of the Academy of Sciences of the USSR, 1970. Theory of Probability and its Applications, 16 (2):389–396, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:RBa

[Ano71d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 16(1): 196–198, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:RBb

[Ano71e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 16(4): 744–746, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:SPPa

[Ano71f] Anonymous. Summary of papers presented at the sessions of the Probability and Statistics Section of the Moscow Mathematical Society (November 1969– February 1971). Theory of Probability and its Applications, 16 (3):559–566, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1971:SPPb

[Ano71g] Anonymous. Summary of papers presented at the sessions of the Seminar in Probability Theory and Mathematical Statistics at the Mathematical Institute of the Academy of Sciences of the USSR. Theory of Probability and its Applications, 16(3):566–572, ???? 1971. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1972:NB

[Ano72a] Anonymous. New books. Theory of Probability and its Applications, 17(1):197–198, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1972:RB

[Ano72b] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 17(1): 194–196, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:NBa

[Ano73a] Anonymous. New books. Theory of Probability and its Applications, 17(2):377, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:NBb

[Ano73b] Anonymous. New books. Theory of Probability and its Applications, 17(3):570–571, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:NBc

[Ano73c] Anonymous. New books. Theory of Probability and its Applications, 17(4):755–756, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:NBd

[Ano73d] Anonymous. New books. Theory of Probability and its Applications, 18(1):205–206, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:RBa

[Ano73e] Anonymous. Review and bibliography. Theory of Probability and its Applications, 17(2):375–376, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:RBb

[Ano73f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 17(3): 568–569, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:SNP

[Ano73g] Anonymous. Scientific news and presentation of the A. A. Markov prize. Theory of Probability and its Applications, 17(2):372, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:SLG

[Ano73h] Anonymous. Summary of lectures given at the sessions of the Seminar on Probability Theory and Mathematical Statistics in the Mathematical Institute of the Academy of Sciences of the U.S.S.R. Theory of Probability and its Applications, 17 (2):365–371, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:SPPc

[Ano73i] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Section of the Moscow Mathematical Society (September 1971– April 1972). Theory of Probability and its Applications, 17 (4):745–754, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:SPPb

[Ano73j] Anonymous. Summary of papers presented at sessions of the Probability Theory and Mathematical Statistics Seminar at the T. G. Shevchenko Kiev State University. Theory of Probability and its Applications, 17(4):734–744, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:SPPa

[Ano73k] Anonymous. Summary of papers presented at the sessions of the Seminar in Probability Theory and Mathematical Statistics at the Mathematical Institute of the Academy of Sciences of the USSR, 1972. Theory of Probability and its Applications, 17 (3):558–567, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1973:YVL

[Ano73l] Anonymous. Yurii Vladimirovich Linnik. Theory of Probability and its Applications, 17(3):510, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:ICP

[Ano74a] Anonymous. International conference on probability theory and mathematical statistics. Theory of Probability and its Applications, 18(4):838–839, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:NBa

[Ano74b] Anonymous. New books. Theory of Probability and its Applications, 18(2):421–422, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:NBb

[Ano74c] Anonymous. New books. Theory of Probability and its Applications, 18(3):646–647, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:NBc

[Ano74d] Anonymous. New books. Theory of Probability and its Applications, 18(4):841, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:NBd

[Ano74e] Anonymous. New books. Theory of Probability and its Applications, 19(1):227–228, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:RB

[Ano74f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 18(3): 643–645, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:SRP

[Ano74g] Anonymous. Summaries of reports presented at sessions of the Section on Theory of Probability and Mathematical Statistics of the Moscow Mathematical Society (February-May 1973). Theory of Probability and its Applications, 18(4):816–826, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:SPPa

[Ano74h] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Section of the Moscow Mathematical Society (October-December, 1972). Theory of Probability and its Applications, 18(2):408–420, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:SPPc

[Ano74i] Anonymous. Summary of papers presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the T. G. Shevchenko Kiev State University. Theory of Probability and its Applications, 18(4):827–837, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1974:SPPb

[Ano74j] Anonymous. Summary of papers presented at the sessions of the Seminar in Probability Theory and Mathematical Statistics at the Mathematical Institute of the Academy of Sciences of the USSR. Theory of Probability and its Applications, 18(3):633–641, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:DPS

[Ano75a] Anonymous. The development of probability and statistics at the USSR Academy of Sciences. Theory of Probability and its Applications, 19(2):229–232, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:LE

[Ano75b] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 20(1):214–215, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:NBa

[Ano75c] Anonymous. New books. Theory of Probability and its Applications, 19(2):428–429, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:NBb

[Ano75d] Anonymous. New books. Theory of Probability and its Applications, 19(3):631, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:NBc

[Ano75e] Anonymous. New books. Theory of Probability and its Applications, 19(4):860–861, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:NBd

[Ano75f] Anonymous. New books. Theory of Probability and its Applications, 20(1):216–217, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:RBa

[Ano75g] Anonymous. Review and bibliography. Theory of Probability and its Applications, 19(3):630, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:RBb

[Ano75h] Anonymous. Review and bibliography. Theory of Probability and its Applications, 19(4):858–859, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:SRM

[Ano75i] Anonymous. Summary of reports made at sessions of the Seminar on Probability Theory and Mathematical Statistics at T. G. Shevchenko Kiev State University. Theory of Probability and its Applications, 19(2):413–427, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:SRPa

[Ano75j] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Section of Moscow Mathematical Society (October-December 1973). Theory of Probability and its Applications, 19(3):614–622, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1975:SRPb

[Ano75k] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Mathematical Institute of the USSR Academy of Sciences. Theory of Probability and its Applications, 19(3):622–629, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:NBa

[Ano76a] Anonymous. New books. Theory of Probability and its Applications, 20(2):444–446, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:NBb

[Ano76b] Anonymous. New books. Theory of Probability and its Applications, 21(1):223–225, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:RB

[Ano76c] Anonymous. Review and bibliography. Theory of Probability and its Applications, 21(1):221–222, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:SPPb

[Ano76d] Anonymous. Summaries of papers presented at the sessions of the Probability and Statistics Section of the Moscow Mathematical Society (December 1974– October 1975). Theory of Probability and its Applications, 21 (1):209–219, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:SPPa

[Ano76e] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Section of the Moscow Mathematical Society (February-October 1974). Theory of Probability and its Applications, 20(2):428–443, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:SRPa

[Ano76f] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences. Theory of Probability and its Applications, 20(3):661–678, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1976:SRPb

[Ano76g] Anonymous. Summary of reports presented at sessions of the Seminar on the theory of Probability and Mathematical Statistics in the Mathematical Institute of the Academy of Sciences of the USSR. Theory of Probability and its Applications, 20(3):679–687, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:LE

[Ano77a] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 21(2):446, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:NBa

[Ano77b] Anonymous. New books. Theory of Probability and its Applications, 21(2):447–448, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:NBb

[Ano77c] Anonymous. New books. Theory of Probability and its Applications, 21(4):877, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:NBc

[Ano77d] Anonymous. New books. Theory of Probability and its Applications, 22(1):201–202, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:RBa

[Ano77e] Anonymous. Review and bibliography. Theory of Probability and its Applications, 21(3):666–668, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:RBb

[Ano77f] Anonymous. Review and bibliography. Theory of Probability and its Applications, 22(1):198–200, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:SPP

[Ano77g] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Section of the Moscow Mathematical Society (November 1975– April 1976). Theory of Probability and its Applications, 21 (4):871–874, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:SRPb

[Ano77h] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1975. Theory of Probability and its Applications, 21(3):648–665, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1977:SRPa

[Ano77i] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Mathematical. Theory of Probability and its Applications, 21(2):442–446, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:NBa

[Ano78a] Anonymous. New books. Theory of Probability and its Applications, 22(2):431–432, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:NBb

[Ano78b] Anonymous. New books. Theory of Probability and its Applications, 22(3):651–652, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:NBc

[Ano78c] Anonymous. New books. Theory of Probability and its Applications, 22(4):884–885, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:NBd

[Ano78d] Anonymous. New books. Theory of Probability and its Applications, 23(1):225–226, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:RB

[Ano78e] Anonymous. Review and bibliography. Theory of Probability and its Applications, 22(2):425–429, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:SVC

[Ano78f] Anonymous. The second vilnius conference on the theory of probability and mathematical statistics. Theory of Probability and its Applications, 22(4):882–883, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:SRPc

[Ano78g] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at T. G. Shevchenko Kiev State University, 1975. Theory of Probability and its Applications, 22(3):647–650, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:SRPb

[Ano78h] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1976. Theory of Probability and its Applications, 22(3):635–646, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:SRPa

[Ano78i] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Mathematical Institute of the USSR Academy of Sciences 1976. Theory of Probability and its Applications, 22 (3):625–635, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:SRPd

[Ano78j] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at T. G. Shevchenko Kiev State University, 1975. Theory of Probability and its Applications, 23 (1):212–217, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1978:SRPe

[Ano78k] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Leningrad Section of the Mathematical Institute of the Academy of Sciences of the U.S.S.R., 1977. Theory of Probability and its Applications, 23 (1):218–224, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:LNB

[Ano79a] Anonymous. Login nikolaevich bol’shev. Theory of Probability and its Applications, 24(1): 145–146, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:NBa

[Ano79b] Anonymous. New books. Theory of Probability and its Applications, 23(2):459–460, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:NBb

[Ano79c] Anonymous. New books. Theory of Probability and its Applications, 23(3):675–676, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:NBc

[Ano79d] Anonymous. New books. Theory of Probability and its Applications, 23(4):855–856, ???? 1979.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:NBd

[Ano79e] Anonymous. New books. Theory of Probability and its Applications, 24(1):238–239, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:SPPb

[Ano79f] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Seminar in the Mathematical Institute of the USSR Academy of Sciences, 1977. Theory of Probability and its Applications, 23 (2):439–457, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:SPPa

[Ano79g] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Seminary at the T. G. Shevchenko Kiev State University, 1975. Theory of Probability and its Applications, 23 (2):431–439, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:SRPb

[Ano79h] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Kiev T. G. Shevchenko State University, 1977. Theory of Probability and its Applications, 23 (4):841–852, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:SRPa

[Ano79i] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1977. Theory of Probability and its Applications, 23(4):829–840, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1979:SRPc

[Ano79j] Anonymous. Summary of reports presented at sessions of the Seminar on the Theory of Probability and Mathematical Statistics at the Mathematical Institute of the Academy of Sciences of the USSR, 1978. Theory of Probability and its Applications, 24 (1):224–236, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:NBa

[Ano80a] Anonymous. New books. Theory of Probability and its Applications, 24(2):451–452, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:NBb

[Ano80b] Anonymous. New books. Theory of Probability and its Applications, 24(4):891–896, ???? 1980.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:NBc

[Ano80c] Anonymous. New books. Theory of Probability and its Applications, 25(1):218–219, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:RB

[Ano80d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 25(1): 216–217, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:SPPa

[Ano80e] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Seminar at the T. G. Shevchenko Kiev State University, 1978. Theory of Probability and its Applications, 24 (3):651–656, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:SPPb

[Ano80f] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Seminar in the Mathematical Institute of the USSR Academy of Sciences, 1978. Theory of Probability and its Applications, 24 (3):656–672, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:SRPc

[Ano80g] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at T. G. Shevchenko Kiev State University, 1979. Theory of Probability and its Applications, 25(1):197–203, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:SRPb

[Ano80h] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Kiev T. G. Chevchenko State University, 1978. Theory of Probability and its Applications, 24 (2):442–447, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:SRPa

[Ano80i] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1978. Theory of Probability and its Applications, 24(2):433–441, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:SRPd

[Ano80j] Anonymous. Summary of reports presented at sessions of the Seminar in Probability Theory and Mathematical Statistics at the Institute of Mathematics of the Siberian Section of the Academy of Sciences of the USSR, 1979. Theory of Probability and its Applications, 25 (1):203–213, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1980:TAU

[Ano80k] Anonymous. The thirteenth allunion colloquium on probability theory and mathematical statistics. Theory of Probability and its Applications, 24(4):888–890, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:FAUb

[Ano81a] Anonymous. The Fourteenth All-Union Colloquium on Brobability Theory and Mathematical Statistics. Theory of Probability and its Applications, 26(1):209–211, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:FAUa

[Ano81b] Anonymous. Fourth all-union seminar on stability problems in stochastic models. Theory of Probability and its Applications, 25(2):409–425, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:NBa

[Ano81c] Anonymous. New books. Theory of Probability and its Applications, 25(2):437–438, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:NBb

[Ano81d] Anonymous. New books. Theory of Probability and its Applications, 26(1):216–217, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:NSL

[Ano81e] Anonymous. News on scientific life. Theory of Probability and its Applications, 25(4):810, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:RBa

[Ano81f] Anonymous. Review and bibliography. Theory of Probability and its Applications, 25(2):433–436, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:RBb

[Ano81g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 26(1): 212–215, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:SPPa

[Ano81h] Anonymous. Summary of papers presented at sessions of the Probability and Mathematical Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1978–1979. Theory of Probability and its Applications, 25(2):426–433, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:SPPb

[Ano81i] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1979. Theory of Probability and its Applications, 26(1):184–195, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:SPPc

[Ano81j] Anonymous. Summary of papers presented at sessions of the Probability and Statistics Seminar at the V. A. Steklov Mathematical Institute of the USSR Academy of Science, 1979. Theory of Probability and its Applications, 26 (1):196–201, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:SRPb

[Ano81k] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at T. G. Shevchenko Liev State University 1979. Theory of Probability and its Applications, 26 (1):202–209, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1981:SRPa

[Ano81l] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Mathematical Institute of the Siberian Section of the USSR Academy of Sciences, 1979. Theory of Probability and its Applications, 25(3):649–657, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:FAU

[Ano82a] Anonymous. The fifteenth allunion colloquium on probability theory and mathematical statistics. Theory of Probability and its Applications, 26(4):867–869, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:FSP

[Ano82b] Anonymous. Fifth seminar on problems of stability of stochastic models. Theory of Probability and its Applications, 26 (3):638–654, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:NBa

[Ano82c] Anonymous. New books. Theory of Probability and its Applications, 26(2):432–433, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:NBb

[Ano82d] Anonymous. New books. Theory of Probability and its Applications, 26(3):655–656, ???? 1982.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:NBc

[Ano82e] Anonymous. New books. Theory of Probability and its Applications, 26(4):870, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:RB

[Ano82f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 26(2): 424–431, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SPPb

[Ano82g] Anonymous. Summary of papers presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Mathematical Institute of the Siberian Section of the USSR Academy of Sciences, 1980. Theory of Probability and its Applications, 26(3):625–633, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPf

[Ano82h] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1980. Theory of Probability and its Applications, 26 (4):856–862, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPg

[Ano82i] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at T. G. Shevchenko Kiev State University, 1980. Theory of Probability and its Applications, 26(4):862–866, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPh

[Ano82j] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Mathematical Institute of the Siberian Section of the USSR Academy of Sciences, 1980. Theory of Probability and its Applications, 27(1):206–210, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPi

[Ano82k] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Mathematical Institute of the Siberian Section of the USSR Academy of Sciences, April–May 1981. Theory of Probability and its Applications, 27 (1):211–214, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPb

[Ano82l] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar in the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1979. Theory of Probability and its Applications, 26(2):408–415, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPe

[Ano82m] Anonymous. Summary of reports presented at sessions of the Seminar in Probability Theory and Mathematical Statistics at the Steklov Institute of Mathematics of the USSR Academy of Sciences, 1980. Theory of Probability and its Applications, 26 (4):852–855, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPc

[Ano82n] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Institute of Mathematics of the Siberian Section of the Academy of Sciences of the USSR, 1979. Theory of Probability and its Applications, 26 (2):416–423, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPd

[Ano82o] Anonymous. Summary of reports presented at sessions of the Seminar on the Probability Theory and Mathematical Statistics at the T. G. Shevchenko Kiev State University, 1980. Theory of Probability and its Applications, 26 (3):633–638, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SRPa

[Ano82p] Anonymous. Summary of reports presented at sessions of the Seminar on the Theory of Probability and Mathematical Statistics at the V. A. Steklov Mathematical Institute of the Academy of Sciences of the USSR, 1980. Theory of Probability and its Applications, 26(2):399–407, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:SPPa

[Ano82q] Anonymous. Summary papers presented at sessions of the Probability and Mathematical Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1980. Theory of Probability and its Applications, 26 (3):617–625, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1982:TA

[Ano82r] Anonymous. A triple anniversary. Theory of Probability and its Applications, 26(4):867, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1983:NBa

[Ano83a] Anonymous. New books. Theory of Probability and its Applications, 27(2):438–440, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1983:NBb

[Ano83b] Anonymous. New books. Theory of Probability and its Applications, 27(3):664–665, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1983:SRPc

[Ano83c] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the Mathematics Institute of the Siberian Section of the USSR Academy of Sciences, September–October 1981. Theory of Probability and its Applications, 27(3):661–663, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1983:SRPb

[Ano83d] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the T. G. Shevchenko Kiev State University, May–December 1981. Theory of Probability and its Applications, 27(3):653–661, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1983:SRPa

[Ano83e] Anonymous. Summary of reports presented at sessions of the Seminar in the Theory of Probability and Mathematical Statistics at the Kiev Shevchenko State University 1981. Theory of Probability and its Applications, 27 (2):430–437, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1983:SRPd

[Ano83f] Anonymous. Summary of reports presented at sessions of the Seminar Probability Theory and Mathematical Statistics at the V. A. Steklov Mathematical Institute of the USSR Academy of Sciences, 1981. Theory of Probability and its Applications, 27 (4):872–886, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:NBa

[Ano84a] Anonymous. New books. Theory of Probability and its Applications, 28(1):216–217, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:NBb

[Ano84b] Anonymous. New books. Theory of Probability and its Applications, 28(4):859–861, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:NM

[Ano84c] Anonymous. News in mathematics. Theory of Probability and its Applications, 28(1):208–209, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:NWM

[Ano84d] Anonymous. News in the world of mathematics. Theory of Probability and its Applications, 28 (4):858, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:NSL

[Ano84e] Anonymous. News on scientific life. Theory of Probability and its Applications, 28(3):651–655, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:RB

[Ano84f] Anonymous. Review and bibliography. Theory of Probability and its Applications, 28(1):210–215, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:SVV

[Ano84g] Anonymous. Seminars VI and VII on problems of continuity and stability of stochastic models. Theory of Probability and its Applications, 28(4):841–855, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:SRPa

[Ano84h] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences 1981. Theory of Probability and its Applications, 28 (1):194–207, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:SRPe

[Ano84i] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the Leningrad Section of the Mathematical Institute of the USSR Academy of Sciences, 1982. Theory of Probability and its Applications, 28 (4):830–840, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:SRPb

[Ano84j] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the Mathematics Institute of the Siberian Section of the USSR Academy of Sciences, February– May 1982. Theory of Probability and its Applications, 28 (3):631–639, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:SRPc

[Ano84k] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the T. G. Shevchenko Kiev State University, February–May 1982. Theory of Probability and its Applications, 28(3):640–647, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1984:SRPd

[Ano84l] Anonymous. Summary of reports presented at sessions of the Probability Theory and Mathematical Statistics Seminar at the Leningrad Section of the Mathematical Institute if the USSR Academy of Sciences, 1982. Theory of Probability and its Applications, 28(3):647–650, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:AEE

[Ano85a] Anonymous. Addendum: On the existence of exact upper sequences. Theory of Probability and its Applications, 29(3):639, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:CE

[Ano85b] Anonymous. Corrections from the Editors. Theory of Probability and its Applications, 29(1):200, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:E

[Ano85c] Anonymous. Erratum. Theory of Probability and its Applications, 29(4):859, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:NBa

[Ano85d] Anonymous. New books. Theory of Probability and its Applications, 29(1):202–204, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:NBb

[Ano85e] Anonymous. New books. Theory of Probability and its Applications, 29(2):425, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:NBc

[Ano85f] Anonymous. New books. Theory of Probability and its Applications, 29(3):644–645, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:NBd

[Ano85g] Anonymous. New books. Theory of Probability and its Applications, 29(4):858, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:NSLa

[Ano85h] Anonymous. News on scientific life. Theory of Probability and its Applications, 29(1):200–201, ???? 1985. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:NSLb

[Ano85i] Anonymous. News on scientific life. Theory of Probability and its Applications, 29(3):640–643, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPc

[Ano85j] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at Kiev State University, 1982 (September to December). Theory of Probability and its Applications, 29(1):190–196, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPe

[Ano85k] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the Mathematics Institute of the Siberian Section of the USSR Academy of Sciences, February– May 1983. Theory of Probability and its Applications, 29 (2):419–424, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPd

[Ano85l] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the T. G.

Shevchenko Kiev State University, February–May 1983. Theory of Probability and its Applications, 29(2):412–419, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPg

[Ano85m] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar at the T. G. Shevchenko Kiev State University, September–December 1983. Theory of Probability and its Applications, 29(3):634–638, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPf

[Ano85n] Anonymous. Summary of reports presented at sessions of the Probability and Mathematical Statistics Seminar of the Mathematics Institute of the Siberian Section of the USSR Academy of Sciences, October–December 1983. Theory of Probability and its Applications, 29(3):628–633, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPh

[Ano85o] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Leningrad Section at the V. A. Steklov Mathematical Institute of the USSR Academy of Sciences, February–December 1983. Theory of Probability and its Applications, 29(4):836–853, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPb

[Ano85p] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the Siberian Section of the Mathematical Institute of the USSR Academy of Sciences, 1982. Theory of Probability and its Applications, 29(1):183–190, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPa

[Ano85q] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the V. A. Steklov Mathematical Institute of the USSR Academy of Sciences, 1982-1983. Theory of Probability and its Applications, 29(1):169–183, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1985:SRPi

[Ano85r] Anonymous. Summary of reports presented at sessions of the Probability and Statistics Seminar at the V. A. Steklov Mathematical Institute of the USSR Academy of Sciences, February-June, 1984. Theory of Probability and its Applications, 29 (4):853–857, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NBa

[Ano86a] Anonymous. New books. Theory of Probability and its Applications, 30(1):224, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NBb

[Ano86b] Anonymous. New books. Theory of Probability and its Applications, 30(2):437–438, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NBc

[Ano86c] Anonymous. New books. Theory of Probability and its Applications, 30(3):658–659, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NBd

[Ano86d] Anonymous. New books. Theory of Probability and its Applications, 30(4):875–876, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NSLb

[Ano86e] Anonymous. News of scientific life. Theory of Probability and its Applications, 30(2):436, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NSLa

[Ano86f] Anonymous. News on scientific life. Theory of Probability and its Applications, 30(1):223, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NSLc

[Ano86g] Anonymous. News on scientific life. Theory of Probability and its Applications, 30(3):656–657, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:NSLd

[Ano86h] Anonymous. News on scientific life. Theory of Probability and its Applications, 30(4):872, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:RB

[Ano86i] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 30(2): 432–435, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:SRPc

[Ano86j] Anonymous. Summary of reports presented at sessions of the Seminar in Probability Theory and Mathematical Statistics at the Shevchenko State University, Kiev. Theory of Probability and its Applications, 30(3):644–655, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:SRPa

[Ano86k] Anonymous. Summary of reports presented at sessions of the Seminar on Probability and Mathematical Statistics at the Leningrad Branch of the Steklov Mathematical Institute of the Academy of Sciences of the USSR. Theory of Probability and its Applications, 30(1):207–217, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:SRPb

[Ano86l] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Shevchenko Kiev State University. Theory of Probability and its Applications, 30(1):217–222, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1986:VSP

[Ano86m] Anonymous. VIII seminar on problems of stability of stochastic models. Theory of Probability and its Applications, 30 (3):633–644, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:IIG

[Ano87a] Anonymous. Iosif il’ich gikhman.

Theory of Probability and its Applications, 31(1):116–117, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:NBa

[Ano87b] Anonymous. New books. Theory of Probability and its Applications, 31(1):176–177, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:NBb

[Ano87c] Anonymous. New books. Theory of Probability and its Applications, 31(2):374, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:NBc

[Ano87d] Anonymous. New books. Theory of Probability and its Applications, 31(3):560–561, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:NBd

[Ano87e] Anonymous. New books. Theory of Probability and its Applications, 31(4):741–742, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:NBN

[Ano87f] Anonymous. New books from nauka Press. Theory of Probability and its Applications, 31 (3):561–562, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:NSL

[Ano87g] Anonymous. News of scientific life. Theory of Probability and its Applications, 31(2):372–373, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:RB

[Ano87h] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 31(2): 368–371, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPe

[Ano87i] Anonymous. Summary of reports presented at sessions of the Probability Theory and Mathematical Statistics Seminar at the T. G. Shevchenko Kiev State University (December 1984–May 1985). Theory of Probability and its Applications, 31(3):534–543, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPc

[Ano87j] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Institute of Mathematics of the Siberian Section of the USSR Academy of Sciences, February– June 1985. Theory of Probability and its Applications, 31 (2):359–363, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPa

[Ano87k] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Institute of Mathematics of the Siberian Section of the USSR Academy of Sciences, September–December 1984. Theory of Probability and its Applications, 31(1):160–165, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPg

[Ano87l] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Leningrad Branch of the V. A. Steklov Mathematical Institute of the Academy of Sciences of the USSR (February-December 1985). Theory of Probability and its Applications, 31 (3):551–559, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPd

[Ano87m] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the V. A. Steklov Mathematics Institute of the USSR Academy of Sciences, November 1984–July 1985. Theory of Probability and its Applications, 31(2):364–367, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPb

[Ano87n] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics in the Leningrad Section of the V. A. Steklov Mathematical Institute of the USSR Academy of Sciences, September–December 1984. Theory of Probability and its Applications, 31(1):166–174, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:SRPf

[Ano87o] Anonymous. Summary of reports presented at the seminar entitled “Probability Methods in Discrete Mathematics” at the V. A. Steklov Mathematical Institute (April–December 1985). Theory of Probability and its Applications, 31(3):544–550, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1987:VVS

[Ano87p] Anonymous. Vyacheslav vasil’evich sazonov. Theory of Probability and its Applications, 31(1):118, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:FWC

[Ano88a] Anonymous. The First World Congress of the Bernoulli Society of Mathematical Statistics and Probability Theory. Theory of Probability and its Applications, 32(2):199–200, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:NBa

[Ano88b] Anonymous. New books. Theory of Probability and its Applications, 32(1):196–197, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:NBb

[Ano88c] Anonymous. New books. Theory of Probability and its Applications, 32(2):387–388, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:NBc

[Ano88d] Anonymous. New books. Theory of Probability and its Applications, 32(3):570–571, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:NBd

[Ano88e] Anonymous. New books. Theory of Probability and its Applications, 32(4):757–758, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:NSLa

[Ano88f] Anonymous. News of scientific life. Theory of Probability and its Applications, 32(2):384–385, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:NSLb

[Ano88g] Anonymous. News of scientific life: XX school-colloquium on probability theory and mathematical statistics. Theory of Probability and its Applications, 32(4):751, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:OAA

[Ano88h] Anonymous. Obituary: Academician Andrei Nikolaevich Kolmogorov. Theory of Probability and its Applications, 32 (4):707, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:RBa

[Ano88i] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 32(1): 191–195, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:RBb

[Ano88j] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 32(4): 752–756, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:SRPa

[Ano88k] Anonymous. Summaries of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Shevchenko State University, Kiev (1985). Theory of Probability and its Applications, 32 (1):179–185, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1988:SRPb

[Ano88l] Anonymous. Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Institute of Mathematics of the Siberian Section of the USSR Academy of Sciences (September-December 1985). Theory of Probability and its Applications, 32(1):186–190, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:AUS

[Ano89a] Anonymous. All-union schoolseminar on ergodic theory of Markov processes. Theory of Probability and its Applications, 33(1):199, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:I

[Ano89b] Anonymous. Information. Theory of Probability and its Applications, 33(3):593, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:ISP

[Ano89c] Anonymous. International seminar on problems of stability of stochastic models. Theory of Probability and its Applications, 33(1):198, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:NBa

[Ano89d] Anonymous. New books. Theory of Probability and its Applications, 33(1):204–205, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:NBb

[Ano89e] Anonymous. New books. Theory of Probability and its Applications, 33(2):403–404, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:NBc

[Ano89f] Anonymous. New books. Theory of Probability and its Applications, 33(3):591–593, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:OSK

[Ano89g] Anonymous. Obituary: Sagdy Khasanovich Sirazhdinov. Theory of Probability and its Applications, 33(4):698–699, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:RBa

[Ano89h] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 33(1):200, ???? 1989. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:RBb

[Ano89i] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 33(1): 201–203, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:RBc

[Ano89j] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 33(3): 586–590, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1989:TFS

[Ano89k] Anonymous. Twenty-First School-Colloquium on Probability Theory and Mathematical Statistics. Theory of Probability and its Applications, 33 (1):197–198, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:BS

[Ano90a] Anonymous. The Bernoulli Society. Theory of Probability and its Applications, 34(2):378–380, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:CSC

[Ano90b] Anonymous. The charter of the Soviet Committee of the Bernoulli Society for Mathematical Statistics, Probability Theory, Combinatorics and Their Applications. Theory of Probability and its Applications, 34 (2):376–377, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:CPM

[Ano90c] Anonymous. The conference “Probability Models of Control and Reliability Processes”. Theory of Probability and its Applications, 34(2):380, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:NBa

[Ano90d] Anonymous. New books. Theory of Probability and its Applications, 34(1):205–206, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:NBb

[Ano90e] Anonymous. New books. Theory of Probability and its Applications, 34(2):382–383, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:NBc

[Ano90f] Anonymous. New books. Theory of Probability and its Applications, 34(3):561–564, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:NSL

[Ano90g] Anonymous. News of scientific life: Two expositions. Theory of Probability and its Applications, 34(4):749–752, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:OSC

[Ano90h] Anonymous. On the organization of the Soviet Committee of the Bernoulli Society. Theory of Probability and its Applications, 34(2):376, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:RB

[Ano90i] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 34(3): 556–560, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1990:TSC

[Ano90j] Anonymous. Twelfth schoolcolloquium on probability theory and mathematical statistics. Theory of Probability and its Applications, 34(2):381, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:A

[Ano91a] Anonymous. Announcement. Theory of Probability and its Applications, 35(3):618, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:NBa

[Ano91b] Anonymous. New books. Theory of Probability and its Applications, 35(1):201–204, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:NBb

[Ano91c] Anonymous. New books. Theory of Probability and its Applications, 35(2):408–409, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:NBc

[Ano91d] Anonymous. New books. Theory of Probability and its Applications, 35(3):622–623, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:NSLa

[Ano91e] Anonymous. News of scientific life. Theory of Probability and its Applications, 35(1):194–200, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:NSLb

[Ano91f] Anonymous. News of scientific life. Theory of Probability and its Applications, 35(4):827–829, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:RBa

[Ano91g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 35(2): 405–407, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:RBb

[Ano91h] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 35(3): 619–621, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1991:RBc

[Ano91i] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 35(4): 825–826, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1992:NB

[Ano92a] Anonymous. New books. Theory of Probability and its Applications, 36(1):207–208, ???? 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1992:NSLa

[Ano92b] Anonymous. News of scientific life. Theory of Probability and its Applications, 36(1):201–206, ???? 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1992:NSLb

[Ano92c] Anonymous. News of scientific life. Theory of Probability and its Applications, 36(3):640–645, ???? 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1992:SIS

[Ano92d] Anonymous. The suzdal international seminar of 1991 on stability problems for stochastic models. Theory of Probability and its Applications, 36(4):803–858, ???? 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1993:WCB

[Ano93a] Anonymous. 2nd World Congress of the Bernoulli Society. Theory of Probability and its Applications, 37(1):1–2, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1993:NB

[Ano93b] Anonymous. New books. Theory of Probability and its Applications, 37(1):206–207, ???? 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1993:NSL

[Ano93c] Anonymous. News of scientific life. Theory of Probability and its Applications, 37(1):194–205, ???? 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1993:XSS

[Ano93d] Anonymous. XVI seminar on stability problems of stochastic models. Theory of Probability and its Applications, 37(4): 707–760, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 763–820.

Anonymous:1994:BAS

[Ano94a] Anonymous. Boris Aleksandrovich Sevast’yanov. Theory of Probability and its Applications, 38(4):721, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1994:E

[Ano94b] Anonymous. Errata. Theory of Probability and its Applications, 38(4):720, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1994:NSLa

[Ano94c] Anonymous. News of scientific life. Theory of Probability and its Applications, 38(1):174, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1994:NSLb

[Ano94d] Anonymous. News of scientific life. Theory of Probability and its Applications, 38(3):572, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1994:NSLc

[Ano94e] Anonymous. News of scientific life. Theory of Probability and its Applications, 38(4):765–767, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1995:NBa

[Ano95a] Anonymous. New books. Theory of Probability and its Applications, 39(3):535, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1995:NBb

[Ano95b] Anonymous. New books. Theory of Probability and its Applications, 39(4):726, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1995:NSL

[Ano95c] Anonymous. News of scientific life: Actuarial and financial center for scientific investigation. Theory of Probability and its Applications, 39(1):195–196, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1995:RB

[Ano95d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 39(4): 720–725, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1996:C

[Ano96a] Anonymous. In commemoration.

Theory of Probability and its Applications, 40(1):202, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1996:NSL

[Ano96b] Anonymous. News of scientific life. Theory of Probability and its Applications, 40(1):201, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1996:RBa

[Ano96c] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 40(2): 403–408, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1996:RBb

[Ano96d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 40(3): 597–598, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:1997:NBa

[Ano97a] Anonymous. New books. Theory of Probability and its Applications, 41(1):196–197, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97509.

Anonymous:1997:NBb

[Ano97b] Anonymous. New books. Theory of Probability and its Applications, 41(3):614, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97542.

Anonymous:1997:NSL

[Ano97c] Anonymous. News of scientific life. Theory of Probability and its Applications, 41 (4):820, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97588.

Anonymous:1997:PLC

[Ano97d] Anonymous. Pafnutiĭ L′vovich Chebyshev (on the 175th anniversary of his birth). Theory of Probability and its Applications, 41(3):519–530, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97527. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 612–627.

Anonymous:1997:RBa

[Ano97e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 41(3): 609–613, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97540.

Anonymous:1997:RBb

[Ano97f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 41(4): 815–819, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97587.

Anonymous:1997:SCa

[Ano97g] Anonymous. Short communications. Theory of Probability and its Applications, 41(1): 137–195, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97508.

Anonymous:1997:SCb

[Ano97h] Anonymous. Short communications. Theory of Probability and its Applications, 41(2): 328–403, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97554.

Anonymous:1997:SCc

[Ano97i] Anonymous. Short communications. Theory of Probability and its Applications, 41(3): 532–608, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97528.

Anonymous:1997:SCd

[Ano97j] Anonymous. Short communications. Theory of Probability and its Applications, 41(4): 741–814, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97575.

Anonymous:1998:AWP

[Ano98a] Anonymous. Award of the Wolf prize. Theory of Probability and its Applications, 42(4): 717–719, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97661.

Anonymous:1998:RBa

[Ano98b] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 42(1): 186–188, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97607.

Anonymous:1998:RBb

[Ano98c] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 42(4): 714–716, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97659.

Anonymous:1998:XIS

[Ano98d] Anonymous. The XVIIIth international seminar on stability problems for stochastic models. Theory of Probability and its Applications, 42(2):315–383, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97621. Translated by

M. V. Khatuntseva, V. Yu. Korolev and others.

Anonymous:1999:LE

[Ano99a] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 43(4): 682, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97726.

Anonymous:1999:NSL

[Ano99b] Anonymous. News of scientific life. Theory of Probability and its Applications, 43(4): 685–686, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97728.

Anonymous:1999:RB

[Ano99c] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 43(4): 683–684, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97727.

Anonymous:1999:ISS

[Ano99d] Anonymous. The 20th International Seminar on Stability Problems for Stochastic Models, Lublin–Naleczow, Poland, September 5–11, 1999. Theory of Probability and its Applications, 43(4):687–688, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97729.

Antonov:1980:PIS

[Ant80] S. N. Antonov. Probability inequalities for series of independent random variables. Theory of Probability and its Applications, 24(3):636–640, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anulova:1980:CWP

[Anu80] S. V. Anulova. On control of a Wiener process in a halfplane with oblique reflection on the boundary. Theory of Probability and its Applications, 25 (1):41–57, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anulova:1981:MST

[Anu81] S. V. Anulova. On Markov stopping times with a given distribution for a Wiener process. Theory of Probability and its Applications, 25(2):362–366, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anulova:1991:CSD

[Anu91] S. V. Anulova. Counterexamples:

a stochastic differential equation with linearly increasing coefficients May have an explosive solution within a domain. Theory of Probability and its Applications, 35(2):336–338, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 329–331.

Arbekov:1997:LBP

[AP97] I. M. Arbekov and G. V. Proskurin. On a lower bound for the probability of the correct classification of observations. Theory of Probability and its Applications, 41(3):532–537, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 628–633.

Apokorin:1967:GMC

[Apo67] D. S. Apokorin. Gaussian measures corresponding to generalized stationary processes. Theory of Probability and its Applications, 12(4):638–646, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Asriev:1986:CRI

[AR86] A. V. Asriev and V. I. Rotar’. On the convergence rate in the infinite-dimensional Central Limit Theorem for probabilities of hitting parallelepipeds. Theory of Probability and its Applications, 30(4):691–701, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Albeverio:1994:SEE

[AR94] S. Albeverio and Yu. A. Rozanov. On stochastic evolution equations with stochastic boundary conditions. Theory of Probability and its Applications, 38(1):1–13, March 1994. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 3–19.

Arato:1961:SSS

[Ara61] M. Arato. On the sufficient statistics for stationary Gaussian random processes. Theory of Probability and its Applications, 6(2):199–201, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arato:1968:ECL

[Ara68] M. Arató. Evaluation of the confidence limits for the “Decay” parameter of a complex stationary Gaussian Markov process. Theory of Probability and its Applications, 13(2):314–320, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1972:LBT

[Ara72] T. V. Arak. On Lévy–Baxter theorems for random fields. Theory of Probability and its Applications, 17(1):153–159, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1975:DMS

[Ara75] T. V. Arak. On the distribution of the maximum of successive partial sums of independent random variables. Theory of Probability and its Applications, 19 (2):245–266, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1976:EB

[Ara76] T. V. Arak. On an estimate of A. A. Borovkov. Theory of Probability and its Applications, 20 (2):372–373, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1981:AFC

[Ara81] T. V. Arak. On the approximation of n-fold convolutions of distributions having non-negative characteristic functions with accompanying laws. Theory of Probability and its Applications, 25(2):221–243, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1982:CRKa

[Ara82a] T. V. Arak. On the convergence rate in Kolmogorov’s uniform limit theorem. I. Theory of Probability and its Applications, 26(2):219–239, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1982:CRKb

[Ara82b] T. V. Arak. On the convergence rate in Kolmogorov’s uniform limit theorem. II. Theory of Probability and its Applications, 26(3):437–451, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arak:1983:ILB

[Ara83] T. V. Arak. An improvement of the lower bound for the rate of convergence in Kolmogorov’s uniform limit theorem. Theory of Probability and its Applications, 27(4):826–832, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arato:1985:PEP

[Ara85] M. Arató. On parameter estimation in the presence of noise. Theory of Probability and its Applications, 29(3):622–628, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arato:1986:SSG

[Ara86] M. Arató. On sufficient statistics of Gaussian processes with rational spectral density function. Theory of Probability and its Applications, 30(1):103–116, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arato:1990:LTG

[Ara90] N. M. Arató. On a limit theorem for generalized Gaussian random fields defined by stochastic partial differential equations. Theory of Probability and its Applications, 34(2):363–366, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 409–411.

Arato:1999:EMV

[Ara99] N. M. Arato. On estimating the mean value of Levy’s Brownian motion. Theory of Probability and its Applications, 43 (1):123–125, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97671.

Arenbaev:1977:ABM

[Are77] N. K. Arenbaev. Asymptotic behavior of the multinomial dstribution. Theory of Probability and its Applications, 21(4):805–810, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arenbaev:1978:IRV

[Are78] N. K. Arenbaev. Inequalities for random vectors. Theory of Probability and its Applications, 22 (3):574–579, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkin:1964:POD

[Ark64a] V. I. Arkin. A problem of optimum distribution of search effort. Theory of Probability and its Applications, 9(1):159–160, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkin:1964:UOS

[Ark64b] V. I. Arkin. Uniformly optimal strategies in search problems. Theory of Probability and its Applications, 9(4):674–680, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkin:1965:SEP

[Ark65] V. I. Arkin. Some extremal problems connected with the theory of search. Theory of Probability and its Applications, 10(3):535–538, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkharov:1971:CIT

[Ark71] L. V. Arkharov. On chebyshev’S inequality in the two-dimensional case. Theory of Probability and its Applications, 16(2):356–361, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkhangelskii:1987:LBP

[Ark87] A. M. Arkhangel’skii. On lower bounds for probabilities of large deviations. Theory of Probability and its Applications, 31 (2):346–351, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkhangelskii:1988:LBP

[Ark88] A. N. Arkhangel’skii. Lower bounds for the probabilities of large deviations in the multidimensional case. Theory of Probability and its Applications, 32 (2):375–380, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkhangelskii:1990:LBP

[Ark90] A. N. Arkhangel′skiĭ. Lower bounds for probabilities of large deviations for sums of independent random variables. Theory of Probability and its Applications, 34(4):565–575, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 625–635.

Arndt:1981:APD

[Arn81] K. Arndt. Asymptotic properties of the distribution of the supremum of a random walk on a Markov chain. Theory of Probability and its Applications, 25 (2):309–324, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Applegate:1997:SPP

[ARS^{+}97] D. Applegate, J. Reeds, S. Scheinberg, L. Shepp, and P. Shor. Some problems in probabilistic tomography. Theory of Probability and its Applications, 41 (2):199–209, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97543.

Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 323–335.

Artamonovskii:1989:MLE

[Art89] V. P. Artamonovskiĭ. On maximum likelihood estimation of the shift and scale parameters based on grouped samples. Theory of Probability and its Applications, 33(4):705–708, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 33(4), (1988), pp. 759–762.

Arjas:1974:NSF

[AS74] E. Arjas and T. P. Speed. A note on the second factorization identity of A. A. Borovkov. Theory of Probability and its Applications, 18(3):576–578, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ayupov:1982:HMC

[AS82] Sh. A. Ayupov and T. A. Sarymsakov. Homogeneous Markov chains on semifields. Theory of Probability and its Applications, 26(3):510–520, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aliev:1983:TPC

[AS83] S. A. Aliev and V. M. Shurenkov.

Transitional phenomena and the convergence of Galton–Watson processes to Jirina processes. Theory of Probability and its Applications, 27(3):472–485, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Antonets:1992:MTG

[AS92] M. A. Antonets and I. A. Shereshevskiĭ. On a model of tree growth. Theory of Probability and its Applications, 36(3): 565–570, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 548–553.

Aleskeviciene:1998:IFC

[AS98] A. K. Aleskeviciene and V. A. Statulevicius. Inversion formulas in the case of a discontinuous limit law. Theory of Probability and its Applications, 42(1):1–16, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97589. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 3–20.

Asarin:1988:SPK

[Asa88] E. A. Asarin. Some properties of Kolmogorov Δ-random finite sequences. Theory of Probability and its Applications, 32 (3):507–508, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Asarin:1991:EBT

[Asa91] E. A. Asarin. On the equivalence of being typical and chaotic for finite objects. Theory of Probability and its Applications, 35 (4):760–761, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 756–758.

Ashurov:1968:FMC

[Ash68] V. M. Ashurov. On filtration of Markov chains on a finite group. Theory of Probability and its Applications, 13(3):490–494, ???? 1968. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Antonets:1986:RRV

[AT86] M. A. Antonets and A. A.

Tsvetkov. On representing random variables as continuous functions. Theory of Probability and its Applications, 30(3):577–580, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anderson:1993:LJD

[AT93] C. W. Anderson and K. F. Turkman. Limiting joint distributions of sums and maxima in a statistical context. Theory of Probability and its Applications, 37 (2):314–316, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 352–355.

Azlarov:1982:LLN

[AV82] T. A. Azlarov and N. A. Volodin.

Laws of large numbers for identically distributed Banach-space valued random variables. Theory of Probability and its Applications, 26(3):573–580, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Averintsev:1972:DMR

[Ave72] M. B. Averintsev. Description of Markovian random fields by Gibbsian conditional probabilities. Theory of Probability and its Applications, 17(1):

20–33, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arbekov:1995:PDR

[AVR95] I. M. Arbekov, V. I. Vinokurov, and B. V. Ryazanov. On the problem of detecting random trajectories. Theory of Probability and its Applications, 39(4): 547–557, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 669–680.

Babkin:1964:DPM

[Bab64] V. I. Babkin. On the distribution of the position of the maxima of a diffusion process with variable parameter. Theory of Probability and its Applications, 9(2):338–342, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Babayan:1985:ABP

[Bab85] N. M. Babayan. On asymptotic behavior of the prediction error in the singular case. Theory of Probability and its Applications, 29(1):147–150, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bagdonavicius:1979:STM

[Bag79] V. B. Bagdonavicius. A statistical test of a model of additive accumulation of damage. Theory of Probability and its Applications, 23(2):385–390, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bagirov:1988:LBD

[Bag88] E. B. Bagirov. A lower bound for distributions of quadratic forms of normally distributed random variables. Theory of Probability and its Applications, 32(4): 744–748, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bagirov:1989:ALB

[Bag89a] È. B. Bagirov. Addendum: a lower bound for the distributions of quadratic forms of normally distributed random variables. Theory of Probability and its Applications, 33(1):196, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 211–211.

Bagirov:1989:LBT

[Bag89b] E. B. Bagirov. Lower bounds for the tails of distributions of cubic forms of normally distributed random variables. Theory of Probability and its Applications, 33(4):711–715, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bagirov:1989:SRM

[Bag89c] E. B. Bagirov. Some remarks on mixtures of normal distributions. Theory of Probability and its Applications, 33(4):709–710, ???? 1989. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Bagirov:1992:NCK

[Bag92] È. B. Bagirov. New consequences of Kolmogorv’s formula for infinitely divisible laws. Theory of Probability and its Applications, 36(1):138–143, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 138–143.

Bakirov:1989:EDS

[Bak89a] N. K. Bakirov. Extrema of the distribution of the S^{2} statistic. Theory of Probability and its Applications, 33(1):171–175, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 184–188.

Bakirov:1989:LE

[Bak89b] N. K. Bakirov. Letter to the editors: “Inequalities for distributions of normalized quadratic forms in random variables” [Teor. Veroyatnost. i Primenen. 23 (1978), no. 3, 676–680; MR 80h:60030] by L. S. Ponomarenko and “A lower bound for distributions of quadratic forms of normally distributed random variables” [ibid. 32 (1987), no. 4, 815–819; MR 90m:62118a] by È. B. Bagirov. Theory of Probability and its Applications, 33(4):765, ???? 1989. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 823–823.

Bakirov:1990:EDQ

[Bak90] N. K. Bakirov. Extrema of the distributions of quadratic forms of Gaussian variables. Theory of Probability and its Applications, 34(2):207–215, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 241–250.

Bakhtin:1996:LDU

[Bak96a] V. I. Bakhtin. Large deviations under power mixing. Theory of Probability and its Applications, 40(4):742–744, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bakirov:1996:CTD

[Bak96b] N. K. Bakirov. Comparison theorems for distribution functions of quadratic forms of Gaussian vectors. Theory of Probability and its Applications, 40(2):340–348, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 404–412.

Bakirov:1996:EDF

[Bak96c] N. K. Bakirov. The extrema of the distribution function for the ratio of quadratic forms in Gaussian variables. Theory of Probability and its Applications, 40 (3):542–546, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 642–646.

Bakirov:1996:RAP

[Bak96d] N. K. Bakirov. A refinement of asymptotics in the Prokhorovdonsker invariance principle for integral functionals. Theory of Probability and its Applications, 40(4):613–634, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 705–730.

Bakirov:1996:UPC

[Bak96e] N. K. Bakirov. A uniqueness property of the chi-square distribution. Theory of Probability and its Applications, 40(1):130–136, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 159–165.

Bakhtin:1998:LDR

[Bak98] V. I. Bakhtin. Large deviations of random variables with a finite number of approximately evaluated cumulants. Theory of Probability and its Applications, 42 (3):513–517, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 603–608.

Balakin:1967:RM

[Bal67] G. V. Balakin. On random matrices. Theory of Probability and its Applications, 12(2):301–307, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Balakin:1968:DRR

[Bal68] G. V. Balakin. The distribution of the rank of random matrices over a finite field. Theory of Probability and its Applications, 13 (4):594–605, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Balakin:1974:DNS

[Bal74] G. V. Balakin. On the distribution of the number of solutions of systems of random Boolean equations. Theory of Probability and its Applications, 18(3):598–602, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Balakrishnan:1994:IDG

[Bal94] A. V. Balakrishnan. An infinitedimensional Gaussian Markov process with nonnuclear steady state covariance. Theory of Probability and its Applications, 38 (3):516–520, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 656–661.

Balakin:1996:GST

[Bal96] G. V. Balakin. Graphs of systems of two-term equations with Boolean unknowns. Theory of Probability and its Applications, 40(2):203–219, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Banjevic:1985:KI

[Ban85] D. Banjević. On a Kolmogorov inequality. Theory of Probability and its Applications, 29 (2):391–394, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 29(2), (1984), pp. 382–384.

Barbarosie:1977:TCM

[Bar77] A. Barbarosie. On the theory of controlled Markov processes. Theory of Probability and its Applications, 22(1): 53–69, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barsov:1985:LTT

[Bar85] D. A. Barsov. Limit theorems for traces of sample covariance matrices. Theory of Probability and its Applications, 29(2):383–391, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barsov:1986:ANA

[Bar86] S. S. Barsov. On the accuracy of the normal approximation of the distribution of a random sum of random vectors. Theory of Probability and its Applications, 30(2):376–379, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barsov:1988:RCN

[Bar88] S. S. Barsov. Rate of convergence to the normal distribution and decrease of the tail of the summand distribution. Theory of Probability and its Applications, 32(2):329–331, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barsov:1989:CP

[Bar89] S. S. Barsov. On a comparison principle. Theory of Probability and its Applications, 33 (3):423–432, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 453–464.

Barabanov:1991:OFS

[Bar91] A. E. Barabanov. Optimal filtering of a stationary process with respect to measured discrepancies. Theory of Probability and its Applications, 35(2):339–341, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 337–340.

Baron:1997:FPT

[Bar97] M. I. Baron. On the first passage time for waiting processes. Theory of Probability and its Applications, 41(2):328–334, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 396–403.

Basharin:1958:FPM

[Bas58] G. P. Basharin. Final probabilities for multi-dimensional Markov processes which describe the action of some two-stage telephone systems with busysignals. Theory of Probability and its Applications, 3(4):420–425, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Basharin:1959:SEE

[Bas59] G. P. Basharin. On a statistical estimate for the entropy of a sequence of independent random variables. Theory of Probability and its Applications, 4 (3):333–336, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Basharin:1960:CTL

[Bas60] G. P. Basharin. The congestion time limit distribution for a fully available group of trunks. Theory of Probability and its Applications, 5(2):223–228, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Basharin:1965:SSF

[Bas65] G. P. Basharin. A single server with a finite queue and items of several types. Theory of Probability and its Applications, 10 (2):261–274, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Basis:1977:IDM

[Bas77] V. Ya. Basis. Infinite-dimensional Markov processes with almost local interaction of components. Theory of Probability and its Applications, 21(4):706–720, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Basalikas:1993:CDM

[Bas93] A. A. Basalikas. On the closeness of distributions of multilinear forms. Theory of Probability and its Applications, 37(3): 405–414, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 449–458.

Baushev:1988:WCG

[Bau88]  A. N. Baushev.  On weak

convergence of Gaussian measures. Theory of Probability and its Applications, 32(4):670–677, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Baushev:1996:WCP

[Bau96] A. N. Baushev. On the weak convergence of probability measures in Orlicz spaces. Theory of Probability and its Applications, 40(3):420–429, ???? 1996.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 495–507.

Baushev:1997:BDD

[Bau97] A. N. Baushev. On the boundedness of the density of the distribution of the norm of a Gaussian random vector. Theory of Probability and its Applications, 41(2):334–340, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 403–409.

Bavarov:1961:TRS

[BB61] E. A. Bavarov and P. F. Belyaev. On testing the randomness of sampling without replacement. Theory of Probability and its Applications, 6(4):419–422, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:1981:ABN

[BB81] I. S. Borisov and A. A. Borovkov. Asymptotic behavior of the number of free servers for systems with refusals. Theory of Probability and its Applications, 25 (3):439–453, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:1987:SOA

[BB87] I. S. Borisov and A. A. Borovkov. Second-order approximation of random polygonal lines in the Donsker–Prokhorov invariance principle. Theory of Probability and its Applications, 31(2): 179–202, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:1991:PSD

[BB91] M. V. Burnashev and I. A. Begmatov. On a problem of signal detection leading to stable distributions. Theory of Probability and its Applications, 35(3): 556–560, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 557–560.

Bakhvalov:1994:MNN

[BBD^{+}94] N. S. Bakhvalov, A. A. Borovkov, R. L. Dobrushin, A. V. Zabrodin, V. M. Zolotarev, I. A. Ibragimov, Yu. V. Prokhorov, B. A. Sevast′yanov, Ya. G. Sinaĭ, A. V. Skorokhod, V. A. Statulyavichus, R. Z. Khas′minskiĭ, A. S. Kholevo, D. M. Chibisov, and A. N. Shiryaev. Memorial: Nikolai Nikolaevitch Chentsov. Theory of Probability and its Applications, 38(3):506–515, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 613–623.

Babkin:1959:SRG

[BBM59] V. I. Babkin, P. F. Belyaev, and Yu. I. Maksimov. Some remarks on Goncharov’s paper from the domain of combinatorics. Theory of Probability and its Applications, 4(4):409–414, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bai:1992:SCM

[BCK^{+}92] Z. D. Bai, X. R. Chen, P. R. Krishnaiah, Y. H. Wu, and L. C. Zhao. Strong consistency of maximum likelihood parameter estimation of superimposed exponential signals in noise. Theory of Probability and its Applications, 36(2):349–355, May 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 392–397.

Balkema:1978:LDO

[BD78] A. A. Balkema and L. De Haan. Limit distributions for order statistics. I. Theory of Probability and its Applications, 23(1):77–92, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Beinicke:1983:PED

[BD83] G. Beinicke and K. O. Dzhaparidze. On parameter estimation by the Davidon–Fletcher– Powell method. Theory of Probability and its Applications, 27 (2):396–402, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Brodskii:1991:AAS

[BD91a] B. E. Brodskiĭ and B. S. Darkhovskiĭ. Asymptotic analysis of some estimates in the A posteriori “Disorder” problem. Theory of Probability and its Applications, 35(3):550–556, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 551–557.

Brodskii:1991:CAS

[BD91b] B. E. Brodskiĭ and B. S. Darkhovskiĭ. Comparative analysis of some nonparametric methods of fastest detection of the moment of “Disorder” of a random sequence. Theory of Probability and its Applications, 35 (4):639–652, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 655–668.

Ball:1993:BPA

[BD93] F. Ball and P. Donnelly. Branching process approximation of epidemic models. Theory of Probability and its Applications, 37 (1):119–121, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 144–147.

Badalbaev:1997:ABN

[BD97a] I. S. Badalbaev and A. V.

Dryakhlov. The asymptotic behavior of nonextinction probability of a branching process with pairwise interaction of particles. Theory of Probability and its Applications, 41(4):617– 631, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97566.

Basu:1997:MMM

[BD97b] S. Basu and A. DasGupta. The mean, median, and mode of unimodal distributions: a characterization. Theory of Probability and its Applications, 41(2):210–223, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97544. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 336–352.

Bretagnolle:1968:PRP

[BDC68] J. Bretagnolle and D. Dacunha-Castelle. Pointwise recurrence of processes with independent increments. Theory of Probability and its Applications, 13 (4):677–682, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Beirlant:1992:BKT

[BDEM92] J. Beirlant, P. Deheuvels, J. H. J. Einmahl, and D. M. Mason. Bahadur-kiefer theorems for uniform spacings processes. Theory of Probability and its Applications, 36(4):647–669, December 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 724–743.

Balkema:1979:LDO

[BdH79] A. A. Balkema and L. de Haan.

Limit distributions for order statistics. II. Theory of Probability and its Applications, 23 (2):341–358, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Begovatov:1969:SIG

[Beg69] E. A. Begovatov. A statistical integral for a Gaussian process. Theory of Probability and its Applications, 14(2):354–356, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Beilinson:1964:SPN

[Bei64] A. A. Beilinson. On the solution of a probabilistic non-equilibrium problem for distributed systems. Theory of Probability and its Applications, 9(3):469–472, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Beilinson:1965:EDS

[Bei65] A. A. Beilinson. On the evolution of distributed systems in the presence of density fluctuations on the boundary. Theory of Probability and its Applications, 10(4):668–673, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1958:USF

[Bel58] Yu. K. Belyaev. On the unboundedness of the sample functions of Gaussian processes. Theory of Probability and its Applications, 3(3):327–329, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1959:ARP

[Bel59] Yu. K. Belyaev. Analytic random processes. Theory of Probability and its Applications, 4 (4):402–409, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1960:LPS

[Bel60] Yu. K. Belyaev. Local properties of the sample functions of stationary Gaussian processes. Theory of Probability and its Applications, 5(1):117–120, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1961:EPM

[Bel61] Yu. K. Belyaev. An example of a process with mixing. Theory of Probability and its Applications, 6(1):93–94, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1963:LTD

[Bel63] Y. K. Belyaev. Limit theorems for dissipative flows. Theory of Probability and its Applications, 8(2):165–173, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1964:PNO

[Bel64] P. F. Belyaev. The probability of non-occurrence of a given number of outcomes. Theory of Probability and its Applications, 9(3):489–494, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1965:PNA

[Bel65] P. F. Belyaev. On the probability of the non-appearance of a given number of s-tuples in compound Markov chains. Theory of Probability and its Applications, 10(3):496–499, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1966:NIL

[Bel66] Yu. K. Belyaev. On the number of intersections of a level by a Gaussian stochastic process. I. Theory of Probability and its Applications, 11(1):106–113, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1967:DSM

[Bel67a] P. F. Belyaev. On the distribution of the statistic of the most powerful test for two simple hypotheses on a compound Markov chain. Theory of Probability and its Applications, 12(4):678–682, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1967:NLC

[Bel67b] Yu. K. Belyaev. On the number of level crossings by a Gaussian stochastic process, II. Theory of Probability and its Applications, 12(3):392–404, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1968:NEA

[Bel68] Yu. K. Belyaev. On the number of exits across the boundary of a region by a vector stochastic process. Theory of Probability and its Applications, 13(2):320–324, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1969:JDF

[Bel69a] P. F. Belyaev. On the joint distribution of frequencies of long scycles in a multinomial scheme with equiprobable events. Theory of Probability and its Applications, 14(3):517–521, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1969:PJD

[Bel69b] P. F. Belyaev. On the problem of the joint distribution of frequencies of s-chains in compound Markov chains with a large number of states. Theory of Probability and its Applications, 14 (2):324–330, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1978:JFD

[Bel78] P. F. Belyaev. On the joint frequency distribution of outcomes in Markov chains with a large number of states. Theory of Probability and its Applications, 22(3):521–532, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belavkin:1994:QDM

[Bel94] V. P. Belavkin. Quantum diffusion, measurement and filtering. I. Theory of Probability and its Applications, 38(4):573–585, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 742–757.

Belavkin:1995:MFQ

[Bel95] V. P. Belavkin. Measurement and filtering of quantum diffusion II. Theory of Probability and its Applications, 39(3):363–378, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 469–487.

Benes:1962:MMS

[Ben62] V. E. Benes. Martingales on metric spaces. Theory of Probability and its Applications, 7(1):81–82, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bendikov:1976:CCO

[Ben76a] A. D. Bendikov. A continuity criterion for one class of Markov processes. Theory of Probability and its Applications, 21(1): 169–171, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bendikov:1976:FHC

[Ben76b] A. D. Bendikov. On functions that are harmonic for a certain class of diffusion processes on a group. Theory of Probability and its Applications, 20(4):759–769, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bentkus:1987:LDB

[Ben87] V. Yu. Bentkus. Large deviations in Banach spaces. Theory of Probability and its Applications, 31(4):627–632, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bergstrom:1970:ACM

[Ber70a] H. Bergström. Addendum: a comparison method for distribution functions of sums of independent and dependent random variables. Theory of Probability and its Applications, 15(4):727, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bergstrom:1970:CMD

[Ber70b] H. Bergström. A comparison method for distribution functions of sums of independent and dependent random variables. Theory of Probability and its Applications, 15(3):430–457, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernstein:1974:APG

[Ber74] A. V. Bernstein. Asymptotic properties of group estimates of a random collection. Theory of Probability and its Applications, 19(1):212–215, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernstein:1976:OAT

[Ber76] A. B. Bernstein. On optimal asymptotic tests for composite hypotheses under non-standard conditions. Theory of Probability and its Applications, 21 (1):34–47, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1978:OAT

[Ber78] A. B. Bernshtein. On optimal asymptotic tests of homogeneity. Theory of Probability and its Applications, 22(2):377–383, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1980:CMT

[Ber80] A. V. Bernshtein. On the construction of majorizing tests. Theory of Probability and its Applications, 25(1):16–26, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1981:VCH

[Ber81] A. V. Bernshtein. On verifying composite hypotheses with nuisance parameters in the multivariate case. Theory of Probability and its Applications, 25 (2):287–298, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1984:SCA

[Ber84] A. V. Bernshtein. The structure of the class of absolutely admissible tests. Theory of Probability and its Applications, 28 (2):426–432, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1985:UCP

[Ber85] A. V. Bernshtein. Uniformity classes in problems of testing composite hypotheses. Theory of Probability and its Applications, 29(4):823–827, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1986:ACC

[Ber86a] A. V. Bernshtein. On asymptotically complete classes of tests in the problem of verification of composite hypotheses under contiguous alternatives. Theory of Probability and its Applications, 30(1):87–102, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1986:RTC

[Ber86b] A. V. Bernshtein. A refinement of theorems on complete classes of tests. Theory of Probability and its Applications, 30(3):613–617, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernshtein:1987:ACC

[Ber87] A. V. Bernshtein. Asymptotically complete classes of tests in the multivariate case with asymptotic expansions. Theory of Probability and its Applications, 31(1):58–71, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bethmann:1986:CPM

[Bet86] J. Bethmann. Characterization of power-mixtures by means of their one-sided asymptotic behavior. Theory of Probability and its Applications, 30(1):153–157, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bethmann:1988:LBS

[Bet88] J. Bethmann. Limiting behavior of sums of a random number of independent and identically distributed random variables. Theory of Probability and its Applications, 32(3):562–565, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bethmann:1989:LFT

[Bet89] J. Bethmann. The Lindberg– Feller theorem for sums of a random number of independent random variables in a triangular array. Theory of Probability and its Applications, 33(2):334–339, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 354–359.

Bevastyanov:1958:BSP

[Bev58] B. A. Bevast’yanov. Branching stochastic processes for particles diffusing in a bounded domain with absorbing boundaries. Theory of Probability and its Applications, 3(2):111–126, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bezhaeva:1971:LTC

[Bez71] Z. I. Bezhaeva. Limit theorems for conditional Markov chains. Theory of Probability and its Applications, 16(3):429–437, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bezaeva:1975:EPC

[Bez75] Z. I. Bezaeva. Ergodic properties of conditional Markov chains. Theory of Probability and its Applications, 19(3):522–531, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bezaeva:1978:CMC

[Bez78]  Z. I. Bezaeva.  Conditional

Markov chains with a countable set of states. Theory of Probability and its Applications, 22 (3):543–553, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bezhaeva:1980:ACM

[Bez80] Z. I. Bezhaeva. Addendum: Conditional Markov chains with a countable set of states. Theory of Probability and its Applications, 24(2):448–449, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bezhaeva:1982:WCM

[Bez82] Z. I. Bezhaeva. Weak convergence of matrices of transition probabilities for conditional Markov chains. Theory of Probability and its Applications, 27 (1):59–68, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Brodskii:1973:COW

[BG73] V. Z. Brodskii and T. I. Golikova. Construction of Doptimal weighing designs with minimal number of observations. Theory of Probability and its Applications, 17(3):547–551, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blekher:1991:PPI

[BG91] P. M. Blekher and N. N.

Ganikhodzhaev. On pure phases of the Ising model on the Bethe lattices. Theory of Probability and its Applications, 35(2): 216–227, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 220–230.

Bentkus:1996:MMA

[BG96] V. Bentkus and F. Götze. On minimal moment assumptions in Berry-Esséen theorems for Ustatistics. Theory of Probability and its Applications, 40(3): 430–445, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 596–614.

Bobkov:1998:MP

[BG98] S. G. Bobkov and F. Götze. On moments of polynomials. Theory of Probability and its Applications, 42(3):518–520, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 638–640.

Bolshev:1961:TCD

[BGS61] L. N. Bol’shev, B. V. Gladkov, and M. V. Shcheglova. Tables for the calculation of B- and Zdistribution functions. Theory of Probability and its Applications, 6(4):410–419, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1990:MRT

[BGS90] Yu. K. Belyaev, B. V. Gnedenko, and A. D. Solov’ev. Mathematical reliability-theoretical research at the probability theory department. Theory of Probability and its Applications, 34(1):165–169, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Beck:1975:RDT

[BGW75] Anatole Beck, Daniel P. Giesy, and Peter Warren. Recent developments in the theory of strong laws of large numbers for vectorvalued random variables. Theory of Probability and its Applications, 20(1):127–134, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bentkus:1998:AQF

[BGZ98] V. Bentkus, F. Götze, and A. Yu. Zaitsev. Approximation of quadratic forms of independent random vectors by accompanying laws. Theory of Probability and its Applications, 42(2):189–212, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97613. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(2), (1997), pp. 308–335.

Bhattacharya:1970:RWC

[Bha70] R. N. Bhattacharya. Rates of weak convergence for the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 15(1):68– 86, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Biggins:1993:MCL

[Big93] J. D. Biggins. Martingale convergence and large deviations in the branching random walk. Theory of Probability and its Applications, 37(2):269–273, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 301–306.

Bikyalis:1969:AEP

[Bik69a] A. Bikyalis. An asymptotic expansion of the product of multivariate characteristic functions. Theory of Probability and its Applications, 14(3):490–492, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bikyalis:1969:AED

[Bik69b] A. Bikyalis. Asymptotic expansions for distributions of sums of identically distributed independent lattice random variables. Theory of Probability and its Applications, 14(3):481–489, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bingham:1990:WKS

[Bin90] N. H. Bingham. The work of A. N. Kolmogorov on strong limit theorems. Theory of Probability and its Applications, 34(1): 129–139, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 34(1), (1989), pp. 152–164.

Berzin-Joseph:1998:WCI

[BJL98] C. Berzin-Joseph and J. R. León.

Weak convergence of the integrated number of level crossings to the local time for Wiener processes. Theory of Probability and its Applications, 42(4): 568–579, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97649.

Borovkov:1965:RAA

[BK65] A. A. Borovkov and V. S. Korolyuk. On the results of asymptotic analysis in problems with boundaries. Theory of Probability and its Applications, 10 (2):236–246, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blyumin:1971:BI

[BK71] S. L. Blyumin and B. D. Kotlyar.

On the S. N. Bernstein inequality. Theory of Probability and its Applications, 16(1):182–184, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinski:1980:LSE

[BK80] A. V. Bulinski and A. N. Kolmogorov. Linear sampling estimates of sums. Theory of Probability and its Applications, 24 (2):241–252, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bratiichuk:1986:RHP

[BK86] N. S. Bratiichuk and V. S. Korolyuk. The resolvent of a homogeneous process with independent increments stopped on a halfline. Theory of Probability and its Applications, 30 (2):395–398, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Braverman:1989:IAP

[BK89] M. Sh. Braverman and V. M. Kruglov. An inequality for absolute pseudomoments. Theory of Probability and its Applications, 33(4):718–722, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 773–777.

Bryandinskaya:1993:E

[BK93] A. A. Bryandinskaya and M. V. Khatunseva. Erratum. Theory of Probability and its Applications, 37(2):404, ???? 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1997:LDP

[BK97] A. A. Borovkov and D. A. Korshunov. Large-deviation probabilities for one-dimensional Markov chains. part 1: Stationary distributions. Theory of Probability and its Applications, 41(1):1–24, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97500.

Original Russian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 3–30.

Brovkovich:1963:RCE

[BL63] G. N. Brovkovich and Yu. V.

Linnik. A remark on confidence estimation by the method of least squares. Theory of Probability and its Applications, 8 (2):200–202, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1966:IEP

[BL66] L. N. Bol’shev and E. A. Loginov.

Interval estimates in the presence of nuisance parameters. Theory of Probability and its Applications, 11(1):82–94, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Beer:1975:STC

[BL75] S. Beer and E. Lukacs. Stability theorems for a characterization of the Poisson distribution. Theory of Probability and its Applications, 19(4):656–668, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blagoveshchenskii:1962:DPD

[Bla62] Yu. N. Blagoveshchenskii. Diffusion processes depending on a small parameter. Theory of Probability and its Applications, 7 (2):130–146, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blagoveshchenskii:1964:CPD

[Bla64] Yu. N. Blagoveshchenskii. Cauchy’s problem for degenerate quasilinear parabolic equations. Theory of Probability and its Applications, 9(2):342–346, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blank:1983:UCF

[Bla83] N. M. Blank. On uniqueness conditions for functions of bounded variation and distribution functions with values on the halfline. Theory of Probability and its Applications, 27(4):844–847, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blank:1989:DPS

[Bla89] M. L. Blank. Deterministic properties of stochastically perturbed dynamic systems. Theory of Probability and its Applications, 33(4):612–623, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 659–671.

Bleher:1986:TLD

[Ble86] P. M. Bleher. Theorem on large deviations in a neighborhood of the critical point of the φ^{4}-hierarchical model. Theory of Probability and its Applications, 30(3):529–541, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1963:APG

[BM63] Yu. K. Belyaev and W. M. Maksimov. Analytic properties of a generating function for the number of renewals. Theory of Probability and its Applications, 8 (1):102–105, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1979:CSG

[BM79] L. N. Bol’shev and M. Mirvaliev. Chi square goodness-of-fit test for the Poisson, binomial and negative binomial distributions. Theory of Probability and its Applications, 23(3):461–474, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Badalbaev:1990:LTS

[BM90] I. S. Badalbaev and A. Mukhitdinov. Limit theorems for some functionals in critical multitype branching processes. Theory of Probability and its Applications, 34(4):690–694, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 753–757.

Badalbaev:1991:LDS

[BM91] I. S. Badalbaev and A. Mukhitdinov. On the limit distributions of some functionals in multi-type branching processes. Theory of Probability and its Applications, 35(4):625–638, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 642–654.

Bulinskii:1992:ANS

[BM92] A. V. Bulinskiĭ and S. A. Molchanov. Asymptotical normality of a solution of Burgers’ equation with random initial data. Theory of Probability and its Applications, 36(2):217–236, May 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 217–235.

Borovkov:1993:LDS

[BM93a] A. A. Borovkov and A. A. Mogul′skiĭ. Large deviations and statistical invariance principle. Theory of Probability and its Applications, 37(1):7–13, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 11–18.

Byambazhav:1993:LTM

[BM93b] D. Byambazhav and A. A. Mogul′skiĭ. A limit theorem for the minimum and minimum point of a “Distorted” function. Theory of Probability and its Applications, 37(3): 514–520, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 37(3), (1992), pp. 548–554.

Borovkov:1999:ILL

[BM99] A. A. Borovkov and A. A.

Mogulskii. Integro-local limit theorems including large deviations for sums of random vectors. Theory of Probability and its Applications, 43(1):1– 12, March 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97662.

Braverman:1996:CPL

[BMS96] M. Braverman, C. Mallows, and L. Shepp. Characterization of probability law by absolute moments of its partial sums. Theory of Probability and its Applications, 40(2):238–249, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belyaev:1969:CEA

[BN69] Yu. K. Belyaev and V. P.

Nosko. Characteristics of excursions above a high level for a Gaussian process and its envelope. Theory of Probability and its Applications, 14(2):296–309, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belopolskaya:1983:CSP

[BN83] Ya. I. Belopol’skaya and Z. I.

Nagolkina. On a class of stochastic partial differential equations.

Theory of Probability and its Applications, 27(3):592–600, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:1994:ALI

[BN94] Ole E. Barndorff-Nielsen. Adjusted likelihood inference about interest parameters. Theory of Probability and its Applications, 38(2):179–193, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 213–232.

Belyaev:1996:LTC

[BN96] K. P. Belyaev and L. V. Nazarov. Limit theorems for characteristics of a queueing system with batch processing. Theory of Probability and its Applications, 40(4):744–747, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:1996:CNE

[BNK96] O. E. Barndorff-Nielsen and A. E. Koudou. Cuts in natural exponential families. Theory of Probability and its Applications, 40(2):220–229, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 361–372.

Barndorff-Nielsen:1966:DNA

[BNS66] O. Barndorff-Nielsen and M. Sobel. On the distribution of the number of admissible points in a vector random sample. Theory of Probability and its Applications, 11(2):249–269, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bagdasarov:1996:RCI

[BO96a] D. R. Bagdasarov and E. I. Ostrovskiĭ. Reversion of Chebyshev’s inequality. Theory of Probability and its Applications, 40 (4):737–742, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 873–879.

Bagdasarova:1996:NEE

[BO96b] I. R. Bagdasarova and E. I. Ostrovskiĭ. Nonuniform exponential estimates for large deviations in Banach space. Theory of Probability and its Applications, 40 (3):538–542, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 638–642.

Bezhaeva:1997:VSF

[BO97] Z. I. Bezhaeva and V. I. Oseledets. On the variance of sums for functions of a stationary Markov process. Theory of Probability and its Applications, 41(3):537–541, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 41(3), (1996), pp. 633–639.

Bagdasarov:1998:ECI

[BO98] D. R. Bagdasarov and E. I. Ostrovskii. Exponential confidence intervals in nonparametric density estimation. Theory of Probability and its Applications, 42(4): 684–689, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97653.

Bobrov:1964:IFV

[Bob64] S. P. Bobrov. An investigation of the free verse in Pushkin’s “Songs of the Western Slavs”. Theory of Probability and its Applications, 9(2):236–247, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boganik:1967:DOP

[Bog67] G. N. Boganik. On the determination of the order of a parabolic regression equation. Theory of Probability and its Applications, 12(4):658–666, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogomolov:1980:CRU

[Bog80] N. A. Bogomolov. Confidence regions for unimodal and symmetric distribution functions. Theory of Probability and its Applications, 24(2):419–423, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogomolov:1982:MMG

[Bog82] N. A. Bogomolov. Minimax measurements in a general statistical decision theory. Theory of Probability and its Applications, 26(4):787–795, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:1989:SDM

[Bog89]  V. I. Bogachëv.  On Sko-

rokhod differentiability of measures. Theory of Probability and its Applications, 33(2):330–334, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 349–354.

Bolshev:1957:NCP

[Bol57] L. N. Bol’shev. A nomogram connecting the parameters of a normal distribution with probabilities for classification into three groups. Theory of Probability and its Applications, 2(1):120–122, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1959:TRV

[Bol59] L. N. Bol’shev. On transformations of random variables. Theory of Probability and its Applications, 4(2):129–141, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1960:EP

[Bol60] L. N. Bol’shev. On estimates of probabilities. Theory of Probability and its Applications, 5 (4):411–415, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1961:RCR

[Bol61] L. N. Bol’shev. A refinement of the Cramér–Rao inequality. Theory of Probability and its Applications, 6(3):295–301, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1962:CPP

[Bol62] L. N. Bol’shev. Comparison of parameters of Poisson distributions. Theory of Probability and its Applications, 7(1):113–114, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1963:APT

[Bol63] L. N. Bol’shev. Asymptotically Pearson transformations. Theory of Probability and its Applications, 8(2):121–146, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1964:DRH

[Bol64a] L. N. Bol’shev. Distributions related to the hypergeometric distribution. Theory of Probability and its Applications, 9(4): 619–624, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1964:MKR

[Bol64b] L. N. Bol’shev. On M. Kadyrov’s random numbers in J. Janko’s statistical tables. Theory of Probability and its Applications, 9 (1):136–138, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1965:CPD

[Bol65a] L. N. Bol’shev. On a characterization of the Poisson distribution and its statistical applications. Theory of Probability and its Applications, 10(3):446–456, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1965:CCL

[Bol65b] L. N. Bol’shev. On the construction of confidence limits. Theory of Probability and its Applications, 10(1):173–177, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1966:QTE

[Bol66] L. N. Bol’shev. On the question of testing for “Exponentiality”. Theory of Probability and its Applications, 11(3):480–482, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolotnikov:1968:LPM

[Bol68] Yu. V. Bolotnikov. Limit processes in a model of distribution of particles into cells with unequal probabilities. Theory of Probability and its Applications, 13(3):504–511, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1981:SMS

[Bol81] L. N. Bol’shev. Seminar on mathematical statistics commemorating. Theory of Probability and its Applications, 25(3):634–649, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boldin:1983:EDN

[Bol83] M. V. Boldin. Estimation of the distribution of noise in an autoregression scheme. Theory of Probability and its Applications, 27(4):866–871, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boldin:1990:THS

[Bol90] M. V. Boldin. On testing hypotheses in the sliding average scheme by the Kolmogorovsmirnov and ω^{2} tests. Theory of Probability and its Applications, 34(4):699–704, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 758–764.

Boldin:1994:DRS

[Bol94] M. B. Boldin. On the distribution of run statistics in a linear regression scheme. Theory of Probability and its Applications, 38(1):129–136, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 154–161.

Boldin:1997:LAD

[Bol97] M. V. Boldin. On the least absolute deviation estimator in nonstationary autoregression, and testing for stationarity. Theory of Probability and its Applications, 41(2):340–348, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 409–417.

Bondesson:1974:CGD

[Bon74] Lennart Bondesson. Characterizations of the gamma distributions. Theory of Probability and its Applications, 18(2):367–369, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bondarev:1992:ENP

[Bon92] B. V. Bondarev. Estimation of nonlinearity parameter in initialboundary value problems with Gaussian perturbations. Theory of Probability and its Applications, 36(3):570–578, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 553–560.

Bondarev:1998:KHR

[Bon98]  B. V. Bondarev.  On the

Kolmogorov–Hajek–Rényi inequality for normed integrals of weakly dependent processes. Theory of Probability and its Applications, 42(2):213–224, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97608.

Borovkov:1960:LTDb

[Bor60a] A. A. Borovkov. Limit theorems on the distribution of maximum of sums of bounded, lattice random variables. II. Theory of Probability and its Applications, 5(4):341–355, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1960:LTDa

[Bor60b] A. A. Borovkov. Limit theorems on the distributions of maxima of sums of bounded lattice random variables. I. Theory of Probability and its Applications, 5(2):125–155, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1961:LTM

[Bor61] A. A. Borovkov. Local theorems and moments for maxima of sums of bounded lattice components. Theory of Probability and its Applications, 6 (1):99–101, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1962:STN

[Bor62] A. A. Borovkov. Some theorems on non-latticed random walk. Theory of Probability and its Applications, 7(2):164–179, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1963:DSS

[Bor63] A. A. Borovkov. On discrete service systems. Theory of Probability and its Applications, 8 (3):232–246, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1964:RWB

[Bor64a] A. A. Borovkov. Remarks on Wiener’s and Blackwell’s theorems. Theory of Probability and its Applications, 9(2):303–312, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1964:SLT

[Bor64b] A. A. Borovkov. Some limit theorems in the theory of mass service. Theory of Probability and its Applications, 9(4):550–565, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1965:CBP

[Bor65a] A. A. Borovkov. Correction to A. A. Borovkov’s paper “Some Limit Theorems in the Theory of Mass Service I”, Volume 9, Number 4, 1964. Theory of Probability and its Applications, 10(3):400, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1965:FPT

[Bor65b] A. A. Borovkov. On the first passage time for one class of processes with independent increments. Theory of Probability and its Applications, 10(2):331–334, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1965:SLT

[Bor65c] A. A. Borovkov. Some limit theorems in the theory of mass service, II multiple channels systems. Theory of Probability and its Applications, 10(3):375–400, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1966:AAS

[Bor66] A. A. Borovkov. Asymptotic analysis of service systems. Theory of Probability and its Applications, 11(4):596–602, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1967:BVP

[Bor67a] A. A. Borovkov. Boundaryvalue problems for random walks and large deviations in function spaces. Theory of Probability and its Applications, 12(4):575–595, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1967:CWD

[Bor67b] A. A. Borovkov. Convergence of weakly dependent processes to the Wiener process. Theory of Probability and its Applications, 12(2):159–186, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1967:CDP

[Bor67c]  A. A. Borovkov.  On the

convergence to diffusion processes. Theory of Probability and its Applications, 12(3):405–431, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1968:SPM

[Bor68a] A. A. Borovkov. On service processes in multichannel systems. Theory of Probability and its Applications, 13(1):179–181, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1968:SIS

[Bor68b] A. A. Borovkov. Some inequalities for sums of multidimensional random variables. Theory of Probability and its Applications, 13(1):156–160, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1970:FIP

[Bor70] A. A. Borovkov. Factorization identities and properties of the distribution of the supremum of sequential sums. Theory of Probability and its Applications, 15 (3):359–402, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1971:PPR

[Bor71] A. A. Borovkov. On the problem of pattern recognition. Theory of Probability and its Applications, 16(1):141–144, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1972:SPS

[Bor72] A. A. Borovkov. Some properties of the supremum of sums of stationary related random variables. Theory of Probability and its Applications, 17(1):149–151, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1973:CTM

[Bor73a] A. A. Borovkov. Continuity theorems for multichannel systems with refusals. Theory of Probability and its Applications, 17 (3):434–444, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1973:NIS

[Bor73b] A. A. Borovkov. Notes on inequalities for sums of independent variables. Theory of Probability and its Applications, 17 (3):556–557, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1974:CCD

[Bor74a] A. A. Borovkov. Conditions for convergence to degenerate processes. Theory of Probability and its Applications, 18(3):423–430, ???? 1974. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1974:RCI

[Bor74b] A. A. Borovkov. On the rate of convergence for the invariance principle. Theory of Probability and its Applications, 18 (2):207–225, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1976:AOT

[Bor76] A. A. Borovkov. Asymptotically optimal tests for compound hypotheses. Theory of Probability and its Applications, 20 (3):447–469, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:1977:RCD

[Bor77a] I. S. Borisov. On the rate of convergence of distributions of functionals of integral type. Theory of Probability and its Applications, 21(2):283–299, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borkikh:1977:LE

[Bor77b] Yu. Borkikh. Letter to the Editor. Theory of Probability and its Applications, 21(4):875–876, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1977:PIN

[Bor77c] A. A. Borovkov. On the power of the chi^{2} test with increasing number of class intervals. Theory of Probability and its Applications, 21(??):366–370, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 22, (1977), pp. 375–379.

Borisov:1978:RCC

[Bor78a] I. S. Borisov. On the rate of convergence in the “Conditional” invariance principle. Theory of Probability and its Applications, 23(1):63–76, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1978:LTS

[Bor78b] A. N. Borodin. A limit theorem for solutions of differential equations with random right-hand side. Theory of Probability and its Applications, 22(3):482–497, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 22, (1977), pp. 498– 512.

Borovkov:1978:PTI

[Bor78c] A. A. Borovkov. On the power of the χ^{2} test with increasing number of class intervals. Theory of Probability and its Applications, 22(2):366–370, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1978:SER

[Bor78d] A. A. Borovkov. Some estimates of the rate of convergence in stability theorems. Theory of Probability and its Applications, 22 (4):668–678, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1979:QM

[Bor79a] A. N. Borodin. Quasi-martingales.

Theory of Probability and its Applications, 23(3):637–641, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1979:SAP

[Bor79b] A. N. Borodin. A stochastic approximation procedure in the case of weakly dependent observations. Theory of Probability and its Applications, 24 (1):34–52, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1979:AET

[Bor79c] A. Borovkov. Addendum: Ergodic theorems and stability for walks in a band and their applications. Theory of Probability and its Applications, 24(1):237, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1979:ETS

[Bor79d] A. A. Borovkov. Ergodic theorems and stability for walks in a strip and their application. Theory of Probability and its Applications, 23(4):677–685, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1979:EST

[Bor79e] A. A. Borovkov. Ergodicity and stability theorems for a class of stochastic equations and their applications. Theory of Probability and its Applications, 23 (2):227–247, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1980:SLT

[Bor80a] A. N. Borodin. Some limit theorems for processes with random time. Theory of Probability and its Applications, 24(4):755–771, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1980:ASE

[Bor80b] A. A. Borovkov. Addendum: Some estimates of the rate of convergence in stability theorems. Theory of Probability and its Applications, 24(2):449–450, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:1981:CEB

[Bor81a] I. S. Borisov. On conditions of existence of bounded density for additive functionals on some random processes. Theory of Probability and its Applications, 25 (3):454–465, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1981:STE

[Bor81b] K. A. Borovkov. Stability theorems and estimates of the rate of convergence of the components of factorizations for walks defined on Markov chains. Theory of Probability and its Applications, 25(2):325–334, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1982:ABL

[Bor82a] A. N. Borodin. On the asymptotic behavior of local times of recurrent random walks with finite variance. Theory of Probability and its Applications, 26 (4):758–772, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borokov:1982:RCB

[Bor82b] K. A. Borokov. Rate of convergence in a boundary problem. Theory of Probability and its Applications, 27(1):148–149, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:1983:CGR

[Bor83a] I. S. Borisov. On a criterion for Gaussian random processes to be Markovian. Theory of Probability and its Applications, 27 (4):863–865, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1983:RCI

[Bor83b] K. A. Borovkov. On the rate of convergence in the invariance principle for generalized renewal processes. Theory of Probability and its Applications, 27(3):

461–471, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:1984:RCD

[Bor84a] I. S. Borisov. On the rate of convergence in the Donsker– Prokhorov invariance principle. Theory of Probability and its Applications, 28(2):388–393, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1984:LTS

[Bor84b] A. N. Borodin. Limit theorems for sums of independent random variables defined on a recurrent random walk. Theory of Probability and its Applications, 28 (1):105–121, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borodin:1985:ABL

[Bor85a] A. N. Borodin. The asymptotic behavior of local times of recurrent random walks with infinite variance. Theory of Probability and its Applications, 29 (2):318–333, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1985:CRI

[Bor85b] A. A. Borovkov. On the convergence rate in the invariance principle. Theory of Probability and its Applications, 29(3):550–553, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1985:NCS

[Bor85c] K. A. Borovkov. A note on the convergence speed in stability theorems. Theory of Probability and its Applications, 29 (1):120–121, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovskikh:1985:ECF

[Bor85d] Yu. V. Borovskikh. Estimates of characteristic functions of certain random variables with application to ω^{2}-statistics. I. Theory of Probability and its Applications, 29(3):488–503, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1986:CBP

[Bor86a] K. A. Borovkov. On the convergence of branching processes to a diffusion process. Theory of Probability and its Applications, 30(3):496–506, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovskikh:1986:ECF

[Bor86b] Yu. V. Borovskikh. Estimates of characteristic functions of some random variables with applications to ω^{2}-statistics. II. Theory of Probability and its Applications, 30(1):117–127, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1987:LTQ

[Bor87] A. A. Borovkov. Limit theorems for queueing networks. I. Theory of Probability and its Applications, 31(3):413–427, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovikov:1988:LTS

[Bor88a] V. P. Borovikov. Limit theorems for statistics that are partial sums of functions of spacings. Theory of Probability and its Applications, 32(1):86– 97, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1988:LTQ

[Bor88b] A. A. Borovkov. Limit theorems for queueing networks. II. Theory of Probability and its Applications, 32(2):257–272, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1989:ESS

[Bor89a] A. A. Borovkov. On the ergodicity and stability of the sequence w_{n+1} = f(w_{n}, ξ_{n}): Applications to communication networks. Theory of Probability and its Applications, 33(4):595–611, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 641–658.

Borovkov:1989:MPL

[Bor89b] K. A. Borovkov. A method of proving limit theorems for branching processes. Theory of Probability and its Applications, 33(1):105–113, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1989:RPA

[Bor89c] K. A. Borovkov. Refinement of Poisson approximation. Theory of Probability and its Applications, 33(2):343–347, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 364–368.

Borodin:1990:DFBb

[Bor90a] A. N. Borodin. Distributions of functionals of Brownian local time, II. Theory of Probability and its Applications, 34 (4):576–590, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 636–649.

Borodin:1990:DFBa

[Bor90b] A. N. Borodin. Distributions of functionals of the Brownian local time I. Theory of Probability and its Applications, 34(3):385–401, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 433–450.

Borovkov:1991:ESM

[Bor91a] A. A. Borovkov. Ergodicity and stability of multidimensional Markov chains. Theory of Probability and its Applications, 35(3): 542–546, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 543–547.

Borovkov:1991:FFE

[Bor91b] K. A. Borovkov. A functional form of the Erdős–Renyi Law of Large Numbers. Theory of Probability and its Applications, 35 (4):762–766, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 758–762.

Borovkov:1991:CPU

[Bor91c] K. A. Borovkov. On the convergence of projections of uniform distributions on balls. Theory of Probability and its Applications, 35(3):546–550, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 547–551.

Borovkov:1992:LFE

[Bor92a] A. A. Borovkov. Lyapunov functions and ergodicity of multidimensional Markov chains. Theory of Probability and its Applications, 36(1):1–18, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 93–110.

Borovkov:1992:NVM

[Bor92b] K. A. Borovkov. On a new variant of the Monte Carlo method. Theory of Probability and its Applications, 36(2): 355–360, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 342–346.

Borovskikh:1992:S

[Bor92c]  Yu. V. Borovskikh.  UB-

Statistics. Theory of Probability and its Applications, 36(3): 427–444, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 417–433.

Borovikov:1993:EFD

[Bor93] V. P. Borovikov. On estimation of functionals of a density. Theory of Probability and its Applications, 37(3):507–514, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 543–548.

Borovkov:1995:ATE

[Bor95] K. A. Borovkov. Analysis of transient effects for branching processes. Theory of Probability and its Applications, 39(3):379–395, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 449–468.

Borovkov:1996:UEB

[Bor96] A. A. Borovkov. Unimprovable exponential bounds for distributions of sums of a random number of random variables. Theory of Probability and its Applications, 40(2):230–237, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 260–269.

Borovkov:1998:PCQ

[Bor98]  K. A. Borovkov.  Propaga-

tion of chaos for queueing networks. Theory of Probability and its Applications, 42(3):385–394, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 449–460.

Borovkov:1999:AOS

[Bor99] A. A. Borovkov. Asymptotically optimal solutions in the changepoint problem. Theory of Probability and its Applications, 43(4): 539–561, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97711.

Belyaev:1970:GFF

[BP70] Yu. K. Belyaev and V. I. Piterbarg. Generalized flows with finite memory. Theory of Probability and its Applications, 15 (2):204–214, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burtin:1973:SMD

[BP73] Yu. D. Burtin and B. G. Pittel’. Semi-Markov decisions in a problem of optimizing a checking procedure of an unreliable queuing system. Theory of Probability and its Applications, 17 (3):472–492, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1978:CTR

[BP78] V. V. Buldygin and N. A. Pidsukha. Comparison theorems for random series in Banach spaces and some schemes of summation. Theory of Probability and its Applications, 23 (1):22–35, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bendikov:1989:SHF

[BP89a] A. D. Bendikov and I. V. Pavlov.

On spaces of harmonic functions with martingale mixed norm. Theory of Probability and its Applications, 33(4):715–718, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 769–772.

Bendikov:1989:WPR

[BP89b] A. D. Bendikov and I. V. Pavlov. A Wiener process with reflection and harmonic functions with finite energy integral. Theory of Probability and its Applications, 33(3):547–550, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 586–589.

Bairamov:1991:SIC

[BP91] I. G. Baĭramov and Yu. I. Petunin. Structure of invariant confidence intervals containing the main distributed mass. Theory of Probability and its Applications, 35(1):15–26, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 15–26.

Brown:1993:TSM

[BP93] T. C. Brown and D. L. Pallant. Time series models for infinitely divisible processes. Theory of Probability and its Applications, 37(1):122–124, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 147–149.

Bertini:1994:DFC

[BPRS94] L. Bertini, E. Presutti, B. Rüdiger, and E. Saada. Dynamical fluctuations at the critical point: Convergence to a nonlinear stochastic PDE. Theory of Probability and its Applications, 38(4):586–629, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 689–741.

Borovskikh:1997:NAS

[BPS97] Yu. V. Borovskikh, L. Madan Puri, and V. V. Sazonov. Normal approximation of U-statistics in Hilbert space. Theory of Probability and its Applications, 41 (3):405–424, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97519.

Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 481–504.

Borovkov:1964:BVP

[BR64] A. A. Borovkov and B. A. Rogozin. Boundary value problems for some two-dimensional random walks. Theory of Probability and its Applications, 9 (3):361–388, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1965:MDC

[BR65] A. A. Borovkov and B. A. Rogozin. On the multi-dimensional Central Limit Theorem. Theory of Probability and its Applications, 10(1):55–62, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Badalbaev:1979:CBP

[BR79] I. S. Badalbaev and I. Rakhimov. Critical branching processes with immigration of decreasing intensity. Theory of Probability and its Applications, 23(2):259–268, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bentkus:1983:DSS

[BR83] R. Yu. Bentkus and R. A. Rudzkis. On the distribution of some statistical estimates of spectral density. Theory of Probability and its Applications, 27 (4):795–814, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Badalbaev:1984:FRB

[BR84] I. S. Badalbaev and I. Rakhimov. Further results on branching random processes with immigration of decreasing intensity. Theory of Probability and its Applications, 28(4):811–816, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bentkus:1985:EDB

[BR85] V. Yu. Bentkus and A. Rachkauskas. Estimates of the distance between sums of independent random elements in Banach spaces. Theory of Probability and its Applications, 29(1):50–65, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bickel:1994:EEU

[BR94] P. J. Bickel and Y. Ritov. Efficient estimation using both direct and indirect observations. Theory of Probability and its Applications, 38(2):194–213, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 233–258.

Brandstadt:1980:PHG

[Bra80] A. Brandstädt. On a property of homogeneous Gaussian Lfields. Theory of Probability and its Applications, 24(3):602–606, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Braverman:1986:CPN

[Bra86] M. Sh. Braverman. The characteristic properties of normal and stable distributions. Theory of Probability and its Applications, 30(3):465–474, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Branitskaya:1987:LDT

[Bra87] L. L. Branitskaya. On limit distributions of third-degree random symmetric polynomials. Theory of Probability and its Applications, 31(4):715–717, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Braverman:1988:WCP

[Bra88] M. Sh. Braverman. On a way of characterizing probability distributions. Theory of Probability and its Applications, 32 (3):503–507, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Braverman:1990:SSS

[Bra90] M. Sh. Braverman. On symmetric spaces and sequences of independent random variables. Theory of Probability and its Applications, 34(3):506–509, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 561–565.

Bratiichuk:1991:ATJ

[Bra91] N. S. Bratiĭchuk. Arrival time and jump size for a random walk on a Markov chain. Theory of Probability and its Applications, 35(2):228–240, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 231–241.

Bratishchenko:1994:AMA

[Bra94]  V. V. Bratishchenko.  An

autoregressive-moving-average process with integer coefficients on a circle. Theory of Probability and its Applications, 38 (1):136–142, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 162–168.

Breiman:1965:SLT

[Bre65] L. Breiman. On some limit theorems similar to the arc-sin law. Theory of Probability and its Applications, 10(2):323–331, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Berti:1998:WCC

[BRR98] P. Berti, E. Regazzini, and P. Rigo. Well calibrated, coherent forecasting systems. Theory of Probability and its Applications, 42(1):82–102, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97598. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 144–168.

Benes:1968:WIA

[BS68a] V. E. Benes and L. A. Shepp.

Wiener integrals associated with diffusion processes. Theory of Probability and its Applications, 13(3):475–478, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borokov:1968:AON

[BS68b] A. A. Borokov and N. M. Sycheva. On asymptotically optimal non-parametric criteria. Theory of Probability and its Applications, 13(3):359–393, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernstein:1973:ESE

[BS73a] A. V. Bernstein and A. A. Sidorov. Estimates of the set of expectations for a normal population. Theory of Probability and its Applications, 17(4):723–726, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bernstein:1973:CPF

[BS73b] A. V. Bernstein and A. A. Sidorov. On characteristic properties of Fisher’s k-statistics. Theory of Probability and its Applications, 17(4):721–723, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1974:RCR

[BS74] A. A. Borovkov and A. I. Sakhanenko. Remarks on convergence of random processes in non-separable metric spaces and on the non-existence of a Borel measure for processes in C(0, ∞). Theory of Probability and its Applications, 18(4): 774–777, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1981:ERC

[BS81] A. A. Borovkov and A. I. Sakhanenko. On estimates of the rate of convergence in the invariance principle for Banach spaces. Theory of Probability and its Applications, 25(4):721–731, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1987:CPS

[BS87] V. V. Buldygin and S. A. Solntsev. The contraction principle and the strong law of large numbers for weighted sums. Theory of Probability and its Applications, 31(3):452–462, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1988:SLL

[BS88] V. V. Buldygin and S. A. Solntsev. The strong law of large numbers for sums of independent random vectors with operator normalizations and null convergence of Gaussian sequences. Theory of Probability and its Applications, 32(2):243–256, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1989:ESS

[BS89] V. V. Buldygin and S. A. Solntsev. Equivalence of sample and sequential continuity of Gaussian processes and the continuity of Gaussian Markov processes. Theory of Probability and its Applications, 33(4):624–637, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 672–686.

Buldygin:1992:CSO

[BS92] V. V. Buldygin and S. A. Solntsev. On convergence of sums of operator-normed independent random vectors. Theory of Probability and its Applications, 36 (2):360–365, May 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Beirlant:1987:ABH

[BT87] J. Beirlant and J. K. Teugels. The asymptotic behavior of Hill’s estimator. Theory of Probability and its Applications, 31 (3):463–469, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1975:CTC

[BU75] L. N. Bol’shev and M. Ubaidullaeva. Chauvenet’s test in the classical theory of errors. Theory of Probability and its Applications, 19(4):683–692, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1984:IRC

[BU84] A. A. Borovkov and S. A. Utev. On an inequality and a related characterization of the normal distribution. Theory of Probability and its Applications, 28 (2):219–228, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:1994:EDS

[BU94] A. A. Borovkov and S. A. Utev. Estimates for distributions of sums stopped at Markov time. Theory of Probability and its Applications, 38(2):214–225, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 259–272.

Buhler:1968:SRB

[Büh68] W. J. Bühler. Some results on the behavior of branching processes. Theory of Probability and its Applications, 13 (1):52–64, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buiculescu:1982:REF

[Bui82] M. Buiculescu. Representation of excessive functions as potentials of additive functionals for regular Markov processes. Theory of Probability and its Applications, 26(2):385–388, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskaya:1961:MNC

[Bul61] E. V. Bulinskaya. On the mean number of crossings of a level by a stationary Gaussian process. Theory of Probability and its Applications, 6(4):435–438, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskaya:1964:SRC

[Bul64a] E. V. Bulinskaya. Some results concerning optimum inventory policies. Theory of Probability and its Applications, 9 (3):389–403, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskaya:1964:SSS

[Bul64b] E. V. Bulinskaya. Steady-state solutions in problems of optimum inventory control. Theory of Probability and its Applications, 9(3):502–507, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskaya:1967:OIP

[Bul67] E. V. Bulinskaya. Optimum inventory policies with a convex ordering cost function. Theory of Probability and its Applications, 12(1):9–21, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1974:RSB

[Bul74a] V. V. Buldygin. On random series in a Banach space. Theory of Probability and its Applications, 18(3):464–475, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1974:LIR

[Bul74b] V. V. Buldygin. On the Lévy inequality for random variables in a Banach space. Theory of Probability and its Applications, 19(1):156–159, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1975:BMC

[Bul75] A. V. Bulinskii. Bounds for mixed cumulants and higherorder covariances of bounded random variables. Theory of Probability and its Applications, 19(4):835–839, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1977:IBE

[Bul77] V. V. Buldygin. On invariance of Bayes estimates for generalized random variables. Theory of Probability and its Applications, 22(1):172–175, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1978:NLI

[Bul78] A. V. Bulinskii. On normalization in the law of the iterated logarithm. Theory of Probability and its Applications, 22(2):398–399, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1981:RET

[Bul81a] V. V. Buldygin. Random elements and their convergence in topological spaces. Theory of Probability and its Applications, 26(1):85–96, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1981:NVF

[Bul81b] A. V. Bulinskii. A new variant of the functional law of the iterated logarithm. Theory of Probability and its Applications, 25 (3):493–503, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1984:COE

[Bul84] A. V. Bulinskii. On a class of optimal estimators of sums of random variables. Theory of Probability and its Applications, 28 (2):419–425, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buldygin:1985:SPM

[Bul85a] V. V. Buldygin. Supports of probability measures in separable Banach spaces. Theory of Probability and its Applications, 29(3):546–549, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1985:AMW

[Bul85b] A. V. Bulinskii. Analogs of the Mann–Whitney statistic in an experiment with symmetric randomization. Theory of Probability and its Applications, 29 (4):795–798, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulychev:1987:MPE

[Bul87] V. A. Bulychev. The Markov property of an evolutionary stochastic equation. Theory of Probability and its Applications, 31(2):322–327, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1994:IMS

[Bul94]  A. V. Bulinskiĭ.  Inequali-

ties for the moments of sums of associated multi-indexed variables. Theory of Probability and its Applications, 38(2):342–349, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 417–425.

Bulinskii:1996:RCC

[Bul96] A. V. Bulinskiĭ. Rate of convergence in the Central Limit Theorem for fields of associated random variables. Theory of Probability and its Applications, 40 (1):136–144, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 165–174.

Bunke:1963:TDE

[Bun63] Olaf Bunke. Theory of domain estimation of statistical parameters. Theory of Probability and its Applications, 8(3): 309–316, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bunke:1970:NPD

[Bun70] O. Bunke. Non-parametric decision functions with asymptotic minimax optimality. Theory of Probability and its Applications, 15(1):148–153, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bunimovic:1974:CLT

[Bun74]  L. A. Bunimovic.  Central

Limit Theorem for a class of billiards. Theory of Probability and its Applications, 19(1): 65–85, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burtin:1975:EMC

[Bur75a] Yu. D. Burtin. On extreme metric characteristics of a random graph. II. Limit distributions. Theory of Probability and its Applications, 20(1):83–101, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burtin:1975:EMP

[Bur75b] Yu. D. Burtin. On extreme metric parameters of a random graph. I. Asymptotic estimates. Theory of Probability and its Applications, 19(4):710–725, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:1979:MDI

[Bur79] M. V. Burnashev. On the minimax detection of an inaccurately known signal in a white Gaussian noise background. Theory of Probability and its Applications, 24(1):107–119, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:1982:SPR

[Bur82] M. V. Burnashev. On a statistical problem related to random walks. Theory of Probability and its Applications, 26(3):554–563, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:1996:LE

[Bur96] M. V. Burnashev. Letters to the Editors. Theory of Probability and its Applications, 40 (3):595–596, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 699–700.

Burnashev:1997:AEM

[Bur97] M. V. Burnashev. Asymptotic expansions for median estimate of a parameter. Theory of Probability and its Applications, 41 (4):632–645, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97567.

Butsan:1977:ORM

[But77] G. P. Butsan. On the orthogonality, with respect to multiplicative shifts, of white noise defined on a ring of operators. Theory of Probability and its Applications, 21(3):624–628, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Butsan:1979:MPN

[But79] G. P. Butsan. On the mixed product of nonhomogeneous stochastic semi-groups. Theory of Probability and its Applications, 24(1):168–175, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Butov:1992:SEO

[But92a] A. A. Butov. Some estimates for a one-dimensional birth and death process in a random environment. Theory of Probability and its Applications, 36(3):

578–583, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 561–565.

Butsan:1992:MPD

[But92b] G. P. Butsan. Mixed product of dependent operator stochastic systems in the absence of continuity and martingale conditions. Theory of Probability and its Applications, 36(2):237–248, May 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 236–246.

Butov:1995:MMR

[But95] A. A. Butov. Martingale methods for random walks in a one-dimensional random environment. Theory of Probability and its Applications, 39(4):558–572, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 681–698.

Bulinskii:1988:ECR

[BV88] A. V. Bulinskii and N. V. Verdenskaya. An estimate of the convergence rate in the functional law of the iterated logarithm. Theory of Probability and its Applications, 32(2):334–337, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bondarenko:1995:RSE

[BV95] M. V. Bondarenko and A. N.

Vishnyakov. On the Russian stock exchange. Theory of Probability and its Applications, 39 (1):188–194, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulcheva:1990:DEF

[BVT90] O. G. Bulcheva and A. D. Vent-Tsel’. Differentiability of expectations of functionals of a Wiener process. Theory of Probability and its Applications, 34 (3):509–512, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 565–568.

Broniatowski:1998:SLS

[BW98] M. Broniatowski and M. Weber.

Strong laws for sums of extreme values. Theory of Probability and its Applications, 42(3):395–404, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 553–563.

Badalbaev:1996:LTC

[BY96] I. S. Badalbaev and T. D.

Yakubov. Limit theorems for the critical Galton–Watson processes with migration. Theory of Probability and its Applications, 40 (4):599–612, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 833–849.

Bai:1988:LED

[BYK88] Z. D. Bai, Y. Q. Yin, and P. R. Krishnaiah. On the limiting empirical distribution function of the eigenvalues of a multivariate F matrix. Theory of Probability and its Applications, 32 (3):490–500, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:1977:CLT

[BZ77] A. V. Bulinskii and I. G. Zhurbenko. A Central Limit Theorem for additive random functions. Theory of Probability and its Applications, 21(4):687–697, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Badalbaev:1984:LTS

[BZ84] I. S. Badalbaev and A. M. Zubkov. Limit theorems for a sequence of branching processes with immigration. Theory of Probability and its Applications, 28(2):404–409, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bolshev:1970:TCF

[BZKR70] L. N. Bol’shev, V. M. Zolotarev, E. S. Kedrova, and M. A. Rybinskaya. Tables of cumulative functions of one-sided stable distributions. Theory of Probability and its Applications, 15(2):299–309, ???? 1970. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Cadre:1999:FVC

[Cad99] B. Cadre. On a functional version of the convergence of a quadratic form in independent martingales to a χ^{2} distribution. Theory of Probability and its Applications, 43(1):13–25, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97663.

Cairns:1993:EER

[Cai93] A. J. G. Cairns. Epidemic estimation with removal time data. Theory of Probability and its Applications, 37(2):317–319, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Capar:1998:FSM

[Cap98] U. Capar. Fusions of stable measures in Banach spaces. Theory of Probability and its Applications, 42(4):580–588, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97650.

Chiang:1996:GSM

[CC96] T. S. Chiang and Y. Chow. On global and sharp Markov properties of random functional series in Sobolev spaces. Theory of Probability and its Applications, 40(2):348–354, ???? 1996.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 464–471.

Coursol:1982:RAL

[CDC82] J. Coursol and D. Dacunha-Castelle. Remarks on the approximation of the likelihood function of a stationary Gaussian process. Theory of Probability and its Applications, 27(1):162–167, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cheng:1997:ADM

[CdHY97] S. Cheng, L. de Haan, and J. Yang. Asymptotic distributions of multivariate intermediate order statistics. Theory of Probability and its Applications, 41(4):646–656, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97573.

Cuesta:1993:CIR

[CDM93] J. A. Cuesta, J. S. Dominguez, and C. Matrán. On the consistency of the L_{p}-isotonic regression. Theory of Probability and its Applications, 37(1):129–132, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 181–184.

Cambanis:1995:PHT

[CFZ95] S. Cambanis and I. Fakhre-Zakeri. On prediction of heavytailed autoregressive sequences: Forward versus reversed time. Theory of Probability and its Applications, 39(2):217–233, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 294–312.

Ceci:1997:SRA

[CGM97] C. Ceci, A. Gerardi, and L. Mazliak. Some results about stopping times of the marked tree space. Theory of Probability and its Applications, 41(3):425–435, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97520. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 578–590.

Chou:1993:RS

[CH93]  R. Chou and F. Hu.  On

the Z-representation of a statistic. Theory of Probability and its Applications, 37(1):124–126, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 219–221.

Chentsov:1956:WCS

[Che56] N. N. Chentsov. Weak convergence of stochastic processes whose trajectories have no discontinuities of the second kind and the “Heuristic” approach to the Kolmogorov-smirnov tests. Theory of Probability and its Applications, 1(1):140–144, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chentsov:1957:LBM

[Che57a] N. N. Chentsov. Lévy Brownian motion for several parameters and generalized white noise. Theory of Probability and its Applications, 2(2):265–266, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherkasov:1957:TDP

[Che57b] I. D. Cherkasov. On the transformation of the diffusion process to a Wiener process. Theory of Probability and its Applications, 2(3):373–377, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherkasov:1960:CCH

[Che60] I. D. Cherkasov. Criteria for coordinate-homogeneity for continuous Markov processes. Theory of Probability and its Applications, 5(2):209–216, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherenkov:1964:CSS

[Che64a] A. P. Cherenkov. On control by a single switching operation when small random perturbations are present. Theory of Probability and its Applications, 9(2):277–290, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherenkov:1964:PPC

[Che64b] A. P. Cherenkov. A problem of parameter control involving small random perturbations. Theory of Probability and its Applications, 9(3):487–488, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chentsov:1966:STE

[Che66] N. N. Chentsov. A systematic theory of exponential families of probability distributions. Theory of Probability and its Applications, 11(3):425–435, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chentsov:1967:EUM

[Che67a] N. N. Chentsov. On estimating an unknown mean of a multivariate normal distribution. Theory of Probability and its Applications, 12(4):560–574, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherenkov:1967:UIF

[Che67b] A. P. Cherenkov. The use of the integral functional for some control methods in the presence of small random perturbations. Theory of Probability and its Applications, 12(1):139–143, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherenkov:1970:TCC

[Che70] A. P. Cherenkov. The twocomponent case in the control by single switching. Theory of Probability and its Applications, 15 (1):147–148, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chernomordik:1980:NPT

[Che80] O. M. Chernomordik. On a nonparametric test of homogeneity of several samples. Theory of Probability and its Applications, 25 (1):194–197, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chentsov:1981:CPS

[Che81] N. N. Chentsov. On correctness of the problem of statistical point estimation. Theory of Probability and its Applications, 26(1):13–29, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chebotarev:1989:SCR

[Che89] A. M. Chebotarev. Sufficient conditions for the regularity of jump Markov processes. Theory of Probability and its Applications, 33(1):22–35, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherkasov:1996:TOD

[Che96] I. D. Cherkasov. Transformation of one-dimensional diffusion fields in the plane. Theory of Probability and its Applications, 40(3):446–455, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 565–577.

Chistyakov:1957:LLT

[Chi57] V. P. Chistyakov. Local limit theorems in the theory of branching random processes. Theory of Probability and its Applications, 2(3):345–363, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1959:GTB

[Chi59] V. P. Chistyakov. Generalization of a theorem for branching processes. Theory of Probability and its Applications, 4 (1):103–106, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1961:TFB

[Chi61a] D. M. Chibisov. On the tests of fit based on sample spacings. Theory of Probability and its Applications, 6(3):325–329, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1961:TPB

[Chi61b] V. P. Chistyakov. Transient phenomena in branching stochastic processes. Theory of Probability and its Applications, 6(1):27–41, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1964:LDO

[Chi64a] D. M. Chibisov. On limit distributions for order statistics. Theory of Probability and its Applications, 9(1):142–148, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1964:APG

[Chi64b] D. M. Chibisov. On the asymptotic power of goodness-of-fit tests for close alternative hypotheses. Theory of Probability and its Applications, 9(3): 507–509, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1964:CPT

[Chi64c] V. P. Chistyakov. On the calculation of the power of the test of empty boxes. Theory of Probability and its Applications, 9 (4):648–653, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1964:TSI

[Chi64d] V. P. Chistyakov. A theorem on sums of independent positive random variables and its applications to branching random processes. Theory of Probability and its Applications, 9 (4):640–648, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1965:IAP

[Chi65a] D. M. Chibisov. An investigation of the asymptotic power of the tests of fit. Theory of Probability and its Applications, 10 (3):421–437, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1965:ALL

[Chi65b] V. P. Chistyakov. Addendum:

Local limit theorems in the theory of branching stochastic processes. Theory of Probability and its Applications, 10(3):538–539, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1967:TAT

[Chi67] D. M. Chibisov. A theorem on admissible tests and its application to an asymptotic problem of testing hypotheses. Theory of Probability and its Applications, 12(1):90–103, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1970:GTD

[Chi70a] G. P. Chistyakov. A generalization of theorems due to H. Cramér and Yu. V. Linnik–V. P. Skitovic. Theory of Probability and its Applications, 15 (2):331–336, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1970:CBP

[Chi70b] V. P. Chistyakov. On convergence of branching processes to diffusion processes. Theory of Probability and its Applications, 15(4):707–710, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1970:SLTa

[Chi70c] V. P. Chistyakov. Some limit theorems for branching processes with immigration. Theory of Probability and its Applications, 15(2):241–257, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1970:SLTb

[Chi70d] V. P. Chistyakov. Some limit theorems for branching processes with particles of final type. Theory of Probability and its Applications, 15(3):515–521, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1971:CCS

[Chi71a] D. M. Chibisov. Certain chisquare type tests for continuous distributions. Theory of Probability and its Applications, 16 (1):1–22, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1971:ABN

[Chi71b] V. P. Chistyakov. The asymptotic behavior of the nonextinction probability for a critical branching process. Theory of Probability and its Applications, 16(4):620–630, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1972:LTA

[Chi72] V. P. Chistyakov. Limit theorems for age-dependent branching processes. Theory of Probability and its Applications, 17 (1):54–71, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1973:AED

[Chi73a] D. M. Chibisov. An asymptotic expansion for the distribution of a statistic admitting an asymptotic expansion. Theory of Probability and its Applications, 17 (4):620–630, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1973:TPB

[Chi73b] V. P. Chistyakov. On transition phenomena in branching stochastic processes with several types of particles. Theory of Probability and its Applications, 17 (4):631–639, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1974:AEC

[Chi74a] D. M. Chibisov. An asymptotic expansion for a class of estimators containing maximum likelihood estimators. Theory of Probability and its Applications, 18 (2):295–303, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1974:AED

[Chi74b] D. M. Chibisov. An asymptotic expansion for the distribution of sums of a special form with an application to minimumcontrast estimates. Theory of Probability and its Applications, 18(4):649–661, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chitashvili:1976:CFM

[Chi76a] R. Ya. Chitashvili. A controlled finite Markov chain with an arbitrary set of decisions. Theory of Probability and its Applications, 20(4):839–847, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chitashvili:1976:FCM

[Chi76b] R. Ya. Chitashvili. A finite controlled Markov chain with small termination probability. Theory of Probability and its Applications, 21(1):158–163, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1980:ABP

[Chi80a] D. M. Chibisov. Asymptotic behavior of partial densities and their derivatives. Theory of Probability and its Applications, 25 (1):1–15, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1980:AEP

[Chi80b] D. M. Chibisov. Asymptotic expansions for partial moments.

Theory of Probability and its Applications, 24(3):453–466, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1981:AEDa

[Chi81a] D. M. Chibisov. An asymptotic expansion for the distribution of a statistic admitting a stochastic expansion. I. Theory of Probability and its Applications, 25 (4):732–744, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1981:AEDb

[Chi81b] D. M. Chibisov. Asymptotic expansion for the distribution of a statistic admitting a stochastic expansion. II. Theory of Probability and its Applications, 26(1):1–12, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 26, (1981), pp. 3–14.

Chistyakov:1983:MFT

[Chi83] V. P. Chistyakov. Matrix functions of trajectories of Markov chains. Theory of Probability and its Applications, 27(2):248–255, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1986:CDA

[Chi86a] D. M. Chibisov. Calculation of the deficiency of asymptotically efficient tests. Theory of Probability and its Applications, 30 (2):289–310, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chitashvili:1986:EIP

[Chi86b] R. Ya. Chitashvili. On the existence of an innovation process for a component of a diffusiontype process. Theory of Probability and its Applications, 30 (1):50–65, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisova:1987:PBV

[Chi87a] E. D. Chibisova. On a problem of B. V. Gnedenko in a scheme of partial summation. Theory of Probability and its Applications, 31(4):712–715, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1987:DSD

[Chi87b] G. P. Chistyakov. Decompositional stability of distribution laws. Theory of Probability and its Applications, 31(3):375–390, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1987:RPR

[Chi87c] S. P. Chistyakov. On a regression problem for random processes. Theory of Probability and its Applications, 31(1):132–135, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisova:1988:SPS

[Chi88] E. D. Chibisova. On a scheme of partial summation. Theory of Probability and its Applications, 32(3):518–521, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1989:ETF

[Chi89a] D. M. Chibisov. The equivalence of two forms of an asymptotic expansion. Theory of Probability and its Applications, 33 (2):373–377, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1989:PFN

[Chi89b] G. P. Chistyakov. On the problem of α-factorizations of the normal distribution. Theory of Probability and its Applications, 33 (1):152–155, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chikin:1990:FLT

[Chi90] D. O. Chikin. A functional limit theorem for stationary processes: a martingale approach. Theory of Probability and its Applications, 34(4):668–678, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 731–741.

Chistyakov:1995:EOE

[Chi95] G. P. Chistyakov. The exactorder estimates of decompositional stability of Linnik’s L class distributions. Theory of Probability and its Applications, 39 (2):315–322, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 438–446.

Chobanyan:1984:CBS

[Cho84] S. A. Chobanyan. Convergence of Bernoulli series and the set of sums of a conditionally convergent series of functions. Theory of Probability and its Applications, 28(2):442–450, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chorny:1987:OSD

[Cho87] V. Chorny. Operator-semistable distributions on R^{d}. Theory of Probability and its Applications, 31(4):703–705, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chow:1993:ODS

[Cho93] K. L. Chow. A-optimal designs for supplementary experiments. Theory of Probability and its Applications, 37(1):126–129, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 216–219.

Chuprunov:1992:GSS

[Chu92a] A. N. Chuprunov. On the generalization of spaces of stable and Rademacher type P. Theory of Probability and its Applications, 36(3):445–458, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Chuprunov:1992:SSB

[Chu92b] A. N. Chuprunov. Spaces of sequences in a Banach space associated with a sequence of independent random variables. Theory of Probability and its Applications, 36(1):143–149, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 186–191.

Chuprunov:1994:LLN

[Chu94] A. N. Chuprunov. The laws of large numbers in Banach spaces of type (F, F_{1}). Theory of Probability and its Applications, 38 (4):733–737, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 906–909.

Chuprunov:1995:RBV

[Chu95] A. N. Chuprunov. On refinement of Banach-valued limit theorems for stable laws. Theory of Probability and its Applications, 39(4):657–662, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 851–856.

Charnes:1968:ZZC

[CKR68] A. Charnes, M. J. L. Kirby, and W. M. Raike. Zero-zero chanceconstrained games. Theory of Probability and its Applications, 13(4):628–646, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Carlstein:1994:NCP

[CL94] E. Carlstein and S. Lele. Nonparametric change-point estimation for data from an ergodic sequence. Theory of Probability and its Applications, 38(4): 726–733, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 910–917.

Clement:1998:SCB

[Cle98] E. Clement. Short communication: Bias correction for the estimation of discretized diffusion processes from an approximated likelihood function. Theory of Probability and its Applications, 42(2):283–288, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97618.

Carlsson:1987:CSP

[CN87] H. Carlsson and O. Nerman. Continuous singular probability measures that are almost lattice. Theory of Probability and its Applications, 31(2):317–319, ???? 1987. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Cohen:1962:DNM

[Coh62] J. W. Cohen. On derived and nonstationary Markov chains. Theory of Probability and its Applications, 7(4):402–423, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cowan:1993:EPF

[Cow93] R. Cowan. Expected pattern frequency in Markov chain realizations conditional upon transition counts: Application of Whittle’s combinatoric formula to problems in DNA sequences. Theory of Probability and its Applications, 37(2):320–322, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 371–374.

Chrysaphinou:1991:OSP

[CP91] O. Chrysaphinou and S. Papastavridis. The occurrence of sequence patterns in repeated dependent experiments. Theory of Probability and its Applications, 35(1):145–152, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cacoullos:1998:SCV

[CPP98] T. Cacoullos, N. Papadatos, and V. Papathanasiou. Short communication:variance inequalities for covariance kernels and applications to Central Limit Theorems. Theory of Probability and its Applications, 42(1):149–155, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97603. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 195–201.

Cacoullos:1993:ACN

[CPU93] T. Cacoullos, V. Papathanasiou, and S. Utev. Another characterization of the normal law and a proof of the Central Limit Theorem connected with it. Theory of Probability and its Applications, 37(4):581–588, December 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Christoph:1996:DQF

[CPU96] G. Christoph, Yu. V. Prokhorov, and V. V. Ulyanov. On distribution of quadratic forms in Gaussian random variables. Theory of Probability and its Applications, 40(2):250–260, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cramer:1956:TOS

[Cra56] Harald Cramer. A theorem on ordered sets of probability distributions. Theory of Probability and its Applications, 1(1):16–21, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cramer:1964:SPC

[Cra64] Harald Cramér. Stochastic processes as curves in Hilbert space. Theory of Probability and its Applications, 9(2):169–179, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cramer:1965:LTM

[Cra65] H. Cramér. A limit theorem for the maximum values of certain stochastic processes. Theory of Probability and its Applications, 10(1):126–128, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chzhi-Shuen:1960:CMP

[CS60] Lyan Chzhi-Shuén. On conditional Markov processes. Theory of Probability and its Applications, 5(2):208–209, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chi-Son:1961:ISM

[CS61] Leung Chi-Son. Invariability of the strong Markov property in the transformations of Dynkin. Theory of Probability and its Applications, 6(2): 209–211, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Csorgo:1971:CBF

[CS71] M. Csörgö and V. Seshadri.

Characterizations of the Behrens– Fisher and related problems (a goodness of fit point of view).

Theory of Probability and its Applications, 16(1):23–35, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Csorgo:1975:AEL

[Csö75] S. Csörgö. Asymptotic expansion for the Laplace transform of the von Mises ω^{2}-test. Theory of Probability and its Applications, 20(1):158–161, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Csorgo:1984:ECF

[Csö84] S. Csörgö. Estimating characteristic functions under random censorship. Theory of Probability and its Applications, 28 (3):615–623, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:1965:ANP

[CV65] V. P. Chistyakov and I. I. Viktorova. Asymptotic normality in a problem of balls when the probabilities of falling into different boxes are different. Theory of Probability and its Applications, 10(1):149–154, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1985:EEL

[CvZ85] D. M. Chibisov and W. R. van Zwet. On the Edgeworth expansion for the logarithm of the likelihood ratio. I. Theory of Probability and its Applications, 29 (3):427–451, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:1989:AEE

[CvZ89] D. M. Chibisov and W. R. van Zwet. Addendum: On the Edgeworth expansion for the logarithm of the likelihood ratio. I. Theory of Probability and its Applications, 33(1):196, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cowan:1979:OSP

[CZ79] R. Cowan and J. Zabczyk. An optimal selection problem associated with the Poisson process. Theory of Probability and its Applications, 23(3):584–592, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dambis:1965:DCS

[Dam65] K. E. Dambis. On the decomposition of continuous submartingales. Theory of Probability and its Applications, 10(3):401–410, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darling:1960:TKS

[Dar60] D. A. Darling. On the theorems of Kolmogorov-smirnov. Theory of Probability and its Applications, 5(4):356–361, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovskh:1976:NMP

[Dar76] B. S. Darkhovskh. A nonparametric method for the a posteriori detection of the “Disorder” time of a sequence of independent random variables. Theory of Probability and its Applications, 21(1):178–183, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovskii:1985:TEP

[Dar85] B. S. Darkhovskii. On two estimation problems for times of change of the probabilistic characteristics of a random sequence. Theory of Probability and its Applications, 29(3): 478–487, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 29(3), (1984), pp. 464–473.

Darkhovskii:1986:NME

[Dar86] B. S. Darkhovskii. A nonparametric method of estimating intervals of homogeneity for a random sequence. Theory of Probability and its Applications, 30 (4):845–849, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovskii:1996:RDC

[Dar96] B. S. Darkhovskii. Retrospective detection of the change-point in some models of regression type. Theory of Probability and its Applications, 40(4):748–753, ???? 1996. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovskii:1999:SRP

[Dar99] B. S. Darkhovskii. On a stochastic renewal problem. Theory of Probability and its Applications, 43(2):282–288, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97688.

Dashyan:1979:LLT

[Das79] Yu. R. Dashyan. On a local limit theorem for a class of random point fields arising in statistical quantum mechanics. Theory of Probability and its Applications, 23(3):559–572, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dasgupta:1994:CED

[Das94] R. Dasgupta. Cauchy equation on discrete domain and some characterization theorems. Theory of Probability and its Applications, 38(3):520–524, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 661–666.

Daud:1976:LTD

[Dau76] S. S. Daud. A limit theorem for a dynamical system corresponding to a d-dimensional ideal gas. Theory of Probability and its Applications, 20(4):836–839, ???? 1976. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Daugavet:1988:BCR

[Dau88] A. I. Daugavet. Bounds for the convergence rate of numerical characteristics in limit theorems with stable limit law. Theory of Probability and its Applications, 32(3):531–535, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1968:CDG

[Dav68]  Yu. A. Davydov.  Conver-

gence of distributions generated by stationary stochastic processes. Theory of Probability and its Applications, 13(4):691–696, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1970:IPS

[Dav70] Yu. A. Davydov. The invariance principle for stationary processes. Theory of Probability and its Applications, 15(3):487–498, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1974:LTF

[Dav74a] Yu. A. Davydov. Limit theorems for functionals of processes with independent increments. Theory of Probability and its Applications, 18(3):431–441, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1974:MCM

[Dav74b] Yu. A. Davydov. Mixing conditions for Markov chains. Theory of Probability and its Applications, 18(2):312–328, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davis:1976:SD

[Dav76a] M. H. A. Davis. On stochastic differentiation. Theory of Probability and its Applications, 20 (4):869–872, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1976:LTRb

[Dav76b] Yu. A. Davydov. Local times for random fields. Theory of Probability and its Applications, 20 (3):652–654, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 21, (1976), pp. 670–671.

Davydov:1976:LTRa

[Dav76c] Yu. A. Davydov. On local times for random processes. Theory of Probability and its Applications, 21(1):171–178, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1979:LTM

[Dav79] Yu. A. Davydov. Local times for multiparameter random processes. Theory of Probability and its Applications, 23(3):573–583, ???? 1979. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1981:SCD

[Dav81] Yu. A. Davydov. On strong convergence of distributions of functionals of random processes. I. Theory of Probability and its Applications, 25(4):772–789, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davis:1982:PSE

[Dav82a] M. H. A. Davis. A pathwise solution of the equations of nonlinear filtering. Theory of Probability and its Applications, 27 (1):167–175, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1982:SCD

[Dav82b] Yu. A. Davydov. On strong convergence of distributions of functionals of random processes. II. Theory of Probability and its Applications, 26(2):258–278, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1989:LLT

[Dav89a] Yu. A. Davydov. Local limit theorems for functionals of random processes. Theory of Probability and its Applications, 33 (4):732–738, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1989:RAC

[Dav89b] Yu. A. Davydov. A remark on the absolute continuity of distributions of Gaussian functionals. Theory of Probability and its Applications, 33(1):158–161, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1991:DMW

[Dav91] Yu. A. Davydov. On distributions of multiple Wiener–Itô integrals. Theory of Probability and its Applications, 35 (1):27–37, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 51–62.

Darkhovskii:1981:PDD

[DB81] B. S. Darkhovskii and B. E.

Brodskii. A posteriori detection of the “Disorder” time of a random sequence. Theory of Probability and its Applications, 25 (3):624–628, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovskii:1988:NMF

[DB88] B. S. Darkhovskii and B. E.

Brodskii. A nonparametric method for fastest detection of a change in the mean of a random sequence. Theory of Probability and its Applications, 32 (4):640–648, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

DSouza:1993:SVE

[DB93] J. C. D’Souza and J. D. Biggins. The supercritical varyingenvironment Galton–Watson process. Theory of Probability and its Applications, 37(1): 171–172, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 206–207.

DelMoral:1999:MIP

[DD99] P. Del Moral and M. Doisy. Maslov idempotent probability calculus, I. Theory of Probability and its Applications, 43(4): 562–576, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97716.

DeDominicis:1982:WLL

[De 82] R. De Dominicis. Weak law of large numbers for random fields of random elements. Theory of Probability and its Applications, 26(4):820–822, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

deValk:1993:ODP

[de 93] Vinsent de Val′k. One-dependent processes and their Hilbert space representations. Theory of Probability and its Applications, 37 (1):176–179, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 37(1), (1992), pp. 149–153.

Dynkin:1977:RCE

[DE77] E. B. Dynkin and I. V.

Evstigneev. Regular conditional expectations of correspondences. Theory of Probability and its Applications, 21(2):325–338, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

DelMoral:1996:NFU

[Del96] P. Del Moral. Nonlinear filtering using random particles. Theory of Probability and its Applications, 40(4):690–701, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 859–872.

Denisov:1984:RWW

[Den84] I. V. Denisov. A random walk and a Wiener process near a maximum. Theory of Probability and its Applications, 28(4):821–824, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Denisov:1989:IDM

[Den89] L. V. Denisov. Infinitely divisible Markov mappings in quantum probability theory. Theory of Probability and its Applications, 33(2):392–395, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 417–420.

Denisov:1992:CSN

[Den92] L. V. Denisov. Convolution semigroups in noncommutative probability theory. Theory of Probability and its Applications, 36(3): 583–588, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 572–577.

Denisova:1997:ABS

[Den97] I. Yu. Denisova. Asymptotic behavior of the solution of a system of stochastic differential equations. Theory of Probability and its Applications, 41(2):348–352, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 425–429.

Deo:1974:TGR

[Deo74] Chandrakant M. Deo. A theorem on Gaussian random fields. Theory of Probability and its Applications, 18(2):369–372, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Deo:1977:NMR

[Deo77] Chandrakant M. Deo. A note on φ-mixing random fields. Theory of Probability and its Applications, 21(4):867–870, ???? 1977.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dmitrovskii:1984:CLT

[DEO84] V. A. Dmitrovskii, S. V. Ermakov, and E. I. Ostrovskii. The Central Limit Theorem for weakly dependent Banach-valued variables. Theory of Probability and its Applications, 28 (1):89–104, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Derquenne:1993:ODB

[Der93] C. Derquenne. Outlier detection before running statistical methods. Theory of Probability and its Applications, 37(2):323–326, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Devyatov:1969:LAN

[Dev69] B. I. Devyatov. Limits of admissibility of normal approximations to the Poisson distribution. Theory of Probability and its Applications, 14(1):170–173, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Daletskii:1965:GMF

[DF65] Yu. L. Daletskii and S. V. Fomin.

Generalized measures in function spaces. Theory of Probability and its Applications, 10(2):304–316, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Durkhem:1993:CRA

[DF93] S. D. Durkhèm and N. Flournoĭ. Convergence results for an adaptive ordinal urn design. Theory of Probability and its Applications, 37(1):14–17, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 53–56.

deHaan:1995:UCD

[dH95] L. de Haan. A unified criterion for the domain of attraction of extreme-value distributions. Theory of Probability and its Applications, 39(2):323–329, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 415–422.

Dhaparidze:1978:EPS

[Dha78] K. O. Dhaparidze. Estimation of parameters of a spectral density with fixed zeroes. Theory of Probability and its Applications, 22(4):708–729, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1971:ABS

[DI71] Yu. A. Davydov and I. A. Ibragimov. On asymptotic behavior of some functionals of processes with independent increments. Theory of Probability and its Applications, 16(1):162–167, ???? 1971. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Danielyan:1974:QLP

[DI74] E. A. Danielyan and G. A.

Ivanov. On the queue length in priority queuing systems. Theory of Probability and its Applications, 18(4):814–816, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Diesperova:1977:CSD

[Die77] M. M. Diesperova. On the convergence of statistical decisions in problems of the type of a “Series-Scheme”. Theory of Probability and its Applications, 21(2): 379–384, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dinges:1993:WG

[Din93] H. Dinges. Wiener germs. Theory of Probability and its Applications, 37(1):132–136, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 165–170.

Direev:1981:ETA

[Dir81] Yu. V. Direev. An ergodic theorem for additive functionals. Theory of Probability and its Applications, 25(3):614–617, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dittrich:1991:CBP

[Dit91] P. Dittrich. A critical branching process in random environment. Theory of Probability and its Applications, 35(3):560– 563, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 612–615.

Delyon:1999:SPS

[DJ99] B. Delyon and A. Juditsky. On small perturbations of stable Markov operators: Unbounded case. Theory of Probability and its Applications, 43(4):577– 587, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97717.

Dharmadhikari:1986:GTI

[DJD86] S. W. Dharmadhikari and K. Joag-Dev. The Gausstchebyshev inequality for unimodal distributions. Theory of Probability and its Applications, 30(4):867–871, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1984:LAF

[DK84a] R. L. Dobrushin and M. Ya. Kel’bert. Local additive functionals of Gaussian random fields. Theory of Probability and its Applications, 28(1):35–47, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1984:SLA

[DK84b] R. L. Dobrushin and M. Ya.

Kel’bert. Stationary local additive functionals of Gaussian random fields. Theory of Probability and its Applications, 28 (3):515–529, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Doksum:1993:MVS

[DK93] K. A. Doksum and A. Khoĭland.

Models for variable-stress accelerated life testing experiments based on Wiener processes and the inverse Gaussian distribution. Theory of Probability and its Applications, 37(1):137–139, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 170–172.

Dudnikova:1997:EPH

[DK97] T. V. Dudnikova and A. I.

Komech. Ergodic properties of hyperbolic equations with mixing. Theory of Probability and its Applications, 41(3):436–448, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97521. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 505–519.

Dynkin:1967:IVG

[DKPF67] E. B. Dynkin, A. N. Kolgomorov, P. T. Polyak, and M. I. Freidlin. Igor’ vladimirovich Girsanov. Theory of Probability and its Applications, 12(3):482–486, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

DiMasi:1995:MVH

[DKR95] G. B. Di Masi, Yu. M. Kabanov, and W. J. Runggaldier. Mean-variance hedging of options on Stocks with Markov volatilities. Theory of Probability and its Applications, 39 (1):172–182, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 211–222.

Dukarski:1986:SOA

[DL86] O. M. Dukarski and B. Ya. Levit. Second-order asymptotic minimax property in linear regression. Theory of Probability and its Applications, 30(4):840–845, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Doerushin:1967:ECP

[DM67] R. L. Doerushin and R. A. Minlos. Existence and continuity of pressure in classical statistical physics. Theory of Probability and its Applications, 12 (4):535–559, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Daley:1968:TSI

[DM68] D. J. Daley and P. A. P. Moran.

Two-sided inequalities for waiting time and queue size distributions in GI/G/1. Theory of Probability and its Applications, 13(2):338–341, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzaparidze:1975:ELR

[DM75] K. O. Dzaparidze and G. I. Marr.

On evaluation of the likelihood ratio for a generalized Gaussian process with rational spectral density. Theory of Probability and its Applications, 19 (2):407–410, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1979:MPG

[DM79] R. L. Dobrushin and R. A. Minlos. The moments and polynomials of a generalized random field. Theory of Probability and its Applications, 23(4):686–699, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dubrovin:1980:CLT

[DM80a] V. T. Dubrovin and D. A.

Moskvin. The Central Limit Theorem for sums of functions of mixing sequences. Theory of Probability and its Applications, 24(3):560–571, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dubrovin:1980:LDP

[DM80b] V. T. Dubrovin and D. A. Moskvin. Large deviation probabilities for sums of functions of mixing sequences. Theory of Probability and its Applications, 24(4):838–845, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1996:LLT

[DM96] Yu. A. Davydov and R. R. Manukyan. A local limit theorem for multiple Wiener-Itô stochastic integrals. Theory of Probability and its Applications, 40 (2):354–361, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzaparidze:1975:MSS

[DN75] K. O. Dzaparidze and M. S. Nikulin. On a modification of the standard statistics of Pearson. Theory of Probability and its Applications, 19(4):851–853, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dragalin:1988:ASK

[DN88] V. P. Dragalin and A. A. Novikov. Asymptotic solution of the Kiefer–Weiss problem for processes with independent increments. Theory of Probability and its Applications, 32 (4):617–627, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 32(4), (1987), pp. 679–690.

Dobrushin:1956:CLTa

[Dob56a] R. L. Dobrushin. Central Limit Theorem for nonstationary Markov chains. I. Theory of Probability and its Applications, 1(1):65–80, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1956:CLTb

[Dob56b] R. L. Dobrushin. Central Limit Theorem for nonstationary Markov chains. II. Theory of Probability and its Applications, 1(4):329–383, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1956:ECH

[Dob56c] R. L. Dobrushin. An example of a countable homogeneous Markov process all states of which are transient. Theory of Probability and its Applications, 1(4): 436–440, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1957:SCH

[Dob57] R. L. Dobrushin. Some classes of homogeneous denumerable Markov processes. Theory of Probability and its Applications, 2(3):366–369, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1958:CCS

[Dob58a] R. L. Dobrushin. The continuity condition for the sample functions of a martingale. Theory of Probability and its Applications, 3(1):92–93, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1958:ITC

[Dob58b] R. L. Dobrushin. Information transmission in a channel with feedback. Theory of Probability and its Applications, 3(4): 367–383, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1958:SME

[Dob58c] R. L. Dobrushin. A simplified method of experimentally evaluating the entropy of a stationary sequence. Theory of Probability and its Applications, 3 (4):428–430, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrusin:1958:SPA

[Dob58d] R. L. Dobrusin. A statistical problem arising in the theory of detection of signals in the presence of noise in a multi-channel system and leading to stable distribution laws. Theory of Probability and its Applications, 3 (2):161–173, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1960:PLU

[Dob60a] R. L. Dobrushin. Passage to the limit under the information and entropy signs. Theory of Probability and its Applications, 5(1):25–32, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1960:PSF

[Dob60b] R. L. Dobrushin. Properties of sample functions of a stationary Gaussian process. Theory of Probability and its Applications, 5(1):120–122, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1962:AEP

[Dob62a] R. L. Dobrushin. Asymptotic estimates of the probability of error for transmission of Messages over a discrete memoryless communication channel with a symmetric transition probability matrix. Theory of Probability and its Applications, 7 (3):270–300, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1962:OBC

[Dob62b] R. L. Dobrushin. Optimal binary codes for small rates of transmission of information. Theory of Probability and its Applications, 7(2):199–204, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1963:AOG

[Dob63] R. L. Dobrushin. Asymptotic optimality of group and systematic codes for some channels. Theory of Probability and its Applications, 8(1):47–60, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1964:ICA

[Dob64] R. L. Dobrushin. Investigation of conditions for the asymptotic existence of the configuration integral of Gibbs’ distribution. Theory of Probability and its Applications, 9(4):566–581, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1965:EPT

[Dob65]  R. L. Dobrushin.  Exis-

tence of a phase transition in two-dimensional and threedimensional Ising models. Theory of Probability and its Applications, 10(2):193–213, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1966:AEP

[Dob66] R. L. Dobrushin. Addendum: Existence of a phase transition in the two-dimensional and threedimensional Ising models. Theory of Probability and its Applications, 11(3):486–487, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobruschin:1968:DRF

[Dob68] P. L. Dobruschin. The description of a random field by means of conditional probabilities and conditions of its regularity. Theory of Probability and its Applications, 13(2):197–224, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1970:PSR

[Dob70] R. L. Dobrushin. Prescribing a system of random variables by conditional distributions. Theory of Probability and its Applications, 15(3):458–486, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1973:GSD

[Dob73] R. L. Dobrushin. Gibbs state describing coexistence of phases for a three-dimensional Ising model. Theory of Probability and its Applications, 17(4):582–600, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dobrushin:1974:IGS

[Dob74] L. R. Dobrushin. Investigation of Gibbsian states for threedimensional lattice systems. Theory of Probability and its Applications, 18(2):253–271, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dohler:1980:SPD

[Döh80] R. Döhler. Some properties of discrete randomized times. Theory of Probability and its Applications, 24(2):378–387, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dohler:1981:CIR

[Döh81]  R. Döhler.  On the condi-

tional independence of random events. Theory of Probability and its Applications, 25(3):628–634, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Doktorov:1969:SEM

[Dok69] B. Z. Doktorov. Some estimators for the multidimensional normal distribution. Theory of Probability and its Applications, 14 (3):526–528, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dokuchaev:1987:FET

[Dok87] N. G. Dokuchaev. On first exit times for homogeneous diffusion processes. Theory of Probability and its Applications, 31 (3):497–498, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dokuchaev:1992:BVP

[Dok92] N. G. Dokuchaev. Boundary value problems for functionals of ItÔ processes. Theory of Probability and its Applications, 36(3): 459–476, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 464–481.

Doksum:1993:MDS

[Dok93] K. A. Doksum. Measures of dependence in semiparametric heteroscedastic regression models. Theory of Probability and its Applications, 37(2):326–328, June 1993. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 367–369.

Dokuchaev:1995:DIP

[Dok95] N. G. Dokuchaev. Distributions of Itô processes: Estimates for the density and for conditional expectations of integral functionals. Theory of Probability and its Applications, 39(4):662–670, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 825–833.

Dokuchaev:1999:LST

[Dok99] N. G. Dokuchaev. Local sojourn time of diffusion and degenerating processes on a mobile surface. Theory of Probability and its Applications, 43 (2):171–188, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97681.

Doob:1957:BMG

[Doo57] J. L. Doob. Brownian motion on a Green space. Theory of Probability and its Applications, 2(1):1–30, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dorogovtsev:1987:AGR

[Dor87] A. A. Dorogovtsev. On application of a Gaussian random operator to random elements. Theory of Probability and its Applications, 31(4):717–720, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dorogovtsev:1990:FTW

[Dor90a] A. A. Dorogovtsev. The Fourier transform of Wiener functionals and various methods of stochastic integration. Theory of Probability and its Applications, 34 (4):705–709, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 769–773.

Dorogovtsev:1990:CRM

[Dor90b] A. A. Dorogovtsev. On the composition of random mappings in Hilbert space. Theory of Probability and its Applications, 34 (2):322–327, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 364–370.

Dorogovtsev:1996:TSD

[Dor96a] A. A. Dorogovtsev. On a type of stochastic differential equations with anticipation. Theory of Probability and its Applications, 40(4):753–756, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 903–907.

Dorogovtsev:1996:SIE

[Dor96b] A. A. Dorogovtsev. On solutions of integral equations with an extended stochastic integral. Theory of Probability and its Applications, 40(3):456–468, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 508–522.

Daletskii:1975:FTG

[DP75] Yu. L. Daletskii and S. N. Paramonova. On a formula from the theory of Gaussian measures and on the estimation of stochastic integrals. Theory of Probability and its Applications, 19 (4):812–817, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Deshayes:1982:TDR

[DP82] J. Deshayes and D. Picard. Tests of disorder of regression: Asymptotic comparison. Theory of Probability and its Applications, 27(1):100–115, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1978:CCF

[DR78] Yu. A. Davydov and A. L. Rozin.

On a characterization of certain families of measures. Theory of Probability and its Applications, 23(1):130–132, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:1979:STF

[DR79] Yu. A. Davydov and A. L. Rozin. On sojourn times for functions and random processes. Theory of Probability and its Applications, 23(3):633–637, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzhneid:1989:PEC

[DRS89] M. O. Dzhneĭd, N. A. Ruzhevich, and V. M. Shurenkov. Potentials of ergodic chains, semi-Markov processes, and sums of nonnegative variables. Theory of Probability and its Applications, 33(3):509–522, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 545–559.

Daletskii:1985:WSC

[DS85] Yu. L. Daletskii and O. G. Smolyanov. On the weak sequential completeness of the spaces of Radon measures. Theory of Probability and its Applications, 29(1):142–147, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dubins:1994:OSR

[DSS94] L. E. Dubins, L. A. Shepp, and A. N. Shiryaev. Optimal stopping rules and maximal inequalities for Bessel processes. Theory of Probability and its Applications, 38(2):226–261, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 288–330.

Dmitriev:1974:EFP

[DT74] Yu. G. Dmitriev and F. P.

Tarasenko. On the estimation of functionals of the probability density and its derivatives. Theory of Probability and its Applications, 18(3):628–633, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dmitriev:1975:CNP

[DT75] Yu. G. Dmitriev and F. P.

Tarasenko. On a class of nonparametric estimates of nonlinear functionals of density. Theory of Probability and its Applications, 19(2):390–394, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dubner:1969:DER

[Dub69] V. M. Dubner. The distribution of the eigenvalues of random cyclic matrices. Theory of Probability and its Applications, 14(2):342–344, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dubrovskii:1974:SRP

[Dub74] V. N. Dubrovskii. On small random perturbations of a secondorder differential equation. Theory of Probability and its Applications, 18(3):476–485, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dubko:1991:DAS

[Dub91] V. A. Dubko. On diffusion approximation of the slow component of a solution of a system of stochastic differential equations. Theory of Probability and its Applications, 35(3):563– 570, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 560–566.

Dai:1996:EUS

[DV96] Dzh. G. Daĭ and R. Dzh. Vil′yams. Existence and uniqueness of semimartingale reflecting Brownian motions in convex polyhedrons. Theory of Probability and its Applications, 40(1):1–40, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 3–53.

Deo:1979:QVG

[DW79] C. M. Deo and S. F. Wong. On quadratic variation of Gaussian random fields. Theory of Probability and its Applications, 23 (3):628–633, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1956:SMP

[DY56] E. B. Dynkin and A. A. Yushkevich. Strong Markov processes. Theory of Probability and its Applications, 1(1):134–139, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1968:SPI

[DY68] E. B. Dynkin and A. A. Yushkevich. On the starting points of incursions of Markov processes. Theory of Probability and its Applications, 13(3):468–470, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dyczka:1973:TFM

[Dyc73] W. Dyczka. A theorem on factorial moments in the multidimensional case. Theory of Probability and its Applications, 17 (4):703–706, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1956:IOM

[Dyn56a] E. B. Dynkin. Infinitesimal operators of Markov processes. Theory of Probability and its Applications, 1(1):34–54, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1956:MPS

[Dyn56b] E. B. Dynkin. Markov processes and semigroups of operators. Theory of Probability and its Applications, 1(1):22–33, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1958:DMP

[Dyn58] E. B. Dynkin. Discontinuous Markov processes. Theory of Probability and its Applications, 3(1):38–57, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1959:ODC

[Dyn59] E. B. Dynkin. One-dimensional continuous strong Markov processes. Theory of Probability and its Applications, 4(1):1–52, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1960:AFW

[Dyn60] E. B. Dynkin. Additive functionals of a Wiener process determined by stochastic integrals. Theory of Probability and its Applications, 5(4):402–410, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1965:CRS

[Dyn65] E. B. Dynkin. Controlled random sequences. Theory of Probability and its Applications, 10(1):1–14, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1967:GBC

[Dyn67] E. B. Dynkin. General boundary conditions for denumberable Markov processes. Theory of Probability and its Applications, 12(2):187–221, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1968:ASS

[Dyn68a] E. B. Dynkin. Addendum: Sufficient statistics for the optimal stopping problem. Theory of Probability and its Applications, 13(4):711, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1968:EMP

[Dyn68b] E. B. Dynkin. On extensions of a Markov process. Theory of Probability and its Applications, 13(4):672–676, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1968:SSO

[Dyn68c] E. B. Dynkin. Sufficient statistics for the optimal stopping problem. Theory of Probability and its Applications, 13(1):152–153, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1970:EFE

[Dyn70] E. B. Dynkin. Excessive functions and the exit space of a Markov process. Theory of Probability and its Applications, 15 (1):37–54, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1971:WMP

[Dyn71] E. B. Dynkin. Wanderings of a Markov process. Theory of Probability and its Applications, 16 (3):401–428, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:1998:SCP

[Dyn98] E. B. Dynkin. Short communication: a probabilistic approach to a nonlinear differential equation on a Riemannian manifold. Theory of Probability and its Applications, 42(2):289–294, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97614. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(2), (1997), pp. 336–341.

Dyubin:1968:GUS

[Dyu68] G. N. Dyubin. On games in the unit square with a “Roof” payoff. Theory of Probability and its Applications, 13(1):140–151, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzaparidze:1970:ESP

[Dza70] K. O. Dzaparidze. On the estimation of the spectral parameters of a Gaussian stationary process with rational spectral density. Theory of Probability and its Applications, 15(3):531–538, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzaparidze:1974:NME

[Dza74] K. O. Dzaparidze. A new method for estimating spectral parameters of a stationary regular time series. Theory of Probability and its Applications, 19(1):122–132, ???? 1974. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Dzhaparidze:1971:MOA

[Dzh71] K. O. Dzhaparidze. On methods for obtaining asymptotically efficient spectral parameter estimates for a stationary Gaussian process with rational spectral density. Theory of Probability and its Applications, 16 (3):550–554, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzhaparidze:1975:SEU

[Dzh75] K. O. Dzhaparidze. On simplified estimators of unknown parameters with good asymptotic properties. Theory of Probability and its Applications, 19(2):347–358, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzhaparidze:1977:TCH

[Dzh77] K. O. Dzhaparidze. Tests of composite hypotheses for random variables and stochastic processes. Theory of Probability and its Applications, 22(1):104–118, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzhvarsheishvili:1989:ADT

[Dzh89]  I. A. Dzhvarsheishvili.  On

Amarts with discrete time. Theory of Probability and its Applications, 33(2):260–268, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzhwarsheishvili:1994:MFH

[Dzh94] I. A. Dzhwarsheishvili. On maximal functions for harmonic functions with respect to Brownian motion. Theory of Probability and its Applications, 38 (1):14–26, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 33–48.

Dzhvarsheishvili:1995:ACT

[Dzh95]  I. A. Dzhvarsheĭshvili.  On

Amarts with continuous time. Theory of Probability and its Applications, 39(3):512–519, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 627–635.

Eagleson:1969:CEB

[Eag69] G. K. Eagleson. Canonical expansions of birth and death processes. Theory of Probability and its Applications, 14(2):209–218, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Eagleson:1977:SSC

[Eag77] G. K. Eagleson. Some simple conditions for limit theorems to be mixing. Theory of Probability and its Applications, 21 (3):637–643, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Eberl:1986:CGD

[Ebe86] W. Eberl. A characterization of the gamma distribution by optimal estimates. Theory of Probability and its Applications, 30 (4):855–860, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ebralidze:1971:IPLb

[Ebr71a] Sh. S. Ebralidze. Inequalities for probabilities of large deviations in terms of pseudomoments. Theory of Probability and its Applications, 16(4): 737–741, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 16, (1971), pp. 760–765.

Ebralidze:1971:IPLa

[Ebr71b] Sh. S. Ebralidze. Inequalities for the probabilities of large deviations in the multidimensional case. Theory of Probability and its Applications, 16(4):733–737, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Efroimovich:1980:SEU

[Efr80] S. Yu. Efroimovich. On sequential estimation under the conditions of local asymptotic normality. Theory of Probability and its Applications, 25 (1):27–40, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Efroimovich:1986:NED

[Efr86] S. Yu. Efroimovich. Nonparametric estimation of a density of unknown smoothness. Theory of Probability and its Applications, 30(3):557–568, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Efroimovich:1990:SNE

[Efr90] S. Yu. Efroĭmovich. Sequential nonparametric estimation of a density. Theory of Probability and its Applications, 34 (2):228–239, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 263–276.

Egorov:1969:LIL

[Ego69] V. A. Egorov. On the law of the iterated logarithm. Theory of Probability and its Applications, 14(4):693–699, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1970:SLL

[Ego70] V. A. Egorov. On the strong law of large numbers and the law of the iterated logarithm for sequences of independent random variables. Theory of Probability and its Applications, 15 (3):509–514, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1972:STS

[Ego72] V. A. Egorov. Some theorems on the strong law of large numbers and law of the iterated logarithm. Theory of Probability and its Applications, 17 (1):86–100, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1973:RCN

[Ego73] V. A. Egorov. On the rate of convergence to normal law which is equivalent to the existence of a second moment. Theory of Probability and its Applications, 18 (1):175–180, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1980:RCS

[Ego80] V. A. Egorov. On the rate of convergence to a stable law. Theory of Probability and its Applications, 25(1):180–187, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1985:WPT

[Ego85] V. A. Egorov. A way of proving theorems on the law of the iterated logarithm. Theory of Probability and its Applications, 29 (1):126–132, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1991:FLI

[Ego91a] V. A. Egorov. A functional law of the iterated logarithm for ordered sums. Theory of Probability and its Applications, 35 (2):342–347, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 343–349.

Egorov:1991:IEO

[Ego91b] V. A. Egorov. On the influence of extreme order statistics on the rate of growth of sums. Theory of Probability and its Applications, 35(3):570–574, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 566–570.

Egorov:1991:SLL

[Ego91c] V. A. Egorov. On the strong law of large numbers and the law of the iterated logarithm for martingales and sums of independent random variables. Theory of Probability and its Applications, 35(4):653–666, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 691–703.

Egorov:1997:ABS

[Ego97] V. A. Egorov. On the asymptotic behavior of self-normalized sums of random variables. Theory of Probability and its Applications, 41(3):542–548, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 643–650.

Elton:1993:FPD

[EH93] J. Elton and T. P. Hill. Fusions of a probability distribution: Preliminary announcement. Theory of Probability and its Applications, 37(2):328–331, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 402–406.

Eidlin:1973:SCT

[Eid73] V. L. Eidlin. On some classes of transformations which preserve normality. Theory of Probability and its Applications, 17 (3):463–471, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Evstigneev:1972:DER

[EK72] I. V. Evstigneev and S. E.

Kuznetsov. On differential equations with random coefficients. Theory of Probability and its Applications, 17(1):183–188, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Evstigneev:1982:EPS

[EK82] I. V. Evstigneev and P. K. Katyshev. Equilibrium paths in stochastic models of economic dynamics. Theory of Probability and its Applications, 27(1): 127–135, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Embrechts:1994:SAI

[EK94] P. Embrechts and C. Klüppelberg. Some aspects of insurance mathematics. Theory of Probability and its Applications, 38(2):262–295, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 374–416.

Eichelsbacher:1999:LDP

[EL99] P. Eichelsbacher and M. Lowe. Large deviations principle for partial sums U-processes. Theory of Probability and its Applications, 43(1):26–41, March 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97664.

Elbakidze:1976:CCO

[Elb76] N. V. Elbakidze. Construction of the cost and optimal policies in a game problem of stopping a Markov process. Theory of Probability and its Applications, 21 (1):163–168, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Elbakidze:1983:OSR

[Elb83] N. V. Elbakidze. Optimal stopping of a random sequence and Hammerstein’s operators. Theory of Probability and its Applications, 27(2):338–356, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Elizarov:1985:VNS

[Eli85] A. I. Elizarov. On the variance of the number of stationary points of a homogeneous Gaussian field. Theory of Probability and its Applications, 29 (3):569–570, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Elizarov:1989:CLT

[Eli89] A. I. Elizarov. Central Limit Theorem for the sojourn time and local time of a stationary process. Theory of Probability and its Applications, 33 (1):161–164, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 173–176.

Egorov:1968:SRV

[EM68] B. A. Egorov and Yu. I. Maksimov. On a sequence of random variables, with values in a compact commutative group. Theory of Probability and its Applications, 13(4):584–593, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egorov:1975:RCL

[EN75] V. A. Egorov and V. B. Nevzorov. On the rate of convergence of linear combinations of absolute order statistics to the normal law. Theory of Probability and its Applications, 20(1):203–211, ???? 1975. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 207– 215.

Egorov:1983:STI

[EN83] V. A. Egorov and V. B. Nevzorov. Some theorems on induced order statistics. Theory of Probability and its Applications, 27 (3):633–639, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Engelbert:1971:TCM

[Eng71] H. J. Engelbert. On the theory of controlled Markov processes. Theory of Probability and its Applications, 16(4):683–689, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 16, (1971), pp. 696– 702.

Engelbert:1974:TOS

[Eng74] H. J. Engelbert. On the theory of optimal stopping rules for Markov processes. Theory of Probability and its Applications, 18(2):304–311, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Engelbert:1975:OSR

[Eng75] H. J. Engelbert. On optimal stopping rules for Markov processes with continuous time. Theory of Probability and its Applications, 19(2):278–296, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Englund:1984:RTE

[Eng84] G. Englund. A remainder term estimate in a random-sum Central Limit Theorem. Theory of Probability and its Applications, 28(1):149–157, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Englund:1985:ART

[Eng85] G. Englund. Addendum: a remainder term estimate in a random-sum Central Limit Theorem. Theory of Probability and its Applications, 29(1):199, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Evstigneev:1979:STR

[EO79] I. V. Evstigneev and A. I. Ovseevich. “Splitting times” for random fields. Theory of Probability and its Applications, 23 (2):415–419, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ermakov:1986:CLT

[EO86] S. V. Ermakov and E. I. Ostrovskii. The Central Limit Theorem for weakly dependent Banach-valued variables. Theory of Probability and its Applications, 30(2):391–394, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Egishyants:1997:LGU

[EO97] S. A. Egishyants and E. I. Ostrovskii. Local and global upper functions for random fields.

Theory of Probability and its Applications, 41(4):657–665, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97568.

Epanechnikov:1968:SLP

[Epa68] V. A. Epanechnikov. The significance level and power of the twosided Kolmogorov test in the case of small sample sizes. Theory of Probability and its Applications, 13(4):686–690, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Epanechnikov:1969:NPE

[Epa69]  V. A. Epanechnikov.  Non-

parametric estimation of a multivariate probability density. Theory of Probability and its Applications, 14(1):153–158, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Epanechnikov:1974:PVS

[Epa74a] V. A. Epanechnikov. On the precise value of the significance level for the truncated one-sided Kolmogorov test. Theory of Probability and its Applications, 18 (4):787–789, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Epanecnikov:1974:POS

[Epa74b] V. A. Epanecnikov. On the power of the one-sided Kolmogorov test when the sample size is small. Theory of Probability and its Applications, 19 (1):207–212, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ephremides:1976:RPR

[Eph76] A. Ephremides. On the rank of the projection of a random process. Theory of Probability and its Applications, 20(2):419–421, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Erickson:1976:TDR

[Eri76] R. V. Erickson. Truncation of dependent random variables. Theory of Probability and its Applications, 20(4):873–880, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ermakov:1966:APA

[Erm66] S. M. Ermakov. Addendum:

Polynomial approximation and the Monte–Carlo method. Theory of Probability and its Applications, 11(4):643, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ermakov:1989:AOE

[Erm89] M. S. Ermakov. On asymptotically optimum equivariant sequential estimation of a location parameter of a non-smooth density. Theory of Probability and its Applications, 33(2):304–315, ???? 1989. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 323–335.

Ermakov:1990:AES

[Erm90] M. S. Ermakov. On asymptotic efficiency of sequential estimation of the location parameter of a nonsmooth density. Theory of Probability and its Applications, 34(3):430–440, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 479–490.

Ermakov:1991:MDS

[Erm91] M. S. Ermakov. Minimax detection of a signal in Gaussian white noise. Theory of Probability and its Applications, 35(4):667–679, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 704–715.

Ermakov:1995:MNT

[Erm95] M. S. Ermakov. Minimax nonparametric testing of hypotheses on the distribution density. Theory of Probability and its Applications, 39(3):396–416, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 488–512.

Ermakov:1996:AMT

[Erm96] M. S. Ermakov. Asymptotic minimaxity of tests of the Kolmogorov and Omega-square types. Theory of Probability and its Applications, 40(1):41–51, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 54–67.

Ermakov:1998:AMC

[Erm98] M. S. Ermakov. Asymptotic minimaxity of chi-square tests. Theory of Probability and its Applications, 42(4):589–610, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97644.

Erokhin:1958:EDR

[Ero58] V. Erokhin. ε-entropy of a discrete random variable. Theory of Probability and its Applications, 3(1):97–100, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ershov:1970:SED

[Ers70] M. P. Ershov. Sequential estimation of diffusion processes. Theory of Probability and its Applications, 15(4):688–700, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ershov:1972:ACM

[Ers72a] M. P. Ershov. On the absolute continuity of measures corresponding to diffusion type processes. Theory of Probability and its Applications, 17(1):169–174, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ershov:1972:RIP

[Ers72b] M. P. Ershov. Representations of Itô processes. Theory of Probability and its Applications, 17 (1):165–169, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ersov:1974:EMG

[Ers74] M. P. Ersov. The existence of a martingale with given diffusion functional. Theory of Probability and its Applications, 19(??):1, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 19, (1974), pp. 665– 688.

Ershov:1975:EMG

[Ers75a] M. P. Ershov. The existence of a martingale with given diffusion functional. Theory of Probability and its Applications, 19 (4):633–655, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ershov:1975:EMS

[Ers75b] M. P. Ershov. Extension of measures and stochastic equations. Theory of Probability and its Applications, 19(3):431–444, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ershov:1979:WCC

[Ers79] M. P. Ershov. On weak compactness of continuous semimartingales. Theory of Probability and its Applications, 24 (1):91–106, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ezhov:1969:MPHa

[ES69a] I. I. Ezhov and A. V. Skorokhod.

Markov processes with homogeneous second component. I. Theory of Probability and its Applications, 14(1):1–13, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ezhov:1969:MPHb

[ES69b] I. I. Ezhov and A. V. Skorokhod. Markov processes with homogeneous second component, II. Theory of Probability and its Applications, 14(4):652–667, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ezhov:1975:CBP

[ES75] I. I. Ezhov and A. A. Shakhbazov.

On a class of branching processes. Theory of Probability and its Applications, 20(1):180–185, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ezhov:1977:ETC

[ES77] I. I. Ezhov and V. M. Shurenkov. Ergodic theorems connected with the Markov property of random processes. Theory of Probability and its Applications, 21(3): 620–624, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Evstigneev:1978:MTR

[Evs78] I. V. Evstigneev. “Markov times” for random fields. Theory of Probability and its Applications, 22(3):563–569, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Evstigneev:1987:RCE

[Evs87] I. V. Evstigneev. Regular conditional expectations of random variables depending on parameters. Theory of Probability and its Applications, 31(3):515–518, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Evstigneev:1989:CRF

[Evs89] I. V. Evstigneev. Controlled random fields on an oriented graph. Theory of Probability and its Applications, 33(3):433–445, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 465–479.

Engelbert:1999:SML

[EW99] H. J. Engelbert and J. Wolf. Strong Markov local Dirichlet processes and stochastic differential equations. Theory of Probability and its Applications, 43 (2):189–202, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97682.

Ermakov:1960:PAM

[EZ60] S. M. Ermakov and V. G. Zolotukhin. Polynomial approximations and the Monte–Carlo method. Theory of Probability and its Applications, 5(4): 428–431, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ermakov:1984:RSG

[EZ84] S. M. Ermakov and A. A.

Zhiglyavskii. On random search for a global extremum. Theory of Probability and its Applications, 28(1):136–141, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fahrmann:1978:OSW

[Fäh78] H. Fährmann. On optimal stopping of a Wiener process with incomplete data. Theory of Probability and its Applications, 23 (1):138–143, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fahrmann:1983:CVO

[Fäh83]  H. Fährmann.  Convergence

of values in optimal stopping of partially observable random sequences with quadratic rewards. Theory of Probability and its Applications, 27(2):386–391, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fainberg:1976:CFS

[Fai76] E. A. Fainberg. On controlled finite state Markov processes with compact control sets. Theory of Probability and its Applications, 20(4):856–862, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fainberg:1979:ESO

[Fai79] E. A. Fainberg. The existence of a stationary ε-optimal policy for a finite Markov chain. Theory of Probability and its Applications, 23(2):297–313, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fainberg:1980:OCF

[Fai80] E. A. Fainberg. An ε-optimal control of a finite Markov chain with an average reward criterion. Theory of Probability and its Applications, 25(1):70– 81, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fainberg:1982:NRM

[Fai82] E. A. Fainberg. Non-randomized Markov and semi-Markov strategies in dynamic programming. Theory of Probability and its Applications, 27(1):116–126, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fainberg:1983:CMP

[Fai83]  E. A. Fainberg.  Controlled

Markov processes with arbitrary numerical criteria. Theory of Probability and its Applications, 27(3):486–503, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fainberg:1987:SCS

[Fai87] E. A. Fainberg. Sufficient classes of strategies in discrete dynamic programming I: Decomposition of randomized strategies and embedded models. Theory of Probability and its Applications, 31 (4):658–668, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fajnberg:1988:SCS

[Faj88] E. A. Fajnberg. Sufficient classes of strategies in discrete dynamic programming. II: Locally stationary strategies. Theory of Probability and its Applications, 32 (3):435–448, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fakeev:1969:NCO

[Fak69] A. G. Fakeev. On necessary conditions for optimality in sequential decision problems. Theory of Probability and its Applications, 14(4):710–713, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fakeev:1970:OSR

[Fak70] A. G. Fakeev. Optimal stopping rules for stochastic processes with continuous parameter. Theory of Probability and its Applications, 15(2):324–331, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fakeev:1971:OSM

[Fak71] A. G. Fakeev. Optimal stopping of a Markov process. Theory of Probability and its Applications, 16(4):694–696, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Faktorovich:1982:SRC

[Fak82] I. Yu. Faktorovich. Some results connected with a waiting process. Theory of Probability and its Applications, 27(1):183–190, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fal:1978:GRO

[Fal78]  A. M. Fal’.  Generalization

of the results obtained by R. L. Dobrushin for additive functionals of a Markov random walk. Theory of Probability and its Applications, 22(3):569–571, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Falin:1988:CQP

[Fal88] G. I. Falin. On the convergence of queueing processes to stationary processes in structurally complex lossy systems. Theory of Probability and its Applications, 32(3):525–527, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Falin:1989:CMP

[Fal89] G. I. Falin. Comparability of migration processes. Theory of Probability and its Applications, 33(2):370–372, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Falin:1992:DAR

[Fal92] G. I. Falin. A diffusion approximation for retrial queueing systems. Theory of Probability and its Applications, 36(1):149–152, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 183–186.

Fatalov:1997:LDG

[Fat97] V. R. Fatalov. Large deviations of Gaussian measures in the spaces l^{p} and L^{p}, p ≥ 2. Theory of Probability and its Applications, 41(3):548–555, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 682–689.

Fedotkin:1969:ASS

[Fed69] M. A. Fedotkin. An analysis of the statistical stability of traffic flow at an intersection controlled by a feedback automaton. Theory of Probability and its Applications, 14(3):470–480, ???? 1969. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Fedotkin:1970:AAS

[Fed70] M. A. Fedotkin. Addendum: an analysis of the statistical stability of traffic flow at an intersection controlled by a feedback automaton. Theory of Probability and its Applications, 15(3):551, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fedorov:1971:DEL

[Fed71a] V. V. Fedorov. Design of experiments for linear optimality criteria. Theory of Probability and its Applications, 16(1):189–195, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fedorov:1971:DEM

[Fed71b] V. V. Fedorov. The design of experiments in the multiresponse case. Theory of Probability and its Applications, 16(2):323–332, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fedotkin:1977:EED

[Fed77] M. A. Fedotkin. On the existence of an ergodic distribution in a system with variable structure for servicing conflicting flows. Theory of Probability and its Applications, 21(4):770–779, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fedotov:1982:IIO

[Fed82] A. M. Fedotov. An information inequality for operator equations in Hilbert space. Theory of Probability and its Applications, 26 (2):371–378, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fedorenko:1985:ESS

[Fed85] I. V. Fedorenko. Existence of a strong solution of the Itô stochastic differential equation. Theory of Probability and its Applications, 29(1):121–123, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1965:ECQ

[Fel65] J. Feldman. Extending complex quasi-measures. Theory of Probability and its Applications, 10 (2):345–348, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1976:GPD

[Fel76] G. M. Fel’dman. On generalized Poisson distributions on groups. Theory of Probability and its Applications, 20(3):641–644, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1977:DGD

[Fel77] G. M. Fel’dman. On the decomposition of Gaussian distributions on groups. Theory of Probability and its Applications, 22(1):133–140, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1979:GDL

[Fel79] G. M. Fel’dman. Gaussian distributions on locally compact Abelian groups. Theory of Probability and its Applications, 23 (3):529–542, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1983:DGP

[Fel83] G. M. Fel’dman. On a decomposition of generalized Poisson distributions on groups. Theory of Probability and its Applications, 27(4):780–794, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1985:GPD

[Fel85] G. M. Fel’dman. Generalized Poisson distribution of class I_{0} on groups. Theory of Probability and its Applications, 29(2):218–230, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1986:IDD

[Fel86] G. M. Fel’dman. On infinitely divisible distributions of the class I_{0} on groups. Theory of Probability and its Applications, 30 (3):475–488, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1987:BGD

[Fel87] G. M. Feld’man. Bernstein Gaussian distributions on groups.

Theory of Probability and its Applications, 31(1):40–49, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1988:CGD

[Fel88] G. M. Fel’dman. On a characterization of the Gaussian distribution on Abelian groups. Theory of Probability and its Applications, 32(3):565–569, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1990:MLT

[Fel90] G. M. Fel′dman. Marcinkiewicz and Lukacs theorems on Abelian groups. Theory of Probability and its Applications, 34(2):290–297, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 330–339.

Feldman:1992:CDA

[Fel92] G. M. Fel’dman. Cauchy distribution on Abelian groups and its characterization. Theory of Probability and its Applications, 36(4):670–681, December 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1993:SDT

[Fel93]  G. M. Fel′dman.  On the

Skitovich—Darmois Theorem on Abelian groups. Theory of Probability and its Applications, 37(4):

621–631, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 695–708.

Feldman:1998:SDT

[Fel98] G. M. Fel’dman. The Skitovich– Darmois theorem for discrete periodic Abelian groups. Theory of Probability and its Applications, 42(4):611–617, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97648.

Feldman:1999:CGD

[Fel99] G. M. Fel’dman. On a characterization of Gaussian distributions on Abelian groups by constancy of regression. Theory of Probability and its Applications, 43(3): 477–480, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97709.

Ferger:1998:SCO

[Fer98] D. Ferger. Short communication: Optimal bounds for the Prokhorov distance of the Miller-Sen process and Brownian motion. Theory of Probability and its Applications, 42(1):155–162, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97604. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 202–208.

Fetisov:1974:IMC

[Fet74] V. N. Fetisov. An inequality in the Monte Carlo method. Theory of Probability and its Applications, 19(1):216–217, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:1982:DCT

[FF82] G. M. Fel’dman and A. E. Fryntov. Decomposition of the convolution of two Poisson distributions on a locally compact Abelian group. Theory of Probability and its Applications, 26 (3):601–607, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Filippov:1961:DSP

[Fil61] A. F. Filippov. On the distribution of the sizes of particles which undergo splitting. Theory of Probability and its Applications, 6(3):275–294, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Filippova:1962:MTA

[Fil62] A. A. Filippova. Mises’ theorem on the asymptotic behavior of functionals of empirical distribution functions and its statistical applications. Theory of Probability and its Applications, 7(1):24–57, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Filonov:1984:ETR

[Fil84] Yu. P. Filonov. Expected times for reaching a level for range type functionals of a Markov chain. Theory of Probability and its Applications, 28(4):825–829, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fintushal:1971:APE

[Fin71] S. M. Fintushal. Admissibility of Pitman’s estimator for a multidimensional location parameter. Theory of Probability and its Applications, 16(4):702–707, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Findley:1993:CNM

[Fin93] D. Findley. Comparing nonnested, misspecified models. Theory of Probability and its Applications, 37(2):274–283, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 342–351.

Fortini:1997:CLP

[FLR97] S. Fortini, L. Ladelli, and E. Regazzini. A central limit problem for partially exchangeable random variables. Theory of Probability and its Applications, 41(2):224–246, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/

article/97545. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 353–379.

Freimer:1992:ACB

[FM92] M. Freimer and G. S. Mudholkar.

An analogue of the Chernoff– Borokov–Utev inequality and related characterization. Theory of Probability and its Applications, 36(3):589–592, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Fuk:1971:PIS

[FN71] D. Kh. Fuk and S. V. Nagaev. Probability inequalities for sums of independent random variables. Theory of Probability and its Applications, 16 (4):643–660, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fellman:1993:REL

[FN93] J. Fellman and K. Nordström.

Robust estimation in linear models with nuisance parameters. Theory of Probability and its Applications, 37(2):332–333, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 393–395.

Fokht:1974:DFJ

[Fok74] A. I. Fokht. On the distribution of the first jump for a generalized Poisson process with drift. Theory of Probability and its Applications, 19(1):159–163, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fokht:1977:DSF

[Fok77] A. I. Fokht. Distribution of the size of the first jump over a level for a class of stochastic processes. Theory of Probability and its Applications, 21(2):419–423, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fomin:1983:ERC

[Fom83] S. V. Fomin. An estimate of the rate of convergence in the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 27(2):365–368, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fortus:1962:FER

[For62] M. I. Fortus. Formulas for extrapolation of random fields. Theory of Probability and its Applications, 7(1):101–108, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fortus:1963:ERF

[For63] M. I. Fortus. On extrapolation of a random field satisfying the wave equation. Theory of Probability and its Applications, 8(2):204–207, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Franken:1963:RLT

[Fra63] P. Franken. A refinement of the limit theorem for the superposition of independent renewal processes. Theory of Probability and its Applications, 8 (3):320–328, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

FragaAlves:1993:ICO

[Fra93] M. I. Fraga Alves. The influence of central observations on discrimination among multivariate extremal models. Theory of Probability and its Applications, 37(2):334–337, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 395–398.

Freidlin:1963:DPR

[Fre63] M. I. Freidlin. Diffusion processes with reflection and problems with a directional derivative on a manifold with a boundary. Theory of Probability and its Applications, 8(1):75–83, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1964:DPE

[Fre64] M. I. Freidlin. Dirichlet’s problem for an equation with periodic coefficients depending on a small parameter. Theory of Probability and its Applications, 9(1):121–125, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1965:NGS

[Fre65] M. I. Freidlin. A note on the generalized solution of the Dirichlet problem. Theory of Probability and its Applications, 10 (1):161–164, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1966:EDP

[Fre66] M. I. Freidlin. The exterior Dirichlet problem in the class of bounded functions. Theory of Probability and its Applications, 11(3):407–414, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1967:SWS

[Fre67] M. I. Freidlin. On small (in the weak sense) perturbations in the coefficients of a diffusion process. Theory of Probability and its Applications, 12(3):487–490, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1968:FNN

[Fre68] M. I. Freidlin. On the factorization of non-negative definite matrices. Theory of Probability and its Applications, 13(2):354–356, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1969:DEE

[Fre69] M. I. Freidlin. On degenerating elliptic equations. Theory of Probability and its Applications, 14(1):137–141, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1973:AFC

[Fre73] M. I. Freidlin. The action functional for a class of stochastic processes. Theory of Probability and its Applications, 17 (3):511–515, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1974:PLD

[Fre74] M. I. Freidlin. Probabilities of large deviations for randomly perturbed dynamic systems and stochastic stability. Theory of Probability and its Applications, 18(4):779–784, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlina:1975:DPS

[Fre75] V. L. Freidlina. On a design problem for screening experiments. Theory of Probability and its Applications, 20(1):102–115, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freidlin:1976:SHR

[Fre76] M. I. Freidlin. On the stability of highly reliable systems. Theory of Probability and its Applications, 20(3):572–583, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Freivald:1988:OTE

[Fre88] R. V. Freivald. On the operating time of errorless probabilistic Turing machines. Theory of Probability and its Applications, 32(3):514–516, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Frid:1969:OSR

[Fri69]  E. B. Frid.  The optimal

stopping rule for a two-person Markov chain with opposing interests. Theory of Probability and its Applications, 14(4):713–716, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Frid:1970:PDL

[Fri70] E. B. Frid. On a problem of D. Lackwell from the theory of dynamic programming. Theory of Probability and its Applications, 15(4):719–722, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Frid:1972:OSC

[Fri72] E. B. Frid. On optimal strategies in control problems with constraints. Theory of Probability and its Applications, 17 (1):188–192, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Frid:1974:SG

[Fri74]  E. B. Frid.  On stochastic

games. Theory of Probability and its Applications, 18(2):389–393, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Frolov:1989:BVP

[Fro89] F. F. Frolov. Boundary-value problems for systems of stochastic partial differential equations. Theory of Probability and its Applications, 33(2):385–388, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 409–412.

Fryntov:1976:FID

[Fry76] A. E. Fryntov. On the factorization of infinitely divisible distributions. Theory of Probability and its Applications, 20 (3):648–652, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fryntov:1989:CGD

[Fry89] A. E. Fryntov. Characterization of a Gaussian distribution by gaps in its sequence of cumulants. Theory of Probability and its Applications, 33(4):638–644, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fryntov:1990:ECG

[Fry90] A. E. Fryntov. Erratum: “Characterization of the Gaussian distribution by lacunae of a sequence of cumulants” [Teor. Veroyatnost. i Primenen. 33 (1988), no. 4, 687–693; MR 90a:60027]. Theory of Probability and its Applications, 34 (3):554–555, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 613–614.

Franken:1979:SGR

[FS79] P. Franken and A. Streller. Stationary generated regenerative processes. Theory of Probability and its Applications, 24 (1):79–90, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Finkelstein:1995:ITI

[FST95] M. Finkelstein, S. Scheiberg, and H. G. Tucker. Integral transforms with infinitely divisible kernels. Theory of Probability and its Applications, 39(4): 670–676, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 856–863.

Finkelstein:1992:CRS

[FTV92] M. Finkelstein, H. G. Tucker, and J. A. Veeh. Convergence of random sums with nonrandom centering. Theory of Probability and its Applications, 36(2):366–371, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 397–402.

Faddeev:1961:CCP

[FV61] D. K. Faddeev and N. N. Vorob’ev. Continuation of conditional probabilities. Theory of Probability and its Applications, 6(1):105–107, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gabovich:1975:SCC

[Gab75] Yu. R. Gabovich. On the stability of certain characteristic properties of the normal distribution. Theory of Probability and its Applications, 19(2):365–371, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gadzhiev:1990:SEL

[Gad90] A. G. Gadzhiev. Statistical estimates in linear regression models with correlated observation errors. Theory of Probability and its Applications, 34(4):694–698, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 764–768.

Galstyan:1971:AEC

[Gal71] F. N. Galstyan. On asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 16 (3):528–533, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1977:RSM

[Gal77] L. I. Gal’chuk. A representation for some martingales. Theory of Probability and its Applications, 21(3):599–605, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1978:GGT

[Gal78a] L. I. Gal’chuk. A generalization of Girsanov’s theorem on substitution of measures to the case of semi-martingales with jumps. Theory of Probability and its Applications, 22(2):271–285, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1978:EOM

[Gal78b] L. I. Gal’chuk. On the existence of optional modifications for martingales. Theory of Probability and its Applications, 22 (3):572–573, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1979:EUS

[Gal79] L. I. Gal’chuk. Existence and uniqueness of a solution for stochastic equations with respect to semimartingales. Theory of Probability and its Applications, 23(4):751–763, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1981:SSD

[Gal81] L. I. Gal’chuk. Stable subspaces and a decomposition theorem for martingales. Theory of Probability and its Applications, 25 (2):366–370, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1983:CTS

[Gal83] L. I. Gal’chuk. A comparison theorem for stochastic equations with integrals with respect to martingales and random measures. Theory of Probability and its Applications, 27(3):450–460, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1985:OSP

[Gal85a] L. I. Gal’chuk. Optional semimartingales and processes with independent increments. Theory of Probability and its Applications, 29(2):293–306, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1985:SIR

[Gal85b] L. I. Gal’chuk. Stochastic integrals with respect to optional semimartingales and random measures. Theory of Probability and its Applications, 29 (1):93–108, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galtchouk:1994:SPI

[Gal94] L. I. Galtchouk. Semimartingales of processes with independent increments and enlargement of filtration. Theory of Probability and its Applications, 38 (3):395–404, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 491–502.

Galambos:1995:DMT

[Gal95] J. Galambos. The development of the mathematical theory of extremes in the past half century. Theory of Probability and its Applications, 39(2):234–248, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 272–293.

Gamkrelidze:1964:LLT

[Gam64] N. G. Gamkrelidze. On a local limit theorem for lattice random variables. Theory of Probability and its Applications, 9 (4):662–664, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1966:SCL

[Gam66] N. G. Gamkrelidze. On the speed of convergence in the local limit theorem for lattice distributions. Theory of Probability and its Applications, 11(1):114–125, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1968:CBL

[Gam68] N. G. Gamkrelidze. On the connection between the local and the integral theorem for lattice distributions. Theory of Probability and its Applications, 13 (1):174–179, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1974:EMP

[Gam74] N. G. Gamkrelidze. Estimation of the maximum probability for sums of lattice random variables. Theory of Probability and its Applications, 18(4):799–803, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1978:EIM

[Gam78] N. G. Gamkrelidze. Esseen’s inequality for multi-dimensional distribution functions. Theory of Probability and its Applications, 22(4):877–880, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1981:MPC

[Gam81] N. G. Gamkrelidze. A method for proving the Central Limit Theorem. Theory of Probability and its Applications, 25(3):609–614, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1982:SPS

[Gam82] N. G. Gamkrelidze. On “Smoothing” the probabilities for sums of independent integer-valued variables. Theory of Probability and its Applications, 26(4):823–828, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1984:ECV

[Gam84] N. G. Gamkrelidze. On an estimate of the closeness in variation of distributions. Theory of Probability and its Applications, 28(2):467–469, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1986:MSM

[Gam86] N. G. Gamkrelidze. A measure of “smoothness” of multidimensional distributions of integervalued random vectors. Theory of Probability and its Applications, 30(2):427–431, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1989:ASF

[Gam89] N. G. Gamkrelidze. On the application of a smoothness function in proving a local limit theorem. Theory of Probability and its Applications, 33(2):352–355, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 373–376.

Gamkrelidze:1991:ACB

[Gam91] N. G. Gamkrelidze. Addendum: On the connection between the local and the integral theorem for lattice distributions. Theory of Probability and its Applications, 35(2):404, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1992:ACB

[Gam92a] N. G. Gamkrelidze. Addendum: On the connection between the local and integral theorems for lattice distributions. Theory of Probability and its Applications, 36(4):859, ???? 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 36(4), (1991), pp. 822–822.

Gamkrelidze:1992:IMC

[Gam92b] N. G. Gamkrelidze. An inequality for multidimensional characteristic function. Theory of Probability and its Applications, 36(3):594–596, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 602–604.

Gaposhkin:1965:LIL

[Gap65] V. F. Gaposhkin. The law of the iterated logarithm for Cesaro’s and Abel’s methods of summation. Theory of Probability and its Applications, 10(3):411–420, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1968:SCG

[Gap68] V. F. Gaposhkin. On the speed of convergence to the Gaussian law of weighted sums of gap series. Theory of Probability and its Applications, 13(3):421–437, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1969:LIL

[Gap69]  V. F. Gaposhkin.  On the

law of the iterated logarithm for strongly multiplicative systems. Theory of Probability and its Applications, 14(3):493–496, ???? 1969. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1970:CLT

[Gap70] V. F. Gaposhkin. The Central Limit Theorem for some weakly dependent sequences. Theory of Probability and its Applications, 15(4):649–666, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposkin:1973:CLT

[Gap73] V. F. Gaposkin. Convergence and limit theorems for sequences of random variables. Theory of Probability and its Applications, 17(3):379–400, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1974:SLL

[Gap74] V. F. Gaposhkin. On the strong law of large numbers for second order stationary processes and sequences. Theory of Probability and its Applications, 18 (2):372–375, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1978:CSL

[Gap78] V. F. Gaposhkin. Criteria for the strong law of large numbers for some classes of second-order stationary processes and homogeneous random fields. Theory of Probability and its Applications, 22(2):286–310, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1980:ASC

[Gap80] V. F. Gaposhkin. Almost sure convergence of estimates for the spectral density of a stationary process. Theory of Probability and its Applications, 25 (1):169–176, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1982:DCR

[Gap82] V. F. Gaposhkin. On the dependence of the convergence rate in the strong law of large numbers for stationary processes on the rate of decay of the correlation function. Theory of Probability and its Applications, 26 (4):706–720, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1986:RSC

[Gap86] V. F. Gaposhkin. A remark on strong consistency of LS estimates under weakly correlated observation errors. Theory of Probability and its Applications, 30(1):177–181, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(1), (1985), pp. 160–163.

Gaposhkin:1988:SAE

[Gap88] V. F. Gaposhkin. On the summation almost everywhere of stationary sequences. Theory of Probability and its Applications, 32(1):128–131, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:1989:RBE

[Gap89a] V. F. Gaposhkin. On a relationship between ergodicity of a continuous-time stationary process and a quantized process. Theory of Probability and its Applications, 33(2):377–381, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 401–405.

Gaposhkin:1989:SSI

[Gap89b] V. F. Gaposhkin. Summability of sequences of independent random variables. Theory of Probability and its Applications, 33 (1):62–74, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 68–82.

Gaposhkin:1992:MBI

[Gap92]  V. F. Gaposhkin.  Moment

bounds for integrals of ρ-mixing fields. Theory of Probability and its Applications, 36(2):249–260, May 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 262–273.

Gaposhkin:1995:SLL

[Gap95]  V. F. Gaposhkin.  On the

strong law of large numbers for blockwise-independent and blockwise-orthogonal random variables. Theory of Probability and its Applications, 39 (4):677–684, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 804–812.

Gaposhkin:1997:SCE

[Gap97] V. F. Gaposhkin. Spectral criteria for existence of generalized ergodic transforms. Theory of Probability and its Applications, 41(2):247–264, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97546. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 251–271.

Gaposhkin:1998:AEC

[Gap98] V. F. Gaposhkin. On almost everywhere convergence of the Riesz averages of homogeneous random fields. Theory of Probability and its Applications, 42 (3):405–415, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 461–472.

Gartner:1975:ABT

[Gär75] Jürgen Gärtner. On the asymptotic behavior of the time of first exit from a region. Theory of Probability and its Applications, 20(1):169–174, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gartner:1976:TLD

[Gär76] Jürgen Gärtner. Theorems on large deviations for a certain class of random processes. Theory of Probability and its Applications, 21(1):96–107, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gartner:1977:LDI

[Gär77] Jürgen Gärtner. On large deviations from the invariant measure. Theory of Probability and its Applications, 22(1):24– 39, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gasparov:1963:SIR

[Gas63] M. Gasparov. Statistical investigation of Russian dolnik trimeter. Theory of Probability and its Applications, 8(1): 96–102, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goltyaeva:1987:ECP

[GC87] O. N. Gol’tyaeva and V. P. Chistyakov. On estimating the criticality parameter of a branching process with immigration. Theory of Probability and its Applications, 31(4):693–697, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gnedenko:1988:J

[GC88] B. V. Gnedenko and D. M.

Chibisov. Jubilees. Theory of Probability and its Applications, 32(2):386, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gatkin:1971:ODS

[GD71] N. G. Gatkin and Yu. L. Daletskii. On optimal detection of a signal in arbitrary noise. Theory of Probability and its Applications, 16(4):728–732, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gushchin:1991:EAP

[GE91] A. A. Gushchin and A. I.

Ekushov. Estimation of the average payoff from an open-loop control for one type of controlled Markov sequences. Theory of Probability and its Applications, 35(3):431–442, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 438–448.

Greenwood:1993:MER

[GE93] P. Greenwood and I. Evstigneev.

A Markov evolving random field and splitting random elements. Theory of Probability and its Applications, 37(1):40–42, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 46–48.

Gelfond:1964:ERT

[Gel64] A. O. Gel’fond. An estimate for the remainder term in a limit theorem for recurrent events. Theory of Probability and its Applications, 9(2):299–303, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gelberg:1974:RBM

[Gel74] M. G. Gel’berg. The relation between Mann–Whitney’s statistic and Kendall’s correlation coefficient τ. Theory of Probability and its Applications, 19(1):205–207, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gendenko:1959:RB

[Gen59] B. V. Gendenko. Reviews and bibliographies. Theory of Probability and its Applications, 4 (3):336–338, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Geodakov:1980:LMP

[Geo80] N. A. Geodakov. A local Markov property for Gaussian stationary processes. Theory of Probability and its Applications, 24 (4):852–856, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gertsbach:1967:EPD

[Ger67] I. B. Gertsbach. On an estimation of parameters of a distribution with an exponential factor. Theory of Probability and its Applications, 12(1):110–111, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gercbah:1973:PCO

[Ger73] I. B. Gercbah. On a problem of choosing the order of dependent trials. Theory of Probability and its Applications, 17(4):709–712, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gershanov:1979:OED

[Ger79] A. M. Gershanov. The optimal estimation of a discrete parameter. Theory of Probability and its Applications, 24(1):220–224, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gerencser:1996:SSA

[Ger96] L. Gerencsér. Slow stochastic approximation processes are good for estimating slope. Theory of Probability and its Applications, 40(1):145–151, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 192–199.

Gorin:1991:GKI

[GF91] E. A. Gorin and S. Yu. Favorov.

Generalizations of Khinchin’s inequality. Theory of Probability and its Applications, 35(4):

766–771, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 762–767.

Girotto:1993:WCM

[GH93] B. Girotto and S. Holzer. Weak convergence of masses on normal topological spaces. Theory of Probability and its Applications, 37(2):337–340, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 364–366.

Giang:1996:MZL

[Gia96] Nguyen Van Giang. Marcinkiewicz– Zygmund laws for Banach space valued random variables with multidimensional parameters. Theory of Probability and its Applications, 40(1):175–181, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 213–219.

Gihman:1991:AQL

[Gih91] Il. I. Gihman. Addendum: Quasilinear evolutionary stochastic systems with averaging. Theory of Probability and its Applications, 35(1):192, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gikhman:1956:APS

[Gik56] I. I. Gikhman. On asymptotic properties of some statistics similar to χ^{2}. Theory of Probability and its Applications, 1 (3):312–315, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gikhman:1957:NHC

[Gik57] I. I. Gikhman. On a nonparametric homogeneity criterion for k samples. Theory of Probability and its Applications, 2 (3):369–373, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gikhman:1958:LTN

[Gik58] I. I. Gikhman. A limit theorem for the number of maxima in the sequence of random variables in a Markov chain. Theory of Probability and its Applications, 3(2):154–160, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gikhman:1990:QES

[Gik90] Il. I. Gikhman. Quasilinear evolutionary stochastic systems with averaging. Theory of Probability and its Applications, 34 (3):416–429, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 465–478.

Gilfanov:1982:NUE

[Gil82] R. M. Gil’fanov. A non-uniform estimate of the rate of convergence in the Central Limit Theorem for the sums of vector-valued functions of independent variables. Theory of Probability and its Applications, 26(4):773–786, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gill:1993:MSAa

[Gil93a] R. D. Gill. Multivariate survival analysis. I. Theory of Probability and its Applications, 37 (1):18–31, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 19–35.

Gill:1993:MSAb

[Gil93b] R. D. Gill. Multivariate survival analysis. II. Methods. Theory of Probability and its Applications, 37(2):284–301, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 307–328.

Gurevich:1997:RLD

[GIM^{+}97] B. M. Gurevich, I. A. Ibragimov, R. A. Minlos, I. A. Ovseevich, V. I. Oseledets, M. S. Pinsker, V. V. Prelov, Yu. V. Prokhorov, Ya. G. Sinaĭ, A. S. Kholevo, and A. N. Shiryaev. Roland L′vovich Dobrushin (July 20, 1929–December 12, 1995). Theory of Probability and its Applications, 41(1):132–136, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 164–169.

Ginzburg:1966:PBT

[Gin66] G. M. Ginzburg. On a problem in S. N. Bernshtein’s theory of stochastic differential equations. Theory of Probability and its Applications, 11(2):291–294, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ginovyan:1989:AEN

[Gin89a] M. S. Ginovyan. Asymptotically efficient nonparametric estimation of functionals of a spectral density having zeros. Theory of Probability and its Applications, 33(2):296–303, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 315–322.

Ginovyan:1989:EVL

[Gin89b] M. S. Ginovyan. On estimating the value of a linear functional of the spectral density of a Gaussian stationary process. Theory of Probability and its Applications, 33(4):722–726, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girsanov:1959:PND

[Gir59] I. V. Girsanov. On a property of non-degenerate diffusion processes. Theory of Probability and its Applications, 4(3):

329–333, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girsanov:1960:TCC

[Gir60a] I. V. Girsanov. On transforming a certain class of stochastic processes by absolutely continuous substitution of measures. Theory of Probability and its Applications, 5(3):285–301, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girsanov:1960:SFP

[Gir60b] I. V. Girsanov. Strongly-Feller processes I. General properties. Theory of Probability and its Applications, 5(1):5–24, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girsanov:1962:ENU

[Gir62] I. V. Girsanov. An example of non-uniqueness of the solution of the stochastic equation of K. Ito. Theory of Probability and its Applications, 7 (3):325–331, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1976:LTP

[Gir76] V. L. Girko. A limit theorem for products of random matrices. Theory of Probability and its Applications, 21(1):197–199, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1977:GLT

[Gir77] V. L. Girko. General limit theorems for spectral functions of random matrices. Theory of Probability and its Applications, 22(1):156–160, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1980:CLT

[Gir80] V. L. Girko. The Central Limit Theorem for random determinants. Theory of Probability and its Applications, 24(4):729–740, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1982:CLT

[Gir82] V. L. Girko. The Central Limit Theorem for random determinants. Theory of Probability and its Applications, 26(3):521–531, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1983:LTP

[Gir83] V. L. Girko. Limit theorems for products of independent random matrices with positive elements. Theory of Probability and its Applications, 27(4):837–844, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1985:CL

[Gir85] V. L. Girko. Circular law. Theory of Probability and its Applications, 29(4):694–706, ???? 1985.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1986:EL

[Gir86] V. L. Girko. Elliptic law. Theory of Probability and its Applications, 30(4):677–690, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1988:IGS

[Gir88] V. L. Girko. Introduction to general statistical analysis. Theory of Probability and its Applications, 32(2):229–242, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1991:LTM

[Gir91] V. L. Girko. Limit theorems for maximal and minimal eigenvalues of random matrices. Theory of Probability and its Applications, 35(4):680–695, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 677–690.

Girko:1992:SLA

[Gir92] V. L. Girko. Systems of linear algebraic equations with random coefficients. Theory of Probability and its Applications, 36 (2):261–271, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 36(2), (1991), pp. 274–284.

Girko:1995:CSE

[Gir95] V. L. Girko. The canonical spectral equation. Theory of Probability and its Applications, 39 (4):685–691, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 812–820.

Girko:1996:MER

[Gir96] V. L. Girko. A matrix equation for resolvents of random matrices with independent blocks. Theory of Probability and its Applications, 40(4):635–644, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 741–753.

Girko:1998:RCL

[Gir98] V. L. Girko. A refinement of the Central Limit Theorem for random determinants. Theory of Probability and its Applications, 42(1):121–129, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97593. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 63–73.

Gusak:1968:FPT

[GK68] D. V. Gusak and V. S. Koralyuk.

On the first passage time across a given level for processes with independent increments. Theory of Probability and its Applications, 13(3):448–456, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gusak:1969:JDP

[GK69] D. V. Gusak and V. S. Korolyuk. On the joint distribution of a process with stationary increments and its maximum. Theory of Probability and its Applications, 14(3):400–409, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposkin:1975:LIL

[GK75] V. F. Gaposkin and T. P. Krasulina. On the law of the iterated logarithm in stochastic approximation processes. Theory of Probability and its Applications, 19(4):844–850, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Genis:1976:EOD

[GK76] I. L. Genis and N. V. Krylov. An example of a one-dimensional controlled process. Theory of Probability and its Applications, 21(1):148–152, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golubev:1980:EGW

[GK80] G. K. Golubev and R. Z. Khas’minskij. Estimation in Gaussian white noise using a finite number of linear statistics. Theory of Probability and its Applications, 24(??):274–286, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golubev:1981:EGW

[GK81] G. K. Golubev and R. Z.

Khas’minskii. Estimation in Gaussian white noise using a finite number of linear statistics. Theory of Probability and its Applications, 25(2):274–286, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Godovanchuk:1984:CLC

[GK84a] V. V. Godovanchuk and A. P.

Korostelev. Conditions for local convergence of recursive stochastic procedures. Theory of Probability and its Applications, 28 (1):142–149, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golubev:1984:STS

[GK84b] G. K. Golubev and R. Z.

Khas’minskii. Sequential testing for several signals in Gaussian white noise. Theory of Probability and its Applications, 28 (3):573–584, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goldstein:1996:EES

[GK96a] L. Goldstein and R. Khas′minskiĭ.

On efficient estimation of smooth functionals. Theory of Probability and its Applications, 40 (1):151–156, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 199–205.

Grinevich:1996:DAS

[GK96b] I. V. Grinevich and Yu. S. Khokhlov. The domains of attraction of semistable laws. Theory of Probability and its Applications, 40(2):361–366, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 417–422.

Gnedenko:1960:WVS

[GKPS60] B. V. Gnedenko, A. N. Kolmogorov, Yu. V. Prokhorov, and O. V. Sarmanov. On the works of N. V. Smirnov in mathematical statistics. Theory of Probability and its Applications, 5 (4):397–401, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gladyshev:1958:MDS

[Gla58] E. G. Gladyshev. On multidimensional stationary random processes. Theory of Probability and its Applications, 3(4): 425–428, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gladyshev:1961:NLT

[Gla61] E. G. Gladyshev. A new limit theorem for stochastic processes with Gaussian increments. Theory of Probability and its Applications, 6(1):52–61, ???? 1961.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gladyshev:1963:PAP

[Gla63]  E. G. Gladyshev.  Periodi-

cally and almost-periodically correlated random processes with a continuous time parameter. Theory of Probability and its Applications, 8(2):173–177, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gladyshev:1965:SA

[Gla65] E. G. Gladyshev. On stochastic approximation. Theory of Probability and its Applications, 10 (2):275–278, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glonti:1970:SFC

[Glo70] O. A. Glonti. Sequential filtering of components of a Markov chain with degenerate diffusion matrix. Theory of Probability and its Applications, 15(4):715–718, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glonti:1979:ROF

[Glo79a] O. A. Glonti. On a representation of an optimal filtration estimate of Markov processes controlled by stochastic integro-differential equations in partial derivatives. Theory of Probability and its Applications, 23(4):824–829, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glonti:1979:MIS

[Glo79b] O. A. Glonti. On mutual information of signals in the case of transmission over a “noisy” feedback channel. Theory of Probability and its Applications, 23 (2):377–379, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glover:1982:EMP

[Glo82] J. Glover. Energy and the maximum principle for nonsymmetric hunt processes. Theory of Probability and its Applications, 26 (4):745–757, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gluskin:1973:NCT

[Glu73] A. M. Gluskin. A new criterion for testing nearness to BTL-Systems in the method of pairwise comparisons. Theory of Probability and its Applications, 17(2):338–342, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grigelionis:1984:SWC

[GM84] B. Grigelionis and R. Mikulyavichus. On stably weak convergence of semi-martingales and of point processes. Theory of Probability and its Applications, 28 (2):337–350, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Greenwood:1987:OSB

[GN87] P. E. Greenwood and A. A.

Novikov. One-sided boundary crossing for processes with independent increments. Theory of Probability and its Applications, 31(2):221–232, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golubev:1993:ASE

[GN93] G. K. Golubev and M. Nussbaum. Adaptive spline estimates for nonparametric regression models. Theory of Probability and its Applications, 37(3): 521–529, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 554–561.

Gnedenko:1960:AYK

[Gne60] B. V. Gnedenko. Alexander Yakovlevich Khinchin — obituary. Theory of Probability and its Applications, 5(1):1–4, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gnedenko:1990:DPT

[Gne90] B. V. Gnedenko. The Department of Probability Theory at Moscow State University. Theory of Probability and its Applications, 34(1):100–107, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaivoronskii:1992:NBD

[GO92] E. I. Gaivoronskii and E. I. Ostrovskii. A nonasymptotic bound for deviation of a multidimensional empirical distribution function. Theory of Probability and its Applications, 36(3): 592–594, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Godovanchuk:1979:PLD

[God79] V. V. Godovanchuk. Probabilities of large deviations for sums of independent random variables attracted to a stable law. Theory of Probability and its Applications, 23(3):602–608, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Godovanchuk:1982:APL

[God82] V. V. Godovan’chuk. Asymptotic probabilities of large deviations due to large jumps of a Markov process. Theory of Probability and its Applications, 26 (2):314–327, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goldin:1965:LDD

[Gol65] S. V. Gol’din. Limit distributions for the deviation of the zero of one stationary Gaussian process from the zero of another such process which is stationarily connected with the former. Theory of Probability and its Applications, 10(1):164–169, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golinskii:1975:ABP

[Gol75] B. L. Golinskii. On the asymptotic behavior of the prediction error. Theory of Probability and its Applications, 19(4):693–709, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goldshtein:1983:CLT

[Gol83] B. G. Goldshtein. A Central Limit Theorem of “Non-Commutative” probability theory. Theory of Probability and its Applications, 27(4):703–714, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golosov:1985:LTR

[Gol85a] A. O. Golosov. Limit theorems for random walks in symmetric random environments. Theory of Probability and its Applications, 29(2):266–280, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golubev:1985:LDS

[Gol85b] G. K. Golubev. Limit distributions of stopping times in sequential testing of hypotheses. Theory of Probability and its Applications, 29(2):307–317, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golubev:1992:LPN

[Gol92a] G. K. Golubev. LAN in problems of nonparametric estimation of functions and lower bounds for quadratic risks. Theory of Probability and its Applications, 36 (1):152–157, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 143–149.

Golubev:1992:QLF

[Gol92b] G. K. Golubev. Quasi-linear filtering of stationary Gaussian sequences. Theory of Probability and its Applications, 36(4): 793–800, December 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 764–770.

Golubev:1994:NES

[Gol94] G. K. Golubev. Nonparametric estimation of smooth spectral densities of Gaussian stationary sequences. Theory of Probability and its Applications, 38 (4):630–639, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 775–786.

Gomez:1993:EID

[Góm93] G. Gómez. Estimation of induction distributions with doubly censored data and application to AIDS. Theory of Probability and its Applications, 37 (1):32–39, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 36–45.

Goodman:1959:NST

[Goo59] Leo A. Goodman. A note on Stepanov’s tests for Markov chains. Theory of Probability and its Applications, 4(1):89–92, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gordin:1971:BVS

[Gor71] M. I. Gordin. On the behavior of the variances of sums of random variables forming a stationary process. Theory of Probability and its Applications, 16 (3):474–484, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorodeckii:1976:RCM

[Gor76]  V. V. Gorodeckii.  On the

rate of convergence for the multidimensional invariance principle. Theory of Probability and its Applications, 20(3):631–638, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorodetskii:1978:CSS

[Gor78a] V. V. Gorodetskii. On convergence to semi-stable Gaussian processes. Theory of Probability and its Applications, 22 (3):498–508, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorodetskii:1978:SMP

[Gor78b] V. V. Gorodetskii. On the strong mixing property for linear sequences. Theory of Probability and its Applications, 22 (2):411–413, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorodetskii:1980:ACS

[Gor80] V. V. Gorodetskii. Addendum: On convergence to semi-stable Gaussian processes. Theory of Probability and its Applications, 24(2):449, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorbunov:1981:MSI

[Gor81] N. I. Gorbunov. Minimax statistical inference procedures with a possible refusal to make a decision. Theory of Probability and its Applications, 26(1): 30–42, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorodetskii:1983:IPS

[Gor83] V. V. Gorodetskii. The invariance principle for stationary random fields with a strong mixing condition. Theory of Probability and its Applications, 27 (2):380–385, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gordienko:1984:UEC

[Gor84a] E. I. Gordienko. Uniform exponential convergence of Markov processes in metrics corresponding to weak topology. Theory of Probability and its Applications, 28(3):601–603, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorodetskii:1984:CRI

[Gor84b] V. V. Gorodetskii. On the convergence rate in the invariance principle for strongly mixing sequences. Theory of Probability and its Applications, 28 (4):816–821, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gordienko:1985:ASC

[Gor85] E. I. Gordienko. Adaptive strategies for certain classes of controlled Markov processes. Theory of Probability and its Applications, 29(3):504–518, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 29(3), (1985), pp. 488–501.

Goryainov:1989:LBT

[Gor89] V. B. Goryainov. On Lévy– Baxter theorems for stochastic elliptic equations. Theory of Probability and its Applications, 33 (1):164–167, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 176–179.

Gamkrelidze:1971:ADS

[GP71] N. G. Gamkrelidze and Yu. V. Prokhorov. On the approximation of the distributions of sums of lattice random variables when the number of summands is small. Theory of Probability and its Applications, 16(1):144–147, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gerlejn:1979:GFT

[GP79] O. V. Gerlejn and R. Pincus. Goodness-of-fit tests based on a measure of discrepancy between two scalar products. Theory of Probability and its Applications, 23(2):248–258, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Greenwood:1985:LTE

[GP85] P. Greenwood and E. Perkins. Limit theorems for excursions from a moving boundary. Theory of Probability and its Applications, 29(4):731–743, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goryainov:1997:LPD

[GP97] V. V. Goryaĭnov and A. A. Polkovnikov. On limit probability distributions for subcritical branching processes. Theory of Probability and its Applications, 41(2):352–359, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 417–425.

Graversen:1998:ROE

[GP98] S. E. Graversen and G. Peškir.

On the Russian option: the expected waiting time. Theory of Probability and its Applications, 42(3):416–425, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 564–575.

Gotze:1998:SBP

[GPU98] F. Gotze, Yu. V. Prokhorov, and V. V. Ulyanov. On smooth behavior of probability distributions under polynomial mappings. Theory of Probability and its Applications, 42(1):28– 38, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97592.

Garabedyan:1967:ACC

[GR67] S. N. Garabedyan and I. V. Romanovskii. Approximate construction of confidence intervals for parameters of Markov chains. Theory of Probability and its Applications, 12(1):112–115, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Galchuk:1971:DPP

[GR71] L. I. Gal’chuk and B. L. Rozovskii. The “Disorder” problem for a Poisson process. Theory of Probability and its Applications, 16(4):712–716, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:1978:RCL

[GR78] N. G. Gamkrelidze and V. I. Rotar’. On the rate of convergence in the limit theorem for quadratic forms. Theory of Probability and its Applications, 22(2):394–397, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gafurov:1984:ERW

[GR84] M. U. Gafurov and V. I. Rotar’. On the exit of a random walk from a curvilinear boundary. Theory of Probability and its Applications, 28(1):179–184, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Germogenov:1985:SCS

[GR85] A. P. Germogenov and A. F. Ronzhin. A sequential chi-square test. Theory of Probability and its Applications, 29(2):397–403, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grigelionis:1963:CSR

[Gri63] B. Grigelionis. On the convergence of sums of random step processes to a Poisson process. Theory of Probability and its Applications, 8(2):177–182, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grincevicjus:1974:CDS

[Gri74] A. K. Grincevicjus. On the continuity of the distribution of a sum of dependent variables connected with independent walks on lines. Theory of Probability and its Applications, 19(1): 163–168, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grin:1975:SRI

[Gri75] A. G. Grin’. On small random impulsive perturbations of dynamical systems. Theory of Probability and its Applications, 20(1):150–158, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grin:1978:APD

[Gri78] A. G. Grin’. Addendum: On perturbations of dynamical systems by regular Gaussian processes. Theory of Probability and its Applications, 22(4):881, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grigorenko:1979:ASE

[Gri79] L. A. Grigorenko. On the σalgebra of symmetric events for a countable Markov chain. Theory of Probability and its Applications, 24(1):199–204, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grin:1983:CRS

[Gri83] A. G. Grin’. On a condition for regularity of stationary random processes. Theory of Probability and its Applications, 27 (4):850–855, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grishechkin:1987:RBP

[Gri87] S. A. Grishechkin. On the regularity of branching processes with several types of particles. Theory of Probability and its Applications, 31(2):233–243, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gribkova:1989:RCN

[Gri89] N. V. Gribkova. On the rate of convergence to the normal law of finitely truncated linear combinations of order statistics. Theory of Probability and its Applications, 33(4):726–729, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grin:1991:DAS

[Gri91a] A. G. Grin′. On domains of attraction for sums of dependent variables. Theory of Probability and its Applications, 35 (2):241–257, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 255–270.

Grishechkin:1991:BPQ

[Gri91b] S. A. Grishechkin. Branching processes and queueing systems with repeated orders or with random discipline. Theory of Probability and its Applications, 35 (1):38–53, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 35–50.

Grin:1992:NSL

[Gri92a] A. G. Grin′. Norming sequences in the limit theorems for weakly dependent variables. Theory of Probability and its Applications, 36(2):272–288, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 285–300.

Grishechkin:1992:CMJ

[Gri92b] S. A. Grishechkin. Crumpmode-Jagers branching processes as a method of investigating M/G/1 systems with processor sharing. Theory of Probability and its Applications, 36 (1):19–35, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 16–33.

Grishechkin:1992:UBP

[Gri92c] S. A. Grishechkin. On the use of branching processes for finding steady-state distributions in queueing theory. Theory of Probability and its Applications, 36(3):477–493, September 1992.

CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 446–463.

Gribkova:1994:ABE

[Gri94a] N. V. Gribkova. On analogues of Berry–Esseen inequality for truncated linear combinations of order statistics. Theory of Probability and its Applications, 38 (1):142–149, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 176–184.

Grin:1994:ANS

[Gri94b] A. G. Grin’. Addendum: Norming sequences in the limit theorems for weakly dependent variables. Theory of Probability and its Applications, 38(1):175, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grinevich:1994:DAM

[Gri94c] I. V. Grinevich. Domains of attraction of the max-semistable laws under linear and power normalizations. Theory of Probability and its Applications, 38 (4):640–650, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 787–799.

Grishechkin:1995:ITQ

[Gri95] S. A. Grishechkin. Interoutput times in an M/G/1/P S queue with instant feedback. Theory of Probability and its Applications, 39(4):692–696, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 820–824.

Grin:1996:LTS

[Gri96a] A. G. Grin′. Limit theorems for the scheme of series for weakly dependent variables. Theory of Probability and its Applications, 40(4):756–764, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 888–897.

Grinevich:1996:RCL

[Gri96b] I. V. Grinevich. The rate of convergence in limit theorems for max-semistable laws. Theory of Probability and its Applications, 40(3):546–549, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 654–657.

Grusho:1973:RMB

[Gru73] A. A. Grusho. Random mappings with bounded multiplicity. Theory of Probability and its Applications, 17(3):416–425, ???? 1973. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Grusho:1976:LDN

[Gru76] A. A. Grusho. Limit distribution of the number of cyclic points in a random mapping of bounded leight. Theory of Probability and its Applications, 20(3):533–544, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grusho:1986:CCP

[Gru86] A. A. Grusho. On convergence of counting processes associated with U-statistics. Theory of Probability and its Applications, 30(3):626–630, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grusho:1988:PDS

[Gru88] A. A. Grusho. On the problem of the distributions of Ustatistics. Theory of Probability and its Applications, 32(2): 369–373, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gnedenko:1963:KWP

[GS63] B. V. Gnedenko and Kh. V. Smirnov. A. N. Kolmogorov’s work in probability theory. Theory of Probability and its Applications, 8(2):157–164, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grigelionis:1965:CTO

[GS65] B. I. Grigelionis and A. N.

Shiryaev. Criteria of “Truncation” for the optimal stopping time in sequential analysis. Theory of Probability and its Applications, 10(4):541–552, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grigelionis:1966:SPO

[GS66] B. I. Grigelionis and A. N.

Shiryaev. On Stefan’s problem and optimal stopping rules for Markov processes. Theory of Probability and its Applications, 11(4):541–558, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gershanov:1976:REP

[GS76] A. M. Gershanov and S. K.

Shamroni. On randomized estimation in problems with a discrete parameter space. Theory of Probability and its Applications, 21(1):192–197, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grigorchuk:1985:GSC

[GS85] R. I. Grigorchuk and A. M.

Stepin. Gibbs states on countable groups. Theory of Probability and its Applications, 29 (2):359–362, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gafurov:1988:SPE

[GS88] M. U. Gafurov and A. D. Slastnikov. Some problems of the exit of a random walk beyond a curvilinear boundary and large deviations. Theory of Probability and its Applications, 32(2):299–321, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gorgadze:1977:FGM

[GT77] Z. G. Gorgadze and Nguyen Xuy Tien. Families of Gaussian measures in Orlicz sequence spaces. Theory of Probability and its Applications, 21(3):614–620, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gurevich:1965:CIP

[Gur65] B. M. Gurevich. Construction of increasing partitions for special flows. Theory of Probability and its Applications, 10(4):627–645, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gurevich:1968:LLT

[Gur68] B. M. Gurevich. The local limit theorem for Markov chains and regularity conditions. Theory of Probability and its Applications, 13(1):182–188, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gurevich:1989:EPR

[Gur89a] V. A. Gurevich. On estimating the parameter of a regression function without the condition of unbiasedness. Theory of Probability and its Applications, 33(4):729–732, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 785–788.

Gurevich:1989:WEN

[Gur89b] V. A. Gurevich. On weighted M-estimates in nonlinear regression. Theory of Probability and its Applications, 33(2):395–399, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 421–424.

Gurevich:1992:JDR

[Gur92] B. M. Gurevich. On joint distribution of random variables with given cross conditional distributions: Discrete case. Theory of Probability and its Applications, 36(2):371–375, May 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 352–355.

Gusak:1969:JDF

[Gus69] D. V. Gusak. On the joint distribution of the first exit time and exit value for homogeneous processes with independent increments. Theory of Probability and its Applications, 14 (1):14–23, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gusev:1976:AEAa

[Gus76a] S. I. Gusev. Asymptotic expansions associated with some statistical estimators in the smooth case. 1. Expansions of random variables. Theory of Probability and its Applications, 20(3): 470–498, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gusev:1976:AEAb

[Gus76b] S. I. Gusev. Asymptotic expansions associated with some statistical estimators in the smooth case. II. Expansions of moments and distributions. Theory of Probability and its Applications, 21(1):14–33, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gusak:1984:DST

[Gus84] D. V. Gusak. Distribution of the sojourn time of a homogeneous process with independent increments over an arbitrary level. Theory of Probability and its Applications, 28(3):503–514, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gushchin:1993:RFS

[Gus93] A. A. Gushchin. A regular filtered statistical model and families of martingales. Theory of Probability and its Applications, 37(1):43–46, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 37(1), (1992), pp. 49–52.

Gushchin:1996:AOE

[Gus96] A. A. Gushchin. On asymptotic optimality of estimators of parameters under the LAQ condition. Theory of Probability and its Applications, 40(2):261–272, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 286–300.

Guter:1964:PDR

[Gut64] R. S. Guter. On the probability of detecting a region by a linear search. Theory of Probability and its Applications, 9 (2):331–333, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gertsenshtein:1959:WRI

[GV59] M. E. Gertsenshtein and V. B.

Vasil’ev. Waveguides with random inhomogeneities and Brownian motion in the Lobachevsky plane. Theory of Probability and its Applications, 4(4):391–398, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Girko:1986:LTN

[GV86] V. L. Girko and V. V. Vasil’yev.

Limit theorems for normalized spectral functions of nonselfadjoint random Jacobi matrices. Theory of Probability and its Applications, 30(1):1–8, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gutierrez:1993:IPE

[GV93] R. Gutiérrez and M. J. Valderrama. An improved procedure for evaluating the statistical characteristics of a doubly stochastic Poisson process. Theory of Probability and its Applications, 37 (2):340–342, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 358–360.

Greenwood:1993:PSF

[GW93] P. E. Greenwood and W. Wefelmeyer. Partially specified filtered models and efficiency. Theory of Probability and its Applications, 37(1):139–142, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 162–165.

Greenwood:1998:MLE

[GW98] P. E. Greenwood and W. Wefelmeyer. Maximum likelihood estimator and Kullback– Leibler information in misspecified Markov chain models. Theory of Probability and its Applications, 42(1):103–111, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/

article/97599. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 169–178.

Gyachyauskas:1964:SQ

[Gya64a] E. Gyachyauskas. On statistical quadratures. Theory of Probability and its Applications, 9 (4):637–640, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gyachyauskas:1964:SO

[Gya64b] E. Gyachyauskas. Search by an oval. Theory of Probability and its Applications, 9(4): 634–637, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gyongy:1995:SSS

[Gyö95] I. Gyöngy. On the support of the solutions of stochastic differential equations. Theory of Probability and its Applications, 39 (3):519–523, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 649–653.

Gusein-Zade:1966:PCS

[GZ66] S. M. Gusein-Zade. The problem of choice and the sptimal stopping rule for a sequence of independent trials. Theory of Probability and its Applications, 11 (3):472–476, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gusein-Zade:1969:GCW

[GZ69] S. M. Gusein-Zade. On a game connected with the Wiener process. Theory of Probability and its Applications, 14(4):701–704, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hadjiev:1978:FSC

[Had78] D. I. Hadjiev. On the filtering of semimartingales in case of observations of point processes. Theory of Probability and its Applications, 23(1):169–178, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hafner:1973:CPC

[Haf73] Robert Hafner. Convergence of point complexes in Z^{n} to Poisson processes. Theory of Probability and its Applications, 18 (1):131–149, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hagberg:1974:ASP

[Hag74] Jan Hagberg. Approximation of the summation process obtained by sampling from a finite population. Theory of Probability and its Applications, 18(4):753–766, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hajek:1961:LET

[Háj61] J. Hájek. On linear estimation theory for an infinite number of observations. Theory of Probability and its Applications, 6 (2):166–177, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hammersley:1966:SFE

[Ham66] J. M. Hammersley. Subadditive functional expectations. Theory of Probability and its Applications, 11(2):311–313, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hartung:1977:SIR

[Har77] Joachim G. Hartung. Some inequalities for a random function. Theory of Probability and its Applications, 21(3):643–648, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Harlamov:1991:RCI

[Har91] B. P. Harlamov. Random curvilinear integrals and their applications. Theory of Probability and its Applications, 35 (1):54–65, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 118–130.

Hasminskii:1966:SPD

[Has66] R. Z. Has’minskii. On stochastic processes defined by differential equations with a small parameter. Theory of Probability and its Applications, 11(2):211–228, ???? 1966. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Hatunceva:1975:EQC

[Hat75] M. V. Hatunceva. Estimates for quantities in the Central Limit Theorem. Theory of Probability and its Applications, 19(3): 598–604, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Haughton:1993:CCI

[Hau93] D. Haughton. Consistency of a class of information criteria for model selection in nonlinear regression. Theory of Probability and its Applications, 37 (1):47–53, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 105–112.

Hazod:1995:LBP

[Haz95] W. Hazod. On the limit behavior of products of a random number of group valued random variables. Theory of Probability and its Applications, 39(2):249–263, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 374–394.

Hazod:1996:GST

[Haz96] W. Hazod. A generalization of E. Siebert’s theorem on convergence of convolution semigroups and accompanying laws. Theory of Probability and its Applications, 40(4):765–769, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 929–934.

Herrmann:1991:PPC

[Her91]  H. Herrmann.  On a Pa-

ditz’s problem concerning the restricted convergence on a halfaxis in the local Central Limit Theorem. Theory of Probability and its Applications, 35(2): 347–352, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 379–384.

Herrmann:1992:CCC

[Her92] H. Herrmann. Continuation of convergence in the Central Limit Theorem for dependent summands. Theory of Probability and its Applications, 36(2):375–378, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Hida:1970:HAS

[Hid70] Takeyuki Hida. Harmonic analysis on the space of generalized functions. Theory of Probability and its Applications, 15 (1):117–122, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hida:1983:REF

[Hid83] T. Hida. The role of exponential functions in the analysis of generalized Brownian functionals. Theory of Probability and its Applications, 27(3):609–613, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hensz:1978:CLL

[HJ78] E. Hensz and R. Jajte. On a class of limit laws. Theory of Probability and its Applications, 23(1):206–211, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hoffmann-Jorgensen:1994:SD

[HJ94] J. Hoffmann-Jørgensen. Stable densities. Theory of Probability and its Applications, 38 (2):350–355, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 470–476.

Hudson:1995:OSP

[HMV95] W. N. Hudson, J. D. Mason, and J. A. Veeh. Operator stable probability measures: an overview. Theory of Probability and its Applications, 39(2): 264–276, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 357–373.

Hoeffding:1964:TVM

[Hoe64] W. Hoeffding. On a theorem of V. M. Zolotarev. Theory of Probability and its Applications, 9(1):89–91, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hognas:1993:SRT

[Hög93] G. Högnäs. Sequences of random transformations. Theory of Probability and its Applications, 37(1):54–60, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 113–120.

Harel:1994:WCS

[HP94] M. Harel and M. L. Puri. Weak convergence of the simple linear rank statistic under mixing conditions in the nonstationary case. Theory of Probability and its Applications, 38(3):405–422, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 579–599.

Hurd:1997:SAP

[HR97] H. L. Hurd and A. Russek.

Stepanov almost periodically correlated and almost periodically unitary processes. Theory of Probability and its Applications, 41(3):449–467, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97522. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 591–611.

Harison:1993:JMF

[HS93] V. Harison and S. N. Smirnov. Jonction maximale. (French) [Maximal Junction]. Theory of Probability and its Applications, 37(2):372–376, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 439–444.

Heathcote:1967:RTT

[HSVJ67] C. R. Heathcote, E. Seneta, and D. Vere-Jones. A refinement of two theorems in the theory of branching processes. Theory of Probability and its Applications, 12(2):297–301, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hurd:1975:PCP

[Hur75] H. L. Hurd. Periodically correlated processes with discontinuous correlation functions. Theory of Probability and its Applications, 19(4):804–807, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1956:CUD

[Ibr56a] I. A. Ibragimov. On the composition of unimodal distributions. Theory of Probability and its Applications, 1(2):255–260, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1956:TTI

[Ibr56b] I. A. Ibragimov. A theorem in the theory of infinitely divisible laws. Theory of Probability and its Applications, 1 (4):440–444, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1957:RPD

[Ibr57] I. A. Ibragimov. A remark on probability distributions of class L. Theory of Probability and its Applications, 2(1):117–119, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1962:SLT

[Ibr62] I. A. Ibragimov. Some limit theorems for stationary processes. Theory of Probability and its Applications, 7(4):349–382, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1963:CLT

[Ibr63a] I. A. Ibragimov. A Central Limit Theorem for a class of dependent random variables. Theory of Probability and its Applications, 8(1):83–89, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1963:ESF

[Ibr63b] I. A. Ibragimov. On estimation of the spectral function of a stationary Gaussian process. Theory of Probability and its Applications, 8(4):366–401, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1964:ABP

[Ibr64] I. A. Ibragimov. On the asymptotic behavior of the prediction error. Theory of Probability and its Applications, 9(4):627–634, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1965:CES

[Ibr65a] I. A. Ibragimov. A class of estimates for the spectral function of a stationary sequence. Theory of Probability and its Applications, 10(1):123–126, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1965:SSG

[Ibr65b] I. A. Ibragimov. On the spectrum of stationary Gaussian sequences satisfying the strong mixing condition I. Necessary conditions. Theory of Probability and its Applications, 10 (1):85–106, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1966:AGA

[Ibr66] I. A. Ibragimov. On the accuracy of Gaussian approximation to the distribution functions of sums of independent variables. Theory of Probability and its Applications, 11(4):559–579, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1967:CLT

[Ibr67a] I. A. Ibragimov. The Central Limit Theorem for sums of functions of independent variables and sums of the form ∑ f(2^{k}t). Theory of Probability and its Applications, 12(4): 596–607, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1967:MLE

[Ibr67b] I. A. Ibragimov. On maximum likelihood estimation of parameters of the spectral density of stationary time series. Theory of Probability and its Applications, 12(1):115–119, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1967:CCA

[Ibr67c]  I. A. Ibragimov.  On the

Chebyshev-Cramér asymptotic expansions. Theory of Probability and its Applications, 12 (3):455–469, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1970:SSG

[Ibr70] I. A. Ibragimov. On the spectrum of stationary Gaussian sequences satisfying the strong mixing condition. II. Sufficient conditions. mixing rate. Theory of Probability and its Applications, 15 (1):23–36, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1974:PSF

[Ibr74] I. A. Ibragimov. Properties of sample functions for stochastic processes and embedding theorems. Theory of Probability and its Applications, 18(3):442–453, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1975:NCL

[Ibr75a] I. A. Ibragimov. A note on the Central Limit Theorems for dependent random variables. Theory of Probability and its Applications, 20(1):135–141, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1975:RET

[Ibr75b] I. A. Ibragimov. A remark on the ergodic theorem for Markov chains. Theory of Probability and its Applications, 20(1):174–176, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1978:DID

[Ibr78] I. A. Ibragimov. On determining an infinitely divisible distribution function by its values on a halfline. Theory of Probability and its Applications, 22(2):384–390, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1984:SCT

[Ibr84] I. A. Ibragimov. On smoothness conditions for trajectories of random functions. Theory of Probability and its Applications, 28 (2):240–262, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1989:PAT

[Ibr89] I. A. Ibragimov. A probabilistic analogue of a theorem of Córdoba and Fefferman. Theory of Probability and its Applications, 33 (1):137–139, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 147–149.

Ibragimov:1994:CAE

[Ibr94] I. A. Ibragimov. On a characteristic of the accuracy of estimating a distribution density. Theory of Probability and its Applications, 38(2):355–360, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 425–431.

Ibragimov:1959:UGS

[IC59] I. A. Ibragimov and K. E.

Chernin. On the unimodality of geometric stable laws. Theory of Probability and its Applications, 4(4):417–419, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ilinskii:1998:PCP

[IC98] A. I. Il′inskiĭ and G. P. Chistyakov. On the problem of characterizing probability distributionsby absolute moments of partial sums. Theory of Probability and its Applications, 42 (3):426–432, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 473–481.

Ignatov:1982:CAA

[Ign82] Ts. Ignatov. On a constant arising in the asymptotic theory of symmetric groups, and on Poisson–Dirichlet measures. Theory of Probability and its Applications, 27(1):136–147, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1975:SE

[IH75a] I. A. Ibragimov and R. Z. Has’minskii. On sequential estimation. Theory of Probability and its Applications, 19(2): 233–244, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1975:SEL

[IH75b] I. A. Ibragimov and R. Z. Has’minskii. On sequential estimation of the location parameter for families of distributions with discontinuous densities. Theory of Probability and its Applications, 19(4):669–682, ???? 1975.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1976:LAN

[IH76a] I. A. Ibragimov and R. Z.

Has’minskii. Local asymptotic normality for non-identically distributed observations. Theory of Probability and its Applications, 20(2):246–260, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 251–266.

Ibragimov:1976:PML

[IH76b] I. A. Ibragimov and R. Z.

Has’minskii. Properties of maximum likelihood and Bayes’ estimators for non-identically distributed observations. Theory of Probability and its Applications, 20(4):689–697, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanova:1977:ST

[II77] T. V. Ivanova and G. I. Ivchenko.

A sequential test. Theory of Probability and its Applications, 21(4):833–836, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Inglot:1991:AMG

[IJL91] T. Inglot, T. Jurlewicz, and T. Ledwina. Asymptotics for multinomial goodness of fit tests for a simple hypothesis. Theory of Probability and its Applications, 35(4):771–777, ???? 1991.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 797–803.

Ilin:1964:EBM

[IK64] A. M. Il’in and R. Z. Khas’minskii. On equations of Brownian motion. Theory of Probability and its Applications, 9(3):421–444, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1973:ABSb

[IK73a] I. A. Ibragimov and R. Z. Khas’minskii. Asymptotic behavior of some statistical estimators II. Limit theorems for the a posteriori density and Bayes’ estimators. Theory of Probability and its Applications, 18 (1):76–91, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1973:ABSa

[IK73b] I. A. Ibragimov and R. Z. Khas’minskii. Asymptotic behavior of statistical estimators in the smooth case. I. Study of the likelihood ratio. Theory of Probability and its Applications, 17 (3):445–462, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1974:AAS

[IK74a] I. A. Ibragimov and R. Z. Khas’minskii. Asymptotic analysis of statistical estimators for the “Almost smooth” case. Theory of Probability and its Applications, 18(2):241–252, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1974:MGB

[IK74b] I. A. Ibragimov and R. Z.

Khas’minskii. On moments of generalized Bayesian estimators and maximum likelihood estimators. Theory of Probability and its Applications, 18(3):508–520, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1979:SAR

[IK79] I. A. Ibragimov and N. M.

Khalfina. Some asymptotic results related to the Chauvenet test. Theory of Probability and its Applications, 23(3):593–597, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1981:ESD

[IK81] I. A. Ibragimov and R. Z.

Khas’minskii. Estimates of the signal, its derivatives and point of maximum for Gaussian distributions. Theory of Probability and its Applications, 25(4):703–720, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1982:BRN

[IK82] I. A. Ibragimov and R. Z.

Khas’minskii. Bounds for the risks of non-parametric regression estimates. Theory of Probability and its Applications, 27 (1):84–99, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1983:EDD

[IK83] I. A. Ibragimov and R. Z. Khas’minskii. Estimation of distribution density belonging to a class of entire functions. Theory of Probability and its Applications, 27(3):551–562, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1985:NEV

[IK85] I. A. Ibragimov and R. Z. Khas’minskii. On nonparametric estimation of the value of a linear functional in Gaussian white noise. Theory of Probability and its Applications, 29 (1):18–32, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1988:ELF

[IK88] I. A. Ibragimov and R. Z. Khas’minskii. Estimation of linear functionals in Gaussian noise. Theory of Probability and its Applications, 32(1):30– 39, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1993:EPU

[IK93] I. A. Ibragimov and R. Z. Khas′minskiĭ. Estimating a parameter under LAN conditions with restrictions. Theory of Probability and its Applications, 37(4):589–598, December 1993.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 609–620.

Ibragimov:1999:EPC

[IK99] I. A. Ibragimov and R. Z.

Khas’minskii. Estimation problems for coefficients of stochastic partial differential equations. Part I. Theory of Probability and its Applications, 43(3): 370–387, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97698.

Ilinskii:1976:ZAC

[Il’76a] A. I. Il’inskii. On the zeros and the argument of a characteristic function. Theory of Probability and its Applications, 20 (2):410–415, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1976:ANC

[IL76b] G. I. Ivchenko and V. V. Levin.

Asymptotic normality of a class of statistics in the multinomial scheme. Theory of Probability and its Applications, 21(1): 188–192, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1978:ANS

[IL78] G. I. Ivchenko and V. V. Levin.

Asymptotic normality in the scheme of simple random sampling without replacement. Theory of Probability and its Applications, 23(1):93–105, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 23, (1978), pp. 93–105.

Inglot:1994:CTL

[IL94] T. Inglot and T. Ledwina. Cramér-Type large deviations for some U-statistics. Theory of Probability and its Applications, 38(4):651–659, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 858–868.

Ivchenko:1965:ARF

[IM65a] G. I. Ivchenko and Yu. I. Medvedev. Asymptotic representations of finite differences of a power function at an arbitrary point. Theory of Probability and its Applications, 10(1):139–144, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1965:SMT

[IM65b] G. I. Ivchenko and Yu. I. Medvedev. Some multidimensional theorems on a classical problem of permutations. Theory of Probability and its Applications, 10(1):144–149, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1966:ABN

[IM66] G. I. Ivchenko and Yu. J.

Medvedev. Asymptotic behavior of the number of particle complexes in a classical allocation problem. Theory of Probability and its Applications, 11 (4):619–626, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1971:ENRa

[IM71a] I. A. Ibragimov and N. B.

Maslova. On the expected number of real zeros of random polynomials I. Coefficients with zero means. Theory of Probability and its Applications, 16(2):228–248, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1971:ENRb

[IM71b] I. A. Ibragimov and N. B.

Maslova. On the expected number of real zeros of random polynomials. II. Coefficients with non-zero means. Theory of Probability and its Applications, 16 (3):485–493, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1979:SSH

[IM79] G. I. Ivchenko and Yu. I.

Medvedev. Separable statistics and hypothesis testing. The case of small samples. Theory of Probability and its Applications, 23 (4):764–775, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1981:DSH

[IM81] G. I. Ivchenko and Yu. I. Medvedev. Decomposable statistics and hypothesis testing for grouped data. Theory of Probability and its Applications, 25 (3):540–551, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1987:ERC

[IM87] G. I. Ivchenko and Yu. I. Medvedev. Estimation of the rate of convergence in limit theorems for divisible statistics. Theory of Probability and its Applications, 31(1):81–87, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1989:ANR

[IM89] G. I. Ivchenko and Sh. A. Mirakhmedov. On the asymptotic normality of randomized separable statistics in a generalized allocation scheme. Theory of Probability and its Applications, 33 (4):749–755, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanova:1993:SIS

[IM93] B. G. Ivanova and E. Mertsbakh. Set-indexed stochastic processes and predictability. Theory of Probability and its Applications, 37(1):61–66, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 57–63.

Ignatyuk:1993:CSQ

[IMS93a] I. A. Ignatyuk, V. A. Malyshev, and V. Sidoravichius. Convergence of the stochastic quantization method. I. Theory of Probability and its Applications, 37 (2):209–221, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 241–253.

Ignatyuk:1993:CMS

[IMS93b] I. A. Ignatyuk, V. A. Malyshev, and V. Sidoravicius. Convergence of a method of the stochastic quantization II. Theory of Probability and its Applications, 37(4): 599–620, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 621–647.

Ivanov:1978:DTF

[IN78] V. A. Ivanov and A. E. Novikov.

On the distribution of the time up to the first occurrence of a given number of different ℓ-tuple series. Theory of Probability and its Applications, 22(3):533–542, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ingster:1983:AET

[Ing83] Yu. I. Ingster. On the asymptotic efficiency of tests for a simple hypothesis against a composite alternative. Theory of Probability and its Applications, 27 (3):628–633, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ingster:1984:AOB

[Ing84] Yu. I. Ingster. Asymptotically optimal Bayes tests for composite hypotheses. Theory of Probability and its Applications, 28 (4):775–794, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 28(4), (1983), pp. 738–757.

Ingster:1986:AOT

[Ing86] Yu. I. Ingster. Asymptotically optimal tests for verifying composite finite-parameter hypotheses. Theory of Probability and its Applications, 30(2):311–332, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ingster:1987:MTN

[Ing87] Yu. I. Ingster. Minimax testing of nonparametric hypotheses on a distribution density in the L_{p} metrics. Theory of Probability and its Applications, 31 (2):333–337, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ingster:1988:CMP

[Ing88] Yu. I. Ingster. On the comparison of the minimax properties of Kolmogorov’s ω^{2} test and χ^{2} test. Theory of Probability and its Applications, 32(2):346–350, ???? 1988. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Ingster:1989:MTS

[Ing89] Yu. I. Ingster. Minimax testing of a simple hypothesis against a weakly nonparametric set of alternatives. Theory of Probability and its Applications, 33 (3):552–556, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 592–596.

Ingster:1995:MTH

[Ing95] Yu. I. Ingster. Minimax testing of hypotheses on the distribution density for ellipsoids in l_{p}. Theory of Probability and its Applications, 39(3):417–436, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 530–553.

Ibragimov:1987:SPN

[INK87] I. A. Ibragimov, A. S. Nemirovskii, and R. Z. Khas’minskii. Some problems on nonparametric estimation in Gaussian white noise. Theory of Probability and its Applications, 31(3):391–406, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1966:ERL

[IO66] I. A. Ibragimov and L. V. Osipov. On an estimate of the remainder in Lindeberg’s theorem. Theory of Probability and its Applications, 11(1):125–128, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Iosifescu:1968:LIL

[Ios68]  M. Iosifescu.  The law of

the iterated logarithm for a class of dependent random variables. Theory of Probability and its Applications, 13(2):304–313, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Iosifesu:1970:ALI

[Ios70] M. Iosifesu. Addendum: The law of the iterated logarithm for a class of dependent random variables. Theory of Probability and its Applications, 15(1):160, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1974:RAD

[IP74] I. A. Ibragimov and E. L. Presman. On the rate of approach of the distributions of sums of independent random variables to accompanying distributions. Theory of Probability and its Applications, 18 (4):713–727, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivkovic:1989:TCG

[IP89] Z. Ivković and P. M. Perunicić. Transformation of a continuous Gaussian process by rarefying its innovation process. Theory of Probability and its Applications, 33(4):738–743, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1970:CBT

[IR70] I. A. Ibragimov and Yu. A.

Rozanov. On the connection between two characteristics of dependence of Gaussian random vectors. Theory of Probability and its Applications, 15(2): 295–299, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivkovich:1971:CHC

[IR71] Z. Ivkovich and Yu. A. Rozanov.

On the canonical Hida-Craméer representation for random processes. Theory of Probability and its Applications, 16(2):351–356, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1968:ABP

[IS68] I. A. Ibragimov and V. N. Solev.

The asymptotic behavior of the prediction error of a stationary sequence with a spectral density of special type. Theory of Probability and its Applications, 13 (4):703–707, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibramkhalilov:1974:DMW

[IS74] I. Sh. Ibramkhalilov and A. V.

Skorokhod. Determination of the mean of a Wiener process observed on an infinite interval. Theory of Probability and its Applications, 18(4):767–771, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:1998:SCE

[IS98] R. Ibragimov and Sh. Sharakhmetov. Short communications: On an exact constant for the Rosenthal inequality. Theory of Probability and its Applications, 42 (2):294–302, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97615.

Isakova:1985:LTD

[Isa85] T. I. Isakova. Limit theorem for diffusion processes with switchings. Theory of Probability and its Applications, 29(3):554–558, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ismagilov:1973:TAR

[Ism73] R. S. Ismagilov. On the tail algebra for random sequences. Theory of Probability and its Applications, 18(1):180–186, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanov:1979:SLT

[IT79] V. A. Ivanov and S. Yu. Terebulin. Some limit theorems in a non-equiprobable scheme for the allocation of particles in groups. Theory of Probability and its Applications, 23(3):621–627, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1983:ESF

[IT83] G. I. Ivchenko and E. E. Timonina. Estimating the size of a finite population. Theory of Probability and its Applications, 27 (2):403–406, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1989:NST

[IT89] G. I. Ivchenko and S. V.

Tsukanov. On new statistical tests in the multinomial scheme and their asymptotic properties. Theory of Probability and its Applications, 33(1):187–191, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 201–205.

Ibarrola:1993:HTM

[IV93] P. Ibarrola and R. Vélez. Hypothesis testing for multidimensional Gaussian process. Theory of Probability and its Applications, 37(1):142–145, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 175–178.

Ivanov:1960:ANC

[Iva60] V. A. Ivanov. On the average number of crossings of a level by the sample functions of a stochastic process. Theory of Probability and its Applications, 5 (3):319–323, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivankov:1965:NRM

[Iva65] S. A. Ivankov. On the number of repetitions in Markov chains. Theory of Probability and its Applications, 10(3):505–507, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivankov:1970:CCR

[Iva70] S. A. Ivankov. Carleman’s classes for random processes. Theory of Probability and its Applications, 15(1):115–117, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanov:1977:AED

[Iva77] A. V. Ivanov. An asymptotic expansion for the distribution of the least squares estimator of the non-linear regression parameter. Theory of Probability and its Applications, 21(3):557–570, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanov:1986:GSS

[Iva86] V. A. Ivanov. On a generalization of separable statistics. Theory of Probability and its Applications, 30(4):835–840, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanovskaya:1992:C

[Iva92] Regina I. Ivanovskaya. Congratulations. Theory of Probability and its Applications, 36 (3):555, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Ivchenko:1971:LTO

[Ivc71a] G. I. Ivchenko. Limit theorems in an occupancy problem. Theory of Probability and its Applications, 16(2):293–307, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1971:LDO

[Ivc71b] G. I. Ivchenko. On limit distributions for the order statistics of the multinomial distribution. Theory of Probability and its Applications, 16(1):102–115, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1973:ABD

[Ivc73] G. I. Ivchenko. On the asymptotic behavior of degrees of vertices in a random graph. Theory of Probability and its Applications, 18(1):188–195, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1974:SCR

[Ivc74a] G. I. Ivchenko. The strength of connectivity of a random graph. Theory of Probability and its Applications, 18(2):396–403, ???? 1974. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1974:VSS

[Ivc74b] G. I. Ivchenko. Variational series for a scheme of summing independent variables. Theory of Probability and its Applications, 18(3):531–545, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivcenko:1975:LDM

[Ivc75a] G. I. Ivcenko. On limit distributions of middle order statistics for a double sequence. Theory of Probability and its Applications, 19(2):267–277, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivchenko:1975:CBP

[Ivc75b] G. I. Ivchenko. On comparison of binomial and Poisson distributions. Theory of Probability and its Applications, 19(3):584–587, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 19, (1974), pp. 612– 615.

Ivchenko:1975:WTT

[Ivc75c] G. I. Ivchenko. Waiting time and testing hypotheses in a multinomial scheme. Theory of Probability and its Applications, 19 (4):807–812, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 19, (1974), pp. 839–844.

Ivchenko:1976:SDC

[Ivc76] G. I. Ivchenko. On some distributions connected with waiting time in a polynomial scheme. Theory of Probability and its Applications, 20(3):545–559, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 557– 570.

Ivchenko:1985:SSI

[Ivc85] G. I. Ivchenko. Some statistical inference problems for stratified populations. Theory of Probability and its Applications, 29 (1):114–119, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivkovich:1968:LAR

[Ivk68] Z. A. Ivkovich. On linear approximation of random processes in the case of random observation. Theory of Probability and its Applications, 13(2):324–326, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivkovic:1986:STH

[Ivk86] Z. A. Ivković. Spectral type of Hermite polynomial of a Wiener process. Theory of Probability and its Applications, 30(1): 162–164, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivkovic:1994:SMN

[Ivk94]  Z. A. Ivković.  A stochas-

tic measure and nonlinear approximation of some random fields. Theory of Probability and its Applications, 38(3):525–528, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 666–670.

Ivnitskii:1977:RCS

[Ivn77] V. A. Ivnitskii. Reconstruction of characteristics of a GI|M|1 system from observations on the outgoing flow. Theory of Probability and its Applications, 22 (1):184–187, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivnitskii:1982:CIS

[Ivn82] V. A. Ivnitskii. On a condition of invariance of stationary probabilities for queueing networks. Theory of Probability and its Applications, 27(1):196–201, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivnitskii:1990:ICS

[Ivn90] V. A. Ivnitskiĭ. An invariance condition for the stationary probabilities of the states of networks of single-line queueing systems. Theory of Probability and its Applications, 34(3):519–523, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 576–580.

Ivnitskii:1993:ISS

[Ivn93] V. A. Ivnitskiĭ. The invariance of the steady state distribution of a multichannel queueing system serving a finite number of different sources of customers under the absolute priority of an arriving customer. Theory of Probability and its Applications, 37 (2):376–379, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 407–410.

Ivnitskii:1998:SCI

[Ivn98] V. A. Ivnitskii. Short communication: On the invariance of stationary probabilities of states of a closed star-shaped queueing network with state-dependent transition probabilities. Theory of Probability and its Applications, 42(1):162–167, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97600.

Jagers:1968:ADB

[Jag68]  P. Jagers.  Age-dependent

branching processes allowing immigration. Theory of Probability and its Applications, 13 (2):225–236, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jakubowski:1998:SCA

[Jak98] A. Jakubowski. Short communication: The almost sure Skorokhod representation for subsequences in nonmetric spaces. Theory of Probability and its Applications, 42(1):167–174, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97605. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 209–216.

Jamukov:1978:EGD

[Jam78] G. I. Jamukov. Estimates of generalized Dudley metrics in spaces of finite-dimensional distributions. Theory of Probability and its Applications, 22 (3):579–583, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Janev:1976:SBP

[Jan76] N. M. Janev. On the statistics of branching processes. Theory of Probability and its Applications, 20(3):612–622, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jankovic:1993:PTU

[Jan93] S. Jankovic. Preservation of type under mixing. Theory of Probability and its Applications, 37(3): 529–535, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jesiak:1980:ADF

[Jes80] B. Jesiak. On analytic distribution functions and analytic properties of infinitely divisible distribution functions. Theory of Probability and its Applications, 24(4):824–830, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jesiak:1981:UDL

[Jes81] B. Jesiak. On the unique determination of Lévy’s spectral functions and parameters of infinitely divisible distributions. Theory of Probability and its Applications, 26(1):156–161, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Johnson:1990:SRZ

[JK90] G. W. Johnson and G. Kallianpur. Some remarks on Y. Z. Hu and P.-A. Meyer’s paper and infinite-dimensional calculus on finitely additive canonical Hilbert space: “Wiener chaos and the Feynman integral” [in Séminaire de Probabilités, XXII, 51–71, Lectures Notes in Math., 1321, Springer, Berlin, 1988; MR 89m:60193]. Theory of Probability and its Applications, 34 (4):679–689, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 742–752.

James-Levy:1956:NIL

[JL56] U. E. James-Levy. A nomogram for the integral law of Student’s distribution. Theory of Probability and its Applications, 1(2):246–248, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jakubowski:1994:FCL

[JM94] A. Jakubowski and J. Mémin. Functional Central Limit Theorems for a class of quadratic forms in independent random variables. Theory of Probability and its Applications, 38(3): 423–432, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 600–612.

Joshi:1975:SSP

[JP75] S. W. Joshi and G. P. Patil. Sum-symmetric power series distributions and minimum variance unbiased estimation. Theory of Probability and its Applications, 19(3):587–594, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

James:1997:CEE

[JP97] N. James and Y. Peres. Cutpoints and exchangeable events for random walks. Theory of Probability and its Applications, 41(4):666–677, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97574.

Jesiak:1982:NVL

[JR82] B. Jesiak and H.-J. Rossberg.

New versions of the Lindeberg– Feller theorem. Theory of Probability and its Applications, 26 (4):845–849, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurek:1989:CCR

[JR89] Z. J. Jurek and J. Rosiński. Continuity of certain random integral mappings and the uniform integrability of infinitely divisible measures. Theory of Probability and its Applications, 33 (3):523–535, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 560–572.

Joshi:1991:AEV

[JR91] S. N. Joshi and A. L. Rukhini.

Asymptotic estimation of variance. Theory of Probability and its Applications, 35(4):777–784, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 803–809.

Jurinskii:1972:BCF

[Jur72] V. V. Jurinskii. Bounds for characteristic functions of certain degenerate multidimensional distributions. Theory of Probability and its Applications, 17 (1):101–113, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurinskii:1974:EBL

[Jur74a]  V. V. Jurinskii.  Exponen-

tial bounds for large deviations. Theory of Probability and its Applications, 19(1):154–155, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurinskii:1974:EBS

[Jur74b] V. V. Jurinskii. Exponential bounds for smooth fields. Theory of Probability and its Applications, 19(1):221–226, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurinskii:1975:SIE

[Jur75a] A. A. Jurinskii. A smoothing inequality for estimates of the Lévy–Prokhorov distance. Theory of Probability and its Applications, 20(1):1–10, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurinskii:1975:IDD

[Jur75b] V. V. Jurinskii. On infinitely divisible distributions. Theory of Probability and its Applications, 19(2):297–308, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurek:1983:DNA

[Jur83] Z. J. Jurek. Domains of normal attraction for G-stable measures on R^{d}. Theory of Probability and its Applications, 27 (2):419–424, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1983:ABI

[KA83] V. M. Kruglov and S. N.

Antonov. On the asymptotic behavior of infinitely divisible distributions in Banach space. Theory of Probability and its Applications, 27(4):667–687, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1985:OMA

[KA85] V. M. Kruglov and S. N.

Antonov. Once more on the asymptotic behavior of infinitely divisible distributions in Banach space. Theory of Probability and its Applications, 29(4):766–775, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1987:AOM

[KA87] V. M. Kruglov and S. N.

Antonov. Addendum: Once more on the asymptotic behavior of infinitely divisible distributions in Banach space. Theory of Probability and its Applications, 31(1):175, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khashimov:1999:ECR

[KA99] A. Khashimov and G. R. Abdurakhmanov. Estimates of the convergence ratein the Central Limit Theorem for generalized U-statistics. Theory of Probability and its Applications, 43 (1):63–74, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97666.

Kabanov:1974:RFW

[Kab74] Yu. M. Kabanov. Representation of functionals of Wiener and Poisson processes in the form of stochastic integrals. Theory of Probability and its Applications, 18(2):362–365, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1975:IRF

[Kab75] Yu. M. Kabanov. Integral representations of functionals of processes with independent increments. Theory of Probability and its Applications, 19(4):853–857, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1976:ESI

[Kab76] Yu. M. Kabanov. On extended stochastic intervals. Theory of Probability and its Applications, 20(4):710–722, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1978:CCP

[Kab78] Yu. M. Kabanov. The capacity of a channel of the Poisson type. Theory of Probability and its Applications, 23(1):143–147, ???? 1978. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1986:ECV

[Kab86] Yu. M. Kabanov. An estimate of closeness in variation of probability measures. Theory of Probability and its Applications, 30 (2):413–417, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1993:PRS

[Kab93] Yu. M. Kabanov. On the probabilistic representation of a solution of the telegraph equation. Theory of Probability and its Applications, 37(2):379–380, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 425–426.

Kabanovich:1996:TDS

[Kab96] V. I. Kabanovich. On time dependency in a simple random walk. Theory of Probability and its Applications, 40(1):156–158, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 175–177.

Kagan:1986:IPE

[Kag86a] A. M. Kagan. Information property of exponential families. Theory of Probability and its Applications, 30(4):831–835, ???? 1986. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:1986:SMP

[Kag86b] A. M. Kagan. A simple modification of Pitman estimates for a location parameter. Theory of Probability and its Applications, 30(3):598–603, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:1988:CTD

[Kag88] A. M. Kagan. A class of twodimensional distributions arising in connection with the Cramér and Darmois-Skitovich Theorems. Theory of Probability and its Applications, 32(2):323–325, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:1989:NCD

[Kag89] A. M. Kagan. New classes of dependent random variables and a generalization of the Darmois– Skitovich Theorem to several forms. Theory of Probability and its Applications, 33(2):286–295, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 305–314.

Kagan:1990:GCR

[Kag90] A. M. Kagan. A generalized condition for random vectors to be identically distributed related to the analytical theory of linear forms of independent random variables. Theory of Probability and its Applications, 34 (2):327–332, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 370–375.

Kahle:1993:IPC

[Kah93] W. Kahle. Improved parametric confidence estimation — the calculation of Bartlett adjustment for the two-parameter exponential family. Theory of Probability and its Applications, 37 (2):343–345, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 369–371.

Kaimanovich:1986:UDC

[Kai86] V. A. Kaimanovich. The uniform distribution on compact homogeneous spaces and the Kantorovich-rubinshtein metric. Theory of Probability and its Applications, 30(4):828–831, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kakosyan:1978:ACL

[Kak78] A. V. Kakosyan. An admissibility condition for a least squares estimator in the class of polynomial estimators. Theory of Probability and its Applications, 23 (1):201–206, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalmykov:1962:POO

[Kal62] G. I. Kalmykov. On the partial ordering of one-dimensional Markov processes. Theory of Probability and its Applications, 7(4):456–459, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalinin:1967:CAE

[Kal67] V. M. Kalinin. Convergent and asymptotic expansions for probability distributions. Theory of Probability and its Applications, 12(1):22–35, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalmykov:1969:SPF

[Kal69] G. I. Kalmykov. Semiordering of the probabilities of the first passage time of Markov processes. Theory of Probability and its Applications, 14(4):704–710, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalasnikov:1976:CPP

[Kal76] V. V. Kalasnikov. Certain properties of piecewise-linear Markov processes. Theory of Probability and its Applications, 20(3): 560–571, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalashnikov:1977:SAQ

[Kal77a] V. V. Kalashnikov. Stability analysis in queueing problems by a method of trial functions. Theory of Probability and its Applications, 22(1):86–103, ???? 1977.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 22, (1977), pp. 89–105.

Kalenskii:1977:LTP

[Kal77b] L. A. Kalenskii. Limit theorems for products of independent triangular matrices. Theory of Probability and its Applications, 22(1):160–166, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalashnikov:1978:UEC

[Kal78] V. V. Kalashnikov. Uniform estimation of the convergence rate in a renewal theorem for the case of discrete time. Theory of Probability and its Applications, 22 (2):390–394, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalinkin:1982:EPB

[Kal82] A. V. Kalinkin. Extinction probability of a branching process with interaction of particles. Theory of Probability and its Applications, 27(1):201–205, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kallenberg:1993:OSR

[Kal93a] O. Kallenberg. From optional skipping to random time change — on some recent advances in exchangeability theory. Theory of Probability and its Applications, 37(1):67–75, March 1993.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 64–74.

Kallenberg:1993:IME

[Kal93b] W. C. M. Kallenberg. Interpretation and manipulation of Edgeworth expansions. Theory of Probability and its Applications, 37(1):145–147, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 179–181.

Kalinkin:1999:FPB

[Kal99] A. V. Kalinkin. Final probabilities for a branching process with interaction of particles and an epidemic process. Theory of Probability and its Applications, 43(4):633– 640, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97720.

Kandelaki:1988:DCO

[Kan88] N. P. Kandelaki. Distributions connected with orthogonal Gaussian systems. Theory of Probability and its Applications, 32 (1):167–168, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kapitonov:1974:RCS

[Kap74] V. M. Kapitonov. On the rate of convergence of a sequence of distributions defined by an autoregression scheme on a compact group. Theory of Probability and its Applications, 18 (3):582–587, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kaplan:1991:DPR

[Kap91] E. I. Kaplan. On the disorder problem for random fields. Theory of Probability and its Applications, 35(2):352–358, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 353–358.

Kartashov:1965:DFR

[Kar65] G. D. Kartashov. On determining a functional relation between random variables. Theory of Probability and its Applications, 10(3):528–535, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kartashov:1969:SPP

[Kar69] G. D. Kartashov. On some probabilistic problems in reliability theory with constraints. Theory of Probability and its Applications, 14(4):595–611, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Karmanov:1975:CRS

[Kar75] V. G. Karmanov. On convergence of a random search method in convex minimization problems. Theory of Probability and its Applications, 19(4):788–794, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kartashov:1980:PEC

[Kar80] N. V. Kartashov. Power estimates for the convergence rate in a renewal theorem. Theory of Probability and its Applications, 24(3):606–612, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kartashov:1981:UAR

[Kar81]  N. V. Kartashov.  A uni-

form asymptotic renewal theorem. Theory of Probability and its Applications, 25(3):589–592, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kartashov:1985:OML

[Kar85] N. V. Kartashov. Operator methods in limit theorems for Markov processes. Theory of Probability and its Applications, 29(4): 828–829, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 29(4), (1984), pp. 792–793.

Kartashov:1986:ITEa

[Kar86a] N. V. Kartashov. Inequalities in theorems of ergodicity and stability for Markov chains with common phase space. I. Theory of Probability and its Applications, 30(2):247–259, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(2), (1985), pp. 230–240.

Kartashov:1986:ITEb

[Kar86b] N. V. Kartashov. Inequalities in theorems of ergodicity and stability for Markov chains with common phase space. II. Theory of Probability and its Applications, 30(3):507–515, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(3), (1985), pp. 478–485.

Karpelevich:1995:EMS

[Kar95] F. I. Karpelevich. On estimation of maxima of sums of random variables indexed by edges of graphs. Theory of Probability and its Applications, 39(4):696–702, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 833–840.

Katyshev:1975:TGV

[Kat75] P. K. Katyshev. Transmission of a Gaussian vector through a channel with noiseless feedback. Theory of Probability and its Applications, 20(1):185–193, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Katona:1978:IDL

[Kat78] G. O. H. Katona. Inequalities for the distribution of the length of random vector sums. Theory of Probability and its Applications, 22(3):450–464, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kaverin:1991:RLT

[Kav91] S. V. Kaverin. A refinement of limit theorems for critical branching processes with emigration. Theory of Probability and its Applications, 35(3): 574–580, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 570–575.

Korolyuk:1986:ANS

[KB86] V. S. Korolyuk and Yu. V. Borovskikh. Approximation of nondegenerate U-statistics. Theory of Probability and its Applications, 30(3):439–450, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1989:CRD

[KB89] V. S. Korolyuk and Yu. V. Borovskikh. Convergence rate for degenerate von Mises functionals. Theory of Probability and its Applications, 33(1):125–135, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 136–146.

Korolyuk:1992:WCD

[KB92] V. S. Korolyuk and Yu. V.

Borovskikh. Weak convergence of distributions of UH-Statistics to stable laws. Theory of Probability and its Applications, 36(3): 596–600, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Kruglov:1997:LTM

[KB97] V. M. Kruglov and Tchzhan Bo. Limit theorems for the maximal random sums. Theory of Probability and its Applications, 41(3):468–478, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97524. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 520–532.

Kolchin:1974:CLT

[KC74] V. F. Kolchin and V. P.

Chistyakov. On a combinatorial limit theorem. Theory of Probability and its Applications, 18 (4):728–739, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1975:LDN

[KC75a] V. F. Kolchin and V. P.

Chistyakov. Limit distributions of the number of non-occurring s-tuples in a multinomial scheme. Theory of Probability and its Applications, 19(4):822–830, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1975:LDS

[KC75b] V. F. Kolchin and V. P. Chistyakov. On limit distributions of a statistic. Theory of Probability and its Applications, 19(2):359–365, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kartashov:1979:FIE

[KC79] G. D. Kartashov and N. M. Chiganova. On finding interval estimates by the method of L. N. Bol’shev and É. A. Loginov. Theory of Probability and its Applications, 23(3):649–652, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kotelnikova:1983:CPE

[KC83] V. F. Kotel’nikova and E. V. Chmaladze. On computing the probability of an empirical process not crossing a curvilinear boundary. Theory of Probability and its Applications, 27(3): 640–648, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Koryakin:1991:EFG

[KC91] A. I. Koryakin and N. N. Chentsov. On estimating a function given at random nodes by the method of least squares. Theory of Probability and its Applications, 35(4):788–792, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kassalygo:1987:UGF

[KD87] L. A. Kassalygo and R. L. Dobrushin. Uniqueness of a Gibes field with random potential — an elementary approach. Theory of Probability and its Applications, 31(4):572–589, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kelmans:1973:AFP

[Kel73] A. K. Kel’mans. Asymptotic formulas for the probability of kconnectedness of random graphs. Theory of Probability and its Applications, 17(2):243–254, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kelbert:1977:CDS

[Kel77] M. Ya. Kel’bert. On convergence of discrete schemes to continuous ones in some problems of sequential analysis. Theory of Probability and its Applications, 21 (3):605–614, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kelbert:1990:CLT

[Kel90] M. Ya. Kel′bert. On the Central Limit Theorem for Jackson networks on countable graphs. Theory of Probability and its Applications, 34(4):713–714, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 778–780.

Kendall:1964:SRW

[Ken64] David G. Kendall. Some recent work and further problems in the theory of queues. Theory of Probability and its Applications, 9(1):1–13, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kendall:1987:FDA

[Ken87] David G. Kendall. Further developments and applications of the statistical theory of shape. Theory of Probability and its Applications, 31(3):407–412, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kendall:1995:HLO

[Ken95] David G. Kendall. How to look at objects in a 5-dimensional shape space. I: Looking at distributions. Theory of Probability and its Applications, 39(2):277–281, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 242–247.

Klimov:1967:SSR

[KF67] G. P. Klimov and L. S. Frank. The structure of a stationary renewal process. Theory of Probability and its Applications, 12 (1):120–124, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kargaev:1991:WFE

[KF91] P. P. Kargaev and S. V. Fomin.

The Walsh–Fourier expansions with random coefficients. Theory of Probability and its Applications, 35(2):258–268, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 271–281.

Kirstein:1977:CMM

[KFS77] B. M. Kirstein, P. Franken, and D. Stoyan. Comparability and monotonicity of Markov processes. Theory of Probability and its Applications, 22 (1):40–52, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kubilius:1991:VAS

[KGPS91] J. Kubilius, B. Grigelionis, V. Paulauskas, and L. Saulis. Vitautas Antanovich Statulyavichus (on the occasion of his sixtieth birthday). Theory of Probability and its Applications, 35 (3):539–541, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 541–542.

Kulinich:1981:CLB

[KH81] G. L. Kulinich and Le Thieng Huong. On the construction and limit behavior of a multiple stochastic integral of a diffusion process. Theory of Probability and its Applications, 25 (3):597–604, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1958:DPE

[Kha58] R. Z. Khas’minskii. Diffusion processes and elliptic differential equations degenerating at the boundary of the domain. Theory of Probability and its Applications, 3(4):400–419, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1959:PSE

[Kha59] R. Z. Khas’minskii. On positive solutions of the equation AU + V u = 0. Theory of Probability and its Applications, 4 (3):309–318, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1960:EPR

[Kha60] R. Z. Khas’minskii. Ergodic properties of recurrent diffusion processes and stabilization of the solution to the Cauchy problem for parabolic equations. Theory of Probability and its Applications, 5(2):179–196, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1961:LDS

[Kha61a] R. Z. Khas’minskii. On limit distributions of sums of conditionally independent random variables. Theory of Probability and its Applications, 6(1):108–113, ???? 1961. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Khazen:1961:EDP

[Kha61b] E. M. Khazen. Estimating the density of the probability distributions for random processes in systems with non-linearities of piecewise-linear type. Theory of Probability and its Applications, 6(2):214–220, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khazen:1961:EOD

[Kha61c] E. M. Khazen. Evaluation of the one-dimensional probability densities and moments of a random process in the output of an essentially non-linear system. Theory of Probability and its Applications, 6(1):117–123, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1963:PAP

[Kha63] R. Z. Khas’minskii. Principle of averaging for parabolic and elliptic differential equations and for Markov processes with small diffusion. Theory of Probability and its Applications, 8(1):1–21, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khakhubiya:1965:LRD

[Kha65] Ts. G. Khakhubiya. A lemma on random determinants and its application to the characterization of multivariate distributions. Theory of Probability and its Applications, 10(4):685–689, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khakhubiya:1966:EDD

[Kha66a] Ts. G. Khakhubiya. On the effectiveness of distinguished detween normal and uniform distributions on the basis of small samples. Theory of Probability and its Applications, 11(1):166–172, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1966:LTS

[Kha66b] R. Z. Khas’minskii. A limit theorem for the solutions of differential equations with random righthand sides. Theory of Probability and its Applications, 11 (3):390–406, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1967:NSC

[Kha67]  R. Z. Khas’minskii.  Neces-

sary and sufficient conditions for the asymptotic stability of linear stochastic systems. Theory of Probability and its Applications, 12(1):144–147, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:1968:PBP

[Kha68] B. P. Kharlamov. On properties of branching processes with an arbitrary set of particle types. Theory of Probability and its Applications, 13 (1):84–98, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:1969:NGB

[Kha69a] B. P. Kharlamov. The numbers of generations in a branching process with an arbitrary set of particle types. Theory of Probability and its Applications, 14 (3):432–449, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:1969:GNP

[Kha69b] B. P. Kharlamov. On the generation numbers of particles in a branching process with overlapping generations. Theory of Probability and its Applications, 14(1):44–50, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:1977:CSM

[Kha77] B. P. Kharlamov. On convergence of semi-Markov walks to a continuous semi-Markov process. Theory of Probability and its Applications, 21(3):482–498, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:1979:LBR

[Kha79a] R. Z. Khas’minskii. A lower bound on the risks of nonparametric estimates of densities in the uniform metric. Theory of Probability and its Applications, 23(4):794–798, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskij:1979:LBR

[Kha79b] R. Z. Khas’minskij. A lower bound on the risks of nonparametric estimates of densities in the uniform metric. Theory of Probability and its Applications, 23(??):794–798, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:1981:MPT

[Kha81] B. P. Kharlamov. A Markov property test for semi-Markov processes. Theory of Probability and its Applications, 25(3): 526–539, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khakhubia:1984:SND

[Kha84a] Ts. G. Khakhubia. On some nonparametric density estimators for a statistical classification problem. Theory of Probability and its Applications, 28(4):727–735, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharin:1984:IRS

[Kha84b] Yu. S. Kharin. The investigation of risk for statistical classifiers using minimum estimators. Theory of Probability and its Applications, 28(3):623–630, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:1984:RSM

[Kha84c] B. P. Kharlamov. Representation of a semi-Markov process as a time-change Markov process. Theory of Probability and its Applications, 28(4):688–702, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khashimov:1985:CRD

[Kha85] Sh. A. Khashimov. On the convergence rate of the distribution of a quadratic measure of deviation of nonparametric density estimates. Theory of Probability and its Applications, 29 (1):163–169, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khakhubia:1987:LTM

[Kha87] Ts. G. Khakhubia. A limit theorem for a maximum-likelihood estimate of the disorder time. Theory of Probability and its Applications, 31(1):141–144, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khashimov:1988:ADG

[Kha88] Sh. A. Khashimov. On the asymptotic distribution of the generalized U-statistics for dependent variables. Theory of Probability and its Applications, 32(2):373–375, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khashimov:1989:LTD

[Kha89a] Sh. A. Khashimov. On limit theorems for the distributions of generalized von Mises functionals. Theory of Probability and its Applications, 33(3):566–571, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 607–612.

Khashimov:1989:LDT

[Kha89b] Sh. A. Khashimov. On the limit distribution of a two-sample von Mises functional with variable kernel. Theory of Probability and its Applications, 33(1):179–182, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khashimov:1990:APF

[Kha90a] Sh. A. Khashimov. Asymptotic properties of functions of spacings. Theory of Probability and its Applications, 34(2):298–306, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 340–348.

Khasminskii:1990:SOA

[Kha90b] R. Z. Khas′minskiĭ. Some optimal algorithms of random multiple access. Theory of Probability and its Applications, 34 (3):454–463, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 505–515.

Khashimov:1993:LBG

[Kha93] Sh. A. Khashimov. Limiting behavior of generalized U-statistics of weakly dependent stationary processes. Theory of Probability and its Applications, 37 (1):148–150, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 214–216.

Khashimov:1994:CLT

[Kha94] Sh. A. Khashimov. The Central Limit Theorem for generalized U-statistics for weakly dependent vectors. Theory of Probability and its Applications, 38 (3):456–469, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 563–578.

Khafizov:1996:WMS

[Kha96] M. U. Khafizov. Weakly measurable seminorms and sufficient topologies. Theory of Probability and its Applications, 40 (2):366–371, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 453–458.

Khinchin:1956:PSC

[Khi56a] A. Ya. Khinchin. On Poisson sequences of chance events. Theory of Probability and its Applications, 1(3):291–297, ???? 1956.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khinchin:1956:SCE

[Khi56b] A. Ya. Khinchin. Sequences of chance events without aftereffects. Theory of Probability and its Applications, 1(1):1–15, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khinchin:1962:EFT

[Khi62] A. Ya. Khinchin. Erlang’s formulas in the theory of mass service. Theory of Probability and its Applications, 7(3):320–325, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khmaladze:1975:ENN

[Khm75] E. V. Khmaladze. Estimation of the necessary number of observations for discriminating simple close hypotheses. Theory of Probability and its Applications, 20(1):116–126, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 115–125.

Khmaladze:1980:UTT

[Khm80] E. V. Khmaladze. The use of ω^{2} tests for testing parametric hypotheses. Theory of Probability and its Applications, 24 (2):283–301, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khmaladze:1981:RWC

[Khm81] E. V. Khmaladze. A remark on the weak convergence of linear decomposable statistics. Theory of Probability and its Applications, 25(3):622–624, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khmaladze:1982:MAT

[Khm82] E. V. Khmaladze. Martingale approach in the theory of goodnessof-fit tests. Theory of Probability and its Applications, 26 (2):240–257, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khmaladze:1984:MLT

[Khm84] E. V. Khmaladze. Martingale limit theorems for divisible statistics. Theory of Probability and its Applications, 28 (3):530–548, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1967:SEP

[Kho67] A. S. Kholevo. Sequential estimates of parameters. Theory of Probability and its Applications, 12(1):147–152, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1969:ERC

[Kho69] A. S. Kholevo. On estimates of regression coefficients. Theory of Probability and its Applications, 14(1):79–104, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1971:APS

[Kho71a] A. S. Kholevo. Almost-periodic signal estimation. Theory of Probability and its Applications, 16(2):249–263, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1971:AEP

[Kho71b] A. S. Kholevo. On the asymptotic efficiency of pseudobest estimates. Theory of Probability and its Applications, 16(3):516–527, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1971:ANE

[Kho71c] A. S. Kholevo. On the asymptotic normality of estimates of regression coefficients. Theory of Probability and its Applications, 16(4):707–711, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1973:EFM

[Kho73a] A. S. Kholevo. On estimating functions of the mean. Theory of Probability and its Applications, 17(3):543–547, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1973:SSP

[Kho73b] A. S. Kholevo. Some statistical problems for quantum fields. Theory of Probability and its Applications, 17(2):347–351, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1974:GRC

[Kho74] A. S. Kholevo. A generalization of the Rao–Cramer inequality. Theory of Probability and its Applications, 18(2):359–362, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1979:AOH

[Kho79] A. S. Kholevo. On asymptotically optimal hypothesis testing in quantum statistics. Theory of Probability and its Applications, 23(2):411–415, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khokhlov:1983:CMD

[Kho83] Yu. S. Khokhlov. On the convergence to a multi-dimensional stable law of the distribution of a shift parameter for the composition of random motions in Euclidean space. Theory of Probability and its Applications, 27 (2):363–365, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1987:IDM

[Kho87] A. S. Kholevo. Infinitely divisible measurements in quantum probability theory. Theory of Probability and its Applications, 31 (3):493–497, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1988:RLK

[Kho88] A. S. Kholevo. A representation of Lévy-Khinchine type in quantum probability theory. Theory of Probability and its Applications, 32(1):131–136, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholevo:1994:LKF

[Kho94] A. S. Kholevo. On the Lévy– Khinchin formula in noncommutative probability theory. Theory of Probability and its Applications, 38(4):660–672, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 842–857.

Khrennikov:1986:ETS

[Khr86] A. Yu. Khrennikov. An existence theorem for a solution of a stochastic differential equation in a locally convex space. Theory of Probability and its Applications, 30(1):174–177, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khrennikov:1991:CLT

[Khr91]  A. Yu. Khrennikov.  The

Central Limit Theorem for a quasi-Gaussian distribution on an infinite-dimensional superspace. Theory of Probability and its Applications, 35(3):580– 583, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 599–602.

Khrennikov:1994:WFS

[Khr94] A. Yu. Khrennikov. The Wiener– Feynman stochastic process on a superspace. Theory of Probability and its Applications, 38 (3):533–537, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 652–656.

Khrennikov:1996:PDF

[Khr96a] A. Yu. Khrennikov. On probability distributions on the field of p-adic numbers. Theory of Probability and its Applications, 40 (1):159–162, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 189–192.

Khrennikov:1996:EMF

[Khr96b] A. Yu. Khrennikov. On the extension of the von Mises frequency approach and Kolmogorov axiomatic approach to the p-adic probability theory. Theory of Probability and its Applications, 40(2):371–376, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khrennikov:1998:ABB

[Khr98]  A. Yu. Khrennikov.  p-adic

behavior of Bernoulli probabilities. Theory of Probability and its Applications, 42(4):689– 694, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97658.

Kifer:1969:OSG

[Kif69] Yu. I. Kifer. Optimal strategy in games with an unbounded sequence of moves. Theory of Probability and its Applications, 14 (2):279–286, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kifer:1971:OSGa

[Kif71a] Yu. I. Kifer. Optimal stopped games. Theory of Probability and its Applications, 16(1):185–189, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kifer:1971:OSGb

[Kif71b] Yu. I. Kifer. Optimal stopping in games with continuous time. Theory of Probability and its Applications, 16(3):545–550, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kifer:1973:BBC

[Kif73] Yu. I. Kifer. Boundary behavior of conditional diffusion processes. Theory of Probability and its Applications, 18(1):158–161, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kifer:1975:CRC

[Kif75]  Yu. I. Kifer.  Certain re-

sults concerning small random perturbations of dynamical systems. Theory of Probability and its Applications, 19(3):487–505, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kifer:1976:BMH

[Kif76] Yu. I. Kifer. Brownian motion and harmonic functions on manifolds of negative curvature. Theory of Probability and its Applications, 21(1):81–95, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kifer:1977:ATD

[Kif77] Yu. I. Kifer. On the asymptotics of the transition density of processes with small diffusion. Theory of Probability and its Applications, 21(3):513–522, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kiryanova:1991:NEA

[Kir91] L. V. Kir′yanova. Nonuniform estimates of the accuracy of normal approximation for martingales. Theory of Probability and its Applications, 35(2):359–360, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 397–398.

Kirillov:1994:BMD

[Kir94] A. I. Kirillov. Brownian motion with drift in a Hilbert space and its application in integration theory. Theory of Probability and its Applications, 38(3):529–533, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 629–634.

Kislitsyn:1962:ALM

[Kis62] S. S. Kislitsyn. The average length of a minimum redundancy binary code for probabilities of coded symbols not differing greatly from each other. Theory of Probability and its Applications, 7(3):331–333, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kitaev:1982:EOH

[Kit82] M. Yu. Kitaev. The existence of optimal homogeneous strategies of controllable semicontinuous semi-Markov models with respect to the average cost criterion. Theory of Probability and its Applications, 26(3):614–616, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kitayev:1986:SMJ

[Kit86] M. Yu. Kitayev. Semi-Markov and jump Markov controlled models: Average cost criterion. Theory of Probability and its Applications, 30(2):272–288, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khalfina:1975:RVL

[KK75] N. M. Khalfina and L. A. Khalfin.

On a robust version of the likelihood ratio test. Theory of Probability and its Applications, 20 (1):199–202, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klimov:1976:IDP

[KK76] G. P. Klimov and A. D. Kuz’min.

Invariant decision procedures. Theory of Probability and its Applications, 20(2):302–322, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Karpelevich:1982:TPQ

[KK82] F. I. Karpelevich and A. Ya.

Kreinin. Two-phase queuing system GI/G/1 → G′/1/∞ under heavy traffic conditions. Theory of Probability and its Applications, 26(2):293–313, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1995:LFM

[KK95a] Yu. M. Kabanov and D. O.

Kramkov. Large financial markets: Asymptotic arbitrage and contiguity. Theory of Probability and its Applications, 39 (1):182–187, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 222–229.

Kabanov:1995:NAE

[KK95b] Yu. M. Kabanov and D. O. Kramkov. No-arbitrage and equivalent martingale measures: an elementary proof of the Harrison–Pliska theorem. Theory of Probability and its Applications, 39(3):523–527, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 635–640.

Kulinich:1996:LTS

[KK96] G. L. Kulinich and M. V. Kharkova. Limit theorems for stochastic differential equations without after-effect. Theory of Probability and its Applications, 40(3):469–483, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 523–541.

Kallianpur:1999:CES

[KK99] G. Kallianpur and Y. T. Kim. A curious example from statistical differential geometry. Theory of Probability and its Applications, 43(1):42–62, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97665.

Kelbert:1989:JND

[KKR89] M. Ya. Kel′bert, M. L. Kontsevich, and A. N. Rybko. On Jackson networks on denumerable graphs. Theory of Probability and its Applications, 33 (2):358–361, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 379–382.

Krylov:1995:BDO

[KKSY95] N. V. Krylov, S. E. Kuznetsov, A. V. Skorokhod, and A. A. Yushkevich. E. B. Dynkin (on the occasion of his seventieth birthday). Theory of Probability and its Applications, 39(4):654–656, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 796–798.

Koshevnik:1977:NPA

[KL77] Yu. A. Koshevnik and B. Ya.

Levit. On a non-parametric analogue of the information matrix. Theory of Probability and its Applications, 21(4):738–753, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuhn:1979:SPH

[KL79] T. Kühn and F. Liese. A short proof of the Hájek–Feldman theorem. Theory of Probability and its Applications, 23(2):429–431, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kanevskii:1983:PSI

[KL83] V. A. Kanevskii and G. Sh. Lev. Point sequences in infinitedimensional spaces and integration of functions. Theory of Probability and its Applications, 27 (2):375–379, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1993:AFD

[KL93] A. P. Korostelëv and S. L. Leonov. An action functional for a diffusion process with discontinuous drift. Theory of Probability and its Applications, 37(3): 543–550, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 570–576.

Klass:1985:IFM

[Kla85] M. J. Klass. An interpretation of the finite mean LIL norming constants. Theory of Probability and its Applications, 29(3):599–601, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1971:RIT

[Kle71] L. V. Klebanov. A remark on the independence of a tube statistic and the sample mean. Theory of Probability and its Applications, 16(4):732–733, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1974:ASM

[Kle74a] L. B. Klebanov. The admissibility of the sample mean as an estimate of the location parameter for non-quadratic risk functions. Theory of Probability and its Applications, 18(2):329–339, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1974:CZR

[Kle74b] L. B. Klebanov. On conditions for the zero regression of one linear statistic with respect to another. Theory of Probability and its Applications, 19(1):200–204, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1974:CFD

[Kle74c] L. B. Klebanov. On the characterization of a family of distributions by the property of independence of statistics. Theory of Probability and its Applications, 18(3):608–611, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1975:UES

[Kle75] L. B. Klebanov. Unbiased estimators and sufficient statistics. Theory of Probability and its Applications, 19(2):379–383, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1977:GDU

[Kle77] L. B. Klebanov. A general definition of unbiasedness. Theory of Probability and its Applications, 21(3):571–585, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1979:SPC

[Kle79] L. B. Klebanov. Some problems of characterizing distributions arising in reliability theory. Theory of Probability and its Applications, 23(4):798–801, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1980:WDQ

[Kle80] L. B. Klebanov. When does a quadratic statistic have a constant regression on a sample mean? Theory of Probability and its Applications, 24(3):649–651, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1981:SRC

[Kle81] L. B. Klebanov. Some results connected with a characterization of the exponential distribution. Theory of Probability and its Applications, 25(3):617–622, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kleptsyna:1985:SSSb

[Kle85] M. L. Kleptsyna. On strong solutions of stochastic equations with degenerate coefficients. Theory of Probability and its Applications, 29(2):403–407, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kleptsyna:1986:CEU

[Kle86] M. I. Kleptsyna. Comparison, existence and uniqueness theorems for stochastic differential equations. Theory of Probability and its Applications, 30(1):164–169, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klesov:1987:LIL

[Kle87] O. I. Klesov. The law of the iterated logarithm for weighted sums of independent identically distributed random variables. Theory of Probability and its Applications, 31(2):337–342, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klesov:1996:ASC

[Kle96] O. I. Klesov. Almost sure convergence of multiple series of independent random variables. Theory of Probability and its Applications, 40(1):52–65, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 68–83.

Kleptsyna:1997:APD

[Kle97] M. L. Kleptsyna. The averaging principle and diffusion approximation for Volterra equations. Theory of Probability and its Applications, 41(2):359–367, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 429–438.

Klimov:1957:KEM

[Kli57] V. N. Klimov. The kinetic equation for mixtures. Theory of Probability and its Applications, 2(2):253–260, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klimov:1962:ACS

[Kli62] V. N. Klimov. On asymptotically continuous stochastic processes. Theory of Probability and its Applications, 7(1):58–80, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klimov:1975:TSS

[Kli75] G. P. Klimov. Time-sharing service systems. I. Theory of Probability and its Applications, 19 (3):532–551, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 19, (1974), pp. 558–576.

Klimov:1977:ETR

[Kli77] G. P. Klimov. An ergodic theorem for regenerating processes. Theory of Probability and its Applications, 21(2):392–395, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 21, (1976), pp. 402– 405.

Klimov:1979:LTH

[Kli79a] G. P. Klimov. Limit theorems for homogeneous Markov processes. Theory of Probability and its Applications, 23(4):700–711, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klimov:1979:TSS

[Kli79b] G. P. Klimov. Time-sharing service systems. II. Theory of Probability and its Applications, 23 (2):314–321, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kloss:1959:PDB

[Klo59] B. M. Kloss. Probability distributions on bicompact topological groups. Theory of Probability and its Applications, 4 (3):237–270, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kloss:1961:LDB

[Klo61] B. M. Kloss. Limiting distributions on bicompact Abelian groups. Theory of Probability and its Applications, 6(4):361–389, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kloss:1962:SDC

[Klo62] B. M. Kloss. Stable distributions on a class of locally compact groups. Theory of Probability and its Applications, 7 (3):237–257, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kloss:1964:TGC

[Klo64] B. M. Kloss. Topology in a group and convergence of distributions. Theory of Probability and its Applications, 9(1): 111–114, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krichagina:1989:DAS

[KLP89] E. V. Krichagina, R. Sh. Liptser, and A. A. Pukhal’skii. Diffusion approximation of systems with arrival depending on queue, and with arbitrary servicing. Theory of Probability and its Applications, 33(1):114–124, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:1984:WSC

[KLS84] Yu. M. Kabanov, R. Sh. Liptser, and A. N. Shiryaev. Weak and strong convergence of the distributions of counting processes. Theory of Probability and its Applications, 28(2):303–336, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kallianpur:1965:MRT

[KM65] G. Kallianpur and V. Mandrekar. Multiplicity and representation theory of purely non-deterministic stochastic processes. Theory of Probability and its Applications, 10(4):553–581, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1980:EOR

[KM80] N. V. Krylov and T. P. Miroshnichenko. On extending the optimality region of Wald’s test. Theory of Probability and its Applications, 25(1):165–169, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krivenko:1989:PEE

[KM89] M. P. Krivenko and I. B. Matskevich. Properties of elements in estimating monotone trend. Theory of Probability and its Applications, 33(2):316–329, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 336–348.

Kramkov:1995:IO

[KM95] D. O. Kramkov and E. Mordecki.

An integral option. Theory of Probability and its Applications, 39(1):162–172, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 201–211.

Klebanov:1985:PZA

[KMM85] L. B. Klebanov, G. M. Maniya, and I. A. Melamed. A problem of Zolotarev and analogs of infinitely divisible and stable distributions in a scheme for summing a random number of random variables. Theory of Probability and its Applications, 29 (4):791–794, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krivyakov:1978:DSM

[KMT78] E. N. Krivyakov, G. V. Martynov, and Yu. N. Tyurin. On the distribution of the ω^{2} statistics in the multi-dimensional case. Theory of Probability and its Applications, 22(2):406–410, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kim:1975:APL

[KN75] L. V. Kim and A. V. Nagaev. On the asymmetrical problem of large deviations. Theory of Probability and its Applications, 20 (1):57–68, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1990:CDP

[KN90] L. B. Klebanov and M. V. Neupokoeva. A characterization of a distribution by a property of the mean values of order statistics. Theory of Probability and its Applications, 34(4):715–720, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 780–785.

Karpenko:1994:LTT

[KN94] A. V. Karpenko and S. V. Nagaev. Limit theorems for the total number of descendants for the Galton–Watson branching process. Theory of Probability and its Applications, 38(3):433–455, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 503–528.

Knight:1966:PCN

[Kni66] F. Knight. A pair of conditional normal convergence theorems. Theory of Probability and its Applications, 11(2):191–210, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kesten:1966:GWP

[KNS66] H. Kesten, P. Ney, and F. Spitzer.

The Galton–Watson process with mean one and finite variance. Theory of Probability and its Applications, 11(4):513–540, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Koralov:1994:ABT

[KNS94] L. B. Koralov, S. K. Nechaev, and Ya. G. Sinaĭ. Asymptotic behavior of a two-dimensional random walk with topological constraints. Theory of Probability and its Applications, 38(2): 296–306, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 331–344.

Kochubei:1990:DIG

[Koc90] A. N. Kochubeĭ. On diffusions with invariant generating operators. Theory of Probability and its Applications, 34(4):597–603, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 656–663.

Kochetov:1998:GOS

[Koc98] V. V. Mazalov E. A. Kochetov. A game with optimal stopping of random walks. Theory of Probability and its Applications, 42(4): 697–701, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97655.

Kogan:1965:ASE

[Kog65] Ya. A. Kogan. Application of stochastic equations to the study of the second boundary value problem for parabolic differential equations with a small parameter. Theory of Probability and its Applications, 10(2):279–286, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kogan:1966:AAS

[Kog66] Ya. A. Kogan. Addendum: Application of stochastic equations to the study of the second boundary value problem for parabolic differential equations with a small parameter. Theory of Probability and its Applications, 11(2):321, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kogan:1968:OCR

[Kog68] Ya. A. Kogan. On optimal control by a random time substitution in a continuous Markov process. Theory of Probability and its Applications, 13(2):343–345, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kogan:1969:OCN

[Kog69] Ya. A. Kogan. On optimal control of a non-terminating diffusion process with reflection. Theory of Probability and its Applications, 14(3):496–502, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kogan:1970:AOC

[Kog70] Ya. A. Kogan. Addendum: On optimal control of a non-stopped diffusion process with reflection. Theory of Probability and its Applications, 15(1):161, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kokotuskin:1975:GPK

[Kok75] V. A. Kokotuskin. A generalization of the Palm–Khinchin theorem. Theory of Probability and its Applications, 19(3):594–597, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolmogorov:1956:SC

[Kol56a] A. N. Kolmogorov. On Skorokhod convergence. Theory of Probability and its Applications, 1(2):215–222, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolmogorov:1956:TUL

[Kol56b] A. N. Kolmogorov. Two uniform limit theorems for sums of independent random variables. Theory of Probability and its Applications, 1(4):384–394, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolmogorov:1962:BVG

[Kol62] A. N. Kolmogorov. B. V. Gnedenko’s work in probability theory (on the occasion of his fiftieth birthday). Theory of Probability and its Applications, 7 (3):313–319, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1966:SCL

[Kol66] V. F. Kolchin. The speed of convergence to limit distributions in the classical ball problem. Theory of Probability and its Applications, 11(1):128–140, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1967:ULL

[Kol67] V. F. Kolchin. Uniform local limit theorem in the classical ball problem for a case with varying lattice. Theory of Probability and its Applications, 12 (1):57–67, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1969:LBE

[Kol69a] V. F. Kolchin. On the limiting behavior of extreme order statistics in a polynomial scheme. Theory of Probability and its Applications, 14(3): 458–469, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolmanovskii:1969:EPS

[Kol69b] V. B. Kolmanovskii. Ergodic properties of stationary solutions of stochastic differential equations. Theory of Probability and its Applications, 14(1):141–146, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1971:PAP

[Kol71] V. F. Kolchin. A problem of the allocation of particles in cells and cycles of random permutations. Theory of Probability and its Applications, 16 (1):74–90, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1976:PAP

[Kol76] V. F. Kolchin. A problem of the allocation of particles in cells and random mappings. Theory of Probability and its Applications, 21(1):48–63, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchin:1987:BRG

[Kol87] V. F. Kolchin. On the behavior of a random graph near a critical point. Theory of Probability and its Applications, 31 (3):439–451, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolomiets:1988:ABP

[Kol88] E. I. Kolomiets. On the asymptotic behavior of probabilities of type I and type II errors in the Neyman–Pearson test. (case of total asymptotically distinguishable hypotheses). Theory of Probability and its Applications, 32(3):458–476, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolchinskii:1992:EOO

[Kol92] V. I. Kolchinskiĭ. Entropic order of operators in Banach spaces and the Central Limit Theorem. Theory of Probability and its Applications, 36(2):303–315, May 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 313–325.

Kolmogorov:1994:PPT

[Kol94] A. N. Kolmogorov. Problems of probability theory. Theory of Probability and its Applications, 38(2):177–178, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolesnikov:1999:TPS

[Kol99] A. V. Kolesnikov. On topological properties of the Skorokhod space. Theory of Probability and its Applications, 43(4): 640–644, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97721.

Kondo:1970:CRM

[Kon70] Ryoji Kondo. On a construction of recurrent Markov chains. II. Theory of Probability and its Applications, 15(3):499–507, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1973:NPE

[Kon73] V. D. Konakov. Non-parametric estimation of density functions. Theory of Probability and its Applications, 17(2):361–362, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1974:ANM

[Kon74a] V. D. Konakov. On the asymptotic normality of the mode of multidimensional distributions. Theory of Probability and its Applications, 18(4):794–799, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 18, (1973), pp. 836–

Konakov:1974:SHR

[Kon74b] V. D. Konakov. Some Hajek– Rènyi type inequalities. Theory of Probability and its Applications, 18(2):393–394, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1978:GMD

[Kon78] V. D. Konakov. On a global measure of deviation for an estimate of the regression line. Theory of Probability and its Applications, 22(4):858–868, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1979:CAE

[Kon79]  V. D. Konakov.  Complete

asymptotic expansions for the maximum deviation of the empirical density function. Theory of Probability and its Applications, 23(3):475–489, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konovalov:1982:LLT

[Kon82] A. G. Konovalov. A local limit theorem on convergence to a diffusion process. Theory of Probability and its Applications, 26 (3):581–587, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1985:TDE

[Kon85] V. D. Konakov. A theorem on deviations of the empirical measure and its applications. Theory of Probability and its Applications, 29(1):158–163, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kopytsev:1996:DNS

[Kop96] V. A. Kopytsev. On the distribution of the number of solutions of random systems of equations which are known to be consistent. Theory of Probability and its Applications, 40(2):376–383, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Koronkevich:1957:SRE

[Kor57] A. I. Koronkevich. Some remarks on evaluating the accuracy of linear extrapolation and filtration. Theory of Probability and its Applications, 2(1): 112–117, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kordonskii:1958:DND

[Kor58] H. B. Kordonskii. On a distribution of the number of defective items in manufactured lots. Theory of Probability and its Applications, 3(3):329–333, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1959:AAD

[Kor59] V. S. Korolyuk. Asymptotic analysis of the distribution of the maximum deviation in the Bernoulli scheme. Theory of Probability and its Applications, 4(4):339–366, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kordonskii:1961:DND

[Kor61a] Kh. B. Kordonskii. The distribution of the number of defective items in manufactured lots. Theory of Probability and its Applications, 6(3):314–320, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1961:ABS

[Kor61b] V. S. Korolyuk. On the asymptotic behavior of some functionals in lattice random walk. Theory of Probability and its Applications, 6(3):308–314, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1962:AAD

[Kor62] V. S. Korolyuk. Asymptotic analysis of distributions of maximum deviations in a lattice random walk. Theory of Probability and its Applications, 7 (4):383–401, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korzhik:1966:MOE

[Kor66] I. V. Korzhik. A method of obtaining estimates for the mathematical expectation of some function of the components of an n-dimensional random vector. Theory of Probability and its Applications, 11(3):476–480, ???? 1966. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Korzhik:1968:AMO

[Kor68] V. I. Korzhik. Addendum: a method for obtaining estimates for the mathematical expectation of some function of the components of an n-dimensional random vector. Theory of Probability and its Applications, 13(4):711, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1973:PRS

[Kor73] A. P. Korostelev. A probabilistic representation of the solution of the directional derivative problem. Theory of Probability and its Applications, 18(1):169–172, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1974:BPC

[Kor74] V. S. Korolyuk. Boundary problems for a compound Poisson process. Theory of Probability and its Applications, 19(1):1–13, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1976:RPC

[Kor76] V. S. Korolyuk. On ruin problems for a compound Poisson process. Theory of Probability and its Applications, 20(2):374–376, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1978:CCC

[Kor78a] A. P. Korostelev. A convergence criterion of continuous stochastic approximation procedures. Theory of Probability and its Applications, 22(3):584–591, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1978:PRD

[Kor78b] A. P. Korostelev. On a probabilistic representation of a discontinuous solution of a parabolic equation. Theory of Probability and its Applications, 22(2): 413–418, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1980:DPD

[Kor80] A. P. Korostelev. Damping perturbations of dynamical systems and convergence conditions for recursive stochastic procedures. Theory of Probability and its Applications, 24(2):302–321, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1984:UFS

[Kor84] A. P. Korostelev. On upper functions for stochastic approximation procedures. Theory of Probability and its Applications, 28 (4):806–811, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kordecki:1986:SRF

[Kor86a] W. Kordecki. Some recurrence formulas for probability of correctedness of random graphs. Theory of Probability and its Applications, 30(3):630–632, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:1986:CMR

[Kor86b] V. Yu. Korolev. Convergence of moments of random sums of independent random variables. Theory of Probability and its Applications, 30(2):386–390, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1986:MPM

[Kor86c] A. P. Korostelev. Minimax properties of the maximum likelihood estimate for a drifting parameter. Theory of Probability and its Applications, 30(4):720–735, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1988:MFL

[Kor88a] A. P. Korostelev. Minimax form of the law of the iterated logarithm for statistical estimates. Theory of Probability and its Applications, 32(3):449–457, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korostelev:1988:MED

[Kor88b] A. P. Korostelev. On minimax estimation of a discontinuous signal. Theory of Probability and its Applications, 32(4):727–730, ???? 1988. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:1989:NEA

[Kor89a] V. Yu. Korolev. Nonuniform estimates of the accuracy of normal approximation of the distributions of random sums of independent random variables. Theory of Probability and its Applications, 33(4):755–759, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:1989:ANA

[Kor89b] V. Yu. Korolev. On the accuracy of normal approximation for the distributions of sums of a random number of independent random variables. Theory of Probability and its Applications, 33 (3):540–544, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 577–581.

Korostelev:1989:RNE

[Kor89c] A. P. Korostelëv. Recursive nonparametric estimation by observing a trajectory of a process with independent increments. Theory of Probability and its Applications, 33(3):544–547, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 581–585.

Korolev:1990:ADR

[Kor90a] V. Yu. Korolev. Approximation of distributions of random sums of independent random variables by mixtures of normal laws. Theory of Probability and its Applications, 34(3):523–531, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 581–588.

Korolyuk:1990:BLA

[Kor90b] V. S. Korolyuk. The boundary layer in the asymptotic analysis of random walks. Theory of Probability and its Applications, 34 (1):179–186, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(1), (1989), pp. 208–215.

Korostelev:1991:EAN

[Kor91] A. P. Korostelëv. On estimation accuracy for nonsmooth functionals of regression. Theory of Probability and its Applications, 35(4):784–787, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 768–770.

Korostelev:1992:MTD

[Kor92] A. P. Korostelëv. Minimax twodimensional image reconstruction. Theory of Probability and its Applications, 36(1):158–165, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 153–159.

Korolev:1993:LDR

[Kor93] V. Yu. Korolev. On limit distributions of randomly indexed random sequences. Theory of Probability and its Applications, 37(3):535–542, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 564–570.

Korostelev:1994:AMR

[Kor94a] A. P. Korostelëv. An asymptotically minimax regression estimator in the uniform norm up to exact constant. Theory of Probability and its Applications, 38 (4):737–743, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 875–882.

Korshunov:1994:TPR

[Kor94b] D. A. Korshunov. Transient phenomena for real-valued Markov chains. Theory of Probability and its Applications, 38(1):149–152, March 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 184–187.

Korolev:1995:CRS

[Kor95] V. Yu. Korolev. Convergence of random sequences with independent random indices. I. Theory of Probability and its Applications, 39(2):282–297, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 313–333.

Korolev:1996:CRS

[Kor96] V. Yu. Korolev. Convergence of random sequences with independent random indices. II. Theory of Probability and its Applications, 40(4):770–772, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 907–910.

Korolev:1998:CDR

[Kor98] V. Yu. Korolev. On the convergence of distributions of random sums of independent random variables to stable laws. Theory of Probability and its Applications, 42(4):695–696, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97654.

Korolev:1999:CDC

[Kor99] V. Yu. Korolev. On convergence of distributions of compound Cox processes to stable laws. Theory of Probability and its Applications, 43(4):

644–650, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97722.

Koshevnik:1985:SLP

[Kos85] Yu. A. Koshevnik. On some limit properties of nonparametric estimates of a distribution function. Theory of Probability and its Applications, 29(4):807–813, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Koshkin:1988:MRB

[Kos88] G. M. Koshkin. On a method of removing bias of estimates. Theory of Probability and its Applications, 32(1):136–138, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Koshkin:1989:INN

[Kos89] G. M. Koshkin. Improved nonnegative kernel estimate of a density. Theory of Probability and its Applications, 33(4):759–764, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 817–822.

Koshkin:1999:NIA

[Kos99] V. A. Vasiliev G. M. Koshkin. Nonparametric identification of autoregressions. Theory of Probability and its Applications, 43 (3):507–517, September 1999.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97706.

Kotelnikov:1964:NCP

[Kot64] V. P. Kotel’nikov. A nomogram connecting the parameters of Weibull’s distribution with probabilities. Theory of Probability and its Applications, 9 (4):670–674, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalenko:1959:COD

[Kov59] I. N. Kovalenko. On a class of optimal decision functions for a binomial family of distributions. Theory of Probability and its Applications, 4(1):95–99, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalenko:1960:LDF

[Kov60] I. N. Kovalenko. On the limit distribution of the first jump. Theory of Probability and its Applications, 5(4):425–428, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalenko:1961:SQP

[Kov61] I. N. Kovalenko. Some queuing problems with restrictions. Theory of Probability and its Applications, 6(2):204–208, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalenko:1967:LDN

[Kov67] I. N. Kovalenko. On the limit distribution of the number of solutions of a random system of linear equations in the class of Boolean functions. Theory of Probability and its Applications, 12(1):47–56, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalenko:1973:DLR

[Kov73] I. N. Kovalenko. Distribution of the linear rank of a random matrix. Theory of Probability and its Applications, 17(2):342–346, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 17, (1972), pp. 354– 359.

Kovyazin:1986:LBC

[Kov86] S. A. Kovyazin. On the limit behavior of a class of empirical means of a random set. Theory of Probability and its Applications, 30(4):814–820, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1971:DRN

[Koz71a] G. A. Kozlov. On the distribution of random numbers produced by sequential physical generators. Theory of Probability and its Applications, 16(2): 369–378, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1971:LTC

[Koz71b] M. V. Kozlov. A limit theorem for a characteristic of a random Boolean matrix. Theory of Probability and its Applications, 16(1):91–101, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1973:EEM

[Koz73] G. A. Kozlov. Estimation of the error of the method of statistical tests (Monte-Carlo) due to imperfections in the distribution of random numbers. Theory of Probability and its Applications, 17(3):493–509, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1974:RWO

[Koz74a] M. V. Kozlov. Random walk in a one-dimensional random medium. Theory of Probability and its Applications, 18(2): 387–388, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1974:ODP

[Koz74b] V. P. Kozlov. On an optimum design problem for a statistical experiment. Theory of Probability and its Applications, 19 (1):218–221, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1977:ABP

[Koz77a] M. V. Kozlov. On the asymptotic behavior of the probability of non-extinction for critical branching processes in a random environment. Theory of Probability and its Applications, 21 (4):791–804, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1977:GDP

[Koz77b] O. K. Kozlov. Gibbsian description of point random fields. Theory of Probability and its Applications, 21(2):339–356, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1978:LTQ

[Koz78] V. V. Kozlov. A limit theorem for a queueing system. Theory of Probability and its Applications, 23(1):182–187, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:1981:ECI

[Koz81] G. A. Kozlov. The error in calculating an integral by the Monte– Carlo method using a random code generator with independent digits. Theory of Probability and its Applications, 25(2):401–408, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozin:1984:LES

[Koz84] I. V. Kozin. Lebesgue expansion for spherically invariant measures. Theory of Probability and its Applications, 28(3):606–610, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:1967:IEF

[KP67] A. M. Kagan and V. P. Palamodov. Incomplete exponential families and minimum-variance unbiased estimators, I. Theory of Probability and its Applications, 12(1):36–46, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kushnir:1971:AOT

[KP71] A. F. Kushnir and A. I. Pinskii. Asymptotically optimal tests for testing hypothesis for interdependent samples of observations. Theory of Probability and its Applications, 16(2):281–292, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1983:CRM

[KP83] V. D. Konakov and V. I. Piterbarg. Convergence rate of maximum deviation distributions of Gaussian processes and empirical densities. I. Theory of Probability and its Applications, 27 (4):759–779, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konakov:1984:RCD

[KP84] V. D. Konakov and V. I. Piterbarg. Rate of convergence of distributions of maximal deviations of Gaussian processes and empirical density functions. II. Theory of Probability and its Applications, 28(1):172–178, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konev:1997:PPE

[KP97] V. V. Konev and S. M. Pergamenshchikov. The prescribed precision estimators of the autoregression parameter using the generalized least square method. Theory of Probability and its Applications, 41(4):678–694, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97569.

Kozlov:1993:WF

[KPĬ93] S. M. Kozlov, J. B. Pitman, and M. Ĭor. Wiener football. Theory of Probability and its Applications, 37(3):550–553, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 562–564.

Kolmogorov:1960:SMC

[KR60] A. N. Kolmogorov and Yu. A. Rozanov. On strong mixing conditions for stationary Gaussian processes. Theory of Probability and its Applications, 5 (2):204–208, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:1967:ESP

[KR67] A. M. Kagan and A. L. Rukhin. On the estimation of a scale parameter. Theory of Probability and its Applications, 12 (4):672–678, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1975:FDL

[KR75] L. B. Klebanov and A. L.

Rukhin. On families of distributions with location parameter admitting a sufficient statistic of rank not greater than two. Theory of Probability and its Applications, 19(3):577–584, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kordonskii:1976:UEM

[KR76] Kh. B. Kordonskii and P. Ya.

Rozenblit. Unbiased estimation of moment polynomials. Theory of Probability and its Applications, 21(1):199–204, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kiryanova:1987:LCD

[KR87] L. V. Kir’yanova and V. I. Rotar’. A lemma on “Corrected” distributions. Theory of Probability and its Applications, 31 (4):708–712, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krishnaiah:1990:NWC

[KR90] Y. S. R. Krishnaiah and M. S.

Rao. A note on weak convergence of empirical processes based on mixing random variables subject to perturbation and scale factors. Theory of Probability and its Applications, 34(2):370–375, ???? 1990. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 417–421.

Kiryanova:1992:ERC

[KR92] L. V. Kir′yanova and V. I. Rotar′. Estimates for the rate of convergence in the Central Limit Theorem for martingales. Theory of Probability and its Applications, 36(2):289–302, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 301–312.

Krasulina:1962:NSS

[Kra62] T. P. Krasulina. A note on some stochastic approximation processes. Theory of Probability and its Applications, 7(1): 108–113, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krasulina:1967:RMP

[Kra67] T. P. Krasulina. On the Robbinsmonro procedure for the case of several roots. Theory of Probability and its Applications, 12 (2):333–337, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krasulina:1969:SAP

[Kra69] T. P. Krasulina. On stochastic approximation processes with infinite variance. Theory of Probability and its Applications, 14 (3):522–526, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krasulina:1971:SAR

[Kra71] T. P. Krasulina. On stochastic approximation for random processes with continuous time. Theory of Probability and its Applications, 16(4):674–682, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krasnitskii:1974:CEP

[Kra74] S. M. Krasnitskii. On conditions of equivalence and perpendicularity of measures corresponding to homogeneous Gaussian fields. Theory of Probability and its Applications, 18 (3):588–592, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krasnitskii:1979:SEE

[Kra79] S. M. Krasnitskii. Some examples of equivalent and orthogonal Gaussian distributions. Theory of Probability and its Applications, 24(1):161–165, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krasnitskii:1980:ASE

[Kra80]  S. M. Krasnitskii.  Adden-

dum: Some examples of equivalent and orthogonal Gaussian distributions. Theory of Probability and its Applications, 25(1):214, ???? 1980. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Krasnitskii:1990:CED

[Kra90] S. M. Krasnit·skiĭ. Conditions for the equivalence of distributions of homogeneous Gaussian fields. Theory of Probability and its Applications, 34(4):720–725, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 786–791.

Kramkov:1994:CSS

[Kra94] D. O. Kramkov. On comparison of some statistical models of “Signal + noise” type. Theory of Probability and its Applications, 38(3):537–540, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 634–638.

Krickeberg:1964:CCE

[Kri64] K. Krickeberg. Convergence of conditional expectation operators. Theory of Probability and its Applications, 9(4):538–549, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruopis:1969:ETI

[Kru69] Yu. I. Kruopis. On estimates of transition intensities with migration. Theory of Probability and its Applications, 14(2):219–228, ???? 1969. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1970:NID

[Kru70] V. M. Kruglov. A note on infinitely divisible distributions. Theory of Probability and its Applications, 15(2):319–324, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1971:ANI

[Kru71a] V. M. Kruglov. Addendum: a note on infinitely divisible distributions. Theory of Probability and its Applications, 16(2):397, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1971:CDS

[Kru71b] V. M. Kruglov. Convergence of distributions of sums of independent random variables with values in Hilbert space. Theory of Probability and its Applications, 16(2):350–351, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1973:LTS

[Kru73a] V. M. Kruglov. Limit theorems for sums of independent random variables with values in Hilbert space. Theory of Probability and its Applications, 17(2):199–217, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1973:ECS

[Kru73b] V. M. Kruglov. On the extension of the class of stable distributions. Theory of Probability and its Applications, 17(4):685–694, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1974:CNC

[Kru74] V. N. Kruglov. Convergence of numerical characteristics of sums of independent random variables with values in a Hilbert space. Theory of Probability and its Applications, 18(4):694–712, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1975:BSI

[Kru75] V. M. Kruglov. The behavior of sums of independent random variables. Theory of Probability and its Applications, 19 (2):374–379, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1976:GLT

[Kru76] V. M. Kruglov. Global limit theorems. Theory of Probability and its Applications, 21(1):125–129, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1980:SOS

[Kru80] V. M. Kruglov. The sum and the order statistics of independent random variables. Theory of Probability and its Applications, 24(4):712–728, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1982:CLT

[Kru82] V. M. Kruglov. Central Limit Theorem and law of large numbers in the mean. Theory of Probability and its Applications, 26 (4):813–815, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1986:WCR

[Kru86] V. M. Kruglov. Weak convergence of random polygonal lines to a Wiener process. Theory of Probability and its Applications, 30(2):225–234, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1988:PDL

[Kru88] V. M. Kruglov. Parameterdependent law of the iterated logarithm. Theory of Probability and its Applications, 32 (1):68–76, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:1989:CMR

[Kru89] V. M. Kruglov. The convergence of moments of random sums. Theory of Probability and its Applications, 33(2):339–342, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 360–363.

Kruglov:1992:CDR

[Kru92] V. M. Kruglov. Convergence of the distributions of random sums to the normal and Poisson laws. Theory of Probability and its Applications, 36(2):378–380, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 361–363.

Kruglov:1993:CCF

[Kru93] V. M. Kruglov. Criteria for concentration functions to belong to the domains of attraction of stable laws. Theory of Probability and its Applications, 37(3): 424–432, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 459–467.

Kruglov:1998:IP

[Kru98] V. M. Kruglov. To the invariance principle. Theory of Probability and its Applications, 42 (2):225–243, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97609.

Kruglov:1999:WCR

[Kru99] V. M. Kruglov. Weak compactness of random sums of independent random variables. Theory of Probability and its Applications, 43(2):203–220, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97683.

Krylov:1965:COS

[Kry65] N. V. Krylov. The construction of an optimal strategy for a finite controlled chain. Theory of Probability and its Applications, 10(1):45–54, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1966:QDP

[Kry66a] N. V. Krylov. On quasi diffusion processes. Theory of Probability and its Applications, 11 (3):373–389, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1966:RBP

[Kry66b] N. V. Krylov. On regular boundary points for Markov processes. Theory of Probability and its Applications, 11(4):609–614, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kryglov:1968:LLT

[Kry68] V. M. Kryglov. A local limit theorem for non-identically distributed random variables. Theory of Probability and its Applications, 13(2):332–334, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1969:ISI

[Kry69] N. V. Krylov. On Itô’s stochastic integral equations. Theory of Probability and its Applications, 14(2):330–336, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1971:ITS

[Kry71] N. V. Krylov. An inequality in the theory of stochastic integrals. Theory of Probability and its Applications, 16(3):438–448, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1972:CSS

[Kry72] N. V. Krylov. Control of a solution of a stochastic integral equation. Theory of Probability and its Applications, 17(1):114–130, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1973:AIS

[Kry73a] N. V. Krylov. Addendum: On Itô’s stochastic integral equations. Theory of Probability and its Applications, 17(2):373–374, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1973:RCP

[Kry73b] N. V. Krylov. The regularity of conditional probabilities for stochastic processes. Theory of Probability and its Applications, 18(1):150–153, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1973:SET

[Kry73c] N. V. Krylov. Some estimates in the theory of stochastic integrals. Theory of Probability and its Applications, 18 (1):54–63, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1980:EAF

[Kry80] N. V. Krylov. On the equivalence of σ-algebras in the filtering problem of diffusion processes. Theory of Probability and its Applications, 24(4):772–781, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1985:EPS

[Kry85] N. V. Krylov. Extremal properties of solutions of stochastic equations. Theory of Probability and its Applications, 29 (2):205–217, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1987:ACD

[Kry87a] N. V. Krylov. An approach to controlled diffusion processes. Theory of Probability and its Applications, 31(4):604–626, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1987:SAM

[Kry87b] N. V. Krylov. The sufficiency of adjoined Markov strategies for controlled diffusion processes. Theory of Probability and its Applications, 31(2):304–309, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1991:SPE

[Kry91] N. V. Krylov. A simple proof of the existence of a solution of Itô’s equation with monotone coefficients. Theory of Probability and its Applications, 35(3): 583–587, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 576–580.

Klepikov:1957:NLC

[KS57] N. P. Klepikov and S. N. Sokolov. Non-linear confluence analysis. Theory of Probability and its Applications, 2(4):465–468, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolmogorov:1960:WBT

[KS60] A. N. Kolmogorov and O. V. Sarmanov. The work of S. N. Bernshtein on the theory of probability. Theory of Probability and its Applications, 5 (2):197–203, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kemeny:1961:FCT

[KS61] John G. Kemeny and J. Laurie Snell. Finite continuous time Markov chains. Theory of Probability and its Applications, 6 (1):101–105, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kandelaki:1964:CLT

[KS64] N. P. Kandelaki and V. V.

Sozanov. On a Central Limit Theorem for random elements with values in Hilbert space. Theory of Probability and its Applications, 9(1):38–46, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:1967:CNL

[KS67] A. M. Kagan and O. V. Shalaevskii. A characterization of the normal law by the property of partial sufficiency. Theory of Probability and its Applications, 12(3):512–514, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kallianpur:1969:SDE

[KS69] G. Kallianpur and C. Striebel.

Stochastic differential equations occurring in the estimation of continuous parameter stochastic processes. Theory of Probability and its Applications, 14(4): 567–594, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kaplan:1980:TIP

[KS80] E. I. Kaplan and D. S. Sil’vestrov.

Theorems of the invariance principle type for recurrent semi-Markov processes with arbitrary phase space. Theory of Probability and its Applications, 24 (3):536–547, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1984:ETA

[KS84] D. V. Korolyuk and D. S. Sil’vestrov. Entry times into asymptotically receding domains for ergodic Markov chains. Theory of Probability and its Applications, 28(2):432–442, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:1985:ETA

[KS85a] D. V. Korolyuk and D. S. Sil’vestrov. Entry times into asymptotically receding regions for processes with semi-Markov switchings. Theory of Probability and its Applications, 29 (3):558–563, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 29(3), (1984), pp. 539–544.

Kuritsyn:1985:SD

[KS85b] Yu. G. Kuritsyn and A. V. Shestakov. On α-symmetric distributions. Theory of Probability and its Applications, 29(4):804–806, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konig:1986:LTS

[KS86] D. König and V. Schmidt. Limit theorems for single-server feedback queues controlled by a general class of marked point processes. Theory of Probability and its Applications, 30(4):712–719, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klyachko:1987:CCF

[KS87] A. A. Klyachko and Yu. V.

Solodyannikov. Calculation of characteristic functions of some functionals of a Wiener process and a Brownian bridge. Theory of Probability and its Applications, 31(3):500–504, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kostava:1988:SRT

[KS88] B. A. Kostava and V. M.

Shurenkov. Some renewal type theorems. Theory of Probability and its Applications, 32 (1):98–106, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kubacki:1989:CMR

[KS89] K. S. Kubacki and D. Szynal.

The convergence of moments in a random limit theorem of H. Robbins’ type. Theory of Probability and its Applications, 33 (1):75–85, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 83–93.

Kovalenko:1991:SSG

[KS91] V. F. Kovalenko and Yu. A. Semenov. C_{0}-semigroups in L^{p}(R^{d}) and Ĉ(R^{d}) spaces generated by the differential expression Δ + b · ∇. Theory of Probability and its Applications, 35(3): 443–453, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 449–458.

Kramkov:1995:RPR

[KS95] D. O. Kramkov and A. N. Shiryaev. On the rational pricing of the “Russian Option” for the symmetrical binomial model of a (B, S)-market. Theory of Probability and its Applications, 39(1):153–162, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 191–200.

Krivenko:1997:DNS

[KS97] M. P. Krivenko and K. V. Streltsov. On the distribution of the number of steps in an estimate of monotone trend. Theory of Probability and its Applications, 41(1):25– 34, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97501.

Korolyuk:1977:MPB

[KSS77] V. S. Korolyuk, B. I. Suprun, and V. M. Shurenkov. Method of potential in boundary problems for processes with independent increases and jumps of the same sign. Theory of Probability and its Applications, 21(2):243–249, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1960:QMF

[KST60] P. I. Kuznetsov, R. L. Stratonovich, and V. I. Tikhonov. Quasimoment functions in the theory of random processes. Theory of Probability and its Applications, 5(1):80–97, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1961:SPC

[KST61] P. I. Kuznetsov, P. L. Stratonovich, and V. I. Tikhonov. Some problems with conditional probability and quasimoment functions. Theory of Probability and its Applications, 6(4):422–427, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Karpelevicii:1959:LTC

[KT59] F. I. Karpelevicii and V. N.

Tutubalin. Limit theorems for the compositions of distributions in the Lobachevsky plane and space. Theory of Probability and its Applications, 4(4):399–402, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klyatskin:1969:PDA

[KT69] V. I. Klyatskin and V. I.

Tatarskii. Probability distribution of the area bounded by a Gaussian random contour. Theory of Probability and its Applications, 14(2):344–349, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kvartskheliya:1977:CLT

[KT77] V. V. Kvartskheliya and Nguyen Xuy Tien. The Central Limit Theorem and the strong law of large numbers in the spaces l_{p}(x), 1 ≤ p < +∞. Theory of Probability and its Applications, 21 (4):780–790, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolmogorov:1988:AR

[KU88a] A. N. Kolmogorov and V. A. Uspenskii. Algorithms and randomness. Theory of Probability and its Applications, 32(3): 389–412, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
tvp/resource/1/tprbau/v32/
i3/p389_s1.

Kruglov:1988:MPD

[KU88b] V. M. Kruglov and A. M. Ulanovskii. Mixtures of probability distributions uniquely determined by their behavior on a half-line. Theory of Probability and its Applications, 32 (4):608–616, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kubacki:1996:CMM

[Kub96] K. S. Kubacki. On the convergence of moments in a martingale Central Limit Theorem. Theory of Probability and its Applications, 40(2):273–284, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudlaev:1974:EDP

[Kud74] E. M. Kudlaev. On estimating distribution parameters from segments of a variational series. Theory of Probability and its Applications, 18(3):622–628, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudina:1976:DRS

[Kud76] L. S. Kudina. On decompositions of radially symmetric distributions. Theory of Probability and its Applications, 20(3):644–648, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudlaev:1978:DWS

[Kud78] E. M. Kudlaev. The distribution of a weighted Sherman statistic under approaching alternatives. Theory of Probability and its Applications, 22(4):794–804, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudlaev:1984:CGF

[Kud84] E. M. Kudlaev. A class of goodness-of-fit tests constructed by means of a central part of the variational series. Theory of Probability and its Applications, 28(3):585–595, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudlaev:1985:LCD

[Kud85] E. M. Kudlaev. Limiting conditional distributions for sums of random variables. Theory of Probability and its Applications, 29(4):776–786, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudlaev:1986:LTC

[Kud86] E. M. Kudlaev. Limit theorems for a class of separable statistics. Theory of Probability and its Applications, 30(1):187–192, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudlaev:1998:WCD

[Kud98] È. M. Kudlaev. Weak convergence of distributions of separable statistics. Theory of Probability and its Applications, 42 (1):130–148, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97595.

Original Russian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 85–107.

Kuhn:1981:SOB

[Küh81] T. Kühn. γ-summing operators in Banach spaces of type p, 1 < p ≤ 2, and cotype q, 2 ≤ q < ∞. Theory of Probability and its Applications, 26(1):118–129, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinic:1967:LBD

[Kul67] G. L. Kulinic. On the limit behavior of the distribution of the solution of a stochastic diffusion equation. Theory of Probability and its Applications, 12 (3):497–499, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinich:1968:LDS

[Kul68] G. L. Kulinich. Limit distributions of a solution of a stochastic diffusion equation. Theory of Probability and its Applications, 13(3):478–482, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinich:1974:EUS

[Kul74] G. L. Kulinich. On existence and uniqueness of a solution of a stochastic differential equation with martingale differential. Theory of Probability and its Applications, 19(1):168–171, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinich:1976:EDP

[Kul76] G. L. Kulinich. On the estimation of the drift parameter of a stochastic diffusion equation. Theory of Probability and its Applications, 20(2):384–387, ???? 1976. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinich:1983:NSC

[Kul83] G. L. Kulinich. On necessary and sufficient conditions for convergence of solutions to onedimensional stochastic diffusion equations with a nonregular dependence of the coefficients on a parameter. Theory of Probability and its Applications, 27 (4):856–862, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinich:1985:LIL

[Kul85a] G. L. Kulinich. On the law of the iterated logarithm for one-dimensional diffusion processes. Theory of Probability and its Applications, 29(3):563–566, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinskaya:1985:CES

[Kul85b] E. V. Kulinskaya. A consistent estimate for a scale parameter. Theory of Probability and its Applications, 29(4):798–803, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulinskaya:1987:LMP

[Kul87]  E. V. Kulinskaya.  On lo-

cally most powerful tests in rank ANOVA. Theory of Probability and its Applications, 31(3): 512–515, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulik:1998:RGR

[Kul98] A. M. Kulik. Regular Gaussian random operatorsand Strook– Varadhan theorem for symmetric stochastic Fredholm-type equations. Theory of Probability and its Applications, 42(2):244–253, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97610. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(2), (1997), pp. 262–273.

Kuritsyn:1985:MLL

[Kur85] Yu. G. Kuritsyn. On monotonicity in the law of large numbers for exchangeable random variables. Theory of Probability and its Applications, 29(1):150–153, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuritsyn:1987:LSM

[Kur87] Yu. G. Kuritsyn. On the leastsquares method for elliptically contoured distributions. Theory of Probability and its Applications, 31(4):738–740, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuritsyn:1988:SMA

[Kur88] Yu. G. Kuritsyn. On strong monotonicity of arithmetic means of exchangeable random variables. Theory of Probability and its Applications, 32(1):165–166, ???? 1988. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Kushnir:1968:AOT

[Kus68] A. F. Kushnir. Asymptotically optimal tests for a regression problem of testing hypotheses. Theory of Probability and its Applications, 13(4):647–666, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kutoyants:1976:PTH

[Kut76] Yu. A. Kutoyants. On a problem of testing hypotheses and asymptotic normality of stochastic integrals. Theory of Probability and its Applications, 20(2):376–384, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 385– 393.

Kutoyant:1978:ETP

[Kut78] Yu. A. Kutoyant. Estimation of the trend parameter of a diffusion process in the smooth case. Theory of Probability and its Applications, 22(2):399–405, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kutoyants:1979:EPD

[Kut79] Yu. A. Kutoyants. Estimation of a parameter of a diffusion process. Theory of Probability and its Applications, 23(3):641–649, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kutoyants:1985:EML

[Kut85] Yu. A. Kutoyants. Expansion of a maximum likelihood estimate by diffusion powers. Theory of Probability and its Applications, 29(3):465–477, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 29(3), (1984), pp. 452–463.

Kutoyants:1989:EEP

[Kut89] Yu. A. Kutoyants. An example of estimating a parameter of a nondifferentiable drift coefficient. Theory of Probability and its Applications, 33(1):175–179, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1974:CMP

[Kuz74] S. E. Kuznetsov. Construction of Markov processes with random times of birth and death. Theory of Probability and its Applications, 18(3):571–575, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1975:EFC

[Kuz75] S. E. Kuznetsov. Excessive functions connected with a regular Markov split process. Theory of Probability and its Applications, 20(1):145–150, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1978:CRS

[Kuz78] S. E. Kuznetsov. Construction of a regular split process. Theory of Probability and its Applications, 22(4):773–793, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1980:BEF

[Kuz80] S. Ye. Kuznetsov. Behavior of excessive functions along trajectories of a Markov process. Theory of Probability and its Applications, 24(3):492–508, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1981:MPB

[Kuz81] S. E. Kuznetsov. Any Markov process in a Borel space has a transition function. Theory of Probability and its Applications, 25(2):384–388, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1983:MTB

[Kuz83] V. P. Kuznetsov. Minimax tests for bounded families of distribution densities. Theory of Probability and its Applications, 27 (2):299–309, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1985:SST

[Kuz85] S. E. Kuznetsov. Specifications and a stopping theorem for random fields. Theory of Probability and its Applications, 29 (1):66–78, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1987:EHT

[Kuz87] S. E. Kuznetsov. On the existence of a homogeneous transition function. Theory of Probability and its Applications, 31 (2):244–254, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kleptsyna:1985:SSSa

[KV85] M. L. Kleptsyna and A. Yu.

Veretennikov. On strong solutions of stochastic Itô-Volterra equations. Theory of Probability and its Applications, 29 (1):153–157, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kvashko:1996:FIE

[Kva96] N. V. Kvashko. A forward interpolation equation of a semimartingale by observations over a point process. Theory of Probability and its Applications, 40 (1):162–165, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 177–180.

Kawazu:1971:BPI

[KW71] Kiyoshi Kawazu and Shinzo Watanabe. Branching processes with immigration and related limit theorems. Theory of Probability and its Applications, 16 (1):36–54, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylov:1964:MRS

[KY64] N. V. Krylov and A. A. Yushkevich. On Markov random sets. Theory of Probability and its Applications, 9(4):666–670, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kartashov:1965:EPP

[KY65] G. D. Kartashov and A. N. Yavriyan. On an extremal problem in probability theory. Theory of Probability and its Applications, 10(3):523–528, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:1974:SAE

[KY74] P. I. Kuznetsov and A. S. Yudina. Some asymptotic expansions for an incomplete probability integral. Theory of Probability and its Applications, 18 (2):355–359, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1984:ISM

[KY84] L. B. Klebanov and R. V. Yanushkyavichyus. ε-independence of the sample mean and tubular statistics. Theory of Probability and its Applications, 28 (1):165–172, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebanov:1986:ESB

[KY86] I. B. Klebanov and R. V.

Yanushkyavichyus. Estimation of stability in S. N. Bernshtein’s theorem. Theory of Probability and its Applications, 30(2): 383–386, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kashtanov:1997:JDC

[KY97] V. A. Kashtanov and I. M. Yanishevskiĭ. Joint distribution of catastrophe time and of the additive functional of profits. Theory of Probability and its Applications, 41(3):555–560, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 650–655.

Kruk:1996:AAS

[KZ96] L. Kruk and W. Ziȩba. On (B, ρ)- Amarts and almost sure convergence. Theory of Probability and its Applications, 40(3):549–556, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 677–685.

Labkovskii:1972:LTG

[Lab72] V. A. Labkovskii. A limit theorem for generalized random branching processes depending on the size of the population. Theory of Probability and its Applications, 17(1):72–

85, ????  1972.  CODEN

TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Labkovskii:1973:PWP

[Lab73] V. A. Labkovskii. On a property of a “Waiting process”. Theory of Probability and its Applications, 18(1):196–198, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Labkovskii:1979:APW

[Lab79] V. A. Labkovskii. Addendum: On a property of a “Waiting process”. Theory of Probability and its Applications, 23(3):672–674, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lamperti:1972:RMB

[Lam72] John Lamperti. Remarks on maximal branching processes. Theory of Probability and its Applications, 17(1):44–53, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lanery:1988:BSD

[Lan88] E. Lanery. Bayesian solutions without disintegration. Theory of Probability and its Applications, 32(1):168–175, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lanery:1990:ASB

[Lan90] É. Lanery. Addendum: Solutions bayesiennes sans désintegration. Theory of Probability and its Applications, 34(3):554, ???? 1990.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 613–613.

Larin:1994:CDS

[Lar94] Yu. V. Larin. On concentration of distributions of sums of independent random vectors on bounded sets. Theory of Probability and its Applications, 38 (4):743–751, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 882–891.

Lauter:1974:ROS

[Läu74] H. Läuter. Remarks on optimal synthesis of systems. Theory of Probability and its Applications, 18(3):592–598, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lax:1963:RSD

[Lax63] Peter D. Lax. On the regularity of spectral densities. Theory of Probability and its Applications, 8(3):316–319, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1968:RAT

[LE68] Yu. V. Linnik and V. L. Eidlin.

Remark on analytic transformations of normal vectors. Theory of Probability and its Applications, 13(4):707–710, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Le:1993:GGG

[Le93] Huiling Le. The global geodesic geometry of shape spaces and an application to collinearity testing. Theory of Probability and its Applications, 37(1): 76–79, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 75–80.

Lebedev:1971:FPT

[Leb71] V. A. Lebedev. On the first passage time of diffusion processes through time-dependent boundaries. Theory of Probability and its Applications, 16(3):541–545, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1973:TSS

[Leb73] V. A. Lebedev. On a transformation of systems of stochastic differential equations. Theory of Probability and its Applications, 17(4):706–709, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1974:BIE

[Leb74] V. A. Lebedev. On the backward interpolation equations for the jump component of a Markov process. Theory of Probability and its Applications, 18(2): 405–408, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1977:MES

[Leb77a] V. A. Lebedev. Moment estimates of the solution of a system of stochastic differential equations. Theory of Probability and its Applications, 21(3):586–593, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1977:CUS

[Leb77b] V. M. Lebedev. On a condition for uniqueness of the solution of a system of stochastic differential equations. Theory of Probability and its Applications, 21 (2):412–419, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1978:URS

[Leb78] V. A. Lebedev. On the uniqueness of a relaxed solution of a system of stochastic differential equations. Theory of Probability and its Applications, 23 (1):147–155, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1981:RCF

[Leb81] V. A. Lebedev. On relative compactness of families of distributions of semimartingales. Theory of Probability and its Applications, 26(1):140–148, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1982:WCF

[Leb82] V. A. Lebedev. On the weak compactness of families of distributions of general semimartingales. Theory of Probability and its Applications, 27(1):14– 23, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1984:RSS

[Leb84] V. A. Lebedev. On the regularity of a solution of a stochastic equation with respect to a martingale and a random measure. Theory of Probability and its Applications, 28(3):610–615, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1986:USS

[Leb86] V. A. Lebedev. On the uniqueness of a solution to a stochastic differential equation with driving martingale and random measure. Theory of Probability and its Applications, 30(1):169–174, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1987:SRP

[Leb87] V. A. Lebedev. On sequences of random processes with tight jump majorization. Theory of Probability and its Applications, 31(3):529–533, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1990:SIR

[Leb90] V. A. Lebedev. Stochastic integration with respect to semimartingale random measures. Theory of Probability and its Applications, 34(4):725–727, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 792–794.

Lebedev:1996:BRM

[Leb96a] V. A. Lebedev. Behavior of random measures under filtration change. Theory of Probability and its Applications, 40 (4):645–652, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 754–763.

Lebedev:1996:FTS

[Leb96b] V. A. Lebedev. The Fubini theorem for stochastic integrals with respect to L^{0}-valued random measures depending on a parameter. Theory of Probability and its Applications, 40 (2):285–293, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lehmann:1965:NVC

[Leh65] E. L. Lehmann. On the nonverifiability of certain parametric functions. Theory of Probability and its Applications, 10 (4):689–691, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leipnik:1982:LDS

[Lei82]  R. Leipnik.  The lognormal

distribution and strong nonuniqueness of the moment problem. Theory of Probability and its Applications, 26(4):850–852, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leonov:1961:DTD

[Leo61] V. P. Leonov. On the dispersion of time-dependent means of a stationary stochastic process. Theory of Probability and its Applications, 6(1):87–93, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leonteva:1978:QSA

[Leo78] N. P. Leont’eva. Queueing systems with arbitrary initial conditions. Theory of Probability and its Applications, 22(4):813–818, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leonteva:1982:ASR

[Leo82] N. T. Leont’eva. Addendum: Summary of reports presented at sessions of the Seminar on Probability Theory and Mathematical Statistics at the Mathematical Institute of the Siberian Section of the USSR Academy of Sciences, 1979, p. 652. Theory of Probability and its Applications, 26(2):434–435, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leontovich:1986:CSG

[Leo86] A. M. Leontovich. On a condition for slowly growing number of dependencies. Theory of Probability and its Applications, 30(1):196–201, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lepskii:1988:AMP

[Lep88]  O. V. Lepskii.  Asymptotic

minimax parameter estimation for nonsymmetric loss functions. Theory of Probability and its Applications, 32(1):160–164, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lepskii:1990:AME

[Lep90] O. V. Lepskiĭ. Asymptotic minimax estimation with prescribed properties. Theory of Probability and its Applications, 34 (4):604–615, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 664–676.

Lepskii:1991:PAE

[Lep91] O. V. Lepskiĭ. On a problem of adaptive estimation in Gaussian white noise. Theory of Probability and its Applications, 35(3): 454–466, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 459–470.

Lepskii:1992:AMA

[Lep92] O. V. Lepskiĭ. Asymptotically minimax adaptive estimation. I. Upper bounds. Optimally adaptive estimates. Theory of Probability and its Applications, 36(4): 682–697, December 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 645–659.

Lepskii:1993:AMA

[Lep93]  O. V. Lepskiĭ.  Asymptoti-

cally minimax adaptive estimation. II. Schemes without optimal adaptation: Adaptive estimators. Theory of Probability and its Applications, 37(3): 433–448, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 468–481.

Lepskii:1994:EMN

[Lep94] O. V. Lepskiĭ. Estimation of the maximum of a nonparametric signal to within a constant. Theory of Probability and its Applications, 38(1):152–158, March 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 187–194.

Letchikov:1989:LTR

[Lët89] A. V. Lëtchikov. A limit theorem for a random walk in a random environment. Theory of Probability and its Applications, 33 (2):228–238, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 246–256.

Letchikov:1991:SKP

[Lët91] A. V. Lëtchikov. On a solution of Kesten’s problem. Theory of Probability and its Applications, 35(2):360–363, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 358–361.

Letchikov:1993:CAC

[Lët93] A. V. Lëtchikov. A criterion for applicability of the Central Limit Theorem to one-dimensional random walks in random environments. Theory of Probability and its Applications, 37(3):553– 557, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 576–580.

Lev:1972:SMP

[Lev72] G. Sh Lev. Semi-Markov processes of multiplication with drift. Theory of Probability and its Applications, 17(1):159–164, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lev:1973:CSM

[Lev73] G. Sh. Lev. On convergence of semi-Markov multiplication processes with drift to a diffusion process. Theory of Probability and its Applications, 17(3): 551–556, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lev:1974:APD

[Lev74] G. Sh. Lev. Asymptotic properties of the degeneracy probability for semi-Markov multiplication processes. Theory of Probability and its Applications, 18 (4):740–752, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lev:1975:APP

[Lev75a] G. Sh. Lev. Asymptotic properties of the probability of degeneration after time t for Markov multiplication processes. Theory of Probability and its Applications, 20(1):161–169, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1975:BGB

[Lev75b] V. Ja. Levit. On the behavior of generalized Bayes estimators for Markov observations. Theory of Probability and its Applications, 19(2):332–346, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1976:ECN

[Lev76] B. Ya. Levit. On the efficiency of a class of non-parametric estimates. Theory of Probability and its Applications, 20(4):723–740, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lev:1978:APE

[Lev78] G. Sh. Lev. Asymptotic properties of the extinction probability for a Markov multiplication process. Theory of Probability and its Applications, 22(4):825–831, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1979:IDI

[Lev79] B. Ya. Levit. Infinite-dimensional informational inequalities. Theory of Probability and its Applications, 23(2):371–377, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lev:1980:DAT

[Lev80] G. Sh. Lev. On the distribution of the absorption time for a semi-Markov multiplication process. Theory of Probability and its Applications, 24(4):876–882, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1981:AME

[Lev81] B. Ya. Levit. On asymptotic minimax estimates of the second order. Theory of Probability and its Applications, 25(3):552–568, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1983:MEP

[Lev83] B. Ya. Levit. Minimax estimation and positive solutions of elliptic equations. Theory of Probability and its Applications, 27 (3):563–586, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1986:SOA

[Lev86a] B. Ya. Levit. Second-order asymptotic optimality and positive solutions of Schrödinger’s equation. Theory of Probability and its Applications, 30(2): 333–363, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levitskaya:1986:PCS

[Lev86b] A. A. Levitskaya. The probability of consistency of a system of random linear equations over an arbitrary finite ring. Theory of Probability and its Applications, 30(2):364–375, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lev:1988:GMP

[Lev88] G. Sh. Lev. Generalized multiplication processes. Theory of Probability and its Applications, 32(4):686–695, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Licea:1969:SPO

[Lic69] G. Licea. On supermartingales with partially ordered parameter set. Theory of Probability and its Applications, 14(1):135–137, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1976:AAC

[Lif76a] B. A. Lifshits. On the accuracy of approximation in the Central Limit Theorem. Theory of Probability and its Applications, 21 (1):108–124, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1976:CMC

[Lif76b] B. A. Lifshits. On the convergence of moments in the Central Limit Theorem for nonhomogeneous Markov chains. Theory of Probability and its Applications, 20(4):741–758, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 755– 772.

Lifshits:1979:CLT

[Lif79] B. A. Lifshits. On the Central Limit Theorem for Markov chains. Theory of Probability and its Applications, 23(2):279–296, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1980:RLF

[Lif80] M. A. Lifshits. On representation of Lévy’s fields by indicators. Theory of Probability and its Applications, 24(3):629–633, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1982:FMA

[Lif82]  M. A. Lifshits.  The fiber

method and its application to the study of functionals of stochastic processes. Theory of Probability and its Applications, 27 (1):69–83, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1983:ACD

[Lif83] M. A. Lifshits. On the absolute continuity of distributions of functionals of random processes. Theory of Probability and its Applications, 27(3):600–607, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1985:IPW

[Lif85a] B. A. Lifshits. Invariance principle for weakly dependent variables. Theory of Probability and its Applications, 29(1):33– 40, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1985:ASM

[Lif85b] M. A. Lifshits. An application of the stratification method to the study of functionals of processes with independent increments. Theory of Probability and its Applications, 29(4):753–765, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:1987:DMG

[Lif87] M. A. Lifshits. On the distribution of the maximum of a Gaussian process. Theory of Probability and its Applications, 31 (1):125–132, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1956:QFG

[Lin56a] Yu. V. Linnik. On the question of finding a general distribution from the distribution of a statistic. Theory of Probability and its Applications, 1 (4):422–434, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1956:RCT

[Lin56b] Yu. V. Linnik. A remark on Cramer’s theorem on the decomposition of the normal law. Theory of Probability and its Applications, 1(4):435–436, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1957:DCG

[Lin57a] Yu. V. Linnik. On the decomposition of the convolution of Gaussian and Poissonian laws. Theory of Probability and its Applications, 2(1):31–57, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1957:SRM

[Lin57b] Yu. V. Linnik. Some remarks on the method of least squares, with application to direct and inverse location problems. Theory of Probability and its Applications, 2(3):332–344, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1958:GTF

[Lin58] Yu. V. Linnik. General theorems on the factorization of infinitely divisible laws. Theory of Probability and its Applications, 3(1):1–37, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1959:ITP

[Lin59a] Ju. V. Linnik. An informationtheoretic proof of the Central Limit Theorem with Lindeberg conditions. Theory of Probability and its Applications, 4(3): 288–299, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1959:GTFa

[Lin59b] Yu. V. Linnik. General theorems on the factorization of infinitely divisible laws II. Sufficient conditions (case of a finite Poisson spectrum). Theory of Probability and its Applications, 4(1):53–82, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1959:GTFb

[Lin59c] Yu. V. Linnik. General theorems on the factorization of infinitely divisible laws III. Sufficient conditions (countable bounded Poisson spectrum. unbounded spectrum. “Stability”). Theory of Probability and its Applications, 4(2):142–163, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1961:LTSa

[Lin61a] Yu. V. Linnik. Limit theorems for sums of independent variables taking into account large deviations. I. Theory of Probability and its Applications, 6 (2):131–148, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1961:LTSb

[Lin61b] Yu. V. Linnik. Limit theorems for sums of independent variables taking into account large deviations. II. Theory of Probability and its Applications, 6 (4):345–360, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1962:LTS

[Lin62] Yu. V. Linnik. Limit theorems for sums of independent variables taking into account large deviations. III. Theory of Probability and its Applications, 7 (2):115–129, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1964:WTC

[Lin64] Yu. V. Linnik. On A. Wald’s test comparing two normal samples. Theory of Probability and its Applications, 9(1):14–27, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1966:AMD

[Lin66] Yu. V. Linnik. Approximate minimax detection of a vector signal in Gaussian noise. Theory of Probability and its Applications, 11(4):497–512, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linde:1990:GMT

[Lin90] W. Linde. Gaussian measure of translated balls in a Banach space. Theory of Probability and its Applications, 34(2):307–317, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 349–359.

Lin:1992:IPS

[Lin92] Z. Y. Lin. On the increments of partial sums of a φ-mixing sequence. Theory of Probability and its Applications, 36 (2):316–328, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1966:CLN

[Lip66] R. Sh. Liptser. A comparison of linear and nonlinear filtration of some Markov processes. Theory of Probability and its Applications, 11(3):467–472, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1967:FPC

[Lip67] R. Sh. Liptser. On filtration and prediction of components in diffusion Markov processes. Theory of Probability and its Applications, 12(4):697–698, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1976:GMG

[Lip76] R. Sh. Liptser. Gaussian martingales and a generalization of the Kalman-bucy filter. Theory of Probability and its Applications, 20(2):285–301, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1977:RLM

[Lip77] R. Sh. Liptser. On the representation of local martingales. Theory of Probability and its Applications, 21(4):698–705, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1997:LDO

[Lip97] R. Sh. Liptser. Large deviations for occupation measures of Markov processes: Discrete time, noncompact case. Theory of Probability and its Applications, 41(1):35–54, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97502.

Original Russian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 65–88.

Liskovets:1976:NMV

[Lis76] V. A. Liskovets. On the number of maximal vertices of a random acyclic digraph. Theory of Probability and its Applications, 20(2):401–409, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lisitskii:1994:OPD

[Lis94] A. D. Lisitskii. One property of M-divisibility of strictly stable vectors. Theory of Probability and its Applications, 38 (1):158–161, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 195–198.

Lisyanoi:1996:GAT

[Lis96] S. V. Lisyanoĭ. On a group analogue of a theorem of A. M. Kagan. Theory of Probability and its Applications, 40(1):165–167, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 180–183.

Lisitskii:1998:EPM

[Lis98] A. D. Lisitskii. The Eaton problem and multiplicative properties of multivariate distributions. Theory of Probability and its Applications, 42(4):618– 632, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97645.

Livshits:1975:IPB

[Liv75] A. N. Livshits. On the isomorphism problem for Bernoulli schemes. Theory of Probability and its Applications, 19(2): 394–399, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Livshits:1980:AIP

[Liv80] A. N. Livshits. Addendum:

On the isomorphism problem for Bernoulli schemes. Theory of Probability and its Applications, 25(1):214–215, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levit:1974:SEP

[LK74] B. Ya. Levit and R. Z. Khas’minskii.

Sequential estimation of the parameter of a Markov chain. Theory of Probability and its Applications, 18(3):546–558, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:1994:GEU

[LL94] A. A. Lebedev and V. A. Lebedev. Generalization of the expectation for a unified description of random and indefinite factors. Theory of Probability and its Applications, 38(3):470–478, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 529–539.

Lloyd:1962:MSD

[Llo62] S. P. Lloyd. On a measure of stochastic dependence. Theory of Probability and its Applications, 7(3):301–312, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ladokhin:1971:ERT

[LM71] V. I. Ladokhin and D. A. Moskvin. An estimate for the remainder term in the Central Limit Theorem for sums of functions of independent variables and sums of the form ∑ f(t2^{k}). Theory of Probability and its Applications, 16(1): 116–125, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lamperti:1968:CBP

[LN68] John Lamperti and Peter Ney. Conditioned branching processes and their limiting diffusions. Theory of Probability and its Applications, 13(1):128–139, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lyashenko:1991:PAR

[LN91] N. N. Lyashenko and M. S. Nikulin. Probability analysis of round-off errors in floating-point arithmetic. Theory of Probability and its Applications, 35 (1):66–74, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lyashenko:1989:FPR

[LNF89] N. N. Lyashenko, M. S. Nikulin, and S. I. Fedorov. The forecasting problem for random processes with allowance for the duration of calculation. Theory of Probability and its Applications, 33 (3):571–574, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leonenko:1996:LTS

[LO96] N. Leonenko and E. Orsingher.

Limit theorems for solutions of the Burgers equation with Gaussian and non-Gaussian initial conditions. Theory of Probability and its Applications, 40 (2):294–308, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 387–403.

Lodkin:1971:ACM

[Lod71] A. A. Lodkin. Absolute continuity of measures corresponding to Markov processes with discrete time. Theory of Probability and its Applications, 16(4):690–694, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Logunov:1991:ECR

[Log91] P. L. Logunov. Estimates for the convergence rate to the Poisson distribution for random sums of independent indicators. Theory of Probability and its Applications, 35(3):587–590, September 1991. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 580–582.

Lotov:1980:AADa

[Lot80a] V. I. Lotov. Asymptotic analysis of distributions in problems with two boundaries. I. Theory of Probability and its Applications, 24(3):480–491, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lotov:1980:AADb

[Lot80b] V. I. Lotov. Asymptotic analysis of distributions in problems with two boundaries. II. Theory of Probability and its Applications, 24(4):869–876, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lotov:1986:ABC

[Lot86] V. I. Lotov. On the asymptotic behavior of characteristics of the sequential likelihood ratio test. Theory of Probability and its Applications, 30(1):181–187, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lotov:1988:AES

[Lot88] V. I. Lotov. Asymptotic expansions in a sequential likelihood ratio test. Theory of Probability and its Applications, 32 (1):57–67, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lotov:1989:AAS

[Lot89] V. I. Lotov. On the accuracy of approximation in the sequential Wald test. Theory of Probability and its Applications, 33 (2):276–285, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 295–304.

Lotov:1992:RWS

[Lot92] V. I. Lotov. On random walks in a strip. Theory of Probability and its Applications, 36(1):165–170, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 160–165.

Lototskii:1996:RAT

[Lot96] S. V. Lototskii. Robust algorithms of the type of stochastic approximation (continuous time). Theory of Probability and its Applications, 40(2):309–328, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linde:1975:MGC

[LP75] W. Linde and A. Pietsch. Mappings of Gaussian cylindrical measures in Banach spaces. Theory of Probability and its Applications, 19(3):445–460, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1967:TAM

[LPS67] Yu. V. Linnik, Yu. V. Prokhorov, and O. V. Shalaevskii. On the theory of approximately minimax detection of signals in Gaussian noise. Theory of Probability and its Applications, 12(3):347–363, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lyubarskij:1963:TDE

[LR63] G. Ya. Lyubarskij and Yu. L. Rabotnikov. On the theory of differential equations with random coefficients. Theory of Probability and its Applications, 8 (3):290–298, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linnik:1965:RTF

[LRS65] Yu. V. Linnik, I. L. Romanovskaya, and O. V. Shalaevskii. A remark on the theory of the Fisher–Welch–Wald test. Theory of Probability and its Applications, 10(4):659–662, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leonov:1959:MCS

[LS59] V. P. Leonov and A. N. Shiryaev. On a method of calculation of semi-invariants. Theory of Probability and its Applications, 4 (3):319–329, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Leonov:1960:SPS

[LS60] V. P. Leonov and A. N. Shiryaev.

Some problems in the spectral theory of higher-order moments, II. Theory of Probability and its Applications, 5(4):417–421, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1965:BPS

[LS65] R. Sh. Liptser and A. N.

Shiryaev. A Bayesian problem of sequential search in diffusion approximation. Theory of Probability and its Applications, 10 (1):178–186, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1968:EMM

[LS68a] R. Sh. Liptser and A. N.

Shiryaev. The extrapolation of multidimensional Markov processes from incomplete data. Theory of Probability and its Applications, 13(1):15–38, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1968:NLI

[LS68b] R. Sh. Liptser and A. N.

Shiryaev. Non-linear interpolation of components of Markov diffusion processes (direct equations, effective formulas). Theory of Probability and its Applications, 13(4):564–583, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lumelskii:1969:UED

[LS69] Ya. P. Lumel’skii and P. N. Sapozhnikov. Unbiased estimates of density functions. Theory of Probability and its Applications, 14(2):357–364, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Legostaeva:1971:MWT

[LS71] I. L. Legostaeva and A. N. Shiryaev. Minimax weights in a trend detection problem of a random process. Theory of Probability and its Applications, 16 (2):344–349, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1981:NSC

[LS81a] R. Sh. Liptser and A. N. Shiryaev. On necessary and sufficient conditions in the functional Central Limit Theorem for semimartingales. Theory of Probability and its Applications, 26 (1):130–135, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1981:FCL

[LS81b] R. Sh. Liptser and A. N. Shiryayev. A functional Central Limit Theorem for semimartingales. Theory of Probability and its Applications, 25(4):667–688, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1982:RCC

[LS82] R. Sh. Liptser and A. N. Shiryaev. On the rate of convergence in the Central Limit Theorem for semimartingales. Theory of Probability and its Applications, 27(1):1–13, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liptser:1984:IPS

[LS84] R. Sh. Liptser and A. N.

Shiryaev. On the invariance principle for semi-martingales: the “Nonclassical” case. Theory of Probability and its Applications, 28(1):1–34, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lo:1996:MSB

[LS96] A. Y. Lo and V. V. Sazonov.

The von Mises ω^{2}-statistic and the bootstrap. Theory of Probability and its Applications, 40 (4):653–659, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 850–858.

Lazakovitch:1997:AIS

[LS97] N. V. Lazakovitch and S. P.

Stashulenok. An approximation of the Ito and Stratonovich stochastic integrals by elements of a direct product of algebras of generalized random processes. Theory of Probability and its Applications, 41(4): 695–715, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97570.

Lazakovitch:1999:ASE

[LSS99] N. V. Lazakovitch, S. P. Stashulenok, and T. V. Stemkovskaia. Associated solutions of equations in differentials in the direct product of algebras of generalized random processes. Theory of Probability and its Applications, 43 (2):221–238, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97684.

Linde:1980:CCC

[LTC80] V. Linde, V. I. Tarieladze, and S. A. Chobanyan. Characterization of certain classes of Banach spaces by properties of Gaussian measures. Theory of Probability and its Applications, 25 (1):159–164, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Luu:1993:FCM

[Lu’93] – Dinh Quang Lu’u. On further classes of martingale-like sequences. Theory of Probability and its Applications, 37 (2):366–371, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 428–434.

Lukacs:1968:CPD

[Luk68] E. Lukacs. Contributions to a problem of D. van Dantzig. Theory of Probability and its Applications, 13(1):116–127, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lumelskii:1969:CLL

[Lum69] Ya. P. Lumel’skii. Confidence limits for linear functions of unknown parameters. Theory of Probability and its Applications, 14(2):364–367, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lumelskii:1971:AEC

[Lum71] Ya. P. Lumel’skii. Addendum:

Estimates of certain functions of known parameters in problems of random walk. Theory of Probability and its Applications, 16(2):397, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Luschgy:1996:LER

[Lus96] H. Luschgy. Linear estimators and Radonifying operators. Theory of Probability and its Applications, 40(1):167–175, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 205–213.

Luschgy:1999:MEI

[Lus99] H. Luschgy. Minimaxity and equivariance in infinite dimension. Theory of Probability and its Applications, 43(3):388– 404, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97704.

Lutsenko:1978:PMD

[Lut78] M. M. Lutsenko. The problem of δ-meeting on a 3-dimensional ball. Theory of Probability and its Applications, 23(1):190–195, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lutsenko:1990:GTM

[Lut90] M. M. Lutsenko. Game-theoretic method of estimating the parameter of a binomial law from results of one observation. Theory of Probability and its Applications, 34(3):531–535, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 589–593.

Lutsenko:1991:GTM

[Lut91] M. M. Lutsenko. A game theoretic method for the estimation of a parameter of the binomial law. Theory of Probability and its Applications, 35(3): 467–477, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 471–481.

Leonenko:1979:IPH

[LY79] N. N. Leonenko and M. I. Yadrenko. On the invariance principle for homogeneous and isotropic random fields. Theory of Probability and its Applications, 24(1):175–181, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lyashenko:1987:GCR

[Lya87a] N. N. Lyashenko. The geometric convergence of random processes with arbitrary discontinuities. Theory of Probability and its Applications, 31(3):518–520, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lyashenko:1987:GRP

[Lya87b] N. N. Lyashenko. Graphs of random processes as random sets. Theory of Probability and its Applications, 31 (1):72–80, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1978:SLD

[MA78] T. L. Malevich and B. Abdalimov. Stable limit distributions for U-statistics. Theory of Probability and its Applications, 22 (2):370–377, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1979:LDP

[MA79] T. L. Malevich and B. Abdalimov. Large deviation probabilities for U-statistics. Theory of Probability and its Applications, 24(1):215–219, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1983:RCL

[MA83] T. L. Malevich and B. Abdalimov. A refinement of the Central Limit Theorem for U-statistics of m-dependent variables. Theory of Probability and its Applications, 27(2):391–396, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Machis:1969:SDU

[Mac69] Yu. Yu. Machis. On the stability of decompositions of the unit distribution function. Theory of Probability and its Applications, 14(4):688–690, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Machis:1971:EST

[Mac71a] J. J. Machis. Estimates of the stability theorem for decompositions of the Poisson distributions. Theory of Probability and its Applications, 16(2):215–227, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Machis:1971:NCL

[Mac71b] Yu. Yu. Machis. Non-classical limit theorems. Theory of Probability and its Applications, 16 (1):175–182, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maimin:1975:DDC

[Mai75a] Z. G. Maimin. On detection of the deterministic component of a random field. Theory of Probability and its Applications, 19 (3):506–521, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maister:1975:BDB

[Mai75b] P. I. Maister. Branching diffusion in a bounded domain with absorbing boundary. Theory of Probability and its Applications, 19(3):563–569, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maister:1979:ETC

[Mai79] P. I. Maister. Expectation of a time-continuous branching diffusion process. Theory of Probability and its Applications, 23 (4):801–805, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maiboroda:1990:CLT

[Maĭ90] R. E. Maĭboroda. The Central Limit Theorem for empirical moment generating functions. Theory of Probability and its Applications, 34(2):332–335, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 375–379.

Maiboroda:1991:NMC

[Maĭ91] R. E. Maĭboroda. A nonparametric method in the changepoint problem for multivariate observations. Theory of Probability and its Applications, 35(3):

590–593, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 582–586.

Maiboroda:1995:SPS

[Maĭ95] R. E. Maĭboroda. Some problems of statistics of distribution mixtures with variable concentrations. Theory of Probability and its Applications, 39(4):703–707, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 840–845.

Maksimov:1965:RBL

[Mak65] V. M. Maksimov. On the relation between limit theorems of finite groups and an ergodic theorem for Markov chains with doubly stochastic transition matrices. Theory of Probability and its Applications, 10(3):493–496, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1966:SCC

[Mak66] V. M. Maksimov. On the speed of convergence of the convolution of identical distributions defined on a finite group. Theory of Probability and its Applications, 11 (3):458–463, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1967:NSS

[Mak67a] V. M. Maksimov. Necessary and sufficient statistics for the family of shifts of probability distributions on continuous BiCompact groups. Theory of Probability and its Applications, 12 (2):267–280, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1967:CPI

[Mak67b] V. M. Maksimov. On the convergence of products of independent random variables taking on values from an arbitrary finite group. Theory of Probability and its Applications, 12(4):619–637, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1968:NSC

[Mak68] V. M. Maksimov. Necessary and sufficient convergence conditions for the convolution of nonidentical distributions given on a finite group. Theory of Probability and its Applications, 13 (2):287–298, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1970:CNH

[Mak70a] V. M. Maksimov. Convergence of non-homogeneous bistochastic Markov chains. Theory of Probability and its Applications, 15 (4):604–618, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1970:RPI

[Mak70b] V. M. Maksimov. Random processes with independent increments with values in an arbitrary finite group. Theory of Probability and its Applications, 15 (2):215–228, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1971:CCS

[Mak71a] V. M. Maksimov. Composition convergent sequences of measures on compact groups. Theory of Probability and its Applications, 16(1):55–73, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1971:DDF

[Mak71b] V. M. Maksimov. Divisible distributions on finite groups. Theory of Probability and its Applications, 16(2):308–322, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1973:NSM

[Mak73] V. M. Maksimov. Nonhomogeneous semigroups of measures on compact Lie groups. Theory of Probability and its Applications, 17(4):601–619, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1975:IES

[Mak75] V. M. Maksimov. Invariant estimates of the shift for distributions on finite groups. Theory of Probability and its Applications, 20(1):69–82, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maksimov:1980:LDR

[Mak80] K. N. Maksimov. Limit distributions of a random variable defined on a Markov chain. Theory of Probability and its Applications, 24(3):633–636, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Makarov:1981:LDC

[Mak81] G. D. Makarov. Large deviations for a critical Galton–Watson process. Theory of Probability and its Applications, 25(3):481–492, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Makarov:1982:EDF

[Mak82a] G. D. Makarov. Estimates for the distribution function of a sum of two random variables when the marginal distributions are fixed. Theory of Probability and its Applications, 26(4):803–806, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Makarov:1982:LDB

[Mak82b] G. D. Makarov. Large deviations in branching processes close to critical ones. Theory of Probability and its Applications, 27 (1):149–154, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Makarov:1988:ABL

[Mak88] A. A. Makarov. The asymptotic behavior of the limit distribution of the Kolmogorov–Smirnov statistic in the case of a composite hypothesis for the class of projecting estimates of an unknown parameter. Theory of Probability and its Applications, 32(2):380–383, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 32(2), (1987), pp. 412–416.

Makarichev:1991:RLC

[Mak91] A. V. Makarichev. Reliability of a large complex of singletype renewable redundant systems. Theory of Probability and its Applications, 35(4):792–797, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 775–779.

Makhno:1995:TLD

[Mak95] Ya. S. Makhno. A theorem on large deviations for one class of diffusion processes. Theory of Probability and its Applications, 39(3):437–447, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 554–566.

Makarichev:1996:OAE

[Mak96a] A. V. Makarichev. Optimal allocation of elements in a complex of renewable standby systems. Theory of Probability and its Applications, 40(1):66–75, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 84–95.

Makhno:1996:LDS

[Mak96b] S. Ya. Makhno. Large deviations for solutions of stochastic equations. Theory of Probability and its Applications, 40(4):660–678, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 764–785.

Maksimov:1997:CSM

[Mak97] V. M. Maksimov. Cubic stochastic matrices and their probability interpretation. Theory of Probability and its Applications, 41(1):55–69, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97503.

Malevich:1964:ABE

[Mal64] T. L. Malevich. The asymptotic behavior of an estimate for the spectral function of a stationary Gaussian process. Theory of Probability and its Applications, 9(2):350–353, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1965:SPE

[Mal65a] T. L. Malevich. Some properties of estimators of the spectrum of a stationary process. Theory of Probability and its Applications, 10(3):457–465, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maloshevskii:1965:LTS

[Mal65b] S. G. Maloshevskii. Limit theorems for a special kind of random walk. Theory of Probability and its Applications, 10(3):507–512, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maloshevskii:1968:SES

[Mal68] S. G. Maloshevskii. Sharpness of an estimate of N. A. Sapogov in the stability problem of Cramér’s theorem. Theory of Probability and its Applications, 13(3):494–496, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1969:ANN

[Mal69a] T. L. Malevich. Asymptotic normality of the number of crossings of level zero by a Gaussian process. Theory of Probability and its Applications, 14(2):287–295, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1969:AEL

[Mal69b] T. L. Malevich. On the asymptotic efficiency of least-squares estimates of regression coefficients. Theory of Probability and its Applications, 14(1):173–175, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevic:1974:SBP

[Mal74a] T. L. Malevic. Some bounds for the probabilities of events generated by Gaussian processes and fields, and applications to problems of level crossings. Theory of Probability and its Applications, 19(1):142–153, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyutov:1974:AMD

[Mal74b] M. B. Malyutov. Addendum:

Minimax designs for testing the degree of a polynomial. Theory of Probability and its Applications, 18(4):840, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevic:1975:LTL

[Mal75] T. L. Malevic. A limit theorem for the length of contours generated by crossings of the zero level by Gaussian fields. Theory of Probability and its Applications, 19(3):474–486, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1977:ASB

[Mal77] T. L. Malevich. Addendum:

Some bounds for the probabilities of events generated by Gaussian processes and fields, and applications to problems of level crossings. Theory of Probability and its Applications, 22(1):196–197, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1980:CMN

[Mal80] T. L. Malevich. On conditions for the moments of the number of zeros of Gaussian stationary processes to be finite. Theory of Probability and its Applications, 24(4):741–754, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1985:CFF

[Mal85a] T. L. Malevich. On conditions for finiteness of the factorial moments of the number of zeros of Gaussian stationary processes. Theory of Probability and its Applications, 29(3):534–545, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malkov:1985:NUE

[Mal85b] A. S. Mal’kov. On non-uniform estimates of the convergence rate in one- and multi-dimensional local limit theorems. Theory of Probability and its Applications, 29(3):582–586, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malinovskii:1986:SOE

[Mal86] V. K. Malinovskii. On secondorder efficiency of an asymptotically efficient test based on Markov observations. Theory of Probability and its Applications, 30(3):603–608, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malinovskii:1987:AEC

[Mal87a] V. K. Malinovskii. Asymptotic expansions in the Central Limit Theorem for recurrent Markov renewal processes. Theory of Probability and its Applications, 31(3):523–526, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malinovskii:1987:LTH

[Mal87b] V. K. Malinovskii. Limit theorems for Harris Markov chains, I. Theory of Probability and its Applications, 31(2):269–285, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malinovskii:1990:LTH

[Mal90a] V. K. Malinovskiĭ. Limit theorems for Harris Markov chains, II. Theory of Probability and its Applications, 34(2):252–265, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 289–303.

Malinovskii:1990:PDF

[Mal90b] V. K. Malinovskiĭ. Power and defect functions of asymptotically efficient tests in the case of Markov observed values. Theory of Probability and its Applications, 34(3):441–453, ???? 1990.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 491–504.

Malinkovskii:1991:CPI

[Mal91] Yu. V. Malinkovskiĭ. A criterion for pointwise independence of states of units in an open stationary Markov queueing network with one class of customers. Theory of Probability and its Applications, 35(4):797–802, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 779–784.

Malinovskii:1992:ALT

[Mal92a] V. K. Malinovskii. Addendum: Limit theorems for Harris Markov chains. I. Theory of Probability and its Applications, 36(2):426, ???? 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Malinovskii:1992:LDR

[Mal92b] V. K. Malinovskiĭ. Large deviations for recurrent Markov renewal processes. Theory of Probability and its Applications, 36 (1):170–173, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 165–167.

Malinovskii:1993:AEP

[Mal93] V. K. Malinovskiĭ. Asymptotic expansions in the problem of sequential estimation of an autoregressive parameter. Theory of Probability and its Applications, 37(2):381–383, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 426–428.

Malinovskii:1994:LTS

[Mal94] V. K. Malinovskiĭ. Limit theorems for stopped random sequences. I: Rates of convergence and asymptotic expansions. Theory of Probability and its Applications, 38(4):673–693, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 800–826.

Malinovskii:1995:AAE

[Mal95] V. K. Malinovskiĭ. Addendum:

Asymptotic expansions in the problem of sequential estimation of an autoregressive parameter. Theory of Probability and its Applications, 39(4):719, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 864–864.

Malinovskii:1997:LTS

[Mal97] V. K. Malinovskiĭ. Limit theorems for stopped random sequences II: Large deviations. Theory of Probability and its Applications, 41(1):70–90, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97504. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 107–132.

Mandl:1964:DPW

[Man64a] Petr Mandl. Diffusion processes with a weakly absorbing barrier. Theory of Probability and its Applications, 9(2):258–267, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mandl:1964:CNT

[Man64b] Petr Mandl. On the control of non-terminating diffusion processes. Theory of Probability and its Applications, 9(4):591–603, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mandl:1967:CWP

[Man67] P. Mandl. On the control of a Wiener process for a limited number of switchings. Theory of Probability and its Applications, 12(1):68–76, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maniya:1968:QEE

[Man68] G. M. Maniya. Quadratic error in estimating multidimensional normal distribution densities from sample data. Theory of Probability and its Applications, 13 (2):341–343, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Marushin:1965:SRC

[Mar65] M. N. Marushin. Some remarks on convergence to a limiting Poisson distribution. Theory of Probability and its Applications, 10 (2):341–344, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martynov:1969:LBT

[Mar69] G. V. Martynov. Linear binary two-sample tests for shift alternatives. Theory of Probability and its Applications, 14(1):158–163, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maruyama:1970:IDP

[Mar70] G. Maruyama. Infinitely divisible processes. Theory of Probability and its Applications, 15(1):1–22, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martynov:1976:CDF

[Mar76a] G. V. Martynov. Computation of distribution functions of quadratic forms of normally distributed random variables. Theory of Probability and its Applications, 20(4):782–793, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martynov:1976:CLD

[Mar76b] G. V. Martynov. Computation of limit distributions of statistics for normality tests of type ω^{2}. Theory of Probability and its Applications, 21(1):1–13, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martynov:1978:GSF

[Mar78] G. V. Martynov. A generalization of Smirnov’s formula for the distribution functions of quadratic forms. Theory of Probability and its Applications, 22(3):602–607, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1979:LIL

[Mar79] A. I. Martikainen. On the law of the iterated logarithm for perturbed and weighted sums. Theory of Probability and its Applications, 23(2):362–365, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1980:NSC

[Mar80] A. I. Martikainen. On necessary and sufficient conditions for the strong law of large numbers. Theory of Probability and its Applications, 24(4):813–820, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1981:CLI

[Mar81] A. I. Martikainen. A converse to the law of the iterated logarithm for a random walk. Theory of Probability and its Applications, 25(2):361–362, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1984:OSV

[Mar84] A. I. Martikainen. One-sided versions of strong limit theorems. Theory of Probability and its Applications, 28(1):48– 64, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1985:CSC

[Mar85] A. I. Martikainen. Criteria for strong convergence of normalized sums of independent random variables and their applications. Theory of Probability and its Applications, 29(3):519–533, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1986:OSL

[Mar86a] A. I. Martikainen. On the onesided law of the iterated logarithm. Theory of Probability and its Applications, 30(4):736–749, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1986:RMS

[Mar86b] A. I. Martikainen. Regular methods of summing random terms. Theory of Probability and its Applications, 30(1):9–18, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1987:CYV

[Mar87a] A. I. Martikainen. On a conjecture of Yu. V. Prokhorov. Theory of Probability and its Applications, 31(4):697–703, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1987:CLN

[Mar87b] A. I. Martikainen. On Chung’s law for nonidentically distributed independent summands. Theory of Probability and its Applications, 31(3):498–500, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Markus:1993:LRG

[Mar93] L. Markus. Likelihood ratio of Gaussian, globally Markovian random fields. Theory of Probability and its Applications, 37 (1):150–153, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 188–190.

Maslov:1960:SLT

[Mas60] K. V. Maslov. On some limit theorems of probability theory. Theory of Probability and its Applications, 5(1):50–79, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Masani:1972:HHS

[Mas72] P. Masani. On helixes in Hilbert space. I. Theory of Probability and its Applications, 17(1):1–19, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Masimov:1974:GBS

[Mas74a] V. M. Masimov. A generalized Bernoulli scheme and its limit distributions. Theory of Probability and its Applications, 18 (3):521–530, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maslova:1974:VNR

[Mas74b] N. B. Maslova. On the variance of the number of real roots of random polynomials. Theory of Probability and its Applications, 19(1):35–52, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maslova:1975:DNR

[Mas75] N. B. Maslova. On the distribution of the number of real roots of random polynomials. Theory of Probability and its Applications, 19(3):461–473, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Masol:1999:TLD

[Mas99] V. I. Masol. A theorem on the limiting distribution for the number of false solutions of a system of nonlinear random Boolean equations. Theory of Probability and its Applications, 43 (1):75–88, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97667.

Matveev:1960:SMS

[Mat60] R. F. Matveev. On singular multidimensional stationary processes. Theory of Probability and its Applications, 5(1):33–39, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matveev:1961:MDR

[Mat61] R. F. Matveev. On multidimensional regular stationary processes. Theory of Probability and its Applications, 6(2): 149–165, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matveev:1962:PFS

[Mat62] R. F. Matveev. On the problem of filtration of stationary processes. Theory of Probability and its Applications, 7(2): 211–213, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mateev:1965:ASS

[Mat65a] R. F. Mateev. On the asymptotic solution of a system of linear differential equations with small random coefficients. Theory of Probability and its Applications, 10(3):513–515, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matveev:1965:AAS

[Mat65b] R. F. Matveev. Addendum: On the asymptotic solution of a system of linear differential equations with small random coefficients. Theory of Probability and its Applications, 10(3):540, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mateev:1978:EMD

[Mat78] P. Mateev. On the entropy of the multinomial distribution. Theory of Probability and its Applications, 23(1):188–190, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matskyavichyus:1984:LBC

[Mat84]  V. K. Matskyavichyus.  A

lower bound for the convergence rate in the Central Limit Theorem. Theory of Probability and its Applications, 28(3):596–601, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matskyavichus:1985:ALB

[Mat85] V. K. Matskyavichus. Addendum: a lower bound for the convergence rate in the Central Limit Theorem. Theory of Probability and its Applications, 29 (1):196–197, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matskyavichyus:1986:ALB

[Mat86a] V. K. Matskyavichyus. Addendum: a lower bound for the convergence rate in the Central Limit Theorem. Theory of Probability and its Applications, 30 (4):873–874, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matskyavichyus:1986:LBC

[Mat86b] V. K. Matskyavichyus. On a lower bound for the convergence rate in a local limit theorem. Theory of Probability and its Applications, 30(4):810–814, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(4), (1985), pp. 763–766.

Mazo:1969:FTE

[Maz69] V. L. Mazo. A fine topology on the entrance boundary. Theory of Probability and its Applications, 14(1):132–135, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Morozova:1968:PLR

[MC68] E. A. Morozova and N. N. Chentsov. P. Lévy’s random fields. Theory of Probability and its Applications, 13(1):153–156, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1984:STN

[MC84] D. Kh. Mushtari and A. N. Chuprunov. Sufficient topologies and norms. Theory of Probability and its Applications, 28 (4):736–751, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

McDonald:1980:LLT

[McD80] David R. McDonald. On local limit theorem for integer-valued random variables. Theory of Probability and its Applications, 24(3):613–619, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

McKean:1963:BMS

[McK63] H. P. McKean, Jr. Brownian motion with a several-dimensional time. Theory of Probability and its Applications, 8(4):335–354, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

McKendrick:1993:ETD

[McK93] I. J. McKendrick. An exploratory technique to detect departures from complete spatial randomness. Theory of Probability and its Applications, 37(1):153–155, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Medvedev:1977:SSP

[Med77]  Yu. I. Medvedev.  Separa-

ble statistics in a polynomial scheme. I. Theory of Probability and its Applications, 22(1):1–15, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Medvedev:1978:SSP

[Med78] Yu. I. Medvedev. Separable statistics in a polynomial scheme. II. Theory of Probability and its Applications, 22(3):607–615, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mednitskii:1998:ESS

[Med98] Yu. V. Mednitskii. On the existence of strong solutions of linear stochastic differential equations in R^{∞}. Theory of Probability and its Applications, 42(4): 702–706, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97656.

Megrabyan:1986:SSF

[Meg86] R. M. Megrabyan. On the set of sums of function series in the spaces L_{Φ}. Theory of Probability and its Applications, 30 (3):542–556, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Melnikov:1979:SSS

[Mel79] A. V. Mel’nikov. On strong solutions of stochastic differential equations with nonsmooth coefficients. Theory of Probability and its Applications, 24(1):147–150, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Melnikov:1983:MLS

[Mel83] A. V. Mel’nikov. Martingalelike stochastic sequences and processes. Theory of Probability and its Applications, 27(3):587–591, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Melnikov:1989:DPS

[Mel89] A. V. Mel′nikov. The disorder problem for semimartingales. Theory of Probability and its Applications, 33(3):578–582, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 621–625.

Menshikov:1988:QER

[Men88] M. V. Men’shikov. Quantitative estimates and rigorous inequalities for critical points of a graph and its subgraphs. Theory of Probability and its Applications, 32(3):544–547, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Meshalkin:1958:LTM

[Mes58] L. D. Meshalkin. Limit theorems for Markov chains with a finite number of states. Theory of Probability and its Applications, 3(4):335–357, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Meshalkin:1960:APD

[Mes60] L. D. Meshalkin. On the approximation of polynomial distributions by infinitely divisible laws. Theory of Probability and its Applications, 5(1):106–114, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Meshalkin:1961:ADF

[Mes61] L. D. Meshalkin. On approximations of distribution functions of sums by infinitely divisible laws. Theory of Probability and its Applications, 6 (3):233–252, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Meshalkin:1963:GST

[Mes63] L. D. Meshalkin. Generalization of Sperner’s theorem on the number of subsets of a finite set. Theory of Probability and its Applications, 8(2):203–204, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Metivier:1974:SIR

[Met74] Michel Metivier. The stochastic integral with respect to processes with values in a reflexive Banach space. Theory of Probability and its Applications, 19(??):1, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 19, (1974), pp. 787– 816.

Metivier:1975:SIR

[Met75] Michel Metivier. The stochastic integral with respect to processes with values in a reflexive Banach space. Theory of Probability and its Applications, 19 (4):758–787, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Meyer:1963:SMP

[Mey63] P. A. Meyer. The strong Markov property of multiplicative functions. Theory of Probability and its Applications, 8(3):328–334, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyutov:1971:WPR

[MF71] M. B. Malyutov and V. V. Fedorov. On weighted polynomial regression designs with minimum average variance. Theory of Probability and its Applications, 16(4):716–720, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Moskvin:1974:STS

[MFY74] D. A. Moskvin, G. A. Frejman, and A. A. Yudin. Structure theory of set addition and local limit theorems for independent lattice random variables. Theory of Probability and its Applications, 19(1):53–64, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mihoc:1956:LLR

[Mih56] G. Mihoc. On limit laws for random vectors connected in a Markov chain. Theory of Probability and its Applications, 1 (1):92–100, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mihailovskaya:1981:OSC

[Mih81] L. G. Mihailovskaya. On the optimal stopping of a controlled diffusion process. Theory of Probability and its Applications, 25 (2):299–308, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhalevish:1956:SBS

[Mik56] V. S. Mikhalevish. Sequential Bayes solutions and optimal methods of statistical acceptance control. Theory of Probability and its Applications, 1 (4):395–421, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhalevich:1958:SSB

[Mik58] V. S. Mikhalevich. Sequential selection between two solutions for a Poisson process. Theory of Probability and its Applications, 3(4):430–434, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1965:MRV

[Mik65] G. A. Mikhailov. On modelling random variables for one class of distribution laws. Theory of Probability and its Applications, 10(4):681–682, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1970:CMC

[Mik70a] G. A. Mikhailov. On a class of Monte Carlo estimators. Theory of Probability and its Applications, 15(1):137–138, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miklukhin:1970:LTA

[Mik70b] O. G. Miklukhin. Limit theorems for an asymmetric random walk. Theory of Probability and its Applications, 15(4):710–715, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1974:LDR

[Mik74] V. G. Mikhailov. Limit distribution of random variables associated with multiple long duplications in a sequence of independent trials. Theory of Probability and its Applications, 19 (1):180–184, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1975:RMS

[Mik75a] G. A. Mikhailov. On the reproduction method for simulating random vectors and stochastic processes (randomization of correlation matrices). Theory of Probability and its Applications, 19(4):839–843, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1975:AIV

[Mik75b] V. G. Mikhailov. Asymptotic independence of vector components of multivariate extreme order statistics. Theory of Probability and its Applications, 19 (4):817–821, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1975:CPP

[Mik75c] V. G. Mikhailov. Convergence to the Poisson process in a scheme of cumulative sums of dependent random variables. Theory of Probability and its Applications, 19(2):403–407, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1976:CLT

[Mik76] V. G. Mikhailov. A Central Limit Theorem for the number of partial long repetitions. Theory of Probability and its Applications, 20(4):862–866, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1977:PLT

[Mik77] V. G. Mikhailov. A Poisson limit theorem in the scheme of group disposal paricles. Theory of Probability and its Applications, 22 (1):152–156, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1978:ERC

[Mik78] V. G. Mikhailov. An estimate of the rate of convergence to the Poisson distribution in group allocation of particles. Theory of Probability and its Applications, 22(3):554–562, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:1980:ANN

[Mik80] V. G. Mikhailov. Asymptotic normality of the number of empty cells for group allocation of particles. Theory of Probability and its Applications, 25 (1):82–90, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikosch:1989:LIL

[Mik89] T. Mikosch. On the law of the iterated logarithm for quadratic forms in independent Gaussian variables. Theory of Probability and its Applications, 33(1): 144–148, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikosch:1991:LBL

[Mik91] T. Mikosch. On a lower bound in the law of the iterated logarithm for Gaussian quadratic forms. Theory of Probability and its Applications, 35(2):363–367, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 387–391.

Mikhailov:1992:TJ

[Mik92] V. G. Mikhaĭlov. On a theorem of Janson. Theory of Probability and its Applications, 36 (1):173–176, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 168–170.

Mikhailov:1994:RCL

[Mik94] V. G. Mikhaĭlov. On a refinement of the Central Limit Theorem for sums of independent random indicators. Theory of Probability and its Applications, 38 (3):479–489, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 540–552.

Mikhailov:1996:RSR

[Mik96a] V. G. Mikhaĭlov. Repetition of states of a random-number generator under multiple access. Theory of Probability and its Applications, 40(4):679–689, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 786–797.

Mikosch:1996:SLL

[Mik96b] T. Mikosch. On the strong law of large numbers for random quadratic forms. Theory of Probability and its Applications, 40(1):76–91, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 125–142.

Mikhailov:1997:LTR

[Mik97] V. G. Mikhaĭlov. Limit theorems for a random covering of a finite set and for the number of solutions of a system of random equations. Theory of Probability and its Applications, 41(2):265–274, ???? 1997. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97547. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 272–283.

Mikhailov:1998:SCP

[Mik98] V. G. Mikhailov. Short communication: On a Poisson approximation for the distribution of the number of empty cells in a nonhomogeneous allocation scheme. Theory of Probability and its Applications, 42(1): 174–179, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97601.

Mikhailov:1999:LTN

[Mik99] V. G. Mikhailov. Limit theorems for the number of nonzero solutions of a system of random equations over GF(2). Theory of Probability and its Applications, 43(3):480–487, September 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97708.

Milshtein:1972:IMP

[Mil72] G. N. Mil’shtein. Interaction of Markov processes. Theory of Probability and its Applications, 17(1):34–43, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtejn:1975:AIS

[Mil75] G. N. Mil’shtejn. Approximate integration of stochastic differential equations. Theory of Probability and its Applications, 19 (3):557–562, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtein:1979:MSO

[Mil79a] G. N. Mil’shtein. A method of second-order accuracy integration of stochastic differential equations. Theory of Probability and its Applications, 23 (2):396–401, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtejn:1979:PTS

[Mil79b]  G. N. Mil’shtejn.  On the

probability-theoretic solution of linear systems of elliptic and parabolic equations. Theory of Probability and its Applications, 23(4):820–824, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtein:1982:QLF

[Mil82] G. N. Mil’shtein. Quadratic Lyapunov functionals and second moments for stochastic systems with after-effect. Theory of Probability and its Applications, 26 (4):721–731, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtein:1986:WAS

[Mil86] G. N. Mil’shtein. Weak approximation of solutions of systems of stochastic differential equations. Theory of Probability and its Applications, 30(4):750–766, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtein:1988:TOC

[Mil88] G. N. Mil’shtein. A theorem on the order of convergence of meansquare approximations of solutions of systems of stochastic differential equations. Theory of Probability and its Applications, 32(4):738–741, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milshtein:1996:SFB

[Mil96] G. N. Mil′shteĭn. Solution of the first boundary value problem for parabolic equations by integration of stochastic differential equations. Theory of Probability and its Applications, 40 (3):556–563, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 657–665.

Milshtein:1997:ANI

[Mil97] G. N. Mil′shteĭn. Application of the numerical integration of stochastic equations for the solution of boundary value problems with Neumann boundary conditions. Theory of Probability and its Applications, 41(1):170–177, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 210–218.

Mindlin:1991:RCC

[Min91] D. S. Mindlin. The rate of convergence of convolutions of random measures on a compact group. Theory of Probability and its Applications, 35(2):368–373, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 361–367.

Miroshin:1969:AEP

[Mir69] R. N. Miroshin. An asymptotic estimate for the probability that a Gaussian stochastic process remains below the line kt + a. Theory of Probability and its Applications, 14(2):349–354, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1974:SCF

[Mir74] R. N. Miroshin. A sufficient condition for finiteness of moments of the number of zeros of a differentiable stationary Gaussian process. Theory of Probability and its Applications, 18(3):454–463, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1975:NCF

[Mir75] R. N. Miroshin. A necessary condition for finiteness of the moments of the number of zeros of a differentiable Gaussian stationary process. Theory of Probability and its Applications, 19 (3):570–577, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshnichenko:1976:OSI

[Mir76] T. P. Miroshnichenko. Optimal stopping of the integral of a Wiener process. Theory of Probability and its Applications, 20 (2):387–391, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1977:CRL

[Mir77] R. N. Miroshin. Convergence of Rice Longuet-Higgins series for a Wong process. Theory of Probability and its Applications, 21 (4):863–866, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1978:CFM

[Mir78a] R. N. Miroshin. Conditions for finiteness of moments of the number of zeros of stationary Gaussian processes. Theory of Probability and its Applications, 22 (3):615–625, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1978:CLN

[Mir78b] R. N. Miroshin. Conditions of local nondeterminism of differentiable Gaussian stationary processes. Theory of Probability and its Applications, 22(4):831–836, ???? 1978. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshnichenko:1978:STT

[Mir78c] T. P. Miroshnichenko. A sequential test of two simple hypotheses about the mean of a Wiener process with delayed observations. Theory of Probability and its Applications, 23 (1):195–201, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirvaliev:1979:ROO

[Mir79a] M. Mirvaliev. The rejection of outlying observations in regression analysis. Theory of Probability and its Applications, 23 (3):598–602, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirvaliev:1979:ROR

[Mir79b] M. Mirvaliev. Rejection of outlying results of angular measurements. Theory of Probability and its Applications, 23(4):814–819, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1980:SOM

[Mir80a] R. N. Miroshin. Second-order Markov and reciprocal stationary Gaussian processes. Theory of Probability and its Applications, 24(4):845–852, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshnichenko:1980:TTS

[Mir80b] T. P. Miroshnichenko. Testing of two simple hypotheses in the presence of delayed observations. Theory of Probability and its Applications, 24(3):467–479, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1981:CLH

[Mir81] R. N. Miroshin. Convergence of Longuet-Higgins series for stationary Gaussian Markov processes of first order. Theory of Probability and its Applications, 26(1):97–117, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirakhmedov:1984:AES

[Mir84a] Sh. A. Mirakhmedov. An asymptotic expansion for a sample sum from a finite population. Theory of Probability and its Applications, 28(3):492–502, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1984:URS

[Mir84b] R. N. Miroshin. The use of Rice series. Theory of Probability and its Applications, 28(4):714–726, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirvaliev:1984:SRA

[Mir84c] M. Mirvaliev. Smoothing of the results of angular measurements made according to Tomilin’s scheme. Theory of Probability and its Applications, 28(2): 455–462, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshin:1985:SCF

[Mir85] R. N. Miroshin. A simple criterion for finiteness of the moments of the number of zeros of a Gaussian stationary process. Theory of Probability and its Applications, 29(3):566–569, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirakhmedov:1986:ECD

[Mir86a] Sh. A. Mirakhmedov. Estimates of closeness of the distribution of a randomized separable statistic to the normal law in a multinomial scheme. Theory of Probability and its Applications, 30 (1):192–196, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirakhmedov:1986:EPN

[Mir86b] Sh. A. Mirakhmedov. Estimates of proximity to the normal distribution in sampling without replacement. Theory of Probability and its Applications, 30 (3):451–464, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mirakhmedov:1988:ADM

[Mir88] Sh. A. Mirakhmedov. Approximation of the distribution of multidimensional randomized divisible statistics by normal distributions (multinomial scheme). Theory of Probability and its Applications, 32(4):696–706, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshnikov:1990:BML

[Mir90a] A. L. Miroshnikov. Bounds for the multidimensional Lévy concentration function. Theory of Probability and its Applications, 34(3):535–540, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 593–598.

Mirvaliev:1990:CSG

[Mir90b] M. Mirvaliev. Chi-square goodness-of-fit tests for a family of multidimensional discrete distributions. Theory of Probability and its Applications, 34 (4):728–732, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 794–799.

Mishura:1989:EFD

[Mis89] Yu. S. Mishura. Exponential formulas and Doléans’ equation for discontinuous two-parameter processes. Theory of Probability and its Applications, 33(2): 388–392, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mishura:1991:MCD

[Mis91] Yu. S. Mishura. A martingale characterization of diffusion random fields on the plane. Theory of Probability and its Applications, 35(1):152–157, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 143–147.

Mitalauskas:1962:LLT

[Mit62] A. A. Mitalauskas. Local limit theorems for stable limit distributions. Theory of Probability and its Applications, 7(2): 180–185, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mitrofanova:1967:AEM

[Mit67] N. M. Mitrofanova. An asymptotic expansion for the maximum likelihood estimate of a vector parameter. Theory of Probability and its Applications, 12 (3):364–372, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mityushin:1988:MCM

[Mit88] L. G. Mityushin. On mixing in complex Markov chains with variable memory depth. Theory of Probability and its Applications, 32(4):678–685, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mkrtchyan:1981:SCD

[Mkr81] S. T. Mkrtchyan. Stability of a characterization of distributions, and some wide-sense characteristic properties of the normal distribution. Theory of Probability and its Applications, 26 (1):161–167, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martin-Lof:1965:CTL

[ML65] P. Martin-Löf. The continuity theorem on a locally compact group. Theory of Probability and its Applications, 10(2):338–341, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martin-Lof:1966:CRS

[ML66] P. Martin-Löf. On the concept of a random sequence. Theory of Probability and its Applications, 11(1):177–179, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Migai:1977:LTF

[MN77] N. I. Migai and V. B. Nevzorov. Limit theorems for the first passage time of a certain level. Theory of Probability and its Applications, 21(2):406–410, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Melnikov:1989:SIP

[MN89] A. V. Mel′nikov and A. A.

Novikov. Sequential inferences with prescribed accuracy for semimartingales. Theory of Probability and its Applications, 33 (3):446–459, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 480–494.

Melnikov:1999:MVH

[MN99] A. V. Melnikov and M. L. Nechaev. On the mean-variance hedging problem. Theory of Probability and its Applications, 43(4):588–603, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97713.

Mnatsakanov:1986:FLT

[Mna86] R. M. Mnatsakanov. A functional limit theorem for additively separable statistics in the case of very rare events. Theory of Probability and its Applications, 30 (3):622–626, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mnatsakanov:1988:CDS

[Mna88a] R. M. Mnatsakanov. On the convergence of divisible statistics to the Poisson process. Theory of Probability and its Applications, 32(2):361–365, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mnatsakanov:1988:CSS

[Mna88b] R. M. Mnatsakanov. On the convergence of separable statistics to a Wiener process. Theory of Probability and its Applications, 32(1):152–157, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1969:SSP

[MO69] S. A. Molchanov and E. Ostrovskii. Symmetric stable processes as traces of degenerate diffusion processes. Theory of Probability and its Applications, 14 (1):128–131, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maltsev:1983:CLT

[MO83] V. V. Mal’tsev and E. I. Ostrovskii. Central Limit Theorem for stationary processes in Hilbert space. Theory of Probability and its Applications, 27 (2):357–359, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matukhin:1991:NDE

[MO91] P. G. Matukhin and E. I. Ostrovskiĭ. Nonparametric density estimation by results of multifactor testing. Theory of Probability and its Applications, 35 (1):75–86, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 72–82.

Mogulskii:1974:AEM

[Mog74] A. A. Mogul’skii. Absolute estimates for moments of certain boundary functionals. Theory of Probability and its Applications, 18(2):340–347, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mogulskii:1975:SDS

[Mog75] A. A. Mogul’skii. Small deviations in a space of trajectories. Theory of Probability and its Applications, 19(4):726–736, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mogulskii:1977:LDT

[Mog77a] A. A. Mogul’skii. Large deviations for trajectories of multidimensional random walks. Theory of Probability and its Applications, 21(2):300–315, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mogulskii:1977:SFJ

[Mog77b] A. A. Mogul’skii. On the size of the first jump for a process with independent increments. Theory of Probability and its Applications, 21(3):470–481, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mogulskii:1977:RLD

[Mog77c] A. A. Mogul’skii. Remarks on large deviations for the ω^{2} statistics. Theory of Probability and its Applications, 22(1):166–171, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mogulskii:1980:LIL

[Mog80]  A. A. Mogul’skii.  On the

law of the iterated logarithm in Chung’s form for functional spaces. Theory of Probability and its Applications, 24(2):405–413, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mokkadem:1991:SRK

[Mok91] A. Mokkadem. Study of risks of kernel estimators. Theory of Probability and its Applications, 35(3):478–486, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 531–538.

Molchanov:1964:PTD

[Mol64] S. A. Molchanov. On a problem in the theory of diffusion processes. Theory of Probability and its Applications, 9(3): 472–477, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchan:1967:SPC

[Mol67a] G. M. Molchan. On some problems concerning Brownian motion in Lévy’s sense. Theory of Probability and its Applications, 12(4):682–690, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1967:MBI

[Mol67b] S. A. Molchanov. Martin boundaries for invariant Markov processes on a solvable group. Theory of Probability and its Applications, 12(2):310–314, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1967:MBD

[Mol67c] S. A. Molchanov. On Martin boundaries for the direct product of Markov chains. Theory of Probability and its Applications, 12(2):307–310, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1968:SFP

[Mol68] S. A. Molchanov. Strong Feller property of diffusion processes on smooth manifolds. Theory of Probability and its Applications, 13(3):471–475, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1988:ULL

[Mol88] I. S. Molchanov. Uniform laws of large numbers for empirical associated functionals of random closed sets. Theory of Probability and its Applications, 32 (3):556–559, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1991:EED

[Mol91] I. S. Molchanov. Empirical estimation of distribution quantiles of random closed sets. Theory of Probability and its Applications, 35(3):594–600, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 586–592.

Molchanov:1992:CEP

[Mol92] I. S. Molchanov. A consistent estimate of parameters of Boolean models of random closed sets. Theory of Probability and its Applications, 36(3):600– 607, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Molchanov:1994:LTU

[Mol94] I. S. Molchanov. Limit theorems for unions of random sets under multiplicative normalization. Theory of Probability and its Applications, 38(3):541–547, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 638–645.

Molchanov:1996:CRP

[Mol96] I. S. Molchanov. On the convergence of random processes generated by polyhedral approximation of convex compacts. Theory of Probability and its Applications, 40(2):383–390, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Monin:1956:SIS

[Mon56] A. S. Monin. A statistical interpretation of the scattering of microscopic particles. Theory of Probability and its Applications, 1(3):298–311, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Monin:1958:SAT

[Mon58] A. S. Monin. The structure of atmospheric turbulence. Theory of Probability and its Applications, 3(3):266–296, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Morozenskii:1975:ALC

[Mor75] L. Yu. Morozenskii. On the asymptotic length of a commercial Traveller’s path when towns are randomly allocated. Theory of Probability and its Applications, 19(4):798–801, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mori:1993:RWB

[Mór93] T. F. Móri. Random walks on de Bruijn graphs. Theory of Probability and its Applications, 37 (1):158–160, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 194–197.

Moskvin:1970:ABL

[Mos70] D. A. Moskvin. On the asymptotic behavior of large deviation probabilities for the sums Σf(x2^{n}). Theory of Probability and its Applications, 15(2): 229–240, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Moskvin:1973:LLT

[Mos73] D. A. Moskvin. A local limit theorem for large deviations in the case of differently distributed lattice summands. Theory of Probability and its Applications, 17 (4):678–684, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 17, (1972), pp. 716–722.

Mosyagin:1991:RCM

[Mos91] V. E. Mosyagin. On the rate of convergence of a maximum likelihood process: The case of discontinuous density. Theory of Probability and its Applications, 35(3):600–602, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 607–609.

Movshovich:1977:SRP

[Mov77] S. M. Movshovich. On the stability of a random price regulation process. Theory of Probability and its Applications, 21 (2):430–435, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maslov:1958:IOC

[MP58] K. V. Maslov and A. Ya. Povzner.

The infinitesimal operators of a class of Markov processes. Theory of Probability and its Applications, 3(1):66–78, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyshev:1970:OHC

[MP70] V. A. Malyshev and A. I. Pinskii. Optimality of homogeneous control in a problem of reliability theory. Theory of Probability and its Applications, 15(1):142–146, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1977:NSC

[MP77] A. I. Martikainen and V. V. Petrov. On necessary and sufficient conditions for the law of the iterated logarithm. Theory of Probability and its Applications, 22(1):16–23, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1978:ANS

[MP78a] A. I. Martikainen and V. V. Petrov. Addendum: On necessary and sufficient conditions for the law of the iterated logarithm. Theory of Probability and its Applications, 22(2):430, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mogulskii:1978:FET

[MP78b] A. A. Mogul’skii and E. A. Pecerskii. On the first exit time from a semigroup in R^{m} for a random walk. Theory of Probability and its Applications, 22(4):818–825, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Moskvin:1979:LLT

[MP79] D. A. Moskvin and A. G. Postnikov. A local limit theorem for the distribution of fractional parts of an exponential function. Theory of Probability and its Applications, 23(3):521–528, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Martikainen:1980:TF

[MP80] A. I. Martikainen and V. V.

Petrov. On a theorem of Feller. Theory of Probability and its Applications, 25(1):191–193, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Monrad:1991:PEV

[MP91] D. Monrad and W. Philipp. The problem of embedding vectorvalued martingales in a Gaussian process. Theory of Probability and its Applications, 35 (2):374–377, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 384–387.

Matsak:1992:SMB

[MP92] I. K. Matsak and A. N. Plichko.

On the support of a measure in a Banach space and finite representability. Theory of Probability and its Applications, 36 (2):381–385, May 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 363–367.

Martikainen:1993:SLT

[MP93a] A. I. Martikaĭnen and V. V. Petrov. On strong limit theorems for sums of independent random variables. Theory of Probability and its Applications, 37(1):80–84, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 81–86.

Menendez:1993:UEM

[MP93b] M. L. Menéndez and J. A. Pardo. Unified (r, s)-entropy measures in a Bayesian context. Theory of Probability and its Applications, 37(1):85–88, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 87–90.

Mikhaleva:1997:DMG

[MP97] T. L. Mikhaleva and V. I. Piterbarg. On the distribution of the maximum of a Gaussian field with constant variance on a smooth manifold. Theory of Probability and its Applications, 41(2):367–379, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 438–451.

Moskvin:1970:AMO

[MPY70] D. A. Moskvin, L. P. Postnikova, and A. A. Yudin. An arithmetic method for obtaining local limit theorems for lattice random variables. Theory of Probability and its Applications, 15 (1):87–97, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshnikov:1980:ICF

[MR80] A. L. Miroshnikov and B. A. Rogozin. Inequalities for the concentration function. Theory of Probability and its Applications, 25(1):176–180, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miroshnikov:1983:RIC

[MR83a] A. L. Miroshnikov and B. A. Rogozin. Remarks on an inequality for the concentration function of sums of independent variables. Theory of Probability and its Applications, 27(4):848–850, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1983:LTR

[MR83b] S. A. Molchanov and A. Ya.

Reznikova. Limit theorems for random partitions. Theory of Probability and its Applications, 27(2):310–323, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mindlin:1989:CRM

[MR89] D. S. Mindlin and B. A. Rubshteĭn. Convolutions of random measures on a compact group. Theory of Probability and its Applications, 33(2):355–357, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 376–379.

Mohan:1993:MDA

[MR93] N. R. Mohan and S. Ravi. Max domains of attraction of univariate and multivariate p-max stable laws. Theory of Probability and its Applications, 37(4): 632–643, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 709–721.

Mattner:1996:OUE

[MR96] L. Mattner and M. Reinders. Optimal unbiased estimators in additive models with bounded errors are deterministic. Theory of Probability and its Applications, 40(4):772–777, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 934–938.

Mori:1994:MSK

[MRS94] T. F. Móri, V. K. Rohatgi, and G. J. Székely. On multivariate skewness and kurtosis. Theory of Probability and its Applications, 38(3):547–551, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 675–679.

Meshalkin:1979:ECM

[MS79] L. D. Meshalkin and V. I. Serdobol’skii. Errors in the classification of multi-variate observations. Theory of Probability and its Applications, 23(4):741–750, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Muminov:1986:ISR

[MS86] M. Muminov and E. G. Samandarov. Inequalities for some random walk processes. Theory of Probability and its Applications, 30(3):489–495, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Menshikov:1988:CCP

[MS88] M. V. Men’shikov and A. F.

Sidorenko. The coincidence of critical points in Poisson percolation models. Theory of Probability and its Applications, 32 (3):547–550, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyshev:1990:SRE

[MS90] V. A. Malyshev and Ya. G. Sinai.

Some research in ergodic theory and mathematical problems of statistical mechanics in the Moscow University department of probability theory. Theory of Probability and its Applications, 34(1):186–193, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Michaletzky:1999:CFJ

[MSV99] Gy. Michaletzky, L. Szeidl, and P. Varlaki. On the characteristic function of the joint limit distribution of the first- and secondorder power sums. Theory of Probability and its Applications, 43(1):126–134, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97672.

Malyshev:1983:QSL

[MT83] V. A. Malyshev and I. P. Tsaregradskii. A queueing system with local interaction. Theory of Probability and its Applications, 27 (3):619–624, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:1985:LMH

[MT85] S. A. Molchanov and V. N. Tutubalin. A linear model of a hydromagnetic dynamo and products of random matrices. Theory of Probability and its Applications, 29(2):231–246, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyutov:1990:SRM

[MT90] M. B. Malyutov and Yu. N. Tyurin. Some research in mathematical statistics by the department of probability theory of Moscow University. Theory of Probability and its Applications, 34(1):193–197, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Morales:1993:HID

[MTP93] D. Morales, I. J. Taneja, and L. Pardo. Hypoentropy as an index of diversity. Theory of Probability and its Applications, 37 (1):155–158, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 191–194.

Molchanov:1991:APP

[MTV91] S. A. Molchanov, V. N. Tutubalin, and A. D. Ventzel’. Asymptotic problems in probability theory and the theory of random media. Theory of Probability and its Applications, 35 (1):87–93, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 27–34.

Malkov:1983:NUE

[MU83] A. S. Mal’kov and V. V.

Ul’yanov. On a non-uniform estimate for the rate of convergence in a local limit theorem for a stable limit distribution. Theory of Probability and its Applications, 27(3):607–609, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Muchnik:1988:LLF

[Muc88] An. A. Muchnik. Lower limits of frequencies in computable sequences and relativized a priori probability. Theory of Probability and its Applications, 32 (3):513–514, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mukhin:1977:WFL

[Muk77] A. B. Mukhin. On a weak form of a local limit theorem. Theory of Probability and its Applications, 21(3):632–636, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mukantseva:1978:TNO

[Muk78] L. A. Mukantseva. Testing normality in one-dimensional and multi-dimensional linear regression. Theory of Probability and its Applications, 22(3):591–602, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mukhin:1985:LLT

[Muk85] A. B. Mukhin. Local limit theorems for distributions of sums of independent random vectors. Theory of Probability and its Applications, 29(2):369–375, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mukhin:1990:LPS

[Muk90] A. B. Mukhin. Local probabilities for sums of independent random variables. Theory of Probability and its Applications, 34 (4):616–624, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 677–685.

Mukhin:1992:LLT

[Muk92] A. B. Mukhin. Local limit theorems for lattice random variables. Theory of Probability and its Applications, 36(4):698– 713, December 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 660–674.

Mukhin:1996:RBL

[Muk96] A. B. Mukhin. A relationship between local and integral limit theorems. Theory of Probability and its Applications, 40 (1):92–103, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 96–110.

Mukhin:1998:ALP

[Muk98] A. B. Mukhin. Approximation of local probabilities of sums of independent random variables. Theory of Probability and its Applications, 42(4): 633–645, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97646.

Mushtari:1969:PLS

[Mus69] D. Kh. Mushtari. On a problem of Laurent Schwartz and on the realization of l_{p} spaces by spaces of random variables. Theory of Probability and its Applications, 14(4):699–701, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mustari:1970:ASC

[Mus70] D. Kh. Mustari. On almost sure convergence in linear spaces of random variables. Theory of Probability and its Applications, 15(2):337–342, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1973:CGQ

[Mus73] D. Kh. Mushtari. Certain general questions of the theory of probability measures in linear spaces. Theory of Probability and its Applications, 18 (1):64–75, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1977:PCH

[Mus77] D. Kh Mushtari. On a probabilistic characterization of Hilbert space. Theory of Probability and its Applications, 21(2):410–412, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1980:LTC

[Mus80a] D. Kh. Mushtari. Lévy type criteria for weak convergence of probabilities in Fréchet spaces. Theory of Probability and its Applications, 24(3):587–592, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1980:SC

[Mus80b] D. Kh. Mushtari. Spaces of cotype p(0 ≤ p ≤ 2). Theory of Probability and its Applications, 25(1):105–117, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1981:RNS

[Mus81] D. Kh. Mushtari. A remark on necessary and sufficient topologies. Theory of Probability and its Applications, 25(3):592–593, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mushtari:1985:RCF

[Mus85] D. Kh. Mushtari. Remarks on compact families of probability measures in Banach spaces. Theory of Probability and its Applications, 29(2):407–410, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malevich:1989:CFM

[MV89] T. L. Malevich and L. N. Volodina. Conditions for the finiteness of the moments of the number of zeros of a Gaussian vector field. Theory of Probability and its Applications, 33 (1):50–61, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 55–67.

Malevich:1994:SFC

[MV94] T. L. Malevich and L. N. Volodina. Some finiteness conditions for factorial moments of the number of zeros of Gaussian field zeros. Theory of Probability and its Applications, 38 (1):27–45, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 49–70.

Masani:1959:BSP

[MW59] P. Masani and N. Wiener. On bivariate stationary processes and the factorization of matrixvalued functions. Theory of Probability and its Applications, 4 (3):300–308, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mishura:1999:ADI

[MW99] Yu. S. Mishura and F. Weisz. Atomic decompositions and inequalities for vector-valued discrete-time martingales. Theory of Probability and its Applications, 43(3):487–496, September 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97707.

Myers:1993:CVV

[Mye93] D. Myers. Cross-validation and variogram estimation. Theory of Probability and its Applications, 37(2):345–347, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1964:ER

[Nad64a] E. A. Nadaraya. On estimating regression. Theory of Probability and its Applications, 9 (1):141–142, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1964:SNE

[Nad64b] E. A. Nadaraya. Some new estimates for distribution functions. Theory of Probability and its Applications, 9(3):497–500, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1965:NPE

[Nad65]  É. A. Nadaraya.  On non-

parametric estimates of density functions and regression curves. Theory of Probability and its Applications, 10(1):186–190, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1970:RNP

[Nad70] É A. Nadaraya. Remarks on nonparametric estimates for density functions and regression curves. Theory of Probability and its Applications, 15(1):134–137, ???? 1970. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraja:1973:CNP

[Nad73a] E. A. Nadaraja. On convergence in the L_{2}-norm of probability density estimates. Theory of Probability and its Applications, 18(??):1, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 18, (1973), pp. 853–856.

Nadkarni:1973:NPR

[Nad73b] M. G. Nadkarni. Notes on problems relating to vector-valued weakly stationary stochastic processes. Theory of Probability and its Applications, 17(3):517–521, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1974:CNP

[Nad74a] E. A. Nadaraya. On convergence in the L_{2}-norm of probability density estimates. Theory of Probability and its Applications, 18(4):808–811, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1974:IMS

[Nad74b] E. A. Nadaraya. On the integral mean square error of some nonparametric estimates for the density function. Theory of Probability and its Applications, 19 (1):133–141, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1977:QMD

[Nad77] E. A. Nadaraya. On a quadratic measure of deviation of the projection estimate of a distribution density. Theory of Probability and its Applications, 21 (4):843–850, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:1993:LDQ

[Nad93] È. A. Nadaraya. On limit distribution for quadratic deviation of generalized kernel estimates of a density. Theory of Probability and its Applications, 37 (2):383–392, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 413–421.

Nagaev:1957:SLT

[Nag57]  S. V. Nagaev.  Some limit

theorems for stationary Markov chains. Theory of Probability and its Applications, 2(4):378–406, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1961:MES

[Nag61] S. V. Nagaev. More exact statement of limit theorems for homogeneous Markov chains. Theory of Probability and its Applications, 6(1):62–81, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1965:SLT

[Nag65] S. V. Nagaev. Some limit theorems for large deviations. Theory of Probability and its Applications, 10(2):214–235, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1967:EEN

[Nag67] A. V. Nagaev. On estimating the expected number of direct descendants of a particle in a branching process. Theory of Probability and its Applications, 12(2):314–320, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1968:ERC

[Nag68a] S. V. Nagaev. An estimation of the rate of convergence for absorption probability in case of zero expectation. Theory of Probability and its Applications, 13 (1):160–164, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1968:SRT

[Nag68b] S. V. Nagaev. Some renewal theorems. Theory of Probability and its Applications, 13(4):547–563, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1969:AIL

[Nag69a] A. V. Nagaev. Addendum: Integral limit theorems taking large deviations into account when Cramer’s condition does not hold. Theory of Probability and its Applications, 14(3):534, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1969:ILTa

[Nag69b] A. V. Nagaev. Integral limit theorems taking large deviations into account when Cramér’s condition does not hold. I. Theory of Probability and its Applications, 14(1):51–64, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1969:ILTb

[Nag69c] A. V. Nagaev. Integral limit theorems taking large deviations into account when Cramér’s condition does not hold. II. Theory of Probability and its Applications, 14(2):193–208, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1969:ASR

[Nag69d] S. V. Nagaev. Addendum: Some renewal theorems. Theory of Probability and its Applications, 14(4):726, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1970:AEM

[Nag70a] S. V. Nagaev. Asymptotic expansions for the maximum of sums of independent random variables. Theory of Probability and its Applications, 15(3): 514–515, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1970:SCBa

[Nag70b] S. V. Nagaev. On the speed of convergence in a boundary problem. I. Theory of Probability and its Applications, 15(2):163–186, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1970:SCBb

[Nag70c] S. V. Nagaev. On the speed of convergence in a boundary problem. II. Theory of Probability and its Applications, 15(3):403–429, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1970:SCD

[Nag70d] S. V. Nagaev. On the speed of convergence of the distribution of maximum sums of independent random variables. Theory of Probability and its Applications, 15(2):309–314, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1971:LDE

[Nag71a] A. V. Nagaev. Limit distribution of the extreme terms of a variational series under conditions of large deviations for the sample mean. Theory of Probability and its Applications, 16(1):126–140, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1971:ECR

[Nag71b] S. V. Nagaev. An estimate for the convergence rate for the absorption probability. Theory of Probability and its Applications, 16(1):147–154, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1973:NSC

[Nag73]  S. V. Nagaev.  On neces-

sary and sufficient conditions for the strong law of large numbers. Theory of Probability and its Applications, 17(4):573–581, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1975:LTB

[Nag75] S. V. Nagaev. A limit theorem for branching processes with immigration. Theory of Probability and its Applications, 20 (1):176–179, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1976:SLT

[Nag76a] A. V. Nagaev. Some limit theorems of renewal theory. Theory of Probability and its Applications, 20(2):323–336, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagai:1976:IPC

[Nag76b] T. Nagai. On an interpolation problem for a certain random field. Theory of Probability and its Applications, 20(2):418–419, ???? 1976. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1977:API

[Nag77] S. V. Nagaev. Addendum: Probability inequalities for sums of independent random variables. Theory of Probability and its Applications, 21(4):875, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1978:PSI

[Nag78] A. V. Nagaev. On a property of sums of independent random variables. Theory of Probability and its Applications, 22 (2):326–338, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1980:RT

[Nag80] A. V. Nagaev. Renewal theorems in R^{d}. Theory of Probability and its Applications, 24 (3):572–581, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1982:ABO

[Nag82] S. V. Nagaev. On the asymptotic behavior of one-sided large deviation probabilities. Theory of Probability and its Applications, 26(2):362–366, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1984:ASL

[Nag84a] A. V. Nagaev. Addendum: Some limit theorems of renewal theory. Theory of Probability and its Applications, 28(4):856–857, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1984:APL

[Nag84b] A. V. Nagaev. On the asymmetric problem of large deviations when the limit law is stable. Theory of Probability and its Applications, 28(4):670–680, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagav:1985:ASC

[Nag85] S. V. Nagav. Addendum: On the speed of convergence in a boundary problem. Theory of Probability and its Applications, 29 (1):197–198, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1986:MCM

[Nag86a] A. V. Nagaev. On a method of calculating moments of ladder heights. Theory of Probability and its Applications, 30 (3):569–572, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1986:RCN

[Nag86b] S. V. Nagaev. On the rate of convergence to normal law in Hilbert space. Theory of Probability and its Applications, 30 (1):19–37, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1987:BPN

[Nag87] A. V. Nagaev. On a boundary problem for a nonrecurrent random walk. Theory of Probability and its Applications, 31 (2):313–317, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1994:LDT

[Nag94] A. V. Nagaev. Large deviation theorems for the first time of crossing an increasing level in a transient random walk. Theory of Probability and its Applications, 38(1):46–52, March 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 71–78.

Nagaev:1998:PIS

[Nag98a] S. V. Nagaev. Probabilistic inequalities for sums of independent random variables in terms of truncated pseudomoments. Theory of Probability and its Applications, 42(3):520–528, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 615–623.

Nagaev:1998:SRP

[Nag98b] S. V. Nagaev. Some refinements of probabilistic and moment inequalities. Theory of Probability and its Applications, 42(4): 707–713, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97657.

Nagaev:1999:CLD

[Nag99] A. V. Nagaev. Cramer large deviations when the extreme conjugate distribution is heavytailed. Theory of Probability and its Applications, 43(3):405– 421, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97700.

Naimzhanov:1982:CLT

[Nai82] A. Naimzhanov. The Central Limit Theorem for the hierarchical Dyson model for small β. Theory of Probability and its Applications, 27(1):191–196, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nakhapetyan:1981:CWC

[Nak81] B. S. Nakhapetyan. On a criterion of weak convergence. Theory of Probability and its Applications, 25(2):370–377, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nakhapetyan:1988:APL

[Nak88] B. S. Nakhapetyan. An approach to proving limit theorems for dependent random variables. Theory of Probability and its Applications, 32(3):535–539, ???? 1988. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Nasr:1957:CBD

[Nas57] Saad K. Nasr. On the convergence, basic distributions and classification of abstract random variables. Theory of Probability and its Applications, 2(2): 172–180, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nasyrov:1988:DLT

[Nas88] F. S. Nasyrov. On the derivative of local time for the Brownian sheet with respect to a space variable. Theory of Probability and its Applications, 32(4):649–658, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nasyrov:1996:LTF

[Nas96a] F. S. Nasyrov. On local times for functions and stochastic processes I. Theory of Probability and its Applications, 40(4): 702–713, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 798–812.

Nasyrov:1996:RCF

[Nas96b] F. S. Nasyrov. On reflection of continuous functions and random processes having local times. Theory of Probability and its Applications, 40(3):563–567, ???? 1996. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 665–669.

Nasyrov:1997:LTF

[Nas97] F. S. Nasyrov. On local times for functions and stochastic processes. Theory of Probability and its Applications, 41(2):275–287, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97548. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 284–299.

Natarajan:1973:CTE

[Nat73] S. Natarajan. A category theorem in ergodic theory. Theory of Probability and its Applications, 17(2):355–356, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nawrotzki:1982:ETR

[Naw82] K. Nawrotzki. Ergodic theorem for random processes with embedded homogeneous flow. Theory of Probability and its Applications, 26(2):388–392, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1977:LLT

[NE77] S. V. Nagaev and M. S. Eppel’.

On a local limit theorem for the maximum of sums of independent random variables. Theory of Probability and its Applications, 21(2):384–385, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nemirovskii:1991:NCE

[Nem91] A. S. Nemirovskii. On necessary conditions for efficient estimation of functionals of a nonparametric signal in white noise. Theory of Probability and its Applications, 35(1):94–103, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Neumann:1965:SES

[Neu65] Peter Neumann. Spectral estimates of stationary random fields. Theory of Probability and its Applications, 10(1):128–132, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevelson:1964:BIM

[Nev64] M. B. Nevel’son. On the behavior of the invariant measure of a diffusion process with small diffusion on a circle. Theory of Probability and its Applications, 9(1):125–131, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevelson:1969:PRE

[Nev69] M. B. Nevel’son. On the problem of recurrence and ergodicity of a Gaussian Markov process. Theory of Probability and its Applications, 14(1):35–43, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevzorov:1971:SCD

[Nev71] V. B. Nevzorov. On the speed of convergence of the distribution of the maximum cumulative sum of independent random variables to the normal law. Theory of Probability and its Applications, 16(2):378–384, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevzorov:1972:ASC

[Nev72] V. B. Nevzorov. Addendum: On the speed of convergence of the distribution of the maximum cumulative sum of independent random variables to the normal law. Theory of Probability and its Applications, 17(1):193, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevelson:1973:CPC

[Nev73a] M. B. Nevel’son. On certain properties of continuous procedures of stochastic approximation. Theory of Probability and its Applications, 17(2):298–308, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevzorov:1973:JDR

[Nev73b] V. B. Nevzorov. Joint distributions of random variables connected with fluctuations of a stable process. Theory of Probability and its Applications, 18 (1):161–169, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevelson:1979:NAE

[Nev79] M. B. Nevel’son. A note on the asymptotic estimation of the extreme point of a regression function. Theory of Probability and its Applications, 23(2):365–370, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevelson:1981:AER

[Nev81] M. B. Nevel’son. On asymptotically efficient recursive estimation of a location parameter. Theory of Probability and its Applications, 25(3):569–579, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevelson:1986:PMS

[Nev86] M. B. Nevel’son. On properties of moments of stochastic approximation procedures. Theory of Probability and its Applications, 30(2):407–413, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevzorov:1988:DMT

[Nev88a] V. B. Nevzorov. The distribution of the maximum term in a sequence of sample means. Theory of Probability and its Applications, 32(1):117–121, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nevzorov:1988:R

[Nev88b] V. B. Nevzorov. Records. Theory of Probability and its Applications, 32(2):201–228, ???? 1988.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nhu:1968:SCC

[Nhu68] Hoang Huu Nhu. On the stability of certain characterizations of a normal population. Theory of Probability and its Applications, 13(2):299–304, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikolskii:1964:PMP

[Nik64] M. S. Nikol’skii. On the probability of a Markov point falling into a plane region with small diameter. Theory of Probability and its Applications, 9(2): 320–325, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikolskii:1966:PPV

[Nik66] M. S. Nikol’skii. On a problem posed by S. I. Vavilov. Theory of Probability and its Applications, 11(1):140–144, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikulin:1974:CST

[Nik74]  M. S. Nikulin.  Chi-square

test for continuous distributions with shift and scale parameters. Theory of Probability and its Applications, 18(3):559–568, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikulin:1975:QT

[Nik75] M. S. Nikulin. On a quantile test. Theory of Probability and its Applications, 19(2):410–413, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikolaev:1977:GBC

[Nik77] M. L. Nikolaev. On a generalization of the best choice problem. Theory of Probability and its Applications, 22(1):187–190, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikulin:1980:HTP

[Nik80a] M. S. Nikulin. Hypothesis testing for a parameter difference in binomial distributions. Theory of Probability and its Applications, 24(2):392–396, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikunen:1980:WCF

[Nik80b] M. Nikunen. On the weak compactness of a family of measures corresponding to continuous strong Markov processes. Theory of Probability and its Applications, 25(1):155–159, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikitin:1985:LAB

[Nik85] Ya. Yu. Nikitin. Local asymptotic Bahadur optimality and characterization problems. Theory of Probability and its Applications, 29(1):79–92, ???? 1985.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikulin:1987:TEE

[Nik87] V. N. Nikulin. On the theory of equivariant estimation on groups. Theory of Probability and its Applications, 31(4):687–691, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikabadze:1988:MCG

[Nik88a] A. M. Nikabadze. On a method for constructing goodness-of-fit tests for parametric hypotheses in R^{m}. Theory of Probability and its Applications, 32(3):539–544, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikitin:1988:HLA

[Nik88b] Ya. Yu. Nikitin. On the Hodges– Lehmann asymptotic efficiency of nonparametric tests of goodness of fit and homogeneity. Theory of Probability and its Applications, 32(1):77–85, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikulin:1991:TMA

[Nik91] V. N. Nikulin. On two metrics applied in problems of stability characterizations. Theory of Probability and its Applications, 35(2):377–380, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 367–370.

Nikolaev:1999:FMD

[Nik99a] A. F. Nikolaev. On a formulation of the multiple “Disorder” problem. Theory of Probability and its Applications, 43 (2):293–297, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97690.

Nikolaev:1999:OMS

[Nik99b] M. L. Nikolaev. On optimal multiple stopping of Markov sequences. Theory of Probability and its Applications, 43(2): 298–306, June 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97691.

Novikov:1979:TAU

[NK79] A. A. Novikov and A. S. Kholevo. The twelfth all-union seminar on probability theory and mathematical statistics. Theory of Probability and its Applications, 23(4):853, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1981:LTC

[NK81] S. V. Nagaev and L. V. Khan. Limit theorems for a critical Galton–Watson branching process with migration. Theory of Probability and its Applications, 25(3):514–525, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1982:ALT

[NK82] S. V. Nagaev and L. V. Khan.

Addendum: Limit theorems for a critical Galton–Watson branching process with migration. Theory of Probability and its Applications, 26(2):434, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1997:ECF

[NK97] S. V. Nagaev and S. S.

Khodzhabagyan. On an estimate for the concentration function of sums of independent random variables. Theory of Probability and its Applications, 41 (3):560–569, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 655–665.

Nagaev:1976:LDD

[NM76] A. V. Nagaev and T. P. Mukhomor. A limit distribution of the duration of an epidemic. Theory of Probability and its Applications, 20(4):805–818, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Norin:1988:SID

[Nor88] N. V. Norin. Stochastic integrals and differential measures. Theory of Probability and its Applications, 32(1):107–116, ???? 1988.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Norvaisha:1991:RSS

[Nor91] R. Norvaĭsha. A remark on some stability results for empirical processes. Theory of Probability and its Applications, 35(2):380–383, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 370–373.

Norberg:1993:ORS

[Nor93a] T. Norberg. An ordered random set coupling. Theory of Probability and its Applications, 37 (1):161–163, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 197–199.

Norvaisha:1993:CLT

[Nor93b] R. Norvaĭsha. The Central Limit Theorem for empirical and quantile processes in L_{P} spaces, 1 ≤ P < ∞. Theory of Probability and its Applications, 37 (2):302–313, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 329–341.

Norin:1995:IFE

[Nor95] N. V. Norin. Itô formula for an extended stochastic integral with nonanticipating kernel. Theory of Probability and its Applications, 39(4):573–592, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 743–765.

Norvaisa:1998:SLW

[Nor98] R. Norvaisa. Strassen’s law for weighted sup-norms. Theory of Probability and its Applications, 42(4):646–655, December 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97651.

Nosko:1977:LTP

[Nos77] V. P. Nosko. A limit theorem for a probabilistic model of corrosion. Theory of Probability and its Applications, 21(4):811–818, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1979:AND

[Nos79] V. P. Nosko. On the average number of distinctive level curve points of a random field on the plane. Theory of Probability and its Applications, 24(1):181–183, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1980:DNE

[Nos80] V. P. Nosko. On the definition of the number of excursions of a random field above a fixed level. Theory of Probability and its Applications, 24(3):598–602, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1983:HRF

[Nos83a] V. P. Nosko. The horizon of a random field of cones on a plane. The mean number of horizon corners. Theory of Probability and its Applications, 27(2):269–279, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1983:WCH

[Nos83b] V. P. Nosko. Weak convergence of the horizon of a random field of cones in an expanding strip. Theory of Probability and its Applications, 27(4):744–758, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1985:ADC

[Nos85] V. P. Nosko. Asymptotic distributions of curvatures of a homogeneous Gaussian random field at points of high local maxima. Theory of Probability and its Applications, 29(4):817–823, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1986:LSH

[Nos86] V. P. Nosko. The local structure of a homogeneous Gaussian random field in a neighborhood of high level points. Theory of Probability and its Applications, 30(4):767–782, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1987:DAL

[Nos87] V. P. Nosko. On the distribution of the arc length of a high-level excursion of a stationary Gaussian process. Theory of Probability and its Applications, 31 (3):521–523, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1988:ADC

[Nos88] V. P. Nosko. Asymptotic distributions of characteristics of high-level overshoots of a homogeneous Gaussian random field. Theory of Probability and its Applications, 32(4):659–669, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nosko:1991:IHL

[Nos91] V. P. Nosko. The investigation of high-level excursions of Gaussian fields: a fresh approach involving convexity. Theory of Probability and its Applications, 35 (1):157–163, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 148–153.

Novikov:1971:MIS

[Nov71a] A. A. Novikov. On moment inequalities for stochastic integrals. Theory of Probability and its Applications, 16(3):538–541, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1971:STW

[Nov71b] A. A. Novikov. On stopping times for a Wiener process. Theory of Probability and its Applications, 16(3):449–456, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1973:ISI

[Nov73] A. A. Novikov. On an identity for stochastic integrals. Theory of Probability and its Applications, 17(4):717–720, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1975:DM

[Nov75] A. A. Novikov. On discontinuous martingales. Theory of Probability and its Applications, 20 (1):11–26, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1979:RIP

[Nov79] A. A. Novikov. Recurrent interpolation of partially observed random fields with discrete parameter. Theory of Probability and its Applications, 23(2): 390–395, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1980:CUI

[Nov80] A. A. Novikov. On conditions for uniform integrability of continuous non-negative martingales. Theory of Probability and its Applications, 24(4):820–824, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1982:ETS

[Nov82] A. A. Novikov. On the exit time of sums of bounded random variables from a curvilinear strip. Theory of Probability and its Applications, 26(2):279–292, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1983:CTO

[Nov83] A. A. Novikov. The crossing time of a one-sided nonlinear boundary by sums of independent random variables. Theory of Probability and its Applications, 27 (4):688–702, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1988:AOS

[Nov88] An. A. Novikov. Asymptotic optimality of the sequential dguaranty test. Theory of Probability and its Applications, 32 (2):358–361, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:1991:FPT

[Nov91] A. A. Novikov. On the first passage time of an autoregressive process over a level and an application to a “Disorder” problem. Theory of Probability and its Applications, 35(2):269–279, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 282–292.

Novak:1992:DMR

[Nov92] S. Yu. Novak. On the distribution of the maximum of a random number of random variables. Theory of Probability and its Applications, 36(4): 714–721, December 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novak:1995:PAN

[Nov95] Yu. S. Novak. Poisson approximation for the number of long match patterns in random sequences. Theory of Probability and its Applications, 39(4): 593–603, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 731–742.

Novak:1996:APA

[Nov96] S. Yu. Novak. Addendum: Poisson approximation for the number of long match patterns in random sequences. Theory of Probability and its Applications, 40(3):596, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novak:1997:DRS

[Nov97a] S. Yu. Novak. On the distribution of the ratio of sums of random variables. Theory of Probability and its Applications, 41(3):479–503, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97525. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 533–560.

Novikov:1997:MTT

[Nov97b] A. A. Novikov. Martingales, Tauberian theorem, and strategies of gambling. Theory of Probability and its Applications, 41 (4):716–729, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97571.

Novak:1998:ERM

[Nov98] S. Yu. Novak. On the Erdős– Rényi maximum of partial sums. Theory of Probability and its Applications, 42(2):254–270, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97611. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(2), (1997), pp. 274–293.

Novikov:1999:HOG

[Nov99] A. A. Novikov. Hedging of options with a given probability. Theory of Probability and its Applications, 43(1):135– 143, March 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97673.

Nevzorov:1969:DMC

[NP69] V. B. Nevzorov and V. V. Petrov. On the distribution of the maximum cumulative sum of independent random variables. Theory of Probability and its Applications, 14(4):682–687, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1978:SID

[NP78] S. V. Nagaev and I. F. Pinelis. Some inequalities for the distribution of sums of independent random variables. Theory of Probability and its Applications, 22(2):248–256, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1999:ILL

[NP99] S. V. Nagaev and E. L. Presman. On the iterated logarithm lawin a control problem. Theory of Probability and its Applications, 43(2):288–293, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97689.

Nagaev:1973:SLT

[NR73] S. V. Nagaev and V. I. Rotar’. On strengthening Lyapunov type estimates (the case when the distribution of the summands is close to the normal distribution). Theory of Probability and its Applications, 18(1):107–119, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1976:ASL

[NR76] S. V. Nagaev and V. I. Rotor’.

Addendum: On strengthening Lyapunov type estimates. Theory of Probability and its Applications, 21(1):220, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Natarajan:1967:SMK

[NRV67] S. Natarajan, T. E. S. Raghavan, and K. Viswanath. On stochastic matrices and kernels. Theory of Probability and its Applications, 12(2):294–297, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1970:AAS

[NS70] A. V. Nagaev and A. N. Startsev. The asymptotic analysis of a stochastic model of an epidemic. Theory of Probability and its Applications, 15(1):98–107, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1989:SPS

[NS89] A. V. Nagaev and S. M.

Shkol′nik. Some properties of symmetric stable distributions close to the normal distribution. Theory of Probability and its Applications, 33(1):139–144, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 150–154.

Nagaev:1975:EPL

[NV75] S. V. Nagaev and N. V. Vahrusev. Estimation of probabilities of large deviations for a critical Galton–Watson process. Theory of Probability and its Applications, 20(1):179–180, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:1976:SLL

[NV76] S. V. Nagaev and N. A. Volodin. On the strong law of large numbers. Theory of Probability and its Applications, 20(3):626–631, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikunen:1985:NOD

[NV85] M. Nikunen and E. Valkeila. A note on one-dimensional distances between two counting processes. Theory of Probability and its Applications, 29(3):578–581, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nyblom:1986:LDP

[Nyb86] J. Nyblom. Large deviation probabilities for the ratio of two quadratic forms in normal variables. Theory of Probability and its Applications, 30(1):157–160, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nyssbaum:1987:NER

[Nys87] M. Nyssbaum. Nonparametric estimation of a regression function that is smooth in a domain in R^{k}. Theory of Probability and its Applications, 31(1):108–115, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Obreja:1993:DDS

[OCG93] G. Obreja, I. Chitescu, and F. C.

Gheorghe. On the distribution of the decisive sample number of the invariant T^{2}-sequential test. Theory of Probability and its Applications, 37(2):348–350, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 355–357.

Orsinger:1997:EDM

[OK97] È. Orsinger and A. Kolesnik. The exact distribution in a model of random motion in a plane, controlled by a fourth-order hyperbolic equation. Theory of Probability and its Applications, 41 (2):379–386, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 451–459.

Olman:1989:PBP

[Ol′89] V. N. Ol′man. Preservation of the Bayes property in the problem of estimating functions of the mean of the normal law. Theory of Probability and its Applications, 33(2):399–402, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 424–428.

Oliveira:1993:LDS

[Oli93] M. F. Oliveira. Limiting distribution of sums of certain types of nonstationary sequences. Theory of Probability and its Applications, 37(2):350–353, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 380–383.

Omey:1997:OPT

[Ome97] E. Omey. Operators of probabilistic type. Theory of Probability and its Applications, 41 (1):178–185, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 219–226.

Onishchenko:1990:EPM

[Oni90] V. F. Onishchenko. Estimation of the parameters of the moving average in the case of infinite variance. Theory of Probability and its Applications, 34(2):335–340, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 379–384.

Osipov:1967:ERT

[OP67] L. V. Osipov and V. V. Petrov.

On an estimate of the remainder term in the Central Limit Theorem. Theory of Probability and its Applications, 12(2):281–286, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1985:IDC

[OR85] L. V. Osipov and V. I. Rotar’.

On an infinite-dimensional Central Limit Theorem. Theory of Probability and its Applications, 29(2):375–383, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Omey:1989:RCE

[OR89] E. Omey and S. Rachev. On the rate of convergence in extreme value theory. Theory of Probability and its Applications, 33 (3):560–566, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 601–607.

Orlov:1973:TSD

[Orl73] A. I. Orlov. On testing the symmetry of distributions. Theory of Probability and its Applications, 17(2):357–361, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Orlov:1975:RCD

[Orl75] A. I. Orlov. The rate of convergence of the distribution of the Smirnov-von Mises statistic. Theory of Probability and its Applications, 19(4):737–757, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Oseledets:1965:MCS

[Ose65] V. I. Oseledets. Markov chains, skew products and ergodic theorems for “General” dynamic systems. Theory of Probability and its Applications, 10(3):499–504, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Oseledets:1990:KTT

[Ose90] V. I. Oseledets. Kolmogorov’s theory of turbulence and the evolution of states of a system of cellular automata of the type of lattice gas invariant states. Theory of Probability and its Applications, 34(1):140–152, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(1), (1989), pp. 165–177.

Osipov:1966:RLT

[Osi66] L. V. Osipov. Refinement of Lindeberg’s theorem. Theory of Probability and its Applications, 11(2):299–302, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1969:AED

[Osi69] L. V. Osipov. Asymptotic expansions of the distribution function of a sum of independent lattice random variables. Theory of Probability and its Applications, 14(3):450–457, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1971:AED

[Osi71] L. V. Osipov. On asymptotic expansions for distribution functions of sums of independent random variables. Theory of Probability and its Applications, 16 (2):333–343, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1973:PLD

[Osi73] L. V. Osipov. On probabilities of large deviations for sums of independent random variables. Theory of Probability and its Applications, 17(2):309–331, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1974:APL

[Osi74] L. V. Osipov. Addendum: On probabilities of large deviations for sums of independent random variables. Theory of Probability and its Applications, 18(3):642, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1975:MLT

[Osi75] L. V. Osipov. Multidimensional limit theorems for large deviations. Theory of Probability and its Applications, 20 (1):38–56, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osipov:1979:LDP

[Osi79] L. V. Osipov. On large deviation probabilities for sums of independent random vectors. Theory of Probability and its Applications, 23(3):490–506, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ososkov:1956:LTF

[Oso56] G. A. Ososkov. A limit theorem for flows of similar events. Theory of Probability and its Applications, 1(2):248–255, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1965:MAY

[Ost65] I. V. Ostrovskii. The multidimensional analogue of Yu. V. Linnik’s theorem on decompositions of a convolution of Gaussian and Poisson laws. Theory of Probability and its Applications, 10 (4):673–677, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1967:AMA

[Ost67] I. V. Ostrovskii. Addendum: Multidimensional analogue of Yu. V. Linnik’s theorem on decompositions of a convolution of Gaussian and Poisson laws. Theory of Probability and its Applications, 12(3):529–530, ????

1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1969:PCR

[Ost69] I. V. Ostrovskii. On a problem of C. R. Rao. Theory of Probability and its Applications, 14 (2):312–313, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1983:DID

[Ost83] I. V. Ostrovskii. On divisors of infinitely divisible distributions representable as Cartesian products. Theory of Probability and its Applications, 27(4):832–837, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1987:APD

[Ost87] I. V. Ostrovskii. The arithmetic of probability distributions. Theory of Probability and its Applications, 31(1):1–24, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1989:ZEC

[Ost89] I. V. Ostrovskii. On the zeros of entire characteristic functions of distributions concentrated on the half-line. Theory of Probability and its Applications, 33 (1):167–171, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:1991:EBD

[Ost91] E. I. Ostrovskiĭ. Exponential bounds for the distribution of the maximum of a non-Gaussian random field. Theory of Probability and its Applications, 35(3): 487–499, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 482–493.

Ostrovskii:1998:SCE

[Ost98] E. I. Ostrovskii. Short communications: Estimates of the distribution of the maximum of a random field. Theory of Probability and its Applications, 42 (2):302–310, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97616.

Ostrovskii:1995:APD

[OT95] E. I. Ostrovskiĭ and S. Yu. Tsykunova. Asymptotic properties of distributions of the maximum of a Gaussian nonstationary process occurring in covariance statistic. Theory of Probability and its Applications, 39 (3):527–534, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 641–649.

Ouknine:1991:SBM

[Ouk91] Y. Ouknine. “Skew-Brownian motion” and derived processes.

Theory of Probability and its Applications, 35(1):163–169, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ozawa:1986:CAF

[Oza86] M. Ozawa. Continuous affine functions on the space of Markov kernels. Theory of Probability and its Applications, 30(3): 516–528, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1965:EQD

[Pag65] V. I. Pagurova. On the evaluation of quantiles of the Γdistribution. Theory of Probability and its Applications, 10 (4):677–680, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1968:CMT

[Pag68] V. I. Pagurova. On a comparison of means of two normal samples. Theory of Probability and its Applications, 13(3):527–534, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1983:CEG

[Pag83] V. I. Pagurova. On confidence estimation in a general linear regression model with nonhomogeneous variances. Theory of Probability and its Applications, 27 (2):407–411, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1986:CTF

[Pag86] V. I. Pagurova. On Chauvenet’s test for finding several outliers. Theory of Probability and its Applications, 30(3): 594–598, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1998:APT

[Pag98] V. I. Pagurova. On the asymptotic power of a test for detecting outliers. Theory of Probability and its Applications, 42 (3):433–443, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 482–495.

Pakshirajan:1959:MPS

[Pak59] R. P. Pakshirajan. On the maximum partial sums of sequences of independent random variables. Theory of Probability and its Applications, 4(4): 367–372, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pancheva:1987:GLT

[Pan87] E. Pancheva. General limit theorems for maximum of independent random variables. Theory of Probability and its Applications, 31(4):645–657, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pancheva:1989:MS

[Pan89] E. Pancheva. Max-stability. Theory of Probability and its Applications, 33(1):155–158, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pancheva:1991:SDM

[Pan91] E. Pancheva. Stability of decomposition in multiplicative semigroup of multidimensional distributions. Theory of Probability and its Applications, 35(3): 602–604, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 609–611.

Pancheva:1994:CTU

[Pan94] E. Pancheva. Convergence of types under monotonous mappings. Theory of Probability and its Applications, 38(3):551–556, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 679–684.

Pancheva:1995:PKT

[Pan95] E. Pancheva. On a problem of a Khinchin-type decomposition theorem for extreme values. Theory of Probability and its Applications, 39(2):329–336, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 395–402.

Parthasarathy:1964:CLT

[Par64] K. R. Parthasarathy. The Central Limit Theorem for the rotation group. Theory of Probability and its Applications, 9 (2):248–257, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Parthasarathy:1966:PRW

[Par66] K. R. Parthasarathy. On a problem in random walk. Theory of Probability and its Applications, 11(4):591–596, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Parthasarathy:1967:IID

[Par67] K. R. Parthasarathy. On the imbedding of an infinitely divisible distribution in a oneparameter convolution semigroup. Theory of Probability and its Applications, 12(3):373–380, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Paramonov:1970:EEB

[Par70] Yu. M. Paramonov. Explicit expression for the Bayesian risk in two cases of sequential analysis. Theory of Probability and its Applications, 15(2):349–354, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Partzsch:1971:ETM

[Par71] Lothar Partzsch. On an ergodic theorem for Markov processes with finite lifetime. Theory of Probability and its Applications, 16(4):697–700, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pardzhanadze:1987:FLT

[Par87] A. M. Pardzhanadze. Functional limit theorems in the problem of a posteriori detection of disorder. Theory of Probability and its Applications, 31(2):355–358, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pashko:1999:UCS

[Pas99] A. A. Pashko. Uniform convergence of sub-Gaussian integrals. Theory of Probability and its Applications, 43(4):650– 655, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97723.

Paulauskas:1977:RCC

[Pau77] V. I. Paulauskas. On the rate of convergence in the Central Limit Theorem in certain Banach spaces. Theory of Probability and its Applications, 21(4):754–769, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 21, (1976), pp. 775– 791.

Paulauskas:1979:ARC

[Pau79] V. I. Paulauskas. Addendum:

On the rate of convergence in the Central Limit Theorem in certain Banach spaces. Theory of Probability and its Applications, 23(2):458, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Paulauskas:1991:NGM

[Pau91] V. I. Paulauskas. A note on Gaussian measure of balls in Banach spaces. Theory of Probability and its Applications, 35 (4):802–805, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 784–787.

Pavlov:1978:ADM

[Pav78] Yu. L. Pavlov. The asymptotic distribution of maximum tree size in a random forest. Theory of Probability and its Applications, 22(3):509–520, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1980:CCD

[Pav80] I. V. Pavlov. A counterexample to the conjecture that H^{∞} is dense in BMO. Theory of Probability and its Applications, 25 (1):152–154, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1981:CBC

[Pav81a] I. V. Pavlov. Confidence bounds for convex functions of many unknown parameters. Theory of Probability and its Applications, 25(2):388–393, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1981:ERI

[Pav81b] I. V. Pavlov. On extreme rays and the integral representation in a cone of supermartingales. Theory of Probability and its Applications, 25(3):593–596, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1982:LTC

[Pav82] Yu. L. Pavlov. Limit theorems for a characteristic of a random mapping. Theory of Probability and its Applications, 26(4):829–834, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1984:LDH

[Pav84] Yu. L. Pavlov. Limit distributions of the height of a random forest. Theory of Probability and its Applications, 28(3):471–480, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1988:SPT

[Pav88] I. V. Pavlov. A sequential procedure for testing many composite hypotheses. Theory of Probability and its Applications, 32 (1):138–142, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:1989:LLT

[Pav89a] A. I. Pavlov. Local limit theorems for the number of components of random permutations and mappings. Theory of Probability and its Applications, 33 (1):183–187, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 196–200.

Pavlov:1989:DNV

[Pav89b] Yu. L. Pavlov. Distributions of the number of vertices in strata of a random forest. Theory of Probability and its Applications, 33(1):96–104, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 105–114.

Pavlov:1991:AOQ

[Pav91a] A. V. Pavlov. Asymptotically optimal queue disciplines in heavy traffic. Theory of Probability and its Applications, 35(4):696–706, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 716–726.

Pavlov:1991:SPT

[Pav91b] I. V. Pavlov. Sequential procedure of testing composite hypotheses with applications to the Kiefer-Weiss problem. Theory of Probability and its Applications, 35(2):280–292, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 293–304.

Pavlov:1992:AOQ

[Pav92]  A. V. Pavlov.  Asymptoti-

cally optimal queueing disciplines for some multichannel networks. Theory of Probability and its Applications, 36(1):185–190, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 170–175.

Pavlov:1997:ISM

[Pav97] I. V. Pavlov. On the invariance of some martingale inequalities for the processes of more general kind. Theory of Probability and its Applications, 41(2):288–295, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97549. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 300–309.

Pavlov:1998:HSS

[Pav98] I. V. Pavlov. Haar systems and some results on the basis in a martingale space with mixed norm. Theory of Probability and its Applications, 42(3):528–531, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 623–626.

Pazman:1968:SDE

[Páz68] A. Pázman. The sequential design of experiments performed at several measurement points. Theory of Probability and its Applications, 13(3):457–467, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1996:TDO

[PC96] V. I. Pagurova and I. L. Chizhikova. Tests for detection of outliers based on robust estimators of nuisance parameters. Theory of Probability and its Applications, 40(2):390–397, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 445–452.

Pecerskii:1974:SIR

[Pec74a] E. A. Pecerskii. Some identities related to the exit of a random walk from a segment and a semi-interval. Theory of Probability and its Applications, 19 (1):106–121, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pechinkin:1974:CRS

[Pec74b] A. V. Pechinkin. On convergence of random sums of random variables to the normal law. Theory of Probability and its Applications, 18(2):366–367, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pechinkin:1977:SLD

[Pec77] A. V. Pechinkin. Some limit distributions for processes with independent increments. Theory of Probability and its Applications, 22(1):175–183, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pechinkin:1981:LDR

[Pec81] A. V. Pechinkin. The limit distribution for a random walk with absorption. Theory of Probability and its Applications, 25 (3):580–584, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Penkov:1977:ADP

[Pen77] B. I. Penkov. Asymptotic distribution of Pyke’s statistic. Theory of Probability and its Applications, 21(2):370–374, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Penev:1993:SNP

[Pen93] S. Penev. Stability of nonparametric procedures against dependence. Theory of Probability and its Applications, 37(2):353–355, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 383–386.

Perez:1959:ITA

[Per59] Albert Perez. Information theory with an abstract alphabet (generalized forms of McMillan’s Limit Theorem) for the case of discrete and continuous times. Theory of Probability and its Applications, 4(1):99–102, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pergamenshchikov:1992:APS

[Per92] S. M. Pergamenshchikov. Asymptotic properties of sequential design for estimating the parameter of a first-order autoregression. Theory of Probability and its Applications, 36(1):36–49, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 42–53.

Peskir:1994:MCS

[Pes94] G. Peskir. Measure compact sets of functions and consistency of statistical models. Theory of Probability and its Applications, 38(2):360–367, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 431–438.

Petrov:1956:VSH

[Pet56a] A. A. Petrov. Verification of statistical hypotheses on the type of a distribution based on small samples. Theory of Probability and its Applications, 1(2): 223–245, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1956:LTD

[Pet56b] V. V. Petrov. A local theorem for densities of sums of independent random variables. Theory of Probability and its Applications, 1(3):316–322, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1959:AED

[Pet59a] V. V. Petrov. Asymptotic expansions for distributions of sums of independent random variables. Theory of Probability and its Applications, 4(2):208–211, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1959:CLT

[Pet59b] V. V. Petrov. On a class of limit theorems for independent random variables. Theory of Probability and its Applications, 4(2):211–214, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1959:SPP

[Pet59c] V. V. Petrov. Summary of papers presented at the sessions of the Probability Research Seminar, Moscow University (February-May 1959). Theory of Probability and its Applications, 4 (4):432–443, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1962:LLT

[Pet62] V. V. Petrov. Local limit theorems for non-identical lattice distributions. Theory of Probability and its Applications, 7 (3):333–335, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1964:LLT

[Pet64] V. V. Petrov. On local limit theorems for sums of independent random variables. Theory of Probability and its Applications, 9(2):312–320, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1965:PLD

[Pet65] V. V. Petrov. On the probabilities of large deviations for sums of independent random variables. Theory of Probability and its Applications, 10(2): 287–298, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1966:ARK

[Pet66a] V. V. Petrov. Addendum: a remark on the Katz–Petrov theorem. Theory of Probability and its Applications, 11(2):320, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1966:RBE

[Pet66b] V. V. Petrov. On a relation between an estimate of the remainder in the Central Limit Theorem and the law of the iterated logarithm. Theory of Probability and its Applications, 11 (3):454–458, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1968:ABP

[Pet68] V. V. Petrov. Asymptotic behavior of probabilities of large deviations. Theory of Probability and its Applications, 13(3):408–420, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1969:SLL

[Pet69] V. V. Petrov. On the strong law of large numbers. Theory of Probability and its Applications, 14(2):183–192, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1970:ECF

[Pet70] V. V. Petrov. On an estimate of the concentration function of a sum of independent random variables. Theory of Probability and its Applications, 15(4):701–703, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1971:TLI

[Pet71] V. V. Petrov. A theorem on the law of the iterated logarithm. Theory of Probability and its Applications, 16(4):700–702, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1974:MDA

[Pet74a] V. V. Petrov. Moments of distributions attracted to stable laws. Theory of Probability and its Applications, 18(3):569–571, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1974:OGS

[Pet74b] V. V. Petrov. On the order of growth and sums of dependent variables. Theory of Probability and its Applications, 18 (2):348–350, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1975:GIL

[Pet75] V. V. Petrov. A generalization of an inequality of Lévy. Theory of Probability and its Applications, 20(1):141–145, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:1976:IMR

[Pet76] V. V. Petrov. An inequality for moments of a random variable. Theory of Probability and its Applications, 20(2):391–392, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pettselberger:1993:UM

[Pët93] K. Pëttsel′berger. Updating of moments. Theory of Probability and its Applications, 37 (2):355–357, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 386–388.

Prokhorov:1957:CND

[PF57] Yu. V. Prokhorov and M. Fish.

A characterization of normal distributions in Hilbert space. Theory of Probability and its Applications, 2(4):468–470, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1980:CIC

[PG80] V. I. Pagurova and V. V. Gurskii.

A confidence interval for the common mean of several normal distributions. Theory of Probability and its Applications, 24 (4):882–888, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinsky:1969:NDD

[Pin69] M. Pinsky. A note on degenerate diffusion processes. Theory of Probability and its Applications, 14(3):502–506, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1979:DSI

[Pin79] I. F. Pinelis. On the distribution of sums of independent random variables with values in a Banach space. Theory of Probability and its Applications, 23 (3):608–615, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1981:PLD

[Pin81] I. F. Pinelis. A problem on large deviations in a space of trajectories. Theory of Probability and its Applications, 26 (1):69–84, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1991:IDS

[Pin91a] I. F. Pinelis. Inequalities for distributions of sums of independent random vectors and their application to estimating a density. Theory of Probability and its Applications, 35(3): 605–607, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 592–594.

Pinelis:1991:MER

[Pin91b] I. F. Pinelis. On minimax estimation of regression. Theory of Probability and its Applications, 35(3):500–512, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1991:MR

[Pin91c] I. F. Pinelis. On minimax risk. Theory of Probability and its Applications, 35(1):104–109, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piranashvili:1967:PIR

[Pir67] Z. A. Piranashvili. On the problem of interpolation of random processes. Theory of Probability and its Applications, 12(4):

647–657, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pirogov:1968:PCE

[Pir68] S. A. Pirogov. Probabilities of compound events and linear programming. Theory of Probability and its Applications, 13 (2):329–332, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pisarenko:1965:CLR

[Pis65] V. F. Pisarenko. On the computation of the likelihood ratio for Gaussian processes with a rational spectrum. Theory of Probability and its Applications, 10 (2):299–303, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pisanets:1980:SCS

[Pis80] S. I. Pisanets. On the strong convergence of a sequence of diffusion type processes. Theory of Probability and its Applications, 24(4):863–869, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pisanets:1982:LTD

[Pis82] S. I. Pisanets. Limit theorems for diffusion-type processes in R^{m}. Theory of Probability and its Applications, 26(3):587–594, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pisanets:1995:TLT

[Pis95] S. I. Pisanets. Two limit theorems for diffusion type stochastic equations. Theory of Probability and its Applications, 39 (4):708–713, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 845–851.

Pisanets:1998:SCE

[Pis98] S. I. Pisanets. Short communication: On the existence of strong solutions of diffusiontype stochastic equations in R^{m∗}. Theory of Probability and its Applications, 42(1):179–184, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97602. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 189–194.

Pittel:1969:AED

[Pit69a] B. G. Pittel. On asymptotic estimates in a dynamic decisionmaking problem. Theory of Probability and its Applications, 14 (2):244–262, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pittenger:1969:CDB

[Pit69b] A. O. Pittenger. Comments on Dynkin’s boundary conditions for continuous-time Markov chains. Theory of Probability and its Applications, 14(4):627–639, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pittel:1971:LPP

[Pit71a] B. G. Pittel’. A linear programming problem connected with optimal stationary control in a dynamic decision problem. Theory of Probability and its Applications, 16(4):724–728, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pittel:1971:MAW

[Pit71b] B. G. Pittel. A minimax analogue of the weak law of large numbers. Theory of Probability and its Applications, 16(2):361–367, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1973:LTM

[Pit73] L. I. Piterbarg. Limit theorems for Markov random sets. Theory of Probability and its Applications, 17(3):426–433, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 17, (1972), pp. 450–457.

Pittel:1975:GRW

[Pit75] B. G. Pittel’. On a game random walk on a straight line. Theory of Probability and its Applications, 19(3):604–614, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1978:CLT

[Pit78] V. I. Piterbarg. The Central Limit Theorem for the number of level crossings of a stationary Gaussian process. Theory of Probability and its Applications, 23(1):178–182, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1982:CDF

[Pit82] V. I. Piterbarg. Comparison of distribution functions of maxima of Gaussian processes. Theory of Probability and its Applications, 26(4):687–705, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1983:LDR

[Pit83] V. I. Piterbarg. Large deviations of random processes close to Gaussian ones. Theory of Probability and its Applications, 27 (3):504–524, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1984:PCR

[Pit84] L. I. Piterbarg. On prediction of a class of random fields. Theory of Probability and its Applications, 28(1):184–191, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1987:SIA

[Pit87a] L. I. Piterbarg. On the structure of the infinitesimal σ-algebra of Gaussian processes and fields. Theory of Probability and its Applications, 31(3):484–492, ????

1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1987:SSS

[Pit87b] L. I. Piterbarg. On the structure of the smallest splitting σ-algebra for a generalized Gaussian field with rational spectrum. Theory of Probability and its Applications, 31(2):351–355, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1992:LJR

[Pit92] V. I. Piterbarg. On large jumps of a random walk. Theory of Probability and its Applications, 36(1):50–62, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 54–64.

Piunovskii:1994:CRS

[Piu94] A. B. Piunovskiĭ. Control of random sequences in problems with constraints. Theory of Probability and its Applications, 38 (4):751–762, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 891–903.

Piunovskii:1998:CJD

[Piu98] A. B. Piunovskii. A controlled jump discounted model with constraints. Theory of Probability and its Applications, 42(1):

51–72, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97596.

Parsadanishvili:1982:TSH

[PK82] E. G. Parsadanishvili and E. V. Khmaladze. The testing of statistical hypotheses on unidentifiable objects. Theory of Probability and its Applications, 27 (1):175–182, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pardzhanadze:1987:ATS

[PK87] A. M. Pardzhanadze and É. V. Khmaladze. On the asymptotic theory of statistics of sequential ranks. Theory of Probability and its Applications, 31(4):669–682, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrovskaya:1983:CLT

[PL83] M. B. Petrovskaya and A. M. Leontovich. The Central Limit Theorem for a sequence of random variables with a slowly growing number of dependences. Theory of Probability and its Applications, 27(4):815–825, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Plaksin:1992:SPS

[Pla92] V. A. Plaksin. A statistical property of the sieve of Eratosthenes. Theory of Probability and its Applications, 36(3):

608–614, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 587–593.

Plachky:1996:CCR

[Pla96] D. Plachky. Characterizations of completion regularity of measures. Theory of Probability and its Applications, 40(1):181–186, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 220–225.

Plucinska:1981:RPE

[Plu81]  A. Plucinska.  Remarks on

prospective equations. Theory of Probability and its Applications, 25(2):347–355, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Plucinska:1985:CGD

[Plu85] A. Plucińska. A characterization of Gaussian distributions based on the Lapunow condition. Theory of Probability and its Applications, 29(3):591–595, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Plucinska:1992:SLP

[Plu92] A. Plucińska. Some limit properties of processes with the linear regression. Theory of Probability and its Applications, 36(3): 615–619, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 612–615.

Pagurova:1992:WCC

[PN92] V. I. Pagurova and S. A. Nesterova. On the weak convergence of counting processes with nuisance parameters. Theory of Probability and its Applications, 36(1):176–185, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 176–183.

Polishchuk:1961:CMS

[Pol61] E. M. Polishchuk. On continual means and singular distributions. Theory of Probability and its Applications, 6(4): 427–430, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pollyak:1963:CTP

[Pol63] Yu. G. Pollyak. On a comparison of two probabilities. Theory of Probability and its Applications, 8(2):182–183, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Polin:1978:LTB

[Pol78] A. K. Polin. Limit theorems for branching processes with immigration. Theory of Probability and its Applications, 22 (4):746–754, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ponomarenko:1979:IDQa

[Pon79a] L. S. Ponomarenko. Inequalities for distributions of quadratic forms in normal random variables. Theory of Probability and its Applications, 23(3):652–656, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ponomarenko:1979:IDQb

[Pon79b] L. S. Ponomarenko. An inequality for distributions of quadratic forms of normally distributed random variables. Theory of Probability and its Applications, 23(3):656–660, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ponomarenko:1986:EDN

[Pon86] L. S. Ponomarenko. On estimating distributions of normalized quadratic forms of normally distributed random variables. Theory of Probability and its Applications, 30(3):580–584, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Popova:1968:LTM

[Pop68] T. Yu. Popova. Limit theorems in a model of distribution of particles of two types. Theory of Probability and its Applications, 13(3):511–516, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Popov:1980:RCC

[Pop80] N. N. Popov. On the rate of convergence for countable Markov chains. Theory of Probability and its Applications, 24(2):401–405, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Portenko:1967:IEL

[Por67] N. I. Portenko. Integral equations and limit theorems for additive functionals of a Markov process. Theory of Probability and its Applications, 12(3):500–505, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Portenko:1975:DPU

[Por75a] N. I. Portenko. Diffusion processes with unbounded drift coefficient. Theory of Probability and its Applications, 20(1): 27–37, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Portenko:1975:ESS

[Por75b] N. I. Portenko. On the existence of solutions of stochastic differential equations with integrable drift coefficient. Theory of Probability and its Applications, 19 (3):552–557, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Portenko:1979:DPG

[Por79] N. I. Portenko. Diffusion processes with generalized drift coefficients. Theory of Probability and its Applications, 24 (1):62–78, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Portenko:1980:SDE

[Por80] N. I. Portenko. Stochastic differential equations with generalized drift vector. Theory of Probability and its Applications, 24 (2):338–353, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Postnikova:1961:CBC

[Pos61] L. P. Postnikova. On the connection between the concepts of collectives of Mises–Church and normal Bernoulli sequences of symbols. Theory of Probability and its Applications, 6(2): 211–213, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:1982:EAB

[PP82] V. I. Piterbarg and V. P.

Prisyazhnyuk. Exact asymptotic behavior of the probability of a large span of a stationary Gaussian process. Theory of Probability and its Applications, 26 (3):468–484, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Polasek:1993:RBM

[PP93] W. Polasek and K. Pötzelberger.

Robust Bayesian methods in simple ANOVA models. Theory of Probability and its Applications, 37(1):163–166, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 200–202.

Prabhu:1995:SMD

[PP95] N. U. Prabhu and A. Pacheco. A storage model for data communication systems. Theory of Probability and its Applications, 39(4):604–627, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 766–795.

Papadatos:1996:DVB

[PP96] N. Papadatos and V. Papathanasiou. Distance in variation between two arbitrary distributions via the associated wfunctions. Theory of Probability and its Applications, 40(3): 567–575, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 685–694.

Pecherskii:1969:JDR

[PR69a] E. A. Pecherskii and B. A. Rogozin. On joint distributions of random variables associated with fluctuations of a process with independent increments. Theory of Probability and its Applications, 14(3):410–423, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Polovin:1969:PDE

[PR69b] R. V. Polovin and V. V. Rozhkov. Partial differential equations with random characteristics. Theory of Probability and its Applications, 14(1):

146–149, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Puri:1985:CSR

[PR85] M. L. Puri and S. S. Ralescu.

Centering of signed rank statistics with a continuous scoregenerating function. Theory of Probability and its Applications, 29(3):601–607, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pagurova:1987:APT

[PR87] V. I. Pagurova and K. D. Rodionov. On asymptotic properties of a test for detecting outliers. Theory of Probability and its Applications, 31(4):705–708, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1997:LLN

[PR97] È. L. Presman and V. I. Rotar′.

On the law of large numbers for nonlinear functions of several random variables. Theory of Probability and its Applications, 41(3):586–592, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 665–672.

Pazizin:1999:ANL

[PR99] S. V. Pazizin and B. V.

Ryazanov. Asymptotic normality of the log-likelihood ratio for a class of m-dependent random variables. Theory of Probability and its Applications, 43(3): 422–433, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97701.

PrakasaRao:1993:SIC

[Pra93] B. L. S. Prakasa Rao. On some inequalities of Chernofftype. Theory of Probability and its Applications, 37(2):392–398, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 434–439.

Prakasa:1999:CSP

[Pra99] B. L. S. Rao Prakasa. On a characterization of stochastic processes by the absolute moments of stochastic integrals. Theory of Probability and its Applications, 43(1):144–145, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97674.

Presman:1965:WTM

[Pre65] E. L. Presman. On waiting time for many-server queueing systems. Theory of Probability and its Applications, 10(1): 63–73, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1967:BPS

[Pre67] E. L. Presman. Boundary problems for sums of lattice random variables, defined on a finite regular Markov chain. Theory of Probability and its Applications, 12(2):323–328, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1974:MVK

[Pre74] E. L. Presman. On a multidimensional version of the Kolmogorov uniform limit theorem. Theory of Probability and its Applications, 18(2):378–384, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1982:TIS

[Pre82] É L. Presman. Two inequalities for symmetric processes and symmetric distributions. Theory of Probability and its Applications, 26(4):815–819, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1984:ABD

[Pre84] E. L. Presman. Approximation of binomial distributions by infinitely divisible ones. Theory of Probability and its Applications, 28(2):393–403, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1986:AVD

[Pre86] E. L. Presman. Approximation in variation of the distribution of a sum of independent Bernoulli variables with a Poisson law. Theory of Probability and its Applications, 30 (2):417–422, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prelov:1989:CCC

[Pre89] V. V. Prelov. Communication channel capacity with almost Gaussian noise. Theory of Probability and its Applications, 33(3):405–422, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 433–452.

Presman:1991:PVT

[Pre91] È. L. Presman. Poisson version of the “two-armed bandit” problem with discounting. Theory of Probability and its Applications, 35(2):307–317, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 318–328.

Presman:1998:OAS

[Pre98] È. L. Presman. Optimality almost surely and in probability for a stochastic linearly quadratic regulator. Theory of Probability and its Applications, 42(3): 531–535, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 627–632.

Prokhorov:1956:CRP

[Pro56] Yu. V. Prokhorov. Convergence of random processes and limit theorems in probability theory. Theory of Probability and its Applications, 1(2):157–214, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1958:SSS

[Pro58] Yu. V. Prokhorov. Strong stability of sums and infinitely divisible distributions. Theory of Probability and its Applications, 3(2):141–153, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1959:EPP

[Pro59a] Yu. V. Prokhorov. An extremal problem in probability theory. Theory of Probability and its Applications, 4(2):201–203, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1959:SRS

[Pro59b] Yu. V. Prokhorov. Some remarks on the strong law of large numbers. Theory of Probability and its Applications, 4(2): 204–208, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1960:ULT

[Pro60] Yu. V. Prokhorov. On a uniform limit theorem of A. N. Kolmogorov. Theory of Probability and its Applications, 5 (1):98–106, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1965:CCP

[Pro65a] Yu. V. Prokhorov. On a characterization of a class of probability distributions by distributions of some statistics. Theory of Probability and its Applications, 10(3):438–445, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1965:LND

[Pro65b] Yu. V. Prokhorov. On the longnormal distribution in geochemical problems. Theory of Probability and its Applications, 10 (1):169–173, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1968:IBF

[Pro68a] A. V. Prokhorov. Inequalities for Bessel functions of a purely imaginary argument. Theory of Probability and its Applications, 13(3):496–501, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1968:MAB

[Pro68b] A. V. Prokhorov. On multidimensional analogues of S. N. Bernstein’s inequalities. Theory of Probability and its Applications, 13(2):268–280, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1968:BIM

[Pro68c] A. V. Prokhorov. S. N. Bernstein’s inequalities in the multidimensional case. Theory of Probability and its Applications, 13 (3):438–447, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1968:EBI

[Pro68d] Yu. V. Prokhorov. An extension of S. N. Bernstein’s inequalities to multidimensional distributions. Theory of Probability and its Applications, 13(2):260–267, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prohorov:1973:SRV

[Pro73] A. V. Prohorov. On sums of random vectors. Theory of Probability and its Applications, 18 (1):186–188, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Proskurin:1974:DNV

[Pro74] G. V. Proskurin. On the distribution of the number of vertices in strata of a random mapping. Theory of Probability and its Applications, 18(4):803–808, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1983:SRA

[Pro83] Yu. V. Prokhorov. Summary of reports of the all-union seminar on scientific methods: “Computational Methods of Mathematical Statistics”. Theory of Probability and its Applications, 27 (3):649–653, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1984:SRV

[Pro84] Yu. V. Prokhorov. On sums of random vectors with values in Hilbert space. Theory of Probability and its Applications, 28 (2):375–379, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1991:LTS

[Pro91a] Yu. V. Prokhorov. Limit theorems for sums of random vectors, whose dimension tends to infinity. Theory of Probability and its Applications, 35(4):755–757, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 751–753.

Prokhorov:1991:SBC

[Pro91b] Yurii Vasil’evich Prokhorov. Sixtieth birthday congratulations. Theory of Probability and its Applications, 35(4):754, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1992:ETV

[Pro92] Yu. V. Prokhorov. An estimate of the total variation of signed measures arising in connection with the Central Limit Theorem. Theory of Probability and its Applications, 36(2):388–390, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 370–372.

Prokhorov:1993:PND

[Pro93] Yu. V. Prokhorov. On polynomials in normally distributed random values. Theory of Probability and its Applications, 37(4): 692–695, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 747–750.

Prokhorov:1994:PRV

[Pro94] Yu. V. Prokhorov. On polynomials in random variables with gamma-distribution. Theory of Probability and its Applications, 38(1):162–166, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 198–202.

Pagurova:1982:DST

[PRR82] V. I. Pagurova, K. D. Rodionov, and M. V. Rodionova. The distribution of the studentized twodimensional range. Theory of Probability and its Applications, 26(2):366–371, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Puri:1987:GRS

[PRR87] M. L. Puri, D. A. Ralescu, and S. S. Ralescu. Gaussian random sets in Banach space. Theory of Probability and its Applications, 31(3):526–529, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:1961:SRA

[PS61] Yu. V. Prokhorov and V. V. Sazonov. Some results associated with Bochner’s theorem. Theory of Probability and its Applications, 6(1):82–87, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Parthasarathy:1964:RID

[PS64] K. R. Parthasarathy and V. V. Sazonov. On the representation of infinitely divisible distributins on locally compact Abelian groups. Theory of Probability and its Applications, 9 (1):108–111, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Puri:1969:ANO

[PS69] M. L. Puri and P. K. Sen. On the asymptotic normality of one sample rank order test statistics. Theory of Probability and its Applications, 14(1):163–167, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1973:BCP

[PS73] E. L. Presman and I. M. Sonin. The best choice problem for a random number of objects. Theory of Probability and its Applications, 17(4):657–668, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Proschan:1975:SMI

[PS75] F. Proschan and J. Sethuraman.

Simple multivariate inequalities using association. Theory of Probability and its Applications, 20(1):193–195, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Presman:1976:EPG

[PS76] E. L. Presman and I. M. Sonin.

Equilibrium points in a game related to the best choice problem. Theory of Probability and its Applications, 20(4):770–781, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokopenko:1977:FES

[PS77] L. N. Prokopenko and A. V. Skorokhod. On flows of events in semi-ordered spaces. Theory of Probability and its Applications, 21(3):523–533, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Panaretos:1985:SCP

[PS85] J. Panaretos and R. Shimizu.

On the stability of a characterization of the Poisson distribution. Theory of Probability and its Applications, 29(4):787–791, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1986:RIL

[PS86] I. F. Pinelis and A. I. Sakhanenko. Remarks on inequalities for large deviation probabilities. Theory of Probability and its Applications, 30(1):143–148, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Perelmuter:1988:EOP

[PS88] M. A. Perel’muter and Yu. A. Semenov. Elliptic operators preserving probability. Theory of Probability and its Applications, 32(4):718–721, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 32(4), (1987), pp. 786–789.

Puri:1989:RCN

[PS89] M. L. Puri and M. Seoh. On the rate of convergence in normal approximation and large deviation probabilities for a class of statistics. Theory of Probability and its Applications, 33(4):682–697, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 735–750.

Pozin:1992:TSE

[PS92a] S. M. Pozin and L. A. Sakhnovich. Two-sided estimation of the smallest eigenvalue of an operator characterizing stable processes. Theory of Probability and its Applications, 36(2):385–388, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 368–370.

Puri:1992:HSV

[PS92b] M. L. Puri and V. V. Sazonov.

On Hilbert-space-valued Ustatistics. Theory of Probability and its Applications, 36(3): 619–621, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 604–605.

Pergamenshchikov:1993:SEP

[PS93] S. M. Pergamenshchikov and A. N. Shiryaev. Sequential estimation of the parameter of a stochastic difference equation with random coefficients. Theory of Probability and its Applications, 37(3):449–470, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 482–501.

Peskir:1998:BFP

[PS98] G. Peškir and A. N. Shiryaev.

On the Brownian first-passage time over a one-sided stochastic boundary. Theory of Probability and its Applications, 42 (3):444–453, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 42(3), (1997), pp. 591–602.

Polyak:1991:AOT

[PT91] B. T. Polyak and A. B. Tsybakov. Asymptotic optimality of the C_{p}-test for the orthogonal series estimation of regression. Theory of Probability and its Applications, 35(2):293–306, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 305–317.

Polyak:1993:FAO

[PT93] B. T. Polyak and A. B. Tsybakov. A family of asymptotically optimal methods for choosing the order of a projective regression estimate. Theory of Probability and its Applications, 37(3):471–481, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 502–512.

Pavlov:1985:UED

[PU85a] I. V. Pavlov and I. A. Ushakov. Unbiased estimation of a distribution function from a multiply truncated sample. Theory of Probability and its Applications, 29(3):617–620, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1985:EMS

[PU85b] I. F. Pinelis and S. A. Utev. Estimates of the moments of sums of independent random variables. Theory of Probability and its Applications, 29(3): 574–577, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:1990:EEB

[PU90] I. S. Pinelis and S. A. Utev.

Exact exponential bounds for sums of independent random variables. Theory of Probability and its Applications, 34(2): 340–346, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 384–390.

Pukhalsky:1982:MSS

[Puk82] E. A. Pukhal’sky. Minimal sufficient statistics for normal models with algebraic structure. Theory of Probability and its Applications, 26(3):564–572, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pukhalskii:1994:TLD

[Puk94] A. A. Pukhal′skiĭ. On the theory of large deviations. Theory of Probability and its Applications, 38(3):490–497, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 553–562.

Purtukhiya:1989:FEM

[Pur89] O. G. Purtukhiya. On the filtering equations of a multidimensional diffusion process (increasing coefficients). Theory of Probability and its Applications, 33 (3):574–578, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 616–621.

Pushkin:1992:BNV

[Pus92] L. N. Pushkin. Borel normal vectors on a manifold in R^{N}. Theory of Probability and its Applications, 36(2):391–395, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Pusz:1993:CED

[Pus93] J. Pusz. Characterization of exponential distributions by conditional moments. Theory of Probability and its Applications, 37(3):557–560, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 599–601.

Pushkin:1997:EPS

[Pus97] L. N. Pushkin. Ergodic properties of sets defined by frequencies of numbers. Theory of Probability and its Applications, 41 (3):593–597, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 672–677.

Parthasarathy:1964:ESS

[PV64] K. R. Parthasarathy and S. R. S.

Varadhan. Extension of stationary stochastic processes. Theory of Probability and its Applications, 9(1):65–71, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pap:1997:AIN

[PvZ97] G. Pap and M. van Zuijlen.

Asymptotic inference for nearly unstable multidimensional AR processes. Theory of Probability and its Applications, 41(3): 578–586, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 703–710.

Puri:1987:ONA

[PW87] M. L. Puri and T.-J. Wu. The order of normal approximation for signed linear rank statistics. Theory of Probability and its Applications, 31(1):145–151, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Popov:1969:SQS

[PY69] Yu. D. Popov and M. I. Yadrenko. Some questions of spectral theory for homogeneous and isotropic random fields. Theory of Probability and its Applications, 14(3):510–517, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Postnikova:1978:CF

[PY78] L. P. Postnikova and A. A. Yudin. On the concentration function. Theory of Probability and its Applications, 22(2): 362–366, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Postnikova:1979:SFI

[PY79] L. P. Postnikova and A. A. Yudin. A sharper form of the inequality for the concentration function. Theory of Probability and its Applications, 23(2): 359–362, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Postnikov:1981:ECF

[PY81] A. G. Postnikov and A. A. Yudin. On estimating the concentration function for a sum of independent identically distributed twodimensional lattice random vectors. Theory of Probability and its Applications, 26(1):152–156, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Postnikov:1988:EMP

[PY88] A. G. Postnikov and A. A. Yudin. Estimating the maximum probability of a sum of independent vectors. Theory of Probability and its Applications, 32 (2):331–334, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Quang:1997:SLL

[QT97] Nguyen Van Quang and Nguyen Duy Tien. The strong law of large numbers for d-dimensional arrays in von Neumann algebras. Theory of Probability and its Applications, 41(3):569–578, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 693–702.

Quas:1999:EEC

[Qua99] A. N. Quas. An entropy estimator for a class of infinite alphabet processes. Theory of Probability and its Applications, 43(3): 496–507, September 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97710.

Rachev:1985:MKM

[Rac85a] S. T. Rachev. The Monge– Kantorovich mass transference problem and its stochastic applications. Theory of Probability and its Applications, 29 (4):647–676, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rachev:1985:CMF

[Rac85b] S. T. Rachev. On a class of minimal functionals on a space of probability measures. Theory of Probability and its Applications, 29(1):41–49, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rachev:1986:UWV

[Rac86] S. T. Rachev. Uniformity in weak and vague convergence. Theory of Probability and its Applications, 30(3):573–576, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Radchenko:1992:IRR

[Rad92] V. N. Radchenko. Integrals with respect to random measures and random linear functionals. Theory of Probability and its Applications, 36(3):621–623, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 594–596.

Radchenko:1997:DIR

[Rad97] V. N. Radchenko. On a definition of the integral of a random function. Theory of Probability and its Applications, 41(3):597–601, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 677–682.

Radchenko:1998:SCC

[Rad98] V. N. Radchenko. Short communications: Convergence of integrals of unbounded real functions in random measures. Theory of Probability and its Applications, 42(2):310–314, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97617.

Rafajlowicz:1990:MCE

[Raf90] E. Rafajłowicz. Minimum cost experimental design with a prescribed information matrix. Theory of Probability and its Applications, 34(2):367–370, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 412–416.

Ragimov:1986:DFE

[Rag86] F. G. Ragimov. On the density of the first exit time of a stable process across a nonlinear boundary. Theory of Probability and its Applications, 30 (2):422–427, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ragimov:1991:LPI

[Rag91] Ph. G. Ragimov. On the local probabilities of intersection of nonlinear boundaries by sums of independent random variables. Theory of Probability and its Applications, 35(2):384–387, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 373–376.

Ragimov:1993:AED

[Rag93] F. G. Ragimov. Asymptotic expansion for the distribution of nonlinear boundary crossing time. Theory of Probability and its Applications, 37(3):560– 564, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 580–583.

Rajput:1973:SGM

[Raj73] Balram S. Rajput. The support of Gaussian measures on Banach spaces. Theory of Probability and its Applications, 17(4):728–734, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rakhimov:1987:CBP

[Rak87] I. Rakhimov. Critical branching processes with infinite variance and decreasing immigration. Theory of Probability and its Applications, 31(1):88–100, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rakhimov:1988:LTR

[Rak88] I. Rakhimov. A limit theorem for random sums of dependent indicators and its applications in the theory of branching processes. Theory of Probability and its Applications, 32(2):290–298, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rakhimov:1989:LLT

[Rak89] I. Rakhimov. Local limit theorems for critical Galton–Watson processes with decreasing immigration. Theory of Probability and its Applications, 33(2): 365–369, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rakhimov:1990:ALT

[Rak90] I. Rakhimov. Addendum: a limit theorem for random sums of dependent indicators and its applications in the theory of branching processes. Theory of Probability and its Applications, 34(3):554, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 613–613.

Rakhimov:1993:GBP

[Rak93] I. Rakhimov. General branching processes with reproduction dependent immigration. Theory of Probability and its Applications, 37(3):482–496, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 513–525.

Ralescu:1997:BKL

[Ral97] S. S. Ralescu. A Bahadur– Kiefer law for the Nadaraya empiric-quantile processes. Theory of Probability and its Applications, 41(2):296–306, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97550. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 380–392.

Rao:1963:ISP

[Rao63a] M. M. Rao. Inference in stochastic processes. I. Theory of Probability and its Applications, 8 (3):266–281, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rao:1963:LLN

[Rao63b] R. Ranga Rao. The law of large numbers for D[0, 1]-valued random variables. Theory of Probability and its Applications, 8 (1):70–74, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rao:1971:LFG

[Rao71] M. M. Rao. Local functionals and generalized random fields with independent values. Theory of Probability and its Applications, 16(3):457–473, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rao:1976:CSP

[Rao76] B. L. S. Prakasa Rao. Characterization of stochastic processes determined up to shift. Theory of Probability and its Applications, 20(3):623–626, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rao:1989:LIL

[Rao89] Prakasa Rao. Law of iterated logarithm for fluctuation of posterior distributions for a class of diffusion processes and a sequential test of power one. Theory of Probability and its Applications, 33(2):269–275, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rasulov:1977:AEE

[Ras77] N. P. Rasulov. On asymptotically efficient estimates of regression coefficients under spectral density of noise degeneration. Theory of Probability and its Applications, 21(2):316–324, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Razinkova:1970:VPG

[Raz70] T. L. Razinkova. On a variant of the “Petersburg Game”. Theory of Probability and its Applications, 15(3):538–540, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rebolledo:1988:WST

[Reb88] R. Rebolledo. Weakening the Skorokhod topology for measure valued processes. Theory of Probability and its Applications, 32 (1):157–159, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Renyi:1956:CPS

[Rén56] A. Rényi. On conditional probability spaces generated by a dimensionally ordered set of measures. Theory of Probability and its Applications, 1(1):55–64, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Revesz:1994:CP

[Rév94]  P. Révész.  Covering prob-

lems. Theory of Probability and its Applications, 38(2):367–379, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 439–453.

Reznik:1968:LIL

[Rez68] M. Kh. Reznik. The law of the iterated logarithm for some classes of stationary processes. Theory of Probability and its Applications, 13(4):606–621, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Reznik:1971:ALI

[Rez71] M. Kh. Reznik. Addendum: The law of the iterated logarithm for some classes of stationary processes. Theory of Probability and its Applications, 16(2):397, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Reznichenko:1977:ISC

[Rez77a] S. V. Reznichenko. Investigation of a system of coupled-mode equations with small random perturbations. Theory of Probability and its Applications, 22 (1):124–128, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Reznichenko:1977:RWD

[Rez77b] S. V. Reznichenko. Radio waveguides and diffusion processes on differentiable manifolds. Theory of Probability and its Applications, 21(2):357–369, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Reznichenko:1978:SFM

[Rez78] S. V. Reznichenko. The structure of the fundamental matrix of a system of coupled-mode equations. Theory of Probability and its Applications, 22(2):311–325, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Reznichenko:1980:AIS

[Rez80] S. V. Reznichenko. Addendum: Investigation of a system of coupled-mode equations with small random perturbations. Theory of Probability and its Applications, 25(1):214, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Reznikova:1981:CLT

[Rez81] A. Ya. Reznikova. The Central Limit Theorem for the spectrum of random Jacobi matrices. Theory of Probability and its Applications, 25(3):504–513, ???? 1981. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1978:ROR

[RF78] B. A. Rogozin and S. G. Foss. Recurrency of an oscillating random walk. Theory of Probability and its Applications, 23(1):155–162, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Richter:1957:LLT

[Ric57] Wolfgang Richter. Local limit theorems for large deviations. Theory of Probability and its Applications, 2(2):206–220, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Richter:1958:MDL

[Ric58] V. Richter. Multi-dimensional local limit theorems for large deviations. Theory of Probability and its Applications, 3(1): 100–106, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Richter:1964:MDL

[Ric64]  Wolfgang Richter.  Multi-

dimensional limit theorems for large deviations and their application to the χ^{2} distribution. Theory of Probability and its Applications, 9(1):28–37, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Richter:1965:LTS

[Ric65] Wolfgang Richter. Limit theorems for sequences of random variables with sequences of random indices. Theory of Probability and its Applications, 10 (1):74–84, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Richter:1968:ESS

[Ric68] Wolfgang Richter. On ergodicity of sequences of sums. Theory of Probability and its Applications, 13(2):334–338, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Riedel:1977:NVC

[Rie77] M. Riedel. A new version of the Central Limit Theorem. Theory of Probability and its Applications, 22(1):183–184, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Riedel:1980:CPQ

[Rie80] M. Riedel. On a characterization problem in the queueing model GI|G|1. Theory of Probability and its Applications, 25 (1):128–138, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Riedel:1981:DSP

[Rie81] M. Riedel. Determination of a stochastic process by means of stochastic integrals. Theory of Probability and its Applications, 25(2):335–346, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rihaoui:1985:TTZ

[Rih85] I. Rihaoui. Two types of zeroone law for a Gaussian probability. Theory of Probability and its Applications, 29(2):411–412, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rivlin:1996:APT

[Riv96] M. I. Rivlin. Asymptotic properties of thinning random sets. Theory of Probability and its Applications, 40(3):484–494, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 542–555.

Rasulov:1979:RPT

[RK79] N. P. Rasulov and A. S. Kholevo. A regression problem for timecontinuous processes. Theory of Probability and its Applications, 23(4):731–740, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Robinson:1963:PWD

[Rob63] Enders A. Robinson. Properties of the Wold decomposition of stationary stochastic processes. Theory of Probability and its Applications, 8(2):187–195, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rodionov:1988:DOC

[Rod88] K. D. Rodionov. On the detection of outliers in the case of an exponential distribution. Theory of Probability and its Applications, 32(3):553–556, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1958:SPF

[Rog58] B. A. Rogozin. Some problems in the field of limit theorems. Theory of Probability and its Applications, 3(2):174–184, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1960:REP

[Rog60]  B. A. Rogozin.  A remark

on Esséen’s paper “A Moment Inequality with an Application to the Central Limit Theorem”. Theory of Probability and its Applications, 5(1):114–117, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1961:ECF

[Rog61a] B. A. Rogozin. An estimate for concentration functions. Theory of Probability and its Applications, 6(1):94–97, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1961:IDS

[Rog61b] B. A. Rogozin. On the increase of dispersion of sums of independent random variables. Theory of Probability and its Applications, 6(1):97–99, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1964:DFJ

[Rog64] B. A. Rogozin. On the distribution of the first jump. Theory of Probability and its Applications, 9(3):450–465, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1965:SCP

[Rog65] B. A. Rogozin. On some classes of processes with independent increments. Theory of Probability and its Applications, 10 (3):479–483, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1966:DFR

[Rog66a] B. A. Rogozin. On distributions of functionals related to boundary problems for processes with independent increments. Theory of Probability and its Applications, 11(4):580–591, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1966:SEP

[Rog66b] B. A. Rogozin. Some extremal problems in the theory of mass service. Theory of Probability and its Applications, 11(1):144–151, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1966:SCD

[Rog66c] B. A. Rogozin. Speed of convergence of the distribution of the maximum of sums of independent random variables to a limit distribution. Theory of Probability and its Applications, 11(3):438–441, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1968:LBP

[Rog68a] B. A. Rogozin. Local behavior of processes with independent increments. Theory of Probability and its Applications, 13 (3):482–486, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1968:EEU

[Rog68b] B. A. Rogozin. On the existence of exact upper sequences. Theory of Probability and its Applications, 13(4):667–672, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1969:RTD

[Rog69] B. A. Rogozin. A remark on a theorem due to W. Feller. Theory of Probability and its Applications, 14(3):529, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1971:DFL

[Rog71] B. A. Rogozin. The distribution of the first ladder moment and height and fluctuation of a random walk. Theory of Probability and its Applications, 16 (4):575–595, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1972:BJP

[Rog72] B. A. Rogozin. The behavior of a jump for processes with independent increments. Theory of Probability and its Applications, 17(1):146–149, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1973:DFH

[Rog73] B. A. Rogozin. The distribution of the first hit for stable and asymptotically stable walks on an interval. Theory of Probability and its Applications, 17 (2):332–338, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1974:ERT

[Rog74] B. A. Rogozin. An estimate of the remainder term in limit theorems of renewal theory. Theory of Probability and its Applications, 18(4):662–677, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1977:ARF

[Rog77a] B. A. Rogozin. Asymptotics of renewal functions. Theory of Probability and its Applications, 21(4):669–686, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1977:RSW

[Rog77b] B. A. Rogozin. Relatively stable walks. Theory of Probability and its Applications, 21(2):375–379, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1988:EMC

[Rog88a] B. A. Rogozin. Estimation of the maximum of a convolution of bounded densities. Theory of Probability and its Applications, 32(1):48–56, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1988:ICF

[Rog88b] B. A. Rogozin. Inequalities for concentration functions of convolutions of arithmetic distributions and distributions with bounded densities. Theory of Probability and its Applications, 32(2):325–329, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:1992:LSP

[Rog92] B. A. Rogozin. Ladder subordinators for processes with independent increments. Theory of Probability and its Applications, 36(3):623–629, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 596–601.

Rogozin:1994:ICD

[Rog94] B. A. Rogozin. Inequalities for concentration of a decomposition. Theory of Probability and its Applications, 38(3):556–562, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 645–652.

Rokhlin:1961:EAC

[Rok61] V. A. Rokhlin. On the entropy of an automorphism of a compact commutative group. Theory of Probability and its Applications, 6(3):322–323, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanovskii:1959:OTR

[Rom59] I. V. Romanovskii. On one theorem of R. Bellman. Theory of Probability and its Applications, 4(4):420–421, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanovskii:1961:GTR

[Rom61] I. V. Romanovskii. Game-type random walks. Theory of Probability and its Applications, 6 (4):393–396, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanovskii:1962:MTG

[Rom62] I. V. Romanovskii. Minimax theorems for games with imperfect transfer of information. Theory of Probability and its Applications, 7(1):87–93, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanovskii:1964:EFS

[Rom64] I. V. Romanovskii. On evaluating final states in inventory processes. Theory of Probability and its Applications, 9(1): 149–156, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanovskii:1965:EOS

[Rom65] I. V. Romanovskii. Existence of an optimal stationary policy in a Markov decision process. Theory of Probability and its Applications, 10(1):120–122, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanovskaya:1969:RWS

[Rom69] I. L. Romanovskaya. A remark on Wald’s sequential probability ratio test. Theory of Probability and its Applications, 14 (1):167–170, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanov:1987:CFL

[Rom87] V. A. Romanov. Conditional functional limit theorems for randomized walks. Theory of Probability and its Applications, 31 (4):726–729, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1980:LTS

[Ron80] A. F. Ronzhin. On limit theorems for sums of dependent random variables. Theory of Probability and its Applications, 24 (3):549–559, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1982:AFM

[Ron82] A. F. Ronzhin. Asymptotic formulas for the moments of Ustatistics with degenerate kernel. Theory of Probability and its Applications, 27(1):49–58, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1984:TSM

[Ron84] A. F. Ronzhin. On two statistics in the multinomial scheme. Theory of Probability and its Applications, 28(4):800–805, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1985:LDC

[Ron85] A. F. Ronzhin. The limit distribution for a chi-square process with disorder. Theory of Probability and its Applications, 29 (3):613–617, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1986:CEG

[Ron86a] A. F. Ronzhin. Chernoff efficiency for goodness-of-fit tests based on empirical distribution functions. Theory of Probability and its Applications, 30(2): 404–407, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1986:FLT

[Ron86b] A. F. Ronzhin. A functional limit theorem for homogeneous U-statistics with a degenerate kernel. Theory of Probability and its Applications, 30(4):806–810, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1986:PES

[Ron86c] A. F. Ronzhin. On the probability of exit from a square-root boundary by the norm of a Brownian motion. Theory of Probability and its Applications, 30 (1):160–162, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1988:LTD

[Ron88] A. F. Ronzhin. Limit theorems for the disruption problem of a sequence of independent random variables. Theory of Probability and its Applications, 32 (2):282–289, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1989:GFT

[Ron89] A. F. Ronzhin. Goodness-of-fit tests for generalized urn schemes based on separable statistics. Theory of Probability and its Applications, 33(1):86–95, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Roos:1999:MMP

[Roo99]  B. Roos.  Metric multivari-

ate Poisson approximation of the generalized multinomial distribution. Theory of Probability and its Applications, 43 (2):306–316, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97692.

Rosenblatt:1964:EMO

[Ros64] M. Rosenblatt. Equicontinuous Markov operators. Theory of Probability and its Applications, 9(2):180–197, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rosanov:1967:GFG

[Ros67] Yu. A. Rosanov. On Gaussian fields with given conditional distributions. Theory of Probability and its Applications, 12 (3):381–391, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rossberg:1975:PKC

[Ros75] H.-J. Rossberg. On a problem of Kolmogorov concerning the normal distribution. Theory of Probability and its Applications, 19 (4):795–798, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rossberg:1980:LTI

[Ros80] H.-J. Rossberg. Limit theorems for identically distributed summands asuming the convergence of the distribution functions on a half axis. Theory of Probability and its Applications, 24 (4):693–711, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rossberg:1991:PDP

[Ros91] H. J. Rossberg. Positive definite probability densities. Theory of Probability and its Applications, 35(1):169–174, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 179–184.

Rossberg:1995:LTI

[Ros95] H.-J. Rossberg. Limit theorems involving restricted convergence. Theory of Probability and its Applications, 39(2):298–314, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 334–356.

Rotar:1967:MVT

[Rot67] V. I. Rotar’. On moments of the value and the time of the first passage over a curvilinear boundary. Theory of Probability and its Applications, 12(4):690–695, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1970:NUE

[Rot70a] V. I. Rotar’. A non-uniform estimate for the convergence speed in the multi-dimensional central theorem. Theory of Probability and its Applications, 15(4):

630–648, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1970:SCM

[Rot70b] V. I. Rotar’. On the speed of convergence in the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 15(2):354–356, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1974:SLT

[Rot74] V. I. Rotar’. Some limit theorems for polynomials of second degree. Theory of Probability and its Applications, 18(3):499–507, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1976:SRS

[Rot76a] R. I. Rotar’. Some remarks on summing independent variables in the non-classical case. Theory of Probability and its Applications, 21(1):130–137, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1976:LTM

[Rot76b] V. I. Rotar’. Limit theorems for multilinear forms and quasipolynomial functions. Theory of Probability and its Applications, 20(3):512–532, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1976:DQF

[Rot76c] V. I. Rotar’. On the distribution of a quadratic form in many random variables. Theory of Probability and its Applications, 20 (4):880–882, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1978:NCEa

[Rot78a] V. I. Rotar’. Non-classical estimates of the rate of convergence in the multi-dimensional Central Limit Theorem. I. Theory of Probability and its Applications, 22(4):755–772, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1978:NCEb

[Rot78b] V. I. Rotar’. Non-classical estimates of the rate of convergence in the multi-dimensional Central Limit Theorem. II. Theory of Probability and its Applications, 23(1):50–62, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Roy:1993:PPS

[Roy93] R. Roy. Percolation of Poisson sticks on the plane. Theory of Probability and its Applications, 37(1):167–168, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 203–204.

Rozanov:1957:LLT

[Roz57] Yu. A. Rozanov. On a local limit theorem for lattice distributions. Theory of Probability and its Applications, 2(2): 260–265, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1959:EGS

[Roz59a] Yu. A. Rozanov. On the extrapolation of generalized stationary random processes. Theory of Probability and its Applications, 4(4):426–431, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1959:SAA

[Roz59b] Yu. A. Rozanov. Spectral analysis of abstract functions. Theory of Probability and its Applications, 4(3):271–287, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1960:CLT

[Roz60a] Yu. A. Rozanov. A Central Limit Theorem for additive random functions. Theory of Probability and its Applications, 5 (2):221–223, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1960:SPM

[Roz60b] Yu. A. Rozanov. Spectral properties of multivariate stationary processes and boundary properties of analytic matrices. Theory of Probability and its Applications, 5(4):362–376, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1961:ACL

[Roz61] Yu. A. Rozanov. On the applicability of the Central Limit Theorem to stationary processes which have passed through a linear filter. Theory of Probability and its Applications, 6 (3):321–322, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1962:DOG

[Roz62] Yu. A. Rozanov. On the density of one Gaussian measure with respect to another. Theory of Probability and its Applications, 7(1):82–87, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1963:PEP

[Roz63] Yu. Rozanov. On the problem of the equivalence of probability measures corresponding to stationary Gaussian processes. Theory of Probability and its Applications, 8(3):223–231, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1964:PMF

[Roz64a] Yu. A. Rozanov. On probability measures in functional spaces corresponding to stationary Gaussian processes. Theory of Probability and its Applications, 9(3):404–420, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1964:SSL

[Roz64b] Yu. A. Rozanov. On the stability of solutions of linear problems for stationary processes. Theory of Probability and its Applications, 9(3):477–479, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1966:DGD

[Roz66a] Yu. A. Rozanov. On the densities of Gaussian distributions and Wiener–Hopf integral equations. Theory of Probability and its Applications, 11(1):152–160, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1966:SRP

[Roz66b] Yu. A. Rozanov. Some remarks on the paper “On Gaussian Distribution Densities and Wiener– Hopf Integral Equations”. Theory of Probability and its Applications, 11(3):483–485, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozental:1967:SGB

[Roz67] A. L. Rozental’. Some generalisations of the boundary value problem with oblique derivative. Theory of Probability and its Applications, 12(2):329–333, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozental:1973:CBM

[Roz73] A. L. Rozental’. On conditional Brownian motions with oblique reflection corresponding to inaccessible singular points. Theory of Probability and its Applications, 17(3):528–535, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1976:AEC

[Roz76] L. V. Rozovskii. Asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 20(4): 794–804, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1977:MER

[Roz77] Yu. A. Rozanov. On Markov extensions of a random process. Theory of Probability and its Applications, 22(1):190–195, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1978:ABR

[Roz78] L. V. Rozovskii. On the asymptotic behavior of the remainder term in the Central Limit Theorem. Theory of Probability and its Applications, 23(1):106–116, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1979:PER

[Roz79] L. V. Rozovskii. On the precision of an estimate of the remainder term in the Central Limit Theorem. Theory of Probability and its Applications, 23(4):712–730, ???? 1979. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1980:CDD

[Roz80] B. L. Rozovskii. On conditional distributions of degenerate diffusion processes. Theory of Probability and its Applications, 25 (1):147–151, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1981:SPL

[Roz81a] Ya. A. Rozanov. Some problems in the linear theory of random functions. Theory of Probability and its Applications, 25 (4):689–702, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1981:CRS

[Roz81b] L. V. Rozovskii. On the convergence rate in the strong law of large numbers. Theory of Probability and its Applications, 26 (1):135–140, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1982:LTL

[Roz82] L. V. Rozovskii. Limit theorems on large deviations in a narrow zone. Theory of Probability and its Applications, 26(4):834–845, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1984:MKP

[Roz84a] Yu. A. Rozanov. On the Markov–Kolmogorov principle for stochastic differential equations. Theory of Probability and its Applications, 28(2):383–388, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1984:FSP

[Roz84b] B. L. Rozovskii. Filtering, smoothing and prediction of degenerate diffusion processes. backward equations. Theory of Probability and its Applications, 28(4):762–774, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovskii:1986:SEP

[Roz86] L. V. Rozovskii. Some estimates for probabilities of onesided large deviations. Theory of Probability and its Applications, 30(4):850–855, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(4), (1985), pp. 800–804.

Rozovskii:1987:AAL

[Roz87] L. V. Rozovskii. On the accuracy of approximation in limit theorems for large deviations. Theory of Probability and its Applications, 31(2):255–268, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1988:MRF

[Roz88] Yu. A. Rozanov. Markov random fields and boundary problems for stochastic partial differential equations. Theory of Probability and its Applications, 32(1):1–29, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1990:TGR

[Roz90a] Yu. A. Rozanov. On the theory of generalized random functions. Theory of Probability and its Applications, 34(1):197–200, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(1), (1989), pp. 228–231.

Rozovskii:1990:CRW

[Roz90b] L. V. Rozovskiĭ. The convergence rate in the weak law of large numbers. Theory of Probability and its Applications, 34(2):348–351, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 392–395.

Rozovskii:1990:PLD

[Roz90c] L. V. Rozovskiĭ. Probabilities of large deviations of sums of independent random variables with common distribution function in the domain of attraction of the normal law. Theory of Probability and its Applications, 34 (4):625–644, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 686–705.

Rozanov:1991:SBV

[Roz91] Yu. A. Rozanov. Some boundary value problems for generalized random fields. Theory of Probability and its Applications, 35(4):707–724, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 625–641.

Rozanov:1994:ERF

[Roz94a] Yu. A. Rozanov. On evolution of random fields with an ultra unbounded stochastic source. Theory of Probability and its Applications, 38(2):316–329, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 356–373.

Rozovskii:1994:PLD

[Roz94b] L. V. Rozovskiĭ. Probabilities of large deviations on the whole axis. Theory of Probability and its Applications, 38 (1):53–79, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 79–109.

Rozanov:1996:SSS

[Roz96] Yu. A. Rozanov. Stochastic Sobolev spaces and their boundary trace. Theory of Probability and its Applications, 40 (1):104–115, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozanov:1997:HSM

[Roz97] Yu. A. Rozanov. On a Hilbert space method for estimating mean values. Theory of Probability and its Applications, 41 (2):387–393, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 459–467.

Rozovskii:1998:GKT

[Roz98a] L. V. Rozovskiĭ. A generalization of the Kolmogorov theorem on the law of the iterated logarithm. Theory of Probability and its Applications, 42(1):112–120, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97597. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 134–143.

Rozovskii:1998:PLD

[Roz98b] L. V. Rozovskiĭ. Probabilities of large deviations for sums of independent random variables with a common distribution functionfrom the domain of attraction of an asymmetric stable law. Theory of Probability and its Applications, 42(3):454–482, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 496–530.

Rozanov:1999:PLA

[Roz99a] Yu. A. Rozanov. On piecewise linear approximation for nonlinear stochastic evolution. Theory of Probability and its Applications, 43(1):146–151, March 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97675.

Rozovsky:1999:CSC

[Roz99b] L. V. Rozovsky. On the Cramér series coefficients. Theory of Probability and its Applications, 43(1):152–157, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97676.

Rosenblatt-Roth:1964:CEP

[RR64] M. Rosenblatt-Roth. The concept of entropy in probability theory and its application in the theory of information transmission through communication channels. Theory of Probability and its Applications, 9(2): 212–235, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rachev:1991:TPM

[RR91] S. T. Rachev and L. Ruschendorf. A transformation property of minimal metrics. Theory of Probability and its Applications, 35(1):110–117, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rachev:1993:RCS

[RR93] S. T. Rachev and L. Rüschendorf. Rate of convergence for sums and maxima and doubly ideal metrics. Theory of Probability and its Applications, 37(2):222–235, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 276–289.

Rachev:1995:MOP

[RR95] S. T. Rachev and L. Ruschendorf. Models for option prices. Theory of Probability and its Applications, 39(1):120–152, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 150–190.

Rinott:1999:SBR

[RR99] Y. Rinott and V. Rotar. Some bounds on the rate of convergence in the CLT for martingales. I. Theory of Probability and its Applications, 43(4): 604–619, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97714.

Ryzhov:1970:HBP

[RS70] Yu. M. Ryzhov and A. V. Skorokhod. Homogeneous branching processes with a finite number of types and continuously varying mass. Theory of Probability and its Applications, 15(4):704–707, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rossberg:1976:CCC

[RS76a] H.-J. Rossberg and G. Siegel.

Continuation of convergence in the Central Limit Theorem. Theory of Probability and its Applications, 20(4):866–868, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rychlik:1976:CRC

[RS76b] Z. Rychlik and D. Szynal. Convergence rates in the Central Limit Theorem for sums of a random number of independent random variables. Theory of Probability and its Applications, 20 (2):351–362, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rychlik:1980:RAR

[RS80] Z. Rychlik and D. Szynal. On the rate of approximation in the random-sum Central Limit Theorem. Theory of Probability and its Applications, 24(3):620–625, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rossberg:1982:OSC

[RS82] H.-J. Rossberg and G. Siegel.

One-sided characterization of the normal distribution in the set of infinitely divisible distributions. Theory of Probability and its Applications, 26(2):392–399, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronzhin:1983:LTM

[RS83] A. F. Ronzhin and B. I. Selivanov. Limit theorems for a Markov time in a disorder problem. Theory of Probability and its Applications, 27(3):615–619, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1986:SED

[RS86] V. I. Rotar’ and T. L. Shervashidze. Some estimates of distributions of quadratic forms. Theory of Probability and its Applications, 30(3):585–590, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rychlik:1988:IPM

[RS88] Z. Rychlik and I. Szyszkowski. The invariance principle for φmixing sequences. Theory of Probability and its Applications, 32(3):559–562, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rotar:1993:CCC

[RS93] V. I. Rotar′ and A. G. Sholomitskiĭ. A condition for convergence of convolutions. Theory of Probability and its Applications, 37(2):398–401, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 410–413.

Rao:1994:PLT

[RS94] M. M. Rao and V. V. Sazonov.

A projective limit theorem for probability spaces and applications. Theory of Probability and its Applications, 38(2):307–315, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 345–355.

Regazzini:1998:CLP

[RS98] E. Regazzini and V. V. Sazonov.

On the central limit problem for partially exchangeable random variables with values in a Hilbert space. Theory of Probability and its Applications, 42(4): 656–670, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97652.

Ruschendorf:1998:CSS

[RT98] L. Rüschendorf and W. Thomsen. Closedness of sum spaces and the generalized Schrödinger problem. Theory of Probability and its Applications, 42(3): 483–494, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 576–590.

Rubinshtein:1957:SEP

[RU57] G. Sh. Rubinshtein and K. Urbanik. Solution of an extremal problem. Theory of Probability and its Applications, 2(3): 364–366, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ronkin:1988:DSC

[RU88] A. L. Ronkin and A. M. Ulanovskii. On determining a sum of close distributions from their convolution values. Theory of Probability and its Applications, 32(4):730–734, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rudzkis:1993:DST

[Rud93] R. Rudzkis. On the distribution of supremum-type functionals of nonparametric estimates of probability and spectral densities. Theory of Probability and its Applications, 37(2):236–249, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 254–267.

Ruggeri:1993:BPP

[Rug93] F. Ruggeri. Bounds on the prior probability of a set and robust Bayesian analysis. Theory of Probability and its Applications, 37(2):358–359, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 37(2), (1992), pp. 389–390.

Rukhin:1967:CND

[Ruk67] A. L. Rukhin. The complex normal distribution and admissibility of the sample average as an estimate of a location parameter. Theory of Probability and its Applications, 12(4):695–697, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rukhin:1976:ISC

[Ruk76] A. L. Rukhin. Invariant statistics and characterizations of probability distributions. Theory of Probability and its Applications, 20(3):584–598, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rusakov:1996:FLT

[Rus96] O. V. Rusakov. A functional limit theorem for random variables with strong residual dependence. Theory of Probability and its Applications, 40(4):714–728, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 813–832.

Ryasin:1966:OSS

[Rya66] V. A. Ryasin. Optimal singleshot correction for a model problem. Theory of Probability and its Applications, 11(4):626–632, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryasin:1976:OTS

[Rya76] V. A. Ryasin. Optimal times and sizes of control impulses for corrections depending on trajectory measurements. Theory of Probability and its Applications, 20 (3):499–511, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryazanov:1979:SRG

[Rya79] B. V. Ryazanov. Spectra of random Gaussian g-cyclic matrices. Theory of Probability and its Applications, 23(3):543–558, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rychlik:1984:NUC

[Ryc84] Z. Rychlik. Non-uniform central limit bounds with applications to probabilities of deviations. Theory of Probability and its Applications, 28(4):681–687, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rykov:1966:MDP

[Ryk66] V. V. Rykov. Markov decision processes with finite state and decision spaces. Theory of Probability and its Applications, 11 (2):302–311, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryzhov:1965:LTS

[Ryz65] Yu M. Ryzhov. Limit theorems for stationary Gaussian processes. Theory of Probability and its Applications, 10(1):132–139, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryzhov:1967:LTF

[Ryz67] Yu. M. Ryzhov. Limit theorems for functionals of sample functions of a stationary Gaussian process. Theory of Probability and its Applications, 12 (2):320–323, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryzhov:1969:LDS

[Ryz69] Yu. M. Ryzhov. Limit distributions of some functionals of a stationary Gaussian process. Theory of Probability and its Applications, 14(2):229–243, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryzhov:1971:EMC

[Ryz71] Yu. M. Ryzhov. Estimation of the mean and correlation function of a stationary process from discrete data. Theory of Probability and its Applications, 16 (2):367–369, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ryzhov:1972:AEM

[Ryz72] Yu. Ryzhov. Addendum: Estimation of the mean and correlation function of a stationary process from discrete data. Theory of Probability and its Applications, 17(1):193, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Statulevicius:1994:LDP

[SA94] V. Statulevicius and A. Aleskeviciene. On large deviations in the Poisson approximation. Theory of Probability and its Applications, 38(2):385–393, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 460–470.

Sachkov:1971:DND

[Sac71a] V. N. Sachkov. Distribution of the number of distinct elements of a symmetric basis in a random mA-Sample. Theory of Probability and its Applications, 16 (3):494–505, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sachkov:1971:DNF

[Sac71b] V. N. Sachkov. Distribution of the number of fixed points of elements of a symmetric semigroup under the condition σ^{h+1} = σ^{h} and the number of trees of height not exceeding h. Theory of Probability and its Applications, 16 (4):661–673, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sachkov:1973:MFS

[Sac73a] V. N. Sachkov. Mappings of a finite set with limitations on contours and height. Theory of Probability and its Applications, 17 (4):640–656, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 17, (1972), pp. 679–694.

Sachkov:1973:RMB

[Sac73b] V. N. Sachkov. Random mappings with bounded height. Theory of Probability and its Applications, 18(1):120–130, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sachkov:1974:RPS

[Sac74] V. N. Sachkov. Random partitions of sets. Theory of Probability and its Applications, 19 (1):184–190, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sadikova:1966:SIC

[Sad66a] S. M. Sadikova. Some inequalities for characteristic functions. Theory of Probability and its Applications, 11(3):441–447, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sadikova:1966:TDA

[Sad66b] S. M. Sadikova. Two-dimensional analogues of an inequality of Esseen with applications to the Central Limit Theorem. Theory of Probability and its Applications, 11(3):325–335, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sadikova:1967:ASI

[Sad67a] S. Sadikova. Addendum: Some inequalities for characteristic functions. Theory of Probability and its Applications, 12(2):343, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sadikova:1967:ATD

[Sad67b] S. Sadikova. Addendum: Twodimensional analogues of an inequality of Esseen with applications to the Central Limit Theorem. Theory of Probability and its Applications, 12(2):343, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sadikova:1968:MCL

[Sad68] S. M. Sadikova. On the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 13(1):164–170, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Safonov:1977:CWP

[Saf77] M. V. Safonov. Control of a Wiener process when the number of switches is bounded. Theory of Probability and its Applications, 21(3):593–599, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Safaryan:1992:CCR

[Saf92] R. G. Safaryan. On a certain class of random processes and related systems of differential equations with a small parameter. Theory of Probability and its Applications, 36(4): 722–737, December 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sagitov:1983:CBP

[Sag83] S. M. Sagitov. Critical branching processes with several particle types and immigration. Theory of Probability and its Applications, 27(2):369–374, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sagitov:1986:RCB

[Sag86] S. M. Sagitov. Reduced critical Bellman-harris branching process with several types of particles. Theory of Probability and its Applications, 30(4):783–796, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sagitov:1991:MCB

[Sag91] S. M. Sagitov. A multidimensional critical branching process generated by a large number of particles of one type. Theory of Probability and its Applications, 35(1):118–130, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 98–109.

Sagitov:1996:EBP

[Sag96] S. M. Sagitov. On an explosive branching process. Theory of Probability and its Applications, 40(3):575–577, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 669–671.

Sahanenko:1975:SCB

[Sah75] A. I. Sahanenko. On the speed of convergence in a boundary problem. Theory of Probability and its Applications, 19(2):399–403, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sakovich:1956:SFC

[Sak56] G. N. Sakovich. A single form for the conditions for attraction to stable laws. Theory of Probability and its Applications, 1 (3):322–325, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sakhanenko:1985:LKI

[Sak85] A. I. Sakhanenko. On Lévy– Kolmogorov inequalities for Banach-space-valued random variables. Theory of Probability and its Applications, 29 (4):830–836, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Saltykova:1971:SQR

[Sal71] Z. N. Saltykova. On some questions related to the joint distribution of functionally dependent random variables. Theory of Probability and its Applications, 16(3):533–538, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Salakhutdinov:1979:RRT

[Sal79] R. Z. Salakhutdinov. On refining the remainder term in the Central Limit Theorem. Theory of Probability and its Applications, 23(3):663–667, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Salov:1980:SAT

[Sal80] G. I. Salov. On a stochastic approximation theorem in a Hilbert space and its applications. Theory of Probability and its Applications, 24(2):413–419, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Salov:1990:EPP

[Sal90] G. I. Salov. An equation for the a posteriori probability of disorder in a sequence of dependent random variables and Shiryaevoptimal detection of the time of occurrence of disorder. Theory of Probability and its Applications, 34(4):732–735, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 799–802.

Saleh:1993:SEP

[Sal93a] A. K. Md. E. Saleh. On shrinkage estimation of the parameters of an autoregressive Gaussian process. Theory of Probability and its Applications, 37(2):250–260, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 290–300.

Salikhov:1993:SCI

[Sal93b] N. P. Salikhov. On strengthening Chernoff’s inequality. Theory of Probability and its Applications, 37(3):564–567, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 583–586.

Salikhov:1997:ECF

[Sal97] N. P. Salikhov. An estimate of the concentration function by the Esseen method. Theory of Probability and its Applications, 41(3):504–518, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97526. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 561–577.

Salikhov:1998:AST

[Sal98] N. P. Salikhov. On the absolute significance test for polynomial distribution. Theory of Probability and its Applications, 42(4): 671–683, December 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97647.

Salikhov:1999:OGC

[Sal99] N. P. Salikhov. On one generalization of Chernov’s distance. Theory of Probability and its Applications, 43(2):239–255, June 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97685.

Samandarov:1963:SSH

[Sam63] E. G. Samandarov. Service systems in heavy traffic. Theory of Probability and its Applications, 8(3):307–309, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Samarova:1982:LLH

[Sam82] S. S. Samarova. On the length of the longest head-run for a Markov chain with two states. Theory of Probability and its Applications, 26(3):498–509, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sanov:1966:AMO

[San66] I. N. Sanov. Addendum: a method of obtaining asymptotic expressions for probabilities of large deviations of nonlinear statistics from the multinornial distribution and its applications, p. 662. Theory of Probability and its Applications, 11(3):487, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sapozhnikov:1979:MUB

[Sap79] P. N. Sapozhnikov. Minimax and uniformly best unbiased estimates based on finite statistical structure with quadratic loss. Theory of Probability and its Applications, 23(4):805–811, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sapozhnikov:1988:HSS

[Sap88] P. N. Sapozhnikov. Hilbert statistical structures admitting a nontrivial optimal algebra. Theory of Probability and its Applications, 32(2):365–368, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sapegin:1992:BPW

[Sap92] I. L. Sapegin. On boundary properties of Wiener fields. Theory of Probability and its Applications, 36(2):395–401, May 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 377–382.

Sarymsakov:1961:IMC

[Sar61] T. A. Sarymsakov. Inhomogeneous Markov chains. Theory of Probability and its Applications, 6(2):178–185, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sarmanov:1967:RUG

[Sar67] O. V. Sarmanov. Remarks on uncorrelated Gaussian dependent random variables. Theory of Probability and its Applications, 12(1):124–126, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sarmanov:1970:PCC

[Sar70] O. V. Sarmanov. On a property of the correlation coefficient. Theory of Probability and its Applications, 15(3):544–546, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sarafyan:1973:DPD

[Sar73]  V. V. Sarafyan.  Diffusion

processes and differential equations degenerating at isolated points. Theory of Probability and its Applications, 17(4):699–703, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sarafyan:1990:WCM

[Sar90] V. V. Sarafyan. Weak convergence of measures and asymptotic behavior of solutions of the Cauchy problem for systems of differential equations with a small parameter. Theory of Probability and its Applications, 34 (2):351–357, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 395–402.

Satyabaldina:1973:AER

[Sat73a] K. I. Satyabaldina. Absolute estimates of the rate of convergence to stable laws. Theory of Probability and its Applications, 17 (4):726–728, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Satyabaldina:1973:ERC

[Sat73b] K. I. Satyabaldina. On the estimation of the rate of convergence in a limit theorem with a stable limit law. Theory of Probability and its Applications, 18 (1):202–204, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sataev:1995:AMS

[Sat95a] E. A. Sataev. Asymptotically Markov stochastic 1. Theory of Probability and its Applications, 39(3):448–464, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 567–587.

Sato:1995:MCU

[Sat95b] K. Sato. Multimodal convolutions of unimodal infinitely divisible distributions. Theory of Probability and its Applications, 39(2):336–347, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 39(2), (1994), pp. 403–415.

Sataev:1996:CDF

[Sat96] E. A. Sataev. Continuous dependence of final distributions on transition probabilities of an asymptotically Markov process. Theory of Probability and its Applications, 40(1):186–191, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 183–188.

Saulis:1992:PLD

[Sau92] L. I. Saulis. Probabilities of large deviations for random vectors. Theory of Probability and its Applications, 36(3):494– 507, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 482–493.

Savchuk:1964:EEL

[Sav64] A. P. Savchuk. On estimates for the entropy of a language according to Shannon. Theory of Probability and its Applications, 9(1):138–141, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savchenko:1999:PEL

[Sav99] S. V. Savchenko. On the probability of the existence of a localized basic state for a discrete Schrödinger equation with random potential, perturbed by a compact operator. Theory of Probability and its Applications, 43(1):158–162, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97677.

Sawyer:1993:IPS

[Saw93] S. Sawyer. Inferring population structure from DNA sequences. Theory of Probability and its Applications, 37(1): 168–169, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazonov:1958:RCF

[Saz58] V. Sazonov. A remark on characteristic functionals. Theory of Probability and its Applications, 3(2):188–192, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazonov:1963:GCT

[Saz63] V. V. Sazonov. On the Glivenko– Cantelli theorem. Theory of Probability and its Applications, 8(3):282–286, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazonov:1965:LBM

[Saz65] V. V. Sazonov. On the limiting behavior of the maximal probability in the infinite dimensional case. Theory of Probability and its Applications, 10(4):684–685, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazonov:1966:MCF

[Saz66] V. V. Sazonov. On multidimensional concentration functions. Theory of Probability and its Applications, 11(4):603–609, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazanov:1967:RCM

[Saz67] V. V. Sazanov. On the rate of convergence in the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 12(1):77– 89, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazanov:1969:ICR

[Saz69] V. V. Sazanov. An improvement of a convergence-rate estimate. Theory of Probability and its Applications, 14(4):640–651, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazanov:1975:EMS

[Saz75] V. V. Sazanov. On the estimation of moments of sums of independent random variables. Theory of Probability and its Applications, 19(2):371–374, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazonov:1990:ARE

[Saz90] V. V. Sazonov. Addendum:

a regular estimate of the accuracy of normal approximation in Hilbert space. Theory of Probability and its Applications, 34(3):554, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 613–613.

Sazonov:1992:AHS

[Saz92] V. V. Sazonov. Addendum: On Hilbert-space-valued Ustatistics. Theory of Probability and its Applications, 36(4):859, ???? 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 822–822.

Sarmanov:1967:PPB

[SB67] O. V. Sarmanov and Z. N. Bratoeva. Probabilistic properties of bilinear expansions of Hermite polynomials. Theory of Probability and its Applications, 12 (3):470–481, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1964:AAN

[SC64a] B. A. Sevast’yanov and V. P. Chistyakov. Addendum: Asymptotic normality in the classical ball problem. Theory of Probability and its Applications, 9 (3):513–514, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1964:ANC

[SC64b] B. A. Sevast’yanov and V. P. Chistyakov. Asymptotic normality in the classical ball problem. Theory of Probability and its Applications, 9(2):198–211, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1971:MRE

[SC71] B. A. Sevast’yanov and V. P.

Chistyakov. Multidimensional renewal equations and moments of branching processes. Theory of Probability and its Applications, 16(2):199–214, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Schaifer:1965:SME

[Sch65] M. P. Schaifer. On the second moments of an estimate for the spectral function. Theory of Probability and its Applications, 10(3):487–489, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Schopf:1979:COT

[Sch79] G. Schopf. On the conjugate orders and types of the rate of decrease of a multi-variate probability law. Theory of Probability and its Applications, 24 (1):204–206, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Schatte:1989:VDS

[Sch89] P. Schatte. On the value distribution of sums of random variables. Theory of Probability and its Applications, 33(4):743–747, ???? 1989. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Schwabe:1993:RED

[Sch93a] Rainer Schwabe. Recursive estimation and difference equations. Theory of Probability and its Applications, 37(2):359–362, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 399–402.

Schweizer:1993:SHI

[Sch93b] M. Schweizer. Semimartingales and hedging in incomplete markets. Theory of Probability and its Applications, 37(1):169–171, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 221–224.

Schladitz:1996:PDF

[SE96] K. Schladitz and H. J. Engelbert. On probability density functions which are their own characteristic functions. Theory of Probability and its Applications, 40 (3):577–581, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 694–698.

Sedunov:1977:CME

[Sed77] E. V. Sedunov. On the choice of a method of estimation in a design procedure for regression experiments which is unbiased in the L_{2} metric. Theory of Probability and its Applications, 21 (2):423–430, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Seidl:1982:QEC

[Sei82] L. Seidl’. Quantitative estimates of the continuity of multi-channel queueing systems. Theory of Probability and its Applications, 26(4):732–744, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Seidl:1987:EMR

[Sei87a] L. Seidl’. An estimate of the moment of the regeneration period in a closed central-server queueing network. Theory of Probability and its Applications, 31 (2):309–313, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Seidl:1987:LDR

[Sei87b] L. Seidl’. On the limit distributions of random symmetric functions. Theory of Probability and its Applications, 31(4):590–603, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Seidl:1989:LDR

[Seĭ89] L. Seĭdl′. On limit distributions of random symmetric polynomials. Theory of Probability and its Applications, 33(2):248–259, ???? 1989. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 266–278.

Seidl:1991:ADL

[Seĭ91] L. Seĭdl′. Analytic description of limit laws for sequences of random symmetric polynomials. Theory of Probability and its Applications, 35(1):174–177, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 153–156.

Selivanov:1969:SEF

[Sel69] B. I. Selivanov. Some explicit formulas in the theory of branching random processes with discrete time and one type of particles. Theory of Probability and its Applications, 14(2):336–342, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Selivanov:1985:LDS

[Sel85] B. I. Selivanov. Limit distributions of the χ^{2} statistic. Theory of Probability and its Applications, 29(1):133–134, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Selivanov:1999:GCA

[Sel99] B. I. Selivanov. On a generalization of the classical allocation problem. Theory of Probability and its Applications, 43 (2):256–268, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97686.

Semovski:1985:BFC

[Sem85] S. V. Semovski. The Bernshtein– Feller Central Limit Theorem in R^{p}. Theory of Probability and its Applications, 29(3):586–591, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Semivskii:1988:DAW

[Sem88a] S. V. Semivskii. Discriminant analysis when the sample is finite and the feature space dimension increases. Theory of Probability and its Applications, 32 (3):521–525, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Semovskii:1988:ONS

[Sem88b] S. V. Semovskii. Operatornormalized sums of random vectors: Convergence to the normal law together with convergence of moments. Theory of Probability and its Applications, 32(4):748–749, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Semakov:1990:PFA

[Sem90] S. L. Semakov. The probability of the first attainment of a level by a component of a multidimensional process on a prescribed interval under restrictions on its other components. Theory of Probability and its Applications, 34(2):357–361, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 402–406.

Senchenko:1968:CNH

[Sen68] D. V. Senchenko. On characteristics of non-homogeneous Markov processes with a finite number of states. Theory of Probability and its Applications, 13(3):517–522, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sen:1974:AST

[Sen74] Pranab Kumar Sen. Asymptotic sequential tests for regular functionals of distribution functions. Theory of Probability and its Applications, 18(2):226–240, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1981:UER

[Sen81] V. V. Senatov. Uniform estimates of the rate of convergence in the multi-dimensional Central Limit Theorem. Theory of Probability and its Applications, 25 (4):745–759, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1982:SNU

[Sen82] V. V. Senatov. Some nonuniform estimates of the speed of convergence in the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 26(4):657–669, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1984:ERC

[Sen84] V. V. Senatov. On estimating the rate of convergence in the Central Limit Theorem over a system of balls in R^{k}. Theory of Probability and its Applications, 28(2):463–467, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1985:ELP

[Sen85] V. V. Senatov. An estimate of the Lévy–Prokhorov metric. Theory of Probability and its Applications, 29(1):109–114, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1986:FEL

[Sen86a] V. V. Senatov. Four examples of lower estimates in the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 30(4):797–805, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1986:DEC

[Sen86b] V. V. Senatov. On the dependency of estimates of the convergence rate in the Central Limit Theorem on the covariance operator of the summands. Theory of Probability and its Applications, 30(2):380–383, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1987:DEC

[Sen87] V. V. Senatov. On the dependence of estimates of the convergence rate in the Central Limit Theorem for balls with center at zero on the covariance operator of the summands. Theory of Probability and its Applications, 31 (1):119–122, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:1989:SRN

[Sen89] V. V. Senatov. Some remarks on nonuniform estimates of the convergence rate. Theory of Probability and its Applications, 33 (3):536–539, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 573–577.

Senatov:1990:ECR

[Sen90] V. V. Senatov. On an estimate of the convergence rate in the Central Limit Theorem in Hilbert space. Theory of Probability and its Applications, 34(2):366, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 412–412.

Senatov:1992:SRE

[Sen92] V. V. Senatov. Some remarks on estimating the rate of convergence in the Central Limit Theorem in the Hilbert space. Theory of Probability and its Applications, 36(2):401–405, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 382–386.

Senatov:1993:ERC

[Sen93] V. V. Senatov. On estimates of the rate of convergence in the Central Limit Theorem in multidimensional spaces. Theory of Probability and its Applications, 37(4):703–706, December 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 759–762.

Seregin:1961:CCS

[Ser61] L. V. Seregin. Continuity conditions for stochastic processes. Theory of Probability and its Applications, 6(1):1–26, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Serdobolskii:1979:CEI

[Ser79] V. I. Serdobol’skii. On classification errors induced by sampling. Theory of Probability and its Applications, 24(1):130–144, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Serfling:1980:SLL

[Ser80] R. J. Serfling. On the strong law of large numbers and related results for quasi-stationary sequences. Theory of Probability and its Applications, 25 (1):187–191, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Serdobolskii:1996:SPS

[Ser96] V. I. Serdobol′skiĭ. Spectral properties of sample covariance matrices. Theory of Probability and its Applications, 40(4): 777–786, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 910–919.

Sevastyanov:1957:ETM

[Sev57a] B. A. Sevast’yanov. An ergodic theorem for Markov processes and its application to telephone systems with refusals. Theory of Probability and its Applications, 2(1):104–112, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1957:LTB

[Sev57b] B. A. Sevast’yanov. Limit theorems for branching stochastic processes of special form. Theory of Probability and its Applications, 2(3):321–331, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1959:TPB

[Sev59] B. A. Sevast’yanov. Transient phenomena in branching stochastic processes. Theory of Probability and its Applications, 4 (2):113–128, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1961:ECB

[Sev61] B. A. Sevast’yanov. Extinction conditions for branching stochastic processes with diffusion. Theory of Probability and its Applications, 6(3):253–263, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1962:ISC

[Sev62] B. A. Sevast’yanov. Influence of storage bin capacity on the average standstill time of a production line. Theory of Probability and its Applications, 7 (4):429–438, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1964:ADB

[Sev64]  B. A. Sevast’yanov.  Age-

dependent branching processes. Theory of Probability and its Applications, 9(4):521–537, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1966:AAD

[Sev66a] B. A. Sevast’yanov. Addendum: Age-dependent branching processes. Theory of Probability and its Applications, 11(2): 321–322, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1966:LTS

[Sev66b] B. A. Sevast’yanov. Limit theorems in a scheme for allocation of particles in cells. Theory of Probability and its Applications, 11(4):614–619, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1967:ABP

[Sev67a] B. A. Sevast’yanov. Asymptotic behavior of the probability of non-extinction of a critical branching process. Theory of Probability and its Applications, 12(1):152–154, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1967:CDN

[Sev67b] B. A. Sevast’yanov. Convergence of the distribution of the number of empty boxes to Gaussian and Poisson processes in the classical ball problem. Theory of Probability and its Applications, 12 (1):126–134, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1968:LTA

[Sev68] B. A. Sevast’yanov. Limit theorems for age-dependent branching processes. Theory of Probability and its Applications, 13 (2):237–259, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1972:RMD

[Sev72] B. A. Sevast’yanov. Random mappings and decompositions of finite sets. Theory of Probability and its Applications, 17 (1):132–145, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1973:PLL

[Sev73]  B. A. Sevast’yanov.  Pois-

son limit law for a scheme of sums of dependent random variables. Theory of Probability and its Applications, 17(4):695–699, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1990:RBP

[Sev90] B. A. Sevast’yanov. Research on branching processes in the probability theory department of Moscow University. Theory of Probability and its Applications, 34(1):200–204, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1994:DHR

[Sev94] B. A. Sevast′yanov. Distribution of the height of a random tree with labeled edges. Theory of Probability and its Applications, 38(2):379–382, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 453–457.

Sevastyanov:1996:EPB

[Sev96] B. A. Sevast′yanov. Extinction probabilities for branching processes bounded from below. Theory of Probability and its Applications, 40(3):495–502, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 556–564.

Sevastyanov:1999:ABE

[Sev99] B. A. Sevastyanov. Asymptotic behavior of the extinction probabilities for stopped branching processes. Theory of Probability and its Applications, 43 (2):315–322, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97693.

Sirazhdinov:1984:ERC

[SF84] S. Kh. Sirazhdinov and Sh. K. Formanov. On estimates of the rate of convergence in the Central Limit Theorem for homogeneous Markov chains. Theory of Probability and its Applications, 28(2):229–239, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savin:1962:STB

[SG62] A. A. Savin and V. P. Ghistyakov. Some theorems for branching processes with several types of particles. Theory of Probability and its Applications, 7(1): 93–100, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sgibnev:1992:ABH

[Sgi92] M. S. Sgibnev. Asymptotic behavior of higher renewal epochs. Theory of Probability and its Applications, 36(3): 508–518, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 494–504.

Sgibnev:1998:AGR

[Sgi98] M. S. Sgibnev. Asymptotics of the generalized renewal functions when the variance is finite. Theory of Probability and its Applications, 42(3):536–541, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 632–637.

Sgibnev:1999:DSR

[Sgi99] M. S. Sgibnev. On the distribution of the supremum of a random walk when the characteristic equation has roots. Theory of Probability and its Applications, 43(2):322–329, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97694.

Shalaevskii:1961:SPC

[Sha61] O. V. Shalaevskii. A short proof of the Cramér–Rao inequality. Theory of Probability and its Applications, 6(3):323–324, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shagdar:1970:UEM

[Sha70] D. Shagdar. An unbiased estimate with minimal variance in the class of linear estimates. Theory of Probability and its Applications, 15(1):132–133, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shalaumov:1980:BDP

[Sha80a] V. A. Shalaumov. On the behavior of a diffusion process with a large drift coefficient in a halfspace. Theory of Probability and its Applications, 24(3):592–598, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shanbhag:1980:DBM

[Sha80b] D. N. Shanbhag. Diagonality of the Bhattacharyya matrix as a characterization. Theory of Probability and its Applications, 24 (2):430–433, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shashiashvili:1993:SIS

[Sha93] M. A. Shashiashvili. On a semimartingale inequality for Snell envelopes. Theory of Probability and its Applications, 37(3): 568–569, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 37(3), (1992), pp. 592–593.

Shao:1996:SDT

[Sha96] Q. M. Shao. A small deviation theorem for independent random variables. Theory of Probability and its Applications, 40(1):191–200, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 225–235.

Shervashidze:1971:UER

[She71] T. L. Shervashidze. On a uniform estimate of the rate of convergence in the multidimensional local limit theorem for densities. Theory of Probability and its Applications, 16(4):741–743, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shergin:1980:CRC

[She80] V. V. Shergin. On the convergence rate in the Central Limit Theorem for m-dependent random variables. Theory of Probability and its Applications, 24 (4):782–796, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shemyakin:1986:PEC

[She86] A. E. Shemyakin. A problem of estimating the coordinates of a radiation source. Theory of Probability and its Applications, 30 (1):128–140, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sherstnev:1987:TPF

[She87] V. I. Sherstnev. Transition phenomena for final probabilities of branching processes. Theory of Probability and its Applications, 31(1):101–107, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shemyakin:1988:RCT

[She88a] A. E. Shemyakin. Rao-Cramér type integral inequalities for estimates of a vector parameter. Theory of Probability and its Applications, 32(3):426–434, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shen:1988:PSF

[She88b] A. Kh. Shen’. Probabilistic strategies in finite games with complete information. Theory of Probability and its Applications, 32(3):516–518, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shemyakin:1993:IIP

[She93a] A. E. Shemyakin. On information inequalities in the parametric estimation. Theory of Probability and its Applications, 37 (1):89–91, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 121–123.

Shen:1993:ACR

[She93b] A. Shen′. Algorithmic complexity and randomness: Recent developments. Theory of Probability and its Applications, 37 (1):92–97, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 124–131.

Sherstnev:1997:JDC

[She97] V. I. Sherstnev. Joint distribution of carries in integer addition. Theory of Probability and its Applications, 41(2):394–399, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 467–473.

Sherstnev:1999:RVU

[She99] V. I. Sherstnev. A random variable uniformly distributed ona finite Abelian group as sum of independent summands. Theory of Probability and its Applications, 43(2):329–335, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97695.

Shiryaev:1960:SPS

[Shi60] A. N. Shiryaev. Some problems in the spectral theory of higherorder moments. I. Theory of Probability and its Applications, 5(3):265–284, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1963:CES

[Shi63a] A. N. Shiryaev. On conditions for ergodicity of stationary processes in terms of higher order moments. Theory of Probability and its Applications, 8 (4):436–439, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1963:OMQ

[Shi63b] A. N. Shiryaev. On optimum methods in quickest detection problems. Theory of Probability and its Applications, 8(1):22–46, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shirokov:1964:DCP

[Shi64a] F. V. Shirokov. On I. D. Cherkasov’s paper “On the Transformation of a Diffusion Process to a Wiener Process”. Theory of Probability and its Applications, 9(1):156–157, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1964:MSS

[Shi64b] A. N. Shiryaev. On Markov sufficient statistics in non-additive Bayes problems of sequential analysis. Theory of Probability and its Applications, 9(4): 604–618, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1965:SEF

[Shi65] A. N. Shiryaev. Some exact formulas in a “Disorder” problem. Theory of Probability and its Applications, 10(2):348–354, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1967:ASE

[Shi67] A. N. Shiryaev. Addendum: On stochastic equations in the theory of conditional Markov processes. Theory of Probability and its Applications, 12(2):342, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiganov:1980:SEA

[Shi80] I. S. Shiganov. Some estimates of the accuracy of approximation in the multi-dimensional Central Limit Theorem. Theory of Probability and its Applications, 24 (3):626–629, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiganov:1982:ASE

[Shi82] I. Shiganov. Addendum: Some estimates of the accuracy of approximation in the multidimensional Central Limit Theorem. Theory of Probability and its Applications, 26(2):434, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shipilov:1986:AGC

[Shi86] O. I. Shipilov. An analogue of the Glivenko–Cantelli theorem for binomial sampling from a finite population. Theory of Probability and its Applications, 30 (3):610–613, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1990:ANK

[Shi90] A. N. Shiryaev. Andrei Nikolaevich Kolmogorov: In memoriam. Theory of Probability and its Applications, 34(1):1–99, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1993:CBB

[Shi93] A. N. Shiryaev. In celebration of the 80th birthday of Boris Vladimirovitch Gnedenko (an interview). Theory of Probability and its Applications, 37(4): 674–691, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 724–746.

Shiryaev:1995:SBC

[Shi95] A. N. Shiryaev. On some basic concepts and some basic stochastic models used in finance. Theory of Probability and its Applications, 39(1):1–13, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 5–22.

Shipilov:1997:EDF

[Shi97] O. I. Shipilov. The empirical distribution function in an experiment with binomial randomization. Theory of Probability and its Applications, 41(2):307–316, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97551. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 310–322.

Shklover:1974:SCR

[Shk74] M. D. Shklover. The simulation of a class of random fields. Theory of Probability and its Applications, 18(2):403–404, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shlosman:1981:LTP

[Shl81] S. B. Shlosman. Limit theorems of probability theory for compact topological groups. Theory of Probability and its Applications, 25(3):604–609, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shlosman:1989:RAC

[Shl89] S. E. Shlosman. Relations among the cumulants of random fields with attraction. Theory of Probability and its Applications, 33 (4):645–655, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 33(4), (1988), pp. 694–705.

Shmulyan:1964:OFN

[Shm64] Yu. L. Shmul’yan. Optimum factorization of non-negative matrix functions. Theory of Probability and its Applications, 9 (2):346–350, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shneiberg:1987:HSR

[Shn87] I. Ya. Shneiberg. Hierarchical sequences of random variables. Theory of Probability and its Applications, 31(1):137–141, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shorgin:1979:NCE

[Sho79] S. Ya. Shorgin. A non-classical estimate of the speed of convergence in the multidimensional Central Limit Theorem taking into account large deviations. Theory of Probability and its Applications, 23(3):667–671, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shorgin:1983:NER

[Sho83] S. Ya. Shorgin. Nonclassical estimates of the rate of convergence in the Central Limit Theorem taking into account large deviations. Theory of Probability and its Applications, 27(2):324–337, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 27, (1982), pp. 308– 318.

Sholomitskii:1996:CCS

[Sho96] A. G. Sholomitskii. On conditions for convergence of sums of independent variables. Theory of Probability and its Applications, 40(4):786–791, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sholomitskii:1997:TWF

[Sho97] A. G. Sholomitskiĭ. A theorem of W. Feller for sums of dependent variables. Theory of Probability and its Applications, 41 (2):399–403, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 473–477.

Sholomitskii:1999:NCN

[Sho99] A. G. Sholomitskii. On the necessary conditions of normal convergence for martingales. Theory of Probability and its Applications, 43(3):434–448, September 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97702.

Shparo:1966:PPT

[Shp66] D. I. Shparo. Principles of potential theory and Markov chains. Theory of Probability and its Applications, 11(3):415–424, ???? 1966. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Shparo:1968:SMC

[Shp68] D. I. Shparo. On some methods of completion of the phase space of a Markov chain. Theory of Probability and its Applications, 13(2):327–329, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shtatland:1965:LPP

[Sht65a] E. S. Shtatland. On local properties of processes with independent increments. Theory of Probability and its Applications, 10 (2):317–322, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shtatland:1965:DMP

[Sht65b] E. S. Shtatland. On the distribution of the maximum of a process with independent increments. Theory of Probability and its Applications, 10(3):483–487, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shtatland:1966:DTM

[Sht66] E. S. Shtatland. The distribution of the time the maximum is achieved for processes with independent increments. Theory of Probability and its Applications, 11(4):637–642, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1958:EPI

[Shu58] M. G. Shur. Ergodic properties of invariant Markov chains on homogeneous spaces. Theory of Probability and its Applications, 3(2):127–140, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1962:LCE

[Shu62] M. G. Shur. Localization of the concept of an excessive function connected with a Markov process. Theory of Probability and its Applications, 7(2): 185–189, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1964:FWS

[Shu64] M. G. Shur. On functions which are superharmonic for Markov processes. Theory of Probability and its Applications, 9(1): 114–121, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1965:MGS

[Shu65] M. G. Shur. On the maximum of a Gaussian stationary process. Theory of Probability and its Applications, 10(2):354–357, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1966:AMP

[Shu66a] M. Shur. Addendum: Markov processes with majorized exit probabilities. Theory of Probability and its Applications, 11 (4):642–643, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1966:MPM

[Shu66b] M. G. Shur. Markov processes with majorized hitting probabilities. Theory of Probability and its Applications, 11(2):229–248, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1967:ETC

[Shu67] M. G. Shur. Ergodic theorems for a class of Markov processes. Theory of Probability and its Applications, 12(3):443–454, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1968:MBC

[Shu68] M. G. Shur. On the Martin boundary for a class of Markov processes. Theory of Probability and its Applications, 13(1): 170–174, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1969:BCFa

[Shu69a] M. G. Shur. Behavior of coexcessive functions near the Martin boundary. I. Theory of Probability and its Applications, 14 (2):263–278, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1969:BCFb

[Shu69b] M. G. Shur. Behavior of coexcessive functions near the Martin boundary. II. Theory of Probability and its Applications, 14 (3):424–431, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1970:RPM

[Shu70a] M. G. Shur. Regular points of the Martin boundary. Theory of Probability and its Applications, 15(4):619–629, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1970:SRA

[Shu70b] M. G. Shur. Some remarks on adjoint Markov processes. Theory of Probability and its Applications, 15(1):109–115, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1973:MCN

[Shu73] M. G. Shur. A Martin compact with a non-negligible irregular boundary point. Theory of Probability and its Applications, 17(2):351–355, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shukri:1976:LLT

[Shu76a] E. M. Shukri. Local limit theorems for sums of weighted independent random variables. Theory of Probability and its Applications, 21(1):137–144, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1976:EPT

[Shu76b] M. G. Shur. Elements of potential theory for nonhomogeneous Markov processes. Theory of Probability and its Applications, 20(2):261–284, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1976:NMD

[Shu76c] V. M. Shurenkov. A note on a multi-dimensional renewal equation. Theory of Probability and its Applications, 20(4):833–836, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1977:ETM

[Shu77a] M. G. Shur. An ergodic theorem for Markov processes. I. Theory of Probability and its Applications, 21(2):400–406, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1977:KEC

[Shu77b] V. M. Shurenkov. Kolmogorov’s equations for a class of transition functions. Theory of Probability and its Applications, 21 (4):854–858, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1977:TRP

[Shu77c] V. M. Shurenkov. Transformations of random processes preserving the Markov property. Theory of Probability and its Applications, 22(1):119–123, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1977:TLT

[Shu77d] V. M. Shurenkov. Two limit theorems for critical branching processes. Theory of Probability and its Applications, 21(3):534–544, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1978:ETM

[Shu78a] M. G. Shur. An ergodic theorem for Markov processes. II. Theory of Probability and its Applications, 22(4):692–707, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1978:DMP

[Shu78b] M. G. Shur. On dual Markov processes. Theory of Probability and its Applications, 22(2):257–270, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1979:BCE

[Shu79a] V. M. Shurenkov. Boundary conditions and an ergodic theorem for processes with independent increments. Theory of Probability and its Applications, 24 (1):53–61, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1979:LDE

[Shu79b] V. M. Shurenkov. Limit distribution of the exit time from an expanding interval and of the position at exit time of a process with independent increments and onesigned jumps. Theory of Probability and its Applications, 23 (2):402–407, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1980:CCM

[Shu80a] M. G. Shur. On continuity criteria for Markov processes. Theory of Probability and its Applications, 25(1):141–147, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1980:AFB

[Shu80b] V. M. Shurenkov. On additive functionals of branching processes. Theory of Probability and its Applications, 24(2):396–401, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1981:RWH

[Shu81] V. M. Shurenkov. Random walks on the half-axis. I. Boundary problems and an ergodic theorem. Theory of Probability and its Applications, 26(1): 43–55, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1982:IMM

[Shu82a] M. G. Shur. Invariant measures for Markov chains and Feller extensions of chains. Theory of Probability and its Applications, 26(3):485–497, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1982:RWH

[Shu82b] V. M. Shurenkov. Random walks on the half-axis. II. Limit distributions of boundary functionals. Theory of Probability and its Applications, 26(3):452–467, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1985:APP

[Shu85a] M. G. Shur. Asymptotic properties of positive operator powers. I. Theory of Probability and its Applications, 29(4):719–730, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1985:PTM

[Shu85b] M. G. Shur. The Poisson theorem and Markov chains. Theory of Probability and its Applications, 29(1):124–126, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1985:TMR

[Shu85c] V. M. Shurenkov. On the theory of Markov renewal. Theory of Probability and its Applications, 29(2):247–265, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1986:APP

[Shu86] M. G. Shur. Asymptotic properties of powers of positive operators. II. Theory of Probability and its Applications, 30(2):260–271, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shurenkov:1987:ATM

[Shu87] V. M. Shurenkov. Addendum: On the theory of Markov renewal. Theory of Probability and its Applications, 31(1):175, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1989:RLT

[Shu89] M. G. Shur. Ratio limits theorems for random walks in homogeneous spaces. I. Theory of Probability and its Applications, 33(4):656–667, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1990:RLT

[Shu90] M. G. Shur. Ratio limit theorems for random walks in homogeneous spaces II. Theory of Probability and its Applications, 34(3):464–473, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 516–527.

Shur:1991:ACS

[Shu91a] M. G. Shur. Absolutely continuously and singularly generated harmonic functions for random walks on groups. Theory of Probability and its Applications, 35 (4):805–810, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 787–793.

Shur:1991:ALT

[Shu91b] M. G. Shur. Addendum: Limit theorems for ratios which are generated by random walks in homogeneous spaces. Theory of Probability and its Applications, 35(3):618, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:1991:ELT

[Shu91c] M. G. Shur. Erratum: “On limit theorems for ratios generated by random walks in homogeneous spaces. II” [Teor. Veroyatnost. i Primenen. 34 (1989), no. 3, 516–527; MR 91d:60165]. Theory of Probability and its Applications, 35(3):618, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 616–616.

Shuteev:1991:SAD

[Shu91d] G. E. Shuteev. Saddlepoint approximation of a distribution of a statistic. Theory of Probability and its Applications, 35(3): 607–612, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 602–607.

Shur:1992:CFF

[Shu92] M. G. Shur. On the compactness of a family of functions that are harmonic for a random walk on a group. Theory of Probability and its Applications, 36(1):193–197, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 194–198.

Shur:1996:AES

[Shu96] M. G. Shur. Asymptotic equidistribution of symmetric random walks on unimodular groups. Theory of Probability and its Applications, 40(2):329–339, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Studnev:1967:RCL

[SI67] Yu. P. Studnev and Yu. I. Ignat. Refinement of the Central Limit Theorem and of its global version. Theory of Probability and its Applications, 12(3):508–512, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sidak:1957:RBS

[Sid57] Zbynek Sidák. On relations between strict-sense and wide-sense conditional expectations. Theory of Probability and its Applications, 2(2):267–272, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sidorenko:1991:PIS

[Sid91] A. F. Sidorenko. A precise inequality for the sum of two random vectors with symmetric distribution. Theory of Probability and its Applications, 35(3): 613–617, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 595–599.

Siegel:1980:USD

[Sie80] G. Siegel. Uniqueness of symmetric distribution functions defined on bounded sets and its applications. Theory of Probability and its Applications, 24(4):830–833, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siegel:1987:VCC

[Sie87] G. Siegel. Vague convergence for convolutions and for infinitely divisible functions. Theory of Probability and its Applications, 31 (1):152–159, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siegel:1990:OSD

[Sie90] G. Siegel. Operator-stable distributions in separable Banach spaces. Theory of Probability and its Applications, 34(3):497–505, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 552–560.

Silvestrov:1973:RLC

[Sil73] D. S. Sil’vestrov. Remarks on the limit of a composite random function. Theory of Probability and its Applications, 17 (4):669–677, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Silvestrov:1974:CCS

[Sil74] D. S. Sil’vestrov. Conditions for convergence of the superposition of random processes in the J-topology. Theory of Probability and its Applications, 18 (3):579–582, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Silvestrov:1979:RWL

[Sil79] D. S. Sil’vestrov. Remarks on the weak limit of the superposition of asymptotically independent random functions. Theory of Probability and its Applications, 24(1):151–156, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

SilVestrov:1992:IPC

[Sīl92] D. S. Sīl′vestrov. Invariance principle for cumulative processes with semi-Markov switch-overs in triangular array. Theory of Probability and its Applications, 36(3):519–535, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 505–520.

Sinai:1957:DFP

[Sin57] Ya. G. Sinai. On the distribution of the first positive sum for a sequence of independent random variables. Theory of Probability and its Applications, 2 (1):122–129, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinai:1960:DSS

[Sin60] Ya. G. Sinai. Dynamic systems and stationary Markov processes. Theory of Probability and its Applications, 5(3):305–308, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinai:1962:LTS

[Sin62] Ya. G. Sinai. On limit theorems for stationary processes. Theory of Probability and its Applications, 7(2):205–210, ???? 1962. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinaj:1963:HOS

[Sin63] Ya. G. Sinaj. On higher order spectral measures of ergodic stationary processes. Theory of Probability and its Applications, 8(4):429–436, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinai:1976:SSP

[Sin76] Yu. G. Sinai. Self-similar probability distributions. Theory of Probability and its Applications, 21(1):64–80, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinai:1983:LBO

[Sin83] Ya. G. Sinai. The limiting behavior of a one-dimensional random walk in a random medium. Theory of Probability and its Applications, 27(2):256–268, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinai:1994:RWR

[Sin94] Ya. G. Sinaĭ. A random walk with random potential. Theory of Probability and its Applications, 38(2):382–385, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 457–460.

Sirazhadinov:1956:EMB

[Sir56] S. Kh. Sirazhadinov. On estimators with minimum bias for a binomial distribution. Theory of Probability and its Applications, 1(1):150–156, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sirazhdinov:1959:EEL

[Sir59] S. Kh. Sirazhdinov. On an exact estimate for a local theorem. Theory of Probability and its Applications, 4(2):215–218, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sirjaev:1963:DDMa

[Sir63a] A. N. Sirjaev. On the detection of disorder in a manufacturing process. I. Theory of Probability and its Applications, 8 (3):247–265, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sirjaev:1963:DDMb

[Sir63b] A. N. Sirjaev. On the detection of disorder in a manufacturing process. II. Theory of Probability and its Applications, 8(4):402–413, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siraya:1973:CRS

[Sir73] T. N. Siraya. On canonical representations for stochastic processes of multiplicities one and two. Theory of Probability and its Applications, 18(1):153–158, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siraya:1977:SP

[Sir77a] T. N. Siraya. On subordinate processes. Theory of Probability and its Applications, 22(1): 129–133, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siraya:1977:MSO

[Sir77b] T. N. Siraya. On the multiplicity of a sum of orthogonal random processes. Theory of Probability and its Applications, 21 (4):858–863, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siraya:1978:CRS

[Sir78] T. N. Siraya. Canonical representations of second-order random processes. Theory of Probability and its Applications, 22 (2):418–424, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sharshanova-Khodareva:1992:NID

[SK92] G. A. Sharshanova-Khodareva. Notes on infinite divisibility of meromorphic characteristic functions with one pole. Theory of Probability and its Applications, 36(1):190–193, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 192–194.

Solovev:1993:REC

[SK93] A. D. Solov′ev and D. G. Kostantinidis. Reliability estimation of a complex renewable system with an unbounded number of repair units. Theory of Probability and its Applications, 37 (1):98–100, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 91–94.

Skitovicn:1956:CBM

[Ski56] V. P. Skitovicn. On a characterization of Brownian motion. Theory of Probability and its Applications, 1(3):326–328, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:1995:TTPa

[SKKM95a] A. N. Shiryaev, Yu. M. Kabanov, D. O. Kramkov, and A. V. Mel′nikov. Toward the theory of pricing of options of both European and American types. I. Discrete time. Theory of Probability and its Applications, 39 (1):14–60, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 23–79.

Shiryaev:1995:TTPb

[SKKM95b] A. N. Shiryaev, Yu. M. Kabanov, D. O. Kramkov, and A. V. Mel′nikov. Toward the theory of pricing of options of both European and American types. II. Continuous time. Theory of Probability and its Applications, 39(1):61–102, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 80–129.

Skorokhod:1956:LTS

[Sko56] A. V. Skorokhod. Limit theorems for stochastic processes. Theory of Probability and its Applications, 1(3):261–290, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1957:LTS

[Sko57a] A. V. Skorokhod. Limit theorems for stochastic processes with independent increments. Theory of Probability and its Applications, 2(2):138–171, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1957:DMW

[Sko57b] A. V. Skorokhod. On the differentiability of measures which correspond to stochastic processes. I. Processes with independent increments. Theory of Probability and its Applications, 2(4): 407–432, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1958:LTM

[Sko58] A. V. Skorokhod. Limit theorems for Markov processes. Theory of Probability and its Applications, 3(3):202–246, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1960:DMC

[Sko60] A. V. Skorokhod. On the differentiability of measures corresponding to random processes. II. Markov processes. Theory of Probability and its Applications, 5(1):40–49, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1961:AFP

[Sko61a] A. V. Skorokhod. Additive functionals of a process of Brownian motion. Theory of Probability and its Applications, 6 (4):396–404, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1961:SED

[Sko61b] A. V. Skorokhod. Stochastic equations for diffusion processes in a bounded region. Theory of Probability and its Applications, 6(3):264–274, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1962:SED

[Sko62] A. V. Skorokhod. Stochastic equations for diffusion processes in a bounded region. II. Theory of Probability and its Applications, 7(1):3–23, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1963:LTH

[Sko63a] A. V. Skorokhod. A limit theorem for homogeneous Markov chains. Theory of Probability and its Applications, 8(1):61–70, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1963:HCM

[Sko63b] A. V. Skorokhod. On homogeneous continuous Markov processes that are martingales. Theory of Probability and its Applications, 8(4):355–365, ???? 1963.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1964:BCC

[Sko64a] A. V. Skorokhod. Boundary conditions for certain Markov processes. Theory of Probability and its Applications, 9(4):582–590, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1964:BDP

[Sko64b] A. V. Skorokhod. Branching diffusion processes. Theory of Probability and its Applications, 9(3):445–449, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1965:ACI

[Sko65] A. V. Skorokhod. Absolute continuity of infinitely divisible distributions under translations. Theory of Probability and its Applications, 10(3):465–472, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1966:LSC

[Sko66] A. V. Skorokhod. On the local structure of continuous Markov processes. Theory of Probability and its Applications, 11(3): 336–372, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1967:HMP

[Sko67] A. V. Skorokhod. Homogeneous Markov processes without discontinuities of the second kind. Theory of Probability and its Applications, 12(2):222–240, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1969:IFF

[Sko69] A. V. Skorokhod. On an inversion formula for Fourier transformation in Hilbert space. Theory of Probability and its Applications, 14(1):124–128, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1970:NGM

[Sko70a] A. V. Skorokhod. A note on Gaussian measures in a Banach space. Theory of Probability and its Applications, 15(3):508, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1970:ATM

[Sko70b] A. V. Skorokhod. On admissible translations of measures in Hilbert space. Theory of Probability and its Applications, 15 (4):557–580, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1974:RCE

[Sko74a] A. V. Skorokhod. A remark on consistent estimates of a Gaussian distribution. Theory of Probability and its Applications, 19(1):190–193, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1974:TCR

[Sko74b] A. V. Skorokhod. A theorem on continuity of a random function on a compact set in Hilbert space. Theory of Probability and its Applications, 18(4):771–774, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1975:RSI

[Sko75] A. V. Skorokhod. A remark on square integrable martingales. Theory of Probability and its Applications, 20(1):195–198, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1976:GSI

[Sko76] A. V. Skorokhod. On a generalization of a stochastic integral. Theory of Probability and its Applications, 20(2):219–233, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1977:RRV

[Sko77] A. V. Skorokhod. On a representation of random variables. Theory of Probability and its Applications, 21(3):628–632, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1979:LAL

[Sko79] A. V. Skorokhod. Linear and almost linear functionals on a measurable Hilbert space. Theory of Probability and its Applications, 23(2):380–385, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1980:MSP

[Sko80a] A. V. Skorokhod. On measurability of stochastic processes. Theory of Probability and its Applications, 25(1):139–141, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1980:CIP

[Sko80b] T. A. Skorokhod. On convergence of infinite products of independent random linear operators in a Hilbert space. Theory of Probability and its Applications, 24(4):809–813, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1981:SDE

[Sko81] A. V. Skorokhod. Stochastic differential equations depending on a parameter. Theory of Probability and its Applications, 25 (4):659–666, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1987:TRM

[Sko87] A. V. Skorokhod. Topologically recurrent Markov chains: Ergodic properties. Theory of Probability and its Applications, 31 (4):563–571, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1992:AEP

[Sko92] A. V. Skorokhod. σ-algebras of events on probability spaces. similarity and factorization. Theory of Probability and its Applications, 36(1):63–73, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 127–137.

Smorodina:1990:DNS

[SL90] N. V. Smorodina and M. A. Lifshits. The distribution of the norm of a stable vector. Theory of Probability and its Applications, 34(2):266–274, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 304–313.

Sedunov:1993:RCR

[SL93] E. V. Sedunov and A. V. Lipin. Randomization of conditions in a regression experiment for functional measurement models. Theory of Probability and its Applications, 37(4):695–702, December 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 750–759.

Slastnikov:1978:DCR

[Sla78] A. D. Slastnikov. On the distribution of certain random determinants. Theory of Probability and its Applications, 22 (4):868–876, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slastnikov:1979:LTM

[Sla79] A. D. Slastnikov. Limit theorems for moderate deviation probabilities. Theory of Probability and its Applications, 23(2):322–340, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slastnikov:1982:LDS

[Sla82] A. D. Slastnikov. On large deviations for sums of nonidentically distributed random variables. Theory of Probability and its Applications, 27 (1):37–48, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slastnikov:1985:NZN

[Sla85] A. D. Slastnikov. Narrow zones of normal convergence for sums of non-identically distributed random variables. Theory of Probability and its Applications, 29 (3):570–574, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slavova:1987:ECR

[Sla87] V. V. Slavova. An estimate of the convergence rate of partial moments. Theory of Probability and its Applications, 31(2):327–332, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slivnyak:1962:SPS

[Sli62] I. M. Slivnyak. Some properties of stationary flows of homogeneous random events. Theory of Probability and its Applications, 7(3):336–341, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slivnyak:1964:ASP

[Sli64] I. Slivnyak. Addendum: Some properties of stationary flows of homogeneous random events. Theory of Probability and its Applications, 9(1):168, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sirazhdinov:1962:CMD

[SM62] S. Kh. Sirazhdinov and M. Mamatov. On convergence in the mean for densities. Theory of Probability and its Applications, 7(4):424–428, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sazonov:1979:CLT

[SM79] V. V. Sazonov and V. G. Mirantsev. On the Central Limit Theorem in R^{K}. Theory of Probability and its Applications, 23 (2):407–411, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Smiriga:1969:IRC

[Smi69] Nora Smiriga. An integral representation of a continuous linear stochastic process with independent pieces. Theory of Probability and its Applications, 14 (1):24–34, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sirazhdinov:1990:PLD

[SMI90a] S. Kh. Sirazhdinov, Sh. A. Mirakhmedov, and Sh. A. Ismatullaev. Probabilities of large deviations for randomized divisible statistics in the multinomial scheme. Theory of Probability and its Applications, 34 (4):645–657, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 706–718.

Smirnov:1990:CET

[Smi90b] S. N. Smirnov. Converse of an ergodic theorem of Sevast’yanov. Theory of Probability and its Applications, 34 (2):346–348, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 390–392.

Smorodina:1985:LIP

[Smo85] N. V. Smorodina. A local invariance principle for processes linearly generated by independent random variables. Theory of Probability and its Applications, 29(4):707–718, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Smorodina:1989:DCC

[Smo89] N. V. Smorodina. Differential calculus in configuration space and stable measures. Theory of Probability and its Applications, 33(3):488–499, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 522–534.

Smorodina:1991:GOF

[Smo91] N. V. Smorodina. The Gaussostrogradsky formula for the space of configurations. Theory of Probability and its Applications, 35(4):725–736, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 727–739.

Smorodina:1997:AED

[Smo97] N. V. Smorodina. An asymptotic expansion of the distribution of a homogeneous functional of a strictly stable vector. Theory of Probability and its Applications, 41(1):91–115, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97505. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 133–163.

Stoyan:1985:CCM

[SO85] D. Stoyan and J. Ohser. Crosscorrelation measures of weighted random measures and their estimation. Theory of Probability and its Applications, 29(2): 345–355, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sobol:1958:PRN

[Sob58] I. M. Sobol’. Pseudo-random numbers for the machine “Strela”. Theory of Probability and its Applications, 3(2):192–197, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sobol:1960:SPI

[Sob60] I. M. Sobol. On the solution of Peierl’s integral equation by the Monte Carlo method. Theory of Probability and its Applications, 5(3):326–331, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sobol:1964:PPR

[Sob64] I. M. Sobol’. On periods of pseudo-random sequences. Theory of Probability and its Applications, 9(2):333–338, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sobolev:1970:GB

[Sob70] A. I. Sobolev. On a game of Borel. Theory of Probability and its Applications, 15(3):540–544, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sobko:1973:DAN

[Sob73a] G. M. Sobko. A diffusion approximation of non-Markov random walks on differentiable manifolds. Theory of Probability and its Applications, 18(1):41– 53, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sobko:1973:FDP

[Sob73b] G. M. Sobko. The first diffusion problem on differentiable manifolds. Theory of Probability and its Applications, 17(3):521–528, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sokolov:1963:CPRa

[Sok63a] S. Sokolov. Continuous planning of regression experiments. Theory of Probability and its Applications, 8(1):89–96, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sokolov:1963:CPRb

[Sok63b] S. Sokolov. Continuous planning of regression Experiments.II. Theory of Probability and its Applications, 8(3):298–304, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sokolova:1984:EGM

[Sok84] S. D. Sokolova. On the equivalence of Gaussian measures corresponding to solutions of stochastic differential equations. Theory of Probability and its Applications, 28(2):451–454, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Solovev:1966:CIA

[Sol66] A. D. Solov’ev. A combinatorial identity and its application to the problem concerning the first occurrence of a rare event. Theory of Probability and its Applications, 11(2):276–282, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Solovev:1967:AMO

[Sol67] A. D. Solov’ev. Addendum: a method for obtaining estimates of the expectation of a function of the components of an ndimensional vector. Theory of Probability and its Applications, 12(3):529, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Solntsev:1994:SCM

[Sol94] S. A. Solntsev. On sample continuity of multidimensional Gaussian Markov processes. Theory of Probability and its Applications, 38(4):694–705, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 827–841.

Soltani:1995:SRM

[Sol95] A. R. Soltani. On spectral representation of multivariate stable processes. Theory of Probability and its Applications, 39 (3):465–475, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 605–617.

Sonin:1967:CDD

[Son67] I. M. Sonin. On a class of degenerate diffusion processes. Theory of Probability and its Applications, 12(3):490–496, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sonin:1992:ANM

[Son92] I. M. Sonin. An arbitrary nonhomogeneous Markov chain with bounded number of states May be decomposed into asymptotically noncommunicating components having the mixing property. Theory of Probability and its Applications, 36(1):74– 85, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 65–77.

Sorgin:1978:AGB

[Sor78] S. Ya. Sorgin. Approximation of a generalized binomial distribution. Theory of Probability and its Applications, 22(4):846–850, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sozanov:1964:AQR

[Soz64] V. V. Sozanov. Answer to a question of R. L. Dobrushin. Theory of Probability and its Applications, 9(1):160–161, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Smirnov:1957:NIF

[SP57] S. V. Smirnov and M. K.

Potapov. A nomogram for the incomplete Γ-function and the χ^{2} probability function. Theory of Probability and its Applications, 2(4):461–465, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Smirnov:1961:NPF

[SP61] S. V. Smirnov and M. K.

Potapov. A nomogram for the χ^{2} probability function. Theory of Probability and its Applications, 6(1):124–126, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Seiffert:1979:UTF

[SR79] I. Seiffert and M. Riedel. Uniqueness of two factorization problems. Theory of Probability and its Applications, 24(1):207–211, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sreehari:1970:CLD

[Sre70] M. Sreehari. On a class of limit distributions for normalized sums of independent random variables. Theory of Probability and its Applications, 15 (2):258–281, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Schweizer:1962:SMS

[SS62] B. Schweizer and A. Sklar. Statistical metric spaces arising from sets of random variables in Euclidean n-space. Theory of Probability and its Applications, 7 (4):447–456, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shi-Sun:1965:ABO

[SS65a] Mao Shi-Sun. Asymptotic behavior of the optimal probability of error for the transmission of information in a channel without memory which is symmetrical on pairs of input symbols for small rates of transmission. Theory of Probability and its Applications, 10(1):154–161, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1965:LTR

[SS65b] A. V. Skorokhod and N. P. Slobodenyuk. Limit theorems for random walks, I. Theory of Probability and its Applications, 10 (4):596–606, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1966:LTR

[SS66] A. V. Skorokhod and N. P. Slobodenyuk. Limit theorems for random walks. II. Theory of Probability and its Applications, 11(1):46–57, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sarafyan:1988:FSD

[SS88a] V. V. Sarafyan and A. V. Skorokhod. On fast-switching dynamical systems. Theory of Probability and its Applications, 32 (4):595–607, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sedunov:1988:UDE

[SS88b] E. V. Sedunov and N. G.

Sidorenko. Unbiased design of experiments in a Hilbert space. Theory of Probability and its Applications, 32(4):734–738, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Siani:1990:FAC

[SS90] Ya. G. Siani and A. N. Shiryaev.

The fiftieth anniversary of the creation of the probability theory department of the mechanics and mathematics faculty of Moscow University, founded by A. N. Kolmogorov. Theory of Probability and its Applications, 34(1):164–165, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Skorokhod:1992:SEP

[SS92] A. V. Skorokhod and V. I.

Stepakhno. On some empirical properties of a multidimensional normal law. Theory of Probability and its Applications, 36(2):405–411, May 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Smolyanov:1993:SPT

[SS93] O. G. Smolyanov and E. T. Shavgulidze. A simple proof of Tarieladze’s theorem on sufficiency of positively sufficient topologies. Theory of Probability and its Applications, 37 (2):402–404, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 421–424.

Shepp:1995:NLP

[SS95] L. A. Shepp and A. N. Shiryaev. A new look at pricing of the “Russian Option”. Theory of Probability and its Applications, 39(1):103–119, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(1), (1994), pp. 130–149.

Sazonov:1966:PDT

[ST66] V. V. Sazonov and V. N. Tutubalin. Probability distributions on topological groups. Theory of Probability and its Applications, 11(1):1–45, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Selezneva:1973:ESC

[ST73] T. V. Selezneva and V. N. Tutubalin. The evaluation of statistical characteristics of a long waveguide in a two-mode model. Theory of Probability and its Applications, 18(1):92–106, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sadovyak:1974:ACF

[ST74] A. M. Sadovyak and E. F.

Tsar’kov. An analogue of the Cauchy formula for stochastic differential equations. Theory of Probability and its Applications, 18(2):394–396, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Staroverov:1963:SP

[Sta63] O. V. Staroverov. On a searching problem. Theory of Probability and its Applications, 8 (2):184–187, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Statulyavichus:1965:LTD

[Sta65] V. A. Statulyavichus. Limit theorems for densities and asymptotic expansions for distributions of sums of independent random variables. Theory of Probability and its Applications, 10(4): 582–595, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Statulyavichus:1984:CAM

[Sta84] V. A. Statulyavichus. On a condition of almost Markov regularity. Theory of Probability and its Applications, 28(2):379–383, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Startsev:1997:DSE

[Sta97] A. N. Startsev. On the distribution of the size of an epidemicin a non-Markovian model. Theory of Probability and its Applications, 41(4):730–740, ???? 1997. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97572.

Stepanov:1969:CAR

[Ste69a] V. E. Stepanov. Combinatorial algebra and random graphs. Theory of Probability and its Applications, 14(3):373–399, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1969:LDC

[Ste69b] V. E. Stepanov. Limit distributions of certain characteristics of random mappings. Theory of Probability and its Applications, 14(4):612–626, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1969:DNV

[Ste69c] V. E. Stepanov. On the distribution of the number of vertices in strata of a random tree. Theory of Probability and its Applications, 14(1):65–78, ???? 1969. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1970:PCR

[Ste70a] V. E. Stepanov. On the probability of connectedness of a random graph G_{⇕}(⊔). Theory of Probability and its Applications, 15(1):55–67, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1970:PTR

[Ste70b] V. E. Stepanov. Phase transitions in random graphs. Theory of Probability and its Applications, 15(2):187–203, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1971:RMS

[Ste71] V. E. Stepanov. Random mappings with a single attracting center. Theory of Probability and its Applications, 16(1):155–162, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1973:SRG

[Ste73] V. E. Stepanov. Structure of the random graphs G_{⇕}(§|⟨). Theory of Probability and its Applications, 17(2):227–242, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stenkin:1984:PSP

[Ste84] D. B. Sten’kin. Preservation of statistical properties for subsequences of random sequences. Theory of Probability and its Applications, 28(4):752–761, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1988:LMI

[Ste88a]  A. V. Stepanov.  On loga-

rithmic moments for interrecord times. Theory of Probability and its Applications, 32(4):708–710, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1988:SFS

[Ste88b]  V. E. Stepanov.  On some

features of the structure of a random graph near a critical point. Theory of Probability and its Applications, 32(4):573–594, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1989:STA

[Ste89] A. K. Stepanov. A similarity theorem and its applications to percolation of random fields on the plane. I. Scaling of percolation. Theory of Probability and its Applications, 33(1): 36–49, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanov:1991:STA

[Ste91] A. K. Stepanov. A similarity theorem and its applications to percolation of random fields on the plane. II. Percolation of random fields on Z^{2}. Theory of Probability and its Applications, 35(3):513–522, September 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 506–514.

Stepanov:1993:LTW

[Ste93] A. V. Stepanov. Limit theorems for weak records. Theory of Probability and its Applications, 37(3):570–574, September 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 586–590.

Stepanov:1994:CTW

[Ste94]  A. V. Stepanov.  A char-

acterization theorem for weak records. Theory of Probability and its Applications, 38(4): 762–764, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 903–905.

Stratonovich:1960:CMP

[Str60] R. L. Stratonovich. Conditional Markov processes. Theory of Probability and its Applications, 5(2):156–178, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Strasser:1978:IBE

[Str78] H. Strasser. Improved bounds for equivalence of Bayes and maximum likelihood estimation. Theory of Probability and its Applications, 22(2):349–361, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Studnev:1960:ADS

[Stu60] Yu. P. Studnev. An approximation to the distribution of sums by infinitely divisible laws. Theory of Probability and its Applications, 5(4):421–425, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Studnev:1965:RKP

[Stu65] Yu. P. Studnev. A remark on the Katz–Petrov theorem. Theory of Probability and its Applications, 10(4):682–684, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Studnev:1967:SGL

[Stu67] Yu. P. Studnev. Some generalizations of limit theorems in probability theory. Theory of Probability and its Applications, 12 (4):668–672, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Studnev:1968:SPA

[Stu68] Yu. P. Studnev. Some properties of accompanying laws for symmetric distribution functions. Theory of Probability and its Applications, 13(4):701–703, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Suchkov:1990:EER

[SU90a] A. P. Suchkov and N. G. Ushakov. Exponential estimates of the rate of decay of the maximum probability in infinitedimensional spaces. Theory of Probability and its Applications, 34(4):735–738, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 802–805.

Suchkov:1990:RDC

[SU90b] A. P. Suchkov and N. G.

Ushakov. The rapid decay of concentration functions of sums of independent random variables. Theory of Probability and its Applications, 34(3):545–548, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 604–607.

Su:1999:WCV

[Su99] Zhong Gen Su. On the weak convergence of vector-valued continuous random processes. Theory of Probability and its Applications, 43(3):463–476, September 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97705.

Sun:1996:CLT

[Sun96] Shan Sun. Central Limit Theorem of the perturbed sample quantile for a sequence of mdependent nonstationary random process. Theory of Probability and its Applications, 40(1):

116–129, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(1), (1995), pp. 143–158.

Sunklodas:1999:LBR

[Sun99]  J. Sunklodas.  On a lower

bound of the rate of convergence in the Central Limit Theorem for m-dependent random fields. Theory of Probability and its Applications, 43(1):162– 169, March 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97678.

Sur:1963:LLN

[Sur63] M. Sur. On the law of large numbers for Markov processes. Theory of Probability and its Applications, 8(2):208–212, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sverchkov:1987:SDD

[Sve87] M. Yu. Sverchkov. On smoothing the distributions of a discretetime renewal process. Theory of Probability and its Applications, 31(4):723–725, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sviridenko:1990:MAL

[Svi90] M. N. Sviridenko. A martingale approach to limit theorems for semi-Markov processes. Theory of Probability and its Applications, 34(3):540–545, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 598–603.

Skorokhod:1973:ACM

[SY73] A. V. Skorokhod and M. I. Yadrenko. On absolute continuity of measures corresponding to homogeneous Gaussian fields. Theory of Probability and its Applications, 18(1):27–40, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sytaya:1967:LTS

[Syt67] G. N. Sytaya. Limit theorems for some functionals of random walks. Theory of Probability and its Applications, 12(3):432–442, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sytaya:1969:ADS

[Syt69] G. N. Sytaya. On admissible displacements of suspended Gaussian measures. Theory of Probability and its Applications, 14 (3):506–509, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sevastyanov:1974:CBP

[SZ74] B. A. Sevast’yanov and A. M.

Zubkov. Controlled branching processes. Theory of Probability and its Applications, 19 (1):14–24, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Szekely:1998:SCR

[SZ98] G. J. Székely and A. Zemplén. Short communication: Remarks on the paper “On a Problem of a Khinchin-Type Decomposition Theorem for Extreme Values” [Teor. Veroyatnost. i Primenen. 39 (1994), no. 2, 395–402; MR 97c:60136] by E. Pancheva. Theory of Probability and its Applications, 42(1):184–185, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97606. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 216–217.

Szasz:1971:LTS

[Sza71] D. Szasz. Limit theorems for stochastic processes stopped at random. Theory of Probability and its Applications, 16(3): 554–555, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Szasz:1973:CLD

[Sza73] D. Szasz. On classes of limit distributions for sums of a random number of identically distributed independent random variables. Theory of Probability and its Applications, 17(3):401–415, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sztrik:1993:AAR

[Szt93] J. Sztrik. Asymptotic analysis of the reliability of a complex renewable standby system with fast repair. Theory of Probability and its Applications, 37(1):101–104, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 132–135.

Szulga:1983:MMS

[Szu83] A. Szulga. On minimal metrics in the space of random variables. Theory of Probability and its Applications, 27(2):424–430, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Szyszkowski:1997:WIP

[Szy97] I. Szyszkowski. Weak invariance principle for integral type functionals. Theory of Probability and its Applications, 41(1):185–192, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 226–233.

Takacs:1956:SES

[Tak56] Lajos Takacs. On the sequence of events, selected by a counter from a recurrent process of events. Theory of Probability and its Applications, 1(1):81–91, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Takacs:1957:SPC

[Tak57] Lajos Takács. On a stochastic process concerning some waiting time problems. Theory of Probability and its Applications, 2(1):90–103, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Takacs:1958:STP

[Tak58] L. Takács. On a sojourn time problem. Theory of Probability and its Applications, 3(1):58–65, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Taksar:1968:PSE

[Tak68] M. I. Taksar. On a property of a sequence of events. Theory of Probability and its Applications, 13(3):501–504, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Taksar:1976:ECR

[Tak76] M. I. Taksar. On events connected with reaching a set by sample paths of a stochastic process. Theory of Probability and its Applications, 21(1):144–148, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Taksar:1977:FWR

[Tak77] M. I. Taksar. A formula for wanderings of a regular Markov process. Theory of Probability and its Applications, 21(4):818–824, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tanaka:1970:PDM

[Tan70] Hiroshi Tanaka. Purely discontinuous Markov processes with nonlinear generators and their propagation of chaos. Theory of Probability and its Applications, 15(4):581–603, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Taraskin:1985:BLR

[Tar85] A. F. Taraskin. On the behavior of the likelihood ratio of semimartingales. Theory of Probability and its Applications, 29 (3):452–464, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tarashchanskii:1990:UEP

[Tar90a] M. T. Tarashchanskiĭ. On the uniqueness of the extension of a probability measure. Theory of Probability and its Applications, 34(3):549–553, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 607–612.

Tarieladze:1990:TDC

[Tar90b] V. I. Tarieladze. A topological description of characteristic functionals in some groups. Theory of Probability and its Applications, 34(4):658–667, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 719–730.

Timonin:1986:MCE

[TC86] V. I. Timonin and O. M. Chernomordik. A method for calculating the exact distribution of Kolmogorov–Smirnov statistics under Lehmann alternatives. Theory of Probability and its Applications, 30(3):608–610, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(3), (1985), pp. 572–573.

Teicher:1966:HMR

[Tei66] H. Teicher. Higher moments of randomly stopped sums. Theory of Probability and its Applications, 11(1):160–165, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tempelman:1973:EHG

[Tem73a] A. A. Tempelman. The ergodicity of homogeneous Gaussian random fields on homogeneous spaces. Theory of Probability and its Applications, 18(1):173–175, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tempelman:1973:GET

[Tem73b] A. A. Tempel’man. A generalization of an ergodic theorem of Hopf. Theory of Probability and its Applications, 17(2):363–365, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

TenVregelaar:1991:NCN

[Ten91] J. M. Ten Vregelaar. Note on the convergence to normality of quadratic forms in independent variables. Theory of Probability and its Applications, 35(1): 177–179, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 184–186.

Teplyaev:1995:CAR

[Tep95] A. V. Teplyaev. Continual analogues of random polynomials that are orthogonal on a circle. Theory of Probability and its Applications, 39(3):476–489, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 588–604.

Terpstra:1993:CAO

[Ter93] T. J. Terpstra. Construction of asymptotically optimal and independent nonparametric experimental design tests in the Cox regression model. Theory of Probability and its Applications, 37 (1):173–175, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 208–210.

Teugels:1975:SCP

[Teu75] J. L. Teugels. The subexponential class of probability distributions. Theory of Probability and its Applications, 19(4):821–822, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1962:SIA

[TF62] V. N. Tutubalin and M. I. Freidlin. On the structure of the infinitesimal σ-algebra of a Gaussian process. Theory of Probability and its Applications, 7 (2):196–199, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tyurin:1991:TRN

[TF91] Yu. N. Tyurin and V. E. Figurnov. On the testing of random number generators. Theory of Probability and its Applications, 35(1):180–184, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Thang:1993:VSR

[Th93] – Dang Hùng Thang. Vector symmetric random measures and random integrals. Theory of Probability and its Applications, 37(3): 415–423, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 526–533.

Thorisson:1991:CCR

[Tho91] Hermann Thorisson. The classical coupling, a refinement. Theory of Probability and its Applications, 35(4):810–819, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 809–817.

Thorburn:1993:CIU

[Tho93a] D. Thorburn. Confidence intervals using a modified bootstrap. Theory of Probability and its Applications, 37(2):363–365, June 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 390–392.

Thorisson:1993:CSC

[Tho93b] H. Thorisson. From coupling to shift-coupling. Theory of Probability and its Applications, 37 (1):105–112, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 95–104.

Thu:1985:PBS

[Thu85] Nguen Van Thu. Prediction of Banach space valued strictly stationary sequences. Theory of Probability and its Applications, 29(2):334–344, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tien:1977:RCS

[Tie77] Nguyen Xuy Tien. Remarks on the convergence of series of independent random elements with values in the spaces l_{p}, 1 ≤ p < +∞. Theory of Probability and its Applications, 21(2):386–388, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tien:1979:SML

[Tie79] Nguen Zuy Tien. The structure of measurable linear functionals on Banach spaces with Gaussian measures. Theory of Probability and its Applications, 24(1):165–168, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tien:1980:KTS

[Tie80] Nguen Zuy Tien. On Kolmogorov’s three series theorem and mean square convergence of martingales in a Banach space. Theory of Probability and its Applications, 24(4):797–808, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhov:1978:OSE

[Tik78] M. S. Tikhov. On optimal sequential estimation plans for non-quadratic loss functions. Theory of Probability and its Applications, 23(1):132–138, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhomirov:1981:CRC

[Tik81] A. N. Tikhomirov. On the convergence rate in the Central Limit Theorem for weakly dependent random variables. Theory of Probability and its Applications, 25(4):790–809, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhov:1982:SED

[Tik82] M. S. Tikhov. On the sequential estimation of the drift coefficient of a diffusion process with quadratic and nonquadratic losses. Theory of Probability and its Applications, 26 (3):607–614, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhov:1985:LDE

[Tik85] M. S. Tikhov. On limit distributions of estimates for two-sided censored samples with a discontinuous density. Theory of Probability and its Applications, 29 (2):363–368, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhomirov:1987:DMS

[Tik87] A. N. Tikhomirov. On the distribution of the maximum sum of weakly dependent variables. Theory of Probability and its Applications, 31(4):733–738, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhov:1988:EPD

[Tik88] S. M. Tikhov. Estimates of parameters of densities that are singular in a quantile neighbourhood, based on some order statistics. Theory of Probability and its Applications, 32(2):353–358, ???? 1988. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhomirov:1992:ANA

[Tik92a] A. N. Tikhomirov. On the accuracy of normal approximation of the probability of hitting a ball of sums of weakly dependent Hilbert space valued random variables I. Theory of Probability and its Applications, 36(4): 738–751, December 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 699–710.

Tikhov:1992:RTD

[Tik92b] M. S. Tikhov. On reduction of test duration in the case of censoring. Theory of Probability and its Applications, 36(3): 629–633, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 606–609.

Tikhov:1993:AE

[Tik93] M. S. Tikhov. Asymptotics of T -estimators. Theory of Probability and its Applications, 37(4): 644–657, December 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tikhomirov:1994:ANA

[Tik94] A. N. Tikhomirov. On the accuracy of normal approximation of the probability of hitting a ball of sums of weakly dependent Hilbert space valued random variables II. Theory of Probability and its Applications, 38(1):80–94, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 110–127.

Tikhov:1998:ASE

[Tik98] M. S. Tikhov. Asymptotics of statistical estimates constructed from censored samples of distributions with regularly varying tails. Theory of Probability and its Applications, 42(3):495–512, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 531–552.

Timonin:1984:EDS

[Tim84] V. I. Timonin. Exact distributions of Smirnov’s statistics for a class of alternatives for complete censored data. Theory of Probability and its Applications, 28(4):795–797, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 28(4), (1983), pp. 758–760.

Timofeev:1985:RMT

[Tim85] E. A. Timofeev. Random minimal trees. Theory of Probability and its Applications, 29 (1):134–141, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Timonin:1988:LDS

[Tim88] V. I. Timonin. On the limit distribution of statistics of a nonparametric test. Theory of Probability and its Applications, 32 (4):721–724, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Timofeev:1989:FEL

[Tim89] E. A. Timofeev. On finding the expected length of a random minimal tree. Theory of Probability and its Applications, 33 (2):361–365, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 383–387.

Timashev:1999:AED

[Tim99] A. N. Timashev. On asymptotic expansions in the domain of large deviations for binomial and Poisson distributions. Theory of Probability and its Applications, 43(1):89–98, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97668.

Ting:1962:AI

[Tin62a] Hu Kuo Ting. On the amount of information. Theory of Probability and its Applications, 7 (4):439–447, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ting:1962:ISS

[Tin62b] Hu Kuo Ting. On the information stability of a sequence of channels. Theory of Probability and its Applications, 7 (3):258–269, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Titov:1982:DCI

[Tit82] A. N. Titov. On determining the convolution of identical distribution functions by its values on the half-axis. Theory of Probability and its Applications, 26 (3):599–600, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsitritskii:1979:RMT

[TM79] O. E. Tsitritskii and L. N.

Mironyuk. On the Riesz– Markov theorem for vectorvalued stochastic measures. Theory of Probability and its Applications, 23(4):811–814, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tolmachev:1999:SSI

[Tol99]  N. A. Tolmachev.  On the

smoothness and singularity of invariant measures and transition probabilities of infinitedimensional diffusions. Theory of Probability and its Applications, 43(4):655–664, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97724.

Tomko:1971:ERT

[Tom71] J. Tomkó. On the estimation of the remainder term in the Central Limit Theorem for sums of a random number of summands. Theory of Probability and its Applications, 16(1):167–175, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tomko:1990:TSM

[Tom90] Ĭ. Tomko. On the theory of semi-Markov processes with common phase space. Theory of Probability and its Applications, 34 (2):275–289, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 314–329.

Topsoe:1967:CBC

[Top67] F. Topsøe. On the connection between P -continuity and P -uniformity in weak convergence. Theory of Probability and its Applications, 12(2):241–250, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Topchii:1983:AEP

[Top83] V. A. Topchii. Asymptotics of the extinction probability for critical general branching processes. Theory of Probability and its Applications, 27(4):715–733, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Topchii:1989:MDT

[Top89] V. A. Topchiĭ. Moderate deviations for the total number of particles in critical branching processes. Theory of Probability and its Applications, 33(2):382–385, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 406–409.

Torondzhadze:1988:CAS

[Tor88] T. A. Torondzhadze. Coincidence of the σ-algebras in a scheme of partially observable diffusion processes. Theory of Probability and its Applications, 32(3): 527–531, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Touati:1992:CLF

[Tou92]  A. Touati.  Sur la conver-

gence en loi fonctionnelle de suites de semimartingales vers un mélange de mouvements browniens. (French) [On the convergence in a functional law of semimartingale sets towards a mixture of Brownian movement]. Theory of Probability and its Applications, 36(4):752– 771, December 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(4), (1991), pp. 744–763.

Touati:1994:TTC

[Tou94] A. Touati. Two theorems on convergence in distribution for stochastic integrals and statistical applications. Theory of Probability and its Applications, 38 (1):95–117, March 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 128–153.

Tse-Pei:1957:LEC

[TP57] Chiang Tse-Pei. On the linear extrapolation of a continuous homogeneous random field. Theory of Probability and its Applications, 2(1):58–89, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tse-Pei:1958:NDA

[TP58] Chiang Tse-Pei. A note on the definition of the amount of information. Theory of Probability and its Applications, 3(1):93–96, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tse-Pei:1959:ERC

[TP59] Chiang Tse-Pei. On the estimation of regression coefficients of a continuous parameter time series with a stationary residual. Theory of Probability and its Applications, 4(4):373–390, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Trigub:1990:CCF

[Tri90] R. M. Trigub. A criterion for a characteristic function and a Polyá-type criterion for radial functions of several variables. Theory of Probability and its Applications, 34(4):738–742, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 805–810.

Trofimov:1991:NSL

[Tro91] E. I. Trofimov. Normed supports and limit theorems for semimartingales. Theory of Probability and its Applications, 35(3): 523–538, September 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(3), (1990), pp. 515–530.

Trofimov:1993:MCM

[Tro93] E. I. Trofimov. On a martingale characterization of measures and on general functional limit theorems. Theory of Probability and its Applications, 37(3): 574–576, September 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 591–592.

Trukhachev:1966:ATR

[Tru66] A. A. Trukhachev. Addendum:

Tables of the Rayleigh–Rice distribution. Theory of Probability and its Applications, 11(4):642, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsaregradskii:1958:NCS

[Tsa58a] I. P. Tsaregradskii. A note on the capacity of a stationary channel with finite memory. Theory of Probability and its Applications, 3(1):79–91, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsaregradskii:1958:UAB

[Tsa58b] I. P. Tsaregradskii. On uniform approximation of the binomial distribution by infinitely divisible laws. Theory of Probability and its Applications, 3 (4):434–438, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsaregradskii:1970:PSN

[Tsa70] I. P. Tsaregradskii. On a problem of search by networks. Theory of Probability and its Applications, 15(2):315–319, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsarkov:1977:AEM

[Tsa77] E. F. Tsar’kov. Asymptotic exponential mean-square stability of the trivial solution of stochastic functional-differential equations. Theory of Probability and its Applications, 21(4):850–854, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsarkov:1979:EST

[Tsa79] E. F. Tsar’kov. Exponential pstability of trivial solutions of stochastic functional-differential equations. Theory of Probability and its Applications, 23 (2):426–429, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsaregradskii:1983:QSI

[Tsa83]  I. P. Tsaregradskii.  On a

queueing system with an infinite number of locally interacting servers. Theory of Probability and its Applications, 27(3):624–628, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsirelson:1976:ESD

[Tsi76a] B. S. Tsirel’son. An example of a stochastic differential equation having no strong solution. Theory of Probability and its Applications, 20(2):416–418, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsirelson:1976:SPL

[Tsi76b] B. S. Tsirel’son. Some properties of lacunary series and Gaussian measures that are connected with uniform versions of the egorov and luzin properties. Theory of Probability and its Applications, 20(3):652–655, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsirelson:1976:DDM

[Tsi76c] V. S. Tsirel’son. The density of the distribution of the maximum of a Gaussian process. Theory of Probability and its Applications, 20(4):847–856, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsitsiashvili:1976:PLM

[Tsi76d] G. Sh. Tsitsiashvili. Piecewiselinear Markov chains and analysis of their stability. Theory of Probability and its Applications, 20(2):337–350, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsitovich:1978:EDD

[Tsi78a] I. I. Tsitovich. On estimating the drift of a diffusion process from observations in a domain. Theory of Probability and its Applications, 22(4):851–858, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsitovich:1978:TFE

[Tsi78b] I. I. Tsitovich. On the time of first exit from a domain. Theory of Probability and its Applications, 23(1):117–129, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsirelson:1983:GAM

[Tsi83] B. S. Tsirel’son. A geometrical approach to maximum likelihood estimation for infinitedimensional Gaussian location.

I. Theory of Probability and its Applications, 27(2):411–418, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsitovich:1985:SDE

[Tsi85] I. I. Tsitovich. Sequential design of experiments for hypothesis testing. Theory of Probability and its Applications, 29 (4):814–817, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsirelson:1986:GAM

[Tsi86] B. S. Tsirel’son. A geometric approach to maximum likelihood estimation for infinitedimensional Gaussian location. II. Theory of Probability and its Applications, 30(4):820–828, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsirelson:1987:GAM

[Tsi87] B. S. Tsirel’son. A geometric approach to maximum likelihood estimation for infinitedimensional location. III. Theory of Probability and its Applications, 31(3):470–483, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsitsiashvili:1991:SMI

[Tsi91]  G. Sh. Tsitsiashvili.  Some

methods of investigating regeneration times in queueing networks. Theory of Probability and its Applications, 35(2):387–390, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 376–379.

Tsigroshvili:1997:MCI

[Tsi97] Z. P. Tsigroshvili. On moduli of continuity of innovation evolvents of empirical fields. Theory of Probability and its Applications, 41(3):601–605, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 689–693.

Tsukanov:1981:TOE

[Tsu81] A. V. Tsukanov. On tests of optimality of experimental designs for a regression model with random variables. Theory of Probability and its Applications, 26 (1):173–177, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsybakov:1988:CBK

[Tsy88] A. B. Tsybakov. On the choice of the bandwidth in kernel nonparametric regression. Theory of Probability and its Applications, 32(1):142–148, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Thang:1980:SSM

[TT80] Dang Hung Thang and Nguyen Duy Tien. On symmetric stable measures on spaces l_{p}(1 ≤ p ≤ +∞). Theory of Probability and its Applications, 25(1):118–127, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Thang:1982:CMG

[TT82] Dang Hung Thang and Nguen Duy Tien. On convergence of martingales and geometric properties of Banach spaces. Theory of Probability and its Applications, 26 (2):379–385, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Thang:1983:MSC

[TT83] Dang Hung Thang and Nguyen Duy Tien. Mappings of stable cylindrical measures in Banach space. Theory of Probability and its Applications, 27(3):525–535, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tien:1978:CFS

[TTC78] Nguyen Zui Tien, V. I. Tarieladze, and S. A. Chobanjan. On compactness of families of secondorder measures in a Banach space. Theory of Probability and its Applications, 22(4):805–810, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tucker:1963:ECD

[Tuc63] Howard G. Tucker. An estimate of the compounding distribution of a compound Poisson distribution. Theory of Probability and its Applications, 8 (2):195–200, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tumanyan:1956:ADC

[Tum56] S. Kh. Tumanyan. Asymptotic distribution of the χ^{2} criterion when the number of observations and number of groups increase simultaneously. Theory of Probability and its Applications, 1 (1):117–131, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Turchin:1971:CMI

[Tur71] V. F. Turchin. On the computation of multidimensional integrals by the Monte–Carlo method. Theory of Probability and its Applications, 16(4): 720–724, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Turner:1991:ANK

[Tur91] D. W. Turner. Addendum: a note on a Kolmogorov inequality for the binomial distribution. Theory of Probability and its Applications, 35(1):193, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1961:IRM

[Tut61] V. N. Tutubalin. Invariant random measures on the sphere. Theory of Probability and its Applications, 6(1):113–117, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1962:LBC

[Tut62] V. N. Tutubalin. On the limit behavior of compositions of measures in the plane and space of Lobachevsky. Theory of Probability and its Applications, 7 (2):189–196, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1964:CMS

[Tut64] V. N. Tutubalin. Compositions of measures on the simplest nilpotent group. Theory of Probability and its Applications, 9 (3):479–487, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1965:LTP

[Tut65] V. N. Tutubalin. On limit theorems for the product of random matrices. Theory of Probability and its Applications, 10 (1):15–27, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1967:MWS

[Tut67] V. N. Tutubalin. On measures whose support is generated by a Lie algebra. Theory of Probability and its Applications, 12 (1):134–138, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1968:APM

[Tut68a] V. N. Tutubalin. Approximation of probability measures in variation and products of random matrices. Theory of Probability and its Applications, 13 (1):65–83, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1968:RWL

[Tut68b] V. N. Tutubalin. On random walk in the Lobachevsky plane. Theory of Probability and its Applications, 13(3):487–490, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1969:ISM

[Tut69a] V. N. Tutubalin. An investigation of some mathematical models for radio waveguides with random irregularities. Theory of Probability and its Applications, 14(4):547–566, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1969:STT

[Tut69b] V. N. Tutubalin. Some theorems of the type of the strong law of large numbers. Theory of Probability and its Applications, 14 (2):313–319, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1971:MWP

[Tut71] V. A. Tutubalin. Multimode waveguides and probability distributions on a symplectic group. Theory of Probability and its Applications, 16(4):631–642, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1973:RRM

[Tut73] V. N. Tutubalin. Representation of random matrices in orispherical coordinates and its application to telegraph equations. Theory of Probability and its Applications, 17(2):255–268, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1974:DSR

[Tut74] V. M. Tutubalin. On dynamic systems with random perturbations. Theory of Probability and its Applications, 18(4):678–693, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1978:VLL

[Tut78] V. N. Tutubalin. A variant of the local limit theorem for products of random matrices. Theory of Probability and its Applications, 22(2):203–214, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tutubalin:1979:NFL

[Tut79] V. N. Tutubalin. A new form of the law of large numbers. Theory of Probability and its Applications, 24(1):1–15, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Topchii:1998:MCG

[TV98] V. A. Topchii and V. A.

Vatutin. Maximum of the critical Galton–Watson processes and left-continuous random walks. Theory of Probability and its Applications, 42(1):17–27, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97590. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 21–34.

Tyurin:1970:PHT

[Tyu70a] Yu. N. Tyurin. On parametric hypothesis testing with nonparametric tests. Theory of Probability and its Applications, 15 (4):722–726, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tyurin:1970:EDF

[Tyu70b] Yu. N. Tyurin. On the estimation of a distribution function. Theory of Probability and its Applications, 15(3):549–550, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tyurin:1974:THL

[Tyu74] Yu. N. Tyurin. Testing the hypothesis that a large-sized multidimensional sample is normal. Theory of Probability and its Applications, 18(3):618–622, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ubaidullaeva:1975:TOO

[Uba75] M. Ubaidullaeva. On testing outlying observations. Theory of Probability and its Applications, 19(4):831–834, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Uglanov:1989:SDF

[Ugl89] A. V. Uglanov. On the smoothness of distributions of functionals of random processes. Theory of Probability and its Applications, 33(3):500–508, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 535–544.

Ulyanov:1977:NUE

[Ul’77] V. V. Ul’yanov. A non-uniform estimate for the speed of convergence in the Central Limit Theorem in R. Theory of Probability and its Applications, 21 (2):270–282, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ulyanov:1979:AMP

[Ul’79a] V. V. Ul’yanov. Addendum:

On more precise convergence rate estimates in the Central Limit Theorem. Theory of Probability and its Applications, 24(1):237, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ulyanov:1979:MPC

[Ul’79b] V. V. Ul’yanov. On more precise convergence rate estimates in the Central Limit Theorem. Theory of Probability and its Applications, 23(3):660–663, ???? 1979.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ulyanov:1987:AED

[Ul’87] V. V. Ul’yanov. Asymptotic expansions for distributions of sums of independent random variables in H. Theory of Probability and its Applications, 31 (1):25–39, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ulyanov:1991:LEN

[Ul′91]  V. V. Ul′yanov.  Letter to

the editors: “Normal approximation in a Hilbert space. I, II” [Teor. Veroyatnost. i Primenen. 33 (1988), no. 2, 225–245; MR 89k:60008; ibid. 33 (1988), no. 3, 508–521; MR 90b:60010] by B. A. Zalesskiĭ, V. V. Sazonov and Ul′yanov. Theory of Probability and its Applications, 35 (1):191–192, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 187–187.

Ulanovskii:1984:DCK

[Ula84] A. M. Ulanovskii. Description of the class I_{0} in Kingman’s semigroup of standard pfunctions. Theory of Probability and its Applications, 28(3): 561–572, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor.

Veroyatnost. i Primenen., 28(3), (1983), pp. 533–543.

Ulanovskii:1988:UDC

[Ula88] A. M. Ulanovskii. On unique determination of convolutions of one-type distribution functions by restriction to a halfline. Theory of Probability and its Applications, 32(1):125–127, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Urbanik:1956:SPW

[Urb56] K. Urbanik. Stochastic processes whose sample functions are distributions. Theory of Probability and its Applications, 1 (1):132–134, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Urinov:1993:THU

[Uri93] I. K. Urinov. Tests of homogeneity under small grouping. martingale limit theorems. Theory of Probability and its Applications, 37(4):658–672, December 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(4), (1992), pp. 676–694.

Ushakov:1978:QSE

[Ush78] V. G. Ushakov. A queueing system with Erlang incoming flow and relative priority. Theory of Probability and its Applications, 22(4):841–846, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakov:1982:SIC

[Ush82] N. G. Ushakov. Some inequalities for characteristic functions of unimodal distributions. Theory of Probability and its Applications, 26(3):595–598, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakov:1983:PR

[Ush83] N. G. Ushakov. On a problem of Rényi. Theory of Probability and its Applications, 27(2):361–362, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakova:1984:SDP

[Ush84] A. P. Ushakova. On the stability of decompositions of probability laws. Theory of Probability and its Applications, 28(3):603–606, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakov:1986:UEM

[Ush86] N. G. Ushakov. Upper estimates of maximum probability for sums of independent random vectors. Theory of Probability and its Applications, 30 (1):38–49, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Utev:1990:CMI

[Ute90] S. A. Utev. Cumulants and moment inequalities. Theory of Probability and its Applications, 34(4):742–747, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 810–815.

Utev:1991:CLT

[Ute91] S. A. Utev. On the Central Limit Theorem for φ-mixing arrays of random variables. Theory of Probability and its Applications, 35(1):131–139, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakov:1985:IPD

[UU85a] V. G. Ushakov and N. G.

Ushakov. Indecomposable probability distributions on groups. Theory of Probability and its Applications, 29(2):356–359, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakov:1985:ILI

[UU85b] V. G. Ushakov and N. G.

Ushakov. On indecomposable laws with infinitely divisible projections. Theory of Probability and its Applications, 29(3): 596–598, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ushakov:1987:DBP

[UU87] V. G. Ushakov and N. G.

Ushakov. On decomposing blends of probability distributions. Theory of Probability and its Applications, 31(2):319–322, ???? 1987. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1997:OBV

[V^{+}97] V. A. Vatutin et al. Obituary: Boris Vladimirovich Gnedenko. Theory of Probability and its Applications, 41(2):326–327, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/97553. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 393–395.

Vakhaniya:1964:ND

[Vak64] N. N. Vakhaniya. On normal distributions in l_{p}. Theory of Probability and its Applications, 9(4):665, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhaniya:1966:NDP

[Vak66] N. N. Vakhaniya. On nondegenerate probability distributions in l_{p} spaces (1 ≤ p < ∞). Theory of Probability and its Applications, 11(3):463–467, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhaniya:1967:PPO

[Vak67] N. N. Vakhaniya. A probabilistic problem for the one-dimensional heat equation. Theory of Probability and its Applications, 12 (4):666–667, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhaniya:1990:KET

[Vak90] N. N. Vakhaniya. A. N. Kolmogorov and the evolution of the theory of probability distributions in linear spaces. Theory of Probability and its Applications, 34(1):170–174, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhania:1994:CFG

[Vak94a] N. N. Vakhania. Canonical factorization of Gaussian covariance operators and some of its applications. Theory of Probability and its Applications, 38(3):498–505, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 481–490.

Vakhania:1994:EPP

[Vak94b] N. N. Vakhania. Elementary proof of Polya’s characterization theorem and of the necessity of second moment in the CLT. Theory of Probability and its Applications, 38(1):166–168, March 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 168–171.

Vakhania:1999:RVV

[Vak99] N. N. Vakhania. Random vectors with values in quaternion Hilbert spaces. Theory of Probability and its Applications, 43 (1):99–115, March 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97669.

Vallander:1973:CWD

[Val73] S. S. Vallander. Calculation of the Wasserstein distance between probability distributions on the line. Theory of Probability and its Applications, 18(??):1, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 18, (1973), pp. 824– 827.

Vallander:1974:RPR

[Val74a] S. S. Vallander. A remark on the papers by R. L. Dobrushin and H. Kesten. Theory of Probability and its Applications, 18 (2):375–378, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vallender:1974:CWD

[Val74b] S. S. Vallender. Calculation of the Wasserstein distance between probability distributions on the line. Theory of Probability and its Applications, 18(4):784–786, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vallender:1982:ACW

[Val82] S. S. Vallender. Addendum: Calculation of the Wasserstein distance between probability distributions on the line. Theory of Probability and its Applications, 26(2):435, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vandemaele:1983:LDP

[Van83] M. Vandemaele. On large deviation probabilities for U-statistics. Theory of Probability and its Applications, 27(3):614, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

vanLambalgen:1994:AR

[van94] M. van Lambalgen. Axioms for randomness. Theory of Probability and its Applications, 38 (3):562–566, ???? 1994. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 670–674.

Vasilev:1970:CES

[Vas70] N. B. Vasil’ev. Correlation equations for the stationary measure of a Markov chain. Theory of Probability and its Applications, 15(3):521–525, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vasudeva:1980:FLI

[Vas80] R. Vasudeva. Functional law of the iterated logarithm for I.I.D. random variables attracted to a completely asymmetric stable law. Theory of Probability and its Applications, 24(4):833–838, ???? 1980. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1974:APF

[Vat74] V. A. Vatutin. The asymptotic probability of the first degeneration for branching processes with immigration. Theory of Probability and its Applications, 19 (1):25–34, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1977:ABN

[Vat77a] V. A. Vatutin. Asymptotic behavior of the non-extinction probability of a critical branching process. Theory of Probability and its Applications, 22 (1):140–146, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1977:DLD

[Vat77b] V. A. Vatutin. Discrete limit distributions of the number of particles in critical Bellman-harris branching process. Theory of Probability and its Applications, 22(1):146–152, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1977:LTC

[Vat77c] V. A. Vatutin. Limit theorems for a critical Bellman-harris branching process with infinite variance. Theory of Probability and its Applications, 21(4): 839–842, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1977:CBH

[Vat77d] V. M. Vatutin. A critical Bellman-harris branching process with immigration and several types of particles. Theory of Probability and its Applications, 21(2):435–442, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1978:CGW

[Vat78] V. A. Vatutin. A critical Galton– Watson branching process with emigration. Theory of Probability and its Applications, 22 (3):465–481, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1979:LTC

[Vat79] V. A. Vatutin. A limit theorem for a critical Bellman-harris branching process with several types of particles and infinite second moments. Theory of Probability and its Applications, 23 (4):776–788, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1980:DLD

[Vat80] V. A. Vatutin. Discrete limit distributions of the number of particles in a Bellman-harris branching process with several types of particles. Theory of Probability and its Applications, 24 (3):509–520, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1981:CCB

[Vat81] V. A. Vatutin. On a class of critical Bellman-harris branching processes with several types of particles. Theory of Probability and its Applications, 25 (4):760–771, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1982:CLT

[Vat82] V. A. Vatutin. A class of limit theorems for a critical Bellmanharris branching process. Theory of Probability and its Applications, 26(4):806–812, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1987:CBH

[Vat87a]  V. A. Vatutin.  A critical

Bellman-harris branching process with particles of final type. Theory of Probability and its Applications, 31(3):428–438, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1987:SCR

[Vat87b] V. A. Vatutin. Sufficient conditions for regularity of Bellmanharris branching processes. Theory of Probability and its Applications, 31(1):50–57, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1994:DDR

[Vat94a] V. A. Vatutin. The distribution of the distance to the root of the minimal subtree containing all the vertices of a given height. Theory of Probability and its Applications, 38(2):330–341, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(2), (1993), pp. 273–287.

Vatutin:1994:TNP

[Vat94b] V. A. Vatutin. The total number of particles in a reduced Bellmanharris branching process. Theory of Probability and its Applications, 38(3):567–571, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(3), (1993), pp. 624–629.

Vatutin:1995:BPF

[Vat95] V. A. Vatutin. Branching processes with final types of particles, and random trees. Theory of Probability and its Applications, 39(4):628–641, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 699–715.

Vatutin:1996:MSR

[Vat96] V. A. Vatutin. On the maximum of a simple random walk. Theory of Probability and its Applications, 40(2):398–402, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(2), (1995), pp. 412–417.

Viktorova:1966:SGT

[VC66] I. I. Viktorova and V. P. Chistyakov. Some generalizations of the test of empty boxes. Theory of Probability and its Applications, 11(2):270–276, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vapnik:1971:UCR

[VC71] V. N. Vapnik and A. Ya. Chervonenkis. On the uniform convergence of relative frequencies of events to their probabilities. Theory of Probability and its Applications, 16(2):264–280, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vapnik:1982:NSC

[VC82] V. N. Vapnik and A. Ya. Chervonenkis. Necessary and sufficient conditions for the uniform convergence of means to their expectations. Theory of Probability and its Applications, 26 (3):532–553, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vasudeva:1993:LIL

[VD93] R. Vasudeva and G. Divanji. The law of the iterated logarithm for delayed sums under a nonidentically distributed setup. Theory of Probability and its Applications, 37(3):497–506, September 1993. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 534–542.

Veits:1993:SHE

[Veĭ93] E. Veĭt·s. A stochastic heat equation for stationary freeway traffic flow. Theory of Probability and its Applications, 37(1):185–188, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 153–156.

Venttsel:1959:BCM

[Ven59] A. D. Venttsel’. On boundary conditions for multidimensional diffusion processes. Theory of Probability and its Applications, 4(2):164–177, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Venttzel:1961:ACT

[Ven61] A. D. Venttzel’. On the absolute continuity of transition probabilities of a one-dimensional diffusion process. Theory of Probability and its Applications, 6 (4):404–410, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Venttsel:1965:FWC

[Ven65]  A. D. Venttsel’.  On func-

tions which are continuous along trajectories of a Wiener process. Theory of Probability and its Applications, 10(4):662–664, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Venttsel:1968:ACT

[Ven68] A. D. Venttsel’. On the absolute continuity of transition probabilities of multidimensional generalized Brownian motion. Theory of Probability and its Applications, 13(1):1–14, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ventsel:1977:RLTa

[Ven77a] A. D. Ventsel’. Rough limit theorems on large deviations for Markov stochastic processes. I. Theory of Probability and its Applications, 21(2):227–242, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ventsel:1977:RLTb

[Ven77b] A. D. Ventsel’. Rough limit theorems on large deviations for Markov stochastic processes, II. Theory of Probability and its Applications, 21(3):499–512, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ventsel:1980:RLT

[Ven80] A. D. Ventsel’. Rough limit theorems on large deviations for Markov stochastic processes. III. Theory of Probability and its Applications, 24(4):675–692, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ventkus:1982:AGM

[Ven82] V. Yu. Ventkus. Analyticity of Gaussian measures. Theory of Probability and its Applications, 27(1):155–161, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Venttsel:1983:RLT

[Ven83] A. D. Venttsel’. Rough limit theorems on large deviations for Markov stochastic processes. IV. Theory of Probability and its Applications, 27(2):215–234, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Venttsel:1986:ICM

[Ven86] A. D. Venttsel’. Infinitesimal characteristics of Markov processes in a function space which describes the past. Theory of Probability and its Applications, 30(4):661–676, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 30(4), (1985), pp. 625–639.

Venttsel:1990:RFC

[Ven90] A. D. Venttsel′. Refinement of the functional Central Limit Theorem for stationary processes. Theory of Probability and its Applications, 34(3):402–415, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 451–464.

Verbitskaya:1964:CSL

[Ver64a] I. N. Verbitskaya. On conditions for the strong law of large numbers to be applicable to second order stationary processes. Theory of Probability and its Applications, 9(2):325–331, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vershik:1964:SCP

[Ver64b] A. M. Vershik. Some characteristic properties of Gaussian stochastic processes. Theory of Probability and its Applications, 9(2):353–356, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Verbickaja:1966:CAS

[Ver66]  I. N. Verbickaja.  On con-

ditions for the applicability of the strong law of large numbers to wide sense stationary processes. Theory of Probability and its Applications, 11(4):632–636, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1980:SSS

[Ver80] A. Yu. Veretennikov. On the strong solutions of stochastic differential equations. Theory of Probability and its Applications, 24(2):354–366, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1982:SWS

[Ver82] A. Yu. Veretennikov. On strong and weak solutions of onedimensional stochastic equations with boundary conditions. Theory of Probability and its Applications, 26(4):670–686, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1983:CES

[Ver83] A. Yu. Veretennikov. On the criteria for existence of a strong solution of a stochastic equation. Theory of Probability and its Applications, 27(3):441–449, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1988:BMR

[Ver88a] A. Yu. Veretennikov. Bounds for the mixing rate in the theory of stochastic equations. Theory of Probability and its Applications, 32(2):273–281, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1988:SSI

[Ver88b] A. Yu. Veretennikov. On strong solutions of Itô stochastic equations with jumps. Theory of Probability and its Applications, 32(1):148–152, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1991:LDS

[Ver91] A. Yu. Veretennikov. On large deviations of the sojourn time for an ergodic process. Theory of Probability and its Applications, 35(2):390–393, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 340–343.

Veretennikov:1992:LDS

[Ver92] A. Yu. Veretennikov. On large deviations for systems of Itô stochastic equations. Theory of Probability and its Applications, 36(4):772–782, December 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:1994:LDA

[Ver94] A. Yu. Veretennikov. On large deviations for additive functionals of Markov processes. I. Theory of Probability and its Applications, 38(4):706–719, ???? 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(4), (1993), pp. 758–774.

Veretennikov:1999:LDS

[Ver99a] A. Yu. Veretennikov. On large deviations for stochastic differential equations with a small diffusion and averaging. Theory of Probability and its Applications, 43(2):335–337, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97696.

Veretennikov:1999:LDA

[Ver99b] A. Yu. Veretennikov. On large deviations in the averaging principle for stochastic differential equations with “Complete dependence”. Theory of Probability and its Applications, 43(4): 664–666, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97718.

Vetrov:1981:LSD

[Vet81] L. G. Vetrov. On linearization of stochastic differential equations of optimal nonlinear filtering. Theory of Probability and its Applications, 25(2):393–401, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vetrov:1982:FIE

[Vet82] L. G. Vetrov. On filtering, interpolation and extrapolation of semimartingales. Theory of Probability and its Applications, 27 (1):24–36, ???? 1982. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradov:1990:LDS

[VG90] V. V. Vinogradov and V. V.

Godovan′chuk. Large deviations of sums of independent random variables without several maximal summands. Theory of Probability and its Applications, 34 (3):512–515, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 569–571.

Vinogradskaya:1984:SMD

[VGR84] T. M. Vinogradskaya, B. A. Geninson, and A. A. Rubchinskii. Semi-Markov decisionmaking processes with vector gains. Theory of Probability and its Applications, 28(1):191–193, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Viand:1991:RIK

[Via91] M.-C. Viand. Random iterations and Kalman filtering. Theory of Probability and its Applications, 35(4):737–745, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 740–748.

Vilenkin:1959:EMS

[Vil59] S. Ya. Vilenkin. On the estimation of the mean in stationary processes. Theory of Probability and its Applications, 4 (4):415–416, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vilkas:1963:ADV

[Vil63] E. I. Vilkas. Axiomatic definition of the value of a matrix game. Theory of Probability and its Applications, 8 (3):304–307, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vilkas:1968:ADE

[Vil68] E. I. Vilkas. The axiomatic definition of equilibrium points and the value of a non-coalitional nperson game. Theory of Probability and its Applications, 13 (3):523–527, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vallander:1969:LDA

[VIL69] S. S. Vallander, I. A. Ibragimov, and N. G. Lindtrop. On limit distributions of absolute values of consecutive differences of independent variables. Theory of Probability and its Applications, 14(4):668–681, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vilchevskii:1989:PIT

[Vil89] N. O. Vil′chevskiĭ. On a process of inventory theory. Theory of Probability and its Applications, 33(3):582–585, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 625–628.

Vinogradov:1964:ADB

[Vin64] O. P. Vinogradov. On an agedependent branching process. Theory of Probability and its Applications, 9(1):131–136, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradov:1968:PDM

[Vin68] O. P. Vinogradov. The problem of the distribution of the maximal queue size and its application. Theory of Probability and its Applications, 13(2):346–353, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradov:1974:DDF

[Vin74] O. P. Vinogradov. The definition of distribution functions with increasing hazard rate in terms of the Laplace transform. Theory of Probability and its Applications, 18(4):811–813, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradov:1976:SRN

[Vin76] O. P. Vinogradov. On summing a random number of random variables having an increasing hazard rate or a strongly unimodal distribution. Theory of Probability and its Applications, 21 (1):205–209, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradov:1977:AIF

[Vin77] O. P. Vinogradov. On applications of an inversion formula for the Laplace transform. Theory of Probability and its Applications, 21(4):836–839, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradova:1986:ANAa

[Vin86a] T. R. Vinogradova. On the accuracy of normal approximation on sets defined by a smooth function. I. Theory of Probability and its Applications, 30(2):235–246, ???? 1986. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradova:1986:ANAb

[Vin86b] T. R. Vinogradova. On the accuracy of normal approximation on sets defined by a smooth function. II. Theory of Probability and its Applications, 30 (3):590–594, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinokurov:1988:CGW

[Vin88] G. V. Vinokurov. On a critical Galton–Watson branching process with emigration. Theory of Probability and its Applications, 32(2):350–353, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradova:1990:ECF

[Vin90] T. R. Vinogradova. Estimation of characteristic functions of functionals of multidimensional Gaussian random variables. Theory of Probability and its Applications, 34(2):318–321, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 360–364.

Vinogradov:1994:RWC

[Vin94] O. P. Vinogradov. Random walk with correction and tandem system with identical service times. Theory of Probability and its Applications, 38(1):169–173, March 1994. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 171–176.

Vinogradov:1995:CAP

[Vin95] O. P. Vinogradov. On certain asymptotic properties of waiting time in a multiserver queueing system with identical times. Theory of Probability and its Applications, 39(4):714–718, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 799–803.

Vinogradov:1996:OSJ

[Vin96] O. P. Vinogradov. On the output stream and the joint distribution of sojourn times in a multiphase system with identical service. Theory of Probability and its Applications, 40(3):581–588, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 646–654.

Vinogradov:1999:RPI

[Vin99] O. P. Vinogradov. Ruin probability of an insurance company in the case where intervals between payments have unequal exponential distributions. Theory of Probability and its Applications, 43(2):337–342, June 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97697.

Virtser:1970:CLT

[Vir70] A. D. Virtser. Central Limit Theorem for semisimple Lie groups. Theory of Probability and its Applications, 15(4):667–687, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Virtser:1974:LTC

[Vir74] A. D. Virtser. Limit theorems for compositions of distributions on certain nilpotent Lie groups. Theory of Probability and its Applications, 19 (1):86–105, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Virtser:1980:PRM

[Vir80] A. D. Virtser. On products of random matrices and operators. Theory of Probability and its Applications, 24(2):367–377, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Virtser:1984:SSL

[Vir84] A. D. Virtser. On the simplicity of the spectrum of the Lyapunov characteristic indices of a product of random matrices. Theory of Probability and its Applications, 28(1):122–135, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Viskov:1964:TAF

[Vis64] O. V. Viskov. Two asymptotic formulas in the theory of queues. Theory of Probability and its Applications, 9(1):158–159, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Viskov:1986:ACN

[Vis86] O. V. Viskov. An algebraic characterization of the normal distribution. Theory of Probability and its Applications, 30(1):141–143, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Viskov:1992:SRH

[Vis92] O. V. Viskov. Some remarks on Hermite polynomials. Theory of Probability and its Applications, 36(3):633–637, September 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Vittver:1970:SDD

[Vit70] Gisela Vittver. The stationary demand distribution under partial return. Theory of Probability and its Applications, 15 (2):343–348, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vere-Jones:1964:RCP

[VJ64] D. Vere-Jones. A rate of convergence problem in the theory of queues. Theory of Probability and its Applications, 9 (1):96–103, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vere-Jones:1959:CPT

[VJK59] D. Vere-Jones and David G.

Kendall. A commutativity problem in the theory of Markov chains. Theory of Probability and its Applications, 4(1):92–95, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhaniya:1967:SIO

[VK67] N. N. Vakhaniya and N. P.

Kandelski. A stochastic integral for operator-valued functions. Theory of Probability and its Applications, 12(3):525–528, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vsekhsvyatskii:1986:EMO

[VK86] S. Yu. Vsekhsvyatskii and V. V.

Kalashnikov. Estimates of the moments of occurrence of rare events in regenerative processes. Theory of Probability and its Applications, 30(3):618–621, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vsekhsvyatskii:1989:ERT

[VK89] S. Yu. Vsekhsvyatskii and V. V.

Kalashnikov. Estimates in Renyi’s theorem in terms of renewal theory. Theory of Probability and its Applications, 33 (2):347–352, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhania:1997:RVV

[VK97] N. N. Vakhania and N. P. Kandelaki. Random vectors with values in complex Hilbert spaces. Theory of Probability and its Applications, 41(1):116–131, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97506. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(1), (1996), pp. 31–52.

Vishik:1964:SPF

[VKR64] S. M. Vishik, A. A. Kobrinskii, and A. L. Rosental’. Separating partitions for a finite family of measures. Theory of Probability and its Applications, 9 (1):148–149, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vasjanin:1976:MRS

[VL76] V. V. Vasjanin and A. I. Linetskii. The mixing rate of stochastic flows and non-vanishing sets. Theory of Probability and its Applications, 20(3):638–641, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirov:1956:MCM

[Vla56] V. S. Vladimirov. Monte Carlo methods as applied to the calculation of the lowest eigenvalue and the associated eigen-function of a linear integral equation. Theory of Probability and its Applications, 1(1):101–116, ???? 1956. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirskii:1979:TPC

[Vla79] Yu. N. Vladimirskii. On topological properties of countablyadditive cylindrical measures. Theory of Probability and its Applications, 24(1):211–215, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirskii:1980:CUW

[Vla80] Yu. N. Vladimirskii. Conditions under which a cylindrical measure on a dual Banach space May be extended to a Radon measure. Theory of Probability and its Applications, 24(3):582–587, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirskii:1981:CMS

[Vla81] Yu. N. Vladimirskii. Cylindrical measures and p-summing operators. Theory of Probability and its Applications, 26 (1):56–68, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirskii:1984:PCS

[Vla84] Yu. N. Vladimirskii. On probabilistic characterization of some classes of locally convex spaces. Theory of Probability and its Applications, 28(3):549–560, ????

1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirskii:1987:TCC

[Vla87]  Yu. N. Vladimirskii.  Tight

components of cylindrical measures. Theory of Probability and its Applications, 31(4):720–723, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vladimirskii:1996:ESC

[Vla96] Yu. N. Vladimirskiĭ. Equimeasurable sets and cylindrical measures. Theory of Probability and its Applications, 40(4):729–736, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 731–740.

Vatutin:1983:LTN

[VM83] V. A. Vatutin and V. G. Mikhailov. Limit theorems for the number of empty cells in an equiprobable scheme for group allocation of particles. Theory of Probability and its Applications, 27(4):734–743, ???? 1983. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinnichenko:1989:GSR

[VM89] S. V. Vinnichenko and V. V. Mazalov. Games with a stopping rule for Wiener processes. Theory of Probability and its Applications, 33(3):550–552, ???? 1989.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 590–591.

Vinnichenko:1991:OSO

[VM91] S. V. Vinnichenko and V. V.

Mazalov. Optimal stopping of observations in control problems for random walks. Theory of Probability and its Applications, 35(4):746–753, ???? 1991. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 669–676.

Vatutin:1997:APM

[VM97] V. A. Vatutin and V. G.

Mikhaĭlov. Asymptotic properties of matrices related to mappings of partitions. Theory of Probability and its Applications, 41(2):318–325, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97552. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 41(2), (1996), pp. 241–250.

Velikij:1995:DPE

[VMS95] Yu. A. Velikij, A. I. Motsa, and D. S. Sil’vestrov. Dual processes and ergodic type theorems for Markov chains in the scheme of series. Theory of Probability and its Applications, 39(4):642–653, ???? 1995. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(4), (1994), pp. 716–730.

Volodin:1978:RSL

[VN78] N. A. Volodin and S. V. Nagaev. A remark on the strong law of large numbers. Theory of Probability and its Applications, 22 (4):810–813, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1994:SEA

[VN94] I. N. Volodin and An. A. Novikov. Statistical estimators with asymptotically minimum drisk. Theory of Probability and its Applications, 38(1):118–128, March 1994. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 38(1), (1993), pp. 20–32.

Volodin:1999:LAE

[VN99] I. N. Volodin and An. A. Novikov. Local asymptotic efficiency of a sequential probability ratio test for d-guarantee discrimination of composite hypotheses. Theory of Probability and its Applications, 43(2): 269–281, June 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97687.

Volkonskii:1957:MDL

[Vol57] V. A. Volkonskii. A multidimensional limit theorem for homogeneous Markov chains with a countable number of states. Theory of Probability and its Applications, 2(2):221–244, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1958:RST

[Vol58a] V. A. Volkonskii. Random substitution of time in strong Markov processes. Theory of Probability and its Applications, 3(3): 310–326, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkov:1958:DSR

[Vol58b] I. S. Volkov. On the distribution of sums of random variables defined on a homogeneous Markov chain with a finite number of states. Theory of Probability and its Applications, 3 (4):384–399, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1959:COD

[Vol59]  V. A. Volkonskii.  Continu-

ous one-dimensional Markov processes and additive functionals derived from them. Theory of Probability and its Applications, 4(2):198–200, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1960:ETD

[Vol60a] V. A. Volkonskii. An ergodic theorem on the distribution of the duration of fades. Theory of Probability and its Applications, 5(3):323–326, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkov:1960:PEV

[Vol60b] I. S. Volkov. On probabilities for extreme values of sums of random variables defined on a homogeneous Markov chain with a finite number of states. Theory of Probability and its Applications, 5(3):308–319, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1961:CNH

[Vol61a] V. A. Volkonskii. Construction of non-homogeneous Markov processes by means of a random substitution of time. Theory of Probability and its Applications, 6(1):42–51, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkov:1961:ASL

[Vol61b] I. S. Volkov. Analysis of some limit theorems for large deviations. Theory of Probability and its Applications, 6(3):302–304, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1964:OEM

[Vol64] V. A. Volkonskii. Optimum estimation of the moment of emergence of a signal in the presence of multiplicative high-frequency Gaussian noise. Theory of Probability and its Applications, 9 (1):72–88, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1965:DBP

[Vol65] I. N. Volodin. On the distinction between the Polya and the Poisson distributions when a large number of small samples is available. Theory of Probability and its Applications, 10(2):335–338, ???? 1965. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1967:NON

[Vol67] I. N. Volodin. The number of observations needed to distinguish between two closely related hypotheses. Theory of Probability and its Applications, 12 (3):519–525, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1969:NOR

[Vol69] I. N. Volodin. On the number of observations required to test two hypotheses for the binomial distribution parameter. Theory of Probability and its Applications, 14(2):320–324, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1970:BDS

[Vol70] I. N. Volodin. Beta-distribution for small parameters. Theory of Probability and its Applications, 15(3):546–549, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1973:DET

[Vol73] I. N. Volodin. Designing experiments for testing the parameters of two normal populations. Theory of Probability and its Applications, 18(1):198–202, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1974:TST

[Vol74] I. N. Volodin. On a two-sample test in the analysis of variance. Theory of Probability and its Applications, 18(4):790–794, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1975:DGW

[Vol75] I. N. Volodin. On the discrimination of gamma and Weibull distributions. Theory of Probability and its Applications, 19 (2):383–390, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1978:BNS

[Vol78a] I. N. Volodin. Bounds for the necessary sample size in statistical classification problems. I. Theory of Probability and its Applications, 22(2):339–348, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1978:ENS

[Vol78b] I. N. Volodin. Estimates of the necessary sample size in problems of statistical classifications. Theory of Probability and its Applications, 22(4):730–745, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1979:LBA

[Vol79] I. N. Volodin. Lower bounds for average sample size and efficiency of statistical inference procedures. Theory of Probability and its Applications, 24 (1):120–129, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1980:LBM

[Vol80] I. N. Volodin. Lower bounds for the mean sample size in goodness-of-fit and homogeneity tests. Theory of Probability and its Applications, 24(3):640–649, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:1981:LBM

[Vol81] I. N. Volodin. Lower bounds for the mean sample size in invariance tests. Theory of Probability and its Applications, 25 (2):356–360, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkova:1986:LTC

[Vol86] E. I. Volkova. Limit theorems for a critical branching process with moving particles. Theory of Probability and its Applications, 30(2):398–403, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volf:1993:EAR

[Vol93] P. Volf. Estimation of additive regression model with proportional intensity. Theory of Probability and its Applications, 37(1):179–181, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 159–162.

Volkova:1996:RCL

[Vol96a] A. Yu. Volkova. A refinement of the Central Limit Theorem for sums of independent random indicators. Theory of Probability and its Applications, 40 (4):791–794, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 885–888.

Volkovich:1996:IDD

[Vol96b] V. Vol′kovich. On V -infinitely divisible distributions. Theory of Probability and its Applications, 40(3):503–517, ???? 1996. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 615–631.

Vorobev:1958:EPB

[Vor58] N. N. Vorob’ev. Equilibrium points in bimatrix games. Theory of Probability and its Applications, 3(3):297–309, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1960:CFS

[Vor60] N. N. Vorob’ev. On coalitionfree strategies. Theory of Probability and its Applications, 5 (4):415–417, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1961:ACI

[Vor61] N. N. Vorob’ev. An analytic characterization of independence and Markov type dependence. Theory of Probability and its Applications, 6(4):390–393, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1962:CFM

[Vor62] N. N. Vorob’ev. Consistent families of measures and their extensions. Theory of Probability and its Applications, 7(2): 147–163, ???? 1962. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1963:MMM

[Vor63a] N. Vorob’ev. Markov measures and Markov extensions. Theory of Probability and its Applications, 8(4):420–429, ???? 1963.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1963:TSI

[Vor63b] N. N. Vorob’ev. On a topologization of the set of interior consistent families of measures. Theory of Probability and its Applications, 8(4):414–419, ???? 1963. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1964:FRT

[Vor64] N. N. Vorob’ev. On families of random transitions. Theory of Probability and its Applications, 9(1):47–64, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobev:1965:ASA

[Vor65] N. M. Vorob’ev. On the application of sequential analysis to the detection of disorder in a manufacturing process. Theory of Probability and its Applications, 10(4):665–668, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vorobyev:1967:CG

[Vor67] N. N. Vorob’yev. Coalition games. Theory of Probability and its Applications, 12(2):251–266, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vostrikova:1982:DDW

[Vos82] L. Yu. Vostrikova. Detection of a “Disorder” in a Wiener process. Theory of Probability and its Applications, 26(2):356–362, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vovk:1988:LIL

[Vov88]  V. G. Vovk.  The law of

the iterated logarithm for random Kolmogorov, or chaotic, sequences. Theory of Probability and its Applications, 32(3): 413–425, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vovk:1992:AEE

[Vov92] V. G. Vovk. Asymptotic efficiency of estimators: Algorithmic approach. Theory of Probability and its Applications, 36 (2):329–343, May 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 247–261.

Vovk:1997:SMV

[Vov97] V. G. Vovk. A strictly martingale version of Kolmogorov’s strong law of large numbers. Theory of Probability and its Applications, 41(3):605–608, ???? 1997. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 41(3), (1996), pp. 639–642.

Viskov:1964:PLC

[VP64] O. V. Viskov and Yu. V.

Prokhorov. The probability of loss of calls in heavy traffic. Theory of Probability and its Applications, 9(1):92–96, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vysochanskii:1983:GIU

[VP83] D. F. Vysochanskii and Yu. I. Petunin. On a Gauss inequality for unimodal distributions. Theory of Probability and its Applications, 27(2):359–361, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1959:SLT

[VR59] V. A. Volkonskii and Yu. A. Rozanov. Some limit theorems for random functions. I. Theory of Probability and its Applications, 4(2):178–197, ???? 1959. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volkonskii:1961:SLT

[VR61] V. A. Volkonskii and Yu. A. Rozanov. Some limit theorems for random functions. II. Theory of Probability and its Applications, 6(2):186–198, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vronskii:1999:RCS

[Vro99] M. A. Vronskii. Rate of convergence in the SLLN for associated sequences and fields. Theory of Probability and its Applications, 43(3):449–462, September 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97699.

Vershik:1977:LMA

[VS77] A. M. Vershik and A. A. Shimdt.

Limit measures arising in the asympyotic theory of symmetric groups. I. Theory of Probability and its Applications, 22 (1):70–85, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vershik:1978:LMA

[VS78] A. M. Vershik and A. A. Shmidt.

Limit measures arising in the asymptotic theory of symmetric groups, II. Theory of Probability and its Applications, 23 (1):36–49, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:1989:DCB

[VS89] V. A. Vatutin and S. M. Sagitov. A decomposable critical Bellman–Harris branching process with two types of particles. I. Theory of Probability and its Applications, 33(3):460–472, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 495–507.

Vatutin:1990:DCB

[VS90] V. A. Vatutin and S. M. Sagitov.

Decomposable critical Bellman– Harris branching process with two particle types. II. Theory of Probability and its Applications, 34(2):216–227, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 251–262.

Volodin:1991:CEW

[VS91] I. N. Volodin and S. V. Simushkin. Confidence estimation within the framework of the d-A Posteriori approach. Theory of Probability and its Applications, 35(2):318–329, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 242–254.

Vatutin:1992:CBP

[VS92] V. A. Vatutin and S. M. Sagitov. A critical branching process: The distant past under favorable present. Theory of Probability and its Applications, 36 (1):86–98, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 3–15.

Vent-Tsel:1973:FEE

[VT73a] A. D. Vent-Tsel’. Formulae for eigenfunctions and eigenmeasures associated with a Markov process. Theory of Probability and its Applications, 18(1):1–26, ???? 1973. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Vent-Tsel:1973:TAF

[VT73b] A. D. Vent-Tsel’. Theorems on the action functional for Gaussian random functions. Theory of Probability and its Applications, 17(3):515–517, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vent-Tsel:1976:ABF

[VT76] A. D. Vent-Tsel’. On the asymptotic behavior of the first eigenvalue of a second-order differential operator with small parameter in higher derivatives. Theory of Probability and its Applications, 20(3):599–602, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhania:1978:COP

[VT78] N. N. Vakhania and V. I.

Tarieladze. Covariance operators of probability measures in locally convex spaces. Theory of Probability and its Applications, 23(1):1–21, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vent-Tsel:1982:ALT

[VT82] A. D. Vent-Tsel. Addendum:

Limit theorems on large deviations for Markov stochastic processes, III. Theory of Probability and its Applications, 26(2):435, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Verdenskaya:1992:AOS

[VT92] N. V. Verdenskaya and A. G. Tartakovskiĭ. Asymptotically optimal sequential testing of multiple hypotheses for nonhomogeneous Gaussian processes in asymmetric case. Theory of Probability and its Applications, 36(3): 536–547, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 434–445.

Vent-Tsel:1973:SPC

[VTF73] A. D. Vent-Tsel’ and M. I. Freidlin. Some problems concerning stability under small random perturbations. Theory of Probability and its Applications, 17 (2):269–283, ???? 1973. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vent-Tsel:1985:PCA

[VTGM85] A. D. Vent-Tsel’, S. A. Gladyshev, and G. N. Mil’shteyn. Piecewise constant approximation for the Monte–Carlo calculation of Wiener integrals. Theory of Probability and its Applications, 29(4):744–752, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhaniya:1981:CRL

[VTT81] N. N. Vakhaniya, V. I. Tarieladze, and A. Tortrat. On the continuity of random linear functions. Theory of Probability and its Applications, 26(1):167–173, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Valkeila:1988:PTC

[VV88] E. Valkeila and L. Yu. Vostrikova.

On predictable tests of (c_{n})- consistency of estimates. Theory of Probability and its Applications, 32(3):477–489, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vvedenskaya:1984:SRV

[Vve84] E. R. Vvedenskaya. Sums of random vectors with values in a Hilbert space. Theory of Probability and its Applications, 28 (4):797–800, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vvedenskaya:1985:EEU

[Vve85a] E. R. Vvedenskaya. On the existence of exact upper sequences. Theory of Probability and its Applications, 29(2): 394–397, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vvedenskaya:1985:LBS

[Vve85b] E. R. Vvedenskaya. On the limit behavior of sums of random vectors with values in Hilbert space. Theory of Probability and its Applications, 29(3):620–622, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vvedenskaya:1989:MEU

[Vve89] E. R. Vvedenskaya. On the monotonicity of exact upper sequences for sums of independent random vectors in Hilbert space. Theory of Probability and its Applications, 33(1):148–151, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(1), (1988), pp. 159–163.

Vatutin:1999:SBP

[VW99] V. Vatutin and A. Wakolbinger. Spatial branching populations with long individual lifetimes. Theory of Probability and its Applications, 43(4): 620–632, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97712.

Vyugin:1977:AMB

[V’y77] O. V. V’yugin. Asymptotic moments for branching processes close to critical ones. Theory of Probability and its Applications, 21(4):825–833, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vyugin:1978:TPBa

[V’y78a]  O. V. V’yugin.  On tran-

sition phenomena in Bellmanharris branching process. Theory of Probability and its Applications, 22(4):836–840, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vyugin:1978:TPBb

[V’y78b] O. V. V’yugin. Transition phenomena in branching processes with final type. Theory of Probability and its Applications, 23 (1):163–168, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vyugin:1988:DRF

[V’y88] V. V. V’yugin. On the defect of randomness of a finite object with respect to measures with given complexity bounds. Theory of Probability and its Applications, 32(3):508–512, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vyugin:1998:ECP

[V′y98a] V. V. V′yugin. Effective convergence in probability and an ergodic theorem for individual random sequences. Theory of Probability and its Applications, 42(1):39–50, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97591.

Original Russian article in Teor. Veroyatnost. i Primenen., 42(1), (1997), pp. 35–50.

Vyugin:1998:LHR

[V′y98b] V. V. V′yugin. On the longest head-run in an individual random sequence. Theory of Probability and its Applications, 42 (3):541–546, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 42(3), (1997), pp. 608–615.

Vorobev:1980:BTT

[VZ80] L. S. Vorob’ev and I. G. Zhurbenko. Bounds for C(α)- tests and their applications. Theory of Probability and its Applications, 24(2):253–268, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Walshaw:1993:AEV

[Wal93] D. Walshaw. An application in extreme value theory of nonregular maximum likelihood estimation. Theory of Probability and its Applications, 37(1):182–184, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 211–213.

Whittle:1958:CGC

[Whi58] P. Whittle. Continuous generalizations of Chebyshev’s inequality. Theory of Probability and its Applications, 3(4): 358–366, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Whittle:1960:BML

[Whi60] P. Whittle. Bounds for the moments of linear and quadratic forms in independent variables.

Theory of Probability and its Applications, 5(3):302–305, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Whittle:1964:CNQ

[Whi64] P. Whittle. On the convergence to normality of quadratic forms in independent variables. Theory of Probability and its Applications, 9(1):103–108, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Whittle:1969:RKI

[Whi69] P. Whittle. Refinements of Kolmogorov’s inequality. Theory of Probability and its Applications, 14(2):310–311, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Whittle:1982:DDE

[Whi82] P. Whittle. A direct derivation of the equilbrium distribution for a polymerization process. Theory of Probability and its Applications, 26(2):344–355, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wittwer:1971:ICP

[Wit71] G. Wittwer. An inventory control problem with partial recovery. Theory of Probability and its Applications, 16(3):506–515, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wolfowitz:1965:AEM

[Wol65] J. Wolfowitz. Asymptotic efficiency of the maximum likelihood estimator. Theory of Probability and its Applications, 10 (2):247–260, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wolfowitz:1976:MPE

[Wol76] J. Wolfowitz. Maximum probability estimators in the classical case and in the “Almost smooth” case. Theory of Probability and its Applications, 20(2):363–371, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Won:1959:TOF

[Won59] Li Heng Won. On the theory of optimal filtering of a signal in the presence of internal random noise. Theory of Probability and its Applications, 4 (4):422–426, ???? 1959. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Woyczynski:1974:RSL

[Woy74] W. A. Woyczynski. Random series and laws of large numbers in some Banach spaces. Theory of Probability and its Applications, 18(2):350–355, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Williamson:1969:RPB

[WR69] J. A. Williamson and G. S. Rinehart. Remarks on the periodic behavior in time of sums of random variables. Theory of Probability and its Applications, 14 (4):690–693, ???? 1969. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wilson:1993:UBP

[WS93] S. R. Wilson and P. J. Solomon.

Use of the back-projection method for predictions of the Australian aids epidemic. Theory of Probability and its Applications, 37(1):188–191, March 1993. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Weiss:1966:GML

[WW66] L. Weiss and J. Wolfowitz. Generalized maximum likelihood estimators. Theory of Probability and its Applications, 11 (1):58–81, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Weiss:1968:GML

[WW68] L. Weiss and J. Wolfowitz. Generalized maximum likelihood estimators in a particular case. Theory of Probability and its Applications, 13(4):622–627, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wu:1992:SRS

[WW92] T.-J. Wu and M. T. Wasan. Some results on sequential unbiased estimation of a binomial parameter. Theory of Probability and its Applications, 36(2):411–419, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(2), (1991), pp. 405–411.

Xekalaki:1989:APY

[XP89] E. Xekalaki and J. Panaretos. On the association of the Pareto and the Yule distribution. Theory of Probability and its Applications, 33(1):191–195, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaglom:1956:AFS

[Yag56] A. M. Yaglom. Application of function-space integrals to the evaluation of the statistical sum of quantum statistics. Theory of Probability and its Applications, 1(1):145–149, ???? 1956. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaglom:1957:SCR

[Yag57] A. M. Yaglom. Some classes of random fields in n-dimensional space, related to stationary random processes. Theory of Probability and its Applications, 2 (3):273–320, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaglom:1960:ESL

[Yag60] A. M. Yaglom. Effective solutions of linear approximation problems for multivariate stationary processes with a rational spectrum. Theory of Probability and its Applications, 5 (3):239–264, ???? 1960. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yagers:1968:AAD

[Yag68] P. Yagers. Addendum: Agedependent branching processes allowing immigration. Theory of Probability and its Applications, 13(4):711, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:1981:RBP

[Yak81] A. L. Yakymiv. Reduced branching processes. Theory of Probability and its Applications, 25 (3):584–588, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:1986:APS

[Yak86] A. L. Yakymiv. Asymptotic properties of subcritical and supercritical reduced branching processes. Theory of Probability and its Applications, 30 (1):201–206, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:1987:APT

[Yak87] A. L. Yakymiv. Asymptotic properties of the times the states change in a random record process. Theory of Probability and its Applications, 31(3):508–512, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:1988:ABC

[Yak88] A. L. Yakymiv. Asymptotic behavior of a class of infinitely divisible distributions. Theory of Probability and its Applications, 32(4):628–639, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakushev:1991:OSC

[Yak91]  Yu. F. Yakushev.  On the

optimal service of conflicting flows. Theory of Probability and its Applications, 35(1):184–190, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(1), (1990), pp. 161–167.

Yakymiv:1996:ATR

[Yak96] A. L. Yakymiv. Asymptotics of the k th record times. Theory of Probability and its Applications, 40(4):794–797, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(4), (1995), pp. 925–928.

Yakshyavichus:1998:MEP

[Yak98] Sh. Yakshyavichus. On a method of expansion of the probabilities of lattice random variables. Theory of Probability and its Applications, 42(2):271–282, ???? 1998. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/97612. Original Rus-

sian article in Teor. Veroyatnost. i Primenen., 42(2), (1997), pp. 294–307.

Yamazato:1985:DES

[Yam85] M. Yamazato. OL distributions on Euclidean spaces. Theory of Probability and its Applications, 29(1):1–17, ???? 1985. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yamazato:1995:ACT

[Yam95] M. Yamazato. Absolute continuity of transition probabilities of multidimensional processes with stationary independent increments. Theory of Probability and its Applications, 39 (2):347–354, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 422–429.

Yamazato:1996:SUI

[Yam96] M. Yamazato. On strongly unimodal infinitely divisible distributions of class CME. Theory of Probability and its Applications, 40(3):518–532, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 578–595.

Yanovskaya:1964:MTG

[Yan64] E. B. Yanovskaya. Minimax theorems for games on the unit square. Theory of Probability and its Applications, 9(3):500–502, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanovskaya:1970:SIZ

[Yan70] E. B. Yanovskaya. The solution of infinite zero-sum twoperson games with finitely additive strategies. Theory of Probability and its Applications, 15 (1):153–158, ???? 1970. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanev:1976:CDB

[Yan76] N. M. Yanev. Conditions for degeneracy of φ-branching processes with random φ. Theory of Probability and its Applications, 20(2):421–428, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanushkyavichyus:1979:ESM

[Yan79] R. V. Yanushkyavichyus. On estimates of the stability measure for decompositions of probability distributions into components. Theory of Probability and its Applications, 23(3):507–520, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanushkyavichene:1985:ESC

[Yan85] O. L. Yanushkyavichene. Estimate of the stability of a characterization of the exponential law. Theory of Probability and its Applications, 29(2):281–292, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanushkyavichyus:1989:CES

[Yan89] R. V. Yanushkyavichyus. Convolution equations in stability problems on characterizations of probability laws. Theory of Probability and its Applications, 33 (4):668–681, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 720–734.

Yanushkevichiene:1996:LDH

[Yan96] O. L. Yanushkevichiene. On the limit distributions of homogeneous random symmetric polynomials of arbitrary degree. Theory of Probability and its Applications, 40(3):588–594, ???? 1996. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 40(3), (1995), pp. 671–677.

Yashin:1984:PSH

[Yas84] A. I. Yashin. On a problem of sequential hypothesis testing. Theory of Probability and its Applications, 28(1):157–165, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevic:1979:HMM

[YF79] A. A. Yushkevic and E. A.

Fainberg. On homogeneous Markov models with continuous time and finite or countable state space. Theory of Probability and its Applications, 24(1):156–161, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1980:AHC

[YF80] A. A. Yushkevich and E. A. Fainberg. Addendum: On homogeneous controlled Markov models with continuous time and finite or countable number of states. Theory of Probability and its Applications, 25(1):215, ???? 1980. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yin:1986:LTE

[YK86] Y. Q. Yin and P. R. Krishnaiah. Limit theorem for the eigenvalues of the sample covariance matrix when the underlying distribution is isotropic. Theory of Probability and its Applications, 30(4):861–867, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yin:1987:LTE

[YK87] Y. Q. Yin and P. R. Krishnaiah. Limit theorems for the eigenvalues of a product of large dimensional random matrices when the underlying distribution is isotropic. Theory of Probability and its Applications, 31(2): 342–346, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanev:1984:LCB

[YM84] N. M. Yanev and K. V. Mitov. Lifetimes of critical branching processes with random migration. Theory of Probability and its Applications, 28(3):481–491, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Young:1989:NKI

[YTS89] D. M. Young, D. W. Turner, and J. W. Seaman. A note on a Kolmogorov inequality for the binomial distribution. Theory of Probability and its Applications, 33(4):747–749, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1970:IDV

[Yur70] V. V. Yurinskii. On an infinitedimensional version of S. N. Bernstein’s inequalities. Theory of Probability and its Applications, 15(1):108–109, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1971:AVC

[Yur71a] V. V. Yurinskii. On application of the van der corput lemma to the estimation of the characteristic functions of certain singular distributions. Theory of Probability and its Applications, 16 (2):387–388, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1971:ILD

[Yur71b] V. V. Yurinskii. On inequalities for large deviations for certain statistics. Theory of Probability and its Applications, 16 (2):385–387, ???? 1971. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1972:AIL

[Yur72] V. V. Yurinskii. Addendum: On inequalities for large deviations for certain statistics. Theory of Probability and its Applications, 17(1):193, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1974:SCR

[Yur74] V. V. Yurinskii. On sample continuity of random fields. Theory of Probability and its Applications, 18(3):602–607, ???? 1974. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 18, (1973), pp. 633–639.

Yurinskii:1978:AID

[Yur78a] V. V. Yurinskii. Addendum: On infinitely divisible distributions. Theory of Probability and its Applications, 22(2):430, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1978:ACN

[Yur78b] V. V. Yurinskii. On the approximation of convolutions by normal laws. Theory of Probability and its Applications, 22(4):653–667, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1978:EGA

[Yur78c] V. V. Yurinskii. On the error of the Gaussian approximation for convolutions. Theory of Probability and its Applications, 22 (2):236–247, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1983:ANA

[Yur83] V. V. Yurinskii. On the accuracy of normal approximation of the probability of hitting a ball. Theory of Probability and its Applications, 27(2):280–289, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 27, (1982), pp. 270– 278.

Yurinskii:1989:EAM

[Yur89] V. V. Yurinskii. On the error of averaging multidimensional diffusions. Theory of Probability and its Applications, 33(1): 11–21, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yurinskii:1992:LDAa

[Yur92a] V. V. Yurinskiĭ. On large deviation asymptotics in Hilbert space I. Theory of Probability and its Applications, 36(1):99–114, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 78–92.

Yurinskii:1992:LDAb

[Yur92b] V. V. Yurinskiĭ. On large deviation asymptotics in Hilbert spaces II. Theory of Probability and its Applications, 36(3): 548–554, September 1992. CO-DEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(3), (1991), pp. 535–541.

Yurinskii:1993:ALD

[Yur93a] V. V. Yurinskiĭ. On asymptotics of large deviations in Hilbert space III. Theory of Probability and its Applications, 37 (2):261–268, June 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(2), (1992), pp. 268–275.

Yurinskii:1993:EBP

[Yur93b] V. V. Yurinskiĭ. On exponential bounds for probabilities of large deviations. Theory of Probability and its Applications, 37 (1):113–114, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 136–138.

Yushkevich:1957:SMP

[Yus57] A. A. Yushkevich. On strong Markov processes. Theory of Probability and its Applications, 2(2):181–205, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1960:DSM

[Yus60] A. A. Yushkevich. On the definition of a strong Markov process. Theory of Probability and its Applications, 5(2):216–220, ???? 1960. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1967:LTA

[Yus67]  A. A. Yushkevich.  On a

limit theorem for an attrition game. Theory of Probability and its Applications, 12(2):286–294, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yusas:1973:DMF

[Yus73] I. S. Yusas. On the distribution of the maximum frequency of a multinomial distribution. Theory of Probability and its Applications, 17(4):712–717, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1974:CSG

[Yus74] A. A. Yushkevich. On a class of strategies in general Markov decision models. Theory of Probability and its Applications, 18 (4):777–779, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1976:RCM

[Yus76] A. A. Yushkevich. Reduction of a controlled Markov model with incomplete data to a problem with complete information in the case of Borel state and control space. Theory of Probability and its Applications, 21(1):153–158, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1978:CMM

[Yus78] A. A. Yushkevich. Controlled Markov models with countable state space and continuous time. Theory of Probability and its Applications, 22(2):215–235, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1980:RJC

[Yus80] A. A. Yushkevich. On reducing a jump controllable Markov model to a model with discrete time. Theory of Probability and its Applications, 25 (1):58–69, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1981:CJM

[Yus81] A. A. Yushkevich. Controlled jump Markov models. Theory of Probability and its Applications, 25(2):244–266, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1982:SMC

[Yus82] A. A. Yushkevich. On semi-Markov controlled models with an average reward criterion. Theory of Probability and its Applications, 26(4):796–803, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1987:NBP

[Yus87] A. P. Yushkevich. Nicholas Bernoulli and the publication of James Bernoulli’s ars conjectandi. Theory of Probability and its Applications, 31(2): 286–303, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1988:ABI

[Yus88] A. A. Yushkevich. Application of Bellman inequalities to control of a queueing system. Theory of Probability and its Applications, 32(1):121–125, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1990:VTM

[Yus90] A. A. Yushkevich. Verification theorems for Markov decision processes with controlled deterministic drift and gradual and impulsive controls. Theory of Probability and its Applications, 34(3):474–496, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 528–551.

Yushkevich:1991:AVT

[Yus91a] A. A. Yushkevich. Addendum:

Verification theorems for Markov decision processes with controlled deterministic shift, and gradual and impulse controls.

Theory of Probability and its Applications, 35(1):192, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yushkevich:1991:SCC

[Yus91b] A. A. Yushkevich. Sensitive criteria in the continuous-time twoarmed bandit problem. Theory of Probability and its Applications, 35(4):819–824, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(4), (1990), pp. 793–797.

Zaitsev:1984:RCK

[ZA84] A. Yu. Zaitsev and T. V. Arak. On the rate of convergence in Kolmogorov’s second uniform limit theorem. Theory of Probability and its Applications, 28 (2):351–374, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zahle:1980:PLT

[Zäh80] U. Zähle. A Poisson limit theorem for rare events of a Gaussian sequence. Theory of Probability and its Applications, 25 (1):91–104, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zahle:1983:EPR

[Zäh83]  M. Zähle.  Ergodic proper-

ties of random fields and images with point imbedded processes. Theory of Probability and its Applications, 27(3):536–550, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zahle:1984:APP

[Zäh84] U. Zähle. Asymptotic properties of point processes of rare events of Gaussian fields. I. Theory of Probability and its Applications, 28(4):703–713, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zahle:1986:APP

[Zäh86] U. Zähle. Asymptotic properties of point processes of rare events of Gaussian fields. II. Theory of Probability and its Applications, 30(1):66–74, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1981:SPF

[Zai81] A. Yu. Zaitsev. Some properties of N-fold convolutions of distributions. Theory of Probability and its Applications, 26 (1):148–152, ???? 1981. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1984:AAD

[Zai84] A. Yu. Zaitsev. On the accuracy of approximation of distributions of sums of independent random variables-which are nonzero with a small probability-by means of accompanying laws. Theory of Probability and its Applications, 28(4):657–669, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1985:ARC

[Zai85] A. Yu. Zaitsev. Addendum: On the rate of convergence in Kolmogorov’s second uniform limit theorem. Theory of Probability and its Applications, 29(1):199, ???? 1985. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1987:ELP

[Zai87a] A. Yu. Zaitsev. Estimates of the Lévy–Prokhorov distance in the multivariate Central Limit Theorem for random variables with finite exponential moments. Theory of Probability and its Applications, 31(2):203–220, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1987:LFS

[Zai87b] A. Yu. Zaitsev. On the logarithmic factor in smoothing inequalities for the Lévy and Lévy– Prokhorov distances. Theory of Probability and its Applications, 31(4):691–693, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1988:AEL

[Zai88a] A. Yu. Zaitsev. Addendum: Estimates of the Lévy-Prokhorov distance in the multivariate Central Limit Theorem for random variables with finite exponential moments. Theory of Probability and its Applications, 32(4):750, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1988:UAD

[Zai88b] A. Yu. Zaitsev. On the uniform approximation of distributions of sums of independent random variables. Theory of Probability and its Applications, 32 (1):40–47, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaitsev:1990:MVS

[Zaĭ90] A. Yu. Zaĭtsev. Multidimensional version of the second uniform limit theorem of Kolmogorov. Theory of Probability and its Applications, 34(1): 108–128, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(1), (1989), pp. 128–151.

Zaitsev:1992:EDW

[Zai92] A. Yu. Zaitsev. An example of a distribution whose set of N-fold convolutions is uniformly separated from the set of infinitely divisible laws in distance in variation. Theory of Probability and its Applications, 36(2):419–425, May 1992. CODEN TPRBAU. ISSN 0040-585x (print), 1095-7219 (electronic).

Zakusilo:1990:GSP

[Zak90a] O. K. Zakusilo. General storage processes with additive inputs. Theory of Probability and its Applications, 34(2):240–251, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(2), (1989), pp. 277–288.

Zakusilo:1990:MSP

[Zak90b] O. K. Zakusilo. Markov semigroups and processes with wide domains of infinitesimal operators. Theory of Probability and its Applications, 34(3):515–519, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(3), (1989), pp. 572–576.

Zakusilo:1991:OML

[Zak91] O. K. Zakusilo. Once more on the local behavior of storage processes with stable input. Theory of Probability and its Applications, 35(2):393–396, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 349–352.

Zalesskii:1983:EAN

[Zal83] B. A. Zalesskii. Estimate of the accuracy of normal approximation in Hilbert space. Theory of Probability and its Applications, 27(2):290–298, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 27, (1982), pp. 279–285.

Zalesskii:1986:CRC

[Zal86] B. A. Zalesskii. On the convergence rate in the Central Limit Theorem on a class of sets in Hilbert space. Theory of Probability and its Applications, 30 (4):702–711, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zalesskii:1989:ANA

[Zal89] B. A. Zalesskiĭ. On the accuracy of normal approximation in Banach spaces. Theory of Probability and its Applications, 33 (2):239–247, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 257–265.

Zalesskii:1990:AGA

[Zal90a] B. A. Zalesskiĭ. The accuracy of Gaussian approximation in Banach spaces. Theory of Probability and its Applications, 34 (4):747–748, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 815–817.

Zalesskii:1990:PLD

[Zal90b] B. A. Zalesskiĭ. Probabilities of large deviations in Hilbert space. Theory of Probability and its Applications, 34(4):591–596, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 650–655.

Zamanov:1970:ASS

[Zam70] T. A. Zamanov. An approximate solution of stochastic differential equations with retarded argument. Theory of Probability and its Applications, 15(1):139–142, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zanzotto:1999:RCS

[Zan99] P. A. Zanzotto. Representation of a class of semimartingales as stable integrals. Theory of Probability and its Applications, 43(4): 666–676, December 1999. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97725.

Zaslavskii:1964:IPS

[Zas64] A. Kh. Zaslavskii. On the isomorphism problem for stationary processes. Theory of Probability and its Applications, 9 (2):291–298, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaslavskii:1967:STS

[Zas67] A. E. Zaslavskii. On statistical tests for a simple hypothesis involving a multi-dimensional normal distribution. Theory of Probability and its Applications, 12 (3):514–519, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaslavskii:1973:ECR

[Zas73] A. E. Zaslavskii. An estimate for the convergence rate in a renewal theorem for random variables defined on a Markov chain. Theory of Probability and its Applications, 17(3):535–543, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zempleni:1993:TMI

[Zem93] A. Zempléni. Testing maxinfinite divisibility. Theory of Probability and its Applications, 37(1):191–193, March 1993. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(1), (1992), pp. 173–175.

Zerakidze:1970:CBP

[Zer70] Z. S. Zerakidze. On the consistency of best and pseudobest estimates. Theory of Probability and its Applications, 15(3):528–531, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zerakidze:1987:CSS

[Zer87] Z. S. Zerakidze. The construction of statistical structures. Theory of Probability and its Applications, 31(3):505–508, ???? 1987. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zezel:1975:MEC

[Zez75] Yu. N. Zezel’. The minimal essentially complete class of discrimination rules for unknown moments of higher orders. Theory of Probability and its Applications, 19(4):801–803, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhdanok:1990:MIR

[Zhd90] T. A. Zhdanok. On the measurability of the inverse of a random linear operator. Theory of Probability and its Applications, 34(4):709–713, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(4), (1989), pp. 774–778.

Zhitomirskii:1963:DFT

[Zhi63] I. S. Zhitomirskii. On the distribution function of the time of first exit. Theory of Probability and its Applications, 8 (2):147–155, ???? 1963. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhuravlev:1966:MSS

[Zhu66] O. G. Zhuravlev. Minimal sufficient statistics for a sequence of independent random variables. Theory of Probability and its Applications, 11(2):282–291, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1968:MRD

[Zhu68] I. G. Zhurbenko. Moments of random determinants. Theory of Probability and its Applications, 13(4):682–686, ???? 1968. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1970:EMS

[Zhu70] I. G. Zhurbenko. On the estimation of mixed semi-invariants for a certain class of random processes. Theory of Probability and its Applications, 15(3):525–528, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1972:SSI

[Zhu72]  I. G. Zhurbenko.  Spectral

semi-invariants of stationary processes with sufficiently strong mixing. Theory of Probability and its Applications, 17(1): 151–153, ???? 1972. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1981:ESD

[Zhu81] I. G. Zhurbenko. On the efficiency of spectral density estimates of a stationary process. Theory of Probability and its Applications, 25(3):466–480, ???? 1981. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1984:ESD

[Zhu84] I. G. Zhurbenko. On the efficiency of spectral density estimates of a stationary process. II. Theory of Probability and its Applications, 28(2):409–419, ???? 1984. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1986:SEH

[Zhu86]  I. G. Zhurbenko.  Statisti-

cal estimation of higher-order spectra. Theory of Probability and its Applications, 30(1): 75–86, ???? 1986. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhulenev:1988:LIL

[Zhu88] S. V. Zhulenev. The law of the iterated logarithm and its indicators. Theory of Probability and its Applications, 32(4):714–717, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhurbenko:1990:TSR

[Zhu90] I. G. Zhurbenko. Time series research in the department of probability theory and mathematical statistics at Moscow University. Theory of Probability and its Applications, 34(1):174–179, ???? 1990. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zakhar-Itkin:1970:MWT

[ZI70] M. Kh. Zakhar-Itkin. A multiwave transmission line with random inhomogeneities and Brownian motion in Siegel’s circle. Theory of Probability and its Applications, 15(2):282–294, ???? 1970. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zakhar-Itkin:1971:MTL

[ZI71] M. Kh. Zakhar-Itkin. Multimode transmission line noises and summation of random vectors. Theory of Probability and its Applications, 16(3):556–559, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zakhar-Itkin:1974:SRC

[ZI74] M. Kh. Zakhar-Itkin. On statistics of the reflection coefficient of a dissipative multiwave transmission line with random section lengths. Theory of Probability and its Applications, 18 (2):281–294, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zorin:1989:DQF

[ZI89] V. A. Zorin and A. K. Iyubimov. The distribution of quadratic forms of Gaussian vectors. Theory of Probability and its Applications, 33(3):557–560, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zieba:1992:SPC

[Zie92] W. Zieba. On some properties of conditional Amarts. Theory of Probability and its Applications, 36(3):637–639, September 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zielinski:1993:MPP

[Zie93] R. Zieliński. A moment preserving probability metric. Theory of Probability and its Applications, 37(3):576–579, September 1993. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 37(3), (1992), pp. 602–604.

Zigangirov:1966:POS

[Zig66] K. Sh. Zigangirov. On a problem in optimal scanning. Theory of Probability and its Applications, 11(2):294–298, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zigel:1982:UEC

[Zig82] G. Zigel’. Upper estimates for the concentration function in Hilbert space. Theory of Probability and its Applications, 26(2):328–343, ???? 1982. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1958:IQP

[Zin58] A. A. Zinger. Independence of quasi-polynomial statistics and analytical properties of distributions. Theory of Probability and its Applications, 3(3):247–265, ???? 1958. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1965:CLD

[Zin65] A. A. Zinger. On a class of limit distributions for normalized sums of independent random variables. Theory of Probability and its Applications, 10 (4):607–626, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1971:LLC

[Zin71]  A. A. Zinger.  Limit laws

for cumulative sums of independent random variables with a finite number of distribution types. Theory of Probability and its Applications, 16(4):596–619, ???? 1971. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 16, (1971), pp. 614– 637.

Zinger:1976:SID

[Zin76] A. A. Zinger. On samples with identically distributed linear statistics. Theory of Probability and its Applications, 20 (3):655–660, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 20, (1975), pp. 668–672.

Zinchenko:1979:LGR

[Zin79] N. M. Zinchenko. Local growth of random fields with independent increments. Theory of Probability and its Applications, 24 (1):184–191, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1980:CMD

[Zin80] A. A. Zinger. On a characterization of the multi-dimensional normal law by the independence of linear statistics. Theory of Probability and its Applications, 24(2):388–392, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinchenko:1986:SIP

[Zin86] N. M. Zinchenko. A strong invariance principle for sums of random variables from the domain of attraction of a stable law. Theory of Probability and its Applications, 30(1):148–152, ???? 1986. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinchenko:1988:ABI

[Zin88] N. M. Zinchenko. Asymptotic behavior of the increments of stable random processes with jumps of one sign. Theory of Probability and its Applications, 32 (4):724–727, ???? 1988. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1961:HPB

[ZK61] V. M. Zolotarev and V. S. Korolyuk. On a hypothesis proposed by B. V. Gnedenko. Theory of Probability and its Applications, 6(4):431–435, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1976:SID

[ZK76] V. M. Zolotarev and V. M. Kruglov. The structure of infinitely divisible distributions on a bicompact Abelian group. Theory of Probability and its Applications, 20(4):698–709, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1977:PRT

[ZK77] A. A. Zinger and A. M. Kagan.

On the problem of reconstruction of the type of a distribution. Theory of Probability and its Applications, 21(2):389–392, ???? 1977. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1992:LSE

[ZK92] A. A. Zinger and A. M. Kagan.

A least squares estimate, nonquadratic losses, and Gaussian distributions. Theory of Probability and its Applications, 36 (1):115–123, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 34–41.

Zhao:1991:ASN

[ZKC91] L. C. Zhao, P. R. Krishnaiah, and X. R. Chen. Almost sure L_{r}-norm convergence for databased histogram density estimates. Theory of Probability and its Applications, 35(2):396–403, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 391–397.

Zinger:1964:CPN

[ZL64a] A. A. Zinger and Yu. V. Linnik.

A characteristic property of the normal distribution. Theory of Probability and its Applications, 9(4):624–626, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zinger:1964:PSN

[ZL64b] A. A. Zinger and Yu. V. Linnik. On polynomial statistics for the normal and related laws. Theory of Probability and its Applications, 9(3):494–497, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1974:LDR

[ZM74] A. M. Zubkov and V. G. Mikhailov. Limit distributions of random variables associated with long duplications in a sequence of independent trials. Theory of Probability and its Applications, 19(1):172–179, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhulenev:1979:SLL

[ZM79a] S. V. Zhulenev and V. S. Medovyj. The strong law of large numbers and normality of Kesten’ procedure. Theory of Probability and its Applications, 23(3):615–621, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1979:EAP

[ZM79b] A. M. Zubkov and V. G. Mikhailov. An estimate of the accuracy of a Poisson approximation in the problem of allocating particles in cells. Theory of Probability and its Applications, 23(4):789–794, ???? 1979. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1980:RTS

[ZM80] A. M. Zubkov and V. G.

Mikhailov. Repetitions of stuples in a sequence of independent trials. Theory of Probability and its Applications, 24 (2):269–282, ???? 1980. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhizhina:1995:LLT

[ZM95] E. A. Zhizhina and R. A. Minlos.

A local limit theorem for nonhomogeneous random walk on a lattice. Theory of Probability and its Applications, 39(3):490–503, ???? 1995. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(3), (1994), pp. 513–529.

Zolotarev:1957:MST

[Zol57a] V. M. Zolotarev. Mellin–Stieltjes transforms in probability theory. Theory of Probability and its Applications, 2(4):433–460, ???? 1957. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1957:MES

[Zol57b] V. M. Zolotarev. More exact statements of several theorems in the theory of branching processes. Theory of Probability and its Applications, 2 (2):245–253, ???? 1957. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1958:DSI

[Zol58] V. M. Zolotarev. Distribution of the superposition of infinitely divisible processes. Theory of Probability and its Applications, 3(2):185–188, ???? 1958. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1961:CCP

[Zol61a]  V. M. Zolotarev.  Concern-

ing a certain probability problem. Theory of Probability and its Applications, 6(2):201–204, ???? 1961. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1961:ABC

[Zol61b] V. M. Zolotarev. On the asymptotic behavior of a class of infinitely divisible laws. Theory of Probability and its Applications, 6(3):304–307, ???? 1961. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1964:FPT

[Zol64a] V. M. Zolotarev. The first passage time of a level and the behavior at infinity for a class of processes with independent increments. Theory of Probability and its Applications, 9(4): 653–662, ???? 1964. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1964:ABC

[Zol64b] V. M. Zolotarev. On asymptotically best constants in refinements of mean limit theorems. Theory of Probability and its Applications, 9(2):268–276, ???? 1964. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1965:ABD

[Zol65a] V. M. Zolotarev. Asymptotic behavior of the distributions of processes with independent increments. Theory of Probability and its Applications, 10 (1):28–44, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1965:CDT

[Zol65b] V. M. Zolotarev. On the closeness of the distributions of two sums of independent random variables. Theory of Probability and its Applications, 10(3): 472–479, ???? 1965. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1966:AER

[Zol66a] V. M. Zolotarev. An absolute estimate of the remainder term in the Central Limit Theorem. Theory of Probability and its Applications, 11(1):95–105, ???? 1966. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1966:AAB

[Zol66b] V. M. Zolotarev. Addendum:

Asymptotic behavior of the distributions of processes with independent increments. Theory of Probability and its Applications, 11(2):320–321, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1966:MAB

[Zol66c] V. M. Zolotarev. A multidimensional analogue of the Berry– Esseen inequality for sets with bounded diameter. Theory of Probability and its Applications, 11(3):447–454, ???? 1966. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1967:GLF

[Zol67a] V. M. Zolotarev. A generalization of the Lindeberg–Feller theorem. Theory of Probability and its Applications, 12(4):608–618, ???? 1967. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1967:DSL

[Zol67b] V. M. Zolotarev. On the divisibility of stable laws. Theory of Probability and its Applications, 12(3):506–508, ???? 1967. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1968:PSD

[Zol68a] V. M. Zolotarev. On the problem of stability of the decomposition of the normal law into components. Theory of Probability and its Applications, 13(4):697–700, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1968:SRM

[Zol68b] V. M. Zolotarev. Some remarks on multidimensional Bernstein– Kolmogorov-type inequalities. Theory of Probability and its Applications, 13(2):281–286, ???? 1968. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1975:QEP

[Zol75] V. M. Zolotarev. Qualitative estimates in problems of continuity of queuing systems. Theory of Probability and its Applications, 20(1):211–213, ???? 1975. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1976:CSS

[Zol76] V. M. Zolotarev. On the continuity of stochastic sequences generated by recurrent processes. Theory of Probability and its Applications, 20(4):819–832, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1977:ADS

[Zol77a] V. M. Zolotarev. Approximation of distributions of sums of independent random variables with values in infinitedimensional spaces. Theory of Probability and its Applications, 21(4):721–737, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1977:SCQ

[Zol77b] V. M. Zolotarev. On stochastic continuity of queueing systems of type G|G|1. Theory of Probability and its Applications, 21(2):250–269, ???? 1977. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1978:AAD

[Zol78a] V. M. Zolotarev. Addendum: Approximation of distributions of sums of random variables with values in infinite-dimensional spaces. Theory of Probability and its Applications, 22(4):881, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1978:AEG

[Zol78b] V. M. Zolotarev. Addendum: On the error of the Gaussian approximation for convolutions. Theory of Probability and its Applications, 22(3):583–584, ???? 1978. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1978:IMP

[Zol78c] V. M. Zolotarev. Ideal metrics in the problem of approximating distributions of sums of independent random variables. Theory of Probability and its Applications, 22(3):433–449, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1978:QEC

[Zol78d] V. M. Zolotarev. Quantitative estimates for the continuity property of queueing systems of type G|G|∞. Theory of Probability and its Applications, 22 (4):679–691, ???? 1978. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1979:P

[Zol79] V. M. Zolotarev. On pseudomoments. Theory of Probability and its Applications, 23(2):269–278, ???? 1979. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1984:APM

[Zol84a] V. M. Zolotarev. Addendum:

Probability metrics. Theory of Probability and its Applications, 28(4):856, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1984:PM

[Zol84b] V. M. Zolotarev. Probability metrics. Theory of Probability and its Applications, 28(2): 278–302, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1990:LTS

[Zol90] V. M. Zolotarev. Limit theorems as stability theorems. Theory of Probability and its Applications, 34(1):153–163, ???? 1990. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 34(1), (1989), pp. 178–189.

Zolotarev:1992:AEP

[Zol92a] V. M. Zolotarev. Asymptotic expansions with probability one for sums of independent random variables. Theory of Probability and its Applications, 36(4): 800–802, December 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zolotarev:1992:RCT

[Zol92b] V. M. Zolotarev. Reflections on the classical theory of limit theorems. Theory of Probability and its Applications, 36(1):124–137, March 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 111–126.

Zolotarev:1995:RDS

[Zol95] V. M. Zolotarev. On representation of densities of stable laws by special functions. Theory of Probability and its Applications, 39(2):354–362, ???? 1995. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 39(2), (1994), pp. 429–437.

Zolotarev:1976:TSE

[ZS76] V. M. Zolotarev and V. V. Senatov. Two-sided estimates of Levy’s metric. Theory of Probability and its Applications, 20 (2):234–245, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zalesskii:1984:CMN

[ZS84] B. A. Zalesskii and V. V. Sazonov. Closeness of moments for normal approximation in a Hilbert space. Theory of Probability and its Applications, 28 (2):263–277, ???? 1984. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zalesskii:1989:NAHa

[ZSU89a] B. A. Zalesskiĭ, V. V. Sazonov, and V. V. Ul′yanov. Normal approximation in Hilbert space. I. Theory of Probability and its Applications, 33(2):207–227, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(2), (1988), pp. 225–245.

Zalesskii:1989:NAHb

[ZSU89b] B. A. Zalesskiĭ, V. V. Sazonov, and V. V. Ul′yanov. Normal approximation in Hilbert space. II. Theory of Probability and its Applications, 33(3):473–487, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(3), (1988), pp. 508–521.

Zalesskii:1989:REA

[ZSU89c] B. A. Zalesskiĭ, V. V. Sazonov, and V. V. Ul′yanov. A regular estimate of the accuracy of normal approximation in Hilbert space. Theory of Probability and its Applications, 33(4):700–701, ???? 1989. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 33(4), (1988), pp. 753–754.

Zalys:1992:SSG

[ZT92] A. J. Zalys and A. A. Tempel’man. Spatially stationary generalized Gaussian fields with local interaction. Theory of Probability and its Applications, 36 (1):197–200, March 1992. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 36(1), (1991), pp. 149–152.

Zubkov:1972:LPB

[Zub72] A. M. Zubkov. Life-periods of a branching process with immigration. Theory of Probability and its Applications, 17(1):174–183, ???? 1972. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1973:DCB

[Zub73] A. M. Zubkov. A degeneracy condition for bounded continuoustime branching processes. Theory of Probability and its Applications, 17(2):284–297, ???? 1973. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1974:LDS

[Zub74] A. M. Zubkov. Limit distributions for a statistical estimate of the entropy. Theory of Probability and its Applications, 18 (3):611–618, ???? 1974. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1975:ABG

[Zub75] A. M. Zubkov. Analogies between Galton–Watson processes and φ-branching processes. Theory of Probability and its Applications, 19(2):309–331, ???? 1975. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1976:LDD

[Zub76] A. M. Zubkov. Limiting distributions of the distance to the closest common ancestor. Theory of Probability and its Applications, 20(3):602–612, ???? 1976. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1983:LBD

[Zub83] A. M. Zubkov. Limit behavior of decomposable critical branching processes with two types of particles. Theory of Probability and its Applications, 27(2):235–247, ???? 1983. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1987:CPM

[Zub87] A. M. Zubkov. On compound probability metrics with fixed minimal ones. Theory of Probability and its Applications, 31 (1):135–137, ???? 1987. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:1999:TSI

[Zub99]  A. M. Zubkov.  Two-sided

inequalities for Laplace transforms. Theory of Probability and its Applications, 43(4):676– 681, December 1999. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97719.

Zuev:1988:BPT

[Zue88] S. A. Zuev. Bounds for the percolation threshold for a square lattice. Theory of Probability and its Applications, 32(3):551–553, ???? 1988. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zuev:1998:EEV

[Zue98] Yu. A. Zuev. On the estimation of efficiency of voting procedures. Theory of Probability and its Applications, 42 (1):73–81, ???? 1998. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97594.

Zulenev:1976:ASA

[Zul76] S. V. Zulenev. On accelerated stochastic approximation. Theory of Probability and its Applications, 20(2):393–401, ???? 1976. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zuparov:1992:RCC

[Zup92] T. M. Zuparov. On the rate of convergence in the Central Limit Theorem for weakly dependent random variables. Theory of Probability and its Applications, 36(4):783–792, December 1992. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).